A Complete TUGboat Bibliography

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 585 1640, +1 801 581 4148

     E-mail: beebe@math.utah.edu, beebe@acm.org,
                 beebe@computer.org (Internet)
   WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 4.20

Abstract This bibliography summarizes the contents of the TEX User Group journal, TUGboat. It is produced semiautomatically by an awk program from the tablesof-contents files (tb*.cnt) released periodically by the journal editors, or since 2017 (volume 38), from journal HTML pages. Hand editing then supplies final tweaks.

Title word cross-reference

#10 [1422]. #12 [1459]. #13 [1534]. #15 [1588]. #7 [1335]. #8 [1335]. #9 [1422].

3 [4501, 2831, 3022]. ^{6} [3219]. ℵ [2862]. ↪ [1421]. Λ [2732]. Ω [1757, 1879, 1933, 1939]. Ω_{2} [3070, 3071]. ε [1758, 2016].

*.bib [4140].

-.02emCAD [3304]. -.03em [3361]. -.03emenberg} [3778, 3997]. -.15em [470]. -.1em [4176, 2800]. -.1emAMS [2704]. -.1emVDU [583]. -1000pt [4954]. -1pt [3848, 2658]. -Babel [1879]. -based [2826]. -TEX [1758, 2016].

.02emCAD [3304]. .03em [3361]. .03emenberg} [3778, 3997]. .05em [2704]. .05em- [2704]. .05emì [2959]. .06em [4215]. .06em-foundry [4215]. .15em [470]. .1em [4176, 2800, 2782]. .1em-TEX [2782]. .1emAMS [2704]. .1emVDU [583]. .dtx [3297]. .dvi [941, 1962]. .ins [3297].

0.4 [3601]. 0.60 [3520]. 0.8 — Prologue [4279]. 0.90 [4135, 4138]. 0.999 [1876]. 000 [1000]. 0pt [4873].

1 [4260, 4476, 944, 3108, 3182, 3307, 3310, 3565, 3605, 3644, 3683, 3719, 3780, 3812, 3928, 4149, 4872, 1851, 3167, 3239, 1811, 4621, 2824, 2696, 3167]. 1.0 [1744]. 1.1 [583]. 1.3 [2738]. 1.3 — the [1932]. 1.40 [3257]. 1.750 [3622]. 1/2011 [3605, 3644]. 1/2012 [3719]. 1/2013 [3812]. 1/2014 [3928]. 1/2015 [4031]. 1/2016 [4149]. 1/2017 [4260]. 1/2018 [4380]. 1/2019 [4476]. 1/2021 [4697]. 1/2022 [4791]. 1/2023 [4913]. 10 [3606]. 10/20 [72]. 100 [37]. 1000 [198]. 1000pt [4954]. 101 [2168]. 11 [1422]. 11#4 [1407]. 11.8MB [4009]. 11/12th [1778]. 1100 [76]. 12 [3681, 1420]. 1200 [219]. 12th [1778]. 13 [3744]. 130 [196]. 14 [2591, 3001, 3779, 1671]. 1400-page [1110]. 14th [4765]. 15 [1184, 3001, 3856]. 158 [2126]. 15th [4960, 593, 655, 700]. 17 [1714]. 19 [1789, 4576, 4594]. 1980 [6, 20]. 1981 [59, 74]. 1982 [185, 61, 163]. 1983 [226]. 1985 [312, 351, 352, 368, 390]. 1986 [442, 389, 433, 443, 556, 834, 555]. 1987 [552, 663, 665, 666, 522, 593, 655, 700]. 1988 [709, 705, 825, 744, 786]. 1989 [879, 881, 830, 840, 874, 908, 953, 925, 882]. 1990 [1151, 1033, 1146]. 1991 [1192, 1182, 1183, 1193, 1337, 1338, 1401, 1507, 1262, 1184]. 1992 [1339, 1391, 1424, 1460, 1464]. 1993 [1494, 1631]. 1994 [1790, 2098, 1685, 1702, 1778]. 1995 [1768, 1883]. 1996 [1886, 1887, 1918]. 1996 — [1990]. 1996 — Knuth [1989, 2015]. 1997 [1957]. 1998 [2157, 2256]. 1999 [2291, 2375, 2499]. 19th [786]. 1pt [3848, 2658].

2 [950, 1713, 2866, 2945, 3183, 3383, 3492, 3604, 3855, 586, 2726, 3459, 2530, 630]. 2-dimensional [2066]. 2.0 [1118]. 2.09 [424, 1455]. 2.0 — an [1368]. 2.0 — New [4244]. 2.10 [968, 1120]. 2.10.0 [4731]. 2.15.0 [4781]. 2.17.1 [4862]. 2.7.0 [4462]. 2/2013 [3855]. 2/2017 [4306]. 2/2022 [4834]. 2/2023 [4961]. 20 [1254, 1912, 3308, 72]. ’2000 [2447, 2472, 2446, 2254, 2430]. 2000 — L^{A}TEX [2348]. 2000 — The [2384]. ’2001 [2481, 2435]. ’2002 [2588, 2675]. ’2003 [2590, 2265, 2676, 2903, 2733, 2753, 2723, 2682]. ’2003 — The [2591]. 2004 [2642, 2717, 2719, 2850, 2848]. 2005 [2845, 2712, 3023, 3017, 3018, 2955]. 2005-1 [3004]. 2005-4 [3004]. 2006 [3048, 3010, 3012, 3013, 3052, 3051, 3087]. 2006-1 [3183]. 2006/1 [3182]. 2007 [3193, 3226, 3228, 3074, 3111, 3147, 3182, 3183, 3186, 3194, 3162, 3149, 3331]. 2007-2 [3183]. 2007-3 [3310]. 2007/1 [3182]. 2007/2 [3307]. 2008 [3187, 3216, 3307, 3310, 3315, 3386, 3323]. 2008-1 [3310]. 2008-2 [3386]. 2008-3 [3386]. 2008/1 [3307]. 2008/2 [3383]. 2009 [3349, 3351, 3383, 3428, 3430, 3491, 3440]. 2009/2 [3383]. 2009/3 [3428]. 2009/4 [3492]. 2010 [3492, 3563, 3565, 3567, 3602, 3604, 3606, 3519, 3590]. 2010-1 [3565]. 2010-2 [3604]. 2010/2 [3492]. 2010/3 [3567]. 2011 [3605, 3644, 3649, 3682, 3683, 3611]. 2011-1 [3683]. 2012 [3719, 3779, 3780, 3813, 3818]. 2012-1 [3780]. 2013 [3866, 3812, 3855, 3890]. 2014 [3928, 3996, 4030, 4072, 4018]. 2015 [4043, 4044, 4151, 4092, 4093]. 2015/3 [4261]. 2016 [4163, 4164, 4149, 4222, 4190, 4166, 4250, 4123, 4124]. 2016/1 [4261]. 2017 [4260, 4273, 4264, 4274, 4305, 4272, 4249, 4295]. 2017/1 [4381]. 2018 [4349, 4385, 4404, 4391, 4374, 4468, 4566, 4393]. 2018/1 [4477]. 2019 [4515, 4488, 4489, 4476, 4490, 4550, 4415, 4447, 4497, 4529, 4482, 4538]. 2019/1 [4580]. 2020 [4592, 4629, 4622, 4563, 4595, 4644, 4567, 4568, 4586, 4594]. 2020/1 [4624]. 2020/3 [4698]. 2021 [4734, 4738, 4710, 4739, 4730, 4759, 4713, 4711, 4627, 4701]. 2021/1 [4792]. 2022 [4838, 4802, 4831, 4803, 4817, 4861, 4804, 4790]. 2022/1 [4915]. 2023 [4958, 4965, 4925, 4994, 4951, 4957, 4905, 4933, 4836, 4878, 4982, 4983, 4926, 4918]. 2023/1 [4962]. 20 — TOPS [1254]. 20th [4240, 4307, 4327]. 21 [2307, 2384, 2393, 2436, 3384]. 21st [4811]. 21st-century [4811]. 22nd [2528]. 23 [2601, 3495]. 24 [1998, 749]. 25 [2048, 3602, 2851]. 268 [1513]. 27 [2099, 3718]. 276 [1414]. 29 [2171, 3813]. 290 [1420]. 2nd [2711, 1702].

3 [1026, 1038, 2053, 3180, 3386, 3428, 3567, 3682, 3996, 4222, 4540, 4649, 4494, 4983, 2100]. 3-across [1381]. 3-disk [2443]. 3.0 [1209, 809]. 3/2011 [3682]. 3/2012 [3743]. 3/2014 [3996]. 3/2015 [4071]. 3/2016 [4222]. 3/2017 [4346]. 3/2018 [4406]. 3/2019 [4545]. 3/2020 [4623]. 3/2021 [4736]. 3/2022 [4876]. 30 [2233, 3323]. 3033 [31]. 3090 [863]. 31 [3927, 2021]. 31/XI [2021]. 33/34 [2379]. 35/36 [2515]. 355 [2015]. 35mm [430, 1085]. 36 [4009]. 37/38 [2515]. 370 [31, 115]. 370/3033 [31]. 37th [4271]. 386-based [695]. 38th [4271, 4390]. 39/40 [2587]. 39th [4390]. 3D [2895, 2095, 3527]. 3DLDF [4871, 4870]. 3dplot [2577]. 3rd [4990, 4964, 2125, 3112, 4365]. 3 [1310].

4 [4261, 1561, 1715, 2049, 2903, 3309, 3442, 3890, 2253, 2742]. 4-volume [1110]. 4/2003 [2903]. 4/2010 [3605, 3644]. 4/2011 [3719, 3778]. 4/2012 [3812]. 4/2013 [3890]. 4/2014 [4031]. 4/2015 [4106]. 4/2016 [4260]. 4/2017 [4380]. 4/2018 [4445]. 4/2019 [4581]. 4/2020 [4697]. 4/2021 [4791]. 4/2022 [4913]. 41st [4594]. 425-433 [1686]. 43 [4872]. 4A [3642]. 4th [2844, 744].

5 [4577]. 5.0 [870]. 500 [379]. 5th [4255].

6 [237, 2098]. 6.0 [3205]. 64 [765]. 64-bit [765]. 640 [107]. 690 [4299]. 6th [3888, 666, 694].

7 [3381, 1384, 2613, 2049]. 7500 [266].

8 [4657, 4695, 2049, 2130]. 8-bit [1517, 4652]. 80th [4361]. 810 [513]. ’89 [879].

9 [3563, 2258]. 9#3 [875]. 9211 [584]. ’94 [1689]. ’94 — Announcement [1688]. ’95 [1899, 1883]. ’95 — St [1725]. ’96 [1914, 1882]. ’97 [1896]. ’97 — Soliciting [1880]. ’98 [2187, 2145]. ’99 [2309, 2345, 2380].

= [4954, 2725, 4028, 4873].

A&M [1133, 375]. A-compliant [4065]. A-in-L^{A}TEX [1577]. A-R [3433]. A-series [588]. A. [50]. a.k.a [4439]. A/S [37]. a2ac — Font [1810]. A4 [1047]. abaci [3371]. ABC [720]. Abort [4020]. Abrahams [1653, 1333]. abroad [4609]. Absolute [2665]. abstract [1992, 1613]. Abstracts [2686, 2937, 3072, 3106, 3145, 3210, 3274, 3346, 3462, 4734, 4515, 4304, 3430, 3534, 4070, 4831, 4404, 4622, 4957, 4305, 1790, 2702, 2705, 2709, 2688, 2693, 2698, 2966, 3669, 3742, 3887, 3960, 4189, 2325]. academics [3142]. accents [4937, 3238, 2694, 2694]. acceptable [2086]. acceptance [624]. accepted [3247]. access [4425, 2691, 640, 4060]. Accessibility [4506, 4952]. accessible [4436, 4565, 4981]. accidentals [4281]. accommodate [4511]. according [2021]. account [4457]. Accountant [834]. accurate [2894]. Acknowledgement [432, 567, 329]. Acknowledgements [1727]. Acknowledgment [447]. ACM [456, 4179]. Acorn [1250]. Acrobat [2330, 2315, 1829]. AcroDVI [2330]. across [1381, 4465]. AcroTEX [2315]. action [1830]. Active [1536, 1879, 4828, 2327]. activities [1150, 3163]. activity [303, 376, 375]. actuarial [4336]. actuarialangle [4336]. actuarialsymbol [4336]. Ada [873]. Adams [2698, 3145]. adapt [1047]. adaptable [534]. adapted [1440, 527, 767, 575]. Adapting [3739, 2797]. Adaptive [1731, 1580]. Addenda [493, 454, 1686]. Addendum [775, 1518]. addicts [2198]. Adding [4859, 915, 2199]. Addison [3218, 4769, 3535, 355, 394, 435, 511, 517]. Addison-Wesley [3218, 4769, 3535]. Addison-Wesley — Computers [511]. Addison-Wesley — MicroTEX [355, 394, 435]. Addison-Wesley — TEX [517]. addition [546]. additional [2160, 4378, 1028, 3983]. additions [3944]. Address [3083, 2851, 2852]. Addresses [1697, 842, 890, 932, 1020, 1062, 1140, 1299, 1397, 1435, 1504, 1546, 1573, 1642, 1666, 1698, 1773, 1799, 1823, 1865, 1895, 1956, 1985, 2011, 2036, 2085, 2118, 2144, 2219, 16, 52, 99, 2247, 2276, 2359, 2392, 2417, 2495, 2564, 2644, 137, 160, 187, 223, 255, 287, 401, 445, 484, 519, 566, 622, 675, 712, 758, 797, 325, 366]. Adelaide [4739]. adequately [3959]. adjunct [1406]. Administrative [4144]. Adobe [4550, 2690, 1843, 1757, 2544]. Adrian [4084]. Advanced [1967, 3343, 2986, 1735, 1675, 2786, 4131, 1113]. Advances [4180]. Advantages [3529]. adventure [4302, 4604, 629]. Adventures [2163, 4058]. Advertisements [616, 928, 979, 4838, 4965]. advertisers [617, 672, 1059, 888, 980, 1017, 1057, 1097, 1137, 1190, 1230, 1347, 1376, 1394, 1429, 1466, 1500, 1540, 1566, 1590, 1634, 1662, 1696, 1723, 1772, 1837, 1890, 1923, 1980, 2006, 2025, 2055, 2133, 480, 559, 618, 673, 708, 754, 794, 838]. Advertising [4325]. æsthetically [3847]. aesthetics [208, 1203]. affecting [648]. Africa [1578]. after [3233]. afternoon [598]. again [1879]. against [3470]. age [3526, 3759, 3327]. agendas [3671, 1556]. Agfa [2554]. Agfa/Eaglefeather [2554]. agony [3525]. ahead [1170]. aid [1474]. Aided [1403, 2545]. Akhmadeeva [3346]. al [2705, 3072, 3346, 1675]. al. [2711, 2866, 2945]. Alamos [1830, 761]. Alan [2152]. Alban [2336]. AlDraTex [4783]. Aldus [3053]. Aleph [2862, 3034]. Alexander [1705]. algebra [2893, 4023]. algorithm [2463, 4569]. Algorithmic [4028]. algorithms [4288]. Alhambra [4980]. ALI [2791]. ALI-BABA [2791]. alia [133]. Aligning [3561]. alignment [2578]. alignments [4986]. alike [4814]. Allen [2937, 3210, 4544]. allocation [88]. Allographic [2768]. Almost [3760]. alone [2690, 1840]. along [3125, 4096]. alphabet [2839, 4855, 1219, 945, 1218, 2985, 2365, 603, 1116, 1515]. Alphabetgeschichten [3165]. alphabets [4601, 1928, 3628, 2804, 3747]. AlQalam [3091, 3246, 3337]. also [1328, 1420]. alternative [1960, 1945]. Alternatives [2165, 2331]. Alto [4491]. amateur [3131]. Amelia [4681]. America [3925]. American [2532, 3647]. Amiga [1251, 728, 497]. AmigaTEX [961]. AmigaTEX… [869]. Amiri [3698]. AmiWeb2c [2123]. AMS [73, 33]. AMSFonts [1559]. amsmath [4874]. amsrefs [2454]. Amsterdam [2015, 1989]. AMS-TEX82 [227]. amsxport [2454]. anaesthesia [2189]. Analogy [2347]. analysis [4539]. analytical [2728]. Anatomy [41, 2511]. Anchored [1554]. anciens [2772]. Ancient [3360, 2765, 1670, 2834, 1514, 3094, 599]. André [1522]. animated [3083]. Animating [4692]. Animations [2821, 2095]. anniversaries [4896]. annotated [4902, 964]. annotations [3423, 2895]. Announcement [2639, 3441, 3442, 3443, 3509, 2846, 2845, 4349, 4385, 3111, 3147, 3148, 3186, 3187, 3216, 3315, 3349, 3434, 3613, 3649, 3687, 3781, 3860, 3936, 3965, 4003, 4036, 570, 61, 72, 1856, 208, 215, 424, 2760, 2972, 2973, 2570, 1576, 2135, 2175, 2210, 2237, 2260, 2445, 2481, 2517, 2556, 2588, 2590, 2638, 2679, 2750, 2908, 3691, 3815]. announcements [53, 100, 138, 161, 188, 224, 256, 288, 402, 326, 367]. Announcing [845, 904, 1044, 1262]. Annual [442, 709, 1192, 1193, 1766, 2307, 2384, 2393, 2436, 2528, 2601, 352, 368, 443, 663, 749, 840, 3112, 4927, 1022, 4492, 4712, 2518, 4190, 4166, 4804, 522, 4793, 4365, 4594, 4393, 4271, 4390, 1372, 1393, 1496, 1635]. Anon [2489]. ANSI [523, 142, 231, 328]. Answer [2499, 155, 1132]. answered [2649]. Answers [11, 132, 1899, 1990, 1800, 937, 1147, 1456, 1320]. Antipolis [609]. Antykwa [2190]. Anwendervereinigung [924, 1119]. anything [3084]. APA [601]. apa6 [3703]. APE — A [873]. APL} [686, 1176]. apnum [4142]. Apocalypse} [2337, 2336]. appearance [3232]. Appendix [50, 3069]. Apple [2713, 2875]. application [2842, 136, 159, 186, 221, 254, 286, 321, 358, 397, 670, 1623, 3242, 3764, 4607, 887, 927, 978, 1015, 1056, 1095, 1136, 1189, 1346, 1428, 1465, 1541, 1568, 1636, 1637, 1663, 1694, 1724, 1771, 1795, 1819, 1921, 1978, 2004, 2030, 2058, 2110, 2138, 2212, 2239, 2268, 2302, 2350, 2406, 2438, 2473, 2484, 2521, 2593, 2677, 2867, 439, 509, 560, 614, 707, 752, 793, 837, 2189]. applications [1757, 1387, 4099, 1229, 1374, 1552, 2289]. Applying [3766]. appreciation [1412, 3642]. Apprendre [3993]. apprentices [3401]. approach [526, 2728, 3445, 575, 1207, 4134, 2701, 1585, 2097, 1942]. approaches [4575]. Approaching [3740, 1458]. April [4579, 1339, 389, 611, 2148, 3856, 4295, 4374]. APS [69]. APS-5 [69]. arabe [2769]. Arabi [3089]. Arabic [1757, 2890, 241, 4754, 2823, 3088, 4722, 2822, 3090, 3091, 1153, 2018, 3092, 3985, 3089, 3768, 3103, 2769, 4511, 3102, 3246, 3337]. arara [4024, 4693]. Arara — TEX [4399]. arbitrary [4142]. archaic [2766]. Archimedes [1250]. Architectural [4929]. architecture [2619]. archivable [3700]. Archive [1630, 1007, 810, 2818, 1278, 908, 953, 1033, 3244, 864, 906, 1200]. archive.cs.ruu.nl [1275]. archives [1277, 1830]. area [3712]. areas [2151]. Arecibo [4979]. argument [3468, 4096]. Arguments [4101, 3673]. arithmetic [3202, 3762]. Arora [4169]. array [827]. array — and [875]. arrowhead [4137]. Arrows [1441]. ArsTEX [3108, 3180, 3309, 3381, 3491, 3563, 3606, 3681, 3744, 3779, 3856]. ArsTeXnica [4958, 4579]. Art [3640, 4577, 3642, 2608, 3230, 4438, 2929, 4330, 2645]. Artful [3922]. Arthur [1146]. articles [135, 158, 4062, 3836]. artificial [4981]. artificial-intelligence [4981]. artificiels [2804]. Artikel [972, 1049, 1089, 1117, 1179]. artist [4893]. Arts [4427, 3133]. Arvind [1521, 1551]. ASCII [724, 4438, 776]. ASCII.sty [1704]. Asian [3493]. ask [62, 62]. asked [3176]. Asking [4326]. aspect [718]. Assemblies [4127]. Assembling [421]. assignments [4128]. Assistant [2461, 562, 563, 619]. asterisk [4221]. asterisks [1323]. AsTEX [2461]. Aston [864, 906, 1496, 1563, 1007, 1265, 1278]. Astrological [4178]. astronomical [1556]. Astronomy [1522, 1338]. Asymptote [3294, 3368, 3527, 3740, 4943]. Atari [803, 1252, 536, 693, 811]. A & [204]. ATILF [2772]. attempt [2796, 4688, 4574]. attendee [4738, 3726]. attendees [60, 1858, 2078, 2188, 2345, 2472]. Attention [571]. attribute [2793]. Attributes [4899]. ATypI [208, 4120]. AUCTEX [3261]. audio [1489, 1852]. Auditing [2800]. Audrey [1905]. August [709, 1016, 1053, 1688, 2145, 2649, 2712, 2719, 2806, 2865, 3010, 352, 368, 663, 749, 840, 2518, 314, 2308]. Aussois [552]. Australian [571]. AUTHIDX [2124]. Author [4621, 2500, 4816, 2124, 1359, 2935, 3739, 3868, 962, 386]. author-to-publisher [1359]. author/editor [2124]. authoring [2610, 2532, 3263, 3663]. Authors [325, 366, 16, 52, 99, 137, 160, 187, 223, 255, 287, 401, 2546, 4504, 1361]. auto [3513]. auto-pst-pdf [3513]. autodoc [920]. Autolayouter [1368, 1103]. AutoLetter [592]. AutoLoan [2071]. Autologic [992]. automata [4900, 1624]. Automated [4950, 3263, 3062, 2830, 3033, 3021, 656]. Automatic [466, 2893, 3059, 4335, 3135, 2695, 88, 747, 1476, 1377, 1733]. Automatically [4784, 4872, 4725]. Automating [4098, 4508]. automation [4399, 3733]. automatisé [2771]. autostereogram [1786]. Availability [77]. available [1092, 315, 353, 393, 428, 1031, 1074, 3697, 1850, 1306, 36, 940]. avanzati [1675]. avez [4145]. Avoid [3701, 1040]. awarded [759, 800, 4408]. aware [3527, 2452]. away [3140, 3550]. AWK [1196]. Axessibility [4436]. Axis [2578].

B [1401]. BABA [2791]. Babel [1870, 1329, 2076, 3093, 2230, 1558, 3178, 2459, 2838, 1879, 2022]. babelbib [2666]. BachoTeX [4274, 4305, 4030, 4529]. Back [4491, 2642, 2717, 2851, 1505, 3340, 4616, 1833]. backend [4135, 3699]. background [4602]. backward [4679]. bad [4171, 1384, 4285, 4683]. bag [1318, 1413, 1450, 1525, 1553, 1655, 1681, 1785, 2401, 2512]. Balancing [94, 747]. Bangla [4572, 4541]. Banquet [4045]. bar [3846, 917]. Barbara [1635, 1688, 4275, 2335]. barcodes [3709, 1782, 2097]. barrier [3325]. Barriers [4028]. Barry [3829]. Bärs [4728, 549, 876]. base [3711, 4208]. based [810, 580, 750, 4721, 3063, 3062, 3058, 2453, 3525, 3844, 4182, 1763, 3882, 4987, 2654, 1614, 2833, 4828, 695, 2826, 2332, 4509, 3290, 2542, 2799, 1550, 3665, 3729, 1937, 2325, 4464, 1490, 1944]. Bashkortostan [4057]. Basic [4330, 1911, 1415, 3045, 916, 4246]. basics [3169, 3211, 3275, 3298, 3487, 3635, 3806, 3884]. Baskerville} [3382, 3497, 1460, 1713]. BaskervilleF [4242]. Basque [2837]. Batch [2934]. Bausum [2523, 2559, 2595, 2625, 2580]. Bay [2747]. Bazargan [2705, 3534, 3669, 3960, 4070, 4189, 4191]. BCS [454]. be [4271, 4390, 1563, 3959, 3552, 4564, 1370, 4285, 1371]. Beamer [2933, 4019, 4934, 4126, 3911, 3912]. beauties [3272]. beautiful [2328, 3471, 4620]. Beccari [1444]. Bechtolsheim [1674]. beckon [1505]. becomes [799, 1711, 1652, 3035]. Becoming [3405]. beer [4280]. Beeton [3106, 2335]. before [4977]. Beginner [4788, 4782, 4894, 3641, 2632, 3137, 4457, 4429, 2618, 4785, 1520]. beginners [4329, 3031, 1940, 606]. beginnings [4323]. Begriffsschrift [4100]. behavior [2768]. Behind [4941, 4433]. belatedly [904]. Beletsky [4256]. Belleek [2191]. belly [3070]. benches [1505]. Benefits [3266, 3468]. Bengali [2653, 2290]. Benson [584]. Benson-Varian [584]. Bernard [3278]. Berry [1653, 3960]. better [4606]. Between [3195, 3256, 2621]. Beyond [4562, 2828, 4599, 3485, 2076, 3468, 4019, 1111]. Bézier [3712]. Bezos [4597]. bib [4140]. bib2gls [4469, 4647, 4732, 4786]. Bibby [2709]. Biber — the [3699]. biblatex [3980, 3699]. biblatex — Request [4117]. Bible [1310, 1730]. Biblet [2936]. bibliographer [2860]. Bibliographic [965, 1900, 1686]. bibliographies [2666, 4733, 3666, 4299, 3366, 1656, 3636]. Bibliography [1610, 1651, 3044, 1075, 1161, 4442, 3206, 3175, 2789, 3332, 2936, 1032]. bibliography-style [4442]. BibTeX [2789, 3068, 3105, 2787]. BibTEX-based [4987, 4509]. BibTEXing [2788]. Bibulous — a [3882]. Bibulus — a [2787]. bibview [1650]. Bids [1896, 1880]. big [206, 37, 2731]. Bigelow [4189, 3839]. bigfoot [2832, 3046, 3266]. bindings [4066]. Biography [4543]. bioinformatics [4288]. biological [3518]. biometrics [2768]. bird [4024]. Birmingham [1563]. birth [2330, 2468, 3171]. birthday [3323, 4361]. bit [1384, 1517, 4652, 1682, 333, 37, 765]. bitmap [2571]. bitmaps [1308, 1385]. Bitnet [640]. bits [1384]. Biuletyn [2904, 3109, 3427]. Blackboard [721]. Blacker [632]. blank [4096]. blight [1384]. blind [311, 4981]. Block [2068]. BLUe [1945, 2182, 2271, 2305, 2354, 2410, 2492, 2560, 2626, 2084, 2117, 2185, 2216, 2244, 2274, 2306, 2356, 2391, 2413, 2444, 2478, 2488, 2524, 2563, 2597, 2629]. blueprint [4645]. BLUes [1379, 1557]. BM2FONT [1385]. Board [1690, 1717, 2265, 2646, 2566, 2335, 1919, 2028, 2136, 4082, 4121, 1783]. Bodoni [4146, 3678, 4147]. Bold [721]. Bonn [4929, 4926]. Book [3639, 3851, 3992, 3993, 2711, 2844, 4260, 4306, 4261, 3714, 3962, 4543, 4544, 4992, 3994, 4659, 4255, 4910, 4990, 4964, 4911, 4788, 3894, 4275, 3995, 4258, 4256, 4259, 4308, 4410, 4475, 4660, 4696, 3599, 3640, 3641, 3678, 3679, 3715, 3747, 3776, 3811, 3849, 3850, 3852, 3853, 3930, 3931, 4029, 3303, 4309, 4257, 4578, 3716, 3777, 3893, 4028, 4345, 4991, 923, 580, 3288, 3411, 4723, 3516, 2856, 1729, 2321, 4929, 1358, 134, 3637, 1739, 4786, 3231, 3956, 4147, 1360, 4139, 1653, 1444, 678, 1445, 573, 1446, 2129, 2377, 3328, 4345]. Bookbinding [4282]. book — A [2731]. bookdb [4139]. booklet [2931]. Books [1160, 3976, 1076, 1409, 3880, 3626, 1207, 1410, 1904, 3083, 3551, 3537, 1528, 2368, 1582]. bookshelf [4602]. bookstore [2291]. Borde [1521, 1551]. Boris [4807]. Boston [1192, 1184, 1193, 665, 422, 3728]. bot [4186]. Both [4461, 3231]. bother [1854]. bottle — Improving [3340]. bottles [3626]. bottom [3979, 4937]. box [1170, 4023, 783, 986]. boxed [1969]. Boxes [3913, 963, 2897, 3774]. brace [4904]. Brackets [3084]. Braille [2961, 4460, 3872, 4727]. Brazilian [4537]. Breaking [2072, 1169, 2455, 4067]. breed [4699]. Breitenlohner [4117, 4118]. bridge [2546, 878, 1083]. Bridging [2704]. Brief [29, 2909, 4544, 4167, 4014, 3243].

Bright [4868]. brilliant [4544]. bring [2529]. Bringhurst [4189, 4308, 4187]. Bringing [4467]. brings [3981]. Brown [34, 120]. browser [4461]. Bruce [4410]. Bubbles [128]. budget [434]. bug [3920, 3923, 647, 1180]. bugs [4561, 4373, 170, 2676, 550, 544]. build [210, 4718, 1607]. builder [3522]. building [3103, 532]. buildrel [423]. Bujdosó [3106]. Bulletin [1623, 2818]. Bumpy [4947]. bundle [2832, 3046, 4298, 4101]. bundle — a [3271]. Burbank [3579]. Burt [2937]. business [163, 2294]. Bylaws [228, 1185, 1345, 1917]. bytes [2965]. Byzantine [2962, 2987]. ByZL^{A}TEX [2987].

C [4805, 852, 737, 733, 1905, 1476, 575, 766, 4371, 2826, 527]. C. [4054]. cabinet [3881]. Cage [4854]. Cahiers [4764, 1912, 1998, 2048, 2099, 2171, 2233, 2296, 2379, 2515, 2587, 2667, 2743, 2905, 3002, 3181, 3494, 3778, 3997, 1335, 1422, 1459, 1534, 1588, 1714, 1789]. Calculating [4930]. calculations [3762, 4471, 4472]. Calculator [2201, 3762]. calculus [2691, 3762]. Calendar [1791, 2840, 4110, 4267, 4312, 4350, 4386, 4416, 4448, 4483, 4520, 4552, 4587, 4626, 4665, 4705, 4741, 4770, 4797, 4837, 4881, 4920, 4968, 4995, 880, 921, 1051, 1090, 1134, 1181, 1223, 1336, 1390, 1423, 1463, 1495, 1535, 1562, 1589, 1633, 1658, 1687, 1716, 1767, 1792, 1815, 1834, 1857, 1881, 1913, 1948, 1973, 1999, 2023, 2054, 2077, 2106, 2131, 2172, 2207, 2236, 2259, 2297, 2346, 2381, 2402, 2431, 2470, 2482, 2516, 2555, 2589, 2622, 2633, 2668, 2710, 2749, 2805, 2864, 2907, 2938, 2967, 3005, 3073]. Calendar [3110, 3146, 3185, 3215, 3314, 3348, 3393, 3435, 3501, 3538, 3568, 3612, 3648, 3686, 3723, 3751, 3783, 3814, 3859, 3897, 3935, 3966, 4004, 4037, 4076, 4111, 4155, 4195, 4225, 310, 350, 388, 431, 446, 507, 551, 612, 664, 704, 748, 789, 828, 973]. California [1688]. Call [709, 1192, 1976, 2105, 453, 543, 522, 3272, 2191, 1053, 1182, 1184, 1193, 1339, 1536, 1689, 1882, 2062, 312, 611, 830, 840]. Called [4682]. Calligraphicum [4258]. Calligraphy [3850, 207, 4275]. Calls [973]. Calma [203, 122]. cals [3632]. CalTech [119]. Cambridge [2301, 2352]. Camel [1900]. Cameron [1560]. Can [3959, 4873, 3552, 1946, 62]. cancelled [1918]. Canon [70]. can’t [2565]. canvas [4161, 4400]. capabilities [3143]. capability [1381]. capitals [4938]. Cappelli [2962]. captions [1969, 4756]. carborundum [4283]. care [4619, 3266]. career [4544]. Carleton [2873, 2941, 3219]. Carol [4544]. Carolina [2808, 2880]. Carter [4890]. cartoon [4265]. Cary [3639, 4427]. Case [3873, 4569, 4823, 610, 4288, 4978, 2818, 2707, 3847, 641, 2837, 2121, 3986, 1555]. Caslon [2815]. Cassetin [2765]. Castro [4204]. Catalan [1671, 1855, 1616, 1672]. catalogs [4611]. catalogue [3869, 4139, 2400, 2399]. catch [3334]. catcher [4068]. Catching [4810, 2614]. category [3533]. causes [682]. CD [2414, 3447, 2291, 2169]. CD-ROMs [3447]. CDC [9, 458, 496, 537, 268, 144]. CD — The [3061]. CDROM [2149]. CDs [4782, 2443]. ce [4145]. celebrate [488]. celebration [4361]. Centaur [4256]. Center [900]. centered [4096]. centering [877]. Centre [2873, 2941, 3219]. century [4240, 4307, 4327, 4811, 3640]. CEO [4716]. ch [4307, 4327]. Chairman [18, 55, 3, 101]. challenge [155, 3247, 4457, 2464, 824, 2766, 858]. challenges [3529, 3771, 3703]. chances [624]. Change [549, 718, 1455, 1720]. Changed [3962, 2500, 4639]. changefiles [468, 526, 629]. changes [469, 503, 4135, 3520, 4511, 4289, 4257]. Changing [3761, 4569, 4823, 1436, 3488]. Chapel [2808, 2880]. Chapter [545]. Character [1066, 3291, 4692, 1731, 949, 3517, 1449, 1524, 4513, 2697]. characteristics [2288]. characterization [29]. Characters [4345, 4978, 1879, 988, 4652, 1704, 915, 2801, 4863, 2094]. characters — [2791]. Charles [1990, 3839]. chart [917, 192, 235, 263, 299, 334, 372]. Charting [175, 129]. charts [493, 154, 4902, 3846, 4641, 860, 1156, 638, 687, 723, 774, 808, 2513, 4178, 410, 581]. ChatGPT [4963]. Cheats [4954]. checkers [735]. checking [4816, 1610, 1651, 3136, 3410]. Chemical [602, 1814, 2572, 1910, 130]. Chemistry [4097, 3880]. chemistry/mathematics [3880]. cherry [4375, 4470]. Cheryl [2874, 2947, 3219, 3432, 3536]. Chess [950, 2466, 821, 947, 964, 1155, 1417, 898]. Chest [2232, 2257, 4254, 4310, 4342, 4379, 4411, 4444, 4478, 4547, 4662, 4582, 4625, 4700, 4737, 4762, 4789, 4830, 4875, 4912, 4988, 3174, 3302, 3380, 3490, 3562, 3600, 3643, 3680, 3717, 3746, 3775, 3809, 3842, 3891, 3929, 3991, 4027, 4103, 4143, 4220, 2901, 2996, 3126, 2374, 2509, 2582, 2662, 2740, 2429]. Cheswick [3534, 3742]. Chi [1628]. Chi-Writer [1628]. child [1946]. Children [4278]. China [2712, 2806, 2865]. Chinese [298, 2959, 1549, 2801, 2957, 3517, 3870, 3299, 3371, 2697]. CHM [4809]. Cho [2702, 2966, 3430, 3887]. choice [1319]. choix [2772]. choose [3769]. chosen [2448]. Christian [3993, 1707]. Christmas [4941]. Church [1871, 4604]. Church-Slavonic [1871]. Cincinnati [157, 61, 104]. circ [2664]. CIRC — A [2068]. circle [1068]. circles [2923, 3372]. circuit [1207, 2664]. CircuiTikZ [4725]. Circular [1068]. Citation [3990, 2100, 4509]. citations [1606, 965, 4140]. CJK [4514, 2827, 3874, 2076]. Clarkson [908, 953, 1033, 1266]. clashes [2900]. class [132, 155, 3671, 3703, 2318, 2340, 3733, 4908, 3667, 4095, 4126, 3977, 4935, 3401, 3177, 2900, 3912, 3140, 2341]. class/package [2900]. class — our [3324]. Classes [3518, 4826, 1742, 4062, 4395]. classical [2652, 3360, 4403]. classics [2684]. classicthesis [3471]. Classified [282]. Claudio [1444, 4189]. cleaning [4289]. Clinic [3245]. clone [421]. close [2773]. closer [3364]. closing [4686]. cloud [4620]. Clumsy [4412]. CM [1902, 2695]. CM-Super [2695]. CMS [1256, 1034, 644, 731, 813, 725, 1073, 1131, 1164, 377, 539]. co [2771]. co-rédaction [2771]. Co. [2815]. COBOL [1013]. Cochineal [4174]. code [3711, 4367, 88, 2830, 3033, 4864, 3623, 2451, 899, 266]. coded [105]. codes [176, 1516, 949]. coding [3668]. collaboration [3293, 2503]. collaborations [2858]. Collaborative [3672, 4247, 4608]. Collecting [4323]. Collection [2971, 1262, 1475, 874, 825, 3031, 593, 655, 700, 744, 786, 3750, 4427, 2863, 907, 1162, 3639, 2443]. collection — submissions [1077]. collections [2772]. college [3133]. Colophon [3820]. Color [1765, 1734, 1442, 4246, 4397, 1739, 3886]. Colorful [4212]. colors [4824]. Colour [1177, 1737, 1738, 1085, 1736]. colourful [3054]. colours [3557]. Column [453, 652, 179, 1012, 1451, 1178, 346, 1043, 1084, 1175, 1325, 1419, 701, 745, 785, 826, 32, 385, 1326]. columns [3979, 4296, 94, 2224]. combination [2449]. combine [1484]. Combining [4499, 1030, 1071]. combs [4942]. come [1449, 1524, 2769, 3297]. come — Finally [2449]. comes [3526, 3759]. comfortable [958]. Comic [4368, 4479, 4583, 4766, 4699, 4548]. Comics [4343, 4412]. Coming [885, 1187, 1342, 1426, 1462, 1538, 1565, 1591, 1660, 1692, 1719, 1817, 486, 954]. comma [2769]. command [428, 4066]. Commander [2934]. commands [2100, 374, 4683]. commedit [4507]. Comment [660, 1326]. commentary [2284]. commented [4507]. Comments [1671, 1110, 485, 1672, 248, 283, 3241, 1235, 1279, 4236, 4322, 4358, 4424, 4456, 4528, 4638, 4560, 4673, 4750, 4778, 4847, 4889, 4975, 846, 892, 935, 1063, 1145, 1301, 1400, 1437, 1506, 1548, 1575, 1644, 1668, 1700, 1775, 1802, 1867, 1898, 1959, 1988, 2014, 2039, 2088, 2120, 2147, 2222, 2250, 2278, 2362, 2395, 2420, 2497, 2567, 2647, 2722, 2884, 2979, 3081, 3121, 3159, 3283, 3357, 3467, 3547, 3578, 3618, 3693, 3758, 3793, 3828, 3904, 3973, 4012, 4083, 4117, 4203, 763, 3213, 4731, 3601, 1144, 2479, 2631, 2598, 489, 1407, 2, 2920]. Comments — A [2187]. Comments — TUG [2446]. commercial [652, 762]. Committee [19, 56, 14, 79, 1157, 291, 228, 290, 861, 903]. commodity [4607]. Commodore [728]. Common [4498]. communication [311]. communities — discussion [3656]. community [2606, 4619, 4244, 905]. commutative [4377, 1784]. Como [351]. comp.text.tex [1280]. Compact [820, 4782]. Companion [2741, 4964, 3716, 2368, 1873, 2866, 2945, 4963, 2091, 2711, 3218, 1705, 4990, 1729]. Company [4029]. comparative [3666, 1165, 1523]. Comparing [1356]. Comparison [4685, 2824, 3067, 3240, 4217]. comparisons [1313, 259, 313, 391]. compatible [1583, 1030, 498]. compatibles [422, 1035]. compatibles — MS [1257]. compatibles — MS-DOS [1257]. competition [845]. compilation [29, 30, 1609, 4020]. compilations [429]. compilers [26]. complement [3591, 2584]. complementi [1675]. Complete [4009, 3930, 958, 3136, 3410, 1035]. completely [4280]. completeness [4816, 1929]. completion [4066]. complex [3097, 3124, 1761, 4500, 4956, 3172]. complex-script [3124]. compliant [4299, 4065]. complicated [1607]. compliments [2149]. component [3201]. composing [3424]. Composition [514, 4620, 1356]. compositor [3131]. compound [1849]. Comprehensive [1630, 3595]. compuscripts [1762]. Computas [37]. computation [4898, 1746, 3373]. computations [1556]. compute [541]. Computer [353, 393, 422, 2331, 2549, 2694, 532, 633, 207, 3811, 4577, 3642, 991, 208, 2451, 2893, 994, 540, 3469, 3952, 2545, 695, 4809, 37, 514, 1403, 1931, 1067, 2165, 858, 4209, 1991]. computer-aided [2545]. computerized [4511]. computerizing [4531]. Computers [2676, 486, 487, 488, 192, 235, 263, 299, 334, 372, 410, 1583, 1191]. Computing [3712, 68, 3661, 4247, 3460]. concept [4610, 4395]. conception [2468]. concepts [4852, 2791, 1675]. conclusion [4332]. concordance [2227]. Concret [4330]. Concrete [944, 855]. ConcTEX [2227]. Concurrent [1747]. Conditional [3487]. conditionals [651]. Conf [4426]. Conference [1600, 3322, 4349, 1727, 2763, 2849, 3050, 3086, 3129, 3225, 3321, 3399, 3507, 3654, 3725, 3865, 3941, 4042, 4162, 389, 2846, 2845, 3020, 3193, 3866, 4043, 4163, 4273, 4271, 4390, 4488, 4489, 4592, 4710, 4802, 4925, 1053, 1339, 2712, 2719, 3441, 3443, 2848, 4563, 4803, 4046, 4713, 4272, 4391, 4711, 3508, 1790, 2158, 2307, 2384, 2393, 2436, 2591, 2683, 2925, 2954, 2956, 351, 451, 2762, 374]. Conference — TUG [4594]. conferences [4596, 3843, 974]. Confessions [3229]. Configuring [1805]. conflicts [3518]. conforming [2452]. Congrès [881, 1183, 1337, 893]. congress [1211]. conjunction [1563]. conservation [2764]. Conserving [2621]. considerations [4621, 688, 2776]. considering [4437]. consistently [4336]. constitutes [697]. constructed [1116]. Constructing [2664]. construction [4871]. consultant [2163]. Consultants [1499, 1638, 2939, 1375]. consulting [2843, 4266, 4313, 4348, 4383, 4413, 4446, 4480, 4518, 4551, 4664, 4585, 4630, 4704, 4740, 4767, 4795, 4839, 4879, 4993, 4917, 4966, 1234, 1430, 1542, 1569, 1594, 1664, 1695, 1722, 1770, 1796, 1820, 1836, 1860, 1889, 1922, 1950, 1979, 2005, 2031, 2059, 2081, 2111, 2139, 2177, 2213, 2240, 2269, 2303, 2351, 2387, 2407, 2439, 2474, 2485, 2522, 2558, 2594, 2624, 2635, 2678, 2715, 2755, 2810, 2869, 2914, 2950, 2970, 3007, 3076, 3116, 3153, 3188, 3220, 3277, 3316, 3350, 3394, 3436, 3464]. consulting [3502, 3539, 3569, 3607, 3646, 3684, 3720, 3727, 3748, 3784, 3816, 3854, 3896, 3934, 4002, 4035, 4075, 4109, 4154, 4194, 4224, 3584]. contacts [1247]. containers [4945]. containing [1277, 2451]. Contemporary [2794]. Content [4728, 3918, 4857, 4981, 3656, 3290, 3095]. Contents [2757, 2812, 2877, 2915, 2951, 2974, 3014, 3077, 3117, 3154, 3189, 3221, 3317, 3352, 3395, 3437, 3503, 3542, 3572, 3614, 3650, 3688, 3752, 3786, 3743, 3821, 3861, 3898, 3937, 3967, 4005, 4077, 4112, 4071, 4157, 4196, 4226, 4261, 4231, 4270, 4317, 4262, 4347, 4353, 4389, 4419, 4407, 4405, 4381, 4451, 4486, 4523, 4546, 4516, 4477, 4579, 4633, 4555, 4590, 4580, 4764, 4668, 4708, 4745, 4698, 4773, 4800, 4842, 4884, 4923, 4971, 2099, 2171, 2233, 2296, 2379, 2515, 2587, 4040, 4222, 2487, 2048, 2667, 2743]. contents [2744, 2745, 2902, 2903, 2904, 2905, 2906, 2976, 3016, 3107, 3108, 3109, 3180, 3181, 3182, 3183, 3306, 3307, 3308, 3309, 3310, 3381, 3382, 3383, 3384, 3385, 3386, 3387, 3388, 3427, 3428, 3429, 3492, 3493, 3494, 3495, 3496, 3497, 3498, 3491, 3563, 3564, 3565, 3566, 3567, 3602, 3603, 3604, 3605, 3606, 3644, 3645, 3681, 3682, 3683, 3718, 3719, 3744, 3745, 3779, 3780, 3813, 3812, 3855, 3856, 3889, 3890, 3927, 3928, 3996, 4031, 4072, 4106, 4149, 4150, 4151, 44, 966, 3774, 4096, 3000]. Contents [3001, 3002, 3003, 3004, 3778, 3997, 556, 874, 825, 593, 655, 700, 744, 786]. Contest [1577, 1800, 2248]. Context [1559, 1606, 4437, 3888, 4765]. context-sensitive [1606]. contextgarden.net [3029]. contextually [1580]. continues [629]. continuing [4119]. Continuous [4719]. contours [1841]. Contra [1357]. Contra-L^{A}TEX [1357]. contradictions [4403]. contributed [2321]. contributions [432, 447, 567, 696, 329]. contributor [4286]. control [1448, 1415, 1704, 1173, 4784, 4872]. controlled [1683, 4060]. Controlling [4756, 780, 961, 3912]. Convenient [2329]. Conventional [3259]. conventions [802, 1855]. conversation [4890]. conversion [3209, 1377, 3256, 2833, 1733, 1811, 2612, 4720, 4437, 2621, 1853]. conversions [1490]. convert [4718]. converter [733]. converters [3957]. Converting [3836, 4337, 2450, 3945, 1841, 2341, 2964]. convincingly [3452]. cookbooks [4090]. cookie [4067]. cookie-cutter [4067]. Cooper [494, 4117]. Cooperating [2348]. coordination [2178, 1024]. coordinator [959]. Coordinators [17]. cope [4564]. Coptic [2889]. copy [2998]. copyediting [3955]. copying [1680]. core [3559, 3647, 4373]. Cork [1151, 3340, 3336]. Coroutines [4372]. Corporation [363, 364]. Corporation — Intelligent [252]. Corporation — Typesetting [514]. Correction [1531]. Corrections [4214]. Corrigenda [330]. Corrigendum [1901]. cost [194]. Counting [3488]. Course [190, 923, 1497, 185, 552, 610, 750, 3451, 297, 92, 227]. Courses [1563, 452, 4117, 1851, 320]. courseware [2453]. Cover [2242, 2272, 2389, 2487, 2411, 2441, 2476, 2493, 2561, 2596, 2627, 481, 755, 1231, 1348, 1352, 1232, 1349, 1353, 1431, 1432, 1467, 1501, 1543, 1570, 1595, 1639, 1665, 1821, 1861, 1891, 1822, 1862, 1892, 1924, 1952, 1981, 2007, 1925, 1953, 1982, 2008, 2032, 2060, 2083, 2115, 2033, 2061, 2116, 2142, 2183, 2215, 2143, 2184, 2243, 2355, 2273, 2390, 2412, 2442, 2477, 2494, 2562, 2636, 2641, 2642, 2628, 2637, 2643, 2716, 2758, 2813, 2717, 2718, 2759, 2814, 2878, 2916, 2879, 2917, 2952, 2975, 3015, 3078, 3118, 3155]. cover [3190, 3222, 3156, 3191, 3223, 3279, 3318, 3353, 3280, 3319, 3354, 3396, 3438, 3397, 3439, 3504, 3543, 3573, 3505, 3544, 3574, 3615, 3651, 3689, 3616, 3652, 3690, 3753, 3787, 3754, 3788, 3822, 3862, 3899, 3823, 3863, 3900, 3938, 3968, 4006, 3939, 3969, 4007, 4078, 4113, 4079, 4114, 4158, 4197, 4227, 4159, 4198, 4228, 4229, 4268, 4315, 4230, 4269, 4316, 4351, 4387, 4417, 4352, 4388, 4418, 4449, 4484, 4521, 4450, 4485, 4522, 4553, 4588, 4631, 4554, 4589, 4632, 4666, 4706, 4743, 4667, 4707, 4744, 4771, 4798, 4840]. cover [4772, 4799, 4841, 4882, 4921, 4969, 4883, 4922, 4970, 4038, 4039]. coverage [4465]. covers [4930]. COVID [4576, 4594]. COVID-19 [4576, 4594]. Cowfont [4213]. CP [237]. CP-6 [237]. CrafTEX [3766]. crafts [3766]. CRAN [4286]. Cray [777]. Crazy [4682]. create [3130, 996, 2796, 2339, 2313]. Creating [4436, 4902, 3846, 4565, 3950, 4906, 2824, 3335, 2799, 4507, 4726, 4656, 4683, 4508, 1623, 4722, 3137, 899, 2894, 4688, 2547, 1014, 2095, 1482, 3127, 2800, 2690]. Creation [3167, 3239, 3327, 3532, 2695, 2795, 1483]. credit [4117]. credits [450]. Cretel [4070]. critical [2585, 2832, 3046, 1174]. critique [2782]. Croatian [1515, 1304, 1903]. cross [3907, 2324, 4647, 4245, 4464]. cross-platform [4464]. cross-references [3907, 4647]. cross-referencing [2324, 4245]. Crossland [3669, 3960]. CrossTEX [3206]. Crossword [3273, 938, 1038]. crosswords [1048]. Crudetype [534]. Cs [1362, 1851]. CS/1 [1851]. CSL [3990]. csname [913]. CSplain [3807]. CSS [4953]. CSS3 [3664]. CTAN [1630, 2291, 2971, 4156, 2552, 3037, 3405, 4244, 4366, 4286, 2443, 2729, 2885, 2257]. CTAN — soul.sty [2232]. CTO [4806]. ctrl [780]. ctrl-M [780]. CTUstyle [4063]. Culture [4308]. cultures [4852]. cuneiform [3335]. Cunningham [3960]. CurExt [2767]. Current [3377, 103, 4292, 2205, 4807, 4331, 3477]. Curvature [4942]. curve [3712]. curved [1388, 2767]. Custom [2071]. customizable [4501, 2936]. Customized [504]. Cut [3673]. Cutout [3773]. cutter [4659, 4067]. cutting [896]. CV [2586]. CVs [4095]. CWEB [2199, 767, 4695, 4720, 3921, 4337]. Cyber [458, 496, 537, 9]. CygnusEd [961]. Cyklop [3234]. cyrillic [371, 409, 2228, 4722, 1938, 1936, 3473, 1928, 1935]. CyrTUG [1238]. Czech [1904]. Czechoslovak [4724]. Czechoslovakia [1225].

D [1287, 4751, 2831, 2772, 4501, 3022, 203]. D.L [4659]. dād [3985]. DadTEX — a [3090]. daemon [2574]. Daly [2686, 2688, 2844]. Daniel [2723]. dans [2772]. DANTE [1182, 4739, 1239, 2291, 4851, 4086, 1119, 2149]. DANTE — Deutschsprachige [924]. d’apprentissage [2803]. Dark [3140]. — [5]. — A [3594]. — addenda [875]. dashed [1844]. — the [1412]. dashundergaps [4441].

— yet [726, 775]. Data [196, 3345, 4650, 3449, 4500, 4335, 4259, 4909, 3478, 1253, 909, 1072, 1128, 270, 414, 457, 587, 642, 692, 729, 778, 812]. database [4611, 2166, 2691, 3064, 1488, 2462, 2323]. database-driven [4611]. databases [3420, 1196]. databases — TEX [2602]. dataset [4509]. datatool [4431]. dateline [1685]. David [50, 2523, 2559, 2595, 2625, 4992, 2580, 4258, 4696, 4805, 1967, 4147]. day [3114].

Daylight [4028]. days [936]. DBI [2602]. dc [1929, 1873]. dc-fonts [1873]. DDA [914]. Deadline [1577]. deaf [311]. Debian [2990, 3255, 4133]. Debrecen [3012]. Debugging [4283, 2202, 2203, 4904]. DEC [1254]. DEC-20 — TOPS-20 [1254]. decade [4780]. December [2296, 2515, 834, 2256, 2375, 4468]. decisions [3204]. DECnet [1273]. decomposition [4692]. Decorating [3447]. decorations [3632]. d’écriture [2777]. DECSystem [72]. DECSystem-10 [72]. DECSystem-10/20 [72]. DECUS [74, 907, 1162, 1262, 376, 1077]. Dedication [4049]. Dednat6 [4438]. Deeply [3476]. Default [2156]. define [3957]. defined [4649]. Defining [2891, 3301]. definitions [741]. degree [800]. DeLand [3210]. delegates [2763, 2849, 2925, 2954, 3050, 3086, 3129, 3225, 3321, 3399, 3507, 3654, 3725, 2683]. delimiters [4185]. delivery — Notes [2378]. demand [2326, 3532]. demander [4145]. demerits [4378]. Demo [2201]. demonstration [4944]. Department [353, 393, 1358]. dependent [1618, 93]. deployment [4948]. depths [780]. description [2448, 2780]. descriptions [465]. descriptive [1754, 3172]. Deseret [2839]. Design [4501, 3219, 609, 1760, 4330, 3204, 2645, 3134, 1536, 208, 4610, 2927, 4433, 4900, 2318, 2856, 3098, 489, 3947, 4810, 4977, 3764, 1737, 4502, 3637, 293, 330, 2129, 2377, 3665, 2503, 2944, 3432, 988, 3715, 610, 2728, 577, 4410, 4544]. Designed [4543]. designer [1202, 4134, 4890, 4365]. designers [2794]. Designing [3288, 3036, 769, 3231, 3337, 2732]. designs [4501]. Desktop [1522, 4578, 4323, 4531, 1338, 2552]. Desmond [4696]. destination [4609]. Detailed [30]. detecting [469]. Determining [211]. DEtool — typesetting [3449]. Deutsche [972, 1049, 1089, 1117, 1179]. Deutschsprachige [1119]. deux [1035]. dev [4505]. Devanāgarī [2609]. develop [3324]. developers [4607]. Developing [4609, 1383, 1614, 2166, 2453]. Development [2910, 4181, 4055, 969, 1933, 81, 3768, 2333, 3246, 3771, 3998, 3008, 3391, 2178]. developments [2898, 3329, 1939, 2861]. developments — autumn [3056]. developments — Spring [3023]. device [493, 534, 69, 2458, 1615, 108, 336]. device-independent [2458]. devices [68, 192, 193, 235, 263, 264, 299, 300, 334, 335, 372, 373, 410, 411, 671, 727, 772, 194, 579, 635, 32, 684, 3456, 860, 1156, 638, 687, 723, 774, 808, 3252, 168, 769, 1263, 581, 476, 960, 902, 2458]. devil [1909]. devnag [2609]. DG [865]. DHDURZ1.Bitnet [1268]. Diablarets [1392]. Diabolic [110]. diacritical [915]. Diag [1519]. Diagnosing [1112]. diagonal [383]. diagram [2992]. diagrams [3561, 4783, 3528, 4377, 1454, 1784, 1307, 1207, 4438, 2664, 1013, 2578]. Diagxy [2992]. Dialog [1380]. dialog.sty [1709]. dictation [4459]. dictator [3301]. Dictionary [605, 2836]. Did [2769]. Didactical [4606]. didn’t [3334, 3868]. Didot [1627]. Differences [1367]. different [3175, 4932, 3252]. difficult [847]. difficulties [2310, 1628]. difficulty [2757, 2812, 2877, 2915, 2951, 2974, 3014, 3077, 3117, 3154, 3189, 3221, 3317, 3352, 3395, 3437, 3503, 3542, 3572, 3614, 3650, 3688, 3752, 3786, 3821, 3861, 3898, 3937, 3967, 4005, 4077, 4112, 4157, 4196, 4226, 4231, 4270, 4317, 4353, 4389, 4419, 4451, 4486, 4523, 4633, 4555, 4590, 4668, 4708, 4745, 4773, 4800, 4842, 4884, 4923, 4971, 3016, 4040, 2976]. Digital [830, 2369, 4052, 2817, 2685, 4891, 1184, 208, 2785, 1627, 636, 4977, 3422, 2793, 4172, 771]. DigitalKhatt [4754]. digits [38]. Diglot [653]. dimensional [4376, 2066, 2577, 3713]. dimensions [3368, 4931, 4599]. dingbat [845, 897]. diplomatic [2019]. Direct [3326, 4727]. direction [3124]. directions [4867, 533, 4896, 515]. directions — [3875]. directly [4402, 4248]. Director [1438]. directories [412]. Directors [2646, 4082, 2566]. dirty [1041, 1196]. Disappearing [38]. disaster [1612]. discardable [2568]. Discovering [1480]. Discussion [13, 3727, 3030]. disk [2443, 816]. Disks [4782]. Display [153, 820, 4650]. displaylines [467]. disquettes [1035]. Dissemination [1536]. dissertation [3869]. Distinguishing [4652]. distribute [956]. Distributed [3586]. distribution [371, 409, 4852, 414, 587, 336, 2332, 3422, 2989, 4656, 4464, 1911]. distributions [3203, 1972, 681]. divergent [248, 247]. diversions [2784]. Diversity [2154]. DK [4089, 4761]. Do [3235, 3336, 3060, 4493, 3552, 1946, 4564, 62, 3868, 1941, 1710, 4910]. doc [919, 4655]. Docbook [2775]. DocCenter — TEXing [4061]. Docker [4533, 4945]. docks [714]. Docs [3672]. docstrip [2156, 4643]. doctoral [4826]. Document [4550, 4739, 4838, 4965, 879, 389, 611, 1611, 1339, 1329, 3559, 1202, 1176, 4333, 2505, 489, 1368, 4095, 3045, 4643, 3324, 1103, 1047, 2531, 2169, 4247, 293, 330, 1367, 2540, 3977, 1550, 3981, 4683, 1742, 3140, 717]. document-level [2540]. Documentation [433, 556, 312, 351, 4243, 4808, 4655, 4498, 451]. documented [1210, 697]. documenting [3951, 3271, 1200]. DocumentNewCommand [3555]. Documents [2104, 552, 2547, 1762, 3931, 4436, 3953, 4539, 2319, 2228, 2451, 3475, 1366, 2822, 503, 4950, 2510, 4611, 3985, 3560, 3261, 1845, 2071, 2817, 3471, 2074, 2463, 2378, 2777, 4828, 3873, 4297, 3558, 1908, 1969, 3422, 4061, 4718, 2339, 3925, 2766, 3455, 4503, 4216, 2737, 2198, 4125, 3532, 1536, 1490, 1474, 1648, 1480, 1216, 996, 2930, 2611, 2775, 1489, 2314, 1070, 1203, 471, 1550, 1483]. Documents — [215]. documents — part [1852]. DocVar [4333]. docx2tex [3331]. Does [3907, 3832, 660, 597, 658, 2459, 1449, 1524, 4014, 1840]. doesn’t [3869]. dog [4699]. doggerel [333]. doing [4238, 2316, 1556]. domain [582, 1210, 1330, 912, 1037]. domaine [1035]. Don [4561, 4408, 2397, 3510, 2503]. Donald [1391, 1494, 1507, 1899, 1990, 3975, 610, 2369, 487, 1576, 3716, 4028, 843, 1023, 2499, 2649, 448, 521, 569, 715, 759, 800, 1678, 4361, 3030, 2421, 2980]. Don’t [124, 3142]. Doob [3960]. dot [337, 2327]. Dotex — integrating [1477]…. [2281]. Dotted [1844]. double [2379, 2587, 3002, 3181]. doublepage [4756]. doublespacing [1411]. Douglas [4475]. Douglass [1905]. Doumont [3534, 3669]. d’ouvrages [2772]. Download [2360]. downloadable [194]. Drama [501, 2896]. dramatist [2896]. draw [1048]. drawdot [3923]. drawer [1327]. Drawing [3528, 3625, 2020, 3139, 1814, 3338, 1207, 2579, 2660, 3172, 3371, 1519, 784, 2328, 2507, 2513, 876]. Drawings [1441]. drawn [3772]. dream [3661]. Dreamboat [96]. Dreams [2310]. drive [4119]. driven [634, 4611]. Driver [1157, 1309, 582, 534, 861, 637, 109, 242, 1208, 863, 903, 688, 1735, 301, 1615]. drivers [336]. drop [3882]. drop-in [3882]. Drucbert [3892, 3908]. drug [1180]. Drümmer [4070]. DSSSL [1611]. DTD [2773]. DTD/schema [2773]. Dual [3198]. Dublin [390]. DuckBoat [4782, 4894]. DuckBoat — Beginner [4682, 4564]. DuckBoat — News [4430, 4326, 4367, 4458]. dummy [3453]. Dürer [1116]. Dutch [1293, 1199, 936, 923, 3040, 829]. DVDs [3447]. DVI [582, 1511, 8, 25, 64, 1550, 1387, 2506, 1735, 862, 3359, 4461, 167, 4338, 861, 1157, 2458, 2330, 266, 903, 688, 1309, 1944, 2170, 1960, 1408, 1359]. DVI-based [1944]. DVI-code [266]. Dvi QMS [301]. DVIasm [3171]. DVII [2506]. DVILASER [472]. DVILASER/PS [472]. DVIPDF [1961, 2330, 2069]. Dvips [4291, 1734, 1964, 4494]. dvisvgm [4338]. DVIto [470, 583]. Dwiggins [4410]. Dynamic [3953, 2823, 3931, 3524, 3515, 4176, 1580, 2324].

E-books [3626]. e-documents [2822]. e-foundry [4292, 4686]. E-interview [2500]. e-math.ams.com [1267]. e-print [1830]. e-reader [3456]. e-Readers [3658]. E-TEX [3803, 1101]. e-textbook [4181]. e.V [2291, 4739]. e.V. [1239, 4086, 1119]. E16 [3449]. each [4362]. Eaglefeather [2554]. EAN [1782]. Early [1995, 48]. Earthshaking [3509]. easier [1711, 3044, 3659]. easily [4336]. Easter [3373]. Eastern [1050]. Easy [2811, 2870, 1487, 1711, 1652, 4399, 1086, 3957, 4464, 2756, 1082, 3256]. easy-to-maintain [4464]. easyTEX [738]. eBook [3421, 3961]. Ebooks [3659, 4209]. Eck [50]. ecosystem [4828]. écriture [2769]. ecstasy [3525]. ed [1522, 952]. ed. [4258]. Edgar [4028]. EDIT [1475, 1937]. Edith [4931]. Editing [1962, 1648, 1475, 2820, 4247, 3986, 2017, 1113]. Edition [2844, 4963, 4696, 2711, 2866, 2945, 1996, 4255, 4990, 4964, 4788, 2734, 2613, 2671, 2741, 2742, 2782, 2772, 2125, 2253, 2019]. editions [2585, 2832, 3046, 4507, 3433, 1174]. Editor [490, 2124, 850, 1583, 4947, 1168, 3262, 2981, 2998, 2284, 680, 716, 148, 173, 205, 331, 294, 369, 86, 91, 126, 151, 177, 984, 408, 2, 54, 2043, 1747]. Editorial [483, 565, 621, 711, 757, 931, 983, 1019, 1061, 1099, 1139, 1298, 1396, 1434, 1469, 1503, 1545, 1572, 1641, 1798, 1864, 1894, 1955, 1984, 2010, 2035, 2114, 2141, 2181, 2218, 2246, 2358, 2409, 2416, 2491, 2526, 2882, 3576, 4081, 4233, 4319, 4355, 4421, 4453, 4525, 4635, 4557, 4670, 4747, 4775, 4844, 4886, 4973, 1726, 4010, 4236, 4322, 4358, 4424, 4456, 4528, 4638, 4560, 4673, 4750, 4778, 4847, 4889, 4975, 846, 892, 935, 1063, 1145, 1301, 1400, 1437, 1506, 1548, 1575, 1644, 1668, 1700, 1775, 1802, 1867, 1898, 1959, 1988, 2014, 2039]. Editorial [2088, 2120, 2147, 2222, 2250, 2278, 2362, 2395, 2420, 2497, 2567, 2647, 2722, 2884, 2979, 3081, 3121, 3159, 3283, 3357, 3467, 3547, 3578, 3618, 3693, 3758, 3793, 3828, 3904, 3973, 4012, 4083, 4117, 4203, 763, 2816, 1598, 3790, 3825, 3902, 3971, 4116, 4200, 139, 2681, 2187, 2479, 2631, 2920, 2598, 2446, 2040, 2225]. editors [474, 504, 4021, 1905]. EDMAC [1174]. ednotes — critical [2734]. EDT [1113]. education [2310]. educational [3560, 4057]. education — a [3445]. EduTEXTUG [3404]. Edward [4259]. effective [2328, 4326]. effectiveness [4512]. effects [1388, 720]. efficiency [1113]. efficient [1014, 2695]. effort [4936, 1199]. efforts [3403]. eight [1384]. eight-bit [1384]. Eighth [1790]. Eijkhout [1446, 1332]. Eisenstein [1995]. Eitan [2020, 3400]. either [1554]. Election [2186, 2520, 1918, 2807, 3074, 3351, 3571, 3785, 3858, 3963, 3999, 4073, 4192, 4223, 2912, 3390, 4264, 4415, 4447, 3611, 3149, 4482, 4627, 4663, 4836, 4878, 4701, 4918, 1957, 2264, 1886, 2012, 2435, 2754, 2674, 3541, 3819]. election — nomination [1887]. electrical [4725]. Electronic [1536, 2104, 611, 1907, 1701, 1339, 1746, 1197, 3586, 1364, 2190, 1944, 1991]. électronique [2776, 2772]. électroniques [2777]. Elefonts [4343]. Elementary [1412, 1452]. Elizabeth [1995]. elle [2769]. ELLHNIKA [1281]. ellipse [4871]. elsewhere [4184]. Emacs [1648, 1747, 2017, 1478, 4302, 4370, 4182, 3261, 4984]. Emacs-based [4182]. Emacs-type [1747]. embedded [4537]. Embedding [2993, 1812]. emergencystretch [4253]. emergent [4852]. emojis [4212]. empirical [577]. emplares} [2745, 3387, 3000]. Empowerment [4605]. enabled [3737]. enables [1761]. Enabling [2691]. encoded [1704]. encoding [2765, 1151, 1066, 3291, 3168, 4429, 3336, 1935]. encodings [2228, 1622, 3303]. encounter [848]. encountered [3703]. EncTEX [2225, 2790]. EncTEX — A [2226]. encyclopaedia [1763]. end [2727, 2534, 3204, 1478]. ended [834]. ending [2265]. endings [4332]. endnotes [1166, 4218]. ends [4463]. engine [2817]. Engineer [3288]. Engineering [4550, 4739, 4838, 4965, 4725, 2545]. engineers [1311]. engines [578, 3622, 4757, 4465, 4909]. engines — experiences [4541]. England [389, 2518]. English [1024, 2836, 4723, 4513]. enhance [2539]. enhanced [850, 2793]. enhancement [1368]. enhancements [4207, 1115, 122, 643]. Enhancing [2694, 4952, 2800]. Enjoying [3178]. enslaved [1001]. Entity [1452, 1202]. entomology [4676]. entries [4786]. Entry [3797, 3845, 3886, 3922, 1832, 3325]. Entry-level [3797, 3845, 3886, 3922]. Enumerations [3232]. envelope [4871]. envelopes [1324]. environment [762, 3702, 862, 295, 1619, 2829, 2779, 1614, 970, 1154, 1947, 661, 971, 876, 2583]. environment — [1064, 1025]. environment — multi [652]. environment — multi-column [652]. environments [3978, 827, 1087]. Envisioning [4811]. envy [869]. EP92 [1339]. Epen [3161]. epigram [2487]. epistulis [3692, 3757, 3792, 3827, 3903, 3972, 4011]. ePiX [2894]. Eplain [2999, 1159]. EPS [4338, 2194, 1964, 1908, 1969, 1359, 1874]. Epspdf [3256]. epstopdf [3513]. ePub [4022, 3660]. ePublication [3975]. eqnarray [3701]. equation [872, 822]. equations [2072, 2313]. equipment [68]. era [4977, 4594]. Erase [4910]. ereader [3627]. ergonomics [803]. Eric [1445, 1677]. Erik [4911, 1676]. Errata [51, 1191, 4271, 4390, 105, 1513, 1340]. Erratum [944, 950, 1026, 2015, 1414, 25, 788, 87, 2126, 1420]. error [1527, 469]. errors [1526, 1684, 993, 1112]. esami [3802]. especially [4132]. Essential [3893, 1584]. Esser [2705]. esthetics [2830, 3033]. estimation [3737]. ETAP [4944]. etc. — A [679]. ET — a [1583]. Ethiopic [948]. Éthique [2772]. ETX [3168]. Eukleides [2579, 2660]. Euler [3293, 2658]. Euler- [2658]. EURO [955, 1288]. EuroBachoTEX [3148, 4105, 3111, 3162, 3227]. EuroBachoTEX2013 [3857]. Europe [555, 1125, 1995, 1050]. European [1790, 2591, 2127, 1873, 1609, 2789]. EuroTEX [974, 3441, 3226, 3228, 1689, 2380, 2591, 2760, 3012, 2762, 3048, 1424, 1464, 2972, 3052, 1883, 3440]. EuroTEX94 [1777, 1800]. Eutypon} [3308, 3384, 3495, 3602, 3718, 3813, 3927, 4072, 4151, 4262, 4407, 4546]. Evaluation [817]. Even [2426, 1711, 2981, 2425]. event [4365]. Everett [2281]. Every [3981, 4928, 3197]. Everyone} [1445]. Everything [3104]. evidence [3665, 3729]. evidence-based [3665, 3729]. evolution [1963, 2988]. Evolutionary [4511]. exact [3489]. examinations [3802]. example [1764, 1749, 143, 3763, 301, 2933, 1521, 1551]. Examples [1042, 1079, 1167, 1679, 3067, 1042]. exams [4508]. Exception [2255, 1512, 2648, 262, 407, 530, 683, 804, 4359, 4409, 4677, 943, 1804, 2427, 2504, 2921, 3285, 3548, 3694, 3830, 4013]. exceptions [607]. Excerpts [113]. exchange [525, 2071]. execJS [2568]. Executing [4372, 3623]. execution [3061]. Executive [1438]. exercise [3064]. Exercises [1147, 2961, 1166, 1065]. exerquiz [2541]. Exeter [705]. exhibit [494]. Exhibition [3850]. existing [4097]. exotic [1515]. expandafter [1909, 739, 651]. Expanded [2999]. expanding [4136]. expansion [1322, 2537, 4141]. experience [2450, 3260, 3324, 3624, 378, 3925, 4530, 2731, 4811, 2798]. Experiences [3945, 3768, 134, 3041, 3461, 3636, 3328, 2706, 3412]. experiment [3954, 4467, 3259, 3665]. experimental [4171, 3254, 3478]. experimentation [4944]. Experiments [4443, 1198, 1177]. expert [3170, 2899]. expl3 [4375, 4646, 4298, 4101, 4897, 4397, 4863, 3486]. Exploiting [2424]. Exploration [4721, 3166]. explorations [4601]. Exploring [4141]. exported [3561]. Exporting [4022, 2198]. expressions [4539, 2890]. expressiveness [1731]. extend [1947]. extended [1579, 3909, 1371, 1151, 872, 1577]. Extending [3202, 4465, 2666, 2426, 1934]. extensible [4904, 4438]. extension [627, 2655, 971, 4298, 4101, 2226]. extensions [1648, 1067, 4136, 472, 1101, 4265, 3102, 3874, 2480, 3041, 2853, 500]. extensions — [3803]. external [4371, 4140, 2783, 4864, 3375]. exTEX [3028]. Extra [799, 4599, 799]. Extracting [4815]. Extravaganza [4941]. extremely [3552]. ey [2294]. Eye [4641].

F [4696]. Face [3984, 2500]. faces [4280, 3040, 771, 807]. Fácil [97]. facilities [2889]. facility [1383, 15, 634]. Facts [2002]. Faculty [4287]. Fahmy [3072]. failure [816]. Fairbairns [4779]. Falling [4028]. families [3837]. family [1907, 1842, 1046, 1088, 3234, 3871, 278, 4687, 582, 1473, 2657]. fancy [1842]. Fandol [3870, 3872]. FAO [3956]. FAQ [4055]. Farewell [4050]. Farmer [3960]. Farsi [3089]. FarsiTEX [2606]. Fascicle [4577]. Fast [1970, 1869, 1901, 695, 2457]. FasTEX [1869, 1901]. Father [2281, 2251]. fc [1578]. fc-fonts [1578]. fear [3231]. FEATPOST [2831]. feature [3424]. features [4694, 3343, 4244, 2986, 2793, 2043, 1113]. February [1182, 6, 1262, 1702, 593, 744, 4566, 4568]. Federico [4204]. feedback [4117, 65, 308]. feeding [3266]. Fehd [2937]. fence [1355]. fences [4339]. Fetamont [3909]. Fetching [1278]. Feuerstack [3106]. few [3194]. fewerfloatpages [4571]. FFI [4371]. Fickle [84]. fields [2611]. FIFO [1557]. Fifth [208]. Fifty [3962, 4992]. figure [3978]. figures [1554, 2690, 1216, 3338, 3952, 1969, 699, 2783, 3800, 2002, 1201, 1969, 2894]. File [4265, 276, 332, 465, 1806, 941, 39, 2506, 1533, 1877, 406, 529, 4502, 4429, 2620, 1741, 2540, 2568, 4787, 1511]. filename [3704]. Filenames [1152, 1407]. files [4186, 718, 862, 3171, 1278, 2340, 1962, 64, 106, 167, 4338, 1387, 455, 4815, 3345, 2330, 4458, 4140, 1936, 3297, 918, 682, 1876, 2227, 8, 25, 27, 1387, 2194, 164, 2170]. files — [4283]. fill — a [2151]. filled [3772]. filling [2151]. Final [2040, 627, 2159, 228]. finally [2397]. Financial [2519, 2911, 3006, 3150, 1343, 834, 4382, 4481, 4584, 4703, 4794, 4919, 3722, 3818, 3932, 4033, 4263, 4153, 3311, 3392, 3499, 3610, 156, 184, 222, 259, 313, 391, 2675, 2753]. Find [4911, 465, 847, 1555]. Fine [3346, 3462, 992]. finite [3907]. fire [714]. First [502, 2340, 293, 246, 4608, 2909, 936, 3516, 4612, 3416, 1873, 3726, 6, 925, 1370, 4530, 2731, 654, 1757, 330, 3729]. First-line [502]. first-time [3726]. Fischer [4990, 4964]. fitting [2781]. Five [1362, 3233]. fix [3923]. fixed [4296, 1508, 1340, 143]. Fixed-point [1340, 143]. Fixes [170]. flash [69]. flavors [4731]. fledgling [2928]. fleurons [3838]. Flexible [2817, 3667]. flight [2928]. float [3978]. Floating [1969, 3202, 1201, 1751, 2650, 699]. floating-point [3202]. floats [2463, 4756]. floats — the [4251]. Florida [196, 1725]. flow [3764]. Flowers [4659]. fly [4620]. fmtutil — past [4289]. FO [3250]. focus [4641]. foils [1486]. FoilTEX [1486]. Folio — a [4503]. Font [2554, 276, 609, 3365, 3677, 79, 27, 3101, 1703, 176, 3066, 3067, 2418, 4396, 3176, 3237, 491, 2798, 3287, 125, 857, 4754, 1151, 332, 4699, 194, 153, 1932, 4855, 4902, 1067, 1580, 2802, 2288, 3235, 2728, 1517, 3168, 3065, 3167, 3239, 4215, 1621, 1622, 1743, 2958, 3197, 947, 2075, 4292, 4686, 3233, 1812, 3768, 213, 88, 916, 1046, 1088, 3103, 3234, 3473, 3102, 3240, 1704, 2986, 406, 2793, 3837, 3944, 4174, 3246, 3871, 1203, 2537, 3761, 854, 1155, 4687, 2657, 1417, 4688, 3336, 3771, 3838, 2503, 1935, 773, 81]. Font [3242, 3169, 3104, 2795, 298, 686, 2796]. Font-specific [3237]. Fontanel [4681]. FontForge [3335, 2795]. Fonts [2450, 3197, 1027, 636, 1407, 3795, 4891, 277, 2823, 2656, 2230, 3292, 3196, 371, 409, 4690, 1152, 2571, 3697, 4173, 578, 2154, 2605, 770, 295, 1619, 1843, 4293, 4540, 4649, 4689, 4867, 3474, 3170, 1933, 3416, 856, 2993, 2193, 3236, 1842, 2047, 1620, 3710, 3043, 4294, 4434, 2800, 4599, 1873, 1929, 2364, 2423, 722, 3870, 3872, 1733, 1811, 608, 3082, 2153, 2796, 2824, 2612, 2964, 3241, 3335, 2360, 1404, 3166, 4494, 2835, 4209, 633, 2658, 805, 3337, 4940, 2086, 2696, 3248, 3364, 2424, 4291, 2799, 768, 3531, 3769, 4939, 2695, 4866, 3838, 991]. fonts [1903, 2550, 1931, 3677, 1578, 946, 532, 3795, 3303, 3838, 2650, 84, 992, 2096, 2127, 1902, 492, 1845, 1812, 2797, 2655, 1406, 2692, 1579, 278, 1549, 991, 2569, 2695, 3777]. fonts — a [2887]. fontspec [4396, 2986]. Fontzie [4766]. Fool [2148]. footnote [3974]. Footnotes [1178]. foreign [1806]. forever [3039]. forlorn [4439]. Form [4753, 1381, 591, 548, 594, 136, 3009, 3115, 3152, 3313, 159, 186, 4000, 221, 254, 286, 321, 322, 358, 359, 397, 615, 153, 592, 3232, 3032, 253, 285, 1887, 2080, 2968, 323, 360, 398, 440, 441, 561, 670, 753, 3817, 220, 510]. formal [2200]. format [1249, 332, 3700, 179, 485, 1012, 2450, 1623, 8, 25, 64, 106, 167, 3168, 1174, 164, 2612, 346, 529, 4502, 873, 1609, 4787, 276, 489, 406, 1960]. format — the [1945]. formats [4500, 4505, 4685, 3252, 4718, 4291, 2802]. formatted [2314, 1483]. Formatting [4854, 505, 4430, 3627, 546, 703, 605, 2692, 2586, 2802, 2632, 134, 2463]. forms [3209, 1382, 4932]. formulae [4436, 602, 4437, 3289, 3378]. formulaires [3032]. formulas [2893, 2572, 1910, 1814]. Fortran [174, 109, 242]. Forum [1703, 491, 3585]. Forward [1041, 4829]. forward — TUG [2851]. found [869, 4059]. foundry [4215, 4292, 4686]. Fourier [4692]. fpTEX [2332]. Fractal [987]. fractions [348]. Fraktur [1217]. Framed [244, 3342]. frames [3596]. framework [4504, 1609, 4503]. Framing [3596]. francais [4818]. France [879, 433, 552, 556, 609, 611, 1052, 553]. Francesco [2782]. Francophones [215]. Frank [2711, 2866, 2945, 1705, 4990, 4964, 2091, 4029, 4578, 4715]. Franklin [759]. Free [911, 2603, 868, 2746, 2796, 2360, 4334]. FreeType [4402, 4510, 4175]. FreeType MF Module [4402]. FreeType MF Module2 [4510]. Frege [4100]. French [295, 2923, 3452, 742, 2769, 2764, 2282]. French-speaking [3452]. frequently [3176]. friendly [3707, 5, 1937, 1942]. friends [3950, 3877, 3736, 3766, 4508, 1510, 3679]. Front [481, 755, 1231, 1348, 1352, 1431, 1821, 1861, 1891, 1924, 1952, 1981, 2007, 2032, 2060, 2083, 2115, 2142, 2183, 2215, 2355, 2636, 2641, 2716, 2758, 2813, 2878, 2916, 3155, 3190, 3222, 3279, 3318, 3353, 3396, 3438, 3504, 3543, 3573, 3615, 3651, 3689, 3753, 3787, 3822, 3862, 3899, 3938, 3968, 4006, 4078, 4113, 4158, 4197, 4227, 4229, 4268, 4315, 4351, 4387, 4417, 4449, 4484, 4521, 4553, 4588, 4631, 4666, 4706, 4743, 4771, 4798, 4840, 4882, 4921, 4969, 4038, 1232, 1349, 1353, 1432, 1467, 1501, 1543, 1570, 1595, 1639]. front [1665, 1822, 1862, 1892, 1925, 1953, 1982, 2008, 2033, 2061, 2116, 2143, 2184, 2243, 2273, 2390, 2412, 2442, 2477, 2494, 2562, 2628, 2637, 2643, 2718, 2759, 2814, 2879, 2917, 2952, 2975, 3015, 3078, 3118, 3156, 3191, 3223, 3280, 3319, 3354, 3397, 3439, 3505, 3544, 3574, 3616, 3652, 3690, 3754, 3788, 3823, 3863, 3900, 3939, 3969, 4007, 4079, 4114, 4159, 4198, 4228, 4230, 4269, 4316, 4352, 4388, 4418, 4450, 4485, 4522, 4554, 4589, 4632, 4667, 4707, 4744, 4772, 4799, 4841, 4883, 4922, 4970, 4039, 111]. front [4463, 3204, 1478]. Frutiger [4084]. FTL [516]. Fujitsu [730]. Full [1938, 4188, 3979, 3090]. full-width [3979]. fullpage [4756]. fully [4614]. Fun [1027, 3625, 3230, 2981, 3584]. function [1937]. functional [3762, 3451, 29]. Functions [4646, 2576]. Fund [2910, 3998, 3008, 3677, 3391, 4549]. fundamentals [4725]. funding [4851]. funds [4702]. Further [807]. Futura [4475]. Future [1884, 1915, 1974, 2000, 2024, 2051, 2107, 2132, 2173, 2234, 2262, 2298, 2382, 2403, 2433, 2672, 2751, 574, 2344, 1142, 1144, 2699, 1197, 4808, 2310, 1866, 3832, 2825, 1605, 1143, 4184, 1101, 3803, 4595, 4289, 2793, 2123, 1509, 1604, 1601, 4398, 2530, 3227, 2089]. futurelet [819]. fvextra [4180]. fx [1313].

G [4054, 1560, 4257, 3069]. Galińska [4275]. Gallen [3637]. galley [4134]. Galleys [1624]. game [965, 3994]. games [1009, 2466]. gap [2704]. Garcia [3145, 4204, 3742]. Garcia-De [4204]. garden [3946]. Garrigues’ [3373]. gate [1943]. Gaulle [3278]. Gdansk [1689, 1790]. GELLMU [3198, 2546]. gems [3876]. genealogy [2736]. General [1235, 1264, 1279, 4927, 909, 1072, 1128, 270, 414, 457, 587, 642, 692, 729, 778, 812, 4492, 4712, 2518, 4190, 4596, 4166, 4804, 4298, 4101, 4793, 3921, 1253]. generalization [651]. generalized [4691]. generalizing [2837]. Generate [3558, 3515, 455, 4725, 3403]. generated [2821]. Generating [2227, 4338, 3709, 741, 3456, 4065, 3095, 4251, 3707, 2936, 4953, 2193, 686, 3414, 2792, 1323, 2737]. generation [656, 175, 3063, 3359, 1731, 3699, 4658, 4509, 4512, 4335, 3135, 3738, 1439]. generator [3844, 4936, 1682, 1786, 1741, 2960]. Generic [1249, 276, 332, 2658, 4953]. gentle [3878]. Geogebra [4376]. Geometric [2784, 1454]. geometry [2579, 2660, 2575, 3172]. George [3992, 4255, 4117, 1673]. Georgian [2192]. German [1287, 666, 781, 1217, 743, 296]. Germany [303, 957, 694, 1302]. Gessler [4070]. get [338, 3907, 3512, 4458]. getoptk [3631]. gets [4122]. Getting [3407, 1320, 1104, 1214]. GF [4510]. ggplot2 [4901]. Ghostscript [4099]. Giambattista [4146]. Gianni [1675]. giddy [1485]. giddy-up [1485]. Gilardi [1675]. Ginger [1676]. give [1485]. given [4361]. gives [2330]. Giving [3550]. Glagolitic [1515]. GLAMP [3532]. glance [4514]. gleanings [3582]. Glisterings [4252, 4332, 2581, 2900, 2994, 3083, 3125, 3173, 3301, 3370, 3425, 3488, 3551, 3596, 3633, 3673, 3704, 3773, 3801, 3838, 3914, 3982, 4020, 4096, 4127, 4217, 2739, 4660]. global [2796]. globaldefs [4907]. Glossaries [4469, 547]. glossary [4104]. GloTEX — indexes [547]. Glue [4682, 1340, 3122, 143, 4023]. glue-penalty [4023]. glyphs [4689, 3914]. GmbH [4994]. gmcontinuo — some [3270]. gmdoc [3271]. gmL^{A}TEXplain — a [3265]. GMOA [4101]. GMS [4298]. gmverse [3270]. GNU [1648, 2017, 4871, 4870, 1478, 4656]. GNU/Linux [4656, 3838]. Gnuplot [3885]. go [801, 1485]. goals [4129, 3459]. Godine [4992, 4147, 4992]. gods [4963]. goes [1037, 4244, 3983, 4392]. Going [3200]. Good [3297, 4171, 1384, 4947, 2285, 2370, 4683]. goodbye [3457]. Google [4550, 4739, 3677, 3672]. Goossens [1705, 2091]. Gotisch [1217]. Gotoh [4288]. Gottlob [4100]. GPL [3870]. gradstudentresume [4095]. graduate [4095]. Graham [2400, 2280]. grammar [735, 3906, 129]. Grand [1027]. Granjon [4659]. Grapevine [736]. Graphic [3625, 4427, 1208]. graphical [2934, 4871, 1650]. GraphicP [2783]. Graphics [3218, 2368, 1965, 3138, 2091, 3776, 3264, 4991, 1109, 900, 4376, 1106, 1069, 3294, 3527, 4377, 859, 1471, 1480, 542, 4338, 3884, 3137, 3707, 2930, 1946, 2329, 3414, 2067, 2126, 1003, 1030, 1071, 1704, 1908, 1969, 3299, 858, 2066, 1070, 1205, 4901, 2465, 2336, 1483, 901, 3713, 4397, 2069, 374, 1972, 2337, 1908, 1002]. GraphicsMagick [4099]. graphicx [1908]. graphiques [881]. Grapholinguistics [4563]. graphs [2328, 2507, 4650]. Grathwohl [3534]. Grätzer [3992, 4255, 1673]. Gray [277]. great [3372, 4941]. Greece [2719]. Greek [1281, 2230, 1105, 949, 1514, 1757, 722, 1406, 4328, 2652, 2834, 2833, 4977, 3287, 4328, 2230]. Greek — An [1029]. Greek… [1305]. Greenwade [4117]. Greetings [86, 91]. Gregg [3341]. grid [2781, 3797]. Griffiths [4696]. Grimm [2336]. Groningen [1053, 3450]. Group [3888, 4517, 4765, 4960, 1886, 2012, 2435, 226, 2518, 925, 4594, 3404, 1702, 524, 1243, 1123, 829, 554, 2121, 4836, 4878]. group — Report [883]. Gudrun [3850]. GUI [4131]. Guide [490, 2253, 1559, 1707, 1362, 1611, 1108, 951, 3137, 1444, 2125, 4429, 4785, 2742, 962, 4788, 2046, 3641, 2844]. Guidelines [1101, 3803, 899, 1044]. guides [4929]. GuIT [2909, 3112]. gulf [175]. Gurari [2693, 2937, 2020, 3400]. GUST} [2904, 3109, 3427, 4105, 4215, 4292, 4686, 3242, 4120, 3857, 2818]. GUT [3778, 3997, 4780, 1035, 1282]. GUTenberg [1282, 1335, 1422, 1459, 1534, 1588, 1714, 1789, 1912, 1998, 2048, 2099, 2171, 2233, 2667, 2743, 2905, 3181, 3002, 4959, 881, 1240, 2296, 2379, 2515, 2587, 893, 956, 1121, 790, 4678, 3494, 3586, 4764, 4763, 4833, 4877, 4914, 2348]. GUTenberg’90 [1052]. GUTenberg’91 [1183, 1337]. GUTenberg’92 [1392]. Gyre [3101, 3416, 3248, 4434]. GyrePagella [3770].

H. [2815]. habits [4301]. hack [3961]. hack-to-learn [3961]. Hacker [1559]. hackers [2044, 3715]. Hacking [3171, 4442, 4302]. hacks [383, 180]. Hagen [2686, 2693, 3072, 3106, 3210, 3274, 3346, 3462, 3534, 3887, 4070]. Hahn [1445]. Hakuta [3887]. Halftone [685]. halftones [1481, 636, 1385]. halftoning [1220]. Halfway [3408]. Hamid [3210, 3430]. Hammersley [4716]. hand [208]. Handbook [3288]. handcraft [3230]. handles [4300]. Handling [3588, 654, 1368, 1810, 4821, 502]. Handmade [3230]. hands — the [4459]. handy [854]. hang [124]. Hanging [182, 214, 4252]. Hangul [4513]. hangup [150]. Hanoi [1418, 425]. happened [570]. Haralambous [3303]. hard [816]. hard-disk [816]. hardened [4497]. HarfBuzz [4467]. Hargreaves [1653]. harmful [4285]. harmonized [4942]. Harnessing [541]. Harris [266]. hashtable [211]. hax [535, 691, 831, 585]. headers [4217, 2156]. headings [3667]. headlines [4479]. headquarters [1093]. heads [3801]. HEARN.Bitnet [1271]. heart [1485]. Hebrew [1730, 2548, 2700, 2252, 46]. Heck [1522]. Hefferon [2693, 3210]. Heft [1561, 1715, 2049, 2130, 2258]. Hegelian [4403]. Heidelberg [1268]. held [1563, 1778]. Helmut [2844, 3358]. Help [606, 4376, 3058, 1383, 2070, 951, 1285]. helping [4362, 3300]. helps [3517]. Helvetica [1933]. Henderson [3742]. Hendrik [4659, 4674]. Henkel [3072, 3346]. Her [4544]. herbarium [4335]. Here [4020, 396]. Hermann [2502, 3747, 3850, 3165, 3293, 3030, 4051, 4052, 4049, 4050, 4543]. Herries [3426]. Herwijnen [1445, 1677]. Hey — it [2295, 2373, 2428, 2041, 2092, 2128, 2155, 2231]. HI [896]. HI-TEX [896]. Hierarchically [3671]. hierarchy [1669]. hieroglyphics [3767]. High [862, 2067, 2126, 1474, 3919, 1108, 2534, 1473, 3092, 3137, 4953, 421, 688, 2617, 2066]. high-level [1474, 2617]. high-performance [421]. high-quality [3092, 3137, 4953]. high-volume [688]. Higham [4696]. higher [2788, 2310]. higher-order [2788]. Highlighting [2583]. Highlights [2848, 2955]. highly [2936]. Hill [2808, 2880]. Hindi [2618, 3093]. HINT [4435, 4502, 4617, 4946]. ‘Hinting’ [2096]. hints [1166]. hip [1387]. histogram [876]. historic [3244]. Historical [4661, 429, 3798]. histories [4531]. History [4240, 4307, 4327, 4090, 4281, 2533, 4256, 4029, 4433, 4173, 1734, 3782, 4211, 4977, 3594, 4548, 3243, 4809, 4187, 3728, 4172, 1728, 4309, 4578]. history — a [4603]. hletter [3667]. Hoax [2148]. Hobby [3534]. Hoekwater [3106, 3274, 3346, 3462, 3534]. Hoenig [2152]. Hoerner [3217]. Hoerner&Sulger [3535]. Hogg [2940]. Høgholm [3210, 3346, 3430]. Holmes [4432]. holy [1730]. home [584, 111]. Hong [2966, 2959]. Hóng-Z [2959]. honor [1941]. hope [2224]. hopefully [627]. Höppner [3145]. horizons [2603]. Horizontal [1679]. horoscop [4178]. Horrors [2998]. host [1880, 1022]. Hour [4850]. hours [2982]. House [1759, 1779]. Houston [4345]. how-to [3843]. Howling [2963]. HP [281, 198]. HP-1000 [198]. HP1000 [588]. HP3000 [114, 145]. HP9000 [379]. HTML [4248, 2854, 4856, 4953, 2170, 2070]. HTML/CSS [4953]. HTML5 [3663, 3664]. HTTP [4607]. HUCC [417]. Hufflen [2688, 3072]. Hug [2208]. Hugill [4681]. Hugill-Fontanel [4681]. Humanist [1732]. Humanities [553, 3469, 522, 1994, 2042]. Hummel [3145]. hundred [1479]. Hungarian [2794, 3051, 2838]. Hungary [3012, 2604]. hvextern [4864]. hvfloat [4756]. hyperactivity [1198]. Hyperlatex [1831]. HyperTEX [1830]. Hypertext [3143, 1832, 1832]. hypertext — The [1197]. Hyphe [4412, 4583]. Hyphe-nation [4412, 4583]. Hyphenation [1512, 262, 407, 530, 683, 804, 1403, 4359, 4409, 4677, 943, 1804, 2427, 2504, 3285, 3548, 3694, 3830, 4013, 295, 2834, 2991, 4138, 1028, 4184, 4496, 4171, 3071, 232, 1304, 3059, 626, 296, 4724, 1849, 607, 2648, 2255, 2921, 261, 1514, 2735, 1848, 2326, 66]. hyphenations [719, 1331].

i-Installer [3203, 2988]. I/O [4787]. ialign [467]. IBM [354, 422, 474, 2071, 1165, 1523, 421, 498, 1030, 115, 863, 911, 1704, 2162, 1255, 1256, 1257, 2475, 2486, 1034, 644, 731, 813, 1073, 1164, 238, 2440]. Icelandic [2019]. Icons [1649]. idea [525, 897]. ideal [1359, 637]. ideas [1025]. ideas — lots [2981]. identification [2418]. identifiers [4056]. IDS [107]. IDS-640 [107]. IdxTEX [547]. Ierusalimschy [3210]. if [1313]. ifcat [1313]. II [390, 2206, 1079, 1717, 4344, 2318, 3275, 3553, 1071, 1852, 3243, 4331]. III [1167, 3597]. Illegitimi [4283]. illuminated [3069, 3363, 1310]. illumination [2685]. illumino [3559]. illustrated [3956]. Illustrator [2690]. image [4753, 2535, 4291, 4210]. Imagen [252, 70]. images [2571, 987, 4533, 3806]. ImageWriter [648]. imaging [1184, 2531]. imaging/digital [1184]. Imagining [830]. Imitation [3994]. Implementation [72, 2838, 1069, 2396, 417, 3036, 76, 379, 875, 827, 77, 1003, 1087, 3134, 7, 2367]. Implementations [477, 4472, 1210, 1165, 1523, 419, 243, 279, 309, 343, 382, 478]. implemented [2452, 3045, 4142]. Implementing [4288, 4401, 2981, 3376, 1579, 2324]. Implementor [59]. Implicit [2728, 2621]. important [162, 1902]. imposition [2542]. impractical [2289]. Impressions [3052, 4529, 4608, 2926]. improve [624]. improved [1155]. Improvements [1873, 4691]. Improving [4602, 3514, 4513, 4437, 1203, 2197]. inaugural [2122]. Inc. [4768, 4880, 4796, 4916, 2082, 2112, 2179, 2214, 2241, 2270, 2304, 2353, 2388, 2872, 2876, 2942, 2944, 2948, 250, 284, 317, 318, 395, 436]. Inc. — A [218]. Inc. — New [515]. Inc. — PCTEX [396, 362, 438]. Inc. — Professional [357]. Inc. — TEX [354, 512, 516]. incarnation [3629]. included [336, 2067, 2126]. including [2930, 1070]. inclusion [859]. Inconsolata [4015]. incorporation [1483]. Increased [1113]. incredible [3868]. incunabula [258]. indent [1603]. indentation [1171, 3298]. independent [590, 2458, 63, 901]. Index [617, 672, 1059, 888, 980, 1017, 1057, 1097, 1137, 1190, 1230, 1347, 1376, 1394, 1429, 1466, 1500, 1540, 1566, 1590, 1634, 1662, 1696, 1723, 1772, 1837, 1890, 1923, 1980, 2006, 2025, 2055, 2133, 193, 264, 300, 335, 373, 411, 480, 559, 618, 671, 673, 708, 754, 727, 772, 794, 838, 505, 4217, 595, 209, 656, 1010, 3327, 108, 127, 152, 178, 245, 1447]. indexes [3424, 1314, 4340, 546]. Indexing [4129, 2124, 1586, 1684, 966, 1328, 1420, 15]. India [3655, 2122]. India — The [2121]. Indian [2603, 2121]. Indic [1748, 1781, 2153, 2229, 2612, 4956, 2616]. Indica [1781]. indices [1992]. individual [1636, 4863]. industry [3676]. inexpensive [736]. infinite [3122]. influence [3978]. INFO [1283]. INFO-TeX [1283]. informal [4166, 164, 4172, 21]. Informatics [4287]. Information [483, 1298, 565, 621, 711, 757, 931, 983, 1019, 1061, 1099, 1139, 1396, 1434, 1469, 1503, 1545, 1572, 1641, 1798, 1864, 1894, 1955, 1984, 2010, 2035, 2114, 2141, 2181, 2218, 2246, 2358, 2409, 2416, 2491, 3193, 3576, 4081, 4233, 4319, 4355, 4421, 4453, 4525, 4488, 4635, 4557, 4592, 4670, 4747, 4710, 4775, 4844, 4802, 4886, 4973, 4925, 3790, 3825, 3902, 3971, 4010, 4116, 4200, 2632, 4815, 2506, 2074, 4952, 4272, 4391, 918, 690, 4060, 4259]. Infrastructure [3092, 3561]. ing [2293, 4370]. INITEX [233]. initials [1217]. Initiation [1316]. Inkster [4189]. Inn [4239, 4364, 4428, 4542, 4573, 4640, 4680, 4853, 4892, 4976, 2090, 2150, 2223, 2287, 2363, 2501, 2888, 2984, 3164, 3286, 3362, 3472, 3696, 3796, 4017, 4091, 4205]. Innovation [1727]. innovations [1759, 1779]. Input [93, 1764, 2327, 2227, 456, 2791, 4620, 3205, 3376]. Input-dependent [93]. Inquiry [429, 430]. INRIA [552]. INRSTEX [782, 499]. insert [3979]. Inserting [2783]. insertions [1167]. Inserts [1012]. Insights [1104, 1214, 4541]. inspired [4182]. install [3735, 2988]. installation [2554, 171, 1621, 3057, 3066, 3450, 3082]. installations [315, 1803, 4657, 4695, 65]. Installer [3203, 2988]. Installing [3590, 3170, 2553, 2552]. Installment [1075, 1161]. Instant [2574]. Institute [922, 4057]. Institutional [2841, 4234, 4314, 4320, 4356, 4414, 4422, 4454, 4519, 4526, 4558, 4628, 4636, 4671, 4742, 4748, 4776, 4835, 4845, 4887, 4967, 4974, 886, 976, 1055, 1094, 1135, 1186, 1228, 1344, 1373, 1427, 1498, 1539, 1567, 1593, 1632, 1661, 1693, 1721, 1769, 1794, 1818, 1835, 1859, 1888, 1920, 1949, 1977, 2003, 2029, 2057, 2079, 2109, 2137, 2176, 2211, 2238, 2267, 2300, 2349, 2385, 2405, 2437, 2471, 2483, 2557, 2592, 2623, 2634, 2669, 2714, 2748, 2809, 2868, 2913, 2949, 2969, 3011, 3075, 3113, 3151]. Institutional [3184, 3214, 4034, 508, 558, 613, 669, 706, 751, 792, 835, 1637, 3276, 3312, 3347, 3389, 3431, 3444, 3500, 3540, 3570, 3609, 3638, 3685, 3721, 3749, 3756, 3791, 3826, 3895, 3933, 3964, 4001, 4074, 4107, 4152, 4193, 4201, 58, 479]. Instructions [135, 158]. Instructors [320]. instrumental [1608]. integrated [2829, 2779, 1763]. Integrating [3418, 1002, 2201, 2531, 3299]. Integration [4510, 4719, 900, 1205, 2793]. intelligence [4981]. Interacting [2292]. interaction [4721, 1709]. Interactive [3527, 4857, 2820, 4944, 4614, 2453, 1747, 2654, 2339]. interactiveworkbook [2654]. interactivity [1750]. interchange [164, 63]. interdisciplinary [3880]. Interessenten [694]. Interesting [4826, 3300, 3232]. interface [1109, 2934, 850, 4251, 3090, 1359, 4655, 1650, 3948, 1046, 1088, 1208, 3025, 1014, 1937, 1756, 1874, 1942]. interfaces [2318, 3631, 3415]. interfaces — Practical [2852]. Intergraph [363, 364]. International [1339, 1247, 3442, 1219, 3233, 1204, 389, 894, 603, 1045]. international — A [801]. Internet [2348]. interplatform [1936]. interpolation [3847]. interpreter [1107, 2458, 3675]. Intertextualité [2776]. Interview [2421, 4597, 4598, 4715, 4716, 4714, 4807, 4806, 2980, 4324, 4237, 4432, 4681, 4169, 4204, 3839, 1507, 2500, 4752]. interviews [3411]. Introducing [3452, 4208, 3587, 3250, 2540, 2568, 487, 1808]. Introduction [1675, 1195, 1826, 3413, 3894, 1901, 2205, 1042, 3557, 3911, 4094, 247, 1100, 1300, 1354, 1399, 4900, 148, 173, 205, 1470, 331, 4870, 2632, 2288, 1827, 294, 3878, 3297, 4897, 4180, 126, 151, 177, 3402, 3373, 984, 4245, 408, 3376, 2663, 985, 1869, 165]. Introductory [4818]. introductory — photos [3020]. inventory [2765, 4978]. Inverted [747]. Invisibility [1404]. invitation [3048]. IoT [4812]. Iranian [2606]. Ireland [390, 413]. Irish [1218]. island [4693]. ISO [1611, 2021, 4299]. isolated [3988]. Israel [116]. Issue [1536, 2048, 2667, 2743, 2157, 2256, 2375, 2430, 4262, 4347, 4407, 4405, 4546, 4516, 4580, 4764, 4735, 4698, 4833, 4877, 4832, 4791, 4876, 4834, 4792, 4958, 4959, 4961, 4915, 4962, 885, 1187, 1342, 1426, 1462, 1538, 1565, 1591, 1660, 1692, 1719, 1817, 2171, 2233, 2379, 2587, 2905, 3000, 3002, 3108, 3181, 3306, 3309, 3382, 3384, 3385, 3387, 3428, 3429, 3495, 3496, 3497, 3498, 3491, 3563, 3564, 3565, 3567, 3602, 3603, 3604, 3606, 3645, 3682, 3683, 3718, 3744, 3745, 3778, 3779, 3780, 3813, 3855]. issue [3856, 3889, 3927, 3928, 3997, 4009, 4072, 4106, 4149, 4150, 4151, 450, 485, 4730, 4759, 4817, 4861, 4933, 3480, 3484, 3554, 3598, 3670, 3731, 3910, 4018, 4249, 4250, 4295, 4374, 4468, 4538, 4644, 4566, 4567, 4568, 4982, 3128, 4092, 4093, 4123, 4124]. Issues [2099, 2296, 4729, 3743, 4071, 4261, 4381, 4477, 4579, 4736, 4914, 4913, 1884, 1915, 1974, 2000, 2024, 2051, 2107, 2132, 2173, 2234, 2262, 2298, 2382, 2403, 2433, 2515, 2672, 2744, 2745, 2751, 2902, 2903, 2904, 2906, 3001, 3003, 3004, 3107, 3109, 3180, 3182, 3183, 3307, 3308, 3310, 3381, 3383, 3386, 3388, 3427, 3492, 3494, 3566, 3605, 3644, 3681, 3719, 3812, 3890, 3996, 4031, 4222, 1068, 2733, 3706, 3374, 3237, 3981, 689]. issues — Basic [3582]. Italian [1403, 1675, 261, 1444]. ITALIC [1241]. Italy [1273, 312, 351]. items [292]. iterations — second [3300]. iTEX [2779]. iTEX — Document [3627]. iTeXMac [2829, 3045]. IV [665, 1679, 3674]. IVRITEX [1163, 1284].

J [1905, 4696]. JabRef [4987]. Jackowski [3462, 4070]. jamais [4145]. Jan [3624]. Janc [2686, 2937]. Jane [1445]. Janie [4104]. January [954, 157, 552, 874, 1033, 19, 140, 61, 104, 1685, 4249, 4567, 4092, 4123]. Japan [1242, 848, 3873, 4426]. Japanese [628, 4723, 1306, 4652, 554, 3371, 3871]. Japanese-language [3871]. Java [2833, 2659]. Java-based [2833]. JavaScript [2201, 2540, 2568]. Javier [4597]. Jean [1759, 1779, 3892, 3908]. Jean-Pierre [3892, 3908]. Jerry [4543, 4256]. Jessica [4910]. jigsaw [3950]. Jill [698]. Jim [3994]. job [3868]. Joe [2940]. John [3962, 4854, 4716, 4806]. Joining [1387]. Joseph [4058, 4059]. Journal [3493, 1044, 3183, 3310, 3386, 3565, 3604, 3683, 3780, 3130, 4895, 3062, 3040, 4062, 3253, 3004, 1992, 4517, 4960, 4765]. journals [1941, 456]. Journées [215]. journey [3130, 4722, 3906, 4286, 4893]. joy [3741, 1960]. JSBox [3949]. JTUG — TUG [1242]. juggling [3774]. July [442, 1192, 1193, 1563, 1688, 2676, 2743, 3010, 3012, 185, 226, 443, 705, 163]. June [922, 1133, 2808, 2880, 433, 556, 2157, 2430, 786, 4563, 4730, 4817, 4933]. june.cs.washington.edu [1269]. Junicode [4855]. Jupyter [4363]. Just [1485, 1078, 1312, 1411, 1222, 3777, 124, 2535, 206, 3262, 2523, 2559, 2595, 2625, 3805, 4087, 4287]. Just-in-time [3805]. justification [4171, 4754, 3088]. justify [4285].

K-talk [817]. Kafkaian [3032]. Kaihsu [3379]. Kaneko [3887]. Kanji [3299]. Karlsruhe [882]. Kastrup [3274]. Kaveh [4191]. keep [4752, 3140]. Kees [4054]. Keith [4345]. Kelly [4256, 4543]. Kennett [4410]. Kerkis [2657]. kernel [4821]. kernel — experiments [4506]. kerning [2537]. Kew [2937, 3210, 3462, 3534]. Key [4821, 2898, 4251, 1105, 2311, 4066, 3376, 3486]. key/value [4251]. keyboard [1869, 1901]. keyfloat [4251]. keying [1869, 1901]. keynote [2851, 2852, 2310]. Keyword [4653, 650]. keywords [3512]. Khanh [2756, 2811, 2870]. Khmer [1617, 1757]. kicking [1900]. Kickstarter [3677]. Kidder [4188]. Kime [4598]. Kirschenbaum [4257]. Kissing [2923]. Kliever [4104]. knitr [4813, 3931]. know [4928, 3104]. knowledge [869, 2780]. known [170]. Knuth [843, 1023, 1391, 1494, 1507, 1768, 2649, 3975, 448, 521, 569, 610, 715, 750, 759, 800, 1678, 2397, 487, 1576, 4028, 1899, 1990, 2499, 791, 4561, 4408, 1440, 2369, 4361, 3030, 2421, 2980, 3510, 2503, 3716, 3893]. Koch [4085]. Kochanowski [3624]. koeieletters [4213]. KOMA [3759]. KOMA-Script [3759]. KöMaL [3061]. Komödie [4476, 1561, 2903, 3182, 3307, 3383, 3428, 3492, 3567, 3605, 3644, 3682, 3719, 3812, 3855, 3890, 3928, 3996, 4149, 4222, 4031, 4106, 4545, 4961, 1715, 2049, 2098, 2130, 2258, 4260, 3743, 4071, 4306, 4346, 4380, 4406, 4445, 4581, 4623, 4736, 4697, 4791, 4876, 4834, 4913]. Kopka [2844, 3358]. Korea [3163]. Kottwitz [4189, 4788, 4991]. kpfonts} [3549]. Krantz [1966]. Kris [4432]. Kurlansky [4309]. Kurzfassungen [972, 1049, 1089, 1117, 1179]. Küster [3346]. kZ [3138, 3414].

L [1995, 640, 1286, 1287]. L.U.G.s [1248]. l3build [4785]. l3build — A [3987]. l3expan [4298, 4101]. Laan [4054]. labeled [2690]. Labelled [1307]. labelling [2329, 1216]. labels [1381, 3267, 2067, 2126, 1533, 4335, 1483, 3914]. laboratoire [2772]. laboratory [4813]. labrea.stanford.edu [1270]. lament [4221]. Lamport [2500]. land [868, 4823]. Language [1611, 490, 2074, 1618, 3990, 4539, 1474, 1764, 2448, 1675, 1806, 2228, 3294, 2199, 2396, 3448, 3451, 3449, 3036, 3197, 4298, 4101, 3983, 1003, 2837, 3664, 2793, 1370, 217, 3871, 2619, 3460, 3663, 2579, 2660, 1115]. Language-dependent [1618]. Language — a [2101]. Languages [3332, 1284, 989, 1670, 2827, 4138, 3533, 3367, 1484, 1938, 2167, 2314, 1805, 599, 945, 1386, 2859, 2735, 1609, 2200]. Lapses [4679]. large [3560, 2935, 2322, 4645, 4125]. Larguier [2281]. Larson [4189]. laser [1071]. LaserJet [1030]. last [4780]. later [4298]. latest [338, 1939]. LaTeX [3961, 4145, 4208, 4238, 3907, 951, 1285, 4788, 4144]. L^{A}TEX-based [2654, 4828]. LaTeX-help [951, 1285]. L^{A}TEX-like [1486]. L^{A}TEX-on-HTTP [4607]. L^{A}TEX-paragraphs [1201]. L^{A}TEX-to-Braille [4727]. L^{A}TEX/TEX [1363]. L^{A}TEX2HTML [2691, 2167]. LaTeX2Nemeth [4874, 4727]. LaTeX2rtf [2178]. L^{A}TEX3 [2103, 2101, 3098, 3731, 3910, 3026, 2073, 4250, 4566, 4567, 1530, 1686, 3484, 3554, 3598, 3706, 3374, 3375, 1421, 1491]. LaTeXacroML [4400]. L^{A}TEX — a [4243]. L^{A}TEX — Part [4621]. L^{A}TEX — the [4248, 4507]. L^{A}TEX 2ε — an [4097]. L^{A}TEXing [3661]. latin [945, 1403, 1514, 3985, 3043, 2694, 4642, 3241, 3166, 3925]. Latin/Arabic [3985]. Laudatio [2502]. Launcher [4290]. Launching [4822]. Lausanne [1339, 830]. layering [3526]. layout [3881, 3096, 1752, 746, 788, 1178, 212, 385, 1455, 3875, 1579, 3488, 3369, 2197]. layouts [1761]. LBP [70]. LBP-10 [70]. le-tex [4994]. leader [4715]. leaders [3774]. Learn [3961, 3995, 3517]. learned [2706, 1729, 4608]. Learning [3849, 1628, 4400, 3468, 4696]. learnlatex.org [4614, 4615]. l’écriture [2803]. lecture [2777]. Lectures [356, 251, 399, 437]. Lee [1146]. Lees [4806]. Lees-Miller [4806]. left [576]. left-to-right [576]. legacy [2510]. legal [1606, 2071]. Lego [2992]. Lego-like [2992]. Lehmanns [2291]. Lehmke [2705, 2693]. Leland [3994]. lemma [1322]. Lemmes [3462]. Leslie [2500]. less [1527, 2982, 3043, 1761]. lessons [1728, 2706, 4729, 4608]. Lester [4146]. Let [3017, 3018, 3995, 124, 3961]. Letter [2037, 592, 4659, 3976, 591, 141, 2224, 46, 4056, 1783]. letter-cutter [4659]. letterforms [208]. Lettering [3851]. Letterpress [4016]. Letters [133, 12, 49, 95, 506, 1381, 206, 3667, 1873, 548, 594, 4088, 1011]. Letterspacing [1587]. Lettre [4763, 4833, 4877, 4914, 4959]. level [1474, 1810, 2617, 2066, 2540, 1408, 3797, 3845, 3886, 3922, 2609]. Lexicography [1005]. LH [2450]. Liaison [328]. liberal [3133]. liberate [3333]. libertine [4122]. LibertinusT1Math [4174]. libraries [4371]. library [4900, 3339, 2458, 189, 3365]. Libre [3677]. License [3594, 3197, 3242, 3242]. Licensing [4496]. Life [2163, 4410, 4146, 4639, 1829, 1729]. LifeLines [2736]. lifetime [3131]. LIFO [1557]. Lifting [3368]. ligatures [1618]. lightweight [4462, 4464]. like [2992, 2320, 4902, 1486, 4943, 990, 3675, 2459, 3869, 3621, 3978, 4014]. Lilypond [3328]. Lilypond-book [3328]. limits [3549]. l’impaziente [1653]. Lindau [1778]. Line [600, 2455, 502, 659, 654, 3058, 383, 918, 246, 901, 1169, 413]. linepenalty [4341]. lines [426, 1844, 4439, 1683, 3922, 4873, 4332]. linguaggio [1675]. lingual [3983, 3327]. linguistic [3848]. Linguistics [1807, 2857]. Lining [3982]. links [2195]. Linotron [69]. Linotype [4029, 2554]. Linux [4656, 3838]. Lion [2208, 1617, 2958, 4024]. LiPPGen [3844]. Lisp [4498]. List [1058, 929, 981, 1096, 795, 839, 43, 2078, 3463, 4094, 427, 1016, 1133, 1233, 1163, 1645, 1457]. listing [1680]. listings [3038]. lists [3906, 4724, 4786, 3083, 427, 1080]. Literate [1472, 1513, 1678, 2659, 2020, 1623, 1851, 3523, 3844, 1148, 4813, 1850, 2703]. literate-programming-based [3844]. literature [3586, 4987, 964, 3665]. Little [4682, 2226, 3297, 4786, 854]. Live [2671, 2971, 4586, 4790, 3519, 3590, 4905, 2863, 4290, 3948, 2613, 3876, 2125, 2253, 3708, 3296, 3406, 3591, 3258, 4463, 4714, 2990, 3255, 4133, 4464, 4948, 4533, 4945, 3450, 4726]. Living [4945]. livre [2776]. livres [2771]. llap [2584]. llmk — Light [4610]. LMTX [4466, 4761, 4829]. LN03 [633]. loading [4396]. Localisation [4216]. localization [2604]. locations [4647]. lockdown [4602]. Lockheed [308]. log [1512, 262, 407, 530, 683, 804, 4359, 4409, 4677, 943, 1804, 2427, 2504, 3285, 3548, 3694, 3830, 4013, 2648, 113, 2921]. Logarithmic [747]. Log — Update [2255]. logic [1728]. Logical [717, 1378]. logo [162, 3167, 3239, 492, 3265, 3909, 3479]. Lojban [2607]. lollipop [1485]. lone [2316]. Long [1166, 1040, 3589, 3914]. Long-winded [1166]. Longest [4096]. look [4679, 2699, 3416, 2459, 3869, 4616, 4896, 4014, 1853, 3364, 4172]. Looking [2851, 1170, 584]. looks [2899]. loop [651]. loops [3300, 3125]. Loopy.TeX [823]. loose [3305]. looseness [3305]. Lotz [3274]. Louis [1759, 1779]. Louis-Jean [1759, 1779]. Low [194, 769, 2609]. Low-cost [194]. Low-level [2609]. low-res [769]. lower [124]. Lowering [3325]. Lozano [4145]. LPEG [4500]. LPPL [3594]. lt3graph [3915]. Ltd [4739, 4838, 4965, 2815]. Ltd. [4550]. ltx2rtf [2198]. ltxsparklines [4286]. Lua [4372, 4369, 3522, 3987, 4784, 4872, 3330]. lua-widow-control [4784, 4872]. LuaCAS [4898]. Luafication [3295]. LuaJITTEX [3804]. LuaMetaTeX [4731, 4694]. luamodulartables [4865]. luaset [4865]. luatruthtable [4825]. Lucida [3629, 3697, 4433, 1728, 4089, 4173, 4868, 3629]. Ludwichowski [3106, 3462]. Luna — my [3593]. lunatic [3446]. L’utilisation [2782]. lwarp [4248]. lxfonts [3292]. LYX — An [2505].

M [780, 204, 1676, 2020, 730]. M-series [730]. Mac [2690, 1153, 2779, 2829, 3948, 912]. MacDraw [471]. Machine [901, 4721, 3032, 63, 728]. Machine-independent [901]. machines [67]. Macintosh [1258, 1517, 1070, 516, 580, 460, 779]. Macintosh — new [533]. MacKay [4048]. MacKichan [2872, 2948, 3218, 3433]. Macro [3631, 90, 3926, 659, 1528, 127, 152, 178, 209, 245, 2101, 40, 1760, 4654, 1104, 1214, 384, 741, 1447, 41, 703, 89, 916, 2511, 93, 697, 871, 1971, 2899]. Macros [85, 650, 1013, 346, 966, 740, 784, 747, 1475, 1947, 1319, 1416, 1415, 1762, 591, 1048, 3211, 3275, 45, 88, 212, 2466, 4906, 4142, 2891, 44, 43, 3371, 873, 4066, 1039, 4691, 2203, 698, 3808, 281, 1968, 1555, 1878, 783, 916, 782]. macros — A [1529]. MacTEX [2971, 4497, 4702, 3203, 3735, 3947, 2989]. MacTeX-2019 [4497]. made [4399, 3659]. madness [90]. Maeda [3887]. magazine [1487]. magic [4641, 728, 3601]. magnetic [135, 158]. magnification [83]. magnifications [770, 805]. mail [810]. mailbox [957]. Mailing [1163]. mainframes — MVS [1255]. mainframes — VM [1256]. mainframes — VM/CMS [1256]. mainstay [4714]. maintain [1629, 4464]. maintenance [4936, 22]. Make [4610, 1806, 1086, 2998, 3592, 4981, 3869, 3290, 824, 871, 3885]. make-up [824, 871]. Makeindex [1420, 1328]. makematch [3268]. makeover [2535]. maker [1204]. makes [1485, 2398]. MakeTeXPK [1803]. Making [4783, 2571, 4282, 1706, 499, 1828, 3906, 1803, 4648, 1556, 2696, 4286, 818, 1314, 3065, 2019, 782, 2931, 2321, 1992, 3956, 1085]. Makor [2700]. Malayalam [3085, 4599, 4956]. Manage [4333]. management [3206, 3915, 4987, 4400, 851, 1932]. manager [2924, 3948, 3876]. Managing [3249, 3367, 3560, 4431, 3057, 4439, 3366, 2819, 3064, 4125, 2551, 2322, 2333]. Mancilla [3346]. MANGLE [630]. Manipulation [879, 215, 389, 611, 4101, 1339]. manner [1851]. man’s [123]. Mansour [3742]. Manual [2523, 2559, 2595, 2625, 3678, 4177, 2580, 2093]. Manuale [3678, 4258]. manuals [1221]. manuscript [2019]. Manutius [3053]. Many [4280, 1526, 1938, 2086, 771]. map [4738, 4583]. Mapping [4863, 2775]. MAPS} [2744, 2906, 3107, 3306, 3385, 3714, 4755, 2819, 3003, 4150, 4405, 4516, 3496, 3564, 3603, 3745, 3889, 1607, 3778, 4347, 4735, 4832]. Marathi [2836]. March [1460, 1989, 1990, 2015, 925, 3161, 4124]. Marching [4096]. Margin [2537, 1554, 3514]. Mark [4309, 2074]. mark-up [2074]. Markdown [4462, 4297, 4648, 4731, 4781, 4862, 4899, 4819, 4983]. Marked [4728]. market [2535]. Marking [503, 2541]. marks [590, 244, 915]. markup [2396, 3533, 4462, 1754, 2820, 1367]. marques [2772]. Marrakesh [3013]. marriage [2607]. Marslide [2538]. Masaryk [4287]. Mass [665, 2535]. Master [1931, 1843, 2655]. Mastering [995]. matching [3268]. Matematyczne — an [4302]. material [1277]. Math [3217, 2547, 1996, 4255, 1622, 3530, 1673, 3628, 2961, 2449, 4173, 2154, 2342, 4570, 3040, 4867, 4829, 4937, 4908, 3588, 4207, 3198, 1620, 2311, 4294, 4460, 3418, 2655, 2611, 1969, 2658, 4940, 3064, 4939, 3336, 3363, 3771, 1039, 2704, 4868, 1379, 3770]. math-fields [2611]. math.ams.com [1267]. mathalfa [3628]. MathBook [3958]. Mathematica [2161, 3561, 3267]. Mathematical [371, 409, 3475, 1004, 1551, 1378, 4539, 2823, 4176, 2822, 3950, 3289, 4122, 3586, 2746, 2890, 4401, 4828, 4437, 3095, 3378, 2780, 3461, 2621, 311]. mathematically [2894]. mathematician [1654]. mathematicians [3847]. Mathematics [944, 1992, 2500, 3515, 2021, 4537, 3880, 2165, 608, 2167, 3343, 3529, 4056, 4403, 855, 2331]. mathematics — [3624]. MathKit} [2331]. MathML [2313, 2820, 2621, 2311, 2692, 2338]. MathPSfrag [3267]. Maths [2725]. mathscape — Combining [2161]. MathType [3219, 2313, 3205]. Matlab [2737]. matrices [1411]. matrix [820, 337, 4564]. matters [4210]. Matthew [4257, 4890]. Mauro [2782]. Maurolico [2782]. May [879, 881, 1052, 2296, 2515, 2587, 312, 351, 609, 908, 59, 2291, 56, 655, 4285, 4958, 4018]. Maya [3767]. mayan [2835]. MayaPS [3767]. McKenna [3960]. me [1485, 4963]. meal [4648]. Meandering [3551, 3195]. meanings [3517]. means [3043]. mechanisms [2898]. Medal [759]. media [3423, 3432]. mediation [4691]. Medical [3415, 3139]. medicine [3132]. Medieval [2019, 3360]. medley [392]. medstarbeamer [4126]. meet [2281, 1690, 1717]. Meeting [442, 709, 1192, 352, 368, 443, 663, 749, 1022, 2518, 4804, 694, 522, 4393, 3888, 4765, 4960, 922, 1016, 1133, 1182, 1372, 1393, 1496, 1635, 157, 185, 3442, 226, 227, 260, 666, 3671, 790, 4086, 679, 3161, 6, 925, 61, 104, 163, 244, 2122, 1702, 1778, 2325, 1193, 1493, 1766, 2528, 2601, 840, 3112, 4927, 4492, 4712, 4190, 4166, 4793, 1631, 2909, 314, 19, 56, 140, 883, 142]. meetingmins [3671]. Meetings [4596, 163]. meets [1989, 2015, 3420, 3068, 2958, 2703]. Melnichuk [1597]. Mem [3034]. member [2037, 2335]. members [2841, 4234, 4314, 4320, 4356, 4414, 4422, 4454, 4519, 4526, 4558, 4628, 4636, 4671, 4742, 4748, 4776, 4835, 4845, 4887, 4967, 4974, 886, 976, 1055, 1094, 1135, 1186, 1228, 1344, 1373, 1427, 1498, 1539, 1567, 1593, 1632, 1661, 1693, 1721, 1769, 1794, 1818, 1835, 1859, 1888, 1920, 1949, 1977, 2003, 2028, 2015, 2029, 2057, 2079, 2109, 2136, 2137, 2176, 2211, 2238, 2267, 2300, 2349, 2385, 2405, 2437, 2471, 2483, 2557, 2592, 2623, 2634, 2669, 2714, 2748, 2809, 2868, 2913, 2949, 2969, 3011, 3075]. members [3113, 3151, 3184, 3214, 3276, 3312, 3347, 3389, 3431, 3444, 3570, 3609, 3685, 3721, 3756, 3826, 3933, 3964, 4001, 4034, 4074, 4152, 4193, 4201, 508, 558, 613, 669, 706, 751, 792, 835, 3510, 2265, 479, 1989, 315, 571, 3500, 3540, 3638, 3749, 3791, 3895, 4107]. Membership [1058, 929, 981, 1096, 1233, 136, 159, 186, 221, 254, 286, 321, 358, 397, 795, 839, 2842, 887, 927, 978, 1015, 1056, 1095, 1136, 1189, 1229, 1346, 1374, 1428, 1465, 1541, 1568, 1636, 1637, 1663, 1694, 1724, 1771, 1795, 1819, 1921, 1978, 2004, 2030, 2058, 2080, 2110, 2138, 2212, 2239, 2268, 2302, 2350, 2406, 2438, 2473, 2484, 2521, 2593, 2677, 2867, 2968, 3009, 3115, 3152, 3313, 3817, 4000, 439, 509, 560, 614, 670, 707, 752, 793, 837, 761, 4119]. Memo [247]. memoir [4167, 3195, 3177]. memoriam [3695, 4047, 3892, 3829, 3908, 4360, 3278]. Memories [4054, 4323, 3926]. memory [166, 851, 477]. memos [1011]. menagerie [89]. menu [3263]. menus.sty [1709]. Merchandising — an [2135]. Merciadri [3634]. merged [645]. merger [1868]. merges [1532]. merging [3521]. mesostic [4854]. mess [4289]. Message [934, 520, 225, 257, 289, 327, 403, 568, 369, 1902, 2513, 4979, 3, 101]. Messing [4218]. Meta [2800, 4298, 3337, 1703, 4330, 3871, 494, 773, 298]. meta-design [4330]. Meta-designing [3337]. meta-family [3871]. Meta-Font [773, 298]. Meta- [494]. Meta-Scenarios [4298]. MetaContour [2858]. Metadata [4895, 4859]. MetaFog [1841]. Metafont [277, 276, 4402, 4956, 4938]. Metafont/MetaPost [4956]. Metafontbook [4691]. METAfonts [771]. Metafoundry [319]. Metafoundry — Introducing [392]. MetaFun [2425, 2426]. METAOBJ [2617]. METAPLOT [901, 2858]. METAPOST [2831, 4137, 4537, 3339, 3459, 3923, 4943, 2993, 3137, 3707, 2507, 3023, 2861, 3056, 3201, 3329, 3622, 3835, 3988, 3479, 1965, 2983, 4956, 3022, 2923, 3371, 4942, 3766, 3797, 3798, 3845, 3886, 3922, 3370, 3369]. MetaType [3167, 3239, 2799]. MetaType1 [3335]. Methodologies [1002]. methods [3666, 3917, 914]. Metric [27]. Mexico [572, 999]. [634, 577, 392, 2824, 3202, 579, 635, 684]. -like [990]. /MetaPost/PostScript [1965]. 84 [588]. MFCONFIG [4175]. MFLua [4279, 3630]. mfnfss [2022]. s [1405]. MFtool [634]. Michael [4751, 1552, 4324]. Michel [2091, 1705]. Michigan [117, 171]. Micro [219, 317, 2480, 4936, 4620, 3041, 2853]. micro-service [4620]. micro-templates [4936]. Micro-typographic [2480, 3041, 2853]. microcomputers [344]. microfiche [85]. Microtypography [3474]. microwave [4648]. Migrating [2818]. Migration [991]. Miklavec [3462]. MiKTEX [3708]. Milan [146]. millennium [2817]. Miller [4806]. million [4061]. mimeTex [2570]. Mimi [3579, 3580]. mind [1009]. mini [37]. miniature [3551]. MiniLaTeX [4618]. Minimal [3473, 4656]. minimum [681]. mining [3345]. Minion [3170]. Minnesota [9, 35]. Minoda [3887]. minority [2735]. Minutes [4190, 2982, 3671, 2518, 6, 1778]. mirror [2552, 3405]. Miscellaneous [2469, 375]. Misha [4256]. mishandle [3289]. Missing [3580]. missing… [1906]. misunderstood [3483, 3553, 3597]. misuse [2998]. Mittelbach [2711, 2866, 2945, 3346, 1705, 4990, 4964, 3534, 3669, 3742, 3887, 3960, 4189, 4715, 2091]. mix [3330, 3448]. mixed [3985, 1203]. mixed-font [1203]. Mixing [576, 2460, 3838]. mixture [3230]. MKII [3099]. MKIV [3099, 3200, 3557]. ml [3105, 3664, 3068]. MlBibTeX [2828, 3044, 4300, 2738, 3249, 3260, 3367]. Ml [2789]. MLTEX [956]. MMIX [3989]. MMTEX [4656]. mode [3876, 4984, 1081, 2017, 4719]. Mode def [806, 579, 635, 684]. model [3045, 2074, 3664, 1613, 1490, 2460]. Model-based [1490]. models [4246, 2314]. moderately [421]. moderately-priced [421]. Modern [1403, 1931, 1995, 2331, 2127, 2694, 1552, 4209, 532, 633, 991, 3919, 3206, 2834, 949, 1029, 3987, 4656, 1067, 1855, 4540, 2165, 2549, 3043, 1406, 3241, 3166]. ModernDvi [3919]. Modernizing [2549]. modes [2378]. modification [3265, 682]. Modifying [2050]. Modula [630]. Modula-2 [630]. Modularity [1846]. module [4402, 4175, 2838]. molecule [1997]. Money [4188]. Monk [2294]. Monk-ey [2294]. Monolithic [201]. monopoly [2833]. monotonic [2833]. Montréal [709, 749]. moon [3593, 2963]. Moore [2686, 3274, 3534, 3669, 3887, 3960]. morning [657]. Morocco [3013]. Morse [4367]. most [3176]. Motorola [1473]. mottos [545]. mouth [1080]. Move [1929, 3845, 2398]. moves [1657, 831]. Moving [1508, 2826, 2965]. Moye [3145]. MP2GL [3022]. MPLIB [4692, 3201, 3329]. MS [852, 737, 3205]. MS-DOS [852, 737]. MSDOS [1583]. TÊX [464, 2778]. much [4714]. Multi [179, 3662, 2228, 1451, 2935, 1614, 1178, 3983, 4056, 1937, 2319, 385, 3327, 1937]. multi-author [2935]. Multi-column [179, 1451, 1178, 385]. Multi-Edit [1937]. multi-function [1937]. multi-language [2228]. multi-letter [4056]. multi-lingual [3983, 3327]. Multi-target [3662]. Multi-use [2319]. multi-windowing [1614]. multibibliography [4064, 3883]. multicolumn [970]. Multidimensional [3344, 2701]. multienumerate [1847]. Multilingual [2958, 2076, 3034, 1329, 1025, 4722, 1150, 370, 627, 3100, 2793, 464, 2666]. multimedia [1550]. Multipage [3632]. Multiple [1931, 590, 1843, 244, 629, 1012, 718, 1319, 2551, 526, 1368, 1484, 3203, 4643, 2378, 2792, 3875, 3876, 1970, 3252, 468, 2655]. multiple-choice [1319]. multiple-column [1012]. multiple-option [1970]. Münster [666]. MuPDF [4535]. muse [3730]. music [3730, 4281, 2987, 1608]. musical [2726, 4211]. MusicTEX [1608]. musings [4593]. MusiXTEX [2086]. MVS [1126, 415, 538, 814]. My [4722, 3728, 3777, 2503, 4639, 2316, 1485, 1877, 4334].

N [2335, 1323]. nail [2320]. name [2285, 2370]. Names [3047, 1448, 3105]. namespace [2452]. Namespaces [3028]. Nancy [4544]. narrow [2224]. Nastaleeq [3247]. nation [4412, 4583]. national [969, 1199]. native [2199, 992]. natural [4539, 964, 4187]. Nature [4308]. Naugle — A [2279]. navigation [4721]. Navigator [2461]. near [868]. necessary [1602]. need [3142, 2535, 3235, 4564, 3336]. needed [2044, 2178]. needs [3482, 4457, 3739, 3851]. neglected [3445]. Nemeth [4460]. Neo [4564]. Neptune — a [4504]. nestable [1081]. nested [3476]. Netherlands [1883, 1053]. Network [1630, 3057, 3450]. Neurology [4238]. never [3530, 1879, 4475]. newbie [4928, 3328]. newbies [4132]. newcomers [4493]. newcommand [4684, 3485, 3555]. NewComputerModern [4687]. NewDocumentCommand [4684]. Newmath [3042]. News [197, 2157, 2256, 2375, 2430, 4946, 572, 147, 338, 4586, 4790, 3406, 3519, 1907, 69, 4730, 4759, 4817, 4861, 4933, 2733, 3480, 3484, 3554, 3598, 3670, 3706, 3731, 3910, 4018, 3026, 4249, 4250, 4295, 4374, 4468, 4538, 4644, 4566, 4567, 4568, 4982, 3128, 3374, 4092, 4093, 4123, 4124, 857, 4934, 2053, 2103, 111, 169, 236, 267, 305, 459, 1121, 849, 463, 517, 414, 587]. newsletter [905]. newspapers [3127]. next [885, 1187, 1342, 1426, 1462, 1538, 1565, 1591, 1660, 1692, 1719, 1817, 781, 4596, 4609, 3699, 4862, 1439, 3738]. NextStep [1708]. NextTEX [1708]. NFSS [1559, 3549]. NFSS2 [1584]. nica} [3108, 3180, 3309, 3381, 3563, 3606, 3681, 3744, 3779, 3856, 3491]. Nice [611, 1196]. nicest [2281]. Nicola [3930, 4144]. nightmares [4894]. NIH [75]. Nijmegen [1271]. Nimbus15 [4174]. Nina [4365]. nine [3385]. nineties [1008]. Ninja [4457]. niord.shsu.edu [1272]. nische [2903, 3182, 3307, 3383, 3428, 3492, 3567, 3605, 3644, 3682, 3719, 3812, 3855, 3890, 3928, 3996, 4031, 4106, 4222, 4476, 4581, 4961, 4149]. NJ [3010]. NL [1293]. nmbib [4064]. No [944, 950, 1912, 4894, 4459, 2727, 1026, 1038, 4140, 1222]. Noblest [4256]. noddy [1108]. node [3647]. nodes [4377]. NOFILL [87, 42]. Nominations [1224]. nomogram [3373]. nomography [3373]. Non [4642, 4283, 2452, 599, 1024, 960, 3059, 2899]. non-English [1024]. non-expert [2899]. Non-Latin [4642]. non-PostScript [960]. non-Roman [599]. non-standard [3059]. noname [1529]. nontrivial [3270]. Noob [4457]. Norbert [4714]. Nord [37]. Nordic [1123, 922, 1243]. normalized [3209]. ‘Normaltext’ [596]. Norman [2279, 1706]. North [2808, 2880, 112, 35, 71, 831]. Northeast [2160, 2159]. Norway [37]. Nos [1713, 1714, 1789, 1998]. nostarch [4935]. Notation [1452, 1378, 2396, 4211, 2464, 130, 964, 2960]. notational [3095]. Note [1598, 4147, 833, 787, 4235, 4202, 232, 4341, 162]. Notes [531, 2526, 1726, 884, 926, 975, 1054, 1091, 1188, 1341, 1425, 1461, 1537, 1564, 1592, 1659, 1691, 1718, 1793, 832, 4303, 4384, 4760, 4903, 4989, 4927, 3810, 3841, 4102, 4148, 1816, 1824, 1885, 1916, 1926, 1975, 2001, 2026, 2052, 2064, 2108, 2134, 2174, 2209, 2235, 2263, 2299, 2383, 2404, 4492, 4712, 1838, 1872, 3476, 2730, 4804, 4813, 3636, 4393, 716, 585, 925, 1849, 3134, 942, 2434, 2673, 2752, 691]. Notice [275, 2012, 2186, 2520]. notices — all [4304]. notices — TUG [4994]. Noticing [4603]. notification [4497]. Noto [4600]. Nottingham [389]. Nouveaux [2777]. Novaya [3984]. novel [1387]. November [1507, 2379, 2667, 3013, 4759, 4861, 4250, 4982]. novice [3925]. Novices [3930]. now — A [3195]. NP [4028]. NTG [1244, 1989, 2015, 1122]. NTS [1702, 1778, 1604, 1647, 2016]. Number [1460, 3907, 3712, 4217, 4471, 4472]. numbering [600, 872, 822]. numbers [2650, 3097, 280, 4142, 3179, 4252]. Numerations [4217]. Numerical [3917, 3622]. numérique — [2776]. numériques [2772]. numérisation [2772]. NYC [494].

O [4787]. Obituary — Yuri [1597]. Object [879, 1755, 1754]. object-oriented [1755, 1754]. objects [3533, 1443, 1518, 3022, 2617]. obscure [3238]. observation [1581]. Observations [248, 4594, 247, 1264, 577, 134]. obsolete [3608, 3531]. obtain [216]. Obyknovennaya [3984]. occasion [3975]. OCLC [278, 80]. OCR [4513]. Octave [3885]. Octavo [2651]. October [4262, 4407, 4546, 4579, 1184, 1338, 390, 665, 666, 830, 1778, 3491, 3563, 3602, 3606, 3718, 3779, 3813, 3927, 4072, 4151, 4538, 4644, 4093]. oddities [3272, 10]. odyssey [2529]. Off [4280, 2282, 1945, 1039, 3673]. off-off [1945]. offerings [4174]. office [998]. Officers [325, 366, 52, 99, 137, 160, 187, 223, 255, 287, 401]. Official [53, 100, 138, 161, 188, 224, 256, 288, 402, 326, 367]. Oh [3840, 3974, 3974]. Ohio [157, 104]. ol [1909]. old [4978, 965, 1217, 3873, 3626]. Olde [1006]. Olsák [2225]. Omega [2727, 1753, 2827, 2825, 2826, 3247, 3035, 2700, 2467]. Omega — Adapting [2609]. Omega — Why [1854]. On-demand [3532]. on-line [3058, 918]. One [1527, 2169, 4009, 3198, 3482, 1933]. One-document [2169]. Ongoing [3403]. Online [4021, 3879, 4593, 4752, 4474, 3583, 4613, 3560, 3656, 3290, 3532, 4104, 2541]. online — a [4655]. only [3945]. onto [730]. ooalign [4985]. Ööc [869]. Op [2726]. Open [3365, 3070, 2505, 4783, 4425, 2707, 4755, 4617, 1783, 3197, 141]. open-access [4425]. Open-belly [3070]. OpenBSD [3296]. OpenDX [2895]. OpenGL [3022]. Opening [1825, 1927, 3055, 2063, 1599, 1547, 1574, 1643, 1667, 1699, 1774, 1801, 562, 563, 619, 986]. OpenOffice.org [3059]. OpenType [2797, 2798, 3770, 4754, 3697, 4937, 4904, 3417, 4207, 4212, 2825, 3236, 3710, 3418, 3082, 4940, 4291, 3027, 4939, 3363, 3629, 3771, 4866]. Operating [229, 1708, 2189, 1440]. operational [4176, 1656, 1686]. operator [997]. opinions [3412]. OPmac [3808, 4140]. OPmac-bib [4140]. opportunities [2616]. OpTeX [4906, 4685]. OpTEX — A [4658]. optical [4176, 4513]. optimal [4253]. optimization [3526]. Optimizing [3455]. option [2230, 1329, 1176, 1047, 1443, 1518, 1704, 1970, 920, 919]. optional [545]. options [4131, 3977]. optotypes [4641]. Oral [1414, 1321]. Order [253, 285, 136, 159, 186, 220, 221, 254, 286, 321, 323, 358, 360, 397, 398, 440, 441, 510, 561, 615, 670, 753, 2788, 3249, 322, 359]. ordered [2757, 2812, 2877, 2915, 2951, 2974, 3014, 3077, 3117, 3154, 3189, 3221, 3317, 3352, 3395, 3437, 3503, 3542, 3572, 3614, 3650, 3688, 3752, 3786, 3821, 3861, 3898, 3937, 3967, 4005, 4077, 4112, 4157, 4196, 4226, 4231, 4270, 4317, 4353, 4389, 4419, 4451, 4486, 4523, 4633, 4555, 4590, 4668, 4708, 4745, 4773, 4800, 4842, 4884, 4923, 4971, 2976, 3016, 4040]. Ordinary [3984]. Oregon [77, 1372, 1393]. O’Reilly [3217, 3432]. ORG [4984, 4719]. ORG-mode [4984, 4719]. organization [4713]. organized [3671]. Oriental [3453, 988, 3124]. Oriented [879, 1755, 1754]. original [4787]. Origination [1536]. Ornaments [3633]. ornements [2772]. orphans [4784, 4872, 4439, 4440]. orthotypographiques [2804]. oser [4145]. Oslo [2649]. OSP [196]. OSP130 [195]. OSX [2829]. Other [1247, 10, 211, 4362, 4611, 1386, 1388, 1387, 4693, 4021, 3471, 2858, 4685, 4718, 1277]. otherwise [3552]. Ottaviani [3994]. Ottoman [1153]. Our [3851, 4613]. outgoing [403]. Outline [3102, 2571, 2096, 2796, 2964, 856]. outlines [2781]. outlining [740]. Output [50, 192, 235, 263, 299, 334, 372, 410, 1263, 574, 108, 168, 502, 68, 193, 196, 264, 300, 335, 373, 411, 493, 581, 671, 727, 772, 652, 179, 476, 579, 635, 862, 685, 726, 775, 2993, 3707, 684, 3835, 960, 860, 902, 1156, 638, 687, 723, 774, 808, 2458, 2791, 970, 109, 242, 1196, 4435, 107, 4932, 475, 3252, 3455, 3956, 4864, 32, 69, 2786, 863, 1042, 1079, 1167, 1679, 3659, 990, 648]. outputs [2792]. ouvrages [2772]. overlays [3912, 4019]. Overleaf [4550, 4739, 4838, 4965, 4947, 4948, 4716, 4806, 4608]. Overview [1174, 3244, 3916, 3520, 3634, 4097, 2153, 3481, 2732]. ovo [1963]. own [3140, 3482, 3324]. ownership [2386]. Oxford [2499, 800, 2518].

P [1653, 4028, 2370]. p. [1420]. p.190 [1524]. Package [2068, 2257, 2900, 528, 625, 1528, 580, 4288, 2992, 4176, 2652, 2124, 3034, 2199, 4820, 3883, 40, 2896, 1760, 2454, 4251, 4248, 2317, 3799, 2282, 2151, 3631, 3915, 3675, 1621, 3266, 3268, 3269, 41, 1847, 3735, 2076, 3802, 2924, 4441, 4440, 4571, 4758, 2538, 4298, 4101, 3763, 2573, 2532, 4142, 4874, 3632, 4394, 4396, 2986, 2835, 3628, 4825, 1871, 2232, 2987, 4507, 4064, 2200, 3595, 4717, 2230, 853, 2666, 2514, 2651, 2465]. Package/package [2900]. Packages [4471, 1742, 4243, 3762, 2831, 3884, 4612, 3512, 3915, 3634, 4395, 4097, 3297, 3426, 4865, 2732, 3481, 2924, 1908]. packages — the [916]. Packed [529, 406]. Packing [1405]. Pad [3736]. Page [3098, 3765, 4620, 1, 98, 324, 365, 400, 444, 482, 518, 564, 620, 674, 710, 756, 796, 841, 889, 930, 982, 1018, 1060, 1098, 1138, 1194, 1297, 1350, 1351, 1395, 1433, 1468, 1502, 1544, 1571, 1596, 1640, 1797, 1863, 1893, 1951, 1954, 1983, 2009, 2034, 2113, 2140, 2180, 2217, 2245, 2275, 2357, 2408, 2415, 2490, 2525, 2630, 2640, 2680, 2720, 2761, 2882, 2847, 2881, 2918, 2953, 2977, 3049, 3079, 3119, 3157, 3192, 3224, 3281, 3320, 3355, 3398, 3465, 3506, 3545, 3575, 3617, 3653, 3724, 3755, 3789, 3824, 3864]. page [3901, 3940, 3970, 4041, 4080, 4115, 4160, 4199, 4349, 4271, 4232, 4318, 4390, 4354, 4420, 4452, 4487, 4524, 4634, 4556, 4591, 4669, 4709, 4746, 4774, 4801, 4843, 4885, 4924, 4972, 4008, 252, 466, 747, 1110, 2455, 1761, 1752, 3667, 3270, 824, 470, 4067, 1204, 871, 746, 788, 2372]. pages [1765, 604, 2167, 4334, 3801, 702, 662, 2261]. Paging [4309]. painful [4826, 3552]. Pakin [4237]. Palatino [4187]. Palo [4491]. Pandoc [3916]. Pandora [986]. Panel [13, 3030, 2543, 2550, 2533, 3510]. Pankow [4258]. Papendal [1883]. Paper [2366, 50, 1671, 3945, 3230, 4431, 624, 1364, 3253, 3659, 4530, 3635, 2935, 1047, 4309]. Papers [1536, 611, 3716, 709, 1192, 1053, 1182, 1184, 1193, 1339, 1689, 1882, 312, 830, 840, 4576, 522, 2321, 973, 2062, 453, 2105, 2160, 1872, 1976]. paperTEX [3127]. paradigms [2610]. Paragraph [1036, 4134, 3522, 3526, 3592, 4439, 246, 4873, 4332]. Paragraphs [183, 3514, 1585, 4219, 3173, 3301, 818, 3173, 1201]. Parallel [4020]. parallels [3372]. Parameterized [3246, 3337]. parameters [650, 4654, 4378, 545]. Parametric [4294]. PARCAT — Applying [3676]. parfillskip [4443, 4873]. Paris [881, 1183, 1337]. Parkosz [4855]. parser [2452]. Parsers [3921, 4509]. parshape [4059]. Parsing [3918, 4500, 2457, 1412, 3704]. parskip [1172]. Part [1690, 1717, 1911, 1972, 1314, 2202, 2203, 2530, 2205, 2206, 1908, 1969, 3333, 1042, 1079, 1167, 1679, 2129, 2377, 2022, 4470, 1071, 4331, 586]. Parthasarathy [3995]. partial [1845, 1812]. Participant [3463]. Participants [1631, 1016, 1133, 3865, 3941, 4042, 4162, 1389, 1493, 1766, 2528, 2601, 227]. participating [4120]. participation [543]. Particle [1452]. particular [3173]. partitioned [1411]. Parts [2318, 787]. party [486]. Pascal [24, 7, 26, 1749, 3457, 29, 30, 42, 105, 26]. Pascal-coded [105]. PassiveTEX [2615, 2456]. Passport [4161]. Past [2825, 4398, 2699, 1022, 4808, 1866, 4810, 4184]. pasting [896]. PATCH [718, 2317]. Patching [4868]. patent [3764]. Path [2366, 4361, 3988]. paths [1388, 4942, 3772, 3227]. patients [3954]. Patrick [2844, 4168]. patrimony [2764]. pattern [1806, 4753, 4980, 3268, 4512, 1204]. patterns [2834, 2991, 1514, 4496, 2735, 4724, 3798, 1028, 2196]. Paul [1333]. Pavneet [4169]. PC [1257, 354, 422, 735, 421, 1030, 1071, 1014, 512, 1035, 1153, 474, 649]. PC-compatible [1030]. PCs [1165, 1523, 1383, 911]. PC TEX [345]. PDF [4990, 4334, 4436, 4009, 4951, 3359, 4728, 3456, 2821, 3256, 2654, 2324, 4401, 4952, 2538, 2611, 3403, 2573, 4065, 4494, 4506, 2542, 2339, 2540, 2568, 3455, 3513, 4860, 2201]. PDF-based [2542]. PDF-only [3945]. PDF/A [4728, 4065]. PDF/A-compliant [4065]. PDF/X [4065]. PDF/X- [4065]. pdf2htmlEX [3879]. PDF — a [4645]. pdfL^{A}TEX [3062, 3263, 3143, 3370]. pdfL^{A}TEX-based [3062]. pdfs [4565, 4620, 4065]. pdfscript [3558]. pdfTEX [2292, 2093, 2853, 2663, 2541, 2542, 2537, 2543, 3257, 3237, 3403, 3364, 3377]. pdfTEX-generated [2821]. pdfTEX — pdfx.sty [4065]. pdftricks [3513]. Pearls [2995, 3272]. Pearson [4550, 4769]. pedigree [3139]. pedigrees [3415]. peek [2089]. PEN [1452]. penalty [4023]. Penrose [2464]. people [2281, 1479, 3411]. perfection [3027]. performance [1473, 421, 4472]. Perl [2465, 2787, 2292, 3328, 2891]. Perl/XML [2787]. PerlTEX [3208, 2891]. permanent [3272]. permission [2425]. permuted [1010]. Persian [1153, 2618, 4511]. persistence [3976]. personal [3130, 1646, 4722, 4602, 540, 2462, 3925, 4603, 4139, 2957, 2876, 2942, 362, 396, 438, 512]. personalized [4725]. perspective [3419, 2603, 2618, 2536, 2957]. perspectives [1878, 3375]. Peter [2937, 3145, 3346, 4660, 2946, 3742, 4117, 4118, 3426]. Petersburg [1725]. Petr [2225]. PGF [3138, 3713]. PGF/TikZ [3713]. PGFPlots [4725, 3846, 3478]. PhD [3930]. philanthropiques [2804]. Philip [4598]. Philological [2889]. philology [2652, 1710, 524]. philosophy [3947]. Phonetic [603, 1219, 1930]. photographs [2785]. photos [3228, 4044, 4164, 4490, 2469, 3941, 4042, 4162, 2956]. phototypesetter [265]. PHP [2691]. phrases [3038]. Physical [2532]. physics [3461]. Pi [4534]. pic} [1003, 2067, 2126, 2066]. picking [4375, 4470]. pict2e [3097]. picture [1003, 876, 2583, 3702, 1947]. Pictures [3020, 206, 1657, 1787, 2376, 1738, 471]. Pierre [4048, 3892, 3908]. pinyin [1549]. Pisa [147]. Piscataway [3010]. Píska [3462]. pitch [3848]. pitfalls [3733]. Pitman [3454]. Pittsburgh [900]. PK [4510, 682, 406]. pkgloader [3915]. placement [3489, 3270]. Placing [3979]. PLAIN [651, 1850, 3417, 872, 4985, 2999, 3474, 3808, 822, 4873, 4787, 4986, 4100, 423, 467, 4658, 1078, 1312, 1411]. Plane [4721]. plans [2729]. plasTEX [3407]. platform [2826, 4464]. please [806]. pleasing [3847]. pleasure [3231]. plot [2576, 2576]. Plots [3885, 3267, 783, 2577, 4650]. Plotting [3478]. plug [4175]. plug-in [4175]. plugin [2544]. plurality [1935]. plus [1708]. PNG [3835]. poems [4854]. Poetica [1757, 1842]. Poetry [2248, 2585]. point [2251, 2650, 3202, 1304, 1508, 1340, 143]. points [2252]. poisons [4786]. Poland [1689, 2145]. Polices [2803]. policy [139, 1720]. Polish [1623, 2190, 4304]. Polished [2708]. polyphonic [1608]. polytonic [2833]. Ponchin [2874, 2947, 3432, 3536, 3219]. Pond [4682, 4782, 4894, 4564]. Poor [123]. pop [1383]. pop-up [1383]. Popineau [2705]. popular [176]. port [852, 737, 2465]. portability [1936]. Portable [1106, 859, 899, 2936, 1471]. portal [3794, 3983]. Portfolio [3639, 4282]. Porting [3296, 200, 536, 766]. Portland [1372, 1393, 3942]. ports [730]. Portuguese [4537, 626]. Position [2266, 3978, 22, 3377]. positioning [2665]. possibilities [4097]. possible [4561, 4950]. possibly [4136, 384]. Post [1211, 505, 3954]. Post-L^{A}TEX [505]. post-stroke [3954]. Postcard [4016]. Postnet [1516]. posts [4430]. PostScript [639, 2195, 2460, 3776, 354, 3239, 3256, 1206]. pour [1035, 2777, 2782]. power [1322]. powerdot — making [2997]. powerful [3915]. PowerShell [3831]. pp [2015, 1671, 1901, 2126, 1518]. pp. [1513, 1414, 1686]. PPCHTEX [1997, 1910]. PracTEX [3004, 3183, 3310, 3386, 3565, 3604, 3683, 3780]. PracTEX’05 [2926]. Practical [3992, 2846, 2850, 2289, 2886, 1445, 1677, 3253, 4575, 1827, 3843, 4349, 2670, 2808, 2880, 3010, 2848, 1220]. practice [1110, 3041, 2377, 2853, 1712, 1674]. Practitioner [1513, 1472]. pragmatic [1585]. Prague [1424, 1464, 1990]. pratique [1707]. pre [1734, 3782]. pre-TEX [3782]. precision [4142]. Prefixation [4368]. Prefixing [3801]. Prehistory [4891]. Preining [3742, 4189, 4714]. Preliminary [1914, 2145, 260, 1688, 2056, 883, 1204, 4171, 61]. preparation [1368, 1103, 966, 2378]. prepare [39]. prepared [1296]. Preparing [4062, 3732, 3843, 4752, 1002, 1481]. prepress [2892]. preprocessing [1749]. preprocessor [1781, 2322, 4438, 1519, 1112]. preprocessor — Sinhalese [1748]. prerelease [647]. prerex [3986]. present [2825, 4398]. presentation [3844, 3198]. Presentations [3853, 2997, 2538, 3911, 3134, 2539]. Presenting [4613, 4259, 2167]. presents [4812]. preservation [3586, 3422]. Preserving [4908]. President [1022, 4807, 4357, 4423, 4637, 4672, 4749, 2266, 934, 1021, 1100, 1141, 1300, 1354, 1399, 2721, 2883, 2919, 2978, 3080, 3120, 3158, 3282, 3356, 3466, 3546, 3577, 3619, 4191, 844, 891, 933, 404, 449, 520, 568, 623, 676, 713, 760, 798, 1470, 4121, 1839, 1897, 1927, 1958, 2013, 2038, 4235, 4202, 2063, 2087, 2119, 2146, 2221, 2277, 2361, 2394, 2419, 2496, 225, 257, 289, 327, 403, 2249, 4321, 4455, 4527, 4559, 4777, 4846, 4888]. press [4511, 2301, 2352, 3426]. Pretprin — a [2200]. pretty [1749, 3921, 2200]. pretty-printing [1749, 3921, 2200]. prettyprinting [1610, 1651]. prevdepth [4954, 4023]. Preventing [4757]. Preview [776, 2574, 533]. previewer [724, 3525, 725, 1747, 1149, 470]. previewers [475]. Previews [4856]. Prez [1398]. priced [421]. Prime [1259, 271, 910, 1129, 304, 339, 589, 3673, 495]. Primer [1966]. primes [541]. primitive [4465]. primitives [4953, 4569]. Primo [4949]. Principiae [3537]. principles [405, 293, 330, 3134]. Print [2321, 4978, 1830, 3290, 2169, 2961]. printed [3560]. Printer [4016, 252]. printers [1071, 1177, 337, 1030]. Printing [702, 2785, 1995, 604, 662, 1759, 964, 1738, 950, 4186, 578, 862, 1749, 1779, 4511, 1533, 688, 898, 3921, 3728, 2200, 1845, 915]. Prize [4408, 4365]. Pro [3170]. problem [425, 1442, 2929, 648, 1935, 598, 132, 155]. problem — solution [657]. Problems [1109, 1870, 131, 4826, 682, 2691, 782, 92, 132, 155, 1811, 244]. procedure [592]. Procedures [3240, 4793, 29, 30, 229]. Proceedings [3888, 556, 4030, 4105, 3017, 3018, 3253, 1790, 3857]. Proceedings — Apologia [1776]. Proceedings — Paperless [2380]. process [3254, 1360, 1369]. Processing [389, 4458, 941, 787, 634, 997, 3487, 4511, 1930, 1370, 1412, 2898, 595]. Processor [454, 422, 1473, 3035, 1377, 3699, 718, 1315, 2505, 631]. processors [3734]. Producing [4248, 4932, 3288, 1319, 3667, 918, 1360]. product [2535, 4611, 3869, 5]. Production [2873, 2941, 3219, 884, 926, 975, 1054, 1091, 1188, 1341, 1425, 1461, 1537, 1564, 1592, 1659, 1691, 1718, 1793, 832, 4303, 4384, 4760, 4903, 4989, 3810, 3841, 4102, 4148, 1816, 1824, 1885, 1916, 1926, 1975, 2001, 2026, 2052, 2064, 2108, 2134, 2174, 2209, 2235, 2263, 2299, 2383, 2404, 1838, 717, 1598, 2843, 4266, 4313, 4348, 4383, 4413, 4446, 4480, 4518, 4551, 4664, 4585, 4630, 4704, 4740, 4767, 4795, 4839, 4879, 4993, 4917, 4966, 1234, 1430, 1542, 1569, 1594, 1664, 1695, 1722, 1726, 1770, 1796, 1820, 1836, 1860, 1889]. production [1922, 1950, 1979, 2005, 2031, 2059, 2081, 2111, 2139, 2177, 2213, 2240, 2269, 2303, 2351, 2387, 2407, 2439, 2474, 2485, 2522, 2558, 2594, 2624, 2635, 2678, 2715, 2755, 2810, 2869, 2914, 2950, 2970, 3007, 3076, 3116, 3153, 3188, 3220, 3277, 3316, 3350, 3394, 3436, 3464, 3502, 3539, 3569, 3607, 3646, 3684, 3720, 3748, 3784, 3816, 3854, 3896, 3934, 4002, 4035, 4075, 4109, 4154, 4194, 4224, 1064, 762, 4852, 3062, 3559, 1619, 1760, 4929, 1358, 1739, 293, 330, 2434, 2673, 2752, 1872, 2798]. production [1108, 998, 2854]. production/commercial [762]. production — lessons [1729]. Prof [1899, 1990, 2336]. Professor [2502, 4238]. Profile [3400, 315]. profiler [1440]. profiles [3533, 3959]. profiling [4816]. profit [3584]. Program [1600, 1727, 610, 3322, 3226, 3866, 4043, 4163, 4273, 4489, 4710, 4802, 4925, 1688, 2683, 2763, 2849, 2925, 2954, 3086, 3129, 3225, 3321, 3399, 3507, 3654, 3725, 3865, 3941, 4042, 4162, 226, 941, 1932, 1810, 4140, 42, 4864, 4787, 2158, 2309, 2447, 2527, 2600, 750, 1961, 2065, 2145, 349, 2199, 87, 1065, 1147, 2056]. program — August [260]. programme [2772]. programmer [3993, 4807]. programmers [3847]. Programming [2101, 1513, 4243, 4733, 3208, 4577, 3642, 3486, 2995, 1755, 1623, 1472, 3523, 3844, 1148, 4694, 3451, 180, 3987, 2703, 1754, 2659, 1370, 217, 3981, 3375, 2800, 2020, 2101, 297, 1678]. programs [1762, 2780, 3989, 1942, 974, 873, 28]. progress [3037, 4860, 1758, 1753, 3771, 117]. Project [2053, 2103, 3242, 3594, 4501, 3737, 3411, 2707, 3459, 3329, 4851, 4645, 4860, 4946, 4025, 2468, 2799, 4331, 3477, 1627, 236, 267, 305, 459, 3101, 3586, 3870, 3872, 2073, 4617, 3510, 4811, 2862, 2778, 2508, 849, 1530, 1753, 4715]. project — An [2796]. Project — From [3196]. project — Towards [2765]. projective [3122]. Projects [291, 4215, 4431, 3197, 2322, 4292, 4686, 3233, 4608, 2333]. prompted [489]. ProofCheck [3410, 3739]. proofing [4504]. Proofmode [83]. proofreader [386]. Proofreaders [3586]. proofs [3136, 3410, 277]. proper [4298]. Proposal [2206, 2100, 63, 58]. Proposed [3875, 2101, 1875, 681, 434]. protection [531]. proTEXt [2971, 390, 665, 3031]. proto [1032, 1075, 1161]. proto-TUG [1032, 1075, 1161]. Prototyping [3022]. Providing [4533]. PS [1579, 472]. ps4pdf [2573]. pseudo [1580]. pseudo-dynamic [1580]. PSfrag [2044]. psnfss [2022]. PSPRINT [809, 583]. pst [3513, 2151, 2664, 3513, 2576, 2577]. pst-3dplot [2577]. pst-circ [2664]. pst-fill — a [2151]. pst-pdf [3513]. pst-plot [2576]. pst2pdf [3513]. PSTricks [2151, 3139, 4331, 3477, 1740, 3338, 4181, 3776, 4376]. PT [3587]. PTEX [65, 4652, 3874]. PUB [3782]. PubliC [851, 1210, 3242, 3594, 1035, 1037, 2612, 940, 582, 1330, 912]. Public-domain [1210]. publication [4752, 39, 1362, 696, 1581, 488, 1364, 3253, 1944, 1991]. Publications [290, 1092, 2045, 1488, 1295, 1296, 322, 359]. publish [1361, 3018]. Published [2523, 2559, 2595, 2625, 3778]. Publisher [1628, 1359, 1361, 4301, 3868, 4147, 4611, 2319]. Publishing [1338, 1536, 611, 1522, 1701, 1339, 1746, 762, 1357, 4852, 4895, 1197, 3880, 4473, 2530, 3662, 283, 678, 3469, 2166, 4401, 3343, 2323, 4819, 573, 2169, 3252, 4949, 1991, 4323, 4531, 3412, 3728, 3879, 2510, 4, 2343, 1213, 4578]. publishing — 1990 [1111]. Pubs [1291]. Punctuation [4479, 182, 214, 4284]. Pune [2746]. Punk [773]. purpose [1764]. Purvis [3994]. Pushing [4829]. Putting [3340]. puzzle [4344, 2220, 1524]. puzzles [4722, 3950, 2466]. puzzling [1971]. Pvt [4550, 4739, 4838, 4965]. PXL [106]. pyramidal [747]. Python [4814, 4499]. PythonTEX [3878, 4180].

Q&A [1411, 1078, 1312, 4983]. Qin [2960]. qstest [3269]. quality [3919, 862, 1108, 3092, 3137, 283, 2067, 2126, 4953, 2339, 219, 317]. quantities [211]. Quarter [610]. quarterly [799]. que [4145]. Quelques [2770]. Question [2499, 3561, 1132]. questionnaire [1064, 473]. Questions [11, 1899, 1990, 3975, 2649, 4326, 4617, 3176]. Quick [3288, 3852, 1196]. QuickDraw [1206]. quickly [3452]. Quickref [4498]. Quintus [4168]. Quirk [2693]. quiz [4366]. quizzes [2541]. Quo [4616, 2508]. Qurānic [3526].

R [4992, 3931, 4147, 3433, 4901, 3524]. R.I.P. [4117]. R.I.P. — S.P.Q.R [4168]. R/Sweave [3524]. Radhakrishnan [3669]. Radó [3072]. ragged [4574]. ragged-right [4574]. Rahilly [3346]. Rahtz [2091, 4117, 4167, 4168]. Raies [3960]. raise [124]. rally [4926]. Rambutan [2659]. Ranade [2709]. Random [4593, 3213, 699, 3582, 1682]. Randomising [4128]. rare [3876, 2801]. Raspberry [4534]. Raster [830, 1184]. rasterizer [4402, 4175]. reaches [4823]. Reaching [3549]. read [1710]. readability [3954]. Reader [2330, 4171, 3421, 3456]. readers [4981, 4285, 3658]. Reading [4576, 4308, 3623, 4332, 1866]. README.1ST [1905]. ready [371, 409, 1045, 4291]. Real [540, 649, 1729, 4217, 4471, 4472, 2316, 4944, 2163, 2781]. real-time [4944, 2781]. reality [3661]. really [1357, 1602]. realm [3518]. reappraisal [1965]. reasoning [1068]. Rebuttal [4849]. RECAU [268]. recipe [1612]. Recipes [348]. recognition [4513, 2713, 2875]. Recollections [4068]. recommendations [4595, 4713]. reconsidered [1215]. record [761]. records [834, 1040]. Recovering [816]. Recreating [3798]. rectangles [1482]. Rectangular [3425]. Recycled [1780]. rédaction [2771]. redefinition [93]. Redistributing [3877]. reduce [4936]. reduction [2572, 4606]. redux [1703]. Reference [2523, 2559, 2595, 2625, 2580, 528, 625, 3288, 4177, 4498, 3852]. references [1041, 3907, 4606, 1686, 4647, 3636]. referencing [2324, 4245]. Refining [1369]. Reflections [3594, 4983, 4308, 4907, 4713]. Réflexions [893]. Reflowing [4435]. réformes [2804, 2804]. Regarding [4848, 2012, 1902]. registers — an [3259]. regression [2102]. regular [3173, 3301]. reimplemented [4474]. related [2765, 4377, 1068, 427, 1942]. relational [3420]. relations [3249]. Relative [1657, 2728]. Relaunch [3794]. relax [280]. release [1873, 4505, 1873]. Released [4276]. Remaking [4179]. remarks [2816, 54, 488, 2681]. remarques [2770]. Remembering [4051]. Remembrance [2279]. remotely [4693]. Removing [3144, 4784, 4872]. rendered [3959]. Rendering [4755, 3919, 2074, 4979]. Rennes [879, 553]. repair [2317]. repeated [4786]. repertoire [3944]. Repetition [3425, 3801]. replacement [3882, 2787]. replica [2190]. Replicating [2766, 1416]. Replies [747]. Reply [2286, 702, 662]. Report [328, 1132, 4191, 4951, 1150, 2746, 37, 3998, 2909, 3390, 3008, 271, 910, 1129, 304, 339, 589, 1034, 644, 731, 813, 3037, 4803, 3087, 3329, 2159, 734, 76, 4046, 4166, 1073, 1131, 1164, 4686, 646, 4860, 977, 4595, 274, 18, 3149, 228, 4711, 4651, 307, 341, 420, 462, 779, 55, 4549, 3391, 3135, 3508, 259, 313, 391, 668, 26, 4701, 4918, 523, 457, 2432, 59, 314, 48, 458, 1123, 35, 861, 1157, 291, 3233, 461, 19, 56, 140, 2674, 163, 142, 231, 903, 378, 628, 1309, 460, 1702, 2264, 290, 2807, 2187, 73, 801]. report [2910, 145, 865, 909, 1072, 1128, 270, 642, 692, 729, 778, 812, 495, 33, 1529, 496, 537, 199, 239, 272, 306, 340, 380, 418, 790, 693, 811, 9, 867, 959, 1127, 866, 732, 815, 679, 121, 149, 172, 202, 240, 238, 1126, 415, 538, 814, 381, 377, 539, 1758, 2016, 20, 57, 102, 230, 2912, 1130]. Reporting [4561, 4373, 3524, 3260, 1331]. Reports [4994, 353, 393, 1710, 736, 4086, 1044, 1479, 4304, 417, 156, 184, 222]. repository [3876]. Representation [4654, 4437]. Representing [4855, 3848]. Reprint [1088]. Reprise [3370]. reproduction [3094]. Republic [1904]. Request [696, 2425]. requirement [1656, 1686]. requirements [3703, 2786, 4301]. reresponse [248]. res [769]. research [1614, 2182, 2271, 2305, 2354, 2410, 2492, 2560, 2626, 2084, 2117, 2185, 2216, 2244, 2274, 2306, 2356, 2391, 2413, 2444, 2478, 2488, 2524, 2563, 2597, 2629]. researchers [3142, 3925]. reset [1788]. Reshaping [3293]. residual [4852]. resisted [869]. resolution [3988]. resources [1031, 2982, 3141, 4021, 2819, 1074, 3104]. respect [3095]. Response [2426, 3379, 248, 3378, 1332, 2283, 1333, 1560]. responses [1064]. responsive [4953]. Restricting [1761]. results [4171, 4425, 2824, 3729, 473]. Resurrecting [3123]. retrospective [3588, 2927]. Return [2684]. returns [2040, 535]. reusable [3201]. Reutenauer [3462, 4189]. Revealing [4284]. reverse [3326]. Review [2711, 2844, 4260, 4306, 4261, 1653, 4240, 4307, 4327, 4255, 4275, 4258, 4256, 4259, 4308, 4410, 4309, 4257, 4130, 3639, 3851, 3992, 1712, 3993, 834, 3714, 3962, 4543, 4544, 4992, 3994, 2368, 4615, 4661, 2831, 4990, 4964, 1444, 3894, 3304, 3995, 1446, 4475, 4660, 4696, 3599, 3640, 3641, 3665, 3678, 3679, 3715, 3747, 3776, 3811, 3849, 3850, 3852, 3853, 3931, 4029, 3303, 4578, 3716, 3777, 3893, 4028, 4345, 4991, 454, 422, 1310, 1311, 1582]. reviews [4659, 4910, 4911, 4788, 1445, 3930, 1004]. revised [409, 503, 242]. revisited [4434, 3803, 3327, 1418]. Revisiting [2610]. revival [2191]. revived [3292]. Revolution [1995]. REVTEX [2532]. Rhapsody — EuroTEX [3051]. Rhatigan [3346]. Rich [3423, 2424]. Richard [4053]. Richter [2686]. riddle [1317]. RIDT [1184]. RIDT’89 — International [830]. right [3907, 576, 699, 4574]. right-to-left [576]. Rio [4392]. Rishi [3669]. rites [1316]. River [2871, 2943, 3433, 3536]. rlap [2584]. road [948, 4947, 4728, 869, 4600, 4536]. roadmap [4955, 3212]. Robert [4308, 4187, 4659]. Robertson [3742, 3960]. Robin [4779]. rocket [4067]. roi [2251]. Rokicki [1734]. role [2319, 3345]. Rolland [1707]. rollback [4395]. Rolling [3140]. ROM [2291]. Romain [2251]. Roman [599, 278, 4256, 4599, 4209]. romance [2923]. Romanian [1806]. Romanized [2229]. romannumeral [4141]. Romano [4029, 4578]. ROMs [3447]. Ronald [1905]. root [541]. roots [4491]. Rosenthol [3210]. Roswitha [2280]. rotated [1443, 1518]. Round [963]. Roundtable [3727]. routine [2786, 3956]. routines [1042, 1079, 1167, 1679, 3659, 3369]. Rowley [2966, 3210, 3430, 3534]. Royal [922]. RPM [2924]. RTF [3024]. R & [203]. Rubricella [4978]. Ruddy [3534]. rules [4373, 2692, 2983, 3589]. Ruling [780]. run [295, 3907]. Rundell [3534]. Running [3831, 421, 695]. runtimes [4497]. rusmv1.rus.uni [1274]. rusmv1.rus.uni-stuttgart.de [1274]. Russia [1963, 1963, 4057]. Russian [1286, 1872, 1993, 2459, 942, 1628, 1303, 3173, 1935]. Russian-speaking [1628]. russianb [1870]. Russians [1093]. RUSTeX [1286]. RUSTeX-L [1286]. rustic [4938]. Ryćko [3274]. RyDArab — Typesetting [2890].

S [1208, 3995, 37, 764, 3589]. safer [1803]. SäferTEX [2830, 3033]. Safety [4252]. SageMathCloud [4247]. SageTEX [4128, 3661]. Salomon [1967]. Samarin [1705]. same [3290]. sample [193, 264, 300, 335, 373, 411, 671, 727, 772, 4725, 196, 107]. samples [502, 50]. Samuel [1146, 1401]. Sanitizing [1173]. sanity [4752]. sans [4145, 3587]. Sanskrit [2618]. Santa [1635, 1688]. SAS [645]. Saturday [598, 657]. savoir [4145]. Sawyer [1966]. Say [4087, 4287, 1222]. says [1398, 3457]. sbMF [1226]. sbTEX [1226]. scalable [4338, 2096]. Scaled [4689, 2376]. scales [747]. scaling [4176]. scanners [2457, 4026]. scanning [4653]. SCANTEX [631]. scatter [783]. scatterplots [3846]. Scenarios [4298]. scenes [4941]. Schedule [3019, 1497, 1914]. schema [2773, 4953]. Schematics [2068]. scheme [1171, 1172, 916, 1452]. scheme — Latin [1151]. Scholar [1576]. scholarly [3124]. Scholarship [1391, 1494, 791, 1576, 843, 1023, 1768, 448, 521, 569, 715]. Schools [4287, 1334, 4087, 1222]. Schröder [3072, 3462]. Schulenburg [2280]. Schwabacher [1217]. Science [2944, 3219, 3537, 353, 393, 4433, 3811, 2021, 3952, 3432]. sciences [1338, 1522]. Scientific [3218, 433, 451, 556, 3925, 312, 351, 4539, 1377, 2935, 624, 3843, 1213, 2169, 4247, 1991]. scientifique [2772]. scientist [994, 3469, 4807]. Scientists [1966, 1311]. scope [1761]. scoping [3259]. score [4185]. Scott [4237]. screen [725]. Scribe [967]. Scribe-to-TEX [967]. script [3124, 634, 4572, 1617, 2653, 4956, 4541, 2793, 2889, 241, 3759]. script-driven [634]. scripting [2859]. scripts [4467, 501, 4642, 4511, 2192]. scrpage2 [3765]. SDDL [3528]. search [4722, 2372, 128, 4494]. searchable [2611]. Searching [1511]. seasonal [4344, 2220]. Seating [154]. Sebastian [4117, 4168, 2091, 4167]. Second [2741, 4696, 1122, 4788, 2790]. secretaries [3262]. secrets [4442]. section [3801]. secure [1970]. See [1328, 1420]. seeks [1438]. seen [3530]. segmentation [4955]. selected [702, 604, 662]. Selection [3920, 916, 3176, 4647]. selection — User [1046, 1088]. Self [1416, 4819, 3412, 2457, 3469, 2541]. self-marking [2541]. self-parsing [2457]. Self-publishing [4819, 3412, 3469]. Self-replicating [1416]. sells [2397]. semantic [4828]. semantic/active [4828]. Semantics [1611, 4284, 2965, 2621]. sémantiques [2776]. semaphore [2365]. semi [4438, 545]. semi- [4438]. semi-parameters [545]. Seminar [208, 4259]. Semitic [1284]. sens [2770]. Sensitive [490, 1606]. separate [1202]. separation [1736, 1442]. September [882, 1151, 1183, 1337, 1424, 1464, 1689, 1790, 2499, 2719, 953, 825, 1883, 700]. sequence [2513]. sequences [1448, 1173]. Serbo [1304]. Serbo-Croatian [1304]. serendiPDF [2611]. Series [495, 2149, 1110, 4692, 910, 1129, 589, 379, 730, 588]. Serif [3587]. Serifed [4328]. Server [3063, 810, 2620]. server — Update [908, 953, 1033]. servers [4851]. service [218, 4620, 4607]. services [2843, 4266, 4313, 4348, 4383, 4413, 4446, 4480, 4518, 4551, 4664, 4585, 4630, 4704, 4740, 4767, 4795, 4839, 4879, 4993, 4917, 4966, 3007, 3076, 3116, 3153, 3188, 3220, 3277, 3316, 3350, 3394, 3436, 3464, 3502, 3539, 3569, 3607, 3646, 3684, 3720, 3748, 3784, 3816, 3854, 3896, 3934, 4002, 4035, 4075, 4109, 4154, 4194, 4224, 4532, 1234, 1430, 1542, 1569, 1594, 1664, 1695, 1722, 1770, 1796, 1820, 1836, 1860, 1889, 1922, 1950, 1979, 2005, 2031, 2059, 2081, 2111, 2139, 2177, 2213, 2240, 2269]. services [2303, 2351, 2387, 2407, 2439, 2474, 2485, 2522, 2558, 2594, 2624, 2635, 2678, 2715, 2755, 2810, 2869, 2914, 2950, 2970, 514]. serving [4655]. servitude [2770]. session [2499, 1132, 4983, 21]. Set [4334, 1066, 4612, 2291, 1629, 873, 1935, 1402]. sets [761]. setting [528, 625, 4945, 1813, 1340, 4821]. setting — an [143]. settings [579, 635, 684, 3095]. setup [3635, 3473]. several [1805]. SGML [1053, 1366, 1612, 1044, 1330, 1458, 1445, 2320, 1476, 4619, 1212, 1365, 1213, 1905, 667, 2070, 1677]. SGML-L^{A}TEX [1044]. SGML — Questions [1456]. shaded [1482]. Shady [4345]. shallow [3066]. shapes [1841, 1036]. Sharpe [4324]. Sheep [4911, 1676]. sheets [1531, 1533]. shell [965]. shift [1105]. Shikano [3887]. shiny [3948]. Shop’s [4066]. Short [4651, 185, 227, 4173, 2288, 3843, 4977, 4873]. shortcuts [1869, 1901]. shorten [4873]. shorthand [3454, 3341]. should [4271, 4390, 4928, 726, 775, 4619, 1370, 1371]. Showing [1039]. Showing-off [1039]. SHSU [1272]. SHSU.bitnet [1272]. SI [3595]. Siamese [1387]. Side [3800, 3063, 3054, 3593, 502, 3140]. Side-by-side [3800, 502]. sides [1355]. Siegman [3695]. sign [3035]. signals [3951]. signes [2777]. SIL [3197]. SILE [4241]. silly [4820]. Silmaril [2939]. simple [4721, 2856, 3985, 3915, 4131, 2252, 1737]. simplifies [1932]. Simplifying [4984]. Simultaneous [1364]. simultaneously [3778]. sing [1557]. Single [4056, 3252]. Single- [4056]. Single-source [3252]. Sinhalese [1781]. Site [457, 458, 461, 460, 73, 145, 865, 909, 1072, 1128, 270, 271, 642, 692, 729, 778, 812, 910, 1129, 304, 339, 495, 589, 33, 496, 537, 199, 239, 272, 306, 340, 380, 418, 1034, 644, 731, 813, 693, 811, 2398, 9, 867, 959, 1127, 734, 1073, 1131, 1164, 866, 1130, 646, 732, 815, 121, 149, 172, 202, 240, 274, 238, 1126, 415, 538, 814, 307, 341, 381, 420, 462, 779, 377, 539, 17, 1880]. Sites [2347]. siunitx [4398, 4822, 3595]. Sixty [3637]. Size [2572, 211, 3455, 3761, 4210]. sized [2767]. Sizes [477, 3659]. sizing [466]. Skoupý [3669]. Sky [2182, 2271, 2305, 2354, 2410, 2492, 2560, 2626, 2084, 2117, 2185, 2216, 2244, 2274, 2306, 2356, 2391, 2413, 2444, 2478, 2488, 2524, 2563, 2597, 2629]. Slanted [1683, 4214]. Slavonic [1871]. slide [3292]. slides [1086, 430, 47, 244, 1085]. slightly [3948, 3774]. slightly-shiny [3948]. SliTEX [430]. slow [1149]. Small [478, 419, 243, 279, 309, 343, 382, 67, 2899, 3801]. Smart [3338, 2958]. smash [2584]. smile [4963]. Smith [3829]. Smoky [4088]. snippets [4731]. Snow [1520]. Society [422, 2532, 3287]. Software [2872, 2948, 77, 3304, 1092, 610, 750, 2603, 853, 2746, 4987, 2333, 4373, 3594, 1208, 2545, 911, 630, 3433, 453, 803, 2707, 2736]. Software-ergonomics [803]. Sojka [3669]. Soliciting [1896]. Solution [1524, 425, 1442, 3875, 4949, 1038, 1317]. solutions [1166, 4725, 1870]. solving [2929]. Some [4690, 1025, 2730, 383, 1048, 802, 3483, 3553, 3597, 3674, 3705, 212, 699, 768, 4308, 1221, 1947, 1314, 1415, 1809, 180, 577, 3728, 782, 65, 2450]. something [1906]. sometimes [2981]. Sonata [2726]. soon [4561]. Sophia [609].

Sophia-Antipolis [609]. Sophisticated [1752]. Sorting [4308, 4732, 1626]. sound [2793]. sound-script-language [2793]. sounds [4276]. Source [2505, 2830, 3033, 502, 412, 4783, 1850, 3734, 2707, 3198, 4815, 4460, 46, 42, 3252, 2568, 3623, 2540]. source/output [502]. sources [4562, 3271, 2824, 2964, 2193, 2017]. Southall [4053]. Souza [3430, 3960]. Space [1522, 2575, 1338, 1624, 4068]. SPAN [1273]. SPAN/DECnet [1273]. Spatial [2961, 1731]. speaking [3452, 1628]. speaks [3093, 4365]. Special [1536, 450, 291, 502, 1764, 654, 2094, 1388, 1735, 720, 3231, 610, 2206, 302, 689]. specialized [3986]. specials [3359, 1875]. specific [3237]. Specification [1611]. specifications [3168, 30]. Specifying [1367]. Specimen [3639]. speech [3209]. speedata [4611]. spelling [735]. Spending [4702]. Sperry [498]. Spheres [3372]. Spidrons [3370]. Spiekermann [1676, 4911]. Spindex — Indexing [2094]. spirit [3414]. Spivak [4751]. spoken [3657]. sponsor [4994]. sponsors [4550, 4629, 4739, 2849, 3050, 3086, 3129, 3225, 3321, 3399, 3507, 3654, 3725, 3865, 3941, 4042, 4162, 3114, 2925, 2954]. spooler [337]. spot [4824]. spreadsheet [4335]. Spring [376, 610, 74]. Springer [3217]. spurious [4068]. SQL [3135]. Squeezed [4728]. ST [693, 811, 1252, 586, 536, 803]. St. [3637]. stability [1929, 4496]. Stacks [2922]. staff [247]. stage [4067]. stand [2690, 3409]. stand-alone [2690]. standalone [4786]. standard [1329, 4299, 63, 1046, 1088, 3059, 1047, 1875, 1408, 768, 1309, 2206]. standardised [2765]. Standardizing [4940]. Standards [894, 1157, 1067, 681, 861, 4401, 903, 3668]. Stanford [1016, 185, 226, 352, 353, 393, 368, 610, 840, 59, 111, 163, 260, 412, 31, 169, 3510]. Stanley [1966]. Star [35, 3738, 112, 71]. starfont [4178]. started [3407]. starters [2885]. StarTEX — a [1940]. starts [4808]. starttext [2983, 2886]. Staszek [4360]. State [4442, 269, 416, 2608, 1987, 2855, 103, 4495, 4651, 4277, 4331, 3477]. Statement [2386, 405, 139, 2519, 4726]. statements [4382, 4481, 4584, 4703, 4794, 4919, 3722, 3818, 3932, 4033, 4263, 4153, 2675, 2753, 2911, 3006, 3150, 3311, 3392, 3499, 3610, 1343]. static [3959]. station [714]. Statistical [3956, 1208, 542]. Statistics [3458, 3582]. Status [31, 3647, 4617, 2508, 80, 34, 120, 4860, 2016, 7, 26]. Stealing [4911, 1676]. Steering [19, 56, 14]. Stefan [4788, 4991]. stems [4214]. Step [1933, 2699, 781]. Stephan [1674]. steps [2340]. Stereographic [1787]. Stern [3210]. Steve [2946]. Steven [1966]. sthetics [3526]. STI [247]. Still [4026, 4136, 1906, 718]. STIX [3196]. STM [4550, 4739, 4838, 4965]. Stock [4544]. Stock-Allen [4544]. Stockholm [922]. Stop [1676, 4069, 4911]. Storing [4909]. story [3549, 4820, 1829, 2202, 2708, 1831]. Stössinger [4365]. strait [1943]. Strasbourg [1338, 433, 556]. Strategies [3470, 4129, 2930]. Stream [427]. streams [1040, 3918]. street [4311, 4755]. stress [4498]. Streszczenia [4304]. stretch — mimicking [3144]. string [4096, 913]. stringing [3125]. stroke [3954]. stroll [4809]. StructElem [4728]. structural [1814, 2572]. structure [166, 1378, 1315, 3045, 1367, 1876]. structured [3528, 1368, 3204, 2074, 4952, 1103]. Structures [552, 4094]. struggles [2928]. strut [210]. Stubborn [3774]. student [311, 4431, 4095, 4063]. students [3624, 2706]. studies [3360]. study [610, 4288, 4978, 2707, 3666, 641, 3873, 2837]. Studying [4531]. Stuttgart [1274]. stuttgart.de [1274]. Style [1611, 3990, 4442, 1329, 874, 825, 3846, 601, 1202, 1176, 3882, 2341, 3667, 3765, 2459, 1044, 1870, 1177, 2936, 1047, 1443, 1518, 1533, 1704, 1741, 1417, 593, 655, 700, 744, 786, 2454]. style-option [1329, 1704]. styles [1266, 3175, 3044, 2789, 3332, 3169, 1047, 3518, 4179, 969, 1329]. stylesheet [3664]. Styling [4901]. sub [1177]. sub-style [1177]. submission [2325]. submissions [4816, 1941]. submit [4530]. submitting [135, 158, 292]. subscriptions — What [189]. subscripts [3489]. subset [4618]. subtle [4284]. suburban [1192, 1193]. subway [4170]. successful [1362]. successor [3036]. Sudokus [3299]. suffice [3907]. sugar [1625]. suggested [770, 1686, 805]. Suggestions [3289, 3241, 1996]. suite [2102, 3987]. Sulger [3217]. summaries [4240, 4307, 4327]. Summary [68, 226, 428, 1622, 117, 1031, 249, 191, 234]. Summer [185, 555, 163]. summons [4108]. Sun [273]. Sunday [2499]. Super [2695]. Supercomputing [900]. superscripts [3489]. Supplement [1191, 1058, 929, 981, 1096, 2414, 795, 839]. Support [3203, 2713, 2875, 4690, 4514, 2199, 4722, 2609, 1743, 2076, 3340, 58, 3983, 2837, 3360, 3876, 1656, 1735, 28, 2772, 2790, 23]. Supporting [4397, 3369, 4176]. surgery [3070]. survey [762, 4425, 4243, 4211, 3837, 4939, 771]. Suspension [4191, 4121]. sustainable [4949]. SUTRA — a [3951]. SVG [2569]. sweat [1828]. Sweave [3524]. Swelled [2983]. SwiExr [2961]. SWIGLIB [4025, 4099]. Switzerland [1339, 830, 1392, 3637]. syllabic [4955]. symbol [4294]. Symbolic [4898, 1746]. symbols [4176, 4336, 371, 409, 1930, 1873, 2767]. symposium [4361, 74]. synchronization [3326]. SyncTEX [3326]. Syntactic [1625]. syntax [1610, 1651, 1013, 2820, 2327]. synthesis [2697, 2768]. synthetic [4539]. synthétiques [2804]. SYSDEP [29, 30]. system [4501, 1755, 500, 3062, 1329, 2101, 3559, 4241, 3880, 1763, 1486, 3235, 3985, 990, 1602, 1869, 1901, 1103, 3290, 1930, 4400, 630, 2619, 4718, 2480, 4245, 1550, 2325, 3838, 1708, 958, 1558, 1440, 1781, 1629, 2531, 1477, 2204]. system — [1368]. system — TEX [3089]. Système [2771]. Systems [317, 609, 516, 4721, 31, 2166, 1030, 1071, 3021, 252, 688, 36]. Systems — Lasergrafix [219]. Szabó [3072].

T [1475]. T-EDIT [1475]. Table [2756, 3016, 44, 4127, 2487, 1711, 153, 1652, 4894, 4296, 3978, 626, 966, 3774, 782, 4096, 2811, 2870, 2976, 419, 243, 279, 309, 343, 382, 499, 1082, 3211, 3275, 1517]. table-of-contents [966]. tables [183, 1476, 3207, 3625, 4458, 3632, 3599, 1613]. tabular [827]. tabular — environments [875]. tabularray [4894]. tactics [4129]. Tagged [4951, 3556, 4645, 4860, 3403, 2170, 4728, 4952, 4506]. Tagging [4621, 4950]. Tai [3379]. Takata [3887]. take [4344, 2928]. takes [3592]. Taking [4614, 4457]. Talaris [361, 316]. Talaris — Talaris [513]. Talbot [3481, 3930, 4144]. tale [3868]. talk [4361, 1388, 4127, 1227, 817]. talks [316]. tame [265]. tamed [1617]. TANGLE [469, 682]. TANSU — a [3881]. Tao [2605]. TAOCP} [3975]. tape [63, 356, 135, 158, 251, 399, 181, 216, 336]. target [3662]. targeting [4182]. tasks [4725, 2786]. Taupin [2723]. taxonomy [3021]. Taylor [2966]. TDS [2551]. teach [4606]. Teaching [1851, 3624, 3844, 2706, 4457, 2764, 4605, 895, 1627]. team [2315, 2053, 2103, 2283]. Team — A [4616]. techexplorer [2475, 2486, 2162, 2440]. Technical [562, 563, 619, 1441, 226, 422, 4808, 1529]. technique [2568, 2252]. techniques [3071, 2824, 1042, 1079, 1167, 1679, 2540, 2670, 1809, 1104, 1214, 1079, 1679]. technologies [4827, 4620, 2871, 2943, 3433, 3536, 2958]. Technology [922, 208, 2021]. Tectonic [4811]. teenage [3229]. TEI [1367, 3025, 4729]. TEI-XML [4729]. TEI/TEX [3025]. Telegram [4186]. Tellechea [3993]. Tempest [4812]. template [1760, 4936, 4063]. Templates [490, 3063, 3882, 2730, 4936, 4057, 995, 1997]. Templating [4470]. Ten [4302, 1654]. Tenth [1182, 840, 836]. Terada [3887]. terdragon [4869]. terminals [1149]. terms [2265, 4104]. test [3949, 2102, 3987, 4498]. testidx.sty [4340]. Testing [4562, 213, 4340, 4098, 699]. tests [3269, 1970, 937, 1319]. teTEX [2332]. teTEX-based [2332]. Tétreault [3960]. teubner [2652]. TEX [465, 451, 459, 470, 471, 3510, 3493, 955, 640, 4740, 4928, 4905, 4985, 3730, 1065, 1147, 4994, 1287, 1293, 1283, 1288, 1291]. TEX& [2020]. TEX&Co. [586]. TEX-alike [4814]. TEX-aware [3527]. TEX-based [580]. TEX-collection [3031]. TEX-compatible [1583]. TeX-D-L [1287]. TEX-Ed [952]. TEX-editor [850]. TEX-enabled [3737]. TeX-EURO [1288]. TEX-formatted [1483]. TEX-Interessenten [694]. TEX-L [640]. TeX-NL [1293]. TEX-Pascal [26]. TEX-PostScript [2195]. TEX-PostScript [960]. TeX-Pubs [1291]. TEX-related [1942]. TEX-to-PostScript [354]. TeX.SE [4326, 4367, 4430, 4458]. TeX.StackExchange [4375, 4470]. TEX/Pascal [29, 30]. TEX/PostScript/graphics [1109]. TEX2.1 [570]. TEX2PDF — Acrobatics [1960]. TEX3 [4661]. TEX370 [170]. TEX4ht [4473, 4858, 4857, 2854]. TEX80 [385, 378]. TEX82 [166, 185, 216, 337, 379, 220, 510, 588, 253, 285, 323, 360, 398, 441, 237]. TEX88 [705]. TEX89 [882]. TEX91 [1183, 1337]. TEX91/Congres [1183, 1337]. TEXarcana [92, 132, 155]. Texas [1133, 204, 375]. TEXbook [1967]. TEXcel [4183]. TEX — A [1965, 4679, 2535, 1946, 858]. TEX — Discussion [374]. TEX — Introduzione [1675]. TEX — Quo [1848]. TEX — the [4288]. TEXer [2316]. TEXhax [374, 338, 376, 191, 234, 1289]. Texinfo [3946, 4498]. TEXline [905]. texlive.info [4532]. TEXMaG [1290]. TEXML [3123, 2366]. Texmlbus [4718]. TEXnical [1104, 1214, 1963, 1114]. TEXnician [113]. TEXnicolour — A [1177]. TeXnische [3743, 4071, 4260, 4306, 4346, 4380, 4406, 4445, 4623, 4736, 4697, 4791, 4876, 4913, 4545, 4834, 1561, 1715, 2049, 2098, 2130, 2258]. TeXnology [4768, 4880, 4916, 4796]. TEXon [3361]. TEXpert [128]. texpic} — Design [1003]. TEXreation — playing [1009]. TEX’s [3323]. TeXShop [4856, 2689, 3947, 2553, 4130, 4429]. TEXspec [2545]. TEXstudio [4132]. Text [389, 4754, 4721, 3561, 3088, 3232, 474, 4434, 1873, 1207, 4458, 3270, 3344, 699, 1370, 3837, 1412, 94, 2337, 1532, 2336, 1169, 3425]. TEXT1 [1037]. textbook [1361, 4181, 1111, 1356]. textbooks [4783]. TEXtension [128]. textpos [2665]. Texts [1310, 1671, 3091, 1616, 1672, 576, 3094, 4020, 2200]. Textset [218, 250, 284, 318, 354, 357, 395, 436, 515]. textual [4437]. texture [4753]. TEXTURES [647, 648]. TEXware [357]. TEXworks [3325]. TEXworks — As [3621]. Thai [989, 2423]. thank [833]. Theatre [2189, 4812]. them [1828]. Thematic [2667]. themes [4731]. theorem [971]. theorems [3714]. Theory [2129]. there [4020, 4059, 2727, 1906]. Theses [3471, 4826, 4062, 4063]. thesis [3733, 3732, 3930]. thickness [1683]. Thiele [3742]. Thing [4682]. Things [4308, 3552, 103, 3297]. think [3142]. 16 [1310]. Third [4963, 541]. Thirty [3523]. Thiruvananthapuram — Report [2599]. Thomas [4085, 4475]. Those [3238, 1386]. Thoughts [3511, 4529, 632]. Three [4376, 3552, 3713, 3368, 4067, 2577, 1161]. Three-dimensional [4376, 3713]. three-stage [4067]. Threshing [2194]. Ti [3138, 3414]. Tiger [557]. Tiger-TEX [557]. Tight [1813]. TikZ [4900, 4979]. TikZ/PSTricks [4376]. tikzducks [4394]. TikZlings [4941, 4717]. tiling [2151]. tilings [2251]. Time [659, 747, 2449, 4602, 3726, 4944, 3805, 2781]. Timelines [3704]. Times [1933, 991, 3907]. timings [1165, 1523]. Timothy [1905]. TinyTeX [4464]. TIPA [1930]. tipografico [3678]. tips [4826, 3674, 3705]. Title [1, 98, 324, 365, 400, 444, 482, 518, 564, 620, 674, 710, 756, 796, 841, 889, 930, 982, 1018, 1060, 1098, 1138, 1194, 1297, 1350, 1351, 1395, 1433, 1468, 1502, 1544, 1571, 1596, 1640, 1797, 1863, 1893, 1951, 1954, 1983, 2009, 2034, 2113, 2140, 2180, 2217, 2245, 2275, 2357, 2408, 2415, 2490, 2525, 2630, 2640, 2680, 2720, 2761, 2847, 2881, 2918, 2953, 2977, 3049, 3079, 3119, 3157, 3192, 3224, 3281, 3320, 3355, 3398, 3465, 3506, 3545, 3575, 3617, 3653, 3724, 3755, 3789, 3824, 3864, 3901, 3940, 3970, 4041]. Title [4080, 4115, 4160, 4199, 4271, 4232, 4318, 4390, 4354, 4420, 4452, 4487, 4524, 4634, 4556, 4591, 4669, 4709, 4746, 4774, 4801, 4843, 4885, 4924, 4972, 4008, 703]. titles [747]. TLaunch [4290]. tlcontrib.metatex.org [3591]. tlmgr [3948]. today [4950, 4493, 2687]. tofu [4757]. together [3661]. together — EuroTEX [3017, 3018]. tokens [4026, 1412]. Tokyo [3867]. Tomaszewski [3274]. tomorrow [2687]. Tony [3695]. Too [1526, 3248]. tool [3206, 741, 3271, 2545, 4727, 3094, 2542]. toolbox [2860]. toolkit [1745]. tools [1709, 1314, 3060, 4535, 735, 4693, 4981, 1330, 2075, 2824, 1972]. topic [4280, 2819, 1446]. Tops [201]. Tops-20 [201]. Toronto [4170, 4165]. Torun [2145]. Toruńska [2190]. totally [384]. touch [2565]. tough [1652, 1711]. toujours [4145]. Toulouse [1052]. Tour [2022, 113, 1911, 1972, 2232]. tourists [342]. tours [2257]. Tout [4145]. Tove [4931]. tower [425, 1418]. TPHON [82]. trace [10, 2514, 2612]. traces [1900]. Tracing [4219, 3949]. Track [4257]. tracklang [4216]. Tracy [4188]. Traditional [2018, 3144, 3091, 2190, 1356, 3838]. traditions [1993]. trail [4431]. Training [3219, 2670, 4614, 525, 641, 679, 3657, 4509, 2874, 2947, 3432, 3536]. traitements [2772]. Tralics [2774]. Transatlantic [387]. transcribing [4727]. transcript [4752]. TransFig [1106]. transformation [3924]. transition [2422]. transitions [4688]. translating [1154]. translation [4513, 2170, 94]. translator [2774]. Translators [455]. transparencies [877]. Transparent [4787]. Trap [4562]. travelogue [2599]. Treasure [2232, 2257, 2509, 2582, 4254, 4310, 4342, 4379, 4411, 4444, 4478, 4547, 4662, 4582, 4625, 4700, 4737, 4762, 4789, 4830, 4875, 4912, 4988, 3174, 3302, 3380, 3490, 3562, 3600, 3643, 3680, 3717, 3746, 3775, 3809, 3842, 3891, 3929, 3991, 4027, 4103, 4143, 4220, 2662, 2740, 2901, 2996, 3126, 2374, 2429]. Treasurer [20, 57, 102, 230, 2432, 977, 668]. treatment [1686]. tree [1327, 2985, 677]. trees [3139, 3714, 347, 2551]. trends [1991]. Trials [4016]. trick [1041]. tricks [1654, 1318, 1413, 1450, 1525, 1553, 1655, 1681, 1785, 3765, 3483, 3553, 3597, 3674, 3705, 2838, 2401, 2512]. Tricky [4339]. Trilingual [4057]. trip [2187, 801, 679, 4562, 3949]. tristesse [1211]. triumvirate [3301]. Trotter [4408]. trouble [2981]. Troubles [10]. Truchet [2251]. Truck [4188]. TrueType [2656, 2364, 3473, 3364, 2799, 4866]. Truth [1586, 423, 467]. trying [4530]. Tsukurimashou [3871]. TTN [1868, 2040, 2040]. TUGboat-Artikel [972, 1049, 1089, 1117, 1179]. Tufte [3846, 4259]. Tufte-style [3846]. Tufts [442, 443]. TUG [845, 974, 1093, 1193, 1493, 1631, 1882, 1887, 2028, 2136, 2265, 2528, 2601, 227, 315, 320, 834, 840, 1032, 1075, 1161, 2187, 801, 2432, 473, 571, 679, 62, 6, 19, 21, 56, 140, 1050, 61, 163, 244, 2122, 2264, 20, 434, 165, 4734, 4515, 442, 709, 1192, 1058, 2642, 2717, 2845, 2842, 3193, 3866, 4043, 4044, 4163, 4164, 4273, 4385, 4488, 4489, 4490, 4550, 4592, 4629, 4738, 4710, 4739, 4838, 4802, 4965, 4925, 887, 927, 929, 978, 981, 1015, 1056, 1092, 1095, 1096, 1136, 1185, 1189, 1229, 1233]. TUG [1343, 1345, 1346, 1372, 1374, 1393, 1428, 1438, 1465, 1496, 1497, 1541, 1563, 1568, 1635, 1636, 1637, 1663, 1688, 1694, 1724, 1725, 1771, 1795, 1819, 1880, 1896, 1899, 1914, 1917, 1918, 1919, 1921, 1957, 1978, 2004, 2030, 2058, 2110, 2138, 2212, 2239, 2268, 2302, 2309, 2350, 2384, 2406, 2438, 2447, 2473, 2484, 2521, 2499, 2527, 2593, 2599, 2600, 2677, 2712, 2719, 2754, 2807, 2867, 2968, 2973, 3009, 3013, 3074, 3115, 3147, 3152, 3186, 3187, 3216, 3276, 3312, 3313, 3315, 3347, 3349]. TUG [3351, 157, 185, 3389, 3430, 3431, 3434, 3443, 3444, 3649, 3685, 3687, 3721, 3756, 3781, 3785, 3826, 3858, 3860, 3933, 3936, 3963, 3964, 3965, 3999, 4000, 4001, 4003, 4036, 4070, 4073, 4074, 4152, 4192, 4193, 4201, 4223, 352, 368, 428, 439, 440, 443, 509, 560, 614, 663, 670, 707, 749, 752, 793, 795, 836, 837, 839, 3194, 4831, 3390, 3511, 3655, 4404, 4382, 4481, 4584, 4703, 4794, 4927, 4919, 2682, 3411, 3697, 3722, 3932, 4033, 4191, 1623, 4951, 4032, 4492, 4712, 4190, 79, 4264, 4415, 4447]. TUG [4362, 4622, 4779, 4848, 3323, 4957, 3611, 4803, 3087, 4263, 4153, 4274, 3943, 4166, 4804, 2955, 3726, 3942, 141, 58, 4305, 4595, 4713, 4807, 2674, 4272, 4391, 3149, 3867, 4165, 4711, 4482, 1783, 1601, 2335, 4983, 4926, 3311, 3392, 57, 102, 156, 184, 222, 230, 259, 391, 4016, 4393, 4392, 4701, 2145, 3669, 3742, 3887, 3960, 4034, 4189, 4046, 2080, 2345, 2472, 2481, 2517, 2556, 2588, 2590, 2638, 2679, 2750, 2806, 2865, 2908, 3500, 3540, 3541, 3609, 3638, 3691, 3749, 3815, 3817, 3819, 3791, 3895, 4107, 2747, 2186, 2912]. TUG [3818, 1987, 2520, 2371, 3610]. TUG2000 — The [2307, 2393, 2436]. TUG’2001 [2445]. TUG2004 [2639]. TUG2019 [4491]. TUG2020 [4593]. TUG’91 [1389]. TUG94 [1776]. TUG’95 — List [1858]. TUG’96 [1856]. TUG’97 [1976, 1986, 2027, 2056, 2065, 2078]. TUG’98 [2105, 2188]. TUG’99 [2378, 2175, 2210, 2237, 2248, 2260]. TUGboat [4872, 4893]. TUG — Job [563, 619, 562]. TUG — TEX [356, 437]. TEX Users Group — TUG [229]. tuneup [4675, 3284, 3905]. Tuning [3482, 578]. Turbo [852]. TurboTEX [737]. Turbulent [2422]. tures [784]. Turkish [1246, 1294, 719]. Turner [1905]. Turtle [1946]. Tutorial [4371, 739, 819, 3414, 2887]. tutorials [3657]. Twelfth [1192, 1193]. Twentieth [3640]. Twenty [3975]. Twice [695]. Two [1355, 47, 4099, 3979, 2368, 4601, 1044, 4896, 4298, 4731, 346, 4873, 1648, 1904, 1075]. two-column [346]. Twombly [4544]. Type [3639, 1402, 2650, 4170, 1843, 4540, 2193, 3167, 3239, 2800, 1845, 1733, 1811, 1812, 4810, 2612, 2695, 4365, 3777, 208, 4721, 4544, 3055, 1747, 4911, 4977, 4890, 4328, 4117, 4649, 2824, 2645, 4494, 2696, 2190]. Type-1 [1733, 1812]. Type1 [2450, 2800]. Typecase [2018]. Typeface [4343, 531, 3698, 4014, 3909, 4187]. Typefaces [3962, 1931, 3475, 3587]. Types [4256, 2765, 2794, 4280, 3747, 427, 1368]. Typeset [2068, 2197, 3989, 2836, 4503]. typesetter [1363, 1745, 4814]. typesetters [3033]. Typesetting [1487, 3085, 944, 821, 2320, 4336, 1671, 1670, 2652, 2839, 2585, 2827, 1616, 1784, 4572, 4611, 1217, 3458, 1388, 2164, 2700, 3207, 2312, 4299, 3952, 3875, 4541, 2204, 3248, 3599, 4864, 4100, 486, 580, 3144, 1805, 3124, 4723, 2893, 992, 4241, 2896, 3091, 2632, 2726, 1176, 1763, 1486, 1910, 1102, 3092, 3985, 1068, 540, 649, 2786, 3100, 1602, 2767, 653, 2734, 3870, 802, 1356, 3102, 2830, 3021, 2619, 1991, 2480, 2252, 1417, 2987, 4944, 3461, 4574, 3468, 3377, 2676, 511, 2021, 2451, 1366, 1672, 949, 1029, 1730, 2548, 855, 878, 1083]. Typesetting [2224, 3873, 2653, 2290, 1382, 2618, 1549, 2801, 2957, 596, 1993, 2534, 2701, 1871, 487, 488, 2962, 1191]. typesetting — Challenges [2616]. typewriter [1517]. typewriting [4511]. typical [4014]. typist [1363]. typists [1114]. Typographer [4045, 4239, 4364, 4428, 4542, 4573, 4640, 4680, 4853, 4892, 4976, 2090, 2150, 2223, 2287, 2363, 2501, 2888, 2984, 3164, 3286, 3362, 3472, 3696, 3796, 4017, 4091, 4205]. Typographers [3847]. Typographic [4583, 4240, 4307, 4327, 3027, 2251, 4753, 2764, 293, 330, 2766, 3094, 2480, 3041, 2781, 4104, 2853]. Typographical [4601, 4604, 1855]. typographie [2776, 2770]. typographiques [2772]. Typography [3954, 611, 830, 3058, 2725, 1184, 1339, 4117, 1710, 3526, 4052, 2817, 4642, 3333, 2793, 4284, 4285, 3665, 3729, 4172, 2104, 2369, 1701, 3415, 1295, 295]. Typography — [3678]. typography — the [1627]. Typohol [2489].

u [4652]. U. [4739]. U.K [1563]. U.K. [690, 2499]. U.S. [74]. Ubuntu [3590]. UCS [2796]. ugly [4683]. UK [810, 413, 883, 925, 1496, 1124, 4779, 4848, 4055]. UK-based [810]. UK-TUG [4848]. uk.ac.tex [1265]. Ukraine [4826, 4529]. UKTEX [1007, 864, 906, 1292]. UKTUG [3114]. Ulrike [4990, 4964]. ultimate [1041]. UML [2703]. Umlaut}s [296]. Unblocking [181]. Unbound [2152]. Unbound — A [2368]. Understanding [4539]. understands [4438]. Undisciplined [38]. unexpected [4183]. Ungler [4978]. unhappy [2998]. unhboxed [1169]. unicase [4601]. Unicode [1854, 2958, 2791, 2449, 3196, 4902, 4212, 2018, 4300, 3100, 1934, 3042, 3418, 3340, 4757, 4548, 2614, 3529, 4056, 4541, 4569, 4823, 4909, 2791]. Unicode-land [4823]. Unicode-like [4902]. unicodefonttable [4758]. Unicodeland [4058]. unified [4015]. Union [494]. unit [4761, 3269, 29, 30]. units [3595]. unity [541]. Univ. [4287]. Univac [76, 36]. universal [4955]. University [1007, 442, 882, 1496, 1563, 1990, 2301, 2352, 2649, 185, 226, 352, 368, 443, 663, 666, 705, 840, 146, 171, 35, 163, 117, 3733, 2706, 4062, 3537, 269, 416, 9, 3450, 201, 204, 3624]. UNIX [336, 1260, 737, 201, 461, 200, 118, 199, 239, 272, 306, 340, 380, 418, 1803, 646, 732, 815]. UnixTEX [866]. unknown [3702, 3906, 3483, 3553, 3597]. unreasonable [4512]. unusual [3774, 1036]. upcoming [4616, 1868]. Update [635, 684, 84, 1558, 3730, 4213, 529, 1685, 125, 649, 24, 464, 1029, 2615, 1491, 1647, 1986, 2027, 625]. updated [4939, 633]. Updates [1158, 4872, 4396]. Updating [4935, 3703, 3711]. Updike [4365]. updmap [4289]. upload [3254]. Uppercase [84, 38]. upTEX — Unicode [3874]. urtext [4562]. USA [1635]. usability [4816]. Use [3424, 4827, 1329, 1024, 4183, 1481, 490, 4333, 176, 1747, 4023, 3082, 1704, 4246, 2658, 4287, 37, 3663, 4475, 2319, 1045]. used [3053]. useful [1648, 1947, 4136, 768]. useless [384]. User [226, 4649, 951, 3707, 925, 303, 2934, 1611, 701, 745, 785, 1628, 3229, 1363, 1650, 3419, 3161, 883, 3402, 2093, 4301, 2536, 1937, 1942, 3876, 48, 190]. User-defined [4649]. User-friendly [3707, 1937, 1942]. Users [1886, 2012, 2435, 2518, 1123, 829, 554, 2121, 4836, 4878, 4594, 3452, 762, 308, 4362, 3060, 3250, 275, 3169, 3211, 3275, 3298, 3487, 3635, 3806, 2129, 2377, 690, 572, 1246, 218, 1031, 1074, 640, 165]. Using [3515, 2836, 4824, 2666, 4943, 3451, 4981, 4643, 3251, 4371, 4297, 3360, 1608, 3460, 4608, 4813, 3930, 4866, 3770, 3085, 4754, 1362, 2893, 4402, 311, 4722, 992, 2690, 1366, 3846, 1787, 1476, 3950, 3198, 3449, 2800, 4857, 425, 678, 2166, 2067, 2126, 2167, 2538, 3403, 130, 4140, 2820, 2891, 2653, 3558, 2964, 1404, 1579, 4396, 2691, 4284, 1516, 3921, 573, 1739, 2339, 4503, 4238, 4335, 3760, 3127, 2366, 4181, 1417, 3264, 1360, 4787, 1113, 1549, 3478, 2697, 1474, 661, 1025, 1488, 639, 3140, 3096, 1843, 2534, 2539, 2982, 996, 722, 2924, 870, 1908]. Using [1969, 2542, 1607, 2313]. UTF [3200, 4657, 4695, 2790, 3327]. UTF-8 [4657, 4695, 2790, 3327]. utilities [3707, 1812]. Utility [45, 2894, 3948, 2506]. Utrecht [1275].

V [3705]. v0.95 [647]. V2 [2089]. V3 [1710]. Vadis [1848, 4616]. Valerie [4146]. validating [2452]. Valley [2871, 2943, 3433, 3536]. value [4251, 4253, 4821, 3376, 3486]. Vancouver [2308]. Varenna [312]. Variable [4293, 4754, 2767, 2897]. variable-sized [2767]. variables [4908, 4333]. Varian [109, 242, 584]. variants [4137, 4643]. variations [965, 3980, 768]. Various [477, 193, 264, 300, 335, 373, 411, 671, 727, 772, 579, 2458, 492]. varying [2224]. VAX/TEX [34, 120]. VAX/UNIX [200]. VAX/VMS [122, 734, 109, 242, 78, 862, 275, 1210, 274, 307, 341, 420, 462, 1261, 867, 121, 149, 172, 202, 240, 381, 77]. VDU [470]. vector [3294, 3527, 4338]. Vega [4363]. Ventrella [4869]. Ventura [1628]. Verbatim [1680, 3038, 1081, 3673, 1087]. verbatim* [1087]. Verbatims [3801]. verification [3067]. Verna [3887]. Versatec [201]. versatile [1615]. versatility [998]. Version [428, 64, 2396, 2572, 2738, 4856, 424, 764, 1035, 3408, 4206, 1368, 174, 2532, 2790, 2467, 2046, 3874, 4822, 1876]. versions [1026, 4089, 736, 3560, 939, 3290, 2321]. versus [4606, 1365, 4301, 4684]. versus/and [1365]. Vertical [877, 4986, 3144, 1102]. Vervliet [4674, 4659]. Very [2617, 2066, 2773, 5, 2320]. Veytsman [3346, 4807]. vf [1387]. VFComb [1932]. via [950, 338, 4212, 878, 1083, 4952, 898, 4932, 2170, 2162, 3879]. Vice [2249]. Vice-President [2249]. Victor [1332, 1446]. videos [4117]. videotape [437]. Videotapes [750]. viendrait [2769]. viendrait-elle [2769]. Vienna [1182]. Vietnamese [915, 2696, 3248]. View [1513, 1472, 1489, 1852, 4603, 1646, 1304]. viewer [3919]. viewers [4461, 3920]. Viewing [2330]. viewpoint [248, 247]. views [2205]. Vincent [4145]. VINITI [1992]. virgule [2769]. Virtual [1619, 2047, 1932, 1404, 851, 3531, 1549, 1027, 2887]. virtuous [2047]. visible [4019]. visit [1093]. Visiting [3943]. visits [3946]. Vistas [3333]. Visual [717, 3986, 4721, 4947, 1962, 2202, 2203]. visual-based [4721]. VM [2658, 1034, 644, 731, 813, 725, 1073, 1131, 1164]. VM/CMS [1034, 644, 731, 813, 725, 1073, 1131, 1164]. VMS [1440, 122, 850, 959, 1127, 275, 734, 1210, 109, 242, 274, 78, 307, 341, 420, 462, 1261, 862, 867, 121, 149, 172, 202, 240, 381, 77]. Vol [944, 950, 1026, 1038].

Volume [1460, 1713, 3493, 4307, 4327, 4577, 3642, 1110, 688, 3243]. Volunteer [677, 1530]. Volunteers [2044, 2178]. voulu [4145]. vous [4145]. vowel [2252]. VPAT [4726]. vs [4852]. vs. [386, 233, 3947, 3265, 4066]. VT [337]. VTEX [1115]. vulcan [2317]. Vulis [1552]. vuTEX [776].

W [2844, 2815, 1653, 2279, 1560]. W. [4410]. Walczak [3462]. Walden [4805]. Walsh [1706]. Walters [3962]. want [4493, 3104, 3868]. Wanted [606, 320, 747, 1996, 703, 1330, 1077]. ware [550]. WARMreader [2544]. was [869]. Washington [663, 416, 201, 269]. watch [280]. Wawrykiewicz [4360]. way [1487, 1086, 2316, 3342, 3585, 3066, 3160, 3957, 2329]. wayfarer [3160]. ways [3338]. wayward [3160]. Web [2547, 3677, 2453, 2310, 2167, 3983, 2321, 2347, 1964, 2691, 2162, 1756, 2325, 630, 2261, 3515, 3525, 4461, 3707, 4655, 3253, 4607, 500, 4562, 4182, 575, 4337, 4720, 4618, 2932, 3096, 2342, 2510, 4473, 2311, 2654, 3343, 3794, 3361, 3290, 2169, 3135, 468, 527, 143]. Web-based [2453, 2325, 3525, 3290]. Web-pages [2167]. Web/CD [2169]. Webless [1148]. Webnotes [4575]. Week [2104, 1701]. Welcome [2850, 2682]. well [697]. well-documented [697]. WEPT [1701]. Wesley [3218, 4769, 3535]. Wesley — Computers [511]. Wesley — MicroTEX [355, 394, 435]. Wesley — TEX [517]. WeTEX [4403]. Wetmore [3145, 3887, 3960]. Whatever [1877]. Where [1620, 1449, 465, 1105, 2529, 3409, 1524]. Which [3585, 1932, 1947, 2773]. Whidden [1401]. white [578]. who [3142, 3868]. whose [2449]. Wiadomości [4302]. Wide [1756]. widow [4784, 4872]. widows [4784, 4872, 3470, 4439, 4440]. widows-and-orphans [4440]. width [3979, 2897]. widths [213, 2224, 4296]. Wierda [2705, 3106]. Wiggly [426]. wiki [3251]. Wikipedia [3836, 3205]. Wikipublisher [3290]. wildcards [3268]. Wiley [1829]. will [956, 2529]. William [2400]. Wilson [3426, 4660]. winded [1166]. winding [3712]. window [1477]. windowing [1614]. Windows [661, 2613, 2536, 660, 597, 3773, 3057, 2332, 3304, 3831, 3708, 4290]. windows — conclusion [658]. wine [3626]. winner [791]. Winter [157, 104]. wish [1645, 1457]. within [1749, 3367, 3451, 1626]. wizard [3401]. Wizards [1027]. Wood [3639]. WOODMAN [879]. Word [3218, 454, 4311, 2198, 422, 3734, 4722, 1377, 997, 4511, 4724, 1849, 2935, 3331, 3205]. WordPerfect [996, 870]. Words [1927, 2063, 1599, 3194, 1825, 1839, 3179, 1547, 1574, 1643, 1667, 1699, 1774, 1801, 2965, 261]. work [4302, 4537, 2459, 1530, 2503, 648, 4144, 2205, 1706]. Work-around [648]. workable [1014]. work — aesthetically [4620]. workflow [3737, 3881, 3951, 4182, 4505, 4724, 4729]. Working [208, 4693, 3404]. WorkPlace [3218]. works [1357, 2041, 2092, 2128, 2155, 2231, 2295, 2373, 2428, 1360, 4911, 4131]. worksheets [3515, 2961]. Workshop [879, 60, 830, 2547, 59, 4951, 4613, 4282, 4818, 1622, 2746, 609, 72, 3990]. Workshops [1492, 2160]. Workstation [557]. workstations [273]. World [4543, 1756, 3627, 1832, 4467, 3032, 3123, 1751, 4286, 3532, 4530, 4146, 1357, 2316, 3962]. worlde [1006]. Would [3053, 3552, 847]. wrapper [3045]. Write [3930, 1608, 578, 1040, 1173]. Write-white [578]. write18 [3424]. Writer [1628, 735]. Writers [1905]. writes [3089]. Writing [3516, 1866, 3168, 3261, 1479, 3136, 3179, 3341, 3454, 2020, 4308, 1474, 1040, 4182, 1104, 1214, 741, 2769, 3297, 3672, 3401, 2340, 3410, 2731]. Writing — The [4240, 4307, 4327]. written [1107, 1621, 3925]. wrong [1877]. Wuhan [2712, 2806, 2865]. Wujastyk [3669]. WWW [2339]. Wynne [4910]. Wynter [1520]. WYSIWYG [2043, 2610, 2367].

X [1477, 1308, 3735, 2690]. X- [4065]. X-window [1477]. X3J6 [142, 231]. X3V1 [523]. X3V1.8 [328]. Xanthi [2719]. XBibTEX [1510]. Xdvipsk [4291]. XE [2778]. Xerox [863]. XETEX — the [2986]. XHTML [2936]. XI [2021]. Xie [3931]. XII [4344, 2220]. Xindy [3327]. xkeyval — new [2898]. XL^{A}TEX [2773, 3664]. XL^{A}TEX-TEXML-HTML5 [3664]. XLingPaper [4827]. XML [2818, 3024, 2608, 2819, 4932, 4718, 4503, 4620, 2366, 4729, 4530, 3958, 2546]. XML-first [4530]. XML-to-L^{A}TEX [3957]. XML/MathML [2338]. xml2tex [3957]. xmltex [2452]. XMP [4859]. xparse [3485, 3555]. XPC [4620]. xpicture [3799]. XSL [3250]. XSL-FO [3250]. xtemplate [3763]. Xy-pic [2067, 2126, 2066].

Y&Y [2082, 2179, 2214, 2241, 2270, 2304, 2353, 2388, 2112]. Ya [2565]. Yabe [3887]. Yannis [3303]. Yatskovsky [3462]. YAWN — a [3737]. Year [2254, 4678, 4596, 4061, 1122]. Yearbook [3956]. Years [2851, 834, 4302, 3523, 4185, 3581, 3233, 4616, 4298, 4133, 3637, 3760, 2335]. Yes [4287, 4087]. yesterday [2687]. yex [4814]. Yihui [3931]. ymir.claremont.edu [1276]. YUNUS [1246, 1294].

Z [2959]. Zapf [4543, 3850, 3165, 3293, 3030, 4051, 4052, 4049, 2502, 4050, 3747, 3850]. Zapfino [2554, 2727]. Zebrackets [1580, 4185]. zero [3840, 3974]. Zpravodaj} [2902, 3388, 3429, 3498, 3566, 3645, 3001, 3778, 4477, 4962, 4261, 4381, 4624, 4580, 4698, 4792, 4915]. ZzTEX [1528].

References

Anonymous:1980:TP

[1] Anonymous. Title page. TUGboat, 1(1):1, October 1980. CODEN

      ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb01-1/
      titlepage.pdf.

Welland:TB1-1-2

[2] Robert Welland. Editor’s Comments. TUGboat, 1(1):2–3, October 1980. ISSN

      0896-3207. URL https://tug.org/TUGboat/tb01-1/tb01edit.
      pdf.

Palais:TB1-1-3

[3] Richard Palais. Message from the Chairman. TUGboat, 1(1):3–7, October

      1980. ISSN 0896-3207. URL https://tug.org/TUGboat/tb01-1/
      tb01pres.pdf.

Swanson:TB1-1-7

[4] Ellen Swanson. Publishing & TEX. TUGboat, 1(1):7–9, October

1980. ISSN 0896-3207. URL https://tug.org/TUGboat/tb01-1/
tb01swanson.pdf.

Spivak:TB1-1-10

[5] Michael Spivak. AMS-TEX — “A very friendly product”. TUGboat, 1(1):10– 11, October 1980. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb01-1/tb01spivak.pdf.

Morris:TB1-1-12

[6] Robert Morris. Minutes of the first TUG meeting, February 1980. TUGboat, 1(1):12–16, October 1980. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb01-1/tb01morris.pdf.

Zabala:TB1-1-16

[7] Ignacio Zabala and Luis Trabb-Pardo. The status of the Pascal implementation of TEX. TUGboat, 1(1):16–17, October 1980. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb01-1/tb01zabala.pdf.

Fuchs:TB1-1-17

[8] David Fuchs. The format of TEX’s DVI files. TUGboat, 1(1):17–19, October

1980. ISSN 0896-3207. URL https://tug.org/TUGboat/tb01-1/
tb01fuchs.pdf.

Hodge:TB1-1-19

[9] Thea Hodge. University of Minnesota CDC Cyber site report. TUGboat, 1 (1):19–20, October 1980. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb01-1/tb01hodge.pdf.

Beeton:TB1-1-20

[10] Barbara Beeton. Troubles with trace and Other oddities. TUGboat, 1(1):20, October 1980. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb01-1/tb01beeton.pdf.

Anonymous:1980:QA

[11] Anonymous. Questions & answers. TUGboat, 1(1):21, October 1980. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb01-1/
tb01misc.pdf.

Emch:TB1-1-22

[12] Gérard Emch and Arnold Pizer. Letters. TUGboat, 1(1):22–23, October

1980. ISSN 0896-3207. URL https://tug.org/TUGboat/tb01-1/
tb01emch.pdf.

Anonymous:1980:APD

[13] Anonymous. AMS-TEX panel discussion. TUGboat, 1(1):c3, October 1980. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb01-1/tb01ams.pdf.

Anonymous:1980:SC

[14] Anonymous. Steering Committee. TUGboat, 1(1):c2, October 1980. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb01-1/
steering.pdf.

Winograd:TB1-1-Appendix-A

[15] Terry Winograd and Bill Paxton. An indexing facility for TEX. TUGboat, 1(1): Appendix A, October 1980. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb01-1/winograd-paxton.pdf.

Anonymous:TB2-1-2

[16] Anonymous. Addresses of authors. TUGboat, 2(1):2, February 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-1/tb02titlepage.
pdf.

Welland:TB2-1-3

[17] Robert Welland. Site Coordinators. TUGboat, 2(1):3, February 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-1/tb02gendel.
pdf.

Palais:TB2-1-3

[18] Richard Palais. Chairman’s report. TUGboat, 2(1):3–5, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02gendel.pdf.

Morris:TB2-1-5

[19] Robert Morris. Report on the January 1981 TUG Steering Committee meeting. TUGboat, 2(1):5–6, February 1981. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb02-1/tb02gendel.pdf.

Whidden:TB2-1-6

[20] Samuel B. Whidden. 1980 TUG Treasurer’s report. TUGboat, 2(1):6, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02gendel.pdf.

Morris:TB2-1-6

[21] Robert Morris. Informal TUG session. TUGboat, 2(1):6–7, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02gendel.pdf.

Morris:TB2-1-7

[22] Robert Morris. A position on TEX maintenance. TUGboat, 2(1):7–8, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02gendel.pdf.

Whidden:TB2-1-9

[23] Samuel B. Whidden. TEX support. TUGboat, 2(1):9–10, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02treas.pdf.

Kim:TB2-1-10

[24] Scott Kim. Update on Pascal . TUGboat, 2(1):10, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02kim.pdf.

Fuchs:TB2-1-11

[25] David Fuchs. Erratum: The format of TEX’s DVI files. TUGboat, 2(1):11, February 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-1/tb02fuchszab.pdf.

Zabala:TB2-1-11

[26] Ignacio Zabala. TEX-Pascal and Pascal compilers (a status report). TUGboat, 2(1):11–12, February 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-1/tb02fuchszab.pdf.

Fuchs:TB2-1-12

[27] David Fuchs. TEX Font Metric files. TUGboat, 2(1):12–16, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02fuchstfm.pdf.

Sherrod:TB2-1-17

[28] Phil Sherrod and Alan Wright. TEX support programs. TUGboat, 2(1):17–19, February 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-1/tb02sherrod.pdf.

Lawson:TB2-1-20

[29] C. L. Lawson, I. Zabala, and M. Díaz. Brief functional characterization of the procedures in the TEX/Pascal compilation unit, SYSDEP. TUGboat, 2 (1):20–31, February 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-1/tb02lawbrief.pdf.

Lawson:TB2-1-32

[30] C. L. Lawson, I. Zabala, and M. Díaz. Detailed specifications of procedures in the TEX/Pascal compilation unit, SYSDEP. TUGboat, 2(1):32–47, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02lawdetail.pdf.

Berns:TB2-1-48

[31] Eagle Berns. Status of TEX on the Stanford 370/3033 systems. TUGboat, 2(1):48, February 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-1/tb02site.pdf.

Doherty:TB2-1-48

[32] Barry Doherty. Output devices: A new column. TUGboat, 2(1):48, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02site.pdf.

Doherty:TB2-1-48-2

[33] Barry Doherty and Barbara Beeton. AMS site report. TUGboat, 2(1):48–49, February 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-1/tb02site.pdf.

Incerpi:TB2-1-49

[34] Janet Incerpi. The status of VAX/TEX at Brown. TUGboat, 2(1):49–50, Febru-

ary 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-
1/tb02site.pdf.

Hodge:TB2-1-51

[35] T. D. Hodge. Report from the North Star, or, TEX at the University of Minnesota. TUGboat, 2(1):51, February 1981. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb02-1/tb02site.pdf.

Stromquist:TB2-1-51

[36] Ralph Stromquist. TEX is available for Univac 1100 systems. TUGboat, 2(1):51, February 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-1/tb02site.pdf.

Totland:TB2-1-51

[37] Helge Totland. Report on the use of TEX at Computas A/S, Norway; Nord 100 computer: 16 bit “big mini”. TUGboat, 2(1):51–52, February 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-1/tb02site.
pdf.

Beeton:TB2-1-53

[38] Barbara Beeton. Disappearing digits; Undisciplined uppercase. TUGboat, 2(1):53, February 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-1/tb02beet.pdf.

Beeton:TB2-1-53-2

[39] Barbara Beeton. How to prepare a file for publication in TUGboat. TUGboat, 2(1):53–54, February 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-1/tb02beet.pdf.

Diaz:TB2-1-55

[40] Max Díaz. TEX macro package. TUGboat, 2(1):55, February 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-1/tb02diaz.
pdf.

Keller:TB2-1-56

[41] Arthur M. Keller. Anatomy of a TEX macro package. TUGboat, 2(1):56–86, February 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-1/tb02kell.pdf.

Price:TB2-1-87

[42] Lynne A. Price and Patrick Milligan. NOFILL program with Pascal source. TUGboat, 2(1):87–97, February 1981. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb02-1/tb02pricenofill.pdf.

Price:TB2-1-98

[43] Lynne A. Price. List macros. TUGboat, 2(1):98–110, February 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-1/tb02pricelist.
pdf.

Price:TB2-1-111

[44] Lynne A. Price. Table of contents macros. TUGboat, 2(1):111–118, February

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02pricetoc.pdf.

Milligan:TB2-1-119

[45] Patrick Milligan and Lynne A. Price. Utility macros. TUGboat, 2(1):119–121, February 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-1/tb02mill.pdf.

Price:TB2-1-122

[46] Lynne A. Price. Hebrew letter (with source). TUGboat, 2(1):122–124, Febru-

ary 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-
1/tb02pricehebrew.pdf.

Price:TB2-1-125

[47] Lynne A. Price. Two slides. TUGboat, 2(1):125–126, February 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-1/tb02priceslides.
pdf.

Eck:TB2-1-127

[48] David Eck. Report from an early AMS-TEX user. TUGboat, 2(1):127, Febru-

ary 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-
1/tb02letters.pdf.

Hirst:TB2-1-127

[49] Graeme Hirst. Letters. TUGboat, 2(1):127, February 1981. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb02-1/tb02letters.pdf.

Anonymous:TB2-1-128

[50] Anonymous. Appendix A. Output samples from a paper by David Eck. TUGboat, 2(1):128–135, February 1981. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb02-1/tb02eck.pdf.

Anonymous:1981:TE

[51] Anonymous. TEX errata. TUGboat, 2(1):991, February 1981. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-1/
tb02workshop.pdf.

Anonymous:TB2-2-2

[52] Anonymous. Addresses of officers, authors and others. TUGboat, 2(2):2, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03titlepage.pdf.

Anonymous:TB2-2-3

[53] Anonymous. Official announcements. TUGboat, 2(2):3, July 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-2/tb03gendel.
pdf.

Welland:TB2-2-3

[54] Robert Welland. Editor’s remarks. TUGboat, 2(2):3, July 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-2/tb03gendel.
pdf.

Spivak:TB2-2-3

[55] Michael Spivak. Chairman’s report. TUGboat, 2(2):3–4, July 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-2/tb03gendel.
pdf.

Morris:TB2-2-4

[56] Robert Morris. Report on the May 1981 TUG Steering Committee meeting. TUGboat, 2(2):4–5, July 1981. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb02-2/tb03gendel.pdf.

Whidden:TB2-2-5

[57] Samuel B. Whidden. TUG Treasurer’s report. TUGboat, 2(2):5, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03gendel.pdf.

Morris:TB2-2-5

[58] Robert Morris. Proposal for institutional support of TUG. TUGboat, 2(2):5–6,

July 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-
2/tb03gendel.pdf.

Doherty:TB2-2-6

[59] Barry Doherty. Report: TEX Implementors’ Workshop, Stanford, 14–15 May 1981. TUGboat, 2(2):6–7, July 1981. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb02-2/tb03gendel.pdf.

Anonymous:TB2-2-8

[60] Anonymous. Workshop attendees. TUGboat, 2(2):8, July 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-2/tb03gendel.
pdf.

Pierce:TB2-2-8

[61] Tom Pierce. Preliminary announcement: TUG meeting, Cincinnati, January 1982. TUGboat, 2(2):8, July 1981. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb02-2/tb03gendel.pdf.

Milligan:TB2-2-9

[62] Patrick Milligan. Ask not what TUG can do for you, ask what you can do for TUG! TUGboat, 2(2):9–10, July 1981. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb02-2/tb03gendel.pdf.

Milligan:TB2-2-10

[63] Patrick Milligan. A proposal for a machine independent tape interchange standard. TUGboat, 2(2):10–12, July 1981. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb02-2/tb03gendel.pdf.

Fuchs:TB2-2-12

[64] David Fuchs. The format of TEX’s DVI files, Version I. TUGboat, 2(2):12–16,

July 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-
2/tb03fuchsdvi.pdf.

Zabala:TB2-2-16

[65] Ignacio Zabala. Some feedback from PTEX installations. TUGboat, 2(2): 16–19, July 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-2/tb03zab.pdf.

Liang:TB2-2-19

[66] Frank M. Liang. TEX and hyphenation. TUGboat, 2(2):19–20, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03liang.pdf.

Harris:TB2-2-21

[67] Kent S. Harris and Robert M. McClure. TEX on small machines. TUGboat, 2(2):21–24, July 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-2/tb03harris.pdf.

Anonymous:TB2-2-25

[68] Anonymous. Summary of computing equipment and output devices. TUGboat, 2(2):25, July 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-2/tb03fuchsout.pdf.

Fuchs:TB2-2-25

[69] David Fuchs. Output device news flash (APS-5 and Linotron 202). TUGboat, 2 (2):25, July 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-2/tb03fuchsout.pdf.

Trabb-Pardo:TB2-2-26

[70] Luis Trabb-Pardo. Imagen (Canon LBP-10). TUGboat, 2(2):26–27, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03imagen.pdf.

Hodge:TB2-2-28

[71] Thea Hodge and Michael Frisch. TEX under the North Star. TUGboat, 2(2):28,

July 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-
2/tb03site.pdf.

Sherrod:TB2-2-28

[72] Phil Sherrod. DECSystem-10/20 Implementation Workshop announcement. TUGboat, 2(2):28–29, July 1981. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb02-2/tb03site.pdf.

Beeton:TB2-2-29

[73] Barbara Beeton. AMS site report. TUGboat, 2(2):29, July 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-2/tb03site.pdf.

Milligan:TB2-2-29

[74] Patrick Milligan. TEX at the 1981 Spring DECUS U.S. Symposium. TUGboat, 2(2):29, July 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-2/tb03site.pdf.

Schwab:TB2-2-29

[75] Rachel Schwab. TEX at NIH. TUGboat, 2(2):29–30, July 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-2/tb03site.pdf.

Kelly:TB2-2-30

[76] Bill Kelly. An implementation report for the Univac 1100. TUGboat, 2(2): 30–33, July 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-2/tb03site.pdf.

Nichols:TB2-2-33

[77] Monte Nichols. Availability of Oregon Software implementation of TEX for the VAX/VMS. TUGboat, 2(2):33–34, July 1981. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb02-2/tb03site.pdf.

Smith:TB2-2-34

[78] Barry Smith. TEX for VAX/VMS. TUGboat, 2(2):34, July 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-2/tb03site.pdf.

Doherty:TB2-2-34

[79] Barry C. W. Doherty. TUG Font Committee. TUGboat, 2(2):34–35, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03fonts.pdf.

Hickey:TB2-2-35

[80] Thomas B. Hickey. The status of at OCLC. TUGboat, 2(2): 35–38, July 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-2/tb03fonts.pdf.

LeVeque:TB2-2-39

[81] William J. LeVeque. Font development at the AMS. TUGboat, 2(2):39–40,

July 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-
2/tb03fonts.pdf.

Whitney:TB2-2-40

[82] Ronald Whitney. TPHON. TUGboat, 2(2):40, July 1981. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb02-2/tb03fonts.pdf.

Doherty:TB2-2-40

[83] Barry Doherty and Ronald Whitney. Proofmode and magnification. TUGboat, 2(2):40–42, July 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-2/tb03fonts.pdf.

Beeton:TB2-2-42

[84] Barbara Beeton. Uppercase update; Fickle fonts. TUGboat, 2(2):42–43, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03beet.pdf.

Anonymous:TB2-2-43

[85] Anonymous. Macros on microfiche. TUGboat, 2(2):43, July 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-2/tb03pricemac.
pdf.

Price:TB2-2-43

[86] Lynne Price. Greetings from the Editor. TUGboat, 2(2):43, July 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-2/tb03mill.
pdf.

Milligan:TB2-2-43

[87] Patrick Milligan. Erratum: NOFILL program. TUGboat, 2(2):43–44, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03mill.pdf.

Milligan:TB2-2-44

[88] Patrick Milligan. TEX macros for automatic font code allocation. TUGboat, 2(2):44–45, July 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-2/tb03mill.pdf.

McKay:TB2-2-46

[89] Brendan McKay. A macro menagerie. TUGboat, 2(2):46–49, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03mckay.pdf.

Spivak:TB2-2-50

[90] Michael Spivak. Macro madness. TUGboat, 2(2):50–54, July 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-2/tb03spivak.
pdf.

Price:TB2-2-54

[91] Lynne Price. Greetings from the Editor. TUGboat, 2(2):54, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03priceprob.pdf.

Anonymous:TB2-2-54

[92] Anonymous. Problems from the TEXarcana course. TUGboat, 2(2):54–56, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03priceprob.pdf.

Spivak:TB2-2-57

[93] Michael Spivak. Input-dependent macro redefinition. TUGboat, 2(2):57, July

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-2/
tb03misc.pdf.

Stovall:TB2-2-57

[94] Johnny Stovall. Balancing columns of text and translation. TUGboat, 2(2):57,

July 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-
2/tb03misc.pdf.

Price:TB2-2-57

[95] Lynne Price. Letters. TUGboat, 2(2):57–58, July 1981. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb02-2/tb03misc.pdf.

Price:TB2-2-58

[96] Lynne Price. Dreamboat. TUGboat, 2(2):58, July 1981. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb02-2/tb03misc.pdf.

Diaz:TB2-2-Appendix-A

[97] Max Díaz. Fácil TEX. TUGboat, 2(2):Appendix A, July 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-2.

Anonymous:1981:TP

[98] Anonymous. Title page. TUGboat, 2(3):1, November 1981. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04titlepage.pdf.

Anonymous:TB2-3-2

[99] Anonymous. Addresses of officers, authors and others. TUGboat, 2(3):2, November 1981. ISSN 0896-3207.

Anonymous:TB2-3-3

[100] Anonymous. Official announcements. TUGboat, 2(3):3, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04gendel.pdf.

Spivak:TB2-3-3

[101] Michael Spivak. Message from the Chairman. TUGboat, 2(3):3–4, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04gendel.pdf.

Whidden:TB2-3-4

[102] Samuel B. Whidden. TUG Treasurer’s report. TUGboat, 2(3):4, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04gendel.pdf.

Knuth:TB2-3-5

[103] Donald Knuth. The current state of things. TUGboat, 2(3):5–6, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04gendel.pdf.

Pierce:TB2-3-7

[104] Tom Pierce. TEX Users Group Winter 1982 meeting, January 11–12, 1982, Cincinnati, Ohio. TUGboat, 2(3):7, November 1981. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb02-3/tb04gendel.pdf.

Samuel:TB2-3-7

[105] Arthur Samuel. Pascal-coded TEX errata. TUGboat, 2(3):7–8, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04samuel.pdf.

Fuchs:TB2-3-8

[106] David Fuchs. The format of PXL files. TUGboat, 2(3):8–12, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04fuchspxl.pdf.

Sauter:TB2-3-13

[107] John Sauter. Sample of output from an IDS-640. TUGboat, 2(3):13, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04sauter.pdf.

Thedford:TB2-3-14

[108] Rilla Thedford. Output device index. TUGboat, 2(3):14, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04output.pdf.

Mooney:TB2-3-14

[109] Jim Mooney. A Varian output driver in VAX/VMS Fortran. TUGboat, 2 (3):14–15, November 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-3/tb04output.pdf.

Murphy:TB2-3-15

[110] Timothy Murphy. Diabolic TEX. TUGboat, 2(3):15–21, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04murdiablo.pdf.

Fuchs:TB2-3-21

[111] David Fuchs. News from the home front (Stanford). TUGboat, 2(3):21–22, November 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-3/tb04site.pdf.

Frisch:TB2-3-22

[112] Michael Frisch. TEX under the North Star. TUGboat, 2(3):22–23, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04site.pdf.

Beeton:TB2-3-23

[113] Barbara Beeton. A TUGboat tour: Excerpts from the TEXnician’s log. TUGboat, 2(3):23–25, November 1981. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb02-3/tb04site.pdf.

Carnes:TB2-3-25

[114] Lance Carnes. TEX for the HP3000. TUGboat, 2(3):25–26, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04site.pdf.

Plass:TB2-3-26

[115] Susan Plass. TEX for the IBM 370. TUGboat, 2(3):26–27, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04site.pdf.

Rosenschein:TB2-3-27

[116] Jeffrey S. Rosenschein. TEX in Israel. TUGboat, 2(3):27–28, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04site.pdf.

Rodgers:TB2-3-28

[117] David Rodgers. TEX at the University of Michigan, summary of progress. TUGboat, 2(3):28–29, November 1981. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb02-3/tb04site.pdf.

Morris:TB2-3-29

[118] Robert Morris. VAX on UNIX. TUGboat, 2(3):29, November 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-3/tb04site.
pdf.

Jackson:TB2-3-29

[119] Calvin Jackson. TEX at CalTech. TUGboat, 2(3):29–32, November 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-3/tb04site.
pdf.

Incerpi:TB2-3-32

[120] Janet Incerpi. The status of VAX/TEX at Brown. TUGboat, 2(3):32, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04site.pdf.

Nichols:TB2-3-32

[121] Monte C. Nichols. VAX/VMS site report. TUGboat, 2(3):32, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04site.pdf.

Blair:TB2-3-32

[122] John Blair. Enhancements of VAX/VMS TEX at Calma. TUGboat, 2(3):32–34, November 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-3/tb04site.pdf.

Sauter:TB2-3-34

[123] John Sauter. “Poor man’s” TEX. TUGboat, 2(3):34–35, November 1981. ISSN

0896-3207. URL https://tug.org/TUGboat/tb02-3/tb04site.
pdf.

Beeton:TB2-3-35

[124] Barbara Beeton. Don’t just \let TEX hang, \raise or \lower it. TUGboat, 2(3):35, November 1981. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb02-3/tb04beet.pdf.

Whitney:TB2-3-35

[125] Ronald Whitney. Font update. TUGboat, 2(3):35, November 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-3/tb04whitney.
pdf.

Price:TB2-3-36

[126] Lynne Price. Editor’s introduction. TUGboat, 2(3):36, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04pricemac.pdf.

Anonymous:TB2-3-36

[127] Anonymous. TUGboat macro index. TUGboat, 2(3):36–37, November

1981. ISSN 0896-3207. URL https://tug.org/TUGboat/tb02-3/
tb04pricemac.pdf.

Murphy:TB2-3-37

[128] Timothy Murphy. Bubbles: A TEXtension in search of a TEXpert. TUGboat, 2(3):37–38, November 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-3/tb04murbub.pdf.

Plass:TB2-3-39

[129] Michael F. Plass. Charting your grammar with TEX. TUGboat, 2(3):39–56, November 1981. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb02-3/tb04plass.pdf.

Nichols:TB2-3-57

[130] Monte Nichols and Barbara Beeton. Chemical notation using TEX. TUGboat, 2(3):57–58, November 1981. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb02-3/tb04nich.pdf.

Price:TB2-3-58

[131] Lynne Price. Problems. TUGboat, 2(3):58–61, November 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-3/tb04priceprob.
pdf.

Anonymous:TB2-3-61

[132] Anonymous. Problems from the TEXarcana class: Answers, and another problem. TUGboat, 2(3):61–65, November 1981. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb02-3/tb04priceprob.pdf.

Cole:TB2-3-66

[133] J. M. Cole. Letters et alia. TUGboat, 2(3):66, November 1981. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb02-3/tb04cole.pdf.

Sannella:TB2-3-66

[134] Michael Sannella. Formatting a book with TEX: Experiences and observations. TUGboat, 2(3):66–74, November 1981. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb02-3/tb04sann.pdf.

Anonymous:TB2-3-xx

[135] Anonymous. Instructions for submitting TUGboat articles on magnetic tape. TUGboat, 2(3):??, November 1981. ISSN 0896-3207.

Anonymous:TB2-3-xx-2

[136] Anonymous. Membership application and order form. TUGboat, 2(3):??, November 1981. ISSN 0896-3207.

Anonymous:TB3-1-2

[137] Anonymous. Addresses of officers, authors and others. TUGboat, 3(1):2, March 1982. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb03-1/tb05complete.pdf.

Anonymous:TB3-1-3

[138] Anonymous. Official announcements. TUGboat, 3(1):3, March 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-1/tb05gendel.
pdf.

Beeton:TB3-1-3

[139] Barbara Beeton. Statement of editorial policy. TUGboat, 3(1):3–4, March

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-1/
tb05gendel.pdf.

Morris:TB3-1-4

[140] Robert Morris. Report on the January 1981 TUG Steering Committee meeting. TUGboat, 3(1):4–5, March 1982. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb03-1/tb05gendel.pdf.

McGaffey:TB3-1-7

[141] Robert McGaffey and Keith Penny. Open letter to TUG. TUGboat, 3(1):7– 9, March 1982. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb03-1/tb05mcgaffey.pdf.

Price:TB3-1-10

[142] Lynne A. Price. Report on the ANSI X3J6 meeting. TUGboat, 3(1):10, March

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-1/
tb05pricex3j6.pdf.

Knuth:TB3-1-10

[143] Donald Knuth. Fixed-point glue setting — an example of WEB. TUGboat, 3(1):10–27, March 1982. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb03-1/tb05knuth.pdf.

Frisch:TB3-1-28

[144] Michael J. Frisch. CDC TEX. TUGboat, 3(1):28, March 1982. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb03-1/tb05site.pdf.

Carnes:TB3-1-28

[145] Lance Carnes. HP3000 site report. TUGboat, 3(1):28–29, March 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-1/tb05site.
pdf.

Canzii:TB3-1-29

[146] G. Canzii, D. Lucarella, and A. Pilenga. TEX at the University of Milan. TUGboat, 3(1):29, March 1982. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb03-1/tb05site.pdf.

Aiello:TB3-1-30

[147] L. Aiello and S. Pavan. TEX news from Pisa. TUGboat, 3(1):30–31, March

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-1/
tb05site.pdf.

Carnes:TB3-1-32

[148] Lance Carnes. Editor’s introduction. TUGboat, 3(1):32, March 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-1/tb05site.
pdf.

Nichols:TB3-1-32

[149] Monte C. Nichols. VAX/VMS site report. TUGboat, 3(1):32, March

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-1/
tb05site.pdf.

Beeton:TB3-1-33

[150] Barbara Beeton. Another hangup. TUGboat, 3(1):33, March 1982. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb03-1/tb05beet.pdf.

Price:TB3-1-33

[151] Lynne Price. Editor’s introduction. TUGboat, 3(1):33, March 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-1/tb05pricemac.
pdf.

Anonymous:TB3-1-33

[152] Anonymous. TUGboat macro index. TUGboat, 3(1):33–35, March

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-1/
tb05pricemac.pdf.

Beeman:TB3-1-35

[153] Roger L. Beeman. Display of a font in table form. TUGboat, 3(1):35–38, March 1982. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb03-1/tb05beemanfont.pdf.

Beeman:TB3-1-39

[154] Roger L. Beeman. Seating charts. TUGboat, 3(1):39–42, March

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-1/
tb05beemanseat.pdf.

Anonymous:TB3-1-43

[155] Anonymous. Problems from the TEXarcana class: Answer to the challenge problem. TUGboat, 3(1):43–44, March 1982. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb03-1/tb05prob.pdf.

Whidden:TB3-1-xx

[156] Samuel B. Whidden. TUG financial reports. TUGboat, 3(1):45–47, March

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-1/
tb05treas.pdf.

Anonymous:TB3-1-xx

[157] Anonymous. TUG Winter meeting, January 11–12, 1982, Cincinnati, Ohio. TUGboat, 3(1):??, March 1982. ISSN 0896-3207.

Anonymous:TB3-1-xx-2

[158] Anonymous. Instructions for submitting TUGboat articles on magnetic tape. TUGboat, 3(1):??, March 1982. ISSN 0896-3207.

Anonymous:TB3-1-xx-3

[159] Anonymous. Membership application and order form. TUGboat, 3(1):??, March 1982. ISSN 0896-3207.

Anonymous:TB3-2-2

[160] Anonymous. Addresses of officers, authors and others. TUGboat, 3(2):2, Octo-

ber 1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-
2/tb06complete.pdf.

Anonymous:TB3-2-3

[161] Anonymous. Official announcements. TUGboat, 3(2):3, October 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-2/tb06announce.
pdf.

Anonymous:TB3-2-4

[162] Anonymous. The TEX logo: An important note. TUGboat, 3(2):4, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06gendel.pdf.

Plass:TB3-2-4

[163] Susan Plass. Report on business meetings, TUG Summer meeting, Stanford University, July 25–27, 1982. TUGboat, 3(2):4–5, October 1982. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb03-2/tb06gendel.
pdf.

MacKay:TB3-2-7

[164] Pierre A. MacKay. An informal interchange format for TEX files. TUGboat, 3(2):7–8, October 1982. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb03-2/tb06mackay.pdf.

Whitney:TB3-2-9

[165] Ron Whitney. Introduction to TEX and TUG for new users. TUGboat, 3(2):9– 12, October 1982. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb03-2/tb06whitney.pdf.

Anonymous:TB3-2-113

[166] Anonymous. TEX82 memory structure. TUGboat, 3(2):13, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06software.pdf.

Fuchs:TB3-2-14

[167] David Fuchs. The format of TEX’s DVI files. TUGboat, 3(2):14–19, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06software.pdf.

Thedford:TB3-2-20

[168] Rilla J. Thedford. Output devices. TUGboat, 3(2):20, October 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-2/tb06site.
pdf.

Fuchs:TB3-2-20

[169] David Fuchs. News from Stanford. TUGboat, 3(2):20–21, October 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-2/tb06site.
pdf.

Plass:TB3-2-21

[170] Susan Plass. Fixes to known bugs in TEX370. TUGboat, 3(2):21–22, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06site.pdf.

Grosso:TB3-2-22

[171] Paul Grosso. TEX installation at the University of Michigan. TUGboat, 3 (2):22–23, October 1982. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb03-2/tb06site.pdf.

Nichols:TB3-2-23

[172] Monte C. Nichols and David Kellerman. VAX/VMS site report. TUGboat, 3(2):23, October 1982. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb03-2/tb06site.pdf.

Carnes:TB3-2-24

[173] Lance Carnes. Editor’s introduction. TUGboat, 3(2):24, October 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-2/tb06carnes.
pdf.

Mong:TB3-2-25

[174] Sao Khai Mong. A Fortran version of . TUGboat, 3(2):25, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06mbj.pdf.

Beeton:TB3-2-25

[175] Barbara Beeton. Charting the generation gulf. TUGboat, 3(2):25–26, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06mbj.pdf.

Jackson:TB3-2-26

[176] Calvin Jackson. Font codes in popular use. TUGboat, 3(2):26, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06mbj.pdf.

Price:TB3-2-27

[177] Lynne Price. Editor’s introduction. TUGboat, 3(2):27, October 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-2/tb06pricemac.
pdf.

Anonymous:TB3-2-27

[178] Anonymous. TUGboat macro index. TUGboat, 3(2):27–28, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06pricemac.pdf.

Beeton:TB3-2-28

[179] Barbara Beeton. Multi-column output format. TUGboat, 3(2):28–33, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06beet.pdf.

Lamport:TB3-2-34

[180] Leslie Lamport. Some TEX programming hacks. TUGboat, 3(2):34–36, Octo-

ber 1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-
2/tb06lamport.pdf.

Beeton:TB3-2-36

[181] Barbara Beeton. Unblocking an AMS-TEX tape. TUGboat, 3(2):36–37, Octo-

ber 1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-
2/tb06beetape.pdf.

Anonymous:TB3-2-38

[182] Anonymous. Hanging punctuation. TUGboat, 3(2):38, October 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-2/tb06prob.
pdf.

Blanford:TB3-2-38

[183] Mark Blanford. Paragraphs in tables. TUGboat, 3(2):38, October 1982. ISSN

0896-3207. URL https://tug.org/TUGboat/tb03-2/tb06prob.
pdf.

Whidden:TB3-2-xx

[184] Samuel B. Whidden. TUG financial reports. TUGboat, 3(2):39–41, October

1982. ISSN 0896-3207. URL https://tug.org/TUGboat/tb03-2/
tb06treas.pdf.

Anonymous:TB3-2-xx

[185] Anonymous. TUG Summer meeting and TEX82 short course, Stanford University, July 25–30, 1982. TUGboat, 3(2):??, October 1982. ISSN 0896-3207.

Anonymous:TB3-2-xx-2

[186] Anonymous. Membership application and order form. TUGboat, 3(2):??, October 1982. ISSN 0896-3207.

Anonymous:TB4-1-2

[187] Anonymous. Addresses of officers, authors and others. TUGboat, 4(1):2, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07complete.pdf.

Anonymous:TB4-1-3

[188] Anonymous. Official announcements. TUGboat, 4(1):3, April 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-1/tb07announce.
pdf.

Anonymous:TB4-1-4

[189] Anonymous. Library subscriptions — What are they? TUGboat, 4(1):4, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07gendel.pdf.

Spivak:TB4-1-4

[190] Michael Spivak. Users’ Course in AMS-TEX. TUGboat, 4(1):4, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07gendel.pdf.

Fuchs:TB4-1-6

[191] David Fuchs. TEXhax summary. TUGboat, 4(1):6–9, April 1983. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb04-1/tb07fuchs.pdf.

Anonymous:TB4-1-10

[192] Anonymous. Chart: Output devices and computers. TUGboat, 4(1):10, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07output.pdf.

Anonymous:TB4-1-10-2

[193] Anonymous. Index to sample output from various devices. TUGboat, 4(1): 10, April 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-1/tb07output.pdf.

Beebe:TB4-1-11

[194] Nelson H. F. Beebe. Low-cost downloadable font devices. TUGboat, 4(1):11– 12, April 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-1/tb07output.pdf.

Ion:TB4-1-12

[195] Patrick Ion, Bill Hall, and Rilla J. Thedford. TEX on the OSP130. TUGboat, 4(1):12, April 1983. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb04-1/tb07output.pdf.

Anonymous:TB4-1-13

[196] Anonymous. Sample output from Florida Data OSP 130. TUGboat, 4(1): 13, April 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-1/tb07output.pdf.

Fuchs:TB4-1-14

[197] David Fuchs. News from all over. TUGboat, 4(1):14–15, April 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-1/tb07site.
pdf.

Bunner:TB4-1-16

[198] Irene J. Bunner and John D. Johnson. TEX on the HP-1000. TUGboat, 4(1): 16, April 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-1/tb07site.pdf.

Furuta:TB4-1-17

[199] Richard Furuta and Pierre MacKay. Unix TEX site report. TUGboat, 4(1):17– 18, April 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-1/tb07site.pdf.

Curtis:TB4-1-18

[200] Pavel Curtis and Howard Trickey. Porting TEX to VAX/UNIX. TUGboat, 4(1):18–20, April 1983. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb04-1/tb07site.pdf.

MacKay:TB4-1-21

[201] Pierre MacKay and Richard Furuta. TEX at the University of Washington: Tops-20, Unix, Versatec, and the Monolithic. TUGboat, 4(1):21–22, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07site.pdf.

Nichols:TB4-1-22

[202] Monte C. Nichols. VAX/VMS site report. TUGboat, 4(1):22, April 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-1/tb07site.
pdf.

Krapp:TB4-1-23

[203] David Krapp. TEX at Calma R & D. TUGboat, 4(1):23, April 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-1/tb07site.
pdf.

Naugle:TB4-1-23

[204] Norman Naugle and Bart Childs. TEX at Texas A & M University. TUGboat, 4(1):23, April 1983. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb04-1/tb07site.pdf.

Carnes:TB4-1-24

[205] Lance Carnes. Editor’s introduction. TUGboat, 4(1):24, April 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-1/tb07carnes.
pdf.

Guthery:TB4-1-25

[206] Scott Guthery. Pictures are just big letters. TUGboat, 4(1):25, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07guth.pdf.

Tobin:TB4-1-26

[207] Georgia K. M. Tobin. Computer calligraphy. TUGboat, 4(1):26–32, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07tobin.pdf.

Anonymous:TB4-1-33

[208] Anonymous. Announcement: Fifth ATypI Working Seminar, The computer and the hand in type design: The aesthetics and technology of digital letterforms. TUGboat, 4(1):33, April 1983. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb04-1/tb07atypi.pdf.

Anonymous:TB4-1-33-2

[209] Anonymous. TUGboat macro index. TUGboat, 4(1):33–35, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07macros.pdf.

Beeton:TB4-1-35

[210] Barbara Beeton. How to build a \strut. TUGboat, 4(1):35–36, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07beet.pdf.

Beeton:TB4-1-36

[211] Barbara Beeton. Determining hashtable size and other quantities. TUGboat, 4(1):36–37, April 1983. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb04-1/tb07beet.pdf.

Mohr:TB4-1-37

[212] August Mohr. Some layout macros. TUGboat, 4(1):37–38, April 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-1/tb07mm.
pdf.

McClure:TB4-1-38

[213] Robert M. McClure. Testing the widths of a font. TUGboat, 4(1):38, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07mm.pdf.

Anonymous:TB4-1-39

[214] Anonymous. Hanging punctuation. TUGboat, 4(1):39, April 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-1/tb07prob.
pdf.

Anonymous:TB4-1-40

[215] Anonymous. Announcement: Manipulation de Documents — Journées Francophones. TUGboat, 4(1):40, April 1983. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb04-1/tb07lett.pdf.

Code:TB4-1-40

[216] Maria Code. How to obtain TEX82 on tape. TUGboat, 4(1):40, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07lett.pdf.

Siegman:TB4-1-41

[217] A. E. Siegman. TEX as a programming language? TUGboat, 4(1):41–42, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07sieg.pdf.

Anonymous:TB4-1-43

[218] Anonymous. Textset, Inc. — A service for TEX users. TUGboat, 4(1):43, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07ads.pdf.

Anonymous:TB4-1-44

[219] Anonymous. Quality Micro Systems — Lasergrafix 1200. TUGboat, 4(1): 44, April 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-1/tb07ads.pdf.

Anonymous:TB4-1-45

[220] Anonymous. TEX82 order form. TUGboat, 4(1):45–46, April 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-1/tb07misc.
pdf.

Anonymous:TB4-1-47

[221] Anonymous. Membership application and order form. TUGboat, 4(1):47–48, April 1983. ISSN 0896-3207. URL https://tug.org/join.html.

Whidden:TB4-1-xx

[222] Samuel B. Whidden. TUG financial reports. TUGboat, 4(1):??, April

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-1/
tb07gendel.pdf.

Anonymous:TB4-2-50

[223] Anonymous. Addresses of officers, authors and others. TUGboat, 4(2):50, September 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-2/tb08addresses.pdf.

Anonymous:TB4-2-551

[224] Anonymous. Official announcements. TUGboat, 4(2):51, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08ann.pdf.

MacKay:TB4-2-52

[225] Pierre MacKay. Message from the President. TUGboat, 4(2):52–53, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08gendel.pdf.

Anonymous:TB4-2-55

[226] Anonymous. Summary of the technical program, TEX User Group meeting, Stanford University, July 11–15, 1983. TUGboat, 4(2):55–57, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08gendel.pdf.

Anonymous:TB4-2-57

[227] Anonymous. Participants, TUG meeting and AMS-TEX82 short course. TUGboat, 4(2):57–59, September 1983. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb04-2/tb08gendel.pdf.

Plass:TB4-2-59

[228] Susan Plass. Final report of the Bylaws Committee. TUGboat, 4(2):59, September 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-2/tb08gendel.pdf.

Anonymous:TB4-2-60

[229] Anonymous. TEX Users Group — TUG Operating Procedures. TUGboat, 4 (2):60–61, September 1983. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb04-2/tb08tugopproc.pdf.

Whidden:TB4-2-62

[230] Samuel B. Whidden. TUG Treasurer’s report. TUGboat, 4(2):62–63, Septem-

ber 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-
2/tb08treas.pdf.

Price:TB4-2-63

[231] Lynne Price. Report on ANSI X3J6. TUGboat, 4(2):63–64, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08price-ansi.pdf.

Knuth:TB4-2-64

[232] Donald Knuth. A note on hyphenation. TUGboat, 4(2):64–65, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08knut.pdf.

Fuchs:TB4-2-65

[233] David Fuchs. TEX vs. INITEX. TUGboat, 4(2):65–66, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08fuchs.pdf.

Fuchs:TB4-2-66

[234] David Fuchs. TEXhax summary. TUGboat, 4(2):66–70, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08fuchs.pdf.

Anonymous:TB4-2-71

[235] Anonymous. Chart: Output devices and computers. TUGboat, 4(2):71, September 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-2/tb08output.pdf.

Fuchs:TB4-2-72

[236] David Fuchs. News from the TEX Project. TUGboat, 4(2):72–73, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08site.pdf.

Mallett:TB4-2-73

[237] Rick Mallett. TEX82 on CP-6. TUGboat, 4(2):73–74, September 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-2/tb08site.
pdf.

Plass:TB4-2-74

[238] Susan Plass. IBM site report. TUGboat, 4(2):74, September 1983. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb04-2/tb08site.pdf.

Furuta:TB4-2-74

[239] Richard Furuta. Unix TEX site report. TUGboat, 4(2):74–75, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08site.pdf.

Nichols:TB4-2-75

[240] Monte C. Nichols. VAX/VMS site report. TUGboat, 4(2):75–76, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08site.pdf.

MacKay:TB4-2-76

[241] Pierre MacKay. TEX for Arabic script. TUGboat, 4(2):76, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08mackay.pdf.

Mooney:TB4-2-76

[242] Jim Mooney. Revised Varian output driver in VAX/VMS Fortran. TUGboat, 4(2):76, September 1983. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb04-2/tb08site.pdf.

Carnes:TB4-2-77

[243] Lance Carnes. Table of “small” TEX implementations. TUGboat, 4(2):77, September 1983. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb04-2/tb08carnes.pdf.

Price:TB4-2-78

[244] Lynne Price. Problems from the TUG meeting: Framed slides; Multiple marks. TUGboat, 4(2):78–79, September 1983. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb04-2/tb08price.pdf.

Anonymous:TB4-2-79

[245] Anonymous. TUGboat macro index. TUGboat, 4(2):79–80, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08tubmacidx.pdf.

Sterken:TB4-2-80

[246] Jim Sterken. First line of paragraph. TUGboat, 4(2):80–81, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08sterken.pdf.

Whidden:1983:LA

[247] Sam Whidden and J. R. Roesser. Observations on TEX from a divergent viewpoint: Introduction and memo to STI staff. TUGboat, 4(2):81–89, September 1983. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb04-2/tb08letters.pdf.

Knuth:1983:OTD

[248] Don Knuth, David Fuchs, Michael Spivak, Richard Palais, Barbara Beeton, and J. R. Roesser. Observations on TEX from a divergent viewpoint: Comments, response, and reresponse. TUGboat, 4(2):90–102, September 1983. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08letters.pdf.

Spivak:TB4-2-103

[249] Michael Spivak. Summary of AMS-TEX. TUGboat, 4(2):103–126, September

1983. ISSN 0896-3207. URL https://tug.org/TUGboat/tb04-2/
tb08spivak.pdf.

Anonymous:TB4-2-127

[250] Anonymous. Textset, Inc. TUGboat, 4(2):127, September 1983. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb04-2/tb08ads.pdf.

Anonymous:TB4-2-127-2

[251] Anonymous. TEX lectures on tape. TUGboat, 4(2):127, September 1983. ISSN

0896-3207. URL https://tug.org/TUGboat/tb04-2/tb08ads.
pdf.

Anonymous:TB4-2-128

[252] Anonymous. Imagen Corporation — Intelligent page printer systems. TUGboat, 4(2):128, September 1983. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb04-2/tb08ads.pdf.

Anonymous:TB4-2-129

[253] Anonymous. TEX82 Order Form. TUGboat, 4(2):129–130, September 1983.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB4-2-131

[254] Anonymous. Membership application and order form. TUGboat, 4(2):131– 132, September 1983. ISSN 0896-3207. URL https://tug.org/join.

html.

Anonymous:TB5-1-2

[255] Anonymous. Addresses of officers, authors and others. TUGboat, 5(1):2, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09complete.pdf.

Anonymous:TB5-1-3

[256] Anonymous. Official announcements. TUGboat, 5(1):3, May 1984. ISSN 0896-3207.

MacKay:TB5-1-4

[257] Pierre MacKay. Message from the President. TUGboat, 5(1):4, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09gendel.pdf.

Knuth:TB5-1-4

[258] Donald E. Knuth. TEX incunabula. TUGboat, 5(1):4–11, May 1984. ISSN

0896-3207. URL https://tug.org/TUGboat/tb05-1/tb09knut.
pdf.

Whidden:TB5-1-12

[259] Samuel B. Whidden. TUG 1983 financial report with comparisons. TUGboat, 5(1):12–13, May 1984. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb05-1/tb09treas.pdf.

Anonymous:TB5-1-13

[260] Anonymous. Preliminary program — August 1984 Stanford meeting. TUGboat, 5(1):13–14, May 1984. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb05-1/tb09meet.pdf.

Canzii:TB5-1-14

[261] G. Canzii, F. Genolini, and D. Lucarella. Hyphenation of Italian words. TUGboat, 5(1):14–15, May 1984. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb05-1/tb09canzii.pdf.

Anonymous:TB5-1-15

[262] Anonymous. Hyphenation exception log. TUGboat, 5(1):15, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09canzii.pdf.

Anonymous:TB5-1-16

[263] Anonymous. Chart: Output devices and computers. TUGboat, 5(1):16, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09output.pdf.

Anonymous:TB5-1-17

[264] Anonymous. Index to sample output from various devices. TUGboat, 5(1): 17, May 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-1/tb09output.pdf.

Haus:TB5-1-17

[265] Goffredo Haus. How to tame your phototypesetter by TEX. TUGboat, 5(1): 17, May 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-1/tb09output.pdf.

Noot:TB5-1-18

[266] Han Noot. DVI-code to the Harris 7500. TUGboat, 5(1):18–21, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09output.pdf.

Fuchs:TB5-1-22

[267] David Fuchs. News from the TEX Project. TUGboat, 5(1):22–23, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09site.pdf.

Lofstedt:TB5-1-23

[268] Benedict Løfstedt. CDC TEX at RECAU. TUGboat, 5(1):23, May 1984. ISSN

0896-3207. URL https://tug.org/TUGboat/tb05-1/tb09site.
pdf.

Guenther:TB5-1-24

[269] Dean Guenther, Alan Hagen-Wittbecker, and Janene Winter. TEX at Washington State University. TUGboat, 5(1):24–5, May 1984. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb05-1/tb09site.pdf.

Childs:TB5-1-26

[270] Bart Childs. Data General site report. TUGboat, 5(1):26, May 1984. ISSN

0896-3207. URL https://tug.org/TUGboat/tb05-1/tb09site.
pdf.

Childs:TB5-1-26-2

[271] Bart Childs. Prime site report. TUGboat, 5(1):26, May 1984. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb05-1/tb09site.pdf.

Furuta:TB5-1-27

[272] Richard Furuta. Unix TEX site report. TUGboat, 5(1):27–28, May 1984. ISSN

0896-3207. URL https://tug.org/TUGboat/tb05-1/tb09site.
pdf.

Rodgers:TB5-1-28

[273] D. L. Rodgers, J. J. Sterken, and P. Grosso. TEX 1.0 on Sun workstations. TUGboat, 5(1):28–29, May 1984. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb05-1/tb09site.pdf.

Nichols:TB5-1-29

[274] Monte C. Nichols. VAX/VMS site report. TUGboat, 5(1):29–30, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09site.pdf.

Kellerman:TB5-1-30

[275] David Kellerman and Barry Smith. Notice to VAX/VMS users. TUGboat, 5 (1):30, May 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-1/tb09site.pdf.

Anonymous:TB5-1-31

[276] Anonymous. Metafont Generic Font file format. TUGboat, 5(1):31–33, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09gf.pdf.

Anonymous:TB5-1-34

[277] Anonymous. Gray fonts for Metafont proofs. TUGboat, 5(1):34–35, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09gray.pdf.

Tobin:TB5-1-36

[278] Georgia K. M. Tobin. The OCLC Roman family of fonts. TUGboat, 5(1):36– 46, May 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-1/tb09tobin.pdf.

Carnes:TB5-1-47

[279] Lance Carnes. Table of “small” TEX implementations. TUGboat, 5(1):47, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09carnes.pdf.

Beeton:TB5-1-48

[280] Barbara Beeton. \relax and watch the numbers. TUGboat, 5(1):48, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09beet.pdf.

Daniels:TB5-1-49

[281] Susan Daniels. The HP TEX macros. TUGboat, 5(1):49–66, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09daniels.pdf.

Anonymous:TB5-1-67

[282] Anonymous. Classified. TUGboat, 5(1):67, May 1984. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb05-1/tb09lett.pdf.

Knuth:TB5-1-67

[283] Don Knuth. Comments on quality in publishing. TUGboat, 5(1):67, May

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-1/
tb09lett.pdf.

Anonymous:TB5-1-68

[284] Anonymous. Textset, Inc. TUGboat, 5(1):68, May 1984. ISSN 0896-3207.

Anonymous:TB5-1-69

[285] Anonymous. TEX82 Order Form. TUGboat, 5(1):69–70, May 1984. ISSN

0896-3207. URL https://tug.org/TUGboat/tb05-1/tb09backm.
pdf.

Anonymous:TB5-1-71

[286] Anonymous. Membership application and order form. TUGboat, 5(1):71–72, May 1984. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB5-2-74

[287] Anonymous. Addresses of officers, authors and others. TUGboat, 5(2):74, November 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-2/tb10complete.pdf.

Anonymous:TB5-2-75

[288] Anonymous. Official announcements. TUGboat, 5(2):75, November 1984. ISSN 0896-3207.

MacKay:TB5-2-76

[289] Pierre MacKay. Message from the President. TUGboat, 5(2):76, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10gendel.pdf.

Welland:TB5-2-77

[290] Robert Welland. Report of the Publications Committee. TUGboat, 5(2):77, November 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-2/tb10gendel.pdf.

Keller:TB5-2-77

[291] Arthur Keller. Report of the Special Projects Committee. TUGboat, 5(2): 77–78, November 1984. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb05-2/tb10gendel.pdf.

Beeton:TB5-2-78

[292] Barbara Beeton. Submitting items to TUGboat. TUGboat, 5(2):78, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10gendel.pdf.

Southall:TB5-2-79

[293] Richard Southall. First principles of typographic design for document production. TUGboat, 5(2):79–90, November 1984. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb05-2/tb10south.pdf.

Jurgensen:TB5-2-91

[294] Helmut Jürgensen. Editor’s introduction. TUGboat, 5(2):91, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10jurg.pdf.

Desarmenien:TB5-2-91

[295] Jacques Désarménien. How to run TEX in a French environment: Hyphenation, fonts, typography. TUGboat, 5(2):91–102, November 1984. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb05-2/tb10desarm.pdf.

Schulze:TB5-2-103

[296] Bernd Schulze. German hyphenation and Umlauts in TEX. TUGboat, 5(2): 103–104, November 1984. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb05-2/tb10schulze.pdf.

Knuth:TB5-2-105

[297] Donald E. Knuth. A course on programming. TUGboat, 5(2): 105–118, November 1984. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb05-2/tb10knut.pdf.

Hobby:TB5-2-119

[298] John D. Hobby and Gu Guoan. A Chinese meta-font. TUGboat, 5(2):119–136, November 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-2/tb10hobby.pdf.

Anonymous:TB5-2-137

[299] Anonymous. Chart: Output devices and computers. TUGboat, 5(2):137, November 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-2/tb10output.pdf.

Anonymous:TB5-2-138

[300] Anonymous. Index to sample output from various devices. TUGboat, 5(2):138, November 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-2/tb10output.pdf.

Rokicki:TB5-2-138

[301] Tom Rokicki. Dvi QMS: an example of a driver. TUGboat, 5(2):138–139, November 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-2/tb10output.pdf.

Rokicki:TB5-2-139

[302] Tom Rokicki. \special. TUGboat, 5(2):139–140, November 1984. ISSN

0896-3207. URL https://tug.org/TUGboat/tb05-2/tb10output.
pdf.

Anonymous:TB5-2-141

[303] Anonymous. TEX users’ activity in Germany. TUGboat, 5(2):141, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10site.pdf.

Crawford:TB5-2-141

[304] John Crawford. Prime site report. TUGboat, 5(2):141, November 1984. ISSN

0896-3207. URL https://tug.org/TUGboat/tb05-2/tb10site.
pdf.

Fuchs:TB5-2-141

[305] David Fuchs. News from the TEX Project. TUGboat, 5(2):141, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10site.pdf.

Furuta:TB5-2-142

[306] Richard Furuta. Unix TEX site report. TUGboat, 5(2):142, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10site.pdf.

Smith:TB5-2-142

[307] Barry Smith. VAX/VMS site report. TUGboat, 5(2):142, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10site.pdf.

Felippa:TB5-2-143

[308] Carlos A. Felippa. Feedback from TEX users at Lockheed. TUGboat, 5(2): 143–144, November 1984. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb05-2/tb10site.pdf.

Carnes:TB5-2-145

[309] Lance Carnes. Table of “small” TEX implementations. TUGboat, 5(2):145, November 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-2/tb10carnes.pdf.

Anonymous:TB5-2-146

[310] Anonymous. Calendar. TUGboat, 5(2):146, November 1984. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb05-2/tb10lettnews.pdf.

Clark:TB5-2-146

[311] Malcolm Clark. Mathematical communication with a deaf and blind student using TEX. TUGboat, 5(2):146, November 1984. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb05-2/tb10lettnews.pdf.

Anonymous:TB5-2-147

[312] Anonymous. Call for papers: TEX for scientific documentation, Varenna, Italy, May 16–17, 1985. TUGboat, 5(2):147, November 1984. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb05-2/tb10lettnews.pdf.

Whidden:TB5-2-148

[313] Samuel B. Whidden. TEX Users Group 1984 financial report with comparisons. TUGboat, 5(2):148–149, November 1984. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb05-2/tb10tugbus.pdf.

Dupree:TB5-2-149

[314] Chuck Dupree. Report of the August 1984 Steering Committee meeting. TUGboat, 5(2):149–152, November 1984. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb05-2/tb10tugbus.pdf.

Anonymous:TB5-2-153

[315] Anonymous. Profile of TEX installations available to TUG members. TUGboat, 5(2):153–155, November 1984. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb05-2/tb10instprof.pdf.

Anonymous:TB5-2-156

[316] Anonymous. Talaris talks TEX. TUGboat, 5(2):156–157, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10ads.pdf.

Anonymous:TB5-2-158

[317] Anonymous. Quality Micro Systems, Inc. TUGboat, 5(2):158, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10ads.pdf.

Anonymous:TB5-2-159

[318] Anonymous. Textset, Inc. TUGboat, 5(2):159, November 1984. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb05-2/tb10ads.pdf.

Anonymous:TB5-2-160

[319] Anonymous. The Metafoundry. TUGboat, 5(2):160, November 1984. ISSN

0896-3207. URL https://tug.org/TUGboat/tb05-2/tb10ads.
pdf.

Anonymous:TB5-2-160-2

[320] Anonymous. Instructors wanted for TUG courses. TUGboat, 5(2):160, November 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-2/tb10ads.pdf.

Anonymous:TB5-2-161

[321] Anonymous. Membership application and order form. TUGboat, 5(2): 161–164, November 1984. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb05-2/tb10backm.pdf.

Anonymous:TB5-2-165

[322] Anonymous. Order form for AMS publications. TUGboat, 5(2):165–166, November 1984. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb05-2/tb10backm.pdf.

Anonymous:TB5-2-167

[323] Anonymous. TEX82 order form. TUGboat, 5(2):167–168, November

1984. ISSN 0896-3207. URL https://tug.org/TUGboat/tb05-2/
tb10backm.pdf.

Anonymous:1985:TPa

[324] Anonymous. Title page. TUGboat, 6(1):1, March 1985. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11titlepage.pdf.

Anonymous:TB6-1-2

[325] Anonymous. Addresses of Officers, Authors and Others. TUGboat, 6(1):2, March 1985. ISSN 0896-3207.

Anonymous:TB6-1-3

[326] Anonymous. Official Announcements. TUGboat, 6(1):3, March 1985. ISSN

0896-3207. URL https://tug.org/TUGboat/tb06-1/tb11ann.
pdf.

MacKay:TB6-1-4

[327] Pierre MacKay. Message from the President. TUGboat, 6(1):4, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11gendel.pdf.

Beck:TB6-1-4

[328] Lawrence A. Beck. ANSI X3V1.8 Liaison Report to the TEX Users Group. TUGboat, 6(1):4–5, March 1985. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb06-1/tb11gendel.pdf.

Anonymous:TB6-1-5

[329] Anonymous. Acknowledgement of Contributions. TUGboat, 6(1):5, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11gendel.pdf.

Southall:TB6-1-6

[330] Richard Southall. First principles of typographic design for document production: Corrigenda. TUGboat, 6(1):6, March 1985. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb06-1/tb11gendel.pdf.

Eppstein:TB6-1-7

[331] Maureen Eppstein. Editor’s introduction. TUGboat, 6(1):7, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11meppstein.pdf.

Anonymous:TB6-1-8

[332] Anonymous. Generic font file format. TUGboat, 6(1):8–12, March 1985. ISSN

0896-3207. URL https://tug.org/TUGboat/tb06-1/tb11gf.
pdf.

Tobin:TB6-1-12

[333] G. K. M. Tobin. A bit of doggerel. TUGboat, 6(1):12, March 1985. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb06-1/tb11tobin.pdf.

Anonymous:TB6-1-13

[334] Anonymous. Chart: Output devices and computers. TUGboat, 6(1):13–15, March 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-1/tb11output.pdf.

Anonymous:TB6-1-15

[335] Anonymous. Index to sample output from various devices. TUGboat, 6(1): 15, March 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-1/tb11output.pdf.

Furuta:TB6-1-15

[336] Richard Furuta. Device drivers included on the UNIX TEX distribution tape. TUGboat, 6(1):15, March 1985. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb06-1/tb11output.pdf.

Luvisetto:TB6-1-15

[337] M. L. Luvisetto and E. Ugolini. A TEX82 spooler for VT and dot matrix printers. TUGboat, 6(1):15–16, March 1985. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb06-1/tb11output.pdf.

Anonymous:TB6-1-16

[338] Anonymous. How to get the latest news via TEXhax. TUGboat, 6(1):16– 17, March 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-1/tb11site.pdf.

Crawford:TB6-1-17

[339] John Crawford. Prime site report. TUGboat, 6(1):17, March 1985. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb06-1/tb11site.pdf.

Furuta:TB6-1-18

[340] Richard Furuta. Unix TEX site report. TUGboat, 6(1):18–20, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11site.pdf.

Smith:TB6-1-20

[341] Barry Smith. VAX/VMS site report. TUGboat, 6(1):20, March 1985. ISSN

0896-3207. URL https://tug.org/TUGboat/tb06-1/tb11site.
pdf.

Cumiskey:TB6-1-21

[342] James A. Cumiskey. TEX for tourists. TUGboat, 6(1):21–25, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11site.pdf.

Carnes:TB6-1-26

[343] Lance Carnes. Table of “small” TEX implementations. TUGboat, 6(1):26, March 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-1/tb11small.pdf.

Company:TB6-1-27

[344] Addison-Wesley Publishing Company. TEX now on microcomputers. TUGboat, 6(1):27–28, March 1985. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb06-1/tb11small.pdf.

Carnes:TB6-1-28

[345] Lance Carnes. PC TEX. TUGboat, 6(1):28–29, March 1985. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb06-1/tb11small.pdf.

Platt:TB6-1-29

[346] Craig Platt. Macros for two-column format. TUGboat, 6(1):29–30, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11platt.pdf.

Eppstein:TB6-1-31

[347] David Eppstein. Trees in TEX. TUGboat, 6(1):31–35, March 1985. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb06-1/tb11deppstein.
pdf.

Knuth:TB6-1-36

[348] Donald E. Knuth. Recipes and fractions. TUGboat, 6(1):36–38, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11knut.pdf.

Anonymous:TB6-1-38

[349] Anonymous. TEX, the program. TUGboat, 6(1):38, March 1985. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb06-1/tb11news.pdf.

Anonymous:TB6-1-39

[350] Anonymous. Calendar. TUGboat, 6(1):39, March 1985. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb06-1/tb11news.pdf.

Anonymous:TB6-1-39-2

[351] Anonymous. Conference: TEX for scientific documentation, Como, Italy, May 16–17, 1985. TUGboat, 6(1):39, March 1985. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb06-1/tb11news.pdf.

Anonymous:TB6-1-39-3

[352] Anonymous. TUG Annual Meeting, Stanford University, August 14–16, 1985. TUGboat, 6(1):39, March 1985. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb06-1/tb11news.pdf.

Anonymous:TB6-1-40

[353] Anonymous. TEX and reports available from Stanford Computer Science Department. TUGboat, 6(1):40, March 1985. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb06-1/tb11news.pdf.

Anonymous:TB6-1-41

[354] Anonymous. Textset, Inc. — TEX and the IBM PC, TEX-to-PostScript. TUGboat, 6(1):41, March 1985. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb06-1/tb11ads.pdf.

Anonymous:TB6-1-42

[355] Anonymous. Addison-Wesley — MicroTEX. TUGboat, 6(1):42, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11ads.pdf.

Anonymous:TB6-1-43

[356] Anonymous. TUG — TEX Lectures on Tape. TUGboat, 6(1):43, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11ads.pdf.

Anonymous:TB6-1-44

[357] Anonymous. Textset, Inc. — Professional TEXware. TUGboat, 6(1):44, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11ads.pdf.

Anonymous:TB6-1-47

[358] Anonymous. Membership application and order form. TUGboat, 6(1):47–48, March 1985. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB6-1-49

[359] Anonymous. Order form for AMS publications. TUGboat, 6(1):49–50, March 1985. ISSN 0896-3207.

Anonymous:TB6-1-51

[360] Anonymous. TEX82 order form. TUGboat, 6(1):51–52, March 1985. ISSN 0896-3207.

Anonymous:TB6-1-445

[361] Anonymous. Talaris. TUGboat, 6(1):445, March 1985. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb06-1/tb11ads.pdf.

Anonymous:TB6-1-446

[362] Anonymous. Personal TEX, Inc. — PC TEX. TUGboat, 6(1):446, March

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-1/
tb11ads.pdf.

Anonymous:1985:IC

[363] Anonymous. Intergraph Corporation. TUGboat, 6(1):c3, March 1985. CODEN ???? ISSN 0896-3207.

Anonymous:TB6-1-cover-3

[364] Anonymous. Intergraph Corporation. TUGboat, 6(1):Cover 3, March 1985. ISSN 0896-3207.

Anonymous:1985:TPb

[365] Anonymous. Title page. TUGboat, 6(2):53, July 1985. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12titlepage.pdf.

Anonymous:TB6-2-54

[366] Anonymous. Addresses of Officers, Authors and Others. TUGboat, 6(2):54, July 1985. ISSN 0896-3207.

Anonymous:TB6-2-55

[367] Anonymous. Official Announcements. TUGboat, 6(2):55, July 1985. ISSN

0896-3207. URL https://tug.org/TUGboat/tb06-2/tb12ann.
pdf.

Anonymous:TB6-2-56

[368] Anonymous. TUG 1985 Annual Meeting, Stanford University, August 12–14, 1985. TUGboat, 6(2):56, July 1985. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb06-2/tb12gendel.pdf.

Jurgensen:TB6-2-56

[369] Helmut Jürgensen. Editor’s message. TUGboat, 6(2):56, July 1985. ISSN

0896-3207. URL https://tug.org/TUGboat/tb06-2/tb12gendel.
pdf.

Ferguson:TB6-2-57

[370] Michael J. Ferguson. A multilingual TÊX. TUGboat, 6(2):57–58, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12ferguson.pdf.

Beeton:TB6-2-59

[371] Barbara Beeton. Mathematical symbols and cyrillic fonts ready for distribution. TUGboat, 6(2):59–63, July 1985. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb06-2/tb12beet.pdf.

Anonymous:TB6-2-64

[372] Anonymous. Chart: Output devices and computers. TUGboat, 6(2):64–66,

July 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-
2/tb12output.pdf.

Anonymous:TB6-2-66

[373] Anonymous. Index to sample output from various devices. TUGboat, 6(2):66,

July 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-
2/tb12output.pdf.

Spragens:TB6-2-66

[374] Alan Spragens. Graphics commands for TEX — Discussion in TEXhax conference. TUGboat, 6(2):66–68, July 1985. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb06-2/tb12spragens.pdf.

Naugle:TB6-2-69

[375] Norman W. Naugle and Tomas G. Rokicki. Miscellaneous activity at Texas A&M. TUGboat, 6(2):69–72, July 1985. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb06-2/tb12naugle.pdf.

Anonymous:TB6-2-73

[376] Anonymous. TEXhax activity at Spring DECUS. TUGboat, 6(2):73, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12site.pdf.

Spragens:TB6-2-73

[377] Alan Spragens. CMS TEX site report. TUGboat, 6(2):73–75, July 1985. ISSN

0896-3207. URL https://tug.org/TUGboat/tb06-2/tb12site.
pdf.

Riesel:TB6-2-76

[378] Hans Riesel. Report on experience with TEX80. TUGboat, 6(2):76–79, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12site.pdf.

Marriott:TB6-2-80

[379] Gregory Marriott. A TEX82 implementation on the HP9000 series 500. TUGboat, 6(2):80, July 1985. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb06-2/tb12site.pdf.

Furuta:TB6-2-81

[380] Richard Furuta. Unix TEX site report. TUGboat, 6(2):81, July 1985. ISSN

0896-3207. URL https://tug.org/TUGboat/tb06-2/tb12site.
pdf.

Smith:TB6-2-81

[381] Barry Smith. VAX/VMS site report. TUGboat, 6(2):81, July 1985. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb06-2/tb12site.pdf.

Carnes:TB6-2-82

[382] Lance Carnes. Table of “small” TEX implementations. TUGboat, 6(2):82, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12small.pdf.

Hendrickson:TB6-2-83

[383] Amy Hendrickson. Some diagonal line hacks. TUGboat, 6(2):83–86, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12hendiag.pdf.

Hendrickson:TB6-2-86

[384] Amy Hendrickson. A (possibly) totally useless macro. TUGboat, 6(2):86, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12henduseless.pdf.

Mohr:TB6-2-87

[385] August Mohr. Multi-column layout in TEX80. TUGboat, 6(2):87–95, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12mohr.pdf.

Bonnetain:TB6-2-96

[386] Jean-Luc Bonnetain. Author vs. proofreader. TUGboat, 6(2):96, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12lett.pdf.

Greene:TB6-2-96

[387] John Greene. Transatlantic TEX. TUGboat, 6(2):96, July 1985. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb06-2/tb12lett.pdf.

Anonymous:TB6-2-97

[388] Anonymous. Calendar. TUGboat, 6(2):97, July 1985. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb06-2/tb12news.pdf.

Anonymous:TB6-2-98

[389] Anonymous. International Conference on Text Processing and Document Manipulation, Nottingham, England, April 14–16, 1986. TUGboat, 6(2):98, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12news.pdf.

Anonymous:TB6-2-98-2

[390] Anonymous. PROTEXT II, Dublin, Ireland, October 21–25, 1985. TUGboat, 6 (2):98, July 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-2/tb12news.pdf.

Whidden:TB6-2-99

[391] Samuel B. Whidden. TUG 1984 financial report with comparisons. TUGboat, 6(2):99–100, July 1985. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb06-2/tb12treas.pdf.

Anonymous:TB6-2-100

[392] Anonymous. The Metafoundry — Introducing mf medley. TUGboat, 6(2): 100, July 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-2/tb12ads.pdf.

Anonymous:TB6-2-101

[393] Anonymous. TEX and reports available from the Stanford Computer Science Department. TUGboat, 6(2):101, July 1985. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb06-2/tb12ads.pdf.

Anonymous:TB6-2-102

[394] Anonymous. Addison-Wesley — MicroTEX. TUGboat, 6(2):102, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12ads.pdf.

Anonymous:TB6-2-103

[395] Anonymous. Textset, Inc. TUGboat, 6(2):103, July 1985. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb06-2/tb12ads.pdf.

Anonymous:TB6-2-104

[396] Anonymous. Personal TEX, Inc. — PC TEX is here. TUGboat, 6(2):104, July

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-2/
tb12ads.pdf.

Anonymous:TB6-2-105

[397] Anonymous. Membership application and order form. TUGboat, 6(2):105– 106, July 1985. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB6-2-107

[398] Anonymous. TEX82 order form. TUGboat, 6(2):107–108, July 1985. ISSN 0896-3207.

Anonymous:TB6-2-cover-3

[399] Anonymous. TEX lectures on tape. TUGboat, 6(2):Cover 3, July 1985. ISSN 0896-3207.

Anonymous:1985:TPc

[400] Anonymous. Title page. TUGboat, 6(3):109, November 1985. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13titlepage.pdf.

Anonymous:TB6-3-110

[401] Anonymous. Addresses of officers, authors and others. TUGboat, 6(3):110, November 1985. ISSN 0896-3207.

Anonymous:TB6-3-111

[402] Anonymous. Official announcements. TUGboat, 6(3):111, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13ann.pdf.

MacKay:TB6-3-112

[403] Pierre MacKay. Message from the outgoing President. TUGboat, 6(3): 112–113, November 1985. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb06-3/tb13gendel.pdf.

Childs:TB6-3-113

[404] Bart Childs. From the President. TUGboat, 6(3):113–114, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13gendel.pdf.

Anonymous:TB6-3-114

[405] Anonymous. Statement of principles. TUGboat, 6(3):114, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13gendel.pdf.

Rokicki:TB6-3-115

[406] Tomas Rokicki. Packed (PK) font file format. TUGboat, 6(3):115–120, November 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-3/tb13pk.pdf.

Anonymous:TB6-3-121

[407] Anonymous. Hyphenation exception log. TUGboat, 6(3):121, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13hyf.pdf.

Tobin:TB6-3-122

[408] Georgia K. M. Tobin. Editor’s introduction. TUGboat, 6(3):122–123, Novem-

ber 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-
3/tb13tobin.pdf.

Beeton:TB6-3-124

[409] Barbara Beeton. Mathematical symbols and cyrillic fonts ready for distribution (revised). TUGboat, 6(3):124–128, November 1985. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb06-3/tb13beetcyr.pdf.

Anonymous:TB6-3-129

[410] Anonymous. Charts: Output devices and computers. TUGboat, 6(3):129–131, November 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-3/tb13output.pdf.

Anonymous:TB6-3-131

[411] Anonymous. Index to sample output from various devices. TUGboat, 6(3):131, November 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-3/tb13site.pdf.

Anonymous:TB6-3-131-2

[412] Anonymous. The TEX directories at the “source” — Stanford. TUGboat, 6(3): 131–132, November 1985. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb06-3/tb13site.pdf.

Anonymous:TB6-3-132

[413] Anonymous. TEX in the UK and Ireland: TEXline. TUGboat, 6(3):132, Novem-

ber 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-
3/tb13site.pdf.

Childs:TB6-3-132

[414] Bart Childs. Data General distribution news. TUGboat, 6(3):132–133, Novem-

ber 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-
3/tb13site.pdf.

Platt:TB6-3-133

[415] Craig Platt. MVS TEX site report. TUGboat, 6(3):133–134, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13site.pdf.

Guenther:TB6-3-134

[416] Dean Guenther. TEX at Washington State University. TUGboat, 6(3):134, November 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-3/tb13site.pdf.

Hayashi:TB6-3-135

[417] Tsunetoshi Hayashi. Reports on TEX implementation at HUCC. TUGboat, 6 (3):135–136, November 1985. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb06-3/tb13site.pdf.

Furuta:TB6-3-137

[418] Richard Furuta. Unix TEX site report. TUGboat, 6(3):137–138, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13site.pdf.

Anonymous:TB6-3-139

[419] Anonymous. Table of “small” TEX implementations. TUGboat, 6(3):139, November 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-3/tb13small.pdf.

Smith:TB6-3-139

[420] Barry Smith. VAX/VMS site report. TUGboat, 6(3):139, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13site.pdf.

Pfeffer:TB6-3-140

[421] Mitch Pfeffer and Alan Hoenig. Assembling a moderately-priced, highperformance clone of the IBM PC for running TEX. TUGboat, 6(3):140–145, November 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-3/tb13pfeffer.pdf.

Beeton:TB6-3-146

[422] Barbara Beeton. Review: The Boston Computer Society’s IBM PC & compatibles technical word processor review. TUGboat, 6(3):146–149, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13beetrev.pdf.

Beeton:TB6-3-150

[423] Barbara Beeton. The PLAIN truth: \buildrel. TUGboat, 6(3):150, Novem-

ber 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-
3/tb13beetplain.pdf.

Lamport:TB6-3-150

[424] Leslie Lamport. Announcement of L^{A}TEX Version 2.09. TUGboat, 6(3): 150–151, November 1985. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb06-3/tb13lamport.pdf.

Leban:TB6-3-151

[425] Bruce Leban. A solution to the tower of Hanoi problem using TEX. TUGboat, 6 (3):151–154, November 1985. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb06-3/tb13leban.pdf.

Cameron:TB6-3-155

[426] A. G. W. Cameron. Wiggly lines. TUGboat, 6(3):155, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13cameron.pdf.

Wonneberger:TB6-3-156

[427] Reinhard Wonneberger. Stream lists and related list types for L^{A}TEX. TUGboat, 6(3):156–157, November 1985. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb06-3/tb13wonneberger.pdf.

Anonymous:TB6-3-158

[428] Anonymous. L^{A}TEX command summary (Version 2.05) available from TUG. TUGboat, 6(3):158, November 1985. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb06-3/tb13lett.pdf.

Hendricks:TB6-3-158

[429] Rick Hendricks. Inquiry: Historical compilations. TUGboat, 6(3):158, Novem-

ber 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-
3/tb13lett.pdf.

Hornbach:TB6-3-158

[430] Kathy Hornbach. Inquiry: 35mm slides with SliTEX. TUGboat, 6(3):158, November 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-3/tb13lett.pdf.

Anonymous:TB6-3-159

[431] Anonymous. Calendar. TUGboat, 6(3):159, November 1985. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb06-3/tb13news.pdf.

Anonymous:TB6-3-159-2

[432] Anonymous. Acknowledgement of contributions. TUGboat, 6(3):159, Novem-

ber 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-
3/tb13news.pdf.

Anonymous:TB6-3-160

[433] Anonymous. TEX for Scientific Documentation, Strasbourg, France, June 19– 21, 1986. TUGboat, 6(3):160, November 1985. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb06-3/tb13news.pdf.

Whidden:TB6-3-161

[434] Samuel B. Whidden. Proposed TUG 1986 budget. TUGboat, 6(3):161, November 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-3/tb13treas.pdf.

Anonymous:TB6-3-162

[435] Anonymous. Addison-Wesley — MicroTEX. TUGboat, 6(3):162, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13ads.pdf.

Anonymous:TB6-3-163

[436] Anonymous. Textset, Inc. TUGboat, 6(3):163, November 1985. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb06-3/tb13ads.pdf.

Anonymous:TB6-3-164

[437] Anonymous. TUG — TEX lectures on videotape. TUGboat, 6(3):164–165, November 1985. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb06-3/tb13ads.pdf.

Anonymous:TB6-3-166

[438] Anonymous. Personal TEX, Inc. — PC TEX. TUGboat, 6(3):166, November

1985. ISSN 0896-3207. URL https://tug.org/TUGboat/tb06-3/
tb13ads.pdf.

Anonymous:TB6-3-167

[439] Anonymous. TUG membership application. TUGboat, 6(3):167–168, November 1985. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB6-3-169

[440] Anonymous. TUG order form. TUGboat, 6(3):169–172, November 1985. ISSN 0896-3207.

Anonymous:TB6-3-173

[441] Anonymous. TEX82 order form. TUGboat, 6(3):173–174, November 1985. ISSN 0896-3207.

Anonymous:1985:TAM

[442] Anonymous. TUG Annual Meeting, Tufts University, July 28–31, 1986. TUGboat, 6(3):c3, November 1985. CODEN ???? ISSN 0896-3207.

Anonymous:TB6-3-cover-3

[443] Anonymous. TUG Annual Meeting, Tufts University, July 28–31, 1986. TUGboat, 6(3):Cover 3, November 1985. ISSN 0896-3207.

Anonymous:1986:TPa

[444] Anonymous. Title page. TUGboat, 7(1):1, March 1986. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14titlepage.pdf.

Anonymous:TB7-1-4

[445] Anonymous. Addresses. TUGboat, 7(1):4–5, March 1986. ISSN 0896-3207.

Anonymous:TB7-1-6

[446] Anonymous. Calendar. TUGboat, 7(1):6, March 1986. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb07-1/tb14calendar.pdf.

Anonymous:TB7-1-7

[447] Anonymous. Acknowledgment of contributions. TUGboat, 7(1):7, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14news.pdf.

Anonymous:TB7-1-7-2

[448] Anonymous. Donald E. Knuth Scholarship. TUGboat, 7(1):7, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14news.pdf.

Childs:TB7-1-7

[449] Bart Childs. From the President. TUGboat, 7(1):7, March 1986. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb07-1/tb14news.pdf.

Anonymous:TB7-1-8

[450] Anonymous. Special credits for this issue. TUGboat, 7(1):8, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14news.pdf.

Anonymous:TB7-1-8-2

[451] Anonymous. Conference: TEX for Scientific Documentation. TUGboat, 7(1): 8, March 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-1/tb14news.pdf.

Anonymous:TB7-1-8-3

[452] Anonymous. TEX courses. TUGboat, 7(1):8–9, March 1986. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb07-1/tb14news.pdf.

Anonymous:TB7-1-9

[453] Anonymous. Software Column: Call for Papers. TUGboat, 7(1):9–10, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14news.pdf.

Anonymous:TB7-1-10

[454] Anonymous. Addenda: The BCS Word Processor Review. TUGboat, 7(1): 10, March 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-1/tb14news.pdf.

Howell:TB7-1-11

[455] Gordon Howell. Translators to generate TEX files. TUGboat, 7(1):11, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14letters.pdf.

Lamport:TB7-1-11

[456] Leslie Lamport. L^{A}TEX input to ACM journals. TUGboat, 7(1):11, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14letters.pdf.

Childs:TB7-1-12

[457] Bart Childs. Data General Site Report. TUGboat, 7(1):12, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14short.pdf.

Fox:TB7-1-12

[458] Jim Fox. CDC Cyber Site Report. TUGboat, 7(1):12, March 1986. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb07-1/tb14short.pdf.

Fuchs:TB7-1-12

[459] David Fuchs. News from the TEX Project. TUGboat, 7(1):12, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14short.pdf.

Smith:TB7-1-13

[460] Barry Smith. Macintosh Site Report. TUGboat, 7(1):13, March 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-1/tb14short.
pdf.

MacKay:TB7-1-13

[461] Pierre MacKay. UNIX Site Report. TUGboat, 7(1):13–14, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14short.pdf.

Smith:TB7-1-15

[462] Barry Smith. VAX/VMS site report. TUGboat, 7(1):15, March 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-1/tb14short.
pdf.

Lamport:TB7-1-15

[463] Leslie Lamport. L^{A}TEX News. TUGboat, 7(1):15–16, March 1986. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb07-1/tb14short.pdf.

Ferguson:TB7-1-16

[464] Michael Ferguson. Multilingual TÊX update. TUGboat, 7(1):16, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14short.pdf.

Anonymous:TB7-1-17

[465] Anonymous. Where to find TEX file descriptions. TUGboat, 7(1):17, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14notes.pdf.

Auerbach:TB7-1-17

[466] Alan Auerbach. Automatic page sizing. TUGboat, 7(1):17, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14notes.pdf.

Beeton:TB7-1-18

[467] B. Beeton. The PLAIN truth: \displaylines, \ialign. TUGboat, 7(1):18–19, March 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-1/tb14beet.pdf.

Appelt:TB7-1-20

[468] W. Appelt and K. Horn. Multiple changefiles in WEB. TUGboat, 7(1):20– 21, March 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-1/tb14appelt.pdf.

Damerell:TB7-1-22

[469] R. M. Damerell. Error detecting changes to TANGLE. TUGboat, 7(1):22– 24, March 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-1/tb14damerell.pdf.

Trevorrow:TB7-1-25

[470] A. Trevorrow. DVItoVDU: A TEX page previewer. TUGboat, 7(1):25–36, March 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-1/tb14trevorrow.pdf.

Varian:TB7-1-37

[471] H. Varian and J. Sterken. MacDraw pictures in TEX documents. TUGboat, 7(1):37–40, March 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-1/tb14varian.pdf.

Maus:TB7-1-41

[472] D. Maus and B. Baker. DVILASER/PS extensions to L^{A}TEX. TUGboat, 7 (1):41–47, March 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-1/tb14maus.pdf.

Fina:TB7-1-48

[473] P. Fina. Results of the 1985 TUG Questionnaire. TUGboat, 7(1):48–50, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14fina.pdf.

Hoenig:TB7-1-51

[474] A. Hoenig and M. Pfeffer. TEX text editors for the IBM PC. TUGboat, 7(1):51– 54, March 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-1/tb14hoenig.pdf.

Senn:TB7-1-55

[475] M. Senn. TEX output previewers. TUGboat, 7(1):55–56, March 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-1/tb14senn.
pdf.

Beeton:TB7-1-57

[476] B. Beeton. TEX output devices. TUGboat, 7(1):57–66, March 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-1/tb14output.
pdf.

Beeton:TB7-1-67

[477] B. Beeton. Memory Sizes in Various Implementations of TEX. TUGboat, 7(1):67–70, March 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-1/tb14mem.pdf.

Carnes:TB7-1-71

[478] L. Carnes. “Small” TEX implementations. TUGboat, 7(1):71–72, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14small.pdf.

Anonymous:TB7-1-73

[479] Anonymous. Institutional Members. TUGboat, 7(1):73–74, March

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
tb14instmem.pdf.

Anonymous:TB7-1-88

[480] Anonymous. Index of advertisers. TUGboat, 7(1):88, March 1986. ISSN 0896-3207.

Anonymous:1986:FC

[481] Anonymous. Front cover. TUGboat, 7(1):c1, March 1986. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-1/
cover1.pdf.

Anonymous:1986:TPb

[482] Anonymous. Title page. TUGboat, 7(2):89, June 1986. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15titlepage.pdf.

Anonymous:1986:EI

[483] Anonymous. Editorial information. TUGboat, 7(2):90, June 1986. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15edit.pdf.

Anonymous:TB7-2-91

[484] Anonymous. Addresses. TUGboat, 7(2):91–92, June 1986. ISSN 0896-3207.

Beeton:TB7-2-93

[485] Barbara Beeton. Comments on the format of this issue. TUGboat, 7(2):93, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15beet.pdf.

Beeton:TB7-2-93-2

[486] Barbara Beeton. Computers & Typesetting coming out party. TUGboat, 7(2): 93, June 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-2/tb15beet.pdf.

Gordon:TB7-2-93

[487] Peter Gordon. Introducing Donald Knuth and Computers & Typesetting. TUGboat, 7(2):93–95, June 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-2/tb15gordon.pdf.

Knuth:TB7-2-95

[488] Donald Knuth. Remarks to celebrate the publication of Computers & Typesetting. TUGboat, 7(2):95–98, June 1986. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb07-2/tb15knut.pdf.

Jackson:TB7-2-98

[489] Cal Jackson. Comments on document design prompted by the new TUGboat format. TUGboat, 7(2):98–99, June 1986. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb07-2/tb15jackson.pdf.

Hornbach:TB7-2-99

[490] Kathy Hornbach. VAX language sensitive editor templates and guide for use with L^{A}TEX. TUGboat, 7(2):99, June 1986. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb07-2/tb15hornbach.pdf.

Tobin:TB7-2-100

[491] Georgia K. M. Tobin. Font forum. TUGboat, 7(2):100–101, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15tobin.pdf.

Knuth:TB7-2-101

[492] Donald E. Knuth. The TEX logo in various fonts. TUGboat, 7(2):101–102, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15knutlogo.pdf.

Anonymous:TB7-2-102

[493] Anonymous. Addenda to the output device charts. TUGboat, 7(2):102, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15output.pdf.

Hoenig:TB7-2-102

[494] Alan Hoenig. Meta- : An exhibit at the Cooper Union, NYC. TUGboat, 7(2):102, June 1986. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb07-2/tb15hoenig.pdf.

Crawford:TB7-2-103

[495] John Crawford. Prime 50 Series site report. TUGboat, 7(2):103, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15site.pdf.

Fox:TB7-2-103

[496] Jim Fox. CDC Cyber site report. TUGboat, 7(2):103, June 1986. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb07-2/tb15site.pdf.

Rokicki:TB7-2-103

[497] Tomas Rokicki. TEX on the Amiga. TUGboat, 7(2):103, June 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-2/tb15site.
pdf.

Pfeffer:TB7-2-104

[498] Mitch Pfeffer and Alan Hoenig. The Sperry IT: An IBM AT compatible. TUGboat, 7(2):104–105, June 1986. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb07-2/tb15pfeffer.pdf.

Ferguson:TB7-2-106

[499] Michael J. Ferguson. Table Making with INRSTEX. TUGboat, 7(2):106–109,

June 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-
2/tb15ferg.pdf.

Becker:TB7-2-109

[500] Helmut Becker. WEB system extensions. TUGboat, 7(2):109, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15queries.pdf.

Kennedy:TB7-2-109

[501] John Kennedy. Drama scripts. TUGboat, 7(2):109, June 1986. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb07-2/tb15queries.
pdf.

Alexander:TB7-2-110

[502] James Alexander. Side-by-side source/output samples; First-line special handling. TUGboat, 7(2):110, June 1986. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb07-2/tb15queries.pdf.

Fernandez:TB7-2-110

[503] Sylvester Fernandez. Marking changes in revised documents. TUGboat, 7(2): 110, June 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-2/tb15queries.pdf.

Kleine:TB7-2-111

[504] Karl Kleine. Customized editors for TEX. TUGboat, 7(2):111, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15queries.pdf.

Ludden:TB7-2-111

[505] Jim Ludden. Post-L^{A}TEX index formatting. TUGboat, 7(2):111, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15queries.pdf.

Spivak:TB7-2-111

[506] Mike Spivak. Letters. TUGboat, 7(2):111, June 1986. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb07-2/tb15spivak.pdf.

Anonymous:TB7-2-112

[507] Anonymous. Calendar. TUGboat, 7(2):112, June 1986. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb07-2/tb15calendar.pdf.

Anonymous:TB7-2-113

[508] Anonymous. Institutional members. TUGboat, 7(2):113–114, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15instmem.pdf.

Anonymous:TB7-2-115

[509] Anonymous. TUG membership application. TUGboat, 7(2):115–116, June

1986. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB7-2-117

[510] Anonymous. TEX82 order form. TUGboat, 7(2):117–118, June 1986. ISSN 0896-3207.

Anonymous:TB7-2-119

[511] Anonymous. Addison-Wesley — Computers & Typesetting. TUGboat, 7(2): 119, June 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-2/tb15ads.pdf.

Anonymous:TB7-2-120

[512] Anonymous. Personal TEX, Inc. — TEX for the PC. TUGboat, 7(2):120, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15ads.pdf.

Anonymous:TB7-2-121

[513] Anonymous. Talaris — Talaris 810. TUGboat, 7(2):121, June 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-2/tb15ads.
pdf.

Anonymous:TB7-2-122

[514] Anonymous. Computer Composition Corporation — Typesetting services. TUGboat, 7(2):122, June 1986. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb07-2/tb15ads.pdf.

Anonymous:TB7-2-123

[515] Anonymous. Textset, Inc. — New directions. TUGboat, 7(2):123, June

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-2/
tb15ads.pdf.

Anonymous:TB7-2-124

[516] Anonymous. FTL systems Inc. — TEX for the Macintosh. TUGboat, 7(2):124,

June 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-
2/tb15ads.pdf.

Anonymous:TB7-2-cover-3

[517] Anonymous. Addison-Wesley — TEX news. TUGboat, 7(2):Cover 3, June 1986. ISSN 0896-3207.

Anonymous:1986:TPc

[518] Anonymous. Title page. TUGboat, 7(3):125, October 1986. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16titlepage.pdf.

Anonymous:TB7-3-127

[519] Anonymous. Addresses. TUGboat, 7(3):127–128, October 1986. ISSN 0896-3207.

Childs:TB7-3-129

[520] Bart Childs. Message from the President. TUGboat, 7(3):129–130, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16gendel.pdf.

Anonymous:TB7-3-130

[521] Anonymous. Donald E. Knuth Scholarship. TUGboat, 7(3):130, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16gendel.pdf.

MacKay:TB7-3-131

[522] Pierre MacKay. 1987 Annual Meeting, Call for papers: TEX for the Humanities. TUGboat, 7(3):131, October 1986. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb07-3/tb16gendel.pdf.

Beck:TB7-3-132

[523] Lawrence A. Beck. Report from ANSI X3V1. TUGboat, 7(3):132, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16gendel.pdf.

Wonneberger:TB7-3-132

[524] Reinhard Wonneberger. Towards a TEX philology group. TUGboat, 7(3): 132–133, October 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-3/tb16gendel.pdf.

Mann:TB7-3-133

[525] Laurie Mann. An idea exchange on TEX training. TUGboat, 7(3):133, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16gendel.pdf.

Guntermann:TB7-3-134

[526] Klaus Guntermann and Wolfgang Rülling. Another approach to multiple changefiles. TUGboat, 7(3):134, October 1986. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb07-3/tb16gunterchange.pdf.

Guntermann:TB7-3-134-2

[527] Klaus Guntermann and Joachim Schrod. WEB adapted to C. TUGboat, 7 (3):134–137, October 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-3/tb16gunterweb.pdf.

Alexander:TB7-3-138

[528] James Alexander. TIb: a reference setting package for TEX. TUGboat, 7 (3):138–140, October 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-3/tb16alexander.pdf.

Rokicki:TB7-3-140

[529] Tomas Rokicki. Packed file format update. TUGboat, 7(3):140–144, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16pk.pdf.

Anonymous:TB7-3-145

[530] Anonymous. Hyphenation exception log. TUGboat, 7(3):145, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16hyf.pdf.

Bigelow:TB7-3-146

[531] Charles Bigelow. Notes on typeface protection. TUGboat, 7(3):146–151, Octo-

ber 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-
3/tb16bigelow.pdf.

Sauter:TB7-3-151

[532] John Sauter. Building Computer Modern Fonts. TUGboat, 7(3):151–152, October 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-3/tb16sauter.pdf.

Jansen:TB7-3-152

[533] Rick Jansen. TEX and Macintosh — new directions in Preview. TUGboat, 7 (3):152–158, October 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-3/tb16jansen.pdf.

Damerell:TB7-3-159

[534] R. M. Damerell. Crudetype: An adaptable device driver. TUGboat, 7(3): 159–163, October 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-3/tb16damerell.pdf.

Brown:TB7-3-164

[535] Malcolm Brown. TEXhax returns. TUGboat, 7(3):164, October 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-3/tb16site.
pdf.

Guntermann:TB7-3-164

[536] Klaus Guntermann. Porting TEX to the Atari ST. TUGboat, 7(3):164, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16site.pdf.

Fox:TB7-3-164

[537] Jim Fox. CDC Cyber site report. TUGboat, 7(3):164–165, October 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-3/tb16site.
pdf.

Platt:TB7-3-165

[538] Craig Platt. MVS TEX site report. TUGboat, 7(3):165–166, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16site.pdf.

Spragens:TB7-3-167

[539] Alan Spragens. CMS TEX site report. TUGboat, 7(3):167, October 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-3/tb16site.
pdf.

Hoenig:TB7-3-167

[540] Alan Hoenig and Mitch Pfeffer. Real typesetting from your personal computer. TUGboat, 7(3):167–168, October 1986. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb07-3/tb16hoenig.pdf.

Thull:TB7-3-169

[541] Klaus Thull. Harnessing TEX to compute third root of unity primes. TUGboat, 7(3):169–170, October 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-3/tb16thull.pdf.

Ehrbar:TB7-3-171

[542] Hans Ehrbar. Statistical graphics with TEX. TUGboat, 7(3):171–175, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16ehrbar.pdf.

Damrau:TB7-3-176

[543] Jackie Damrau and Ken Yap. Call for participation. TUGboat, 7(3):176, Octo-

ber 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-
3/tb16damrau.pdf.

Lamport:TB7-3-176

[544] Leslie Lamport. L^{A}TEX bugs. TUGboat, 7(3):176, October 1986. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb07-3/tb16lamport.
pdf.

Wonneberger:TB7-3-177

[545] Reinhard Wonneberger. Chapter mottos and optional semi-parameters. TUGboat, 7(3):177–185, October 1986. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb07-3/tb16wonneberger.pdf.

Hofmann:TB7-3-186

[546] Thomas Hofmann. A L^{A}TEX addition for formatting indexes. TUGboat, 7 (3):186, October 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-3/tb16hofmann.pdf.

Aurbach:TB7-3-187

[547] Richard L. Aurbach. IdxTEX and GloTEX — indexes and glossaries. TUGboat, 7(3):187, October 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-3/tb16queries.pdf.

Lee:TB7-3-187

[548] John Lee. Form letters. TUGboat, 7(3):187, October 1986. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb07-3/tb16queries.pdf.

Fox:TB7-3-187

[549] Jim Fox. Change bars. TUGboat, 7(3):187–188, October 1986. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb07-3/tb16queries.
pdf.

Sewell:TB7-3-188

[550] Wayne Sewell. Bugs in ware. TUGboat, 7(3):188–189, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16sewell.pdf.

Anonymous:TB7-3-190

[551] Anonymous. Calendar. TUGboat, 7(3):190, October 1986. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb07-3/tb16news.pdf.

Anonymous:TB7-3-191

[552] Anonymous. INRIA course: Structures for Documents, Aussois, France, January 19–23, 1987. TUGboat, 7(3):191, October 1986. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb07-3/tb16news.pdf.

Anonymous:TB7-3-191-2

[553] Anonymous. , TEX and the humanities, Rennes, France. TUGboat, 7(3):191, October 1986. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb07-3/tb16news.pdf.

Ohno:TB7-3-191

[554] Yoshio Ohno. Japanese TEX Users Group. TUGboat, 7(3):191–192, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16news.pdf.

Beeton:TB7-3-192

[555] Barbara Beeton. TEX in Europe, Summer 1986. TUGboat, 7(3):192–193, October 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-3/tb16news.pdf.

Anonymous:TB7-3-193

[556] Anonymous. Contents: Proceedings of “TEX for Scientific Documentation”, Strasbourg, France, June 19–21, 1986. TUGboat, 7(3):193–194, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16news.pdf.

Lachmann:TB7-3-194

[557] Susanne Lachmann. The Tiger-TEX Workstation. TUGboat, 7(3):194–195, October 1986. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb07-3/tb16lach.pdf.

Anonymous:TB7-3-196

[558] Anonymous. Institutional members. TUGboat, 7(3):196–197, October

1986. ISSN 0896-3207. URL https://tug.org/TUGboat/tb07-3/
tb16instmem.pdf.

Anonymous:TB7-3-198

[559] Anonymous. Index of advertisers. TUGboat, 7(3):198, October 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-3/tb16ads.
pdf.

Anonymous:TB7-3-199

[560] Anonymous. TUG membership application. TUGboat, 7(3):199–200, October 1986. ISSN 0896-3207.

Anonymous:TB7-3-201

[561] Anonymous. TEX order form. TUGboat, 7(3):201–202, October 1986. ISSN

0896-3207. URL https://tug.org/TUGboat/tb07-3/tb16news.
pdf.

Anonymous:1986:TJO

[562] Anonymous. TUG — job opening, technical assistant. TUGboat, 7(3):c3, October 1986. CODEN ???? ISSN 0896-3207.

Anonymous:TB7-3-cover-3

[563] Anonymous. TUG — Job opening, Technical assistant. TUGboat, 7(3):Cover 3, October 1986. ISSN 0896-3207.

Anonymous:1987:TPa

[564] Anonymous. Title page. TUGboat, 8(1):1, April 1987. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17titlepage.pdf.

Anonymous:1987:EIa

[565] Anonymous. Editorial information. TUGboat, 8(1):2, April 1987. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17edit.pdf.

Anonymous:TB8-1-3

[566] Anonymous. Addresses. TUGboat, 8(1):3–4, April 1987. ISSN 0896-3207.

Anonymous:TB8-1-5

[567] Anonymous. Acknowledgement of contributions. TUGboat, 8(1):5, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17gendel.pdf.

Childs:TB8-1-5

[568] Bart Childs. President’s message. TUGboat, 8(1):5, April 1987. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb08-1/tb17gendel.
pdf.

Anonymous:TB8-1-6

[569] Anonymous. Donald E. Knuth Scholarship. TUGboat, 8(1):6, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17gendel.pdf.

Knuth:TB8-1-6

[570] Donald Knuth. It happened: announcement of TEX2.1. TUGboat, 8(1):6, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17gendel.pdf.

Krick:TB8-1-7

[571] Linda Krick. Attention, Australian TUG members. TUGboat, 8(1):7, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17gendel.pdf.

Wolf:TB8-1-7

[572] Kurt Bernardo Wolf. News of TEX users in Mexico. TUGboat, 8(1):7, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17gendel.pdf.

Siegman:TB8-1-8

[573] Tony Siegman. Book publishing using TEX. TUGboat, 8(1):8–11, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17siegman.pdf.

Lamport:TB8-1-12

[574] Leslie Lamport. TEX output for the future. TUGboat, 8(1):12, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17lamp.pdf.

Levy:TB8-1-12

[575] Silvio Levy. WEB adapted to C, another approach. TUGboat, 8(1):12–13, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17levy.pdf.

Knuth:TB8-1-14

[576] Donald Knuth and Pierre MacKay. Mixing right-to-left texts with left-to-right texts. TUGboat, 8(1):14–25, April 1987. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb08-1/tb17knutmix.pdf.

Tobin:TB8-1-26

[577] Georgia K. M. Tobin. Some empirical observations on design. TUGboat, 8(1):26–28, April 1987. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb08-1/tb17tobin.pdf.

Billawala:TB8-1-29

[578] Neenie Billawala. Write-white printing engines and tuning fonts with - . TUGboat, 8(1):29–32, April 1987. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb08-1/tb17billawala.pdf.

Beeton:TB8-1-33

[579] Barbara Beeton. mode def settings for various TEX output devices. TUGboat, 8(1):33, April 1987. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb08-1/tb17beet.pdf.

Anonymous:TB8-1-34

[580] Anonymous. A new TEX-based book typesetting package for the Macintosh. TUGboat, 8(1):34, April 1987. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb08-1/tb17output.pdf.

Anonymous:TB8-1-34-2

[581] Anonymous. Charts: TEX output devices. TUGboat, 8(1):34–40, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17output.pdf.

Beebe:TB8-1-41

[582] Nelson H. F. Beebe. Public domain TEX DVI driver family. TUGboat, 8(1):41– 42, April 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-1/tb17beebe.pdf.

Trevorrow:TB8-1-42

[583] Andrew Trevorrow. DVItoVDU 1.7 and PSPRINT 1.1. TUGboat, 8(1):42– 45, April 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-1/tb17trevorrow.pdf.

Anonymous:TB8-1-45

[584] Anonymous. Benson-Varian 9211 looking for a home. TUGboat, 8(1): 45, April 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-1/tb17trevorrow.pdf.

Brown:TB8-1-46

[585] Malcolm Brown. Notes from TEXhax. TUGboat, 8(1):46, April 1987. ISSN

0896-3207. URL https://tug.org/TUGboat/tb08-1/tb17site.
pdf.

Fuss:TB8-1-46

[586] Edgar Fuß. TEX&Co. on the ST, Part 2. TUGboat, 8(1):46–47, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17site.pdf.

Childs:TB8-1-47

[587] Bart Childs. Data General distribution news. TUGboat, 8(1):47, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17site.pdf.

Lillqvist:TB8-1-47

[588] Tor Lillqvist. TEX82 and 84 for the HP1000 A-series. TUGboat, 8(1):47–48, April 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-1/tb17site.pdf.

Crawford:TB8-1-48

[589] John Crawford. Prime 50 series site report. TUGboat, 8(1):48, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17site.pdf.

Fox:TB8-1-49

[590] Jim Fox. Multiple, independent marks. TUGboat, 8(1):49–53, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17fox.pdf.

Garavelli:TB8-1-53

[591] John S. Garavelli. Form letter macros. TUGboat, 8(1):53, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17garavelli.pdf.

Bell:TB8-1-54

[592] Edwin V. Bell, II. AutoLetter: A TEX form letter procedure. TUGboat, 8(1):54– 57, April 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-1/tb17bell.pdf.

Yap:TB8-1-58

[593] Ken Yap. Contents of L^{A}TEX style collection as of 15th February 1987. TUGboat, 8(1):58–59, April 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-1/tb17yap.pdf.

McKinstry:TB8-1-60

[594] Graeme McKinstry. Form letters. TUGboat, 8(1):60–61, April 1987. ISSN

0896-3207. URL https://tug.org/TUGboat/tb08-1/tb17mckinstry.
pdf.

Zocchi:TB8-1-62

[595] Maurizio Zocchi. L^{A}TEX’s Index Processing. TUGboat, 8(1):62, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17zocchi.pdf.

Wonneberger:TB8-1-63

[596] Reinhard Wonneberger. Typesetting ‘Normaltext’. TUGboat, 8(1):63–72, April 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-1/tb17wonneberger.pdf.

Hoenig:TB8-1-73

[597] Alan Hoenig. TEX does windows. TUGboat, 8(1):73, April 1987. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb08-1/tb17knutsat.
pdf.

Knuth:TB8-1-73

[598] Donald Knuth. Problem for a Saturday afternoon. TUGboat, 8(1):73, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17knutsat.pdf.

Celoni:TB8-1-74

[599] James R. Celoni, S.J. Ancient non-Roman languages. TUGboat, 8(1):74, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17queries.pdf.

Cutter:TB8-1-74

[600] Lawrence D. Cutter. Line numbering. TUGboat, 8(1):74, April 1987. ISSN

0896-3207. URL https://tug.org/TUGboat/tb08-1/tb17queries.
pdf.

Dyck:TB8-1-74

[601] Jennifer Dyck. APA style in L^{A}TEX. TUGboat, 8(1):74, April 1987. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb08-1/tb17queries.
pdf.

Hawkins:TB8-1-74

[602] Judy Hawkins. Chemical formulae. TUGboat, 8(1):74, April 1987. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb08-1/tb17queries.
pdf.

Michailovsky:TB8-1-74

[603] Boyd Michailovsky. International Phonetic Alphabet. TUGboat, 8(1):74, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17queries.pdf.

Horstman:TB8-1-75

[604] Helen S. Horstman. Printing out selected pages. TUGboat, 8(1):75, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17queries.pdf.

Nash:TB8-1-75

[605] David Nash. Dictionary formatting. TUGboat, 8(1):75, April 1987. ISSN

0896-3207. URL https://tug.org/TUGboat/tb08-1/tb17queries.
pdf.

Ryan:TB8-1-75

[606] Raymond A. Ryan. Wanted: Help for beginners. TUGboat, 8(1):75, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17letters.pdf.

Thulin:TB8-1-76

[607] Anders Thulin. More hyphenation exceptions. TUGboat, 8(1):76, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17letters.pdf.

McWorter:TB8-1-76

[608] William A. McWorter. New fonts for mathematics? TUGboat, 8(1):76– 77, April 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-1/tb17letters.pdf.

Anonymous:TB8-1-77

[609] Anonymous. Workshop on Font Design Systems (Sophia-Antipolis, France, 18–19 May 1987). TUGboat, 8(1):77, April 1987. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb08-1/tb17news.pdf.

Anonymous:TB8-1-77-2

[610] Anonymous. Special Spring Quarter course (Stanford, Donald Knuth): TEX: The Program: A case study in software design. TUGboat, 8(1):77, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17news.pdf.

Anonymous:TB8-1-78

[611] Anonymous. Call for Papers: Electronic Publishing, Document Manipulation and Typography (Nice, France, 20–22 April 1988). TUGboat, 8(1):78, April

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17news.pdf.

Anonymous:TB8-1-79

[612] Anonymous. Calendar. TUGboat, 8(1):79–80, April 1987. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb08-1/tb17calendar.pdf.

Anonymous:TB8-1-81

[613] Anonymous. Institutional members. TUGboat, 8(1):81–82, April 1987. ISSN

0896-3207. URL https://tug.org/TUGboat/tb08-1/tb17instmem.
pdf.

Anonymous:TB8-1-83

[614] Anonymous. TUG membership application. TUGboat, 8(1):83–84, April 1987.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB8-1-85

[615] Anonymous. TEX order form. TUGboat, 8(1):85–86, April 1987. ISSN 0896-

Anonymous:1987:A

[616] Anonymous. Advertisements. TUGboat, 8(1):87–96, April 1987. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-1/
tb17ads.pdf.

Anonymous:1987:IAa

[617] Anonymous. Index of advertisers. TUGboat, 8(1):c2, April 1987. CODEN ???? ISSN 0896-3207.

Anonymous:TB8-1-cover-3

[618] Anonymous. Index of advertisers. TUGboat, 8(1):Cover 3, April 1987. ISSN 0896-3207.

Anonymous:TB8-1-cover-3-2

[619] Anonymous. TUG — Job opening, Technical assistant. TUGboat, 8(1):Cover 3, April 1987. ISSN 0896-3207.

Anonymous:1987:TPb

[620] Anonymous. Title page. TUGboat, 8(2):97, July 1987. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18titlepage.pdf.

Anonymous:1987:EIb

[621] Anonymous. Editorial information. TUGboat, 8(2):98, July 1987. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18edit.pdf.

Anonymous:TB8-2-99

[622] Anonymous. Addresses. TUGboat, 8(2):99–100, July 1987. ISSN 0896-3207.

Childs:TB8-2-101

[623] Bart Childs. From the President. TUGboat, 8(2):101, July 1987. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb08-2/tb18gendel.
pdf.

Koren:TB8-2-101

[624] Gideon Koren. How to improve the chances for acceptance of your scientific paper. TUGboat, 8(2):101–102, July 1987. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb08-2/tb18gendel.pdf.

Alexander:TB8-2-102

[625] James Alexander. TIb: a reference setting package, update. TUGboat, 8(2): 102, July 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-2/tb18software.pdf.

Rezende:TB8-2-102

[626] Pedro de Rezende. Portuguese hyphenation table for TEX. TUGboat, 8(2):102,

July 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-
2/tb18software.pdf.

Ferguson:TB8-2-102

[627] Michael Ferguson. A (hopefully) final extension of multilingual TÊX. TUGboat, 8(2):102–103, July 1987. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb08-2/tb18software.pdf.

Saito:TB8-2-103

[628] Yasuki Saito. Report on JTEX: A Japanese TEX. TUGboat, 8(2):103–116, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18saito.pdf.

Sewell:TB8-2-117

[629] E. Wayne Sewell. Multiple changefiles: the adventure continues. TUGboat, 8(2):117–118, July 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-2/tb18sewellch.pdf.

Sewell:TB8-2-118

[630] E. Wayne Sewell. How to MANGLE your software: the WEB system for Modula-2. TUGboat, 8(2):118–122, July 1987. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb08-2/tb18sewellmangle.pdf.

Sewell:TB8-2-123

[631] E. Wayne Sewell. The SCANTEX processor. TUGboat, 8(2):123–128, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18sewellscantex.pdf.

Gourlay:TB8-2-128

[632] John S. Gourlay. Blacker Thoughts. TUGboat, 8(2):128–129, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18gourlay.pdf.

Sauter:TB8-2-129

[633] John Sauter. Updated Computer Modern fonts for the LN03. TUGboat, 8 (2):129–130, July 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-2/tb18sauter.pdf.

Crawford:TB8-2-131

[634] John Crawford. MFtool: A script-driven processing facility. TUGboat, 8(2):131, July 1987. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb08-2/tb18crawford.pdf.

Beeton:TB8-2-132

[635] Barbara Beeton. Update: mode def settings for TEX output devices. TUGboat, 8(2):132–134, July 1987. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb08-2/tb18beet.pdf.

Knuth:TB8-2-135

[636] Donald E. Knuth. Fonts for digital halftones. TUGboat, 8(2):135–160, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18knut.pdf.

McGaffey:TB8-2-161

[637] Robert W. McGaffey. The ideal TEX driver. TUGboat, 8(2):161–163, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18mcgaffey.pdf.

Hosek:TB8-2-163

[638] Don Hosek. TEX output devices (with charts). TUGboat, 8(2):163–170, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18hosek.pdf.

Dunn:TB8-2-171

[639] Alec Dunn. Using PostScript with TEX. TUGboat, 8(2):171–173, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18dunn.pdf.

Vanderburg:TB8-2-174

[640] Glenn L. Vanderburg. TEX-L access for Bitnet users. TUGboat, 8(2):174, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18site.pdf.

Mann:TB8-2-174

[641] Laurie Mann. TEX and training: A case study. TUGboat, 8(2):174–176, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18site.pdf.

Childs:TB8-2-176

[642] Bart Childs. Data General site report. TUGboat, 8(2):176–177, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18site.pdf.

Clark:TB8-2-177

[643] Adrian F. Clark. Enhancements to TEX on the VAX. TUGboat, 8(2):177, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18site.pdf.

Guenther:TB8-2-178

[644] Dean Guenther. IBM VM/CMS site report. TUGboat, 8(2):178, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18site.pdf.

Pierson:TB8-2-178

[645] Gil Pierson. SAS merged with TEX. TUGboat, 8(2):178–179, July 1987. ISSN

0896-3207. URL https://tug.org/TUGboat/tb08-2/tb18site.
pdf.

MacKay:TB8-2-179

[646] Pierre MacKay. Unix TEX site report. TUGboat, 8(2):179–181, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18site.pdf.

Kohlmayr:TB8-2-181

[647] Gerhard F. Kohlmayr. A bug in TEXTURES v0.95 prerelease. TUGboat, 8(2): 181, July 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-2/tb18site.pdf.

Smith:TB8-2-181

[648] Barry Smith. Work-around for an ImageWriter problem affecting TEXTURES output. TUGboat, 8(2):181, July 1987. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb08-2/tb18site.pdf.

Hoenig:TB8-2-182

[649] Alan Hoenig and Mitch Pfeffer. Update: Real typesetting from your PC. TUGboat, 8(2):182, July 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-2/tb18site.pdf.

Appelt:TB8-2-182

[650] Wolfgang Appelt. Macros with keyword parameters. TUGboat, 8(2):182–184,

July 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-
2/tb18appelt.pdf.

Kabelschacht:TB8-2-184

[651] Alois Kabelschacht. \expandafter in conditionals; a generalization of PLAIN’s \loop. TUGboat, 8(2):184–185, July 1987. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb08-2/tb18kabel.pdf.

Barnhart:TB8-2-185

[652] Elizabeth Barnhart. TEX in the commercial environment — multi-column output. TUGboat, 8(2):185–189, July 1987. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb08-2/tb18barnhart.pdf.

LeHardy:TB8-2-190

[653] Charles LeHardy. Diglot typesetting. TUGboat, 8(2):190–192, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18lehardy.pdf.

Bruggemann-Klein:TB8-2-193

[654] Anne Brüggemann-Klein. First line special handling with TEX. TUGboat, 8(2):193–197, July 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-2/tb18bruggemann.pdf.

Yap:TB8-2-198

[655] Ken Yap. Contents of L^{A}TEX style collection as of 15th May 1987. TUGboat, 8(2):198–200, July 1987. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb08-2/tb18yap.pdf.

Aurbach:TB8-2-201

[656] Richard Aurbach. Automated index generation for L^{A}TEX. TUGboat, 8(2):201– 209, July 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-2/tb18aurbach.pdf.

Knuth:TB8-2-210

[657] Donald Knuth. Saturday morning problem — solution. TUGboat, 8(2):210,

July 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-
2/tb18knutsat.pdf.

Hoenig:TB8-2-211

[658] Alan Hoenig. TEX does windows — conclusion. TUGboat, 8(2):211–215, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18hoenig.pdf.

Alexander:TB8-2-216

[659] James Alexander. Time line macro. TUGboat, 8(2):216, July 1987. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb08-2/tb18queries.
pdf.

Fox:TB8-2-216

[660] Jim Fox. Comment on “TEX does windows”. TUGboat, 8(2):216, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18fox.pdf.

Black:TB8-2-217

[661] Mike Black. Using the Windows environment. TUGboat, 8(2):217, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18queries.pdf.

Knuth:TB8-2-217

[662] Donald Knuth. Reply: Printing out selected pages. TUGboat, 8(2):217, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18queries.pdf.

Anonymous:TB8-2-218

[663] Anonymous. TUG Annual Meeting, University of Washington, August 24–26, 1987. TUGboat, 8(2):218, July 1987. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb08-2/tb18news.pdf.

Anonymous:TB8-2-219

[664] Anonymous. Calendar. TUGboat, 8(2):219–220, July 1987. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb08-2/tb18news.pdf.

Anonymous:TB8-2-220

[665] Anonymous. PROTEXT IV, Boston, Mass., 22–24 October 1987. TUGboat, 8(2):220, July 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-2/tb18news.pdf.

Anonymous:TB8-2-221

[666] Anonymous. 6th German TEX meeting, University of Münster, 8–9 October 1987. TUGboat, 8(2):221, July 1987. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb08-2/tb18news.pdf.

Price:TB8-2-221

[667] Lynne A. Price. SGML and TEX. TUGboat, 8(2):221–225, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18price.pdf.

Whidden:TB8-2-225

[668] Samuel B. Whidden. Treasurer’s report. TUGboat, 8(2):225–227, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18treas.pdf.

Anonymous:TB8-2-228

[669] Anonymous. Institutional members. TUGboat, 8(2):228–230, July

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-2/
tb18instmem.pdf.

Anonymous:TB8-2-231

[670] Anonymous. TUG membership application; TEX order form. TUGboat, 8(2): 231–232, July 1987. ISSN 0896-3207. URL https://tug.org/join.

html.

Anonymous:TB8-2-237

[671] Anonymous. Index to sample output from various devices. TUGboat, 8(2): 237, July 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-2/tb18indexoutput.pdf.

Anonymous:1987:IAb

[672] Anonymous. Index of advertisers. TUGboat, 8(2):238–252, July 1987. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-
2/tb18ads.pdf.

Anonymous:TB8-2-252

[673] Anonymous. Index of advertisers. TUGboat, 8(2):252, July 1987. ISSN 0896-3207.

Anonymous:1987:TPc

[674] Anonymous. Title page. TUGboat, 8(3):254, November 1987. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19titlepage.pdf.

Anonymous:TB8-3-255

[675] Anonymous. Addresses. TUGboat, 8(3):255–256, November 1987. ISSN

0896-3207. URL https://tug.org/TUGboat/tb08-3/tb19edit.
pdf.

Childs:TB8-3-257

[676] Bart Childs. From the President. TUGboat, 8(3):257, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19gendel.pdf.

Thedford:TB8-3-258

[677] Rilla Thedford. The Volunteer Tree. TUGboat, 8(3):258, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19gendel.pdf.

Marle:TB8-3-259

[678] Charles-Michel Marle. Book publishing using TEX. TUGboat, 8(3):259–260, November 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-3/tb19gendel.pdf.

Mann:TB8-3-260

[679] Laurie Mann. TEX training, etc. — A TUG meeting trip report. TUGboat, 8(3): 260–262, November 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-3/tb19gendel.pdf.

Beeton:TB8-3-262

[680] Barbara Beeton. From the Editor. TUGboat, 8(3):262, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19gendel.pdf.

Childs:TB8-3-262

[681] Bart Childs. Proposed minimum standards for TEX distributions. TUGboat, 8 (3):262–263, November 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-3/tb19childs.pdf.

Reid:TB8-3-264

[682] Thomas J. Reid. TANGLE modification causes problems in and PK files. TUGboat, 8(3):264–265, November 1987. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb08-3/tb19reid-tangle.pdf.

Anonymous:TB8-3-266

[683] Anonymous. Hyphenation exception log. TUGboat, 8(3):266–267, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19hyf.pdf.

Henderson:TB8-3-268

[684] Doug Henderson. Update: mode def settings for TEX output devices. TUGboat, 8(3):268–270, November 1987. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb08-3/tb19henderson.pdf.

Clark:TB8-3-270

[685] Adrian F. Clark. Halftone output from TEX. TUGboat, 8(3):270–274, Novem-

ber 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-
3/tb19clark.pdf.

Hohti:TB8-3-275

[686] Aarno Hohti and Okko Kanerva. Generating an APL font. TUGboat, 8(3): 275–278, November 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-3/tb19hohti.pdf.

Hosek:TB8-3-279

[687] Don Hosek. TEX output devices (with charts). TUGboat, 8(3):279–286, November 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-3/tb19hosek.pdf.

Reid:TB8-3-287

[688] Thomas J. Reid. DVI driver considerations for high-volume printing systems. TUGboat, 8(3):287–291, November 1987. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb08-3/tb19reid-dvi.pdf.

Vanderburg:TB8-3-291

[689] Glenn L. Vanderburg and Thomas J. Reid. \special issues. TUGboat, 8(3): 291–300, November 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-3/tb19vanderburg.pdf.

Toal:TB8-3-301

[690] Graham Toal. TEX information for users in the U.K. TUGboat, 8(3):301, November 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-3/tb19site.pdf.

Brown:TB8-3-301

[691] Malcolm Brown. TEXhax Notes. TUGboat, 8(3):301–303, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19site.pdf.

Childs:TB8-3-304

[692] Bart Childs. Data General site report. TUGboat, 8(3):304, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19site.pdf.

Guntermann:TB8-3-304

[693] Klaus Guntermann. Atari ST site report. TUGboat, 8(3):304, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19site.pdf.

Lammarsch:TB8-3-304

[694] Joachim Lammarsch. 6th Meeting of the “TEX-Interessenten” in Germany. TUGboat, 8(3):304, November 1987. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb08-3/tb19site.pdf.

Pfeffer:TB8-3-305

[695] Mitch Pfeffer and Alan Hoenig. Running TEX on a 386-based computer: Twice as fast as an AT. TUGboat, 8(3):305–306, November 1987. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb08-3/tb19pfeffer.pdf.

Flynn:TB8-3-306

[696] Peter Flynn. Request for contributions to a new publication. TUGboat, 8 (3):306, November 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-3/tb19flynn.pdf.

Thiele:TB8-3-307

[697] Christina Thiele. What constitutes a well-documented macro? TUGboat, 8(3): 307–308, November 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-3/tb19thiele.pdf.

Knuth:TB8-3-309

[698] Donald E. Knuth. Macros for Jill. TUGboat, 8(3):309–314, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19knut.pdf.

Reid:TB8-3-315

[699] Thomas J. Reid. Floating figures at the right, and Some random text for testing. TUGboat, 8(3):315–320, November 1987. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb08-3/tb19reid-float.pdf.

Yap:TB8-3-320

[700] Ken Yap. Contents of L^{A}TEX style collection as of 15th September 1987. TUGboat, 8(3):320–323, November 1987. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb08-3/tb19yap.pdf.

Damrau:TB8-3-323

[701] Jackie Damrau. The L^{A}TEX user’s column. TUGboat, 8(3):323–326, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19damrau.pdf.

Boes:TB8-3-326

[702] Jeffery Boes. Reply: Printing out selected pages. TUGboat, 8(3):326, Novem-

ber 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-
3/tb19boes.pdf.

Lipp:TB8-3-326

[703] Stephen C. Lipp. Title formatting macro wanted. TUGboat, 8(3):326–327, November 1987. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb08-3/tb19lipp.pdf.

Anonymous:TB8-3-328

[704] Anonymous. Calendar. TUGboat, 8(3):328, November 1987. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb08-3/tb19calendar.pdf.

Anonymous:TB8-3-329

[705] Anonymous. Exeter University: TEX88, 18–20 July 1988. TUGboat, 8(3): 329–330, November 1987. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb08-3/tb19calendar.pdf.

Anonymous:TB8-3-331

[706] Anonymous. Institutional members. TUGboat, 8(3):331–332, November

1987. ISSN 0896-3207. URL https://tug.org/TUGboat/tb08-3/
tb19instmem.pdf.

Anonymous:TB8-3-333

[707] Anonymous. TUG membership application. TUGboat, 8(3):333–334, November 1987. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB8-3-352

[708] Anonymous. Index of advertisers. TUGboat, 8(3):352, November 1987. ISSN

0896-3207. URL https://tug.org/TUGboat/tb08-3/tb19ads.
pdf.

Anonymous:1987:CPT

[709] Anonymous. Call for papers: TUG Annual Meeting, Montréal, 22–24 August 1988. TUGboat, 8(3):c3, November 1987. CODEN ???? ISSN 0896-3207.

Anonymous:1988:TPa

[710] Anonymous. Title page. TUGboat, 9(1):1, April 1988. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20titlepage.pdf.

Anonymous:1988:EIa

[711] Anonymous. Editorial information. TUGboat, 9(1):2, April 1988. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20edit.pdf.

Anonymous:TB9-1-3

[712] Anonymous. Addresses. TUGboat, 9(1):3–4, April 1988. ISSN 0896-3207.

Childs:TB9-1-5

[713] Bart Childs. From the President. TUGboat, 9(1):5, April 1988. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb09-1/tb20gendel.
pdf.

Wittbecker:TB9-1-5

[714] Alan Wittbecker. TUGboat docks at fire station. TUGboat, 9(1):5, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20gendel.pdf.

Anonymous:TB9-1-6

[715] Anonymous. Donald E. Knuth Scholarship. TUGboat, 9(1):6, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20gendel.pdf.

Beeton:TB9-1-6

[716] Barbara Beeton. Notes from the Editor. TUGboat, 9(1):6–7, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20gendel.pdf.

Lamport:TB9-1-8

[717] Leslie Lamport. Document Production: Visual or Logical? TUGboat, 9(1): 8–10, April 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-1/tb20lamport.pdf.

Breitenlohner:TB9-1-11

[718] Peter Breitenlohner. Still another aspect of multiple change files: The PATCH processor. TUGboat, 9(1):11–12, April 1988. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb09-1/tb20breitenlohner.pdf.

MacKay:TB9-1-12

[719] Pierre A. MacKay. Turkish hyphenations for TEX. TUGboat, 9(1):12–14, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20mackay.pdf.

Tobin:TB9-1-15

[720] Georgia K. M. Tobin. The ABC’s of special effects. TUGboat, 9(1):15– 18, April 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-1/tb20tobin.pdf.

Messer:TB9-1-19

[721] Robert Messer. Blackboard Bold. TUGboat, 9(1):19–20, April 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-1/tb20messer.
pdf.

Levy:TB9-1-20

[722] Silvio Levy. Using Greek fonts with TEX. TUGboat, 9(1):20–24, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20levy.pdf.

Hosek:TB9-1-25

[723] Don Hosek. TEX output devices (with charts). TUGboat, 9(1):25–33, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20output.pdf.

Brown:TB9-1-34

[724] Marcus Brown. An ASCII previewer for TEX. TUGboat, 9(1):34–36, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20brown.pdf.

Hosek:TB9-1-37

[725] Don Hosek. A screen previewer for VM/CMS. TUGboat, 9(1):37, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20hosek.pdf.

Dunne:TB9-1-37

[726] Shane Dunne. Why TEX should NOT output PostScript — yet. TUGboat, 9(1):37–39, April 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-1/tb20dunne.pdf.

Anonymous:TB9-1-39

[727] Anonymous. Index to sample output from various devices. TUGboat, 9(1): 39, April 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-1/tb20devoutindex.pdf.

Rokicki:TB9-1-40

[728] Tom Rokicki. The Commodore Amiga: A magic TEX machine. TUGboat, 9(1):40–41, April 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-1/tb20site.pdf.

Childs:TB9-1-41

[729] Bart Childs. Data General site report. TUGboat, 9(1):41, April 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-1/tb20site.
pdf.

Anonymous:TB9-1-42

[730] Anonymous. Fujitsu ports TEX 1.0 onto M-series. TUGboat, 9(1):42, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20site.pdf.

Guenther:TB9-1-42

[731] Dean Guenther. IBM VM/CMS site report. TUGboat, 9(1):42, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20site.pdf.

MacKay:TB9-1-42

[732] Pierre MacKay. Unix TEX site report. TUGboat, 9(1):42–44, April 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-1/tb20site.
pdf.

Morgan:TB9-1-44

[733] Tim Morgan. TEX to C converter. TUGboat, 9(1):44–45, April 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-1/tb20site.
pdf.

Kellerman:TB9-1-45

[734] David Kellerman. VAX/VMS site report. TUGboat, 9(1):45, April 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-1/tb20site.
pdf.

Hoenig:TB9-1-46

[735] Alan Hoenig and Mitch Pfeffer. Writers’ tools I: PC spelling and grammar checkers. TUGboat, 9(1):46–47, April 1988. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb09-1/tb20hoenig.pdf.

Hoenig:TB9-1-47

[736] Alan Hoenig and Mitch Pfeffer. Grapevine reports of inexpensive versions of TEX. TUGboat, 9(1):47–48, April 1988. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb09-1/tb20hoenig.pdf.

Kinch:TB9-1-48

[737] Richard J. Kinch and Jennifer L. Vollbrecht. TurboTEX: A new port in C for Unix and MS-DOS. TUGboat, 9(1):48–52, April 1988. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb09-1/tb20kinch.pdf.

Crisanti:TB9-1-52

[738] Ester Crisanti, Alberto Formigoni, and Paco La Bruna. easyTEX. TUGboat, 9(1):52–56, April 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-1/tb20crisanti.pdf.

Bechtolsheim:TB9-1-57

[739] Stephan v. Bechtolsheim. A tutorial on \expandafter. TUGboat, 9(1):57– 61, April 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-1/tb20bechtolsheim.pdf.

Walker:TB9-1-61

[740] James W. Walker. Macros for outlining. TUGboat, 9(1):61–64, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20walker.pdf.

Hendrickson:TB9-1-64

[741] Amy Hendrickson. A macro writing tool: Generating new definitions. TUGboat, 9(1):64, April 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-1/tb20hendrickson.pdf.

Gariepy:TB9-1-65

[742] Alonzo Gariepy. French in TEX. TUGboat, 9(1):65–69, April 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-1/tb20gariepy.
pdf.

Partl:TB9-1-70

[743] Hubert Partl. German TEX. TUGboat, 9(1):70–72, April 1988. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb09-1/tb20partl.pdf.

Yap:TB9-1-73

[744] Ken Yap. Contents of L^{A}TEX style collection as of 4th February 1988. TUGboat, 9(1):73–75, April 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-1/tb20yap.pdf.

Damrau:TB9-1-76

[745] Jackie Damrau. The L^{A}TEX user’s column. TUGboat, 9(1):76–77, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20damrau.pdf.

McPherson:TB9-1-78

[746] Kent McPherson. Page layout in L^{A}TEX. TUGboat, 9(1):78–82, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20mcpherson.pdf.

Bartlett:TB9-1-83

[747] Frederick H. Bartlett. Automatic page balancing macros wanted; Replies: Inverted pyramidal titles; Logarithmic time scales. TUGboat, 9(1):83–85, April

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-1/
tb20bartlett.pdf.

Anonymous:TB9-1-86

[748] Anonymous. Calendar. TUGboat, 9(1):86, April 1988. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb09-1/tb20news.pdf.

Anonymous:TB9-1-87

[749] Anonymous. TUG 1988 Annual Meeting, Montréal, August 22–24. TUGboat, 9(1):87, April 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-1/tb20news.pdf.

Anonymous:TB9-1-87-2

[750] Anonymous. Videotapes of Knuth’s software course based on TEX: The Program. TUGboat, 9(1):87, April 1988. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb09-1/tb20news.pdf.

Anonymous:TB9-1-88

[751] Anonymous. Institutional members. TUGboat, 9(1):88–90, April 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-1/tb20instmem.
pdf.

Anonymous:TB9-1-91

[752] Anonymous. TUG membership application. TUGboat, 9(1):91–93, April 1988.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB9-1-95

[753] Anonymous. TEX order form. TUGboat, 9(1):95–96, April 1988. ISSN 0896-3207.

Anonymous:TB9-1-112

[754] Anonymous. Index of advertisers. TUGboat, 9(1):112, April 1988. ISSN 0896-3207.

Anonymous:1988:FC

[755] Anonymous. Front cover. TUGboat, 9(1):c1, April 1988. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-1/cover1.
pdf.

Anonymous:1988:TPb

[756] Anonymous. Title page. TUGboat, 9(2):113, August 1988. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21titlepage.pdf.

Anonymous:1988:EIb

[757] Anonymous. Editorial information. TUGboat, 9(2):114, August 1988. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-
2/tb21edit.pdf.

Anonymous:TB9-2-115

[758] Anonymous. Addresses. TUGboat, 9(2):115–116, August 1988. ISSN 0896-3207.

Anonymous:TB9-2-117

[759] Anonymous. Donald Knuth awarded Franklin Medal. TUGboat, 9(2):117, Au-

gust 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-
2/tb21gendel.pdf.

Childs:TB9-2-117

[760] Bart Childs. From the President. TUGboat, 9(2):117, August 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-2/tb21gendel.
pdf.

Goucher:TB9-2-117

[761] Ray Goucher. Los Alamos sets new membership record. TUGboat, 9(2): 117, August 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-2/tb21gendel.pdf.

Barnhart:TB9-2-118

[762] Elizabeth Barnhart. TEX in the publishing environment: A survey of production/commercial users. TUGboat, 9(2):118–120, August 1988. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb09-2/tb21gendel.
pdf.

Beeton:TB9-2-121

[763] Barbara Beeton. Editorial comments. TUGboat, 9(2):121, August 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-2/tb21beet-
new.pdf.

Beeton:TB9-2-121-2

[764] Barbara Beeton. New version(s) of TEX and . TUGboat, 9(2):121– 122, August 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-2/tb21beet-new.pdf.

Childs:TB9-2-122

[765] Bart Childs. 64-bit TEX. TUGboat, 9(2):122, August 1988. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb09-2/tb21childs-64.pdf.

Lichtenwalder:TB9-2-123

[766] Klaus Lichtenwalder. Porting TEX to C. TUGboat, 9(2):123–124, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21lichtenwalder.pdf.

Kennedy:TB9-2-124

[767] David Kennedy. TEX adapted to CWEB. TUGboat, 9(2):124–125, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21kennedy.pdf.

Vanderburg:TB9-2-125

[768] Glenn Vanderburg. Some useful variations of standard fonts. TUGboat, 9(2): 125, August 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-2/tb21vanderburg.pdf.

Tobin:TB9-2-126

[769] Georgia K. M. Tobin. Designing for low-res devices. TUGboat, 9(2):126– 128, August 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-2/tb21tobin.pdf.

Childs:TB9-2-129

[770] Bart Childs. TEX fonts and suggested magnifications. TUGboat, 9(2):129– 131, August 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-2/tb21childs-fonts.pdf.

Wujastyk:TB9-2-131

[771] Dominik Wujastyk. The many faces of TEX: A survey of digital fonts. TUGboat, 9(2):131–151, August 1988. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb09-2/tb21wujastyk.pdf.

Anonymous:TB9-2-151

[772] Anonymous. Index to sample output from various devices. TUGboat, 9(2): 151, August 1988. ISSN 0896-3207.

Knuth:TB9-2-152

[773] Donald E. Knuth. A Punk Meta-Font. TUGboat, 9(2):152–168, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21knut.pdf.

Hosek:TB9-2-169

[774] Don Hosek. TEX output devices (with charts). TUGboat, 9(2):169–177, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21output.pdf.

Dunne:TB9-2-178

[775] Shane Dunne. Why TEX should NOT output PostScript — yet: Addendum. TUGboat, 9(2):178, August 1988. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb09-2/tb21dunne.pdf.

Wolfe:TB9-2-178

[776] Warren Wolfe. ASCII Preview with vuTEX. TUGboat, 9(2):178–180, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21wolfe.pdf.

Childs:TB9-2-181

[777] Bart Childs. TEX on the Cray. TUGboat, 9(2):181, August 1988. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb09-2/tb21site.pdf.

Childs:TB9-2-181-2

[778] Bart Childs. Data General site report. TUGboat, 9(2):181, August 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-2/tb21site.
pdf.

Smith:TB9-2-181

[779] Barry Smith. Macintosh site report. TUGboat, 9(2):181, August 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-2/tb21site.
pdf.

Beeton:TB9-2-182

[780] Barbara Beeton. Controlling <ctrl-M>; Ruling the depths. TUGboat, 9 (2):182–183, August 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-2/tb21beet-warn.pdf.

Breitenlohner:TB9-2-183

[781] Peter Breitenlohner. German TEX, a next step. TUGboat, 9(2):183–185, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21breitenlohner.pdf.

Wichura:TB9-2-186

[782] Michael J. Wichura. Some problems with the INRSTEX table making macros. TUGboat, 9(2):186–188, August 1988. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb09-2/tb21wichura-inrstex.pdf.

Haagen:TB9-2-189

[783] A. J. Van Haagen. Box plots and scatter plots with TEX macros. TUGboat, 9(2):189–192, August 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-2/tb21haagen.pdf.

Wichura:TB9-2-193

[784] Michael J. Wichura. PICTEX: Macros for drawing PICtures. TUGboat, 9 (2):193–197, August 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-2/tb21wichura-pictex.pdf.

Damrau:TB9-2-198

[785] Jackie Damrau. The L^{A}TEX user’s column. TUGboat, 9(2):198–199, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21damrau.pdf.

Yap:TB9-2-200

[786] Ken Yap. Contents of L^{A}TEX style collection as of 19th June 1988. TUGboat, 9(2):200–203, August 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-2/tb21yap.pdf.

Bechtolsheim:TB9-2-203

[787] Stephan v. Bechtolsheim. A note on processing parts with L^{A}TEX. TUGboat, 9(2):203–204, August 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-2/tb21bechtolsheim.pdf.

McPherson:TB9-2-204

[788] Kent McPherson. Page layout in L^{A}TEX: Erratum. TUGboat, 9(2):204, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21mcpherson.pdf.

Anonymous:TB9-2-205

[789] Anonymous. Calendar. TUGboat, 9(2):205, August 1988. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb09-2/tb21calendar.pdf.

Gaulle:TB9-2-206

[790] Bernard Gaulle. GUTenberg meeting report. TUGboat, 9(2):206–207, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21guten.pdf.

Anonymous:TB9-2-207

[791] Anonymous. Knuth Scholarship winner. TUGboat, 9(2):207, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21schol.pdf.

Anonymous:TB9-2-208

[792] Anonymous. Institutional members. TUGboat, 9(2):208–210, August

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-2/
tb21instmem.pdf.

Anonymous:TB9-2-211

[793] Anonymous. TUG membership application. TUGboat, 9(2):211–212, August

1988. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB9-2-226

[794] Anonymous. Index of advertisers. TUGboat, 9(2):226, August 1988. ISSN 0896-3207.

Anonymous:TB9-2-M

[795] Anonymous. TUG Membership List Supplement. TUGboat, 9(2):M–1–M–14, August 1988. ISSN 0896-3207.

Anonymous:1988:TPc

[796] Anonymous. Title page. TUGboat, 9(3):227, December 1988. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22titlepage.pdf.

Anonymous:TB9-3-229

[797] Anonymous. Addresses. TUGboat, 9(3):229–230, November 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-3/tb22edit.
pdf.

Childs:TB9-3-231

[798] Bart Childs. From the President. TUGboat, 9(3):231–232, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22gendel.pdf.

Anonymous:TB9-3-232

[799] Anonymous. Extra! Extra! TUGboat becomes a quarterly. TUGboat, 9(3):232, November 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-3/tb22gendel.pdf.

Anonymous:TB9-3-232-2

[800] Anonymous. Donald Knuth awarded Oxford degree. TUGboat, 9(3):232–233, November 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-3/tb22gendel.pdf.

Beeton:TB9-3-233

[801] Barbara Beeton. TEX and TUG go international — A trip report. TUGboat, 9 (3):233–235, November 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-3/tb22gendel.pdf.

McKinstry:TB9-3-236

[802] Graeme McKinstry. Some typesetting conventions. TUGboat, 9(3):236–238, November 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-3/tb22mckinstry.pdf.

Heidrich:TB9-3-238

[803] Klaus Heidrich. Software-ergonomics on the Atari ST. TUGboat, 9(3):238, November 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-3/tb22heidrich.pdf.

Anonymous:TB9-3-239

[804] Anonymous. Hyphenation exception log. TUGboat, 9(3):239–241, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22hyf.pdf.

Schrod:TB9-3-241

[805] Joachim Schrod. L^{A}TEX fonts and suggested magnifications. TUGboat, 9(3): 241–243, November 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-3/tb22schrod.pdf.

Henderson:TB9-3-244

[806] Doug Henderson. Mode def’s please. TUGboat, 9(3):244–245, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22henderson.pdf.

Wujastyk:TB9-3-246

[807] Dominik Wujastyk. Further faces. TUGboat, 9(3):246–251, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22wujastyk.pdf.

Hosek:TB9-3-251

[808] Don Hosek. TEX output devices (with charts). TUGboat, 9(3):251–260, November 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-3/tb22output.pdf.

Abbott:TB9-3-261

[809] Peter Abbott and Andrew Trevorrow. DVItoVDU 3.0 and PSPRINT 3.0. TUGboat, 9(3):261–262, November 1988. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb09-3/tb22abbott.pdf.

Abbott:TB9-3-263

[810] Peter Abbott. A UK-based TEX mail archive server. TUGboat, 9(3):263–264, November 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-3/tb22site.pdf.

Guntermann:TB9-3-264

[811] Klaus Guntermann. Atari ST site report. TUGboat, 9(3):264–265, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22site.pdf.

Childs:TB9-3-265

[812] Bart Childs. Data General site report. TUGboat, 9(3):265, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22site.pdf.

Guenther:TB9-3-265

[813] Dean Guenther. IBM VM/CMS site report. TUGboat, 9(3):265, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22site.pdf.

Platt:TB9-3-266

[814] Craig Platt. MVS site report. TUGboat, 9(3):266–267, November 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-3/tb22site.
pdf.

MacKay:TB9-3-267

[815] Pierre MacKay. Unix TEX site report. TUGboat, 9(3):267–269, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22site.pdf.

Pfeffer:TB9-3-269

[816] Mitch Pfeffer and Alan Hoenig. Recovering from a hard-disk failure. TUGboat, 9(3):269–271, November 1988. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb09-3/tb22pfeffer.pdf.

Laan:TB9-3-271

[817] C. G. van der Laan and J. R. Luyten. Evaluation of K-talk. TUGboat, 9(3): 271–272, November 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-3/tb22laan.pdf.

Wittbecker:TB9-3-272

[818] Alan Wittbecker. Making paragraphs. TUGboat, 9(3):272–276, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22wittbecker.pdf.

Bechtolsheim:TB9-3-276

[819] Stephan v. Bechtolsheim. A tutorial on \futurelet. TUGboat, 9(3): 276–278, November 1988. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb09-3/tb22bechtolsheim.pdf.

Kneser:TB9-3-279

[820] Thomas Kneser. Compact matrix display. TUGboat, 9(3):279–284, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22kneser.pdf.

Appelt:TB9-3-284

[821] Wolfgang Appelt. Typesetting chess. TUGboat, 9(3):284–287, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22appelt.pdf.

Pittman:TB9-3-287

[822] J. E. Pittman. Equation numbering in plain TEX. TUGboat, 9(3):287–288, November 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-3/tb22pittman-eq.pdf.

Pittman:TB9-3-289

[823] J. E. Pittman. Loopy.TeX. TUGboat, 9(3):289–291, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22pittman-loop.pdf.

Rogers:TB9-3-292

[824] David F. Rogers. A page make-up challenge. TUGboat, 9(3):292–293, Novem-

ber 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-
3/tb22rogers.pdf.

DeCorte:TB9-3-294

[825] Michael DeCorte. Contents of L^{A}TEX style collection as of 24 September 1988. TUGboat, 9(3):294–297, November 1988. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb09-3/tb22decorte.pdf.

Damrau:TB9-3-297

[826] Jackie Damrau. The L^{A}TEX column. TUGboat, 9(3):297, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22damrau.pdf.

Mittelbach:TB9-3-298

[827] Frank Mittelbach. A new implementation of the array– and tabular– environments. TUGboat, 9(3):298–314, November 1988. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb09-3/tb22mitt.pdf.

Anonymous:TB9-3-315

[828] Anonymous. Calendar. TUGboat, 9(3):315, November 1988. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb09-3/tb22news.pdf.

Laan:TB9-3-316

[829] C. G. van der Laan. Dutch TEX Users Group. TUGboat, 9(3):316, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22news.pdf.

Anonymous:TB9-3-316

[830] Anonymous. Call for papers: RIDT’89 — International Workshop on Raster Imagining and Digital Typography, Lausanne, Switzerland, 12–13 October 1989. TUGboat, 9(3):316–317, November 1988. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb09-3/tb22news.pdf.

MacKay:TB9-3-317

[831] Pierre MacKay. TEXhax moves north. TUGboat, 9(3):317, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22late.pdf.

Beeton:TB9-3-317

[832] Barbara Beeton. Production notes. TUGboat, 9(3):317–318, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22late.pdf.

Anonymous:TB9-3-318

[833] Anonymous. A thank you note. TUGboat, 9(3):318, November 1988. ISSN

0896-3207. URL https://tug.org/TUGboat/tb09-3/tb22treas.
pdf.

Anonymous:TB9-3-318-2

[834] Anonymous. Accountant’s review of TUG financial records, years ended December 31, 1987 and 1986. TUGboat, 9(3):318–321, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22treas.pdf.

Anonymous:TB9-3-322

[835] Anonymous. Institutional members. TUGboat, 9(3):322–324, November

1988. ISSN 0896-3207. URL https://tug.org/TUGboat/tb09-3/
tb22instmem.pdf.

Anonymous:TB9-3-324

[836] Anonymous. TUG’s Tenth. TUGboat, 9(3):324, November 1988. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb09-3/tb22tenth.pdf.

Anonymous:TB9-3-325

[837] Anonymous. TUG membership application. TUGboat, 9(3):325, November

1988. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB9-3-342

[838] Anonymous. Index of advertisers. TUGboat, 9(3):342, November 1988. ISSN 0896-3207.

Anonymous:TB9-3-M

[839] Anonymous. TUG Membership List Supplement. TUGboat, 9(3):M–1–M–12, November 1988. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb09-3/tb22mem.pdf.

Anonymous:TB9-3-cover-3

[840] Anonymous. Call for papers: TUG Tenth Annual Meeting, Stanford University, 21–23 August 1989. TUGboat, 9(3):Cover 3, November 1988. ISSN 0896-3207.

Anonymous:1989:TPa

[841] Anonymous. Title page. TUGboat, 10(1):1, April 1989. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23titlepage.pdf.

Anonymous:TB10-1-3

[842] Anonymous. Addresses. TUGboat, 10(1):3–4, April 1989. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb10-1/tb23edit.pdf.

Anonymous:TB10-1-5

[843] Anonymous. Donald E. Knuth Scholarship. TUGboat, 10(1):5, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23gendel.pdf.

Childs:TB10-1-5

[844] Bart Childs. From the President. TUGboat, 10(1):5, April 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-1/tb23gendel.
pdf.

Anonymous:TB10-1-5-2

[845] Anonymous. Announcing a TUG dingbat competition. TUGboat, 10(1):5– 6, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23gendel.pdf.

Beeton:TB10-1-6

[846] Barbara Beeton. Editorial comments. TUGboat, 10(1):6–7, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23gendel.pdf.

Knuth:TB10-1-8

[847] Donald Knuth. TEX would find it difficult …. TUGboat, 10(1):8, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23gendel.pdf.

Beeton:TB10-1-8

[848] Barbara Beeton. A TEX encounter in Japan. TUGboat, 10(1):8–10, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23gendel.pdf.

Harrison:TB10-1-11

[849] Michael Harrison. News from the VORTEX project. TUGboat, 10(1):11– 14, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23harrison.pdf.

Clark:TB10-1-14

[850] Adrian Clark. An enhanced TEX-editor interface for VMS. TUGboat, 10 (1):14–15, April 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-1/tb23clark.pdf.

Thull:TB10-1-15

[851] Klaus Thull. The virtual memory management of PubliC TEX. TUGboat, 10(1):15–22, April 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-1/tb23thull.pdf.

Kinch:TB10-1-23

[852] Richard Kinch. Turbo : A new port in C for unix and MS-DOS. TUGboat, 10(1):23–24, April 1989. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb10-1/tb23kinch.pdf.

Bechtolsheim:TB10-1-25

[853] Stephan v. Bechtolsheim. The TEX PostScript software package. TUGboat, 10(1):25–27, April 1989. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb10-1/tb23bechtolsheim.pdf.

Tobin:TB10-1-28

[854] Georgia K. M. Tobin. A handy little font. TUGboat, 10(1):28–30, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23tobin.pdf.

Knuth:TB10-1-31

[855] Donald Knuth. Typesetting Concrete Mathematics. TUGboat, 10(1):31– 36, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23knut.pdf.

Henderson:TB10-1-36

[856] Doug Henderson. Outline fonts with . TUGboat, 10(1):36– 38, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23henderson.pdf.

Wujastyk:TB10-1-39

[857] Dominik Wujastyk. Font news. TUGboat, 10(1):39, April 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-1/tb23wujastyk.
pdf.

Rogers:TB10-1-39

[858] David F. Rogers. Computer graphics and TEX — a challenge. TUGboat, 10(1):39–44, April 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-1/tb23rogers.pdf.

Childs:TB10-1-44

[859] Bart Childs, Alan Stolleis, and Don Berryman. A portable graphics inclusion. TUGboat, 10(1):44–46, April 1989. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb10-1/tb23childs.pdf.

Hosek:TB10-1-46

[860] Don Hosek. TEX output devices (with charts). TUGboat, 10(1):46–55, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23output.pdf.

Hosek:TB10-1-56

[861] Don Hosek. Report from the DVI driver standards committee. TUGboat, 10(1): 56, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23hosek.pdf.

Broeren:TB10-1-56

[862] Marius Broeren and Jan van Knippenberg. High quality printing of TEX DVI output files in the VAX/VMS environment. TUGboat, 10(1):56–58, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23broeren.pdf.

Puente:TB10-1-58

[863] Arturo Puente. Output driver for Xerox 4045 on IBM 3090. TUGboat, 10(1): 58, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23query.pdf.

Abbott:TB10-1-59

[864] Peter Abbott. UKTEX and the Aston archive. TUGboat, 10(1):59–60, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23site.pdf.

Childs:TB10-1-60

[865] Bart Childs. DG site report. TUGboat, 10(1):60, April 1989. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb10-1/tb23site.pdf.

MacKay:TB10-1-60

[866] Pierre MacKay. UnixTEX site report. TUGboat, 10(1):60–61, April 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-1/tb23site.
pdf.

Kellerman:TB10-1-62

[867] David Kellerman. VAX/VMS site report. TUGboat, 10(1):62, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23site.pdf.

Hoenig:TB10-1-62

[868] Alan Hoenig. The land of the free and the near free. TUGboat, 10(1):62– 64, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23hoenig.pdf.

Kubik:TB10-1-65

[869] Kim Kubik. AmigaTEX… or How envy was resisted and knowledge found on the road to Ööç. TUGboat, 10(1):65–67, April 1989. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb10-1/tb23kubik.pdf.

Modest:TB10-1-67

[870] Michael Modest. Using TEX and L^{A}TEX with WordPerfect 5.0. TUGboat, 10(1):67–72, April 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-1/tb23modest.pdf.

Zalmstra:TB10-1-73

[871] Joost Zalmstra and David F. Rogers. A page make-up macro. TUGboat, 10(1):73–81, April 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-1/tb23zalmstra.pdf.

Nearing:TB10-1-82

[872] James Nearing. Extended equation numbering in Plain TEX. TUGboat, 10(1):82–88, April 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-1/tb23nearing.pdf.

Sankar:TB10-1-89

[873] Sriram Sankar. APE — A set of TEX macros to format Ada programs. TUGboat, 10(1):89–97, April 1989. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb10-1/tb23sankar.pdf.

DeCorte:TB10-1-97

[874] Michael DeCorte. Contents of L^{A}TEX style collection as of 16 January 1989. TUGboat, 10(1):97–102, April 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-1/tb23decorte.pdf.

Mittelbach:TB10-1-103

[875] Frank Mittelbach. “A new implementation of the array — and tabular — environments of L^{A}TEX” (TUGboat 9#3) — addenda. TUGboat, 10(1):103– 104, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23mitt.pdf.

Schopf:TB10-1-105

[876] Rainer Schöpf. Drawing histogram bars inside the L^{A}TEX picture–environment. TUGboat, 10(1):105–107, April 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-1/tb23schopf.pdf.

Nagy:TB10-1-108

[877] Dezső Nagy. Vertical centering for transparencies. TUGboat, 10(1):108– 112, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23nagy.pdf.

Laan:TB10-1-113

[878] C. G. van der Laan. Typesetting bridge via L^{A}TEX. TUGboat, 10(1):113– 116, April 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-1/tb23laan.pdf.

Anonymous:TB10-1-116

[879] Anonymous. WOODMAN ’89: Workshop on Object Oriented Document Manipulation, Rennes, France, 29–31 May 1989. TUGboat, 10(1):116, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23news.pdf.

Anonymous:TB10-1-117

[880] Anonymous. Calendar. TUGboat, 10(1):117, April 1989. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb10-1/tb23news.pdf.

Anonymous:TB10-1-118

[881] Anonymous. GUTenberg Congrès: TEX et les graphiques, Paris, 16–17 May 1989. TUGboat, 10(1):118, April 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-1/tb23news.pdf.

Anonymous:TB10-1-118-2

[882] Anonymous. TEX89: Karlsruhe University, 11–13 September 1989. TUGboat, 10(1):118, April 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-1/tb23news.pdf.

Osborne:TB10-1-119

[883] David Osborne. A UK TEX users’ group — Report of a preliminary meeting. TUGboat, 10(1):119–120, April 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-1/tb23news.pdf.

Beeton:TB10-1-120

[884] Barbara Beeton. Production notes. TUGboat, 10(1):120–121, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23late.pdf.

Anonymous:TB10-1-121

[885] Anonymous. Coming next issue. TUGboat, 10(1):121, April 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-1/tb23late.
pdf.

Anonymous:TB10-1-122

[886] Anonymous. Institutional members. TUGboat, 10(1):122–124, April

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-1/
tb23instmem.pdf.

Anonymous:TB10-1-125

[887] Anonymous. TUG membership application. TUGboat, 10(1):125, April 1989.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB10-1-142

[888] Anonymous. Index of advertisers. TUGboat, 10(1):142, April 1989. ISSN 0896-3207.

Anonymous:1989:TPb

[889] Anonymous. Title page. TUGboat, 10(2):143, July 1989. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24titlepage.pdf.

Anonymous:TB10-2-145

[890] Anonymous. Addresses. TUGboat, 10(2):145–146, July 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-2/tb24edit.pdf.

Childs:TB10-2-147

[891] Bart Childs. From the President. TUGboat, 10(2):147, July 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-2/tb24gendel.
pdf.

Beeton:TB10-2-147

[892] Barbara Beeton. Editorial comments. TUGboat, 10(2):147–150, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24gendel.pdf.

Clark:TB10-2-150

[893] Malcolm Clark. Réflexions sur le Congrès GUTenberg. TUGboat, 10(2):150– 153, July 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-2/tb24gendel.pdf.

Clark:TB10-2-153

[894] Malcolm Clark. International Standards and TEX. TUGboat, 10(2):153–156,

July 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-
2/tb24gendel.pdf.

Childs:TB10-2-156

[895] Bart Childs. Teaching TEX. TUGboat, 10(2):156–163, July 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-2/tb24childs.
pdf.

Spivak:TB10-2-164

[896] Michael Spivak, Michael Ballantyne, and Yoke Lee. HI-TEX cutting & pasting. TUGboat, 10(2):164–165, July 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-2/tb24spivak.pdf.

Tobin:TB10-2-166

[897] Georgia K. M. Tobin. Another dingbat idea. TUGboat, 10(2):166–169, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24tobin.pdf.

Rubinstein:TB10-2-170

[898] Zalman Rubinstein. Chess printing via and TEX. TUGboat, 10(2):170–172, July 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-2/tb24rubinstein.pdf.

Hosek:TB10-2-173

[899] Don Hosek. Guidelines for creating portable code. TUGboat, 10(2):173–176, July 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-2/tb24hosek-mf.pdf.

Andrews:TB10-2-177

[900] Phil Andrews. Integration of TEX and graphics at the Pittsburgh Supercomputing Center. TUGboat, 10(2):177–178, July 1989. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb10-2/tb24andrews.pdf.

Wilcox:TB10-2-179

[901] Patricia Wilcox. METAPLOT: Machine-independent line graphics for TEX. TUGboat, 10(2):179–187, July 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-2/tb24wilcox.pdf.

Hosek:TB10-2-188

[902] Don Hosek. TEX output devices. TUGboat, 10(2):188, July 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-2/tb24hosek-
output.pdf.

Reid:TB10-2-188

[903] Tom Reid and Don Hosek. Report from the DVI driver standards committee. TUGboat, 10(2):188–191, July 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-2/tb24hosek-output.pdf.

Hosek:TB10-2-192

[904] Don Hosek. Announcing (belatedly) TEXMAG. TUGboat, 10(2):192, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24resources.pdf.

Clark:TB10-2-193

[905] Malcolm Clark. TEXline: A newsletter of the TEX community. TUGboat, 10(2):193–194, July 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-2/tb24resources.pdf.

Abbott:TB10-2-194

[906] Peter Abbott. UKTEX and the Aston archive. TUGboat, 10(2):194–195, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24resources.pdf.

Nieland:TB10-2-195

[907] M. Edward Nieland. The DECUS TEX collection. TUGboat, 10(2):195–196,

July 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-
2/tb24resources.pdf.

DeCorte:TB10-2-196

[908] Michael DeCorte. The Clarkson archive server — Update as of 1 May 1989. TUGboat, 10(2):196–201, July 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-2/tb24resources.pdf.

Childs:TB10-2-201

[909] Bart Childs. Data General site report. TUGboat, 10(2):201, July 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-2/tb24site.
pdf.

Crawford:TB10-2-201

[910] John Crawford. Prime 50 series site report. TUGboat, 10(2):201, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24site.pdf.

Radel:TB10-2-202

[911] Jon Radel. “Free” TEX software for IBM PCs. TUGboat, 10(2):202, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24pc.pdf.

Trevorrow:TB10-2-202

[912] Andrew Trevorrow. Public domain TEX for the Mac. TUGboat, 10(2):202–203,

July 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-
2/tb24pc.pdf.

Bechtolsheim:TB10-2-203

[913] Stephan v. Bechtolsheim. \csname and \string. TUGboat, 10(2):203– 206, July 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-2/tb24bechtolsheim.pdf.

Salomon:TB10-2-207

[914] David Salomon. DDA methods in TEX. TUGboat, 10(2):207–216, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24salomon.pdf.

Vogel:TB10-2-217

[915] Brother Eric Vogel. Printing Vietnamese characters by adding diacritical marks. TUGboat, 10(2):217–221, July 1989. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb10-2/tb24vogel.pdf.

Mittelbach:TB10-2-222

[916] Frank Mittelbach and Rainer Schöpf. A new font selection scheme for TEX macro packages — the basic macros. TUGboat, 10(2):222–238, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24mitt.pdf.

Nagy:TB10-2-239

[917] Dezső Nagy. A bar chart in L^{A}TEX. TUGboat, 10(2):239–240, July 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-2/tb24nagy.
pdf.

Partl:TB10-2-241

[918] Hubert Partl. Producing on-line information files with L^{A}TEX. TUGboat, 10 (2):241–244, July 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-2/tb24partl.pdf.

Mittelbach:TB10-2-245

[919] Frank Mittelbach. The doc–option. TUGboat, 10(2):245–273, July 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-2/tb24mitt-
doc.pdf.

Kelly:TB10-2-274

[920] B. Hamilton Kelly. The autodoc–option. TUGboat, 10(2):274–284, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24hamiltonkelly.pdf.

Anonymous:TB10-2-285

[921] Anonymous. Calendar. TUGboat, 10(2):285–286, July 1989. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb10-2/tb24news.pdf.

Anonymous:TB10-2-287

[922] Anonymous. Nordic TEX meeting, 12 June 1989, Royal Institute of Technology, Stockholm. TUGboat, 10(2):287, July 1989. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb10-2/tb24news.pdf.

Anonymous:TB10-2-287-2

[923] Anonymous. L^{A}TEX course book in Dutch. TUGboat, 10(2):287, July

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24news.pdf.

Lammarsch:TB10-2-287

[924] Joachim Lammarsch. DANTE — Deutschsprachige Anwendervereinigung TEX. TUGboat, 10(2):287–288, July 1989. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb10-2/tb24news.pdf.

Osborne:TB10-2-288

[925] David Osborne. Notes on first meeting of UK TEX Users’ Group, 15 March 1989. TUGboat, 10(2):288–289, July 1989. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb10-2/tb24news.pdf.

Beeton:TB10-2-289

[926] Barbara Beeton. Production notes. TUGboat, 10(2):289, July 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-2/tb24news.
pdf.

Anonymous:TB10-2-290

[927] Anonymous. TUG membership application. TUGboat, 10(2):290, July 1989.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:1989:A

[928] Anonymous. Advertisements. TUGboat, 10(2):291, July 1989. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-2/
tb24ads.pdf.

Anonymous:TB10-2-M

[929] Anonymous. TUG Membership List Supplement. TUGboat, 10(2):M–1–M– 11, July 1989. ISSN 0896-3207.

Anonymous:1989:TPd

[930] Anonymous. Title page. TUGboat, 10(3):307, November 1989. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25titlepage.pdf.

Anonymous:1989:EIb

[931] Anonymous. Editorial information. TUGboat, 10(3):308, November 1989. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25edit.pdf.

Anonymous:TB10-3-309

[932] Anonymous. Addresses. TUGboat, 10(3):309–310, November 1989. ISSN 0896-3207.

Childs:TB10-3-311

[933] Bart Childs. From the President. TUGboat, 10(3):311–312, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25gendel.pdf.

Beebe:TB10-3-312

[934] Nelson H. F. Beebe. Message from the new President. TUGboat, 10(3): 312–314, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25gendel.pdf.

Beeton:TB10-3-314

[935] Barbara Beeton. Editorial comments. TUGboat, 10(3):314–316, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25gendel.pdf.

Eijkhout:TB10-3-316

[936] Victor Eijkhout and Nico Poppelier. The first Dutch TEX days. TUGboat, 10 (3):316–318, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25gendel.pdf.

Childs:TB10-3-319

[937] Bart Childs. Answers to the TEX tests. TUGboat, 10(3):319–323, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25childs.pdf.

Technites:TB10-3-324

[938] Technites. Crossword. TUGboat, 10(3):324, November 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-3/tb25crossword.
pdf.

Knuth:TB10-3-325

[939] Donald E. Knuth. The new versions of TEX and . TUGboat, 10 (3):325–328, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25knut.pdf.

Thull:TB10-3-328

[940] Klaus Thull. PubliC available. TUGboat, 10(3):328, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25thull.pdf.

Bechtolsheim:TB10-3-329

[941] Stephan von Bechtolsheim. A .dvi file processing program. TUGboat, 10(3): 329–332, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25bechtolsheim.pdf.

Vulis:TB10-3-332

[942] Dimitri Vulis. Notes on Russian TEX. TUGboat, 10(3):332–336, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25vulis.pdf.

Beeton:TB10-3-336

[943] Barbara Beeton. Hyphenation exception log. TUGboat, 10(3):336–341, November 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25hyph.pdf.

Anonymous:TB10-3-342

[944] Anonymous. Erratum: Typesetting Concrete Mathematics, TUGboat Vol. 10, No. 1. TUGboat, 10(3):342, November 1989. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb10-3/tb25hara-latin.pdf.

Haralambous:TB10-3-342

[945] Yannis Haralambous. TEX and latin alphabet languages. TUGboat, 10(3): 342–345, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25hara-latin.pdf.

Parker:TB10-3-346

[946] Mike Parker. Fonts and PostScript. TUGboat, 10(3):346–350, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25parker.pdf.

Larsson:TB10-3-351

[947] Jan Eric Larsson. A chess font for TEX. TUGboat, 10(3):351, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25larsson.pdf.

Andulem:TB10-3-352

[948] Abass Andulem. The road to Ethiopic TEX. TUGboat, 10(3):352–354, Novem-

ber 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-
3/tb25andulem.pdf.

Haralambous:TB10-3-354

[949] Yannis Haralambous and Klaus Thull. Typesetting modern Greek with 128 character codes. TUGboat, 10(3):354–359, November 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-3/tb25hara-
greek.pdf.

Anonymous:TB10-3-359

[950] Anonymous. Erratum: Chess printing via and TEX, TUGboat Vol. 10, No. 2. TUGboat, 10(3):359, November 1989. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb10-3/tb25chess.pdf.

Hailperin:TB10-3-360

[951] Max Hailperin. Users’ guide to LaTeX-help. TUGboat, 10(3):360, Novem-

ber 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-
3/tb25hailperin.pdf.

Kean:TB10-3-360

[952] Yin Kean. TEX-Ed. TUGboat, 10(3):360–361, November 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-3/tb25kean.pdf.

DeCorte:TB10-3-362

[953] Michael DeCorte. The Clarkson archive server — Update as of 22 September 1989. TUGboat, 10(3):362–365, November 1989. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb10-3/tb25decorte.pdf.

Anonymous:TB10-3-365

[954] Anonymous. Coming in January from the AMS. TUGboat, 10(3):365–366, November 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25ams.pdf.

Lammarsch:TB10-3-366

[955] Joachim Lammarsch. TEX EURO. TUGboat, 10(3):366–367, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25site.pdf.

Gaulle:TB10-3-367

[956] Bernard Gaulle. GUTenberg will distribute MLTEX. TUGboat, 10(3):367, November 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25site.pdf.

Glendown:TB10-3-367

[957] Garry Glendown. A TEX mailbox in Germany. TUGboat, 10(3):367–368, November 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25site.pdf.

Lindner:TB10-3-368

[958] Stephan Lindner and Lutz Birkhahn . Towards a complete and comfortable TEX System. TUGboat, 10(3):368–372, November 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-3/tb25site.pdf.

Kellerman:TB10-3-372

[959] David Kellerman. VMS site coordinator’s report. TUGboat, 10(3):372, November 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25site.pdf.

Hoenig:TB10-3-373

[960] Alan Hoenig and Mitch Pfeffer. TEX-PostScript output on non-Post-Script devices. TUGboat, 10(3):373–375, November 1989. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb10-3/tb25hoenig.pdf.

Glendown:TB10-3-376

[961] Garry Glendown. Controlling AmigaTEX from CygnusEd. TUGboat, 10(3): 376–377, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25glendown-pc.pdf.

Whitney:TB10-3-378

[962] Ron Whitney and Barbara Beeton. TUGboat authors’ guide. TUGboat, 10(3): 378–385, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25whitney.pdf.

Glendown:TB10-3-385

[963] Garry Glendown. Round boxes for plain TEX. TUGboat, 10(3):385–387, November 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25glendown-round.pdf.

Rubinstein:TB10-3-387

[964] Zalman Rubinstein. Printing annotated chess literature in natural notation. TUGboat, 10(3):387–390, November 1989. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb10-3/tb25rubinstein.pdf.

Durst:TB10-3-390

[965] Lincoln Durst. Bibliographic citations; or variations on the old shell game. TUGboat, 10(3):390–394, November 1989. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb10-3/tb25durst.pdf.

Salomon:TB10-3-394

[966] David Salomon. Macros for indexing and table-of-contents preparation. TUGboat, 10(3):394–400, November 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-3/tb25salomon.pdf.

Bush:TB10-3-400

[967] Terry Bush. Scribe-to-TEX. TUGboat, 10(3):400, November 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-3/tb25mitt-
latex.pdf.

Mittelbach:TB10-3-400

[968] Frank Mittelbach and Rainer Schöpf. Towards L^{A}TEX 2.10. TUGboat, 10(3): 400–401, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25mitt-latex.pdf.

Braams:TB10-3-401

[969] Johannes Braams, Victor Eijkhout, and Nico Poppelier. The development of national L^{A}TEX styles. TUGboat, 10(3):401–406, November 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-3/tb25braams.
pdf.

Mittelbach:TB10-3-407

[970] Frank Mittelbach. An environment for multicolumn output. TUGboat, 10(3): 407–415, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25mitt-multicol.pdf.

Mittelbach:TB10-3-416

[971] Frank Mittelbach. An extension of the L^{A}TEX theorem environment. TUGboat, 10(3):416–426, November 1989. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb10-3/tb25mitt-theorem.pdf.

Dietsche:TB10-3-427

[972] Luzia Dietsche. Deutsche Kurzfassungen der TUGboat-Artikel. TUGboat, 10 (3):427–428, November 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-3/tb25abstracts.pdf.

Anonymous:TB10-3-429

[973] Anonymous. Calendar and Calls for Papers. TUGboat, 10(3):429–431, November 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25calendar.pdf.

Anonymous:TB10-3-432

[974] Anonymous. Programs of the recent TUG and EuroTEX Conferences. TUGboat, 10(3):432–436, November 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-3/tb25programs.pdf.

Beeton:TB10-3-437

[975] Barbara Beeton. Production notes. TUGboat, 10(3):437, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25prod.pdf.

Anonymous:TB10-3-438

[976] Anonymous. Institutional members. TUGboat, 10(3):438–440, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25instmem.pdf.

Ness:TB10-3-440

[977] David Ness. Treasurer’s report. TUGboat, 10(3):440–443, November

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25treas.pdf.

Anonymous:TB10-3-444

[978] Anonymous. TUG membership application. TUGboat, 10(3):444, November

1989. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB10-3-445

[979] Anonymous. [Advertisements]. TUGboat, 10(3):445–462, November 1989. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25ads.pdf.

Anonymous:TB10-3-461

[980] Anonymous. Index of advertisers. TUGboat, 10(3):461, November 1989. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-
3/tb25ads.pdf.

Anonymous:TB10-3-M

[981] Anonymous. TUG Membership List Supplement. TUGboat, 10(3):M–1–M– 12, November 1989. ISSN 0896-3207.

Anonymous:1989:TPf

[982] Anonymous. Title page. TUGboat, 10(4):463, December 1989. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-3/
tb25titlepage.pdf.

Anonymous:1989:EId

[983] Anonymous. Editorial information. TUGboat, 10(4):464, December 1989. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-3/tb25edit.pdf.

Thiele:TB10-4-465

[984] Christina Thiele. Editor’s introduction. TUGboat, 10(4):465, December

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-4/
tb26thiele.pdf.

Henderson:TB10-4-467

[985] Doug Henderson. Introduction to . TUGboat, 10(4):467–479, December 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-4/tb26henderson.pdf.

Billawala:TB10-4-481

[986] Neenie Billawala. Opening Pandora’s box. TUGboat, 10(4):481–489, Decem-

ber 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-
4/tb26billawala.pdf.

Hoenig:TB10-4-491

[987] Alan Hoenig. Fractal images with TEX. TUGboat, 10(4):491–498, December

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-4/
tb26hoenig.pdf.

Hosek:TB10-4-499

[988] Don Hosek. Design of Oriental characters with . TUGboat, 10(4): 499–502, December 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-4/tb26hosek.pdf.

Batzinger:TB10-4-503

[989] Bob Batzinger. Thai languages and . TUGboat, 10(4):503, December 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-4/tb26batzinger.pdf.

Hobby:TB10-4-505

[990] John D. Hobby. A -like system with PostScript output. TUGboat, 10(4):505–512, December 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-4/tb26hobby.pdf.

Youngen:TB10-4-513

[991] R. E. Youngen, W. B. Woolf, and D. C. Latterner. Migration from Computer Modern fonts to Times fonts. TUGboat, 10(4):513–519, December

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-4/
tb26youngen.pdf.

Conrad:TB10-4-521

[992] Arvin C. Conrad. Fine typesetting with TEX using native Autologic fonts. TUGboat, 10(4):521–528, December 1989. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb10-4/tb26conrad.pdf.

Knuth:TB10-4-529

[993] Donald E. Knuth. The errors of TEX. TUGboat, 10(4):529–531, December

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-4/
tb26knut.pdf.

Doob:TB10-4-533

[994] Michael Doob. Of the computer scientist, by the computer scientist, for the computer scientist. TUGboat, 10(4):533–539, December 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-4/tb26doob.pdf.

Hamilton:TB10-4-541

[995] Hope Hamilton. Mastering TEX with templates. TUGboat, 10(4):541–548, December 1989. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb10-4/tb26hamilton.pdf.

Hoover:TB10-4-549

[996] Anita Z. Hoover. Using WordPerfect 5.0 to create TEX and L^{A}TEX documents. TUGboat, 10(4):549–559, December 1989. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb10-4/tb26hoover.pdf.

Kubek:TB10-4-561

[997] Robin L. Kubek. TEX for the word processing operator. TUGboat, 10(4): 561–566, December 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-4/tb26kubek.pdf.

Rattey-Hicks:TB10-4-567

[998] Jo Ann Rattey-Hicks. TEX and its versatility in office production. TUGboat, 10 (4):567–577, December 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-4/tb26rattey.pdf.

Diaz:TB10-4-579

[999] Max Díaz. TEX in México. TUGboat, 10(4):579–593, December 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-4/tb26diaz.
pdf.

Haskell:TB10-4-595

[1000] James Haskell, Wally Deschene, and Alan Stolleis. TEX for 30,000. TUGboat, 10(4):595–602, December 1989. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb10-4/tb26haskell.pdf.

Wittbecker:TB10-4-603

[1001] Alan E. Wittbecker. TEX enslaved. TUGboat, 10(4):603–606, December

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-4/
tb26wittbecker.pdf.

Renfrow:TB10-4-607

[1002] J. T. Renfrow. Methodologies for preparing and integrating PostScript graphics. TUGboat, 10(4):607–626, December 1989. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb10-4/tb26renfrow.pdf.

Olejniczak-Burkert:TB10-4-627

[1003] Rolf Olejniczak-Burkert. texpic — Design and implementation of a picture graphics language in TEX à la pic. TUGboat, 10(4):627–637, December

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-4/
tb26olejniczak.pdf.

Latterner:TB10-4-639

[1004] D. C. Latterner and W. B. Woolf. TEX at Mathematical Reviews. TUGboat, 10 (4):639–654, December 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-4/tb26latterner.pdf.

Pind:TB10-4-655

[1005] Jörgen L. Pind. Lexicography with TEX. TUGboat, 10(4):655–665, December

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-4/
tb26pind.pdf.

Clark:TB10-4-667

[1006] Malcolm Clark. Olde worlde TEX. TUGboat, 10(4):667–674, December

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-4/
tb26clark.pdf.

Abbott:TB10-4-675

[1007] Peter Abbott. The UKTEX archive at the University of Aston. TUGboat, 10 (4):675–680, December 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-4/tb26abbott.pdf.

Mittelbach:TB10-4-681

[1008] Frank Mittelbach and Rainer Schöpf. With L^{A}TEX into the nineties. TUGboat, 10(4):681–690, December 1989. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb10-4/tb26mitt.pdf.

Greene:TB10-4-691

[1009] Andrew Marc Greene. TEXreation — playing games with TEX’s mind. TUGboat, 10(4):691–705, December 1989. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb10-4/tb26greene.pdf.

Cheswick:TB10-4-707

[1010] Bill Cheswick. A permuted index for TEX and L^{A}TEX. TUGboat, 10(4): 707–713, December 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-4/tb26cheswick.pdf.

Sydoriak:TB10-4-715

[1011] Steve Sydoriak. L^{A}TEX memos and letters. TUGboat, 10(4):715–726, December

1989. ISSN 0896-3207. URL https://tug.org/TUGboat/tb10-4/
tb26sydoriak.pdf.

Benson:TB10-4-727

[1012] Gary Benson, Debi Erpenbeck, and Janet Holmes. Inserts in a multiple-column format. TUGboat, 10(4):727–742, December 1989. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb10-4/tb26benson.pdf.

McClure:TB10-4-743

[1013] Mary McClure. TEX Macros for COBOL syntax diagrams. TUGboat, 10(4): 743–750, December 1989. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb10-4/tb26mcclure.pdf.

Riley:TB10-4-751

[1014] Don L. Riley and Brad L. Halverson. Creating an efficient and workable PC interface for TEX. TUGboat, 10(4):751–759, December 1989. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb10-4/tb26riley.pdf.

Anonymous:TB10-4-760

[1015] Anonymous. TUG membership application. TUGboat, 10(4):760, December

1989. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB10-4-761

[1016] Anonymous. List of participants at the Stanford meeting (August 20–23, 1989). TUGboat, 10(4):761–765, December 1989. ISSN 0896-3207.

Anonymous:TB10-4-765

[1017] Anonymous. Index of advertisers. TUGboat, 10(4):765, December 1989. ISSN

0896-3207. URL https://tug.org/TUGboat/tb10-4/tb26ads.
pdf.

Anonymous:1990:TPa

[1018] Anonymous. Title page. TUGboat, 11(1):1, April 1990. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27titlepage.pdf.

Anonymous:1990:EIa

[1019] Anonymous. Editorial information. TUGboat, 11(1):2, April 1990. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27edit.pdf.

Anonymous:TB11-1-3

[1020] Anonymous. Addresses. TUGboat, 11(1):3–4, April 1990. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb11-1/tb27addresses.pdf.

Beebe:TB11-1-5

[1021] Nelson H. F. Beebe. From the President. TUGboat, 11(1):5–6, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27gendel.pdf.

Childs:TB11-1-6

[1022] Bart Childs. From the past President and Annual Meeting host. TUGboat, 11(1):6–7, April 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-1/tb27gendel.pdf.

Anonymous:TB11-1-7

[1023] Anonymous. Donald E. Knuth Scholarship. TUGboat, 11(1):7–8, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27gendel.pdf.

Ferguson:TB11-1-8

[1024] Michael Ferguson. Coordination of non-English use of TEX. TUGboat, 11(1): 8–9, April 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-1/tb27gendel.pdf.

Breitenlohner:TB11-1-9

[1025] Peter Breitenlohner. Using TEX 3 in a multilingual environment — Some ideas. TUGboat, 11(1):9–12, April 1990. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb11-1/tb27gendel.pdf.

Anonymous:TB11-1-12

[1026] Anonymous. Erratum: The new versions of TEX and , TUGboat vol. 10, no. 3. TUGboat, 11(1):12, April 1990. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb11-1/tb27erratum.pdf.

Knuth:TB11-1-13

[1027] Donald Knuth. Virtual Fonts: More Fun for Grand Wizards. TUGboat, 11(1):13–23, April 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-1/tb27knut.pdf.

Kuiken:TB11-1-24

[1028] Gerard Kuiken. Additional Hyphenation Patterns. TUGboat, 11(1):24–25, April 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-1/tb27kuiken.pdf.

Haralambous:TB11-1-26

[1029] Yannis Haralambous. Typesetting modern Greek — An update. TUGboat, 11(1):26, April 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-1/tb27hara.pdf.

Pickrell:TB11-1-26

[1030] Lee S. Pickrell. Combining graphics with TEX on IBM PC-compatible systems and LaserJet printers. TUGboat, 11(1):26–31, April 1990. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb11-1/tb27pickrell.pdf.

Beeton:TB11-1-32

[1031] Barbara Beeton and Ron Whitney. Summary of resources available to TEX users. TUGboat, 11(1):32–36, April 1990. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb11-1/tb27beet-res.pdf.

Beeton:TB11-1-36

[1032] Barbara Beeton. A proto-TUG bibliography. TUGboat, 11(1):36–37, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27beet-bib.pdf.

DeCorte:TB11-1-38

[1033] Michael DeCorte. The Clarkson archive server — Update as of 26 January 1990. TUGboat, 11(1):38–45, April 1990. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb11-1/tb27decorte.pdf.

Guenther:TB11-1-45

[1034] Dean Guenther. IBM VM/CMS site report. TUGboat, 11(1):45, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27guenther-vmcms.pdf.

Brouard:TB11-1-46

[1035] Nicolas Brouard. Une version complète de TEX du domaine public pour compatibles PC : les “deux disquettes GUT”. TUGboat, 11(1):46–51, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27brouard.pdf.

Eijkhout:TB11-1-51

[1036] Victor Eijkhout. Unusual Paragraph Shapes. TUGboat, 11(1):51–53, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27eijkhout.pdf.

Guenther:TB11-1-54

[1037] Dean Guenther. TEXT1 goes public domain. TUGboat, 11(1):54–56, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27guenther-text1.pdf.

Anonymous:TB11-1-56

[1038] Anonymous. Solution to crossword of TUGboat vol. 10, no. 3. TUGboat, 11(1):56, April 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-1/tb27cross.pdf.

Wichura:TB11-1-57

[1039] Michael J. Wichura. Showing-off math macros. TUGboat, 11(1):57–61, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27wichura.pdf.

Breitenlohner:TB11-1-62

[1040] Peter Breitenlohner. How to avoid writing long records to TEX’s \write streams. TUGboat, 11(1):62, April 1990. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb11-1/tb27breitenlohner.pdf.

Durst:TB11-1-62

[1041] Lincoln Durst. Forward references and the ultimate dirty trick. TUGboat, 11(1):62–68, April 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-1/tb27durst.pdf.

Salomon:TB11-1-69

[1042] David Salomon. Output routines: Examples and techniques. Part I: Introduction and examples. TUGboat, 11(1):69–85, April 1990. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb11-1/tb27salomon.pdf.

Damrau:TB11-1-85

[1043] Jackie Damrau. The L^{A}TEX column. TUGboat, 11(1):85–86, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27damrau.pdf.

Laan:TB11-1-86

[1044] C. G. van der Laan. Announcing two reports: SGML-L^{A}TEX and Journal style guidelines. TUGboat, 11(1):86, April 1990. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb11-1/tb27laan.pdf.

Schrod:TB11-1-87

[1045] Joachim Schrod. International L^{A}TEX is ready to use. TUGboat, 11(1):87– 90, April 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-1/tb27schrod.pdf.

Mittelbach:TB11-1-91

[1046] Frank Mittelbach and Rainer Schöpf. The new font family selection — User interface to standard L^{A}TEX. TUGboat, 11(1):91–97, April 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-1/tb27mitt.pdf.

Poppelier:TB11-1-98

[1047] Nico Poppelier and Johannes Braams. A style option to adapt the standard L^{A}TEX document styles to A4 paper. TUGboat, 11(1):98–103, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27poppelier.pdf.

Kelly:TB11-1-103

[1048] B Hamilton Kelly. Some macros to draw crosswords. TUGboat, 11(1):103– 119, April 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-1/tb27kelly.pdf.

Dietsche:TB11-1-120

[1049] Luzia Dietsche. Deutsche Kurzfassungen der TUGboat-Artikel. TUGboat, 11(1):120–122, April 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-1/tb27dietsche.pdf.

Partl:TB11-1-122

[1050] Hubert Partl. TEX, TUG, and Eastern Europe. TUGboat, 11(1):122–123, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27news.pdf.

Anonymous:TB11-1-123

[1051] Anonymous. Calendar. TUGboat, 11(1):123–124, April 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-1/tb27news.pdf.

Anonymous:TB11-1-125

[1052] Anonymous. GUTenberg’90, 15–17 May 1990, Toulouse, France. TUGboat, 11(1):125–126, April 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-1/tb27news.pdf.

Anonymous:TB11-1-126

[1053] Anonymous. Call for papers: SGML & TEX conference, 31 August 1990, Groningen, The Netherlands. TUGboat, 11(1):126, April 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-1/tb27news.pdf.

Beeton:TB11-1-127

[1054] Barbara Beeton. Production notes. TUGboat, 11(1):127, April 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-1/tb27late.
pdf.

Anonymous:TB11-1-128

[1055] Anonymous. Institutional members. TUGboat, 11(1):128–130, April

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-1/
tb27instmem.pdf.

Anonymous:TB11-1-131

[1056] Anonymous. TUG membership application. TUGboat, 11(1):131, April 1990.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB11-1-148

[1057] Anonymous. Index of advertisers. TUGboat, 11(1):148, April 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-1/tb27ads.
pdf.

Anonymous:1990:TML

[1058] Anonymous. TUG membership list supplement. TUGboat, 11(2):1–9, June 1990. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-2/tb28errata.pdf.

Anonymous:1990:IA

[1059] Anonymous. Index of advertisers. TUGboat, 11(2):10, June 1990. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28ads.pdf.

Anonymous:1990:TPb

[1060] Anonymous. Title page. TUGboat, 11(2):149, June 1990. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28titlepage.pdf.

Anonymous:1990:EIb

[1061] Anonymous. Editorial information. TUGboat, 11(2):150, June 1990. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28edit.pdf.

Anonymous:TB11-2-151

[1062] Anonymous. Addresses. TUGboat, 11(2):151–152, June 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-2/tb28addresses.
pdf.

Beeton:TB11-2-153

[1063] Barbara Beeton. Editorial comments. TUGboat, 11(2):153, June 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-2/tb28gendel.
pdf.

Barnhart:TB11-2-154

[1064] Elizabeth Barnhart. TEX in the production environment — questionnaire responses. TUGboat, 11(2):154–164, June 1990. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb11-2/tb28gendel.pdf.

Knuth:TB11-2-165

[1065] Donald Knuth. Exercises for TEX: The Program. TUGboat, 11(2):165–170,

June 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-
2/tb28knut.pdf.

Beebe:TB11-2-171

[1066] Nelson H. F. Beebe. Character set encoding. TUGboat, 11(2):171–175, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28beebe.pdf.

Bien:TB11-2-175

[1067] Janusz S. Bień. On standards for Computer Modern font extensions. TUGboat, 11(2):175–183, June 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-2/tb28bien.pdf.

Hoenig:TB11-2-183

[1068] Alan Hoenig. Circular reasoning: typesetting on a circle, and related issues. TUGboat, 11(2):183–190, June 1990. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb11-2/tb28hoenig.pdf.

Berendt:TB11-2-190

[1069] Gerhard Berendt. On the implementation of graphics into TEX. TUGboat, 11(2):190–194, June 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-2/tb28berendt.pdf.

Schwer:TB11-2-194

[1070] Len Schwer. Including Macintosh graphics in L^{A}TEX documents. TUGboat, 11(2):194–200, June 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-2/tb28schwer.pdf.

Pickrell:TB11-2-200

[1071] Lee S. Pickrell. Combining graphics with TEX on PC systems with laser printers, part II. TUGboat, 11(2):200–206, June 1990. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb11-2/tb28pickrell.pdf.

Childs:TB11-2-206

[1072] Bart Childs. Data General site report. TUGboat, 11(2):206, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28resources.pdf.

Lammarsch:TB11-2-207

[1073] Joachim Lammarsch. IBM VM/CMS site report. TUGboat, 11(2):207, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28resources.pdf.

Beeton:TB11-2-207

[1074] Barbara Beeton. Resources available to TEX users. TUGboat, 11(2):207–208,

June 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-
2/tb28resources.pdf.

Beeton:TB11-2-208

[1075] Barbara Beeton. A proto-TUG bibliography: Installment two. TUGboat, 11(2):208–209, June 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-2/tb28resources.pdf.

Eijkhout:TB11-2-210

[1076] Victor Eijkhout. New books on TEX. TUGboat, 11(2):210–211, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28resources.pdf.

Nieland:TB11-2-211

[1077] Ted Nieland. DECUS TEX collection — submissions wanted. TUGboat, 11(2): 211, June 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-2/tb28resources.pdf.

Hoenig:TB11-2-212

[1078] Alan Hoenig. Just plain Q&A. TUGboat, 11(2):212, June 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-2/tb28queries.
pdf.

Salomon:TB11-2-212

[1079] David Salomon. Output routines: Examples and techniques. Part II: otr techniques. TUGboat, 11(2):212–236, June 1990. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb11-2/tb28salomon.pdf.

Jeffrey:TB11-2-237

[1080] Alan Jeffrey. Lists in TEX’s mouth. TUGboat, 11(2):237–245, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28jeffrey.pdf.

Taylor:TB11-2-245

[1081] Philip Taylor. A nestable verbatim mode. TUGboat, 11(2):245–249, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28taylor.pdf.

Ha:TB11-2-250

[1082] Khanh Ha. Easy Table. TUGboat, 11(2):250–264, June 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-2/tb28ha.pdf.

Laan:TB11-2-265

[1083] Kees van der Laan. Typesetting bridge via TEX. TUGboat, 11(2):265–276, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28laan.pdf.

Damrau:TB11-2-276

[1084] Jackie Damrau. The L^{A}TEX column. TUGboat, 11(2):276–278, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28damrau.pdf.

Yap:TB11-2-279

[1085] Ken Yap. Making 35mm colour slides with SLiTEX. TUGboat, 11(2):279–280,

June 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-
2/tb28yap.pdf.

Denk:TB11-2-280

[1086] Georg Denk. An easy way to make slides with L^{A}TEX. TUGboat, 11(2):280– 283, June 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-2/tb28denk.pdf.

Schopf:TB11-2-284

[1087] Rainer Schöpf. A new implementation of the L^{A}TEX verbatim and verbatim* environments. TUGboat, 11(2):284–296, June 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-2/tb28schopf.
pdf.

Mittelbach:TB11-2-297

[1088] Frank Mittelbach and Rainer Schöpf. Reprint: The new font family selection — User interface to standard L^{A}TEX. TUGboat, 11(2):297–305, June 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-2/tb28mitt.
pdf.

Dietsche:TB11-2-305

[1089] Luzia Dietsche. Deutsche Kurzfassungen der TUGboat-Artikel. TUGboat, 11(2):305–307, June 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-2/tb28dietsche.pdf.

Anonymous:TB11-2-308

[1090] Anonymous. Calendar. TUGboat, 11(2):308–309, June 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-2/tb28news.pdf.

Beeton:TB11-2-309

[1091] Barbara Beeton. Production notes. TUGboat, 11(2):309–310, June 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-2/tb28news.
pdf.

Anonymous:TB11-2-310

[1092] Anonymous. New publications and software available through TUG. TUGboat, 11(2):310–313, June 1990. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb11-2/tb28news.pdf.

Anonymous:TB11-2-313

[1093] Anonymous. Russians visit TUG headquarters. TUGboat, 11(2):313, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28news.pdf.

Anonymous:TB11-2-314

[1094] Anonymous. Institutional members. TUGboat, 11(2):314–316, June

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-2/
tb28instmem.pdf.

Anonymous:TB11-2-317

[1095] Anonymous. TUG membership application. TUGboat, 11(2):317, June 1990.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB11-2-M-2

[1096] Anonymous. TUG Membership List Supplement. TUGboat, 11(2):M–1–M–9, June 1990. ISSN 0896-3207.

Anonymous:TB11-2-M

[1097] Anonymous. Index of advertisers. TUGboat, 11(2):M–10, June 1990. ISSN 0896-3207.

Anonymous:1990:TPc

[1098] Anonymous. Title page. TUGboat, 11(3):333, September 1990. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29titlepage.pdf.

Anonymous:1990:EIc

[1099] Anonymous. Editorial information. TUGboat, 11(3):334, September 1990. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-3/tb29edit.pdf.

Beebe:TB11-3-335

[1100] Nelson H. F. Beebe. President’s introduction. TUGboat, 11(3):335–336, September 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-3/tb29beebe.pdf.

Mittelbach:TB11-3-337

[1101] Frank Mittelbach. E-TEX: Guidelines for future TEX extensions. TUGboat, 11(3):337–345, September 1990. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb11-3/tb29mitt.pdf.

Hamano:TB11-3-346

[1102] Hisato Hamano. Vertical typesetting with TEX. TUGboat, 11(3):346–352, September 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-3/tb29hamano.pdf.

Miyabe:TB11-3-353

[1103] Yoshiyuki Miyabe, H. Ohta, and K. Tsuga. Structured document preparation system AutoLayouter. TUGboat, 11(3):353–358, September 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-3/tb29miyabe.
pdf.

Hendrickson:TB11-3-359

[1104] Amy Hendrickson. Getting TEXnical: Insights into TEX macro writing techniques. TUGboat, 11(3):359–370, September 1990. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb11-3/tb29hendrickson.pdf.

Fulling:TB11-3-371

[1105] Stephen Alex Fulling. Where’s the Greek shift key? TUGboat, 11(3):371–372, September 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-3/tb29fulling.pdf.

Beck:TB11-3-373

[1106] Micah Beck and A. Siegel. TransFig: Portable graphics for TEX. TUGboat, 11(3):373–380, September 1990. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb11-3/tb29beck.pdf.

Greene:TB11-3-381

[1107] Andrew Marc Greene. BASIX: An interpreter written in TEX. TUGboat, 11(3): 381–392, September 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-3/tb29greene.pdf.

Gibson:TB11-3-393

[1108] Helen Gibson. A noddy’s guide to high quality TEX production. TUGboat, 11(3):393–402, September 1990. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb11-3/tb29gibson.pdf.

Adams:TB11-3-403

[1109] Robert A. Adams. Problems on the TEX/PostScript/graphics interface. TUGboat, 11(3):403–408, September 1990. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb11-3/tb29adams.pdf.

Bechtolsheim:TB11-3-409

[1110] Stephan von Bechtolsheim. TEX in practice: Comments on a 4-volume, 1400page series on TEX. TUGboat, 11(3):409–412, September 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-3/tb29bechtolsheim.
pdf.

LaFrenz:TB11-3-413

[1111] Mimi L. LaFrenz. Textbook publishing — 1990 and beyond. TUGboat, 11(3): 413–416, September 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-3/tb29lafrenz.pdf.

Ness:TB11-3-417

[1112] David Ness. Diagnosing TEX errors with a preprocessor. TUGboat, 11(3): 417–420, September 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-3/tb29ness.pdf.

Williams:TB11-3-421

[1113] Linda Williams and Linda Hall. Increased efficiency using advanced EDT editing features. TUGboat, 11(3):421–424, September 1990. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb11-3/tb29williams.pdf.

Martin:TB11-3-425

[1114] Charles R. Martin. TEX for TEXnical typists. TUGboat, 11(3):425–428, September 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-3/tb29martin.pdf.

Vulis:TB11-3-429

[1115] Michael Vulis. VTEX enhancements to the TEX language. TUGboat, 11(3): 429–434, September 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-3/tb29vulis.pdf.

Hoenig:TB11-3-435

[1116] Alan Hoenig. A constructed Dürer alphabet. TUGboat, 11(3):435–438, September 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-3/tb29hoenig.pdf.

Dietsche:TB11-3-439

[1117] Luzia Dietsche. Deutsche Kurzfassungen der TUGboat-Artikel. TUGboat, 11(3):439–442, September 1990. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb11-3/tb29dietsche.pdf.

Beebe:TB11-3-442

[1118] Nelson H. F. Beebe. TEX 3.0 and 2.0. TUGboat, 11(3):442–443, September 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-3/tb29status.pdf.

Lammarsch:TB11-3-444

[1119] Joachim Lammarsch. dante, Deutschsprachige Anwendervereinigung TEX e.V. TUGboat, 11(3):444, September 1990. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb11-3/tb29reports.pdf.

Mittelbach:TB11-3-444

[1120] Frank Mittelbach. L^{A}TEX 2.10. TUGboat, 11(3):444, September 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-3/tb29status.
pdf.

Gaulle:TB11-3-445

[1121] Bernard Gaulle. News from and about GUTenberg. TUGboat, 11(3):445, September 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-3/tb29reports.pdf.

Laan:TB11-3-446

[1122] Kees van der Laan. NTG’s second year. TUGboat, 11(3):446–447, September

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29reports.pdf.

Graham:TB11-3-447

[1123] Roswitha Graham and Jan Michael Rynning. Report from the Nordic TEX Users Group. TUGboat, 11(3):447–448, September 1990. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb11-3/tb29reports.pdf.

Clark:TB11-3-448

[1124] Malcolm Clark. TEX in the UK. TUGboat, 11(3):448–449, September

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29reports.pdf.

Clark:TB11-3-450

[1125] Malcolm Clark. TEX in Europe. TUGboat, 11(3):450–451, September

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29reports.pdf.

Platt:TB11-3-451

[1126] Craig Platt. MVS site report. TUGboat, 11(3):451, September 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-3/tb29site.
pdf.

Kellerman:TB11-3-451

[1127] David Kellerman. VMS site report. TUGboat, 11(3):451–452, September

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29site.pdf.

Childs:TB11-3-452

[1128] Bart Childs. Data General site report. TUGboat, 11(3):452, September

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29site.pdf.

Crawford:TB11-3-453

[1129] John Crawford. Prime 50 series site report. TUGboat, 11(3):453, September

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29site.pdf.

MacKay:TB11-3-453

[1130] Pierre A. MacKay. unixTEX site report. TUGboat, 11(3):453–454, September

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29site.pdf.

Lammarsch:TB11-3-454

[1131] Joachim Lammarsch. VM/CMS site report. TUGboat, 11(3):454–455, Septem-

ber 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-
3/tb29site.pdf.

Beeton:TB11-3-455

[1132] Barbara Beeton. Report from the question and answer session. TUGboat, 11 (3):455–458, September 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-3/tb29queries.pdf.

Anonymous:TB11-3-458

[1133] Anonymous. List of participants at the Texas A&M meeting (June 17–20, 1990). TUGboat, 11(3):458–461, September 1990. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb11-3/tb29participants.pdf.

Anonymous:TB11-3-462

[1134] Anonymous. Calendar. TUGboat, 11(3):462, September 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-3/tb29calendar.
pdf.

Anonymous:TB11-3-463

[1135] Anonymous. Institutional members. TUGboat, 11(3):463–465, September

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29instmem.pdf.

Anonymous:TB11-3-466

[1136] Anonymous. TUG membership application. TUGboat, 11(3):466, September

1990. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB11-3-479

[1137] Anonymous. Index of advertisers. TUGboat, 11(3):479, September

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-3/
tb29ads.pdf.

Anonymous:1990:TPd

[1138] Anonymous. Title page. TUGboat, 11(4):481, December 1990. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30titlepage.pdf.

Anonymous:1990:EId

[1139] Anonymous. Editorial information. TUGboat, 11(4):482, December 1990. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-4/tb30edit.pdf.

Anonymous:TB11-4-483

[1140] Anonymous. Addresses. TUGboat, 11(4):483–484, November 1990. ISSN 0896-3207.

Beebe:TB11-4-485

[1141] Nelson H. F. Beebe. From the President. TUGboat, 11(4):485–487, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30beebe-pres.pdf.

Anonymous:TB11-4-488

[1142] Anonymous. The future of TEX. TUGboat, 11(4):488, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30futuretex.pdf.

Knuth:TB11-4-489

[1143] Donald E. Knuth. The future of TEX and . TUGboat, 11(4):489, November 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-4/tb30knut.pdf.

Beebe:TB11-4-490

[1144] Nelson H. F. Beebe. Comments on the future of TEX and . TUGboat, 11(4):490–494, November 1990. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb11-4/tb30beebe-future.pdf.

Beeton:TB11-4-494

[1145] Barbara Beeton. Editorial comments. TUGboat, 11(4):494–496, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30beet.pdf.

Knuth:TB11-4-497

[1146] Donald Knuth. Arthur Lee Samuel, 1901–1990. TUGboat, 11(4):497–498, November 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-4/tb30knut-samuel.pdf.

Knuth:TB11-4-499

[1147] Donald Knuth. Answers to Exercises for TEX: The Program. TUGboat, 11(4): 499–511, November 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-4/tb30knut-exercises.pdf.

Fox:TB11-4-511

[1148] Jim Fox. Webless literate programming. TUGboat, 11(4):511–513, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30fox.pdf.

Stokes:TB11-4-513

[1149] Harold T. Stokes. A TEX previewer for “slow” terminals. TUGboat, 11(4): 513–514, November 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-4/tb30stokes.pdf.

Ferguson:TB11-4-514

[1150] Michael Ferguson. Report on multilingual activities. TUGboat, 11(4):514–515, November 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-4/tb30ferguson.pdf.

Anonymous:TB11-4-516

[1151] Anonymous. Extended TEX font encoding scheme — Latin, Cork, September 12, 1990. TUGboat, 11(4):516, November 1990. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb11-4/tb30ferguson.pdf.

Berry:TB11-4-517

[1152] Karl Berry. Filenames for fonts. TUGboat, 11(4):517–520, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30berry.pdf.

Haralambous:TB11-4-520

[1153] Yannis Haralambous. Arabic, Persian and Ottoman TEX for Mac and PC. TUGboat, 11(4):520–524, November 1990. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb11-4/tb30hara.pdf.

Yanai:TB11-4-525

[1154] Shimon Yanai and Daniel M. Berry. Environment for translating to PostScript. TUGboat, 11(4):525–541, November 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-4/tb30yanai.pdf.

Tofsted:TB11-4-542

[1155] David Tofsted. An improved chess font. TUGboat, 11(4):542–544, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30tofsted.pdf.

Hosek:TB11-4-545

[1156] Don Hosek. TEX output devices (with charts). TUGboat, 11(4):545–569, November 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-4/tb30hosek.pdf.

Hosek:TB11-4-570

[1157] Don Hosek. Report from the DVI Driver Standards Committee. TUGboat, 11(4):570, November 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-4/tb30resources.pdf.

Beeton:TB11-4-570

[1158] Barbara Beeton. Updates from all over. TUGboat, 11(4):570–571, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30resources.pdf.

Berry:TB11-4-571

[1159] Karl Berry. eplain. TUGboat, 11(4):571–572, November 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-4/tb30resources.
pdf.

Eijkhout:TB11-4-572

[1160] Victor Eijkhout. New Books on TEX. TUGboat, 11(4):572–573, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30resources.pdf.

Beeton:TB11-4-573

[1161] Barbara Beeton. A proto-TUG bibliography: Installment three. TUGboat, 11 (4):573–576, November 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-4/tb30resources.pdf.

Nieland:TB11-4-576

[1162] Ted Nieland. The 1990 DECUS TEX collection. TUGboat, 11(4):576–577, November 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb11-4/tb30resources.pdf.

Hosek:TB11-4-578

[1163] Don Hosek. The IVRITEX Mailing List. TUGboat, 11(4):578, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30resources.pdf.

Lammarsch:TB11-4-578

[1164] Joachim Lammarsch. IBM VM/CMS site report. TUGboat, 11(4):578, Novem-

ber 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-
4/tb30resources.pdf.

Neuwirth:TB11-4-578

[1165] Erich Neuwirth. TEX implementations for IBM PCs: comparative timings. TUGboat, 11(4):578–580, November 1990. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb11-4/tb30neuwirth.pdf.

Durst:TB11-4-580

[1166] Lincoln Durst. Long-winded endnotes and exercises with hints or solutions. TUGboat, 11(4):580–588, November 1990. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb11-4/tb30durst.pdf.

Salomon:TB11-4-588

[1167] David Salomon. Output routines: Examples and techniques. Part III: insertions. TUGboat, 11(4):588–605, November 1990. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb11-4/tb30salomon.pdf.

Eijkhout:TB11-4-605

[1168] Victor Eijkhout. A new editor. TUGboat, 11(4):605, November 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-4/tb30downes.
pdf.

Downes:TB11-4-605

[1169] Michael Downes. Line breaking in \unhboxed text. TUGboat, 11(4): 605–612, November 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-4/tb30downes.pdf.

Maus:TB11-4-612

[1170] Sonja Maus. Looking ahead for a <box>. TUGboat, 11(4):612–613, Novem-

ber 1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-
4/tb30maus.pdf.

Eijkhout:TB11-4-613

[1171] Victor Eijkhout. An indentation scheme. TUGboat, 11(4):613–616, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30eijkhout-indent.pdf.

Eijkhout:TB11-4-616

[1172] Victor Eijkhout. A \parskip scheme. TUGboat, 11(4):616–619, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30eijkhout-parskip.pdf.

Whitney:TB11-4-620

[1173] Ron Whitney. Sanitizing control sequences under \write. TUGboat, 11(4): 620–622, November 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-4/tb30whitney.pdf.

Lavagnino:TB11-4-623

[1174] John Lavagnino and Dominik Wujastyk. An Overview of EDMAC: A plain TEX format for critical editions. TUGboat, 11(4):623–643, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30lava.pdf.

Damrau:TB11-4-644

[1175] Jackie Damrau. The L^{A}TEX column. TUGboat, 11(4):644, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30damrau.pdf.

Geyer-Schulz:TB11-4-644

[1176] Andreas Geyer-Schulz, Josef Matulka, and Gustaf Neumann. A L^{A}TEX document style option for typesetting APL. TUGboat, 11(4):644–651, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30geyer.pdf.

Love:TB11-4-652

[1177] David Love. Experiments in TEXnicolour — A SLiTEX sub-style for Colour Printers. TUGboat, 11(4):652–656, November 1990. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb11-4/tb30love.pdf.

Mittelbach:TB11-4-657

[1178] Frank Mittelbach. Footnotes in a multi-column layout. TUGboat, 11(4): 657–662, November 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-4/tb30mitt.pdf.

Dietsche:TB11-4-663

[1179] Luzia Dietsche. Deutsche Kurzfassungen der TUGboat-Artikel. TUGboat, 11 (4):663–665, November 1990. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb11-4/tb30abstracts.pdf.

Akwai:TB11-4-665

[1180] Peter C. Akwai. TEX drug bug. TUGboat, 11(4):665, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30letters.pdf.

Anonymous:TB11-4-666

[1181] Anonymous. Calendar. TUGboat, 11(4):666, November 1990. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb11-4/tb30calendar.
pdf.

Anonymous:TB11-4-667

[1182] Anonymous. Call for papers: Tenth meeting, DANTE, Vienna, 20–22 February 1991. TUGboat, 11(4):667, November 1990. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb11-4/tb30calendar.pdf.

Anonymous:TB11-4-667-2

[1183] Anonymous. TEX91/Congres GUTenberg’91, Paris, 23–26 September 1991. TUGboat, 11(4):667, November 1990. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb11-4/tb30calendar.pdf.

Anonymous:TB11-4-668

[1184] Anonymous. Call for papers: RIDT 91 (raster imaging/digital typography), Boston, 15–16 October 1991 . TUGboat, 11(4):668, November 1990. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb11-4/tb30calendar.pdf.

Anonymous:TB11-4-669

[1185] Anonymous. TUG Bylaws. TUGboat, 11(4):669–676, November 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-4/tb30bylaws.
pdf.

Anonymous:TB11-4-676

[1186] Anonymous. Institutional members. TUGboat, 11(4):676–678, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30instmem.pdf.

Anonymous:TB11-4-679

[1187] Anonymous. Coming next issue. TUGboat, 11(4):679, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30notes.pdf.

Beeton:TB11-4-679

[1188] Barbara Beeton. Production notes. TUGboat, 11(4):679, November

1990. ISSN 0896-3207. URL https://tug.org/TUGboat/tb11-4/
tb30notes.pdf.

Anonymous:TB11-4-681

[1189] Anonymous. TUG membership application. TUGboat, 11(4):681, November

1990. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB11-4-693

[1190] Anonymous. Index of advertisers. TUGboat, 11(4):693, November 1990. ISSN

0896-3207. URL https://tug.org/TUGboat/tb11-4/tb30ads.
pdf.

Anonymous:1990:CTE

[1191] Anonymous. Computers & Typesetting: Errata, supplement. TUGboat, 11(4): 991, December 1990. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb11-4/tb30bugs.pdf.

Anonymous:1990:CPT

[1192] Anonymous. Call for papers: TUG Twelfth Annual Meeting, suburban Boston, 15–18 July 1991. TUGboat, 11(4):c3, December 1990. CODEN ???? ISSN 0896-3207.

Anonymous:TB11-4-cover-3

[1193] Anonymous. Call for papers: TUG Twelfth Annual Meeting, suburban Boston, 15–18 July 1991. TUGboat, 11(4):Cover 3, November 1990. ISSN 0896-3207.

Anonymous:1991:TPa

[1194] Anonymous. Title page. TUGboat, 12(1):1, March 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31titlepage.pdf.

Beeton:TB12-1-3

[1195] Barbara Beeton. Introduction. TUGboat, 12(1):3–4, March 1991. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb12-1/tb31beet.pdf.

Neuwirth:TB12-1-5

[1196] Erich Neuwirth. Quick and dirty databases with nice output: AWK and TEX. TUGboat, 12(1):5–7, March 1991. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb12-1/tb31eneuwirth.pdf.

Detig:TB12-1-8

[1197] Christine Detig. TEX & hypertext — The future of electronic publishing? TUGboat, 12(1):8–12, March 1991. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb12-1/tb31detig.pdf.

Carr:TB12-1-13

[1198] Les Carr, Sebastian Rahtz, and Wendy Hall. Experiments with TEX and hyperactivity. TUGboat, 12(1):13–20, March 1991. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb12-1/tb31carr.pdf.

Braams:TB12-1-21

[1199] Johannes Braams, Victor Eijkhout, and Nico Poppelier. The Dutch national L^{A}TEX effort. TUGboat, 12(1):21–24, March 1991. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb12-1/tb31braams.pdf.

Clark:TB12-1-25

[1200] Adrian Clark. Documenting a TEX archive. TUGboat, 12(1):25–27, March

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31aclark-archive.pdf.

Kneser:TB12-1-28

[1201] Thomas Kneser. L^{A}TEX-paragraphs floating around figures. TUGboat, 12 (1):28–30, March 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-1/tb31kneser.pdf.

Eijkhout:TB12-1-31

[1202] Victor Eijkhout. The document style designer as a separate entity. TUGboat, 12(1):31–34, March 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-1/tb31eijkhout.pdf.

Taylor:TB12-1-35

[1203] Philip Taylor. Improving the æsthetics of mixed-font documents. TUGboat, 12(1):35–44, March 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-1/tb31taylor.pdf.

Wittbecker:TB12-1-45

[1204] Alan Wittbecker. ArchıTEX, a preliminary international page pattern maker. TUGboat, 12(1):45–57, March 1991. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb12-1/tb31wittbecker.pdf.

Sowa:TB12-1-58

[1205] Friedhelm Sowa. Integration of graphics into TEX. TUGboat, 12(1):58–63, March 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-1/tb31sowa.pdf.

Murphy:TB12-1-64

[1206] Timothy Murphy. PostScript, QuickDraw, TEX. TUGboat, 12(1):64–65, March

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31murphy.pdf.

Maclenan:TB12-1-66

[1207] M. P. Maclenan and G. M. Burns. An approach to drawing circuit diagrams for text books. TUGboat, 12(1):66–69, March 1991. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb12-1/tb31maclenan.pdf.

Nicole:TB12-1-70

[1208] Olivier Nicole. A graphic driver to interface statistical software S and PICTEX. TUGboat, 12(1):70–73, March 1991. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb12-1/tb31nicole.pdf.

Mittelbach:TB12-1-74

[1209] Frank Mittelbach and Rainer Schöpf. Towards L^{A}TEX 3.0. TUGboat, 12(1):74– 79, March 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-1/tb31mitt.pdf.

Kelly:TB12-1-80

[1210] Brian Hamilton Kelly. Public-domain, documented implementations of TEX and for VAX/VMS. TUGboat, 12(1):80–83, March

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31hamiltonkelly.pdf.

Clark:TB12-1-84

[1211] Malcolm Clark. Post congress tristesse. TUGboat, 12(1):84–89, March

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31mclark.pdf.

Laan:TB12-1-90

[1212] Kees van der Laan. SGML (, TEX and …). TUGboat, 12(1):90–104, March

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31laan.pdf.

Poppelier:TB12-1-105

[1213] N. A. F. M. Poppelier. SGML and TEX in scientific publishing. TUGboat, 12 (1):105–109, March 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-1/tb31poppelier.pdf.

Hendrickson:TB12-1-110

[1214] Amy Hendrickson. Getting TEXnical: Insights into TEX macro writing techniques. TUGboat, 12(1):110, March 1991. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb12-1/tb31hendrickson.pdf.

Wonneberger:TB12-1-111

[1215] Reinhard Wonneberger and Frank Mittelbach. BibTEX reconsidered. TUGboat, 12(1):111–124, March 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-1/tb31wonn.pdf.

Hoenig:TB12-1-125

[1216] Alan Hoenig. Labelling figures in TEX documents. TUGboat, 12(1):125– 128, March 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-1/tb31hoenig.pdf.

Haralambous:TB12-1-129

[1217] Yannis Haralambous. Typesetting old German: Fraktur, Schwabacher, Gotisch and initials. TUGboat, 12(1):129–138, March 1991. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb12-1/tb31hara.pdf.

Searcoid:TB12-1-139

[1218] M/iche/al /O Searc/oid. The Irish alphabet. TUGboat, 12(1):139–148, March

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31searcoid.pdf.

Guenther:TB12-1-149

[1219] Dean Guenther and Janene Winter. An international phonetic alphabet. TUGboat, 12(1):149–156, March 1991. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb12-1/tb31guenther.pdf.

Clark:TB12-1-157

[1220] Adrian F. Clark. Practical halftoning with TEX. TUGboat, 12(1):157–165, March 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-1/tb31clark.pdf.

Barden:TB12-1-166

[1221] Angela Barden. Some TEX manuals. TUGboat, 12(1):166–170, March

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31barden.pdf.

Neuwirth:TB12-1-171

[1222] Konrad Neuwirth. TEX in schools: Just say no. TUGboat, 12(1):171–174, March 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-1/tb31kneuwirth.pdf.

Anonymous:TB12-1-175

[1223] Anonymous. Calendar. TUGboat, 12(1):175, March 1991. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb12-1/tb31announce.pdf.

Anonymous:TB12-1-176

[1224] Anonymous. Nominations. TUGboat, 12(1):176, March 1991. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb12-1/tb31announce.
pdf.

Vesely:TB12-1-176

[1225] Jiří Veselý. TEX in Czechoslovakia. TUGboat, 12(1):176–177, March

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31announce.pdf.

Sullivan:TB12-1-177

[1226] Wayne Sullivan. sbTEX and sbMF. TUGboat, 12(1):177, March 1991. ISSN

0896-3207. URL https://tug.org/TUGboat/tb12-1/tb31announce.
pdf.

Raman:TB12-1-178

[1227] T. V. Raman. TEX TALK. TUGboat, 12(1):178, March 1991. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb12-1/tb31announce.pdf.

Anonymous:TB12-1-179

[1228] Anonymous. Institutional members. TUGboat, 12(1):179–181, March

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
tb31instmem.pdf.

Anonymous:TB12-1-182

[1229] Anonymous. TUG membership applications. TUGboat, 12(1):182–183, March

1991. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB12-1-200

[1230] Anonymous. Index of advertisers. TUGboat, 12(1):200, March 1991. ISSN

0896-3207. URL https://tug.org/TUGboat/tb12-1/tb31ads.
pdf.

Anonymous:1991:FCa

[1231] Anonymous. Front cover. TUGboat, 12(1):c1, March 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
cover1.pdf.

Anonymous:1991:IFCa

[1232] Anonymous. Inside front cover. TUGboat, 12(1):c2, March 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-1/
cover2.pdf.

Anonymous:TB12-2-3

[1233] Anonymous. TUG Membership List. TUGboat, 12(2):3–117, June 1991. ISSN 0896-3207.

Anonymous:TB12-2-118

[1234] Anonymous. TEX consulting and production services. TUGboat, 12(2):118, June 1991. ISSN 0896-3207.

Anonymous:TB12-2-119

[1235] Anonymous. General comments. TUGboat, 12(2):119, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-119-2

[1236] Anonymous. TEX Users Group. TUGboat, 12(2):119–120, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-120

[1237] Anonymous. CSTUG. TUGboat, 12(2):120, June 1991. ISSN 0896-3207. Anonymous:TB12-2-120-2

[1238] Anonymous. CyrTUG. TUGboat, 12(2):120, June 1991. ISSN 0896-3207. Anonymous:TB12-2-120-3

[1239] Anonymous. DANTE e.V. TUGboat, 12(2):120–121, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-121

[1240] Anonymous. GUTenberg. TUGboat, 12(2):121–122, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-122

[1241] Anonymous. ITALIC. TUGboat, 12(2):122, June 1991. ISSN 0896-3207. Anonymous:TB12-2-122-2

[1242] Anonymous. JTUG — TUG Japan. TUGboat, 12(2):122, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-123

[1243] Anonymous. Nordic TEX Group. TUGboat, 12(2):123, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-123-2

[1244] Anonymous. NTG. TUGboat, 12(2):123–124, June 1991. ISSN 0896-3207. Anonymous:TB12-2-124

[1245] Anonymous. UK TUG. TUGboat, 12(2):124, June 1991. ISSN 0896-3207. Anonymous:TB12-2-125

[1246] Anonymous. YUNUS – Turkish TEX users. TUGboat, 12(2):125, June 1991.

ISSN 0896-3207.

Anonymous:TB12-2-125-2

[1247] Anonymous. Other international contacts. TUGboat, 12(2):125, June 1991. ISSN 0896-3207.

Barnhart:TB12-2-126

[1248] Elizabeth M. Barnhart. TEX L.U.G.s. TUGboat, 12(2):126, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-127

[1249] Anonymous. Generic format. TUGboat, 12(2):127–128, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-128

[1250] Anonymous. Acorn Archimedes. TUGboat, 12(2):128, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-128-2

[1251] Anonymous. Amiga. TUGboat, 12(2):128, June 1991. ISSN 0896-3207. Anonymous:TB12-2-128-3

[1252] Anonymous. Atari ST. TUGboat, 12(2):128, June 1991. ISSN 0896-3207. Anonymous:TB12-2-128-4

[1253] Anonymous. Data General. TUGboat, 12(2):128, June 1991. ISSN 0896-3207. Anonymous:TB12-2-128-5

[1254] Anonymous. DEC-20 — TOPS-20. TUGboat, 12(2):128, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-128-6

[1255] Anonymous. IBM mainframes — MVS. TUGboat, 12(2):128–129, June 1991.

ISSN 0896-3207.

Anonymous:TB12-2-129

[1256] Anonymous. IBM mainframes — VM/CMS. TUGboat, 12(2):129, June 1991.

ISSN 0896-3207.

Anonymous:TB12-2-129-2

[1257] Anonymous. IBM PC and compatibles — MS-DOS. TUGboat, 12(2):129,

June 1991. ISSN 0896-3207.

Anonymous:TB12-2-130

[1258] Anonymous. Macintosh. TUGboat, 12(2):130, June 1991. ISSN 0896-3207. Anonymous:TB12-2-130-2

[1259] Anonymous. Prime. TUGboat, 12(2):130, June 1991. ISSN 0896-3207. Anonymous:TB12-2-130-3

[1260] Anonymous. Unix. TUGboat, 12(2):130, June 1991. ISSN 0896-3207.

Anonymous:TB12-2-130-4

[1261] Anonymous. VAX/VMS. TUGboat, 12(2):130–131, June 1991. ISSN 0896-

3207.

Nieland:TB12-2-131

[1262] Ted Nieland. Announcing the DECUS TEX Collection, February 1991. TUG-

boat, 12(2):131–132, June 1991. ISSN 0896-3207.

Hosek:TB12-2-133

[1263] Don Hosek. TEX Output Devices. TUGboat, 12(2):133–158, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-159

[1264] Anonymous. General observations. TUGboat, 12(2):159, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-160

[1265] Anonymous. Aston (uk.ac.tex). TUGboat, 12(2):160, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-160-2

[1266] Anonymous. Clarkson (L^{A}TEX styles). TUGboat, 12(2):160, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-160-3

[1267] Anonymous. e-math.ams.com. TUGboat, 12(2):160, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-161

[1268] Anonymous. Heidelberg (DHDURZ1.Bitnet). TUGboat, 12(2):161, June

1991. ISSN 0896-3207.

Anonymous:TB12-2-161-2

[1269] Anonymous. june.cs.washington.edu. TUGboat, 12(2):161, June

1991. ISSN 0896-3207.

Anonymous:TB12-2-161-3

[1270] Anonymous. labrea.stanford.edu. TUGboat, 12(2):161, June 1991.

ISSN 0896-3207.

Anonymous:TB12-2-161-4

[1271] Anonymous. Nijmegen (HEARN.Bitnet). TUGboat, 12(2):161, June 1991. ISSN 0896-3207.

Anonymous:TB12-2-162

[1272] Anonymous. SHSU (niord.shsu.edu, SHSU.bitnet). TUGboat, 12

(2):162, June 1991. ISSN 0896-3207.

Anonymous:TB12-2-162-2

[1273] Anonymous. SPAN/DECnet (Italy). TUGboat, 12(2):162, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-162-3

[1274] Anonymous. Stuttgart (rusmv1.rus.uni-stuttgart.de). TUGboat,

12(2):162–163, June 1991. ISSN 0896-3207.

Anonymous:TB12-2-163

[1275] Anonymous. Utrecht (archive.cs.ruu.nl). TUGboat, 12(2):163, June

1991. ISSN 0896-3207.

Anonymous:TB12-2-163-2

[1276] Anonymous. ymir.claremont.edu. TUGboat, 12(2):163, June 1991.

ISSN 0896-3207.

Anonymous:TB12-2-163-3

[1277] Anonymous. Other archives containing TEX material. TUGboat, 12(2):163–

164, June 1991. ISSN 0896-3207.

Damerell:TB12-2-164

[1278] R. M. Damerell. Fetching files from the Aston archive. TUGboat, 12(2):164–

169, June 1991. ISSN 0896-3207.

Anonymous:TB12-2-170

[1279] Anonymous. General comments. TUGboat, 12(2):170–171, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-171

[1280] Anonymous. comp.text.tex. TUGboat, 12(2):171, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-171-2

[1281] Anonymous. ELLHNIKA (Greek). TUGboat, 12(2):171, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-172

[1282] Anonymous. GUT (GUTenberg). TUGboat, 12(2):172, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-172-2

[1283] Anonymous. INFO-TeX. TUGboat, 12(2):172, June 1991. ISSN 0896-3207.

Anonymous:TB12-2-172-3

[1284] Anonymous. IVRITEX (Semitic languages). TUGboat, 12(2):172, June 1991.

ISSN 0896-3207.

Anonymous:TB12-2-172-4

[1285] Anonymous. LaTeX-help. TUGboat, 12(2):172, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-173

[1286] Anonymous. RUSTeX-L (Russian). TUGboat, 12(2):173, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-173-2

[1287] Anonymous. TeX-D-L (German). TUGboat, 12(2):173, June 1991. ISSN

0896-3207.

Anonymous:TB12-2-173-3

[1288] Anonymous. TeX-EURO. TUGboat, 12(2):173, June 1991. ISSN 0896-3207. Anonymous:TB12-2-173-4

[1289] Anonymous. TeXhax. TUGboat, 12(2):173–174, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-174

[1290] Anonymous. TEXMaG. TUGboat, 12(2):174, June 1991. ISSN 0896-3207. Anonymous:TB12-2-174-2

[1291] Anonymous. TeX-Pubs. TUGboat, 12(2):174, June 1991. ISSN 0896-3207. Anonymous:TB12-2-174-3

[1292] Anonymous. UKTeX. TUGboat, 12(2):174–175, June 1991. ISSN 0896-3207. Anonymous:TB12-2-175

[1293] Anonymous. TeX-NL (Dutch). TUGboat, 12(2):175, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-175-2

[1294] Anonymous. YUNUS (Turkish). TUGboat, 12(2):175, June 1991. ISSN 0896-

3207.

Anonymous:TB12-2-176

[1295] Anonymous. Publications about TEX and typography. TUGboat, 12(2):176– 183, June 1991. ISSN 0896-3207.

Anonymous:TB12-2-183

[1296] Anonymous. Publications prepared with TEX. TUGboat, 12(2):183–194, June 1991. ISSN 0896-3207.

Anonymous:1991:TPb

[1297] Anonymous. Title page. TUGboat, 12(2):201, June 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32titlepage.pdf.

Anonymous:1991:EI

[1298] Anonymous. Editorial information. TUGboat, 12(2):202, June 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32edit.pdf.

Anonymous:TB12-2-203

[1299] Anonymous. Addresses. TUGboat, 12(2):203–204, June 1991. ISSN 0896-3207.

Beebe:TB12-2-205

[1300] Nelson H. F. Beebe. President’s introduction. TUGboat, 12(2):205–208, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32pres.pdf.

Beeton:TB12-2-208

[1301] Barbara Beeton. Editorial comments. TUGboat, 12(2):208–211, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32beet.pdf.

Obermiller:TB12-2-211

[1302] Walter Obermiller. TEX in Germany. TUGboat, 12(2):211–212, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32obermiller.pdf.

Malyshev:TB12-2-212

[1303] Basil Malyshev, Alexander Samarin, and Dimitri Vulis. Russian TEX. TUGboat, 12(2):212–214, June 1991. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb12-2/tb32malyshev.pdf.

Krstev:TB12-2-215

[1304] Cvetana Krstev. Serbo-Croatian hyphenation: a TEX point of view. TUGboat, 12(2):215–223, June 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-2/tb32krstev.pdf.

Haralambous:TB12-2-224

[1305] Yannis Haralambous. On TEX and Greek. … TUGboat, 12(2):224–226, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32hara.pdf.

Jalbert:TB12-2-227

[1306] François Jalbert. JemTEX 2.00 available for Japanese. TUGboat, 12(2):227,

June 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-
2/tb32jalbert.pdf.

Jeffrey:TB12-2-227

[1307] Alan Jeffrey. Labelled diagrams in . TUGboat, 12(2):227–229,

June 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-
2/tb32jeffrey.pdf.

Fossmeier:TB12-2-229

[1308] Reinhard Fößmeier. X bitmaps in TEX. TUGboat, 12(2):229–232, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32fossmeier.pdf.

Schrod:TB12-2-232

[1309] Joachim Schrod. Report on the DVI Driver Standard. TUGboat, 12(2):232– 233, June 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-2/tb32schrod.pdf.

Berry:TB12-2-233

[1310] Karl Berry and Kathy Hargreaves. Review of 3 : 16 Bible Texts Illuminated. TUGboat, 12(2):233–235, June 1991. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb12-2/tb32reviews.pdf.

Poppelier:TB12-2-235

[1311] Nico Poppelier. Review of L^{A}TEX for engineers and scientists. TUGboat, 12 (2):235–236, June 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-2/tb32reviews.pdf.

Hoenig:TB12-2-237

[1312] Alan Hoenig. Just plain Q&A. TUGboat, 12(2):237–238, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32hoenig.pdf.

Salomon:TB12-2-238

[1313] David Salomon. The \if, \fx and \ifcat comparisons. TUGboat, 12 (2):238–247, June 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-2/tb32salomon.pdf.

Durst:TB12-2-248

[1314] Lincoln Durst. Some tools for making indexes: Part I. TUGboat, 12(2):248– 252, June 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-2/tb32durst.pdf.

Eijkhout:TB12-2-253

[1315] Victor Eijkhout. The structure of the TEX processor. TUGboat, 12(2):253–256,

June 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-
2/tb32eijkhout-structure.pdf.

Beeton:TB12-2-257

[1316] Barbara Beeton. Initiation rites. TUGboat, 12(2):257–258, June 1991. ISSN

0896-3207. URL https://tug.org/TUGboat/tb12-2/tb32beet-
initiation.pdf.

Mittelbach:TB12-2-259

[1317] Frank Mittelbach. Solution to the riddle from TUGboat 1, no. 1(4). TUGboat, 12(2):259, June 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-2/tb32mitt.pdf.

Eijkhout:TB12-2-260

[1318] Victor Eijkhout. The bag of tricks. TUGboat, 12(2):260, June 1991. ISSN

0896-3207. URL https://tug.org/TUGboat/tb12-2/tb32eijkhout-
tricks.pdf.

DeSmet:TB12-2-261

[1319] Don De Smet. TEX macros for producing multiple-choice tests. TUGboat, 12(2):261–269, June 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-2/tb32smet.pdf.

Hefferon:TB12-2-270

[1320] Jim Hefferon. Getting \answers in TEX. TUGboat, 12(2):270–272, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32heff.pdf.

Eijkhout:TB12-2-272

[1321] Victor Eijkhout. Oral TEX. TUGboat, 12(2):272–276, June 1991. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb12-2/tb32eijkhout-
oral.pdf.

Maus:TB12-2-277

[1322] Sonja Maus. An expansion power lemma. TUGboat, 12(2):277, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32maus.pdf.

Russell:TB12-2-278

[1323] George Russell. Generating \n asterisks. TUGboat, 12(2):278–279, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32russell.pdf.

Vulis:TB12-2-279

[1324] Dimitri Vulis. TEX and envelopes. TUGboat, 12(2):279–284, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32vulis.pdf.

Damrau:TB12-2-284

[1325] Jackie Damrau. The L^{A}TEX column. TUGboat, 12(2):284–285, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32damrau.pdf.

Poppelier:TB12-2-285

[1326] Nico Poppelier. A comment on The L^{A}TEX column. TUGboat, 12(2):285–286,

June 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-
2/tb32poppelier.pdf.

Swonk:TB12-2-286

[1327] Glenn L. Swonk. L^{A}TEX tree drawer. TUGboat, 12(2):286–289, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32swonk.pdf.

Thimbleby:TB12-2-290

[1328] Harold Thimbleby. “See also” indexing with Makeindex. TUGboat, 12(2): 290, June 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-2/tb32thim.pdf.

Braams:TB12-2-291

[1329] Johannes Braams. Babel, a multilingual style-option system for use with L^{A}TEX’s standard document styles. TUGboat, 12(2):291–301, June 1991. ISSN

0896-3207. URL https://tug.org/TUGboat/tb12-2/tb32braa.
pdf.

Lankford:TB12-2-302

[1330] Jeff Lankford. Public domain SGML tools wanted. TUGboat, 12(2):302, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32letters.pdf.

Vesely:TB12-2-302

[1331] Jiří Veselý. Reporting TEX’s hyphenations. TUGboat, 12(2):302, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32letters.pdf.

Abrahams:TB12-2-302

[1332] Paul Abrahams. Response to Victor Eijkhout. TUGboat, 12(2):302–303, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32letters.pdf.

Eijkhout:TB12-2-303

[1333] Victor Eijkhout. Response to Paul Abrahams. TUGboat, 12(2):303, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32letters.pdf.

Cuoco:TB12-2-303

[1334] Al Cuoco. TEX in schools: Why not? TUGboat, 12(2):303–304, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32letters.pdf.

Anonymous:TB12-2-305

[1335] Anonymous. Cahiers GUTenberg #7 and #8. TUGboat, 12(2):305–306, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32abstracts.pdf.

Anonymous:TB12-2-307

[1336] Anonymous. Calendar. TUGboat, 12(2):307–308, June 1991. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb12-2/tb32calendar.
pdf.

Anonymous:TB12-2-309

[1337] Anonymous. TEX91/Congres GUTenberg’91, Paris, 23–26 September 1991. TUGboat, 12(2):309–310, June 1991. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb12-2/tb32announce.pdf.

Anonymous:TB12-2-311

[1338] Anonymous. Desktop Publishing in astronomy and space sciences, Strasbourg, 1–3 October 1991. TUGboat, 12(2):311, June 1991. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb12-2/tb32announce.pdf.

Anonymous:TB12-2-311-2

[1339] Anonymous. Call for papers: EP92: International conference on electronic publishing, document manipulation, and typography, Lausanne, Switzerland, 7–10 April 1992. TUGboat, 12(2):311–312, June 1991. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb12-2/tb32announce.pdf.

Knuth:TB12-2-313

[1340] Donald Knuth. Fixed-point glue setting: Errata. TUGboat, 12(2):313, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32knut.pdf.

Beeton:TB12-2-313

[1341] Barbara Beeton. Production notes. TUGboat, 12(2):313–314, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32notes.pdf.

Anonymous:TB12-2-315

[1342] Anonymous. Coming next issue. TUGboat, 12(2):315, June 1991. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb12-2/tb32treas.pdf.

Anonymous:TB12-2-315-2

[1343] Anonymous. TUG financial statements. TUGboat, 12(2):315–318, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32treas.pdf.

Anonymous:TB12-2-319

[1344] Anonymous. Institutional members. TUGboat, 12(2):319–320, June

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
tb32instmem.pdf.

Anonymous:TB12-2-321

[1345] Anonymous. TUG Bylaws. TUGboat, 12(2):321–328, June 1991. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb12-2/tb32bylaws.
pdf.

Anonymous:TB12-2-331

[1346] Anonymous. TUG membership application. TUGboat, 12(2):331, June 1991.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB12-2-348

[1347] Anonymous. Index of advertisers. TUGboat, 12(2):348, June 1991. ISSN

0896-3207. URL https://tug.org/TUGboat/tb12-2/tb32ads.
pdf.

Anonymous:1991:FCb

[1348] Anonymous. Front cover. TUGboat, 12(2):c1, June 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
cover1.pdf.

Anonymous:1991:IFCb

[1349] Anonymous. Inside front cover. TUGboat, 12(2):c2, June 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-2/
cover2.pdf.

Anonymous:1991:TPc

[1350] Anonymous. Title page. TUGboat, 12(3–4):349, December 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33atitlepage.pdf.

Anonymous:1991:TPd

[1351] Anonymous. Title page. TUGboat, 12(3–4):469, December 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33btitlepage.pdf.

Anonymous:1991:FCc

[1352] Anonymous. Front cover. TUGboat, 12(3–4):c1, December 1991. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33bcover1.pdf.

Anonymous:1991:IFCc

[1353] Anonymous. Inside front cover. TUGboat, 12(3–4):c2, December 1991. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-
3-4/tb33bcover2.pdf.

Beebe:TB12-3-351

[1354] Nelson H. F. Beebe. President’s introduction. TUGboat, 12(3/4):351–352, November 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-3-4/tb33beebe.pdf.

Poppelier:TB12-3-353

[1355] Nico A. F. M. Poppelier. Two sides of the fence. TUGboat, 12(3/4):353–358, November 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-3-4/tb33poppelier.pdf.

Petrycki:TB12-3-359

[1356] Laurie J. Petrycki. Comparing TEX and traditional typesetting for the composition of a textbook. TUGboat, 12(3/4):359–366, November 1991.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-4/
tb33petrycki.pdf.

Bartlett:TB12-3-367

[1357] Frederick H. Bartlett. Contra-L^{A}TEX, or what really works in the publishing world. TUGboat, 12(3/4):367–371, November 1991. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb12-3-4/tb33bartlett.pdf.

Ratner:TB12-3-372

[1358] Howard Ratner and Kenneth Dreyhaupt. TEX in a book production department. TUGboat, 12(3/4):372–376, November 1991. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb12-3-4/tb33ratner.pdf.

Horn:TB12-3-377

[1359] Berthold K. P. Horn. dvi and eps: The ideal author-to-publisher interface? TUGboat, 12(3/4):377–381, November 1991. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb12-3-4/tb33horn.pdf.

Weiss:TB12-3-382

[1360] Neil A. Weiss. Producing a book using TEX: How the process works. TUGboat, 12(3/4):382–386, November 1991. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb12-3-4/tb33weiss.pdf.

Rhoads:TB12-3-387

[1361] Samuel E. Rhoads. Authors new to TEX publish a textbook with a publisher new to TEX. TUGboat, 12(3/4):387–392, November 1991. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb12-3-4/tb33rhoads.pdf.

Brosnan:TB12-3-393

[1362] Colleen Brosnan. The “Five Cs”: A guide to successful publication using TEX. TUGboat, 12(3/4):393–396, November 1991. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb12-3-4/tb33brosnan.pdf.

Hoover:TB12-3-397

[1363] Anita Z. Hoover. L^{A}TEX/TEX user: A typist, or typesetter? TUGboat, 12(3/4): 397–400, November 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-3-4/tb33hoover.pdf.

Lavagnino:TB12-3-401

[1364] John Lavagnino. Simultaneous electronic and paper publication. TUGboat, 12(3/4):401–405, November 1991. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb12-3-4/tb33lavagnino.pdf.

McGaffey:TB12-3-406

[1365] Robert W. McGaffey. SGML versus/and TEX. TUGboat, 12(3/4):406–408, November 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-3-4/tb33mcgaffey.pdf.

Dobrowolski:TB12-3-409

[1366] Andrew E. Dobrowolski. Typesetting SGML documents using TEX. TUGboat, 12(3/4):409–414, November 1991. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb12-3-4/tb33dobrowolski.pdf.

Sperberg-McQueen:TB12-3-415

[1367] C. M. Sperberg-McQueen. Specifying document structure: Differences in L^{A}TEX and TEI markup. TUGboat, 12(3/4):415–421, November 1991.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-4/
tb33sperberg.pdf.

Kakiuchi:TB12-3-422

[1368] Takashi Kakiuchi, Yuki Kusumi, Yoshiyuki Miyabe, and Kazu Tsuga. A structured document preparation system — Autolayouter version 2.0 — an enhancement for handling multiple document types. TUGboat, 12(3/4):422–429, November 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-3-4/tb33kakiuchi.pdf.

Williams:TB12-3-430

[1369] Linda Williams. Refining a process. TUGboat, 12(3/4):430–433, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33williams.pdf.

Semenzato:TB12-3-434

[1370] Luigi Semenzato and Edward Wang. A text processing language should be first a programming language. TUGboat, 12(3/4):434–441, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33semenzato.pdf.

Vulis:TB12-3-442

[1371] Michael Vulis. Should TEX be extended? TUGboat, 12(3/4):442–447, Novem-

ber 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-
3-4/tb33vulis.pdf.

Anonymous:TB12-3-448

[1372] Anonymous. TUG 1992 annual meeting, Portland, Oregon. TUGboat, 12 (3/4):448, November 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-3-4/tb33tug92.pdf.

Anonymous:TB12-3-449

[1373] Anonymous. Institutional members. TUGboat, 12(3/4):449–450, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33instmem.pdf.

Anonymous:TB12-3-451

[1374] Anonymous. TUG membership applications. TUGboat, 12(3/4):451–452, November 1991. ISSN 0896-3207. URL https://tug.org/join.

html.

Anonymous:TB12-3-453

[1375] Anonymous. TEX consultants. TUGboat, 12(3/4):453, November 1991. ISSN 0896-3207.

Anonymous:TB12-3-468

[1376] Anonymous. Index of advertisers. TUGboat, 12(3/4):468, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33ads.pdf.

Horstmann:TB12-3-471

[1377] Cay S. Horstmann. Automatic conversion from a scientific word processor to TEX. TUGboat, 12(3/4):471–478, November 1991. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb12-3-4/tb33horstman.pdf.

Arnon:TB12-3-479

[1378] Dennis S. Arnon and Sandra A. Mamrak. On the logical structure of mathematical notation. TUGboat, 12(3/4):479–484, November 1991. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb12-3-4/tb33arnon.pdf.

Laan:TB12-3-485

[1379] Kees van der Laan. Math into BLUes. TUGboat, 12(3/4):485–501, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33laan.pdf.

Downes:TB12-3-502

[1380] Michael J. Downes. Dialog with TEX. TUGboat, 12(3/4):502–509, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33downes.pdf.

Damrau:TB12-3-510

[1381] Jackie Damrau and Michael Wester. Form letters with 3-across labels capability. TUGboat, 12(3/4):510–516, November 1991. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb12-3-4/tb33damrau.pdf.

Roth:TB12-3-517

[1382] Mark A. Roth. Typesetting forms with L^{A}TEX. TUGboat, 12(3/4):517–522, November 1991. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb12-3-4/tb33roth.pdf.

Flynn:TB12-3-523

[1383] Peter Flynn. Developing a pop-up help facility for TEX on PCs. TUGboat, 12(3/4):523–527, November 1991. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb12-3-4/tb33flynn.pdf.

Clark:TB12-3-528

[1384] Malcolm Clark, Brian Hamilton Kelly, and Niel Kempson. 7 bits good, 8 bits bad or “The eight-bit blight”. TUGboat, 12(3/4):528–533, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33clark.pdf.

Sowa:TB12-3-534

[1385] Friedhelm Sowa. Bitmaps and halftones with BM2FONT. TUGboat, 12(3/4): 534–538, November 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-3-4/tb33sowa.pdf.

Haralambous:TB12-3-539

[1386] Yannis Haralambous. TEX and those other languages. TUGboat, 12(3/4): 539–548, November 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-3-4/tb33hara.pdf.

Hosek:TB12-3-549

[1387] Don Hosek. Siamese TEX: Joining dvi files at the hip and other novel applications of vf files. TUGboat, 12(3/4):549–553, November 1991. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb12-3-4/tb33hosek.
pdf.

Hoenig:TB12-3-554

[1388] Alan Hoenig. When TEX and talk: Typesetting on curved paths and other special effects. TUGboat, 12(3/4):554–557, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33hoenig.pdf.

Anonymous:TB12-3-558

[1389] Anonymous. Participants at TUG’91. TUGboat, 12(3/4):558–561, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33names.pdf.

Anonymous:TB12-3-562

[1390] Anonymous. Calendar. TUGboat, 12(3/4):562–563, November 1991.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-4/
tb33calendar.pdf.

Anonymous:TB12-3-563

[1391] Anonymous. The Donald E. Knuth Scholarship for 1992. TUGboat, 12(3/4): 563–564, November 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-3-4/tb33announce.pdf.

Anonymous:TB12-3-564

[1392] Anonymous. GUTenberg’92, Les Diablarets, Switzerland. TUGboat, 12(3/4): 564–565, November 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-3-4/tb33announce.pdf.

Anonymous:TB12-3-570

[1393] Anonymous. TUG 1992 annual meeting, Portland, Oregon. TUGboat, 12 (3/4):570, November 1991. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb12-3-4/tb33announce.pdf.

Anonymous:TB12-3-588

[1394] Anonymous. Index of advertisers. TUGboat, 12(3/4):588, November

1991. ISSN 0896-3207. URL https://tug.org/TUGboat/tb12-3-
4/tb33bads.pdf.

Anonymous:1992:TPa

[1395] Anonymous. Title page. TUGboat, 13(1):1, April 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34titlepage.pdf.

Anonymous:1992:EIa

[1396] Anonymous. Editorial information. TUGboat, 13(1):2, April 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34edit.pdf.

Anonymous:TB13-1-3

[1397] Anonymous. Addresses. TUGboat, 13(1):3–4, April 1992. ISSN 0896-3207.

Clark:TB13-1-5

[1398] Malcolm Clark. Prez says. TUGboat, 13(1):5, April 1992. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb13-1/tb34clark.pdf.

Beebe:TB13-1-6

[1399] Nelson H. F. Beebe. President’s introduction. TUGboat, 13(1):6–10, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34beebe.pdf.

Beeton:TB13-1-10

[1400] Barbara Beeton. Editorial comments. TUGboat, 13(1):10–11, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34beet.pdf.

Anonymous:TB13-1-11

[1401] Anonymous. Samuel B. Whidden, 1930–1991. TUGboat, 13(1):11–12, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34beet.pdf.

Asher:TB13-1-13

[1402] Graham Asher. Inside Type & Set. TUGboat, 13(1):13–22, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34asher.pdf.

Beccari:TB13-1-23

[1403] Claudio Beccari. Computer Aided Hyphenation for Italian and Modern Latin. TUGboat, 13(1):23–33, April 1992. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb13-1/tb34becc.pdf.

Rahtz:TB13-1-34

[1404] Sebastian Rahtz. Invisibility using virtual fonts. TUGboat, 13(1):34–36, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34rahtz.pdf.

Thain:TB13-1-36

[1405] Toby Thain. Packing s into PostScript. TUGboat, 13(1):36– 38, April 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-1/tb34thain.pdf.

Mylonas:TB13-1-39

[1406] C. Mylonas and R. Whitney. Modern Greek with adjunct fonts. TUGboat, 13(1):39–50, April 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-1/tb34mylonas.pdf.

Mittelbach:TB13-1-51

[1407] Frank Mittelbach. Comments on “Filenames for Fonts” (TUGboat 11#4). TUGboat, 13(1):51–53, April 1992. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb13-1/tb34mittfont.pdf.

TDDSC:TB13-1-54

[1408] TUG DVI Driver Standards Committee. DVI driver standard, level 0. TUGboat, 13(1):54–57, April 1992. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb13-1/tb34dvistd.pdf.

Eijkhout:TB13-1-57

[1409] Victor Eijkhout. New books on TEX. TUGboat, 13(1):57–58, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34resources.pdf.

Poppelier:TB13-1-58

[1410] Nico Poppelier. New books on L^{A}TEX. TUGboat, 13(1):58–59, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34resources.pdf.

Hoenig:TB13-1-60

[1411] Alan Hoenig. Just plain Q&A: Of partitioned matrices and doublespacing. TUGboat, 13(1):60–62, April 1992. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb13-1/tb34hoenig.pdf.

Siebenmann:TB13-1-62

[1412] L. Siebenmann. Elementary text processing and parsing in TEX — the appreciation of tokens — . TUGboat, 13(1):62–73, April 1992. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb13-1/tb34siebenmann.pdf.

Eijkhout:TB13-1-74

[1413] Victor Eijkhout. The bag of tricks. TUGboat, 13(1):74, April 1992. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb13-1/tb34eijkhout-
tricks.pdf.

Eijkhout:TB13-1-75

[1414] Victor Eijkhout. Erratum: Oral TEX, TUGboat 1, no. 2(2), pp. 272–276. TUGboat, 13(1):75, April 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-1/tb34fine.pdf.

Fine:TB13-1-75

[1415] Jonathan Fine. Some basic control macros for TEX. TUGboat, 13(1):75– 83, April 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-1/tb34fine.pdf.

Eijkhout:TB13-1-84

[1416] Victor Eijkhout and Ron Sommeling. Self-replicating macros. TUGboat, 13(1):84, April 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-1/tb34eijkhout-selfrepl.pdf.

Tutelaers:TB13-1-85

[1417] Piet Tutelaers. A font and a style for typesetting chess using L^{A}TEX or TEX. TUGboat, 13(1):85–90, April 1992. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb13-1/tb34tutelaers.pdf.

Laan:TB13-1-91

[1418] Kees van der Laan. Tower of Hanoi, revisited. TUGboat, 13(1):91–94, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34laan.pdf.

Damrau:TB13-1-94

[1419] Jackie Damrau. The L^{A}TEX column. TUGboat, 13(1):94–95, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34damrau.pdf.

Thimbleby:TB13-1-95

[1420] Harold Thimbleby. Erratum: “See also” indexing with Makeindex, TUGboat 12, no. (2), p. 290. TUGboat, 13(1):95, April 1992. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb13-1/tb34thim.pdf.

Mittelbach:TB13-1-96

[1421] Frank Mittelbach and Chris Rowley. L^{A}TEX 2.09 ↪ L^{A}TEX3. TUGboat, 13 (1):96–101, April 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-1/tb34mittl3.pdf.

Anonymous:TB13-1-101

[1422] Anonymous. Cahiers GUTenberg #9 and #10–11. TUGboat, 13(1):101– 105, April 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-1/tb34abstracts.pdf.

Anonymous:TB13-1-106

[1423] Anonymous. Calendar. TUGboat, 13(1):106–107, April 1992. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb13-1/tb34calendar.
pdf.

Anonymous:TB13-1-107

[1424] Anonymous. EuroTEX 92, Prague, 14–18 September 1992. TUGboat, 13 (1):107–108, April 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-1/tb34calendar.pdf.

Beeton:TB13-1-108

[1425] Barbara Beeton. Production notes. TUGboat, 13(1):108–109, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34notes.pdf.

Anonymous:TB13-1-109

[1426] Anonymous. Coming next issue. TUGboat, 13(1):109, April 1992. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb13-1/tb34notes.pdf.

Anonymous:TB13-1-110

[1427] Anonymous. Institutional members. TUGboat, 13(1):110–111, April

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34instmem.pdf.

Anonymous:TB13-1-112

[1428] Anonymous. TUG membership application. TUGboat, 13(1):112, April 1992.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB13-1-126

[1429] Anonymous. Index of advertisers. TUGboat, 13(1):126, April 1992. ISSN

0896-3207. URL https://tug.org/TUGboat/tb13-1/tb34ads.
pdf.

Anonymous:TB13-1-127

[1430] Anonymous. TEX consulting and production services. TUGboat, 13(1): 127, April 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-1/tb34ads.pdf.

Anonymous:1992:FC

[1431] Anonymous. Front cover. TUGboat, 13(1):c1, April 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34cover1.pdf.

Anonymous:1992:IFCa

[1432] Anonymous. Inside front cover. TUGboat, 13(1):c2, April 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-1/
tb34cover2.pdf.

Anonymous:1992:TPb

[1433] Anonymous. Title page. TUGboat, 13(2):129, July 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35titlepage.pdf.

Anonymous:1992:EIb

[1434] Anonymous. Editorial information. TUGboat, 13(2):130, July 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35edit.pdf.

Anonymous:TB13-2-131

[1435] Anonymous. Addresses. TUGboat, 13(2):131–132, July 1992. ISSN 0896-3207.

Clark:TB13-2-133

[1436] Malcolm Clark. Changing TEX? TUGboat, 13(2):133–134, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35clark.pdf.

Beeton:TB13-2-134

[1437] Barbara Beeton. Editorial comments. TUGboat, 13(2):134–137, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35beet.pdf.

Anonymous:TB13-2-137

[1438] Anonymous. TUG seeks Executive Director. TUGboat, 13(2):137, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35beet.pdf.

Taylor:TB13-2-138

[1439] Philip Taylor. TEX: The next generation. TUGboat, 13(2):138, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35taylor.pdf.

Damerell:TB13-2-139

[1440] R. M. Damerell. Knuth’s profiler adapted to the VMS operating system. TUGboat, 13(2):139–145, July 1992. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb13-2/tb35damerell.pdf.

Salomon:TB13-2-146

[1441] David Salomon. Arrows for Technical Drawings. TUGboat, 13(2):146–149,

July 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-
2/tb35salomon.pdf.

Levin:TB13-2-150

[1442] Daniel Levin. A solution to the color separation problem. TUGboat, 13 (2):150–155, July 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-2/tb35levin.pdf.

Rahtz:TB13-2-156

[1443] Sebastian Rahtz and Leonor Barroca. A style option for rotated objects in TEX. TUGboat, 13(2):156–180, July 1992. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb13-2/tb35rahtz.pdf.

Luvisetto:TB13-2-181

[1444] Marisa Luvisetto and Massimo Calvani. Book review: An Italian guide to L^{A}TEX (by Claudio Beccari). TUGboat, 13(2):181–182, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35resources.pdf.

Poppelier:TB13-2-182

[1445] Nico Poppelier. Book reviews: Jane Hahn, L^{A}TEX for Everyone; Eric van Herwijnen, Practical SGML. TUGboat, 13(2):182–185, July 1992. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb13-2/tb35resources.pdf.

Taylor:TB13-2-185

[1446] Philip Taylor. Book review: Victor Eijkhout, TEX by Topic. TUGboat, 13 (2):185–188, July 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-2/tb35resources.pdf.

Jones:TB13-2-188

[1447] David M. Jones. A TEX macro index. TUGboat, 13(2):188–189, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35resources.pdf.

Eijkhout:TB13-2-189

[1448] Victor Eijkhout. Names of control sequences. TUGboat, 13(2):189–190, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35eijkhout-names.pdf.

Mittelbach:TB13-2-190

[1449] Frank Mittelbach. Where does this character come from? TUGboat, 13(2): 190, July 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-2/tb35eijkhout-names.pdf.

Eijkhout:TB13-2-191

[1450] Victor Eijkhout. The bag of tricks. TUGboat, 13(2):191, July 1992. ISSN

0896-3207. URL https://tug.org/TUGboat/tb13-2/tb35eijkhout-
tricks.pdf.

Huszar:TB13-2-192

[1451] Péter Huszár. Over the multi-column. TUGboat, 13(2):192–200, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35huszar.pdf.

Goossens:TB13-2-201

[1452] Michel Goossens and Eric van Herwijnen. The elementary Particle Entity Notation (PEN) scheme. TUGboat, 13(2):201–207, July 1992. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb13-2/tb35goossens.pdf.

Luvisetto:TB13-2-208

[1453] Maria Luisa Luvisetto and Enzo Ugolini. From TEX to L^{A}TEX. TUGboat, 13(2):208–214, July 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-2/tb35luvisetto.pdf.

Cameron:TB13-2-215

[1454] Peter J. Cameron. Geometric diagrams in L^{A}TEX. TUGboat, 13(2):215–216,

July 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-
2/tb35cameron.pdf.

Partl:TB13-2-217

[1455] Hubert Partl. How to change the layout with L^{A}TEX 2.09. TUGboat, 13(2):217– 220, July 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-2/tb35partl.pdf.

Wonneberger:TB13-2-221

[1456] Reinhard Wonneberger and Frank Mittelbach. SGML — Questions and answers. TUGboat, 13(2):221–223, July 1992. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb13-2/tb35wonneberger-sgmlqa.pdf.

Barr:TB13-2-223

[1457] Michael Barr. TEX wish list. TUGboat, 13(2):223–226, July 1992. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb13-2/tb35barr.pdf.

Wonneberger:TB13-2-226

[1458] Reinhard Wonneberger. Approaching SGML from TEX. TUGboat, 13(2):226– 227, July 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-2/tb35wonneberger-sgmltex.pdf.

Anonymous:TB13-2-227

[1459] Anonymous. Cahiers GUTenberg #12. TUGboat, 13(2):227–228, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35abstracts.pdf.

Anonymous:TB13-2-228

[1460] Anonymous. Baskerville, Volume 2, Number 1, March 1992. TUGboat, 13(2):228–229, July 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-2/tb35abstracts.pdf.

Beeton:TB13-2-229

[1461] Barbara Beeton. Production notes. TUGboat, 13(2):229–230, July

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35notes.pdf.

Anonymous:TB13-2-230

[1462] Anonymous. Coming next issue. TUGboat, 13(2):230, July 1992. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb13-2/tb35notes.pdf.

Anonymous:TB13-2-231

[1463] Anonymous. Calendar. TUGboat, 13(2):231–232, July 1992. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb13-2/tb35calendar.pdf.

Anonymous:TB13-2-232

[1464] Anonymous. EuroTEX 92, Prague, 14–18 September 1992. TUGboat, 13 (2):232–233, July 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-2/tb35calendar.pdf.

Anonymous:TB13-2-235

[1465] Anonymous. TUG membership application. TUGboat, 13(2):235–236, July

1992. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB13-2-247

[1466] Anonymous. Index of advertisers. TUGboat, 13(2):247, July 1992. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb13-2/tb35ads.pdf.

Anonymous:1992:IFCb

[1467] Anonymous. Inside front cover. TUGboat, 13(2):c2, July 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-2/
tb35cover2.pdf.

Anonymous:1992:TPc

[1468] Anonymous. Title page. TUGboat, 13(3):249, October 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-3/
tb36titlepage.pdf.

Anonymous:1992:EIc

[1469] Anonymous. Editorial information. TUGboat, 13(3):250, October 1992. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-
3/tb36edit.pdf.

Clark:TB13-3-251

[1470] Malcolm Clark. President’s introduction. TUGboat, 13(3):251–252, October

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-3/
intro.pdf.

Clark:TB13-3-253

[1471] Malcolm Clark. Portable graphics in TEX. TUGboat, 13(3):253–260, October

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-3/
clark-keynote.pdf.

Childs:TB13-3-261

[1472] Bart Childs. Literate programming, a practitioner’s view. TUGboat, 13(3): 261–268, October 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-3/childs.pdf.

Hampson:TB13-3-269

[1473] Steve Hampson and Barry Smith. A high performance TEX for the Motorola 68000 processor family. TUGboat, 13(3):269–271, October 1992. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb13-3/smith.pdf.

Baldwin:TB13-3-272

[1474] Harry L. Baldwin, Jr. Using a high-level language as an aid in writing TEX documents. TUGboat, 13(3):272–280, October 1992. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb13-3/baldwin.pdf.

Bennett:TB13-3-281

[1475] Larry F. Bennett. T-EDIT, a collection of editing macros for TEX. TUGboat, 13 (3):281–290, October 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-3/bennett.pdf.

Gaffey:TB13-3-291

[1476] Robert McGaffey. Automatic tables using SGML, C, and TEX. TUGboat, 13 (3):291–294, October 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-3/mcgaffey.pdf.

Starks:TB13-3-295

[1477] Anthony J. Starks. Dotex — integrating TEX into the X-window system. TUGboat, 13(3):295–303, October 1992. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb13-3/starks.pdf.

Thorup:TB13-3-304

[1478] Kresten Krab Thorup. GNU emacs as a front end to L^{A}TEX. TUGboat, 13 (3):304–308, October 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-3/thorup.pdf.

Laan:TB13-3-309

[1479] Walter van der Laan and Johannes Braams. Writing reports with more than a hundred people. TUGboat, 13(3):309–314, October 1992. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb13-3/braams.pdf.

Damrau:TB13-3-315

[1480] Jackie Damrau. Discovering graphics in L^{A}TEX documents. TUGboat, 13 (3):315–321, October 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-3/damrau.pdf.

Harris:TB13-3-322

[1481] Robert L. Harris. Preparing halftones for use in TEX. TUGboat, 13(3):322–326, October 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-3/harris.pdf.

Salomon:TB13-3-327

[1482] David Salomon. Creating shaded rectangles with PostScript. TUGboat, 13 (3):327–329, October 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-3/salomon.pdf.

Weiss:TB13-3-330

[1483] Neil A. Weiss. Creation and incorporation of PostScript graphics with TEX-formatted labels into TEX documents. TUGboat, 13(3):330–334, October

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-3/
weiss.pdf.

Knuutila:TB13-3-335

[1484] Timo Knuutila. How to combine multiple languages, PostScript, and L^{A}TEX. TUGboat, 13(3):335–340, October 1992. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb13-3/knuutila.pdf.

Eijkhout:TB13-3-341

[1485] Victor Eijkhout. Just give me a lollipop (it makes my heart go giddy-up). TUGboat, 13(3):341–346, October 1992. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb13-3/eijkhout.pdf.

Hafner:TB13-3-347

[1486] James L. Hafner. FoilTEX, a L^{A}TEX-like system for typesetting foils. TUGboat, 13(3):347–356, October 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-3/hafner.pdf.

Abbott:TB13-3-357

[1487] Peter Abbott. Typesetting a magazine the easy way. TUGboat, 13(3):357–361, October 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-3/abbott.pdf.

Burbank:TB13-3-362

[1488] Mimi Burbank and Donna Burnette. Using TEX for a publications database. TUGboat, 13(3):362–371, October 1992. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb13-3/burbank.pdf.

Raman:TB13-3-372

[1489] T. V. Raman. An audio view of (La)TEX documents. TUGboat, 13(3):372–379, October 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-3/raman.pdf.

Arnon:TB13-3-380

[1490] Dennis S. Arnon, Isabelle Attali, and Paul Franchi-Zannettacci. Modelbased conversions of L^{A}TEX documents. TUGboat, 13(3):380–389, October

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-3/
arnon.pdf.

Rowley:TB13-3-390

[1491] Chris Rowley. L^{A}TEX3 update. TUGboat, 13(3):390–391, October

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-3/
tb36rowley.pdf.

Anonymous:TB13-3-391

[1492] Anonymous. Workshops. TUGboat, 13(3):391–392, October 1992. ISSN

0896-3207. URL https://tug.org/TUGboat/tb13-3/workshop-
summary.pdf.

Anonymous:TB13-3-393

[1493] Anonymous. Participants at the 1992 TUG Meeting. TUGboat, 13(3):393–394, October 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-3/tb36names.pdf.

Anonymous:TB13-3-395

[1494] Anonymous. The Donald E. Knuth Scholarship for 1992 and 1993. TUGboat, 13(3):395–396, October 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-3/tb36announce.pdf.

Anonymous:TB13-3-396

[1495] Anonymous. Calendar. TUGboat, 13(3):396–397, October 1992. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb13-3/tb36announce.
pdf.

Anonymous:TB13-3-398

[1496] Anonymous. TUG 1993 annual meeting, Aston University, UK. TUGboat, 13(3):398, October 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-3/tb36announce.pdf.

Anonymous:TB13-3-399

[1497] Anonymous. TUG 1993 course schedule. TUGboat, 13(3):399, October

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-3/
tb36announce.pdf.

Anonymous:TB13-3-400

[1498] Anonymous. Institutional members. TUGboat, 13(3):400–401, October

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-3/
tb36instmem.pdf.

Anonymous:TB13-3-402

[1499] Anonymous. Consultants. TUGboat, 13(3):402, October 1992. ISSN 0896-

3207. URL https://tug.org/consultants.html.

Anonymous:TB13-3-411

[1500] Anonymous. Index of advertisers. TUGboat, 13(3):411, October 1992. ISSN

0896-3207. URL https://tug.org/TUGboat/tb13-3/tb36ads.
pdf.

Anonymous:1992:IFCc

[1501] Anonymous. Inside front cover. TUGboat, 13(3):c2, October 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-3/
tb36cover2.pdf.

Anonymous:1992:TPd

[1502] Anonymous. Title page. TUGboat, 13(4):413, December 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37titlepage.pdf.

Anonymous:1992:EId

[1503] Anonymous. Editorial information. TUGboat, 13(4):414, December 1992. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-4/tb37edit.pdf.

Anonymous:TB13-4-415

[1504] Anonymous. Addresses. TUGboat, 13(4):415–416, December 1992. ISSN 0896-3207.

Clark:TB13-4-417

[1505] Malcolm Clark. The back benches beckon. TUGboat, 13(4):417–418, Decem-

ber 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-
4/tb37clark.pdf.

Beeton:TB13-4-418

[1506] Barbara Beeton. Editorial comments. TUGboat, 13(4):418–419, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37beet.pdf.

Anonymous:TB13-4-419

[1507] Anonymous. An interview with Donald Knuth, November 1991. TUGboat, 13 (4):419–425, December 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-4/tb37knut.pdf.

Palais:TB13-4-425

[1508] Richard Palais. Moving a fixed point. TUGboat, 13(4):425–432, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37palais.pdf.

Taylor:TB13-4-433

[1509] Philip Taylor. The future of TEX. TUGboat, 13(4):433–442, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37taylor.pdf.

Kelly:TB13-4-443

[1510] Nickolas J. Kelly and Christian H. Bischof. XBibTEX and friends. TUGboat, 13(4):443–446, December 1992. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb13-4/tb37kelly.pdf.

Chapman:TB13-4-447

[1511] Nigel Chapman. Searching in a DVI file. TUGboat, 13(4):447–451, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37chapman.pdf.

Anonymous:TB13-4-452

[1512] Anonymous. Hyphenation exception log. TUGboat, 13(4):452–457, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37hyf.pdf.

Childs:TB13-4-457

[1513] Bart Childs. Errata: Literate Programming, A Practitioner’s View, TUGboat 1, no. 3(3), pp. 261–268. TUGboat, 13(4):457, December 1992. ISSN

0896-3207. URL https://tug.org/TUGboat/tb13-4/tb37hara-
hyfgl.pdf.

Haralambous:TB13-4-457

[1514] Yannis Haralambous. Hyphenation patterns for ancient Greek and Latin. TUGboat, 13(4):457–469, December 1992. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb13-4/tb37hara-hyfgl.pdf.

Zubrinic:TB13-4-470

[1515] Darko Žubrinić. The exotic Croatian Glagolitic alphabet. TUGboat, 13(4): 470–471, December 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-4/tb37zubrinic.pdf.

Sauter:TB13-4-472

[1516] John Sauter. Postnet codes using . TUGboat, 13(4):472–476, December 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-4/tb37sauter.pdf.

Haralambous:TB13-4-476

[1517] Yannis Haralambous. A typewriter font for the Macintosh 8-bit font table. TUGboat, 13(4):476–477, December 1992. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb13-4/tb37hara-mactt.pdf.

Rahtz:TB13-4-477

[1518] Sebastian Rahtz and Leonor Barroca. Addendum: A style option for rotated objects in TEX (TUGboat 1, no. 3(2), pp. 156–180). TUGboat, 13(4):477, December 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-4/tb37hara-mactt2.pdf.

Seyfarth:TB13-4-478

[1519] Ray Seyfarth. Diag: a drawing preprocessor for L^{A}TEX. TUGboat, 13(4): 478–485, December 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-4/tb37seyfarth.pdf.

Eijkhout:TB13-4-486

[1520] Victor Eijkhout. Wynter Snow, TEX for the Beginner. TUGboat, 13(4): 486–487, December 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-4/tb37reviews.pdf.

Greenwade:TB13-4-487

[1521] George Greenwade. Arvind Borde, TEX by Example. TUGboat, 13(4):487–489, December 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-4/tb37reviews.pdf.

Cameron:TB13-4-489

[1522] A. G. W. Cameron. André Heck, ed., Desktop Publishing in Astronomy & Space Sciences. TUGboat, 13(4):489–490, December 1992. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb13-4/tb37reviews.pdf.

Neuwirth:TB13-4-490

[1523] Erich Neuwirth. TEX implementations for IBM PCs: comparative timings. TUGboat, 13(4):490–492, December 1992. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb13-4/tb37neuwirth.pdf.

Mittelbach:TB13-4-493

[1524] Frank Mittelbach. Where does this character come from? Solution to the puzzle, TUGboat 1, no. 3(2), p.190. TUGboat, 13(4):493, December 1992. ISSN

0896-3207. URL https://tug.org/TUGboat/tb13-4/tb37mitt-
puzzle.pdf.

Eijkhout:TB13-4-494

[1525] Victor Eijkhout. The bag of tricks. TUGboat, 13(4):494–495, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37eijkhout-tricks.pdf.

Fine:TB13-4-495

[1526] Jonathan Fine. Too many errors. TUGboat, 13(4):495–496, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37fine.pdf.

Eijkhout:TB13-4-496

[1527] Victor Eijkhout. One error less. TUGboat, 13(4):496–497, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37eijkhout-error.pdf.

Anagnostopoulos:TB13-4-497

[1528] Paul Anagnostopoulos. ZzTEX: A macro package for books. TUGboat, 13(4): 497–505, December 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-4/tb37anag.pdf.

Fine:TB13-4-505

[1529] Jonathan Fine. The \noname macros — A technical report. TUGboat, 13(4): 505–509, December 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-4/tb37finn.pdf.

Mittelbach:TB13-4-510

[1530] Frank Mittelbach, Chris Rowley, and Michael Downes. Volunteer work for the L^{A}TEX3 project. TUGboat, 13(4):510–515, December 1992. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb13-4/tb37mitt-l3.pdf.

Piff:TB13-4-516

[1531] Mike Piff. Correction sheets in L^{A}TEX. TUGboat, 13(4):516–518, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37piff-corr.pdf.

Piff:TB13-4-518

[1532] Mike Piff. Text merges in TEX and L^{A}TEX. TUGboat, 13(4):518–523, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37piff-merge.pdf.

Rahtz:TB13-4-524

[1533] Sebastian Rahtz. A style file for printing sheets of labels. TUGboat, 13(4): 524–528, December 1992. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb13-4/tb37rahtz.pdf.

Anonymous:TB13-4-528

[1534] Anonymous. Cahiers GUTenberg #13. TUGboat, 13(4):528–529, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37abstracts.pdf.

Anonymous:TB13-4-530

[1535] Anonymous. Calendar. TUGboat, 13(4):530–532, December 1992. ISSN

0896-3207. URL https://tug.org/TUGboat/tb13-4/tb37calendar.
pdf.

Anonymous:TB13-4-532

[1536] Anonymous. Call for Papers: Special Issue of Electronic Publishing: Origination, Dissemination and Design on Active Documents. TUGboat, 13(4):532, December 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-4/tb37calendar.pdf.

Beeton:TB13-4-533

[1537] Barbara Beeton. Production notes. TUGboat, 13(4):533, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37notes.pdf.

Anonymous:TB13-4-534

[1538] Anonymous. Coming next issue. TUGboat, 13(4):534, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37instmem.pdf.

Anonymous:TB13-4-534-2

[1539] Anonymous. Institutional members. TUGboat, 13(4):534–536, December

1992. ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37instmem.pdf.

Anonymous:TB13-4-536

[1540] Anonymous. Index of advertisers. TUGboat, 13(4):536, December 1992. ISSN

0896-3207. URL https://tug.org/TUGboat/tb13-4/tb37ads.
pdf.

Anonymous:TB13-4-537

[1541] Anonymous. TUG membership application. TUGboat, 13(4):537–538, December 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-4/tb37ads.pdf.

Anonymous:TB13-4-544

[1542] Anonymous. TEX consulting and production services. TUGboat, 13(4):544, December 1992. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb13-4/tb37ads.pdf.

Anonymous:1992:IFCd

[1543] Anonymous. Inside front cover. TUGboat, 13(4):c2, December 1992. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb13-4/
tb37cover2.pdf.

Anonymous:1993:TPa

[1544] Anonymous. Title page. TUGboat, 14(1):1, March 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-1/
tb38titlepage.pdf.

Anonymous:1993:EIa

[1545] Anonymous. Editorial information. TUGboat, 14(1):2, March 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-1/
tb38edit.pdf.

Anonymous:TB14-1-3

[1546] Anonymous. Addresses. TUGboat, 14(1):3–4, April 1993. ISSN 0896-3207.

Thiele:TB14-1-5

[1547] Christina Thiele. Opening words. TUGboat, 14(1):5–6, April 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-1/tb38thiele.
pdf.

Beeton:TB14-1-7

[1548] Barbara Beeton. Editorial comments. TUGboat, 14(1):7, April 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-1/tb38beet.
pdf.

Wong:TB14-1-8

[1549] Wai Wong. Typesetting Chinese pinyin using virtual fonts. TUGboat, 14(1): 8–11, April 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-1/tb38wong.pdf.

Vesilo:TB14-1-12

[1550] R. A. Vesilo and A. Dunn. A multimedia document system based on TEX and DVI documents. TUGboat, 14(1):12–16, April 1993. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb14-1/tb38vesilo.pdf.

Taylor:TB14-1-17

[1551] Philip Taylor. Arvind Borde, Mathematical TEX by Example. TUGboat, 14(1):17–20, April 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-1/tb38reviews.pdf.

Radel:TB14-1-20

[1552] Jon Radel. Michael Vulis, Modern TEX and Its Applications. TUGboat, 14(1):20–23, April 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-1/tb38reviews.pdf.

Eijkhout:TB14-1-23

[1553] Victor Eijkhout. The bag of tricks. TUGboat, 14(1):23–24, April

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-1/
tb38eijkhout.pdf.

Comenetz:TB14-1-25

[1554] Daniel Comenetz. Anchored figures at either margin. TUGboat, 14(1):25– 34, April 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-1/tb38comenetz.pdf.

Fine:TB14-1-35

[1555] Jonathan Fine. The \CASE and \FIND macros. TUGboat, 14(1):35–39, April

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-1/
tb38fine.pdf.

Saludes:TB14-1-40

[1556] Jordi Saludes. Doing astronomical computations with TEX: Making agendas. TUGboat, 14(1):40–53, April 1993. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb14-1/tb38saludes.pdf.

Laan:TB14-1-54

[1557] Kees van der Laan. FIFO and LIFO sing the BLUes. TUGboat, 14(1):54– 60, April 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-1/tb38laan.pdf.

Braams:TB14-1-60

[1558] Johannes Braams. An update on the babel system. TUGboat, 14(1):60– 62, April 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-1/tb38braa.pdf.

Zbikowski:TB14-1-62

[1559] Rafa� Żbikowski. Hacker’s Guide to AMSFonts and NFSS in the Context of L^{A}TEX. TUGboat, 14(1):62–69, April 1993. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb14-1/tb38zbikowski.pdf.

Heck:TB14-1-70

[1560] André Heck. Response to A. G. W. Cameron. TUGboat, 14(1):70, April

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-1/
tb38heck.pdf.

Anonymous:TB14-1-71

[1561] Anonymous. Die TEXnische Komödie 1992, Heft 1–4. TUGboat, 14(1):71– 76, April 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-1/tb38abstracts.pdf.

Anonymous:TB14-1-77

[1562] Anonymous. Calendar. TUGboat, 14(1):77–79, April 1993. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb14-1/tb38calendar.pdf.

Anonymous:TB14-1-79

[1563] Anonymous. Courses to be held in conjunction with TUG 93 (Aston University, Birmingham, U.K., 26–30 July 1993). TUGboat, 14(1):79, April

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-1/
tb38calendar.pdf.

Beeton:TB14-1-81

[1564] Barbara Beeton. Production notes. TUGboat, 14(1):81, April 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-1/tb38notes.
pdf.

Anonymous:TB14-1-81

[1565] Anonymous. Coming next issue. TUGboat, 14(1):81–82, April 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-1/tb38notes.
pdf.

Anonymous:TB14-1-82

[1566] Anonymous. Index of advertisers. TUGboat, 14(1):82, April 1993. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb14-1/tb38notes.pdf.

Anonymous:TB14-1-83

[1567] Anonymous. Institutional members. TUGboat, 14(1):83–84, April

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-1/
tb38instmem.pdf.

Anonymous:TB14-1-85

[1568] Anonymous. TUG membership application. TUGboat, 14(1):85–86, April

1993. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB14-1-87

[1569] Anonymous. TEX consulting and production services. TUGboat, 14(1): 87, April 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-1/tb38ads.pdf.

Anonymous:1993:IFCa

[1570] Anonymous. Inside front cover. TUGboat, 14(1):c2, March 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-1/
tb38cover2.pdf.

Anonymous:1993:TPb

[1571] Anonymous. Title page. TUGboat, 14(2):93, June 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39titlepage.pdf.

Anonymous:1993:EIb

[1572] Anonymous. Editorial information. TUGboat, 14(2):94, June 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39edit.pdf.

Anonymous:TB14-2-95

[1573] Anonymous. Addresses. TUGboat, 14(2):95–96, July 1993. ISSN 0896-3207.

Thiele:TB14-2-97

[1574] Christina Thiele. Opening words. TUGboat, 14(2):97–98, July 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-2/tb39thiele.
pdf.

Beeton:TB14-2-99

[1575] Barbara Beeton. Editorial comments. TUGboat, 14(2):99–101, July

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39beet.pdf.

Poppelier:TB14-2-102

[1576] Nico Poppelier. The Donald E. Knuth Scholarship: 1993 Scholar and 1994 announcement. TUGboat, 14(2):102–103, July 1993. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb14-2/tb39knuthscholar.pdf.

Anonymous:TB14-2-103

[1577] Anonymous. The A-in-L^{A}TEX Contest: Deadline extended. TUGboat, 14(2): 103, July 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-2/tb39knuthscholar.pdf.

Knappen:TB14-2-104

[1578] Jörg Knappen. Fonts for Africa: the fc-fonts. TUGboat, 14(2):104–106, July

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39knappen.pdf.

Rahtz:TB14-2-107

[1579] Sebastian Rahtz. Implementing the extended TEX layout using PS fonts. TUGboat, 14(2):107–117, July 1993. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb14-2/tb39rahtz-pscork.pdf.

Cohen:TB14-2-118

[1580] Michael Cohen. Zebrackets: a pseudo-dynamic contextually adaptive font. TUGboat, 14(2):118–122, July 1993. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb14-2/tb39cohen.pdf.

Jurriens:TB14-2-123

[1581] Theo Jurriens. From observation to publication. TUGboat, 14(2):123–126,

July 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-
2/tb39jurriens.pdf.

Poppelier:TB14-2-127

[1582] Nico Poppelier. Review of recent L^{A}TEX books. TUGboat, 14(2):127–128, July

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39reviews.pdf.

Collins:TB14-2-128

[1583] John Collins. ET — a TEX-compatible editor for MSDOS computers. TUGboat, 14(2):128–131, July 1993. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb14-2/tb39collins.pdf.

Rahtz:TB14-2-132

[1584] Sebastian Rahtz. Essential NFSS2. TUGboat, 14(2):132–137, July

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39rahtz-nfss.pdf.

Taylor:TB14-2-138

[1585] Philip Taylor. A pragmatic approach to paragraphs. TUGboat, 14(2):138–140,

July 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-
2/tb39taylor-para.pdf.

Durst:TB14-2-141

[1586] Lincoln Durst. Truth in indexing. TUGboat, 14(2):141, July 1993. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb14-2/tb39durst.pdf.

Taylor:TB14-2-141

[1587] Philip Taylor. Letterspacing in TEX. TUGboat, 14(2):141–145, July

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39taylor-letterspacing.pdf.

Anonymous:TB14-2-146

[1588] Anonymous. Cahiers GUTenberg #15. TUGboat, 14(2):146, July

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39abstracts.pdf.

Anonymous:TB14-2-147

[1589] Anonymous. Calendar. TUGboat, 14(2):147–148, July 1993. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb14-2/tb39calendar.pdf.

Anonymous:TB14-2-148

[1590] Anonymous. Index of advertisers. TUGboat, 14(2):148, July 1993. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb14-2/tb39calendar.
pdf.

Anonymous:TB14-2-149

[1591] Anonymous. Coming next issue. TUGboat, 14(2):149, July 1993. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb14-2/tb39notes.pdf.

Beeton:TB14-2-149

[1592] Barbara Beeton. Production notes. TUGboat, 14(2):149, July 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-2/tb39notes.
pdf.

Anonymous:TB14-2-150

[1593] Anonymous. Institutional members. TUGboat, 14(2):150–151, July

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39instmem.pdf.

Anonymous:TB14-2-152

[1594] Anonymous. TEX consulting and production services. TUGboat, 14(2):152,

July 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-
2/tb39ads.pdf.

Anonymous:1993:IFCb

[1595] Anonymous. Inside front cover. TUGboat, 14(2):c2, June 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-2/
tb39cover2.pdf.

Anonymous:1993:TPc

[1596] Anonymous. Title page. TUGboat, 14(3):157, October 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-3/
tb40titlepage.pdf.

Clark:1993:OYM

[1597] Malcolm Clark. Obituary — Yuri Melnichuk. TUGboat, 14(3):158, October 1993. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-3/tb40edit.pdf.

Rahtz:1993:EPN

[1598] Sebastian Rahtz and Mimi Burbank. Editorial and production note. TUGboat, 14(3):158, October 1993. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb14-3/tb40edit.pdf.

Thiele:1993:OW

[1599] Christina Thiele. Opening words. TUGboat, 14(3):159, October 1993. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-
3/tb40thiele-intro.pdf.

Anonymous:1993:CP

[1600] Anonymous. Conference program. TUGboat, 14(3):160–161, October 1993. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-3/tb40program.pdf.

Thiele:TB14-3-162

[1601] Christina A. L. Thiele. The future of TEX and TUG. TUGboat, 14(3):162–166, October 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-3/tb40thiele-future.pdf.

Lammarsch:TB14-3-167

[1602] Joachim Lammarsch. A new typesetting system: is it really necessary? TUGboat, 14(3):167–170, October 1993. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb14-3/tb40lammarsch.pdf.

Rycko:TB14-3-171

[1603] Marek Ryćko and Bogus�aw Jackowski. TEX from \indent to \par. TUGboat, 14(3):171–176, October 1993. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb14-3/tb40rycko.pdf.

Taylor:TB14-3-177

[1604] Philip Taylor. NTS: the future of TEX? TUGboat, 14(3):177–182, October

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-3/
tb40taylor.pdf.

Hunter:TB14-3-183

[1605] Roger Hunter. A future for TEX. TUGboat, 14(3):183–186, October

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-3/
tb40hunter.pdf.

Bennett:TB14-3-187

[1606] Frank G. Bennett, Jr. LexITEX: context-sensitive legal citations for L^{A}TEX. TUGboat, 14(3):187–195, October 1993. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb14-3/tb40bennett.pdf.

Taupin:TB14-3-196

[1607] Daniel Taupin. Using TEX and to build complicated maps. TUGboat, 14(3):196–202, October 1993. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb14-3/tb40taupin-maps.pdf.

Taupin:TB14-3-203

[1608] Daniel Taupin. MusicTEX: Using TEX to write polyphonic or instrumental music. TUGboat, 14(3):203–211, October 1993. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb14-3/tb40musictex.pdf.

Siebenmann:TB14-3-212

[1609] Laurent Siebenmann. A format compilation framework for European languages. TUGboat, 14(3):212–221, October 1993. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb14-3/tb40siebenmann.pdf.

Beebe:TB14-3-222

[1610] Nelson H. F. Beebe. Bibliography prettyprinting and syntax checking. TUGboat, 14(3):222, October 1993. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb14-3/tb40beebe.pdf.

Bryan:TB14-3-223

[1611] Martin Bryan. A TEX user’s guide to ISO’s Document Style Semantics and Specification Language (DSSSL). TUGboat, 14(3):223–226, October

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-3/
tb40bryan.pdf.

Flynn:TB14-3-227

[1612] Peter Flynn. TEX and SGML: a recipe for disaster? TUGboat, 14(3):227–230, October 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-3/tb40flynn.pdf.

Wang:TB14-3-231

[1613] Xinxin Wang and Derick Wood. An abstract model for tables. TUGboat, 14 (3):231–237, October 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-3/tb40wang.pdf.

Lavaud:TB14-3-238

[1614] Michel Lavaud. Developing a multi-windowing environment for research based on TEX. TUGboat, 14(3):238–244, October 1993. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb14-3/tb40lavaud.pdf.

Kawaguti:TB14-3-245

[1615] Minato Kawaguti. A versatile TEX device driver. TUGboat, 14(3):245–251, October 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-3/tb40kawaguti.pdf.

Feruglio:TB14-3-252

[1616] Gabriel Valiente Feruglio and Robert Fuster. Typesetting Catalan texts with TEX. TUGboat, 14(3):252–259, October 1993. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb14-3/tb40valiente.pdf. See

comments [1671] and reply [1672].

Haralambous:TB14-3-260

[1617] Yannis Haralambous. The Khmer script tamed by the Lion (of TEX). TUGboat, 14(3):260–270, October 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-3/tb40hara.pdf.

Plaice:TB14-3-271

[1618] John Plaice. Language-dependent ligatures. TUGboat, 14(3):271–274, Octo-

ber 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-
3/tb40plaice.pdf.

Doob:TB14-3-275

[1619] Michael Doob and Craig Platt. Virtual fonts in a production environment. TUGboat, 14(3):275–281, October 1993. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb14-3/tb40doob.pdf.

Horn:TB14-3-282

[1620] Berthold K. P. Horn. Where are the math fonts? TUGboat, 14(3):282–284, October 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-3/tb40horn.pdf.

Jeffrey:TB14-3-285

[1621] Alan Jeffrey. A PostScript font installation package written in TEX. TUGboat, 14(3):285–292, October 1993. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb14-3/tb40jeffrey.pdf.

Jeffrey:TB14-3-293

[1622] Alan Jeffrey. Math font encodings: a workshop summary. TUGboat, 14 (3):293–295, October 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-3/tb40mathenc.pdf.

Bzyl:TB14-3-296

[1623] W�odek Bzyl and Tomasz Przechlewski. An application of literate programming: creating a format for the Bulletin of the Polish TUG. TUGboat, 14 (3):296–299, October 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-3/tb40bzyl.pdf.

Fine:TB14-3-300

[1624] Jonathan Fine. Galleys, space, and automata. TUGboat, 14(3):300–309, Octo-

ber 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-
3/tb40fine.pdf.

Laan:TB14-3-310

[1625] Kees van der Laan. Syntactic sugar. TUGboat, 14(3):310–318, October

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-3/
tb40laan-sugar.pdf.

Laan:TB14-3-319

[1626] Kees van der Laan. Sorting within TEX. TUGboat, 14(3):319–328, October

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-3/
tb40laan-sort.pdf.

Dyson:TB14-3-329

[1627] Mary Dyson. Teaching digital typography — the Didot Project. TUGboat, 14 (3):329–332, October 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-3/tb40dyson.pdf.

Gorbunova:TB14-3-333

[1628] Irina V. Gorbunova. Russian-speaking user: from Chi-Writer and Ventura Publisher to TEX; learning difficulties. TUGboat, 14(3):333–334, October

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-3/
tb40gorbunova.pdf.

Reese:TB14-3-335

[1629] R. Allan Reese. How to set up and maintain a TEX system. TUGboat, 14 (3):335–341, October 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-3/tb40reese.pdf.

Greenwade:TB14-3-342

[1630] George D. Greenwade. The Comprehensive TEX Archive Network (CTAN). TUGboat, 14(3):342–351, October 1993. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb14-3/tb40green.pdf.

Anonymous:TB14-3-352

[1631] Anonymous. Participants at the 1993 TUG meeting. TUGboat, 14(3):352–355, October 1993. ISSN 0896-3207.

Anonymous:TB14-3-356

[1632] Anonymous. Institutional members. TUGboat, 14(3):356–357, October

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-3/
tb40instmem.pdf.

Anonymous:TB14-3-358

[1633] Anonymous. Calendar. TUGboat, 14(3):358–359, October 1993. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb14-3/tb40calendar.
pdf.

Anonymous:TB14-3-359

[1634] Anonymous. Index of advertisers. TUGboat, 14(3):359, October 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-3/tb40calendar.
pdf.

Anonymous:TB14-3-360

[1635] Anonymous. TUG 1994 annual meeting, Santa Barbara, USA. TUGboat, 14 (3):360, October 1993. ISSN 0896-3207.

Anonymous:TB14-3-361

[1636] Anonymous. TUG individual membership application. TUGboat, 14(3):361, October 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-3/tb40ads.pdf.

Anonymous:TB14-3-362

[1637] Anonymous. TUG institutional membership application. TUGboat, 14(3):362, October 1993. ISSN 0896-3207.

Anonymous:TB14-3-363

[1638] Anonymous. Consultants. TUGboat, 14(3):363, October 1993. ISSN 0896-

3207. URL https://tug.org/consultants.html.

Anonymous:1993:IFCc

[1639] Anonymous. Inside front cover. TUGboat, 14(3):c2, October 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-3/
tb40cover2.pdf.

Anonymous:1993:TPd

[1640] Anonymous. Title page. TUGboat, 14(4):367, December 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-4/
tb41titlepage.pdf.

Anonymous:1993:EIc

[1641] Anonymous. Editorial information. TUGboat, 14(4):368, December 1993. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-4/tb41edit.pdf.

Anonymous:TB14-4-369

[1642] Anonymous. Addresses. TUGboat, 14(4):369–370, December 1993. ISSN 0896-3207.

Thiele:TB14-4-371

[1643] Christina Thiele. Opening words. TUGboat, 14(4):371, December

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-4/
tb41thiele.pdf.

Beeton:TB14-4-372

[1644] Barbara Beeton. Editorial comments. TUGboat, 14(4):372–373, December

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-4/
tb41beet.pdf.

Anonymous:TB14-4-373

[1645] Anonymous. TUGboat wish list. TUGboat, 14(4):373, December 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-4/tb41beet.
pdf.

Clark:TB14-4-374

[1646] Malcolm Clark. NEXTEX: A personal view. TUGboat, 14(4):374–380, Decem-

ber 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-
4/tb41clark.pdf.

Taylor:TB14-4-381

[1647] Philip Taylor. NTS update. TUGboat, 14(4):381–382, December 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-4/tb41taylor.
pdf.

Becker:TB14-4-382

[1648] Thomas Becker. Two extensions to GNU Emacs that are useful when editing TEX documents. TUGboat, 14(4):382–386, December 1993. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb14-4/tb41becker.pdf.

Knuth:TB14-4-387

[1649] Donald E. Knuth. Icons for TEX and . TUGboat, 14(4):387–389, December 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-4/tb41knut.pdf.

Liebl:TB14-4-390

[1650] Armin Liebl. bibview: A graphical user interface to BibTEX. TUGboat, 14(4): 390–395, December 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-4/tb41liebl.pdf.

Beebe:TB14-4-395

[1651] Nelson H. F. Beebe. Bibliography prettyprinting and syntax checking. TUGboat, 14(4):395–419, December 1993. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb14-4/tb41beebe.pdf.

Carmody:TB14-4-420

[1652] Kevin Carmody. A tough table becomes easy with PICTEX. TUGboat, 14 (4):420, December 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-4/tb41carmody.pdf.

Beccari:TB14-4-421

[1653] Claudio Beccari. Book Review: P. W. Abrahams, K. Berry and K. Hargreaves, TEX per l’impaziente. TUGboat, 14(4):421–422, December 1993. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb14-4/tb41reviews.
pdf.

Aslaksen:TB14-4-423

[1654] Helmer Aslaksen. Ten TEX tricks for the mathematician. TUGboat, 14(4): 423–424, December 1993. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb14-4/tb41aslaksen.pdf.

Eijkhout:TB14-4-424

[1655] Victor Eijkhout. The bag of tricks. TUGboat, 14(4):424, December

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-4/
tb41eijkhout.pdf.

Rhead:TB14-4-425

[1656] David Rhead. The “operational requirement” for support of bibliographies. TUGboat, 14(4):425–433, December 1993. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb14-4/tb41rhead.pdf.

Bland:TB14-4-433

[1657] Richard Bland. Relative moves in L^{A}TEX pictures. TUGboat, 14(4):433–437, December 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-4/tb41bland.pdf.

Anonymous:TB14-4-438

[1658] Anonymous. Calendar. TUGboat, 14(4):438–439, December 1993. ISSN

0896-3207. URL https://tug.org/TUGboat/tb14-4/tb41calendar.
pdf.

Beeton:TB14-4-439

[1659] Barbara Beeton. Production notes. TUGboat, 14(4):439, December

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-4/
tb41notes.pdf.

Anonymous:TB14-4-440

[1660] Anonymous. Coming next issue. TUGboat, 14(4):440, December

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-4/
tb41notes.pdf.

Anonymous:TB14-4-441

[1661] Anonymous. Institutional members. TUGboat, 14(4):441–442, December

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-4/
tb41instmem.pdf.

Anonymous:TB14-4-442

[1662] Anonymous. Index of advertisers. TUGboat, 14(4):442, December

1993. ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-4/
tb41instmem.pdf.

Anonymous:TB14-4-443

[1663] Anonymous. TUG membership application. TUGboat, 14(4):443–444, December 1993. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB14-4-445

[1664] Anonymous. TEX consulting and production services. TUGboat, 14(4):445, December 1993. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb14-4/tb41ads.pdf.

Anonymous:1993:IFCd

[1665] Anonymous. Inside front cover. TUGboat, 14(4):c2, December 1993. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb14-4/
tb41cover2.pdf.

Anonymous:TB15-1-3

[1666] Anonymous. Addresses. TUGboat, 15(1):3–4, March 1994. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb15-1/tb42addresses.pdf.

Thiele:TB15-1-5

[1667] Christina Thiele. Opening words. TUGboat, 15(1):5–6, March 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-1/tb42thiele.
pdf.

Beeton:TB15-1-6

[1668] Barbara Beeton. Editorial comments. TUGboat, 15(1):6–7, March 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-1/tb42beet.
pdf.

Arseneau:TB15-1-7

[1669] Donald Arseneau, Raymond Chen, and Victor Eijkhout. The TEX hierarchy. TUGboat, 15(1):7–9, March 1994. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb15-1/tb42arseneau.pdf.

Beccari:TB15-1-9

[1670] Claudio Beccari. Typesetting of ancient languages. TUGboat, 15(1):9–16, March 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-1/tb42becc-ancient.pdf.

Beccari:TB15-1-17

[1671] Claudio Beccari. Comments on the paper “Typesetting Catalan texts with TEX” (TUGboat 14, no. (3), pp. 252–259). TUGboat, 15(1):17, March 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-1/tb42becc-

catalan.pdf. See [1616, 1672].

Feruglio:TB15-1-17

[1672] Gabriel Valiente Feruglio and Robert Fuster. Comments on the comments: Typesetting Catalan texts with TEX. TUGboat, 15(1):17–18, March

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-1/
tb42valiente.pdf. See [1616, 1671].

Poppelier:TB15-1-18

[1673] Nico Poppelier. George Grätzer, Math into TEX. TUGboat, 15(1):18–19, March

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-1/
tb42poppelier-gratzer.pdf.

Pappas:TB15-1-19

[1674] T. L. (Frank) Pappas. Stephan von Bechtolsheim, TEX in Practice. TUGboat, 15(1):19–21, March 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-1/tb42pappas.pdf.

Beccari:TB15-1-21

[1675] Claudio Beccari. Gianni Gilardi, Il TEX — Introduzione al linguaggio e complementi avanzati (Italian) [TEX: Introduction to the language and advanced concepts]. TUGboat, 15(1):21–22, March 1994. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb15-1/tb42becc-gilardi.pdf.

Sawdey:TB15-1-22

[1676] Merry Obrecht Sawdey. Erik Spiekermann & E. M. Ginger, Stop Stealing Sheep. TUGboat, 15(1):22–24, March 1994. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb15-1/tb42sawdey.pdf.

Poppelier:TB15-1-24

[1677] Nico Poppelier. Eric van Herwijnen, Practical SGML. TUGboat, 15(1):24– 25, March 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-1/tb42poppelier-herwijnen.pdf.

Detig:TB15-1-25

[1678] Christine Detig and Joachim Schrod. Donald E. Knuth, Literate Programming. TUGboat, 15(1):25–27, March 1994. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb15-1/tb42detig.pdf.

Salomon:TB15-1-28

[1679] David Salomon. Output routines: Examples and techniques, Part IV: Horizontal techniques. TUGboat, 15(1):28–40, March 1994. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb15-1/tb42salomon-
output.pdf.

Salomon:TB15-1-40

[1680] David Salomon. Verbatim copying and listing. TUGboat, 15(1):40–54, March

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-1/
tb42salomon-verbatim.pdf.

Eijkhout:TB15-1-55

[1681] Victor Eijkhout. The bag of tricks. TUGboat, 15(1):55–57, March

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-1/
tb42eijkhout.pdf.

Meer:TB15-1-57

[1682] Hans van der Meer. Random bit generator in TEX. TUGboat, 15(1):57–58, March 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-1/tb42meer.pdf.

Salomon:TB15-1-59

[1683] David Salomon and Matthew N. Hendryx. Slanted lines with controlled thickness. TUGboat, 15(1):59–62, March 1994. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb15-1/tb42salomon-slant.pdf.

Fine:TB15-1-62

[1684] Jonathan Fine. On indexing and errors. TUGboat, 15(1):62, March 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-1/tb42finl.
pdf.

Rowley:TB15-1-63

[1685] Chris Rowley. L^{A}TEX 2ε update, dateline: 31 January 1994. TUGboat, 15(1): 63, March 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-1/tb42rowley.pdf.

Rhead:TB15-1-64

[1686] David Rhead. Addenda: A suggested “operational requirement” for L^{A}TEX3’s treatment of bibliographic references (TUGboat 1, no. 4(4), pp. 425-433). TUGboat, 15(1):64–65, March 1994. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb15-1/tb42rhead.pdf.

Anonymous:TB15-1-66

[1687] Anonymous. Calendar. TUGboat, 15(1):66–67, March 1994. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb15-1/tb42calendar.pdf.

Anonymous:TB15-1-68

[1688] Anonymous. TUG ’94 — Announcement and preliminary program (Santa Barbara, California, 31 July–4 August 1994). TUGboat, 15(1):68–69, March

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-1/
tb42ann.pdf.

Anonymous:TB15-1-69

[1689] Anonymous. Call for papers, EuroTEX ’94 (Gdansk, Poland, 26–30 September 1994). TUGboat, 15(1):69–70, March 1994. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb15-1/tb42ann.pdf.

Anonymous:TB15-1-70

[1690] Anonymous. Meet the Board, Part I. TUGboat, 15(1):70–73, March

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-1/
tb42board.pdf.

Beeton:TB15-1-74

[1691] Barbara Beeton. Production notes. TUGboat, 15(1):74, March 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-1/tb42news.
pdf.

Anonymous:TB15-1-75

[1692] Anonymous. Coming next issue. TUGboat, 15(1):75, March 1994. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb15-1/tb42instmem.
pdf.

Anonymous:TB15-1-75-2

[1693] Anonymous. Institutional members. TUGboat, 15(1):75–76, March

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-1/
tb42instmem.pdf.

Anonymous:TB15-1-77

[1694] Anonymous. TUG membership application. TUGboat, 15(1):77–78, March

1994. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB15-1-79

[1695] Anonymous. TEX consulting and production services. TUGboat, 15 (1):79, March 1994. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB15-1-81

[1696] Anonymous. Index of advertisers. TUGboat, 15(1):81, March 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-1/tb42ads.
pdf.

Anonymous:1994:A

[1697] Anonymous. Addresses. TUGboat, 15(2):3–4, June 1994. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43addresses.pdf.

Anonymous:TB15-2-87

[1698] Anonymous. Addresses. TUGboat, 15(2):87–88, June 1994. ISSN 0896-3207.

Thiele:TB15-2-89

[1699] Christina Thiele. Opening words. TUGboat, 15(2):89, June 1994. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb15-2/tb43thiele.
pdf.

Beeton:TB15-2-89

[1700] Barbara Beeton. Editorial comments. TUGboat, 15(2):89–91, June 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-2/tb43beet.
pdf.

Goossens:TB15-2-92

[1701] Michel Goossens. WEPT: A Week on Electronic Publishing and Typography. TUGboat, 15(2):92–95, June 1994. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb15-2/tb43goossens.pdf.

Taylor:TB15-2-96

[1702] Philip Taylor. Report of the 2nd meeting of the NTS group, February 1994. TUGboat, 15(2):96–97, June 1994. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb15-2/tb43taylor.pdf.

Hoenig:TB15-2-97

[1703] Alan Hoenig. {Meta} Font Forum redux. TUGboat, 15(2):97, June

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43hoenig-fonts.pdf.

Ramasubramanian:TB15-2-98

[1704] R. Ramasubramanian, R. W. D. Nickalls, and M. A. Reed. ASCII.sty: A new style-option and encoded font with IBM graphics control characters for use with TEX and L^{A}TEX. TUGboat, 15(2):98–103, June

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43ramasubramanian.pdf.

Jones:TB15-2-103

[1705] David M. Jones and David E. Wald. Michel Goossens, Frank Mittelbach, and Alexander Samarin, The L^{A}TEX Companion. TUGboat, 15(2):103–106, June

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43jones.pdf.

Eijkhout:TB15-2-106

[1706] Victor Eijkhout. Norman Walsh, Making TEX Work. TUGboat, 15(2):106–107,

June 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-
2/tb43eijkhout.pdf.

Andre:TB15-2-107

[1707] Jacques André. Christian Rolland, L^{A}TEX guide pratique. TUGboat, 15(2):107,

June 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-
2/tb43andre.pdf.

Hoenig:TB15-2-107

[1708] Alan Hoenig. NextTEX: TEX plus the NextStep Operating System. TUGboat, 15(2):107–109, June 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-2/tb43hoenig-next.pdf.

Downes:TB15-2-110

[1709] Michael Downes. Interaction tools: dialog.sty and menus.sty. TUGboat, 15 (2):110–130, June 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-2/tb43downes.pdf.

Gaulle:TB15-2-131

[1710] Bernard Gaulle. L^{A}TEX V3: philology & typography, reports to read, reports to do. TUGboat, 15(2):131, June 1994. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb15-2/tb43gaulle.pdf.

Beccari:TB15-2-131

[1711] Claudio Beccari. Tough table becomes easy with PICTEX, but it’s even easier with L^{A}TEX. TUGboat, 15(2):131–132, June 1994. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb15-2/tb43becc.pdf.

Anagnostopolous:TB15-2-132

[1712] Paul Anagnostopolous. On the review of TEX in Practice. TUGboat, 15(2): 132, June 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-2/tb43anag.pdf.

Anonymous:TB15-2-133

[1713] Anonymous. Baskerville, Volume 4, Nos. 1–2. TUGboat, 15(2):133–136, June

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43abstracts.pdf.

Anonymous:TB15-2-136

[1714] Anonymous. Cahiers GUTenberg Nos. 16 and 17. TUGboat, 15(2):136–137,

June 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-
2/tb43abstracts.pdf.

Anonymous:TB15-2-138

[1715] Anonymous. Die TEXnische Komödie 1993, Heft 1–4. TUGboat, 15(2):138– 142, June 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-2/tb43abstracts.pdf.

Anonymous:TB15-2-143

[1716] Anonymous. Calendar. TUGboat, 15(2):143–144, June 1994. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb15-2/tb43calendar.
pdf.

Anonymous:TB15-2-145

[1717] Anonymous. Meet the Board, Part II. TUGboat, 15(2):145–148, June

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43board.pdf.

Beeton:TB15-2-148

[1718] Barbara Beeton. Production notes. TUGboat, 15(2):148–149, June 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-2/tb43news.
pdf.

Anonymous:TB15-2-149

[1719] Anonymous. Coming next issue. TUGboat, 15(2):149, June 1994. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb15-2/tb43news.pdf.

Beeton:TB15-2-150

[1720] Barbara Beeton. Change in TUGboat policy. TUGboat, 15(2):150, June

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43news.pdf.

Anonymous:TB15-2-151

[1721] Anonymous. Institutional members. TUGboat, 15(2):151–152, June

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43instmem.pdf.

Anonymous:TB15-2-152

[1722] Anonymous. TEX consulting and production services. TUGboat, 15(2): 152–153, June 1994. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB15-2-154

[1723] Anonymous. Index of advertisers. TUGboat, 15(2):154, June 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-2/tb43ads.
pdf.

Anonymous:TB15-2-155

[1724] Anonymous. TUG membership application. TUGboat, 15(2):155–156, June

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43ads.pdf.

Anonymous:TB15-2-160

[1725] Anonymous. TUG ’95 — St. Petersburg, Florida. TUGboat, 15(2):160, June

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-2/
tb43news.pdf.

Anonymous:TB15-3-166

[1726] Anonymous. Editorial and production notes. TUGboat, 15(3):166, September

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-3/
tb44edit.pdf.

Anonymous:TB15-3-167

[1727] Anonymous. Acknowledgements and Conference Program: Innovation. TUGboat, 15(3):167–168, September 1994. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb15-3/tb44confinfo.pdf.

Bigelow:TB15-3-169

[1728] Charles Bigelow. Lucida and TEX: lessons of logic and history. TUGboat, 15(3):169, September 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-3/tb44bigelow.pdf.

Mittelbach:TB15-3-170

[1729] Frank Mittelbach and Michel Goossens. Real life book production — lessons learned from The L^{A}TEX Companion. TUGboat, 15(3):170–173, September

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-3/
tb44goossens.pdf.

Haralambous:TB15-3-174

[1730] Yannis Haralambous. Typesetting the holy Bible in Hebrew, with TEX. TUGboat, 15(3):174–191, September 1994. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb15-3/tb44haralambous-hebrew.pdf.

Cohen:TB15-3-192

[1731] Michel Cohen. Adaptive character generation and spatial expressiveness. TUGboat, 15(3):192–198, September 1994. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb15-3/tb44cohen.pdf.

Haralambous:TB15-3-199

[1732] Yannis Haralambous. Humanist. TUGboat, 15(3):199, September 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-3/tb44haralambous-
humanist.pdf.

Malyshev:TB15-3-200

[1733] Basil Malyshev. Automatic conversion of fonts to Type-1 Post-Script. TUGboat, 15(3):200, September 1994. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb15-3/tb44malyshev.pdf.

Hafner:TB15-3-201

[1734] James Hafner. The (pre)history of color in Rokicki’s dvips. TUGboat, 15(3): 201–204, September 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-3/tb44hafner.pdf.

Rokicki:TB15-3-205

[1735] Tom Rokicki. Advanced ‘special’ support in a dvi driver. TUGboat, 15(3): 205–212, September 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-3/tb44rokicki.pdf.

Duggan:TB15-3-213

[1736] Angus Duggan. Colour separation and PostScript. TUGboat, 15(3): 213–217, September 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-3/tb44duggan.pdf.

Rahtz:TB15-3-218

[1737] Sebastian Rahtz and Michel Goossens. Simple colour design with L^{A}TEX 2ε. TUGboat, 15(3):218–222, September 1994. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb15-3/tb44rahtz.pdf.

Sowa:TB15-3-223

[1738] Friedhelm Sowa. Printing colour pictures. TUGboat, 15(3):223–227, Septem-

ber 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-
3/tb44sowa.pdf.

Sofka:TB15-3-228

[1739] Michael Sofka. Color book production using TEX. TUGboat, 15(3):228–238, September 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-3/tb44sofka.pdf.

Zandt:TB15-3-239

[1740] Timothy van Zandt and Denis Girou. Inside PSTricks. TUGboat, 15(3): 239–246, September 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-3/tb44tvz.pdf.

Stenerson:TB15-3-247

[1741] Jon Stenerson. A L^{A}TEX style file generator. TUGboat, 15(3):247–254, Septem-

ber 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-
3/tb44stenerson.pdf.

Braams:TB15-3-255

[1742] Johannes Braams. Document classes and packages in L^{A}TEX 2ε. TUGboat, 15(3):255–262, September 1994. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb15-3/tb44braams.pdf.

Jeffrey:TB15-3-263

[1743] Alan Jeffrey. PostScript font support in L^{A}TEX 2ε. TUGboat, 15(3): 263–268, September 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-3/tb44jeff.pdf.

Patashnik:TB15-3-269

[1744] Oren Patashnik. BibTEX 1.0. TUGboat, 15(3):269–273, September

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-3/
tb44patashnik.pdf.

MacKay:TB15-3-274

[1745] Pierre MacKay. A typesetter’s toolkit. TUGboat, 15(3):274–284, September

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-3/
tb44mackay.pdf.

Barnett:TB15-3-285

[1746] Michael P. Barnett and Kevin R. Perry. Symbolic computation for electronic publishing. TUGboat, 15(3):285–292, September 1994. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb15-3/tb44barnett.pdf.

Kawaguti:TB15-3-293

[1747] Minato Kawaguti and Norio Kitajima. Concurrent use of interactive TEX previewer with an Emacs-type editor. TUGboat, 15(3):293–300, September

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-3/
tb44kawaguti.pdf.

Haralambous:TB15-3-301

[1748] Yannis Haralambous. An Indic TEX preprocessor — Sinhalese TEX. TUGboat, 15(3):301, September 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-3/tb44haralambous-indic.pdf.

Doumont:TB15-3-302

[1749] Jean-Luc Doumont. Pascal pretty-printing: an example of “preprocessing within TEX”. TUGboat, 15(3):302–307, September 1994. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb15-3/tb44doumont.pdf.

Schrod:TB15-3-309

[1750] Joachim Schrod. Towards interactivity for TEX. TUGboat, 15(3):309–317, September 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-3/tb44schrod.pdf.

Rowley:TB15-3-318

[1751] Chris Rowley and Frank Mittelbach. The floating world. TUGboat, 15(3):318, September 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-3/tb44rowl.pdf.

Hosek:TB15-3-319

[1752] Don Hosek. Sophisticated page layout with TEX. TUGboat, 15(3):319, September 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-3/tb44hosek.pdf.

Plaice:TB15-3-320

[1753] John Plaice. Progress in the Omega project. TUGboat, 15(3):320–324, Septem-

ber 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-
3/tb44plaice.pdf.

Ogawa:TB15-3-325

[1754] Arthur Ogawa. Object-oriented programming, descriptive markup, and TEX. TUGboat, 15(3):325–330, September 1994. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb15-3/tb44ogawa.pdf.

Baxter:TB15-3-331

[1755] William Erik Baxter. An object-oriented programming system in TEX. TUGboat, 15(3):331–338, September 1994. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb15-3/tb44baxter.pdf.

Walsh:TB15-3-339

[1756] Norm Walsh. A World Wide Web interface to CTAN. TUGboat, 15(3):339–343, September 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-3/tb44walsh.pdf.

Haralambous:TB15-3-344

[1757] Yannis Haralambous and John Plaice. First applications of Ω: Adobe Poetica, Arabic, Greek, Khmer. TUGboat, 15(3):344–352, September

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-3/
tb44haralambous-omega.pdf.

Taylor:TB15-3-353

[1758] Philip Taylor. ε-TEX & N T S: A progress report. TUGboat, 15(3):353–358, September 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-3/tb44taylor.pdf.

Laugier:TB15-3-359

[1759] Maurice Laugier and Yannis Haralambous. TEX innovations by the Louis-Jean Printing House. TUGboat, 15(3):359, September 1994. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb15-3/tb44laugier.pdf.

Downes:TB15-3-360

[1760] Michael Downes. Design by template in a production macro package. TUGboat, 15(3):360–368, September 1994. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb15-3/tb44downes.pdf.

Hoenig:TB15-3-369

[1761] Alan Hoenig. Less is More: Restricting TEX’s scope enables complex page layouts. TUGboat, 15(3):369–380, September 1994. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb15-3/tb44hoenig.pdf.

Fine:TB15-3-381

[1762] Jonathan Fine. Documents, compuscripts, programs and macros. TUGboat, 15(3):381–385, September 1994. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb15-3/tb44fine.pdf.

Grobelnik:TB15-3-386

[1763] Marko Grobelnik, Dunja Mladenić, Darko Zupanič, and Borut Žnidar. Integrated system for encyclopaedia typesetting based on TEX. TUGboat, 15(3): 386–387, September 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-3/tb44grobelnik.pdf.

Baragar:TB15-3-388

[1764] Henry Baragar and Gail E. Harris. An example of a special purpose input language to L^{A}TEX. TUGboat, 15(3):388–396, September 1994. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb15-3/tb44baragar.pdf.

Anonymous:TB15-3-397

[1765] Anonymous. Color pages. TUGboat, 15(3):397–404, September 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-3/tb44color.
pdf.

Anonymous:TB15-3-405

[1766] Anonymous. Participants at the Annual Meeting. TUGboat, 15(3):405–407, September 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-3/tb44addresses.pdf.

Anonymous:TB15-3-408

[1767] Anonymous. Calendar. TUGboat, 15(3):408–409, September 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-3/tb44calendar.
pdf.

Anonymous:TB15-3-409

[1768] Anonymous. 1995 Knuth Scholarship. TUGboat, 15(3):409–410, September

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-3/
tb44knuthscholar.pdf.

Anonymous:TB15-3-411

[1769] Anonymous. Institutional members. TUGboat, 15(3):411–412, September

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-3/
tb44instmem.pdf.

Anonymous:TB15-3-412

[1770] Anonymous. TEX consulting and production services. TUGboat, 15(3): 412–413, September 1994. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB15-3-415

[1771] Anonymous. TUG membership application. TUGboat, 15(3):415–416, September 1994. ISSN 0896-3207. URL https://tug.org/join.

html.

Anonymous:TB15-3-417

[1772] Anonymous. Index of advertisers. TUGboat, 15(3):417, September

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-3/
tb44ads.pdf.

Anonymous:TB15-4-423

[1773] Anonymous. Addresses. TUGboat, 15(4):423–424, December 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-4/tb45complete.
pdf.

Thiele:TB15-4-425

[1774] Christina Thiele. Opening words. TUGboat, 15(4):425, December 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-4/tb45thie.
pdf.

Beeton:TB15-4-426

[1775] Barbara Beeton. Editorial comments. TUGboat, 15(4):426–427, December

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-4/
tb45beet.pdf.

Goossens:TB15-4-427

[1776] Michel Goossens, Sebastian Rahz, and Barbara Beeton. The TUG94 Proceedings — Apologia. TUGboat, 15(4):427–428, December 1994. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb15-4/tb45goos.pdf.

Bolland:TB15-4-429

[1777] Julita Bolland, Toni Walter, and W�odek Bzyl. EuroTEX94. TUGboat, 15(4): 429–433, December 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-4/tb45boll.pdf.

Taylor:TB15-4-434

[1778] Philip Taylor. Minutes of the NTS meeting held at Lindau on October 11/12th 1994. TUGboat, 15(4):434–437, December 1994. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb15-4/tb45ntsx.pdf.

Laugier:TB15-4-438

[1779] Maurice Laugier and Yannis Haralambous. TEX innovations at the Louis-Jean printing house. TUGboat, 15(4):438–443, December 1994. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb15-4/tb45laugier.pdf.

MacKay:TB15-4-444

[1780] Pierre A. MacKay. Recycled . TUGboat, 15(4):444–446, Decem-

ber 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-
4/tb45mack.pdf.

Haralambous:TB15-4-447

[1781] Yannis Haralambous. Indica, an Indic preprocessor for TEX: A Sinhalese TEX system. TUGboat, 15(4):447–458, December 1994. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb15-4/tb45hari.pdf.

Olsak:TB15-4-459

[1782] Petr Olšák. The EAN barcodes by TEX. TUGboat, 15(4):459–464, December

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-4/
tb45olsa.pdf.

Taylor:TB15-4-464

[1783] Philip Taylor. An open letter to the TUG Board. TUGboat, 15(4):464–465, December 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-4/tb45tayl.pdf.

Feruglio:TB15-4-466

[1784] Gabriel Valiente Feruglio. Typesetting commutative diagrams. TUGboat, 15 (4):466–484, December 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-4/tb45vali.pdf.

Eijkhout:TB15-4-485

[1785] Victor Eijkhout. The bag of tricks. TUGboat, 15(4):485–405, December

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-4/
tb45eijk.pdf.

Richer:TB15-4-486

[1786] Jacques Richer. A TEX autostereogram generator. TUGboat, 15(4):486–492, December 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-4/tb45rich.pdf.

Fossmeier:TB15-4-492

[1787] Reinhard Fößmeier. Stereographic pictures using TEX. TUGboat, 15(4): 492–495, December 1994. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb15-4/tb45foss.pdf.

Braams:TB15-4-496

[1788] Johannes Braams. To reset or not to reset. TUGboat, 15(4):496–497, December

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-4/
tb45braa.pdf.

Anonymous:TB15-4-498

[1789] Anonymous. Cahiers GUTenberg Nos. 18 and 19. TUGboat, 15(4):498–499, December 1994. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb15-4/tb45cahi.pdf.

Anonymous:TB15-4-500

[1790] Anonymous. Abstracts of the Proceedings of the Eighth European TEX Conference, Gdańsk, September 1994. TUGboat, 15(4):500–503, December

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-4/
tb45eurotex8.pdf.

Anonymous:1994:C

[1791] Anonymous. Calendar. TUGboat, 15(4):504, December 1994. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-4/
tb45calendar.pdf.

Anonymous:TB15-4-505

[1792] Anonymous. Calendar. TUGboat, 15(4):505–504, December 1994. ISSN

0896-3207. URL https://tug.org/TUGboat/tb15-4/tb45prod.
pdf.

Beeton:TB15-4-505

[1793] Barbara Beeton. Production notes. TUGboat, 15(4):505, December

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-4/
tb45prod.pdf.

Anonymous:TB15-4-506

[1794] Anonymous. Institutional members. TUGboat, 15(4):506, December

1994. ISSN 0896-3207. URL https://tug.org/TUGboat/tb15-4/
tb45instmem.pdf.

Anonymous:TB15-4-507

[1795] Anonymous. TUG membership application. TUGboat, 15(4):507, December

1994. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB15-4-508

[1796] Anonymous. TEX consulting and production services. TUGboat, 15(4): 508, December 1994. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:1995:TPa

[1797] Anonymous. Title page. TUGboat, 16(1):1, March 1995. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-1/
titlepage.pdf.

Anonymous:1995:EIa

[1798] Anonymous. Editorial information. TUGboat, 16(1):2, March 1995. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-1/
tb46edit.pdf.

Anonymous:TB16-1-3

[1799] Anonymous. Addresses. TUGboat, 16(1):3–4, March 1995. ISSN 0896-3207.

Bzyl:TB16-1-4

[1800] W�odek Bzyl. EuroTEX94 Contest Answers. TUGboat, 16(1):4, March

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-1/
tb46beet-contest.pdf.

Thiele:TB16-1-5

[1801] Christina Thiele and Michel Goossens. Opening words. TUGboat, 16(1):5– 7, March 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-1/tb46pres.pdf.

Beeton:TB16-1-8

[1802] Barbara Beeton. Editorial comments. TUGboat, 16(1):8, March 1995. ISSN

0896-3207. URL https://tug.org/TUGboat/tb16-1/tb46beet.
pdf.

Jaegermann:TB16-1-9

[1803] Michael Jaegermann. Making MakeTeXPK safer for Unix installations. TUGboat, 16(1):9–12, March 1995. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb16-1/tb46jaeg.pdf.

Beeton:TB16-1-12

[1804] Barbara Beeton. Hyphenation exception log. TUGboat, 16(1):12–17, March

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-1/
tb46hyf.pdf.

Beccari:TB16-1-18

[1805] Claudio Beccari. Configuring TEX or L^{A}TEX for typesetting in several languages. TUGboat, 16(1):18–30, March 1995. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb16-1/tb46bec1.pdf.

Beccari:TB16-1-30

[1806] Claudio Beccari, Radu Oprea, and Elena Tulei. How to make a foreign language pattern file: Romanian. TUGboat, 16(1):30–41, March 1995. ISSN

0896-3207. URL https://tug.org/TUGboat/tb16-1/tb46bec2.
pdf.

Thiele:TB16-1-42

[1807] Christina Thiele. TEX and Linguistics. TUGboat, 16(1):42–44, March

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-1/
tb46ling.pdf.

Hoenig:TB16-1-45

[1808] Alan Hoenig. Introducing . TUGboat, 16(1):45, March 1995. ISSN

0896-3207. URL https://tug.org/TUGboat/tb16-1/tb46hoen.
pdf.

Haralambous:TB16-1-46

[1809] Yannis Haralambous. Some techniques. TUGboat, 16(1):46– 53, March 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-1/tb46harm.pdf.

Olsak:TB16-1-54

[1810] Petr Olšák. The program a2ac — Font handling on the PostScript level. TUGboat, 16(1):54–59, March 1995. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb16-1/tb46olsa.pdf.

Malyshev:TB16-1-60

[1811] Basil Malyshev. Problems of the conversion of fonts to Post-Script Type 1. TUGboat, 16(1):60–68, March 1995. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb16-1/tb46malyPT.pdf.

Malyshev:TB16-1-69

[1812] Basil Malyshev and Michel Goossens. Partial font embedding utilities for PostScript Type-1 fonts. TUGboat, 16(1):69–77, March 1995. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb16-1/tb46malyFL.
pdf.

Jeffrey:TB16-1-78

[1813] Alan Jeffrey. Tight setting with TEX. TUGboat, 16(1):78–80, March

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-1/
tb46jeff.pdf.

Fujita:TB16-1-80

[1814] Shinsaku Fujita. XΥMTEX for drawing chemical structural formulas. TUGboat, 16(1):80–88, March 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-1/tb46fuji.pdf.

Anonymous:TB16-1-89

[1815] Anonymous. Calendar. TUGboat, 16(1):89–90, March 1995. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Burbank:TB16-1-90

[1816] Mimi Burbank. Production notes. TUGboat, 16(1):90–91, March 1995. ISSN

0896-3207. URL https://tug.org/TUGboat/tb16-1/tb46prod-
coming.pdf.

Anonymous:TB16-1-91

[1817] Anonymous. Coming next issue. TUGboat, 16(1):91, March 1995. ISSN

0896-3207. URL https://tug.org/TUGboat/tb16-1/tb46prod-
coming.pdf.

Anonymous:TB16-1-92

[1818] Anonymous. Institutional members. TUGboat, 16(1):92, March 1995. ISSN 0896-3207.

Anonymous:TB16-1-93

[1819] Anonymous. TUG membership application. TUGboat, 16(1):93, March 1995.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB16-1-94

[1820] Anonymous. TEX consulting and production services. TUGboat, 16 (1):94, March 1995. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:1995:FCa

[1821] Anonymous. Front cover. TUGboat, 16(1):c1, March 1995. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-1/
cover1.pdf.

Anonymous:1995:IFCa

[1822] Anonymous. Inside front cover. TUGboat, 16(1):c2, March 1995. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-1/
cover2.pdf.

Anonymous:TB16-2-99

[1823] Anonymous. Addresses. TUGboat, 16(2):99–100, June 1995. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb16-2/tb47complete.
pdf.

Burbank:TB16-2-100

[1824] Mimi Burbank and Barbara Beeton. Production notes. TUGboat, 16(2):100,

June 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-
2/tb47prod.pdf.

Goossens:TB16-2-101

[1825] Michel Goossens. Opening words. TUGboat, 16(2):101, June 1995. ISSN

0896-3207. URL https://tug.org/TUGboat/tb16-2/gendel.
pdf.

Clark:TB16-2-101

[1826] Malcolm Clark. Introduction. TUGboat, 16(2):101–102, June 1995. ISSN

0896-3207. URL https://tug.org/TUGboat/tb16-2/gendel.
pdf.

Goossens:TB16-2-103

[1827] Michel Goossens and Janne Saarela. A practical introduction to SGML. TUGboat, 16(2):103–145, June 1995. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb16-2/tb47goo1.pdf.

Flynn:TB16-2-146

[1828] Peter Flynn. HTML & TEX: Making them sweat. TUGboat, 16(2):146–150,

June 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-
2/pf-gg-md-os.pdf.

Granger:TB16-2-151

[1829] Geeti Granger. The inside story of life at Wiley with SGML, L^{A}TEX and Acrobat. TUGboat, 16(2):151–154, June 1995. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb16-2/pf-gg-md-os.pdf.

Doyle:TB16-2-154

[1830] Mark D. Doyle. The Los Alamos e-print archives: HyperTEX in action. TUGboat, 16(2):154–159, June 1995. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb16-2/pf-gg-md-os.pdf.

Schwarzkopf:TB16-2-159

[1831] Otfried Schwarzkopf. The Hyperlatex story. TUGboat, 16(2):159–162, June

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-2/
pf-gg-md-os.pdf.

Haralambous:TB16-2-162

[1832] Yannis Haralambous and Sebastian Rahtz. L^{A}TEX, hypertext and PDF, or the entry of TEX into the world of hypertext. TUGboat, 16(2):162–173, June

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-2/
tb47hara.pdf.

Goossens:TB16-2-174

[1833] Michel Goossens and Janne Saarela. TEX to HTML and back. TUGboat, 16(2):174–214, June 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-2/tb47goo2.pdf.

Anonymous:TB16-2-215

[1834] Anonymous. Calendar. TUGboat, 16(2):215–216, June 1995. ISSN 0896-

3207. URL https://tug.org/calendar.html.

Anonymous:TB16-2-217

[1835] Anonymous. Institutional members. TUGboat, 16(2):217, June 1995. ISSN 0896-3207.

Anonymous:TB16-2-218

[1836] Anonymous. TEX consulting and production services. TUGboat, 16(2):218,

June 1995. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB16-2-218-2

[1837] Anonymous. Index of advertisers. TUGboat, 16(2):218, June 1995. ISSN

0896-3207. URL https://tug.org/consultants.html.

Fairbairns:TB16-3-222

[1838] Robin Fairbairns. Production notes. TUGboat, 16(3):222, September

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-3/
tb48prod.pdf.

Goossens:TB16-3-223

[1839] Michel Goossens. President’s words. TUGboat, 16(3):223–226, Septem-

ber 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/Pres/
tb48pres.html.

Zlatuska:TB16-3-227

[1840] Jiří Zlatuška. When does it alone. TUGboat, 16(3):227–232, September 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-3/tb48zlat.pdf.

Kinch:TB16-3-233

[1841] Richard J. Kinch. MetaFog: converting shapes to contours. TUGboat, 16(3):233–243, September 1995. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb16-3/tb48kinc.pdf.

Hoenig:TB16-3-244

[1842] Alan Hoenig. The Poetica family: fancy fonts with TEX and L^{A}TEX. TUGboat, 16(3):244–252, September 1995. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb16-3/tb48hoen.pdf.

Goossens:TB16-3-253

[1843] Michel Goossens, Sebastian Rahtz, and Robin Fairbairns. Using Adobe Type 1 Multiple Master fonts with TEX. TUGboat, 16(3):253–258, September

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-3/
tb48goos.pdf.

Gibbons:TB16-3-259

[1844] Jeremy Gibbons. Dotted and dashed lines in . TUGboat, 16(3): 259–264, September 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-3/tb48gibb.pdf.

Lesenko:TB16-3-265

[1845] Sergey Lesenko. Printing TEX documents with partial Type 1 fonts. TUGboat, 16(3):265–268, September 1995. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb16-3/tb48lese.pdf.

Swift:TB16-3-269

[1846] Matthew Swift. Modularity in L^{A}TEX. TUGboat, 16(3):269–275, September

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-3/
tb48swif.pdf.

Kletzing:TB16-3-276

[1847] Dennis Kletzing. A multienumerate package. TUGboat, 16(3):276–279, September 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-3/tb48klet.pdf.

Sojka:TB16-3-280

[1848] Petr Sojka and Pavel Ševeček. Hyphenation in TEX — Quo Vadis? TUGboat, 16(3):280–289, September 1995. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb16-3/tb48soj1.pdf.

Sojka:TB16-3-290

[1849] Petr Sojka. Notes on compound word hyphenation in TEX. TUGboat, 16(3): 290–296, September 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-3/tb48soj2.pdf.

Bzyl:TB16-3-297

[1850] W�odek Bzyl. Literate Plain source is available! TUGboat, 16(3):297–299, September 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-3/tb48bzyl.pdf.

Childs:TB16-3-300

[1851] Bart Childs, Deborah Dunn, and William Lively. Teaching CS/1 courses in a literate manner. TUGboat, 16(3):300–309, September 1995. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb16-3/tb48chil.pdf.

Raman:TB16-3-310

[1852] T. V. Raman. An audio view of (La)TEX documents — part II. TUGboat, 16(3): 310–314, September 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-3/tb48rama.pdf.

Rahtz:TB16-3-315

[1853] Sebastian Rahtz. Another look at L^{A}TEX to SGML conversion. TUGboat, 16(3): 315–324, September 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-3/tb48raht.pdf.

Fairbairns:TB16-3-325

[1854] Robin Fairbairns. Omega — Why bother with Unicode? TUGboat, 16(3): 325–328, September 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-3/tb48fair.pdf.

Feruglio:TB16-3-329

[1855] Gabriel Valiente Feruglio. Modern Catalan typographical conventions. TUGboat, 16(3):329–338, September 1995. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb16-3/tb48vali.pdf.

Anonymous:TB16-3-339

[1856] Anonymous. TUG’96 Announcement. TUGboat, 16(3):339, September

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-3/
dubna.pdf.

Anonymous:TB16-3-340

[1857] Anonymous. Calendar. TUGboat, 16(3):340–341, September 1995. ISSN

0896-3207. URL https://tug.org/calendar.html.

Anonymous:TB16-3-341

[1858] Anonymous. TUG’95 — List of attendees. TUGboat, 16(3):341–343, Septem-

ber 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-
3/tb48attn.pdf.

Anonymous:TB16-3-344

[1859] Anonymous. Institutional members. TUGboat, 16(3):344, September 1995. ISSN 0896-3207.

Anonymous:TB16-3-345

[1860] Anonymous. TEX consulting and production services. TUGboat, 16(3): 345, September 1995. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:1995:FCb

[1861] Anonymous. Front cover. TUGboat, 16(3):c1, September 1995. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-3/
cover-color.pdf.

Anonymous:1995:IFCb

[1862] Anonymous. Inside front cover. TUGboat, 16(3):c2, September 1995. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-3/
cover2.pdf.

Anonymous:1995:TPb

[1863] Anonymous. Title page. TUGboat, 16(4):349, December 1995. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-4/
titlepage.pdf.

Anonymous:1995:EIb

[1864] Anonymous. Editorial information. TUGboat, 16(4):350, December 1995. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-4/tb49edit.pdf.

Anonymous:TB16-4-351

[1865] Anonymous. Addresses. TUGboat, 16(4):351–352, December 1995. ISSN 0896-3207.

Goossens:TB16-4-353

[1866] Michel Goossens. Writing the future is reading the past. TUGboat, 16(4): 353–354, December 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-4/tb49pres.pdf.

Beeton:TB16-4-355

[1867] Barbara Beeton. Editorial comments. TUGboat, 16(4):355–357, December

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-4/
tb49beet-flyn.pdf.

Flynn:TB16-4-357

[1868] Peter Flynn. Upcoming TTN merger into TUGboat. TUGboat, 16(4):357, December 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-4/tb49beet-flyn.pdf.

Machi:TB16-4-358

[1869] Filip Machi, Jerrold E. Marsden, and Wendy G. McKay. Introduction to FasTEX: A system of keyboard shortcuts for the fast keying of TEX. TUGboat, 16(4):358–363, December 1995. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb16-4/tb49mars.pdf.

Lapko:TB16-4-364

[1870] Olga Lapko and Irina Makhovaya. The style russianb for Babel: Problems and solutions. TUGboat, 16(4):364–372, December 1995. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb16-4/tb49olga.pdf.

Slepukhin:TB16-4-373

[1871] Andrey Slepukhin. A package for Church-Slavonic typesetting. TUGboat, 16 (4):373–380, December 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-4/tb49slep.pdf.

Goossens:TB16-4-379

[1872] Michel Goossens. Production notes on the Russian papers. TUGboat, 16 (4):379, December 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-4/tb49slep.pdf.

Knappen:TB16-4-381

[1873] Jörg Knappen. Release 1.2 of the dc-fonts: Improvements to the European letters and first release of text companion symbols. TUGboat, 16(4):381–387, December 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-4/tb49knap.pdf.

Jackowski:TB16-4-388

[1874] Bogus�aw Jackowski. A –EPS interface. TUGboat, 16(4): 388–395, December 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-4/tb49jack.pdf.

Rokicki:TB16-4-395

[1875] Tomas G. Rokicki. A proposed standard for specials. TUGboat, 16(4): 395–401, December 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-4/tb49roki.pdf.

TWGTDS:TB16-4-401

[1876] TUG Working Group on a TEX Directory Structure. A directory structure for TEX files (Version 0.999). TUGboat, 16(4):401–413, December 1995. ISSN

0896-3207. URL https://tug.org/TUGboat/tb16-4/tb49tds.
pdf.

Rahtz:TB16-4-413

[1877] Sebastian Rahtz. Whatever is wrong with my L^{A}TEX file? TUGboat, 16(4): 413–416, December 1995. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb16-4/tb49rahtz.pdf.

Fine:TB16-4-416

[1878] Jonathan Fine. New perspectives on TEX macros. TUGboat, 16(4):416–418, December 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-4/tb49fine.pdf.

Haralambous:TB16-4-418

[1879] Yannis Haralambous, John Plaice, and Johannes Braams. Never again active characters! Ω-Babel. TUGboat, 16(4):418–427, December 1995. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb16-4/tb49har.pdf.

Anonymous:TB16-4-427

[1880] Anonymous. TUG ’97 — Soliciting bids for host site. TUGboat, 16(4):427, December 1995. ISSN 0896-3207.

Anonymous:TB16-4-428

[1881] Anonymous. Calendar. TUGboat, 16(4):428, December 1995. ISSN 0896-

3207. URL https://tug.org/calendar.html.

Anonymous:TB16-4-429

[1882] Anonymous. Call for papers: TUG ’96. TUGboat, 16(4):429–431, December

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-4/
tug96.pdf.

Goossens:TB16-4-432

[1883] Michel Goossens. EuroTEX ’95, Papendal, The Netherlands, 4–8 September 1995. TUGboat, 16(4):432–439, December 1995. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb16-4/euro95.pdf.

Anonymous:TB16-4-440

[1884] Anonymous. Future issues. TUGboat, 16(4):440, December 1995. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb16-4/tb49prod.pdf.

Burbank:TB16-4-440

[1885] Mimi Burbank and Michel Goossens. Production notes. TUGboat, 16(4):440, December 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb16-4/tb49prod.pdf.

Anonymous:TB16-4-441

[1886] Anonymous. 1996 TEX Users Group election. TUGboat, 16(4):441, December

1995. ISSN 0896-3207. URL https://tug.org/TUGboat/election.

Anonymous:TB16-4-441-2

[1887] Anonymous. 1996 TUG election — nomination form. TUGboat, 16(4):441, December 1995. ISSN 0896-3207. URL https://tug.org/TUGboat/

election.

Anonymous:TB16-4-442

[1888] Anonymous. Institutional members. TUGboat, 16(4):442, December 1995. ISSN 0896-3207.

Anonymous:TB16-4-443

[1889] Anonymous. TEX consulting and production services. TUGboat, 16(4): 443, December 1995. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB16-4-443-2

[1890] Anonymous. Index of advertisers. TUGboat, 16(4):443, December 1995. ISSN

0896-3207. URL https://tug.org/consultants.html.

Anonymous:1995:FCc

[1891] Anonymous. Front cover. TUGboat, 16(4):c1, December 1995. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-4/
cover1.pdf.

Anonymous:1995:IFCc

[1892] Anonymous. Inside front cover. TUGboat, 16(4):c2, December 1995. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb16-4/
cover2.pdf.

Anonymous:1996:TPa

[1893] Anonymous. Title page. TUGboat, 17(1):1, March 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
titlepage.pdf.

Anonymous:1996:EIa

[1894] Anonymous. Editorial information. TUGboat, 17(1):2, March 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
tb50edit.pdf.

Anonymous:TB17-1-3

[1895] Anonymous. Addresses. TUGboat, 17(1):3–4, March 1996. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb17-1/tug97.pdf.

Anonymous:TB17-1-4

[1896] Anonymous. Soliciting Bids for TUG ’97. TUGboat, 17(1):4, March

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
tug97.pdf.

Goossens:TB17-1-5

[1897] Michel Goossens. From the President. TUGboat, 17(1):5–6, March

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/Pres/
tb50pres.html.

Beeton:TB17-1-6

[1898] Barbara Beeton. Editorial comments. TUGboat, 17(1):6–7, March 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-1/tb50beet.
pdf.

Anonymous:TB17-1-7

[1899] Anonymous. TUG ’95: Questions and answers with Prof. Donald E. Knuth. TUGboat, 17(1):7–22, March 1996. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb17-1/tb50knut.pdf.

Bennett:TB17-1-22

[1900] Frank G. Bennett, Jr. Camel: kicking over the bibliographic traces in BibTEX. TUGboat, 17(1):22–28, March 1996. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb17-1/tb50benn.pdf.

Machi:TB17-1-28

[1901] Filip Machi, Jerrold E. Marsden, and Wendy G. McKay. Corrigendum: Introduction to FasTEX: a system of keyboard shortcuts for the fast keying of TEX (TUGboat 1, no. 6(4), pp. 358–363). TUGboat, 17(1):28, March 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-1/tb50mars.
pdf.

Knuth:TB17-1-29

[1902] Donald E. Knuth. Important message regarding CM fonts. TUGboat, 17(1): 29, March 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-1/tb50cmup.pdf.

Zubrinic:TB17-1-29

[1903] Darko Žubrinić. Croatian fonts. TUGboat, 17(1):29–33, March 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-1/tb50zubr.
pdf.

Vesely:TB17-1-34

[1904] J. Veselý. Two new books on TEX in the Czech Republic. TUGboat, 17(1): 34, March 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-1/tb50revc.pdf.

Price:TB17-1-35

[1905] Lynne A. Price. Ronald C. Turner, Timothy A. Douglass, and Audrey J. Turner, README.1ST: SGML for Writers and Editors. TUGboat, 17(1):35–37, March

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
tb50revs.pdf.

Porrat:TB17-1-37

[1906] Rama Porrat. There’s still something missing. … TUGboat, 17(1):37, March

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
tb50porr.pdf.

Burbank:TB17-1-37

[1907] Mimi Burbank and Michel Goossens. Electronic news from the family. TUGboat, 17(1):37–42, March 1996. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb17-1/tb50lugs.pdf.

Reckdahl:TB17-1-43

[1908] Keith Reckdahl. Using EPS graphics in L^{A}TEX 2ε documents, Part 1: the graphics and graphicx packages. TUGboat, 17(1):43–53, March 1996. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb17-1/tb50reck.pdf.

Hosek:TB17-1-53

[1909] Don Hosek. That ol’ devil \expandafter. TUGboat, 17(1):53–54, March

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
tb50hose.pdf.

Hagen:TB17-1-54

[1910] J. Hagen and A. F. Otten. PPCHTEX: typesetting chemical formulas in TEX. TUGboat, 17(1):54–66, March 1996. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb17-1/tb50hage.pdf.

Carlisle:TB17-1-67

[1911] David Carlisle. A L^{A}TEX tour, Part 1: The basic distribution. TUGboat, 17(1):67–73, March 1996. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb17-1/tb50carl.pdf.

Anonymous:TB17-1-73

[1912] Anonymous. Les Cahiers GUTenberg, No. 20. TUGboat, 17(1):73–74, March

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
tb50cahi.pdf.

Anonymous:TB17-1-74

[1913] Anonymous. Calendar. TUGboat, 17(1):74–75, March 1996. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB17-1-76

[1914] Anonymous. TUG ’96 Preliminary schedule. TUGboat, 17(1):76–77, March

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
tug96.pdf.

Anonymous:TB17-1-78

[1915] Anonymous. Future issues. TUGboat, 17(1):78, March 1996. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb17-1/tb50prod.pdf.

Burbank:TB17-1-78

[1916] Mimi Burbank. Production notes. TUGboat, 17(1):78, March 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-1/tb50prod.
pdf.

Anonymous:TB17-1-79

[1917] Anonymous. TUG Bylaws. TUGboat, 17(1):79–83, March 1996. ISSN 0896-

3207. URL https://tug.org/TUGboat/bylaws/.

Anonymous:TB17-1-84

[1918] Anonymous. 1996 TUG election cancelled. TUGboat, 17(1):84, March

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
board.pdf.

Anonymous:TB17-1-84-2

[1919] Anonymous. TUG Board. TUGboat, 17(1):84, March 1996. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb17-1/board.pdf.

Anonymous:TB17-1-85

[1920] Anonymous. Institutional members. TUGboat, 17(1):85, March 1996. ISSN 0896-3207.

Anonymous:TB17-1-86

[1921] Anonymous. TUG membership application. TUGboat, 17(1):86, March 1996.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB17-1-87

[1922] Anonymous. TEX consulting and production services. TUGboat, 17 (1):87, March 1996. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB17-1-87-2

[1923] Anonymous. Index of advertisers. TUGboat, 17(1):87, March 1996. ISSN

0896-3207. URL https://tug.org/consultants.html.

Anonymous:1996:FCa

[1924] Anonymous. Front cover. TUGboat, 17(1):c1, March 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
cover1.pdf.

Anonymous:1996:IFCa

[1925] Anonymous. Inside front cover. TUGboat, 17(1):c2, March 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-1/
cover2.pdf.

Burbank:TB17-2-90

[1926] Mimi Burbank. Production notes. TUGboat, 17(2):90, June 1996. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb17-2/tb51prod.pdf.

Goossens:TB17-2-91

[1927] Michel Goossens. Opening words by the President. TUGboat, 17(2):91,

June 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/Pres/
tb51pres.html.

Piska:TB17-2-92

[1928] Karel Píška. Cyrillic alphabets. TUGboat, 17(2):92–98, June 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-2/tb51pisk.
pdf.

Knappen:TB17-2-99

[1929] Jörg Knappen. The dc fonts 1.3: Move towards stability and completeness. TUGboat, 17(2):99–101, June 1996. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb17-2/tb51knap.pdf.

Rei:TB17-2-102

[1930] Fukui Rei. TIPA: A system for processing phonetic symbols in L^{A}TEX. TUGboat, 17(2):102–114, June 1996. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb17-2/tb51rei.pdf.

Berdnikov:TB17-2-115

[1931] A. S. Berdnikov. Computer Modern Typefaces as Multiple Master Fonts. TUGboat, 17(2):115–119, June 1996. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb17-2/tb51mff.pdf.

Berdnikov:TB17-2-120

[1932] A. S. Berdnikov. VFComb 1.3 — the program which simplifies virtual font management. TUGboat, 17(2):120–125, June 1996. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb17-2/tb51vfc.pdf.

Haralambous:TB17-2-126

[1933] Yannis Haralambous. ΩTimes and ΩHelvetica fonts under development: Step One. TUGboat, 17(2):126–146, June 1996. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb17-2/tb51omeg.pdf.

Kinch:TB17-2-147

[1934] Richard J. Kinch. Extending TEX for Unicode. TUGboat, 17(2):147–160, June

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-2/
tb51kin.pdf.

Znamenskaya:TB17-2-161

[1935] L. N. Znamenskaya and S. V. Znamenskii. Russian encoding plurality problem and a new Cyrillic font set. TUGboat, 17(2):161–165, June 1996. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb17-2/tb51mens.pdf.

Ovchenkov:TB17-2-166

[1936] Peter A. Ovchenkov. Cyrillic TEX files: interplatform portability. TUGboat, 17(2):166–171, June 1996. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb17-2/tb51ovch.pdf.

Vinogradov:TB17-2-172

[1937] Michael M. Vinogradov. A user-friendly multi-function TEX interface based on Multi-Edit. TUGboat, 17(2):172–174, June 1996. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb17-2/tb51vino.pdf.

Lapko:TB17-2-174

[1938] Olga G. Lapko. Full Cyrillic: How many languages? TUGboat, 17(2):174– 180, June 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-2/tb51olga.pdf.

Plaice:TB17-2-181

[1939] John Plaice and Yannis Haralambous. The latest developments in Ω. TUGboat, 17(2):181–183, June 1996. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb17-2/tb51plaice.pdf.

Langmyhr:TB17-2-184

[1940] Dag Langmyhr. StarTEX — a TEX for beginners. TUGboat, 17(2):184–190,

June 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-
2/tb51dag.pdf.

Feruglio:TB17-2-191

[1941] Gabriel Valiente Feruglio. Do journals honor L^{A}TEX submissions? TUGboat, 17(2):191–199, June 1996. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb17-2/tb51vali.pdf.

Znamenskii:TB17-2-200

[1942] Sergei V. Znamenskii and Denis E. Leinartas. A new approach to the TEXrelated programs: A user-friendly interface. TUGboat, 17(2):200–203, June

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-2/
tb51znam.pdf.

Vsesvetsky:TB17-2-204

[1943] Ivan G. Vsesvetsky. The strait gate to TEX. TUGboat, 17(2):204–205, June

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-2/
tb51ivan.pdf.

Siebenmann:TB17-2-206

[1944] Laurent Siebenmann. DVI-based electronic publication. TUGboat, 17(2):206– 214, June 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-2/tb51sieb.pdf.

Laan:TB17-2-215

[1945] Kees van der Laan. BLUe’s format — the off-off alternative. TUGboat, 17 (2):215–221, June 1996. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb17-2/tb51blue.pdf.

Laan:TB17-2-222

[1946] Kees van der Laan. Turtle graphics and TEX — a child can do it. TUGboat, 17(2):222–228, June 1996. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb17-2/tb51turt.pdf.

Berdnikov:TB17-2-229

[1947] A. S. Berdnikov, O. A. Grineva, and S. B. Turtia. Some useful macros which extend the L^{A}TEX picture environment. TUGboat, 17(2):229–232, June

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-2/
tb51grf.pdf.

Anonymous:TB17-2-233

[1948] Anonymous. Calendar. TUGboat, 17(2):233, June 1996. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB17-2-234

[1949] Anonymous. Institutional members. TUGboat, 17(2):234, June 1996. ISSN 0896-3207.

Anonymous:TB17-2-235

[1950] Anonymous. TEX consulting and production services. TUGboat, 17(2):235,

June 1996. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:1996:TPb

[1951] Anonymous. Title page. TUGboat, 17(2):237, June 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-3/
titlepage.pdf.

Anonymous:1996:FCb

[1952] Anonymous. Front cover. TUGboat, 17(2):c1, June 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-3/
cover1.pdf.

Anonymous:1996:IFCb

[1953] Anonymous. Inside front cover. TUGboat, 17(2):c2, June 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-3/
cover2.pdf.

Anonymous:1996:TPd

[1954] Anonymous. Title page. TUGboat, 17(3):237, September 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
titlepage.pdf.

Anonymous:1996:EIc

[1955] Anonymous. Editorial information. TUGboat, 17(3):238, September 1996. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-4/tb53edit.pdf.

Anonymous:TB17-3-239

[1956] Anonymous. Addresses. TUGboat, 17(3):239, September 1996. ISSN 0896-3207.

Anonymous:TB17-3-240

[1957] Anonymous. 1997 TUG Election. TUGboat, 17(3):240, September 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-3/tb52elec.
pdf.

Goossens:TB17-3-241

[1958] Michel Goossens. From the President. TUGboat, 17(3):241–243, Septem-

ber 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/Pres/
tb52pres.html.

Beeton:TB17-3-243

[1959] Barbara Beeton. Editorial comments. TUGboat, 17(3):243–244, September

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-3/
tb52beet.pdf.

Sojka:TB17-3-244

[1960] Petr Sojka, Hàn Thê´ Thành, and JiříZlatuška. The joy of TEX2PDF — Acrobatics with an alternative to DVI format. TUGboat, 17(3):244–251, September 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-3/tb52sojk.pdf.

Lesenko:TB17-3-252

[1961] Sergey Lesenko. The DVIPDF Program. TUGboat, 17(3):252–254, September

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-3/
tb52lese.pdf.

Fine:TB17-3-255

[1962] Jonathan Fine. Editing .dvi files, or Visual TEX. TUGboat, 17(3):255–259, September 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-3/tb52finv.pdf.

Makhovaya:TB17-3-259

[1963] Irina A. Makhovaya. TEX in Russia: ab ovo, or, About the TEXnical evolution in Russia. TUGboat, 17(3):259–264, September 1996. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb17-3/tb52makh.pdf.

Rahtz:TB17-3-264

[1964] Sebastian Rahtz. L^{A}TEX, dvips, EPS and the Web …. TUGboat, 17(3):264–268, September 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-3/tb52rahtz.pdf.

Laan:TB17-3-269

[1965] Kees van der Laan. Graphics and TEX — a reappraisal of / MetaPost/PostScript. TUGboat, 17(3):269–279, September 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-3/tb52meta.
pdf.

Thompson:TB17-3-280

[1966] David Thompson. Stanley Sawyer and Steven Krantz, A TEX Primer for Scientists. TUGboat, 17(3):280, September 1996. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb17-3/tb52revt.pdf.

Bzyl:TB17-3-281

[1967] W�odek Bzyl. David Salomon, The Advanced TEXbook. TUGboat, 17(3): 281–282, September 1996. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb17-3/tb52revb.pdf.

Fairbairns:TB17-3-282

[1968] Robin Fairbairns. The new (L^{A}TEX 2ε) TUGboat macros. TUGboat, 17(3): 282–288, September 1996. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb17-3/tb52guid.pdf.

Reckdahl:TB17-3-288

[1969] Keith Reckdahl. Using EPS graphics in L^{A}TEX 2ε documents, Part 2: Floating figures, boxed figures, captions, and math in figures. TUGboat, 17(3):288–310, September 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-3/tb52reck.pdf.

Saludes:TB17-3-310

[1970] Jordi Saludes. Fast and secure multiple-option tests. TUGboat, 17(3):310–319, September 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-3/tb52salu.pdf.

Schmitt:TB17-3-319

[1971] Peter Schmitt. A puzzling TEX macro. TUGboat, 17(3):319–320, September

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-3/
tb52schm.pdf.

Carlisle:TB17-3-321

[1972] David Carlisle. A L^{A}TEX tour, Part 2: The tools and graphics distributions. TUGboat, 17(3):321–326, September 1996. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb17-3/tb52carl.pdf.

Anonymous:TB17-3-326

[1973] Anonymous. Calendar. TUGboat, 17(3):326–327, September 1996. ISSN

0896-3207. URL https://tug.org/calendar.html.

Anonymous:TB17-3-327

[1974] Anonymous. Future issues. TUGboat, 17(3):327, September 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-3/tb52prod.
pdf.

Burbank:TB17-3-327

[1975] Mimi Burbank. Production notes. TUGboat, 17(3):327, September 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-3/tb52prod.
pdf.

Anonymous:TB17-3-328

[1976] Anonymous. TUG’97 Call for Papers. TUGboat, 17(3):328, September

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-3/
call97.pdf.

Anonymous:TB17-3-329

[1977] Anonymous. Institutional members. TUGboat, 17(3):329, September 1996. ISSN 0896-3207.

Anonymous:TB17-3-330

[1978] Anonymous. TUG membership application. TUGboat, 17(3):330, September

1996. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB17-3-331

[1979] Anonymous. TEX consulting and production services. TUGboat, 17(3): 331, September 1996. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB17-3-331-2

[1980] Anonymous. Index of advertisers. TUGboat, 17(3):331, September 1996.

ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:1996:FCd

[1981] Anonymous. Front cover. TUGboat, 17(3):c1, September 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
cover1.pdf.

Anonymous:1996:IFCd

[1982] Anonymous. Inside front cover. TUGboat, 17(3):c2, September 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
cover2.pdf.

Anonymous:1996:TPf

[1983] Anonymous. Title page. TUGboat, 17(4):333, December 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
titlepage.pdf.

Anonymous:1996:EIe

[1984] Anonymous. Editorial information. TUGboat, 17(4):334, December 1996. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-4/tb53edit.pdf.

Anonymous:TB17-4-335

[1985] Anonymous. Addresses. TUGboat, 17(4):335, December 1996. ISSN 0896-3207.

Anonymous:TB17-4-336

[1986] Anonymous. TUG’97 Update. TUGboat, 17(4):336, December 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-4/tug97update.
pdf.

Goossens:TB17-4-337

[1987] Michel Goossens. The state of TUG. TUGboat, 17(4):337–340, Decem-

ber 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/Pres/
tb53pres.html.

Beeton:TB17-4-340

[1988] Barbara Beeton. Editorial comments. TUGboat, 17(4):340–342, December

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
tb53beet.pdf.

Anonymous:TB17-4-342

[1989] Anonymous. Amsterdam, 13 March 1996 — Knuth meets NTG members. TUGboat, 17(4):342–355, December 1996. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb17-4/tb53knun.pdf.

Anonymous:TB17-4-355

[1990] Anonymous. CSTUG, Charles University, Prague, March 1996 — Questions and answers with Prof. Donald E. Knuth. TUGboat, 17(4):355–367, December

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
tb53knuc.pdf.

Taylor:TB17-4-367

[1991] Philip Taylor. Computer typesetting or electronic publishing? New trends in scientific publication. TUGboat, 17(4):367–381, December 1996. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb17-4/tb53tayl.pdf.

Strelkov:TB17-4-382

[1992] S. A. Strelkov and G. R. Epshtein. Making indices for VINITI’s “Mathematics” abstract journal. TUGboat, 17(4):382–385, December 1996. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb17-4/tb53stre.pdf.

Grinchuk:TB17-4-385

[1993] Mikhail Ivanovich Grinchuk. TEX and Russian traditions of typesetting. TUGboat, 17(4):385–388, December 1996. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb17-4/tb53grin.pdf.

Thiele:TB17-4-388

[1994] Christina Thiele. TEX and the Humanities. TUGboat, 17(4):388–393, Decem-

ber 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-
4/tb53thie.pdf.

Fine:TB17-4-393

[1995] Jonathan Fine. The Printing Revolution in Early Modern Europe, by Elizabeth L. Eisenstein. TUGboat, 17(4):393–396, December 1996. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb17-4/tb53fine.pdf.

Gratzer:TB17-4-396

[1996] George Grätzer. Suggestions wanted for new edition of Math into L^{A}TEX. TUGboat, 17(4):396, December 1996. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb17-4/tb53query.pdf.

Kuypers:TB17-4-396

[1997] Dirk Kuypers. PPCHTEX molecule templates. TUGboat, 17(4):396, December

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
tb53query.pdf.

Anonymous:TB17-4-397

[1998] Anonymous. Les Cahiers GUTenberg, Nos. 21–24. TUGboat, 17(4):397–403, December 1996. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb17-4/tb53cahi.pdf.

Anonymous:TB17-4-403

[1999] Anonymous. Calendar. TUGboat, 17(4):403–405, December 1996. ISSN

0896-3207. URL https://tug.org/calendar.html.

Anonymous:TB17-4-405

[2000] Anonymous. Future issues. TUGboat, 17(4):405, December 1996. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb17-4/tb53treas.pdf.

Burbank:TB17-4-405

[2001] Mimi Burbank. Production notes. TUGboat, 17(4):405, December 1996. ISSN

0896-3207. URL https://tug.org/TUGboat/tb17-4/tb53prod.
pdf.

Burbank:TB17-4-405-2

[2002] Mimi Burbank. Facts and Figures. TUGboat, 17(4):405–408, December

1996. ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
tb53treas.pdf.

Anonymous:TB17-4-409

[2003] Anonymous. Institutional members. TUGboat, 17(4):409, December 1996. ISSN 0896-3207.

Anonymous:TB17-4-410

[2004] Anonymous. TUG membership application. TUGboat, 17(4):410, December

1996. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB17-4-411

[2005] Anonymous. TEX consulting and production services. TUGboat, 17(4): 411, December 1996. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB17-4-411-2

[2006] Anonymous. Index of advertisers. TUGboat, 17(4):411, December 1996. ISSN

0896-3207. URL https://tug.org/consultants.html.

Anonymous:1996:FCf

[2007] Anonymous. Front cover. TUGboat, 17(4):c1, December 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
cover1.pdf.

Anonymous:1996:IFCf

[2008] Anonymous. Inside front cover. TUGboat, 17(4):c2, December 1996. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb17-4/
cover2.pdf.

Anonymous:1997:TPa

[2009] Anonymous. Title page. TUGboat, 18(1):1, March 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-1/
titlepage.pdf.

Anonymous:1997:EIa

[2010] Anonymous. Editorial information. TUGboat, 18(1):2, March 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-1/
tb54edit.pdf.

Anonymous:TB18-1-3

[2011] Anonymous. Addresses. TUGboat, 18(1):3, March 1997. ISSN 0896-3207.

Anonymous:TB18-1-4

[2012] Anonymous. Notice regarding 1997 TEX Users Group election. TUGboat, 18(1):4, March 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-1/tb54beet.pdf.

Goossens:TB18-1-5

[2013] Michel Goossens. From the President. TUGboat, 18(1):5, March 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/Pres/tb54pres.
html.

Beeton:TB18-1-5

[2014] Barbara Beeton. Editorial comments. TUGboat, 18(1):5–6, March 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-1/tb54beet.
pdf.

Anonymous:TB18-1-6

[2015] Anonymous. Erratum: Amsterdam, 13 March 1996 — Knuth meets NTG members, TUGboat 1, no. 7(4), pp. 342–355. TUGboat, 18(1):6, March 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-1/tb54beet.
pdf.

Taylor:TB18-1-6

[2016] Philip Taylor. NTS & ε-TEX: a status report. TUGboat, 18(1):6–12, March

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-1/
tb54tayl.pdf.

Vieth:TB18-1-12

[2017] Ulrik Vieth. A GNU Emacs editing mode for and sources. TUGboat, 18(1):12–16, March 1997. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb18-1/tb54viet.pdf.

Haralambous:TB18-1-17

[2018] Yannis Haralambous. The traditional Arabic typecase, Unicode, TEX and . TUGboat, 18(1):17–29, March 1997. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb18-1/tb54arab.pdf.

Weenen:TB18-1-30

[2019] Andrea de Leeuw van Weenen. A Medieval Icelandic manuscript: The making of a diplomatic edition. TUGboat, 18(1):30–36, March 1997. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb18-1/tb54leew.pdf.

Sofka:TB18-1-37

[2020] Michael D. Sofka. Writing with TEX, and TEX& L^{A}TEX: Drawing & Literate Programming, by Eitan M. Gurari. TUGboat, 18(1):37–38, March 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-1/tb54revg.
pdf.

Beccari:TB18-1-39

[2021] Claudio Beccari. Typesetting mathematics for science and technology according to ISO 31/XI. TUGboat, 18(1):39–48, March 1997. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb18-1/tb54becc.pdf.

Carlisle:TB18-1-48

[2022] David Carlisle. A L^{A}TEX Tour, part 3: mfnfss, psnfss and babel. TUGboat, 18(1):48–55, March 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-1/tb54carl.pdf.

Anonymous:TB18-1-56

[2023] Anonymous. Calendar. TUGboat, 18(1):56–67, March 1997. ISSN 0896-3207.

Anonymous:TB18-1-57

[2024] Anonymous. Future issues. TUGboat, 18(1):57, March 1997. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb18-1/tb54prod.pdf.

Anonymous:TB18-1-57-2

[2025] Anonymous. Index of advertisers. TUGboat, 18(1):57, March 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-1/tb54prod.
pdf.

Burbank:TB18-1-57

[2026] Mimi Burbank. Production notes. TUGboat, 18(1):57, March 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-1/tb54prod.
pdf.

Anonymous:TB18-1-58

[2027] Anonymous. TUG’97 Update. TUGboat, 18(1):58, March 1997. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb18-1/tug97-update.
pdf.

Anonymous:TB18-1-59

[2028] Anonymous. New members of the TUG Board. TUGboat, 18(1):59–60, March

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-1/
tb54boar.pdf.

Anonymous:TB18-1-61

[2029] Anonymous. Institutional members. TUGboat, 18(1):61, March 1997. ISSN 0896-3207.

Anonymous:TB18-1-62

[2030] Anonymous. TUG membership application. TUGboat, 18(1):62, March 1997.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB18-1-63

[2031] Anonymous. TEX consulting and production services. TUGboat, 18 (1):63, March 1997. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:1997:FCa

[2032] Anonymous. Front cover. TUGboat, 18(1):c1, March 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-1/
cover1.pdf.

Anonymous:1997:IFCa

[2033] Anonymous. Inside front cover. TUGboat, 18(1):c2, March 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-1/
cover2.pdf.

Anonymous:1997:TPb

[2034] Anonymous. Title page. TUGboat, 18(2):65, June 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
titlepage.pdf.

Anonymous:1997:EIb

[2035] Anonymous. Editorial information. TUGboat, 18(2):66, June 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
tb55edit.pdf.

Anonymous:TB18-2-67

[2036] Anonymous. Addresses. TUGboat, 18(2):67, June 1997. ISSN 0896-3207.

Kopriva:TB18-2-68

[2037] Phil Kopriva. Letter from a member. TUGboat, 18(2):68, June 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-2/tb55kopr.
pdf.

Goossens:TB18-2-69

[2038] Michel Goossens. From the President. TUGboat, 18(2):69–71, June

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/Pres/
tb55pres.html.

Beeton:TB18-2-72

[2039] Barbara Beeton. Editorial comments. TUGboat, 18(2):72–73, June 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-2/tb55beet.
pdf.

Flynn:TB18-2-74

[2040] Peter Flynn. TTN returns: Final TTN editorial. TUGboat, 18(2):74, June

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
tb55flyn.pdf.

Gibbons:TB18-2-75

[2041] Jeremy Gibbons. ‘Hey — it works!’. TUGboat, 18(2):75–78, June

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
tb55works.pdf.

Thiele:TB18-2-78

[2042] Christina Thiele. TEX and the Humanities. TUGboat, 18(2):78, June

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
tb55huma.pdf.

Chow:TB18-2-79

[2043] Sidney Chow. Features for a WYSIWYG L^{A}TEX editor. TUGboat, 18(2):79, June

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
tb55query.pdf.

Grant:TB18-2-79

[2044] Michael C. Grant. Volunteers needed: PSfrag hackers. TUGboat, 18(2):79,

June 1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-
2/tb55query.pdf.

Anonymous:TB18-2-80

[2045] Anonymous. New publications. TUGboat, 18(2):80, June 1997. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb18-2/tb55newpubs.
pdf.

Rahtz:TB18-2-81

[2046] Sebastian Rahtz and Michel Goossens. The TEXLive Guide, version 2. TUGboat, 18(2):81–112, June 1997. ISSN 0896-3207. URL ftp://tug.org/

historic/systems/texlive/1997/tb55tlguide2.pdf.

Hoenig:TB18-2-113

[2047] Alan Hoenig. Virtual fonts, virtuous fonts. TUGboat, 18(2):113–121, June

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
hoenig.html.

Anonymous:TB18-2-121

[2048] Anonymous. Les Cahiers GUTenberg, contents of Issue 25. TUGboat, 18 (2):121–122, June 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-2/tb55cahi.pdf.

Anonymous:TB18-2-122

[2049] Anonymous. Die TEXnische Komödie 7, 1995, Heft 1–4; 8, 1996, Heft 1–4. TUGboat, 18(2):122–126, June 1997. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb18-2/tb55komo.pdf.

L3PT:TB18-2-127

[2050] L^{A}TEX 3 Project Team. Modifying L^{A}TEX. TUGboat, 18(2):127–130, June

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
tb55guid.pdf.

Anonymous:TB18-2-130-2

[2051] Anonymous. Future issues. TUGboat, 18(2):130, June 1997. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb18-2/tb55prod.pdf.

Burbank:TB18-2-130

[2052] Mimi Burbank. Production notes. TUGboat, 18(2):130, June 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-2/tb55prod.
pdf.

Anonymous:TB18-2-130

[2053] Anonymous. News from the L^{A}TEX 3 Project Team. TUGboat, 18(2):130–137,

June 1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-
2/tb55prod.pdf.

Anonymous:TB18-2-138

[2054] Anonymous. Calendar. TUGboat, 18(2):138, June 1997. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB18-2-138-2

[2055] Anonymous. Index of advertisers. TUGboat, 18(2):138, June 1997. ISSN

0896-3207. URL https://tug.org/calendar.html.

Anonymous:TB18-2-139

[2056] Anonymous. TUG’97 preliminary program. TUGboat, 18(2):139–140, June

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
tug97prog.pdf.

Anonymous:TB18-2-141

[2057] Anonymous. Institutional members. TUGboat, 18(2):141, June 1997. ISSN 0896-3207.

Anonymous:TB18-2-142

[2058] Anonymous. TUG membership application. TUGboat, 18(2):142, June 1997.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB18-2-143

[2059] Anonymous. TEX consulting and production services. TUGboat, 18(2):143,

June 1997. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:1997:FCb

[2060] Anonymous. Front cover. TUGboat, 18(2):c1, June 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
cover1.pdf.

Anonymous:1997:IFCb

[2061] Anonymous. Inside front cover. TUGboat, 18(2):c2, June 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-2/
cover2.pdf.

Anonymous:TB18-3-146

[2062] Anonymous. Call for Papers. TUGboat, 18(3):146, September 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-3/tug98.pdf.

Jett:TB18-3-147

[2063] Mimi Jett. Opening words by the President. TUGboat, 18(3):147–148, Septem-

ber 1997. ISSN 0896-3207. URL https://tug.org/TUGboat/Pres/
tb56pres.html.

Burbank:TB18-3-148

[2064] Mimi Burbank. Production notes. TUGboat, 18(3):148, September 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-3/tb56prod.
pdf.

Anonymous:TB18-3-149

[2065] Anonymous. TUG’97 program. TUGboat, 18(3):149, September 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-3/tug97prog.
pdf.

Rose:TB18-3-151

[2066] Kristoffer Høgsbro Rose. Very high level 2-dimensional graphics with TEX and Xy-pic. TUGboat, 18(3):151–158, September 1997. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb18-3/tb56rose.pdf.

Moore:TB18-3-159

[2067] Ross Moore. High quality labels on included graphics, using Xy-pic. TUGboat, 18(3):159–163, September 1997. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb18-3/tb56moore.pdf.

Tannert:1997:CPT

[2068] Sebastian Tannert. CIRC — a package to typeset block schematics. TUGboat, 18(3):164–165, September 1997. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb18-3/tb56circ.pdf.

Lesenko:TB18-3-166

[2069] Sergey Lesenko. DVIPDF and Graphics. TUGboat, 18(3):166–169, September

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-3/
tb56lese.pdf.

Hamlin:TB18-3-170

[2070] Christopher B. Hamlin. From SGML to HTML with help from TEX. TUGboat, 18(3):170–174, September 1997. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb18-3/tb56haml.pdf.

Lovell:TB18-3-175

[2071] Douglas Lovell. Custom legal documents for the IBM AutoLoan Exchange. TUGboat, 18(3):175–181, September 1997. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb18-3/tb56love.pdf.

Downes:TB18-3-182

[2072] Michael Downes. Breaking equations. TUGboat, 18(3):182–194, September

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-3/
tb56down.pdf.

Mittelbach:TB18-3-195

[2073] Frank Mittelbach and Chris Rowley. The L^{A}TEX3 Project. TUGboat, 18(3): 195–198, September 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-3/l3project.pdf.

Mittelbach:TB18-3-199

[2074] Frank Mittelbach and Chris Rowley. Language information in structured documents: a model for mark-up and rendering. TUGboat, 18(3):199–205, Septem-

ber 1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-
3/tb56lang.pdf.

Lemberg:TB18-3-206

[2075] Werner Lemberg. New font tools for TEX. TUGboat, 18(3):206–213, Septem-

ber 1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-
3/fonttool600.pdf.

Lemberg:TB18-3-214

[2076] Werner Lemberg. The CJK package: Multilingual support beyond Babel. TUGboat, 18(3):214–224, September 1997. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb18-3/cjkintro600.pdf.

Anonymous:TB18-3-225

[2077] Anonymous. Calendar. TUGboat, 18(3):225, September 1997. ISSN 0896-

3207. URL https://tug.org/calendar.html.

Anonymous:TB18-3-226

[2078] Anonymous. TUG’97 list of attendees. TUGboat, 18(3):226–228, September 1997. ISSN 0896-3207.

Anonymous:TB18-3-229

[2079] Anonymous. Institutional members. TUGboat, 18(3):229, September 1997. ISSN 0896-3207.

Anonymous:TB18-3-230

[2080] Anonymous. TUG membership form. TUGboat, 18(3):230, September 1997.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB18-3-231

[2081] Anonymous. TEX consulting and production services. TUGboat, 18(3): 231, September 1997. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB18-3-232

[2082] Anonymous. Y&Y Inc. TUGboat, 18(3):232, September 1997. ISSN 0896-

3207. URL https://tug.org/twg/yy/.

Anonymous:1997:FCc

[2083] Anonymous. Front cover. TUGboat, 18(3):c1, September 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-3/
proccover.pdf.

Anonymous:TB18-3-cover3

[2084] Anonymous. Blue Sky Research. TUGboat, 18(3):Cover 3, September 1997. ISSN 0896-3207.

Anonymous:TB18-4-235

[2085] Anonymous. Addresses. TUGboat, 18(4):235–236, December 1997. ISSN 0896-3207.

Taupin:TB18-4-236

[2086] Daniel Taupin. MusiXTEX: How many fonts are acceptable? TUGboat, 18 (4):236, December 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-4/tb57query.pdf.

Jett:TB18-4-237

[2087] Mimi Jett. From the President. TUGboat, 18(4):237, December 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/Pres/tb57pres.
html.

Beeton:TB18-4-238

[2088] Barbara Beeton. Editorial comments. TUGboat, 18(4):238, December

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-4/
tb57beet.pdf.

Taylor:TB18-4-239

[2089] Phil Taylor. eTEXV2: a peek into the future. TUGboat, 18(4):239–242, Decem-

ber 1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-
4/tb57tayl.pdf.

Flynn:TB18-4-242

[2090] Peter Flynn. Typographers’ Inn. TUGboat, 18(4):242–245, December

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-4/
tb57flyn.pdf.

Rogers:TB18-4-246

[2091] David F. Rogers. The L^{A}TEX Graphics Companion, by Michel Goossens, Sebastian Rahtz and Frank Mittelbach. TUGboat, 18(4):246, December 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-4/tb57revl.
pdf.

Gibbons:TB18-4-246

[2092] Jeremy Gibbons. ‘Hey — it works!’. TUGboat, 18(4):246–248, December

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-4/
tb57work.pdf.

Thanh:TB18-4-249

[2093] Hàn Thê´ Thành and Sebastian Rahtz. The pdfTEX user manual. TUGboat, 18 (4):249–254, December 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-4/tb57than.pdf.

Finston:TB18-4-255

[2094] Laurence Finston. Spindex — Indexing with special characters. TUGboat, 18 (4):255–273, December 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-4/tb57fins.pdf.

Roegel:TB18-4-274

[2095] Denis Roegel. Creating 3D animations with . TUGboat, 18(4): 274–283, December 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-4/tb57roeg.pdf.

Horn:TB18-4-284

[2096] Berthold K. P. Horn. ‘Hinting’ of scalable outline fonts. TUGboat, 18(4): 284–286, December 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-4/tb57horn.pdf.

Willadt:TB18-4-286

[2097] Peter Willadt. Another approach to barcodes. TUGboat, 18(4):286–290, December 1997. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb18-4/tb57will.pdf.

Anonymous:TB18-4-290

[2098] Anonymous. Die TEXnische Komödie 6, 1994. TUGboat, 18(4):290–295, December 1997. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb18-4/tb57komo.pdf.

Anonymous:TB18-4-295

[2099] Anonymous. Les Cahiers GUTenberg, Contents of Issues 26 and 27. TUGboat, 18(4):295–297, December 1997. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb18-4/tb57cahi.pdf.

Aphalo:TB18-4-297

[2100] Pedro J. Aphalo. A proposal for citation commands in L^{A}TEX3. TUGboat, 18 (4):297–302, December 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-4/tb57apha.pdf.

Carlisle:TB18-4-303

[2101] David Carlisle, Chris Rowley, and Frank Mittelbach. The L^{A}TEX3 Programming Language — a proposed system for TEX macro programming. TUGboat, 18(4): 303–308, December 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-4/tb57rowl.pdf.

Mittelbach:TB18-4-309

[2102] Frank Mittelbach. A regression test suite for L^{A}TEX 2ε. TUGboat, 18(4): 309–311, December 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-4/tb57mitt.pdf.

Anonymous:TB18-4-312

[2103] Anonymous. News from the L^{A}TEX3 Project Team. TUGboat, 18(4):312, December 1997. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb18-4/ltnews08.pdf.

Anonymous:TB18-4-313

[2104] Anonymous. A Week on Electronic Documents and Typography. TUGboat, 18(4):313, December 1997. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb18-4/wept98.pdf.

Anonymous:TB18-4-314

[2105] Anonymous. TUG’98 Call for papers. TUGboat, 18(4):314, December

1997. ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-4/
tug98call.pdf.

Anonymous:TB18-4-315

[2106] Anonymous. Calendar. TUGboat, 18(4):315–316, December 1997. ISSN

0896-3207. URL https://tug.org/calendar.html.

Anonymous:TB18-4-316

[2107] Anonymous. Future issues. TUGboat, 18(4):316, December 1997. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb18-4/tb57prod.pdf.

Burbank:TB18-4-316

[2108] Mimi Burbank. Production notes. TUGboat, 18(4):316, December 1997. ISSN

0896-3207. URL https://tug.org/TUGboat/tb18-4/tb57prod.
pdf.

Anonymous:TB18-4-317

[2109] Anonymous. Institutional members. TUGboat, 18(4):317, December 1997. ISSN 0896-3207.

Anonymous:TB18-4-318

[2110] Anonymous. TUG membership application. TUGboat, 18(4):318, December

1997. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB18-4-319

[2111] Anonymous. TEX consulting and production services. TUGboat, 18(4): 319, December 1997. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB18-4-320

[2112] Anonymous. Y&Y Inc. TUGboat, 18(4):320, December 1997. ISSN 0896-

3207. URL https://tug.org/twg/yy/.

Anonymous:1997:TPc

[2113] Anonymous. Title page. TUGboat, 18(4):345, December 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-4/
titlepage.pdf.

Anonymous:1997:EIc

[2114] Anonymous. Editorial information. TUGboat, 18(4):346, December 1997. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb18-4/tb57edit.pdf.

Anonymous:1997:FCd

[2115] Anonymous. Front cover. TUGboat, 18(4):c1, December 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-4/
cover1.pdf.

Anonymous:1997:IFCc

[2116] Anonymous. Inside front cover. TUGboat, 18(4):c2, December 1997. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb18-4/
cover2.pdf.

Anonymous:TB18-4-cover3

[2117] Anonymous. Blue Sky Research. TUGboat, 18(4):Cover 3, December 1997. ISSN 0896-3207.

Anonymous:TB19-1-3

[2118] Anonymous. Addresses. TUGboat, 19(1):3, March 1998. ISSN 0896-3207.

Jett:TB19-1-4

[2119] Mimi Jett. From the President. TUGboat, 19(1):4, March 1998. ISSN 0896-

3207. URL https://tug.org/TUGboat/Pres/tb58pres.html.

Beeton:TB19-1-5

[2120] Barbara Beeton. Editorial comments. TUGboat, 19(1):5, March 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-1/tb58beet.
pdf.

Radhakrishnan:TB19-1-6

[2121] C. V. Radhakrishnan. A case for TEX in India — The Indian TEX Users Group. TUGboat, 19(1):6–9, March 1998. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb19-1/tb58radh.pdf.

Rahtz:TB19-1-9

[2122] Sebastian Rahtz. The inaugural meeting of TUG India. TUGboat, 19(1):9– 11, March 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-1/tb58raht.pdf.

Scherer:TB19-1-11

[2123] Andreas Scherer. The future of AmiWeb2c. TUGboat, 19(1):11–12, March

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-1/
tb58sche.pdf.

Beebe:TB19-1-12

[2124] Nelson H. F. Beebe. AUTHIDX: An author/editor indexing package. TUGboat, 19(1):12–18, March 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-1/tb58beeb.pdf.

Rahtz:TB19-1-19

[2125] Sebastian Rahtz and Michel Goossens. The TEX Live guide, 3rd edition. TUGboat, 19(1):19–60, March 1998. ISSN 0896-3207. URL ftp://tug.org/

historic/systems/texlive/1998/TeXLive3/doc/live.pdf.

Moore:TB19-1-61

[2126] Ross Moore. Erratum: High quality labels on included graphics, using Xy-pic, TUGboat 1, no. 8(3), pp. 151–158. TUGboat, 19(1):61, March

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-1/
tb58moore.pdf.

Horn:TB19-1-62

[2127] Berthold K. P. Horn. The European Modern fonts. TUGboat, 19(1):62–63, March 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-1/tb58horn.pdf.

Gibbons:TB19-1-63

[2128] Jeremy Gibbons. ‘Hey — it works!’. TUGboat, 19(1):63–65, March

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-1/
tb58works.pdf.

Taylor:TB19-1-65

[2129] Philip Taylor. Book design for TEX users, Part 1: Theory. TUGboat, 19(1):65– 74, March 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-1/tb58tay1.pdf.

Anonymous:TB19-1-75

[2130] Anonymous. Die TEXnische Komödie 8 (1996, Heft 1–4). TUGboat, 19(1):75– 80, March 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-1/tb58komo.pdf.

Anonymous:TB19-1-81

[2131] Anonymous. Calendar. TUGboat, 19(1):81–82, March 1998. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB19-1-82

[2132] Anonymous. Future issues. TUGboat, 19(1):82, March 1998. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb19-1/tb58prod.pdf.

Anonymous:TB19-1-82-2

[2133] Anonymous. Index of advertisers. TUGboat, 19(1):82, March 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-1/tb58prod.
pdf.

Burbank:TB19-1-82

[2134] Mimi Burbank. Production notes. TUGboat, 19(1):82, March 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-1/tb58prod.
pdf.

Schroder:TB19-1-83

[2135] Martin Schröder. TEX Merchandising — an announcement. TUGboat, 19(1): 83, March 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-1/tb58schroder.pdf.

Anonymous:TB19-1-84

[2136] Anonymous. New members of the TUG Board. TUGboat, 19(1):84–85, March

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-1/
tb58elec.pdf.

Anonymous:TB19-1-86

[2137] Anonymous. Institutional members. TUGboat, 19(1):86, March 1998. ISSN 0896-3207.

Anonymous:TB19-1-87

[2138] Anonymous. TUG membership application. TUGboat, 19(1):87, March 1998.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB19-1-88

[2139] Anonymous. TEX consulting and production services. TUGboat, 19 (1):88, March 1998. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:1998:TPa

[2140] Anonymous. Title page. TUGboat, 19(1):345, March 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-1/
titlepage.pdf.

Anonymous:1998:EIa

[2141] Anonymous. Editorial information. TUGboat, 19(1):346, March 1998. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-
1/tb58edit.pdf.

Anonymous:1998:FCa

[2142] Anonymous. Front cover. TUGboat, 19(1):c1, March 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-1/
cover1.pdf.

Anonymous:1998:IFCa

[2143] Anonymous. Inside front cover. TUGboat, 19(1):c2, March 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-1/
cover2.pdf.

Anonymous:TB19-2-91

[2144] Anonymous. Addresses. TUGboat, 19(2):91, June 1998. ISSN 0896-3207.

Anonymous:TB19-2-92

[2145] Anonymous. TUG 98 – Torun, Poland, 17–21 August 1998, Preliminary program. TUGboat, 19(2):92, June 1998. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb19-2/tug98prog.pdf.

Jett:TB19-2-93

[2146] Mimi Jett. From the President. TUGboat, 19(2):93, June 1998. ISSN 0896-

3207. URL https://tug.org/TUGboat/Pres/tb59pres.html.

Beeton:TB19-2-94

[2147] Barbara Beeton. Editorial comments. TUGboat, 19(2):94, June 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-2/tb59beet.
pdf.

Anonymous:TB19-2-95

[2148] Anonymous. April Fool’s hoax. TUGboat, 19(2):95–96, June 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-2/tb59hoax.
pdf.

Anonymous:TB19-2-97

[2149] Anonymous. CTAN CDROM series, compliments of DANTE. TUGboat, 19(2): 97–98, June 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-2/tb59dant.pdf.

Flynn:TB19-2-98

[2150] Peter Flynn. Typographers’ Inn. TUGboat, 19(2):98–100, June 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-2/tb59inn.
pdf.

Girou:TB19-2-101

[2151] Denis Girou. pst-fill — a PSTricks package for filling and tiling areas. TUGboat, 19(2):101–113, June 1998. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb19-2/tb59girou.pdf.

Doob:TB19-2-113

[2152] Michael Doob. TEX Unbound, by Alan Hoenig. TUGboat, 19(2):113–114, June

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
tb59revh.pdf.

Pandey:TB19-2-115

[2153] Anshuman Pandey. An overview of Indic fonts for TEX. TUGboat, 19(2):115– 120, June 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-2/tb59pand.pdf.

Bouche:TB19-2-120

[2154] Thierry Bouche. Diversity in math fonts. TUGboat, 19(2):120–134, June

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
tb59bouc.pdf.

Gibbons:TB19-2-135

[2155] Jeremy Gibbons. ‘Hey — it works!’. TUGboat, 19(2):135–136, June

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
tb59gibb.pdf.

LPT:TB19-2-137

[2156] L^{A}TEX Project Team. Default docstrip headers. TUGboat, 19(2):137–138, June

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
tb59ltdocstrip.pdf.

LPT:TB19-2-139

[2157] L^{A}TEX Project Team. L^{A}TEX News, Issue 9, June 1998. TUGboat, 19(2):139,

June 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-
2/tb59news.pdf.

Anonymous:TB19-2-140

[2158] Anonymous. Conference Program. TUGboat, 19(2):140–141, June 1998. ISSN 0896-3207.

Hogue:TB19-2-142

[2159] Stephanie Hogue. Final report: TEXNortheast. TUGboat, 19(2):142–143, June

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
tb59hogue.pdf.

Anonymous:TB19-2-144

[2160] Anonymous. TEXNortheast: Workshops and additional papers. TUGboat, 19(2):144–146, June 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-2/tb59abs.pdf.

Barnett:TB19-2-147

[2161] Michael P. Barnett. mathscape — Combining Mathematica and TEX. TUGboat, 19(2):147–156, June 1998. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb19-2/tb59barn.pdf.

Sutor:TB19-2-157

[2162] Robert S. Sutor and Samuel S. Dooley. TEX and L^{A}TEX on the Web via IBM techexplorer. TUGboat, 19(2):157–161, June 1998. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb19-2/tb59dool.pdf.

Hendrickson:TB19-2-162

[2163] Amy Hendrickson. Real Life L^{A}TEX: Adventures of a TEX consultant. TUGboat, 19(2):162–167, June 1998. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb19-2/tb59amy.pdf.

Hoenig:TB19-2-168

[2164] Alan J. Hoenig. Typesetting with TEX and L^{A}TEX. TUGboat, 19(2):168–175,

June 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-
2/tb59hintro.pdf.

Hoenig:TB19-2-176

[2165] Alan J. Hoenig. Alternatives to Computer Modern mathematics. TUGboat, 19(2):176–187, June 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-2/tb59hmath.pdf.

McArthur:TB19-2-188

[2166] Jeffrey McArthur. Developing database publishing systems using TEX. TUGboat, 19(2):188–194, June 1998. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb19-2/tb59mcar.pdf.

Moore:TB19-2-195

[2167] Ross Moore. Presenting mathematics and languages in Web-pages, using L^{A}TEX2HTML. TUGboat, 19(2):195–203, June 1998. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb19-2/tb59moore.pdf.

Patashnik:TB19-2-204

[2168] Oren Patashnik. BibTEX 101. TUGboat, 19(2):204–207, June 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-2/tb59patash.
pdf.

Signell:TB19-2-208

[2169] Peter Signell. One-document scientific publishing for print and Web/CD. TUGboat, 19(2):208–213, June 1998. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb19-2/tb59sign.pdf.

Sofka:TB19-2-214

[2170] Michael D. Sofka. TEX to HTML translation via tagged DVI files. TUGboat, 19(2):214–222, June 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-2/tb59sofka.pdf.

Anonymous:TB19-2-223

[2171] Anonymous. Les Cahiers GUTenberg, Contents of issue 28–29. TUGboat, 19(2):223–226, June 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-2/tb59cahi.pdf.

Anonymous:TB19-2-227

[2172] Anonymous. Calendar. TUGboat, 19(2):227, June 1998. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB19-2-228

[2173] Anonymous. Future issues. TUGboat, 19(2):228, June 1998. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb19-2/tb59prod.pdf.

Burbank:TB19-2-228

[2174] Mimi Burbank. Production notes. TUGboat, 19(2):228, June 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-2/tb59prod.
pdf.

Anonymous:TB19-2-229

[2175] Anonymous. TUG’99 Announcement. TUGboat, 19(2):229, June 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-2/tug99.pdf.

Anonymous:TB19-2-230

[2176] Anonymous. Institutional members. TUGboat, 19(2):230, June 1998. ISSN 0896-3207.

Anonymous:TB19-2-231

[2177] Anonymous. TEX consulting and production services. TUGboat, 19(2):231,

June 1998. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Hennings:TB19-2-231

[2178] Wilfred Hennings. Volunteers needed for LaTeX2rtf coordination and development. TUGboat, 19(2):231, June 1998. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb19-2/tb59henn.pdf.

Anonymous:TB19-2-232

[2179] Anonymous. Y&Y Inc. TUGboat, 19(2):232, June 1998. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb19-2/yandycolor.pdf.

Anonymous:1998:TPb

[2180] Anonymous. Title page. TUGboat, 19(2):345, June 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
titlepage.pdf.

Anonymous:1998:EIb

[2181] Anonymous. Editorial information. TUGboat, 19(2):346, June 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
tb59edit.pdf.

Anonymous:1998:BSR

[2182] Anonymous. Blue Sky Research. TUGboat, 19(2):c3, June 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
bluesky98.pdf.

Anonymous:1998:FCb

[2183] Anonymous. Front cover. TUGboat, 19(2):c1, June 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
cover1.pdf.

Anonymous:1998:IFCb

[2184] Anonymous. Inside front cover. TUGboat, 19(2):c2, June 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-2/
cover2.pdf.

Anonymous:TB19-2-Cover-3

[2185] Anonymous. Blue Sky Research. TUGboat, 19(2):Cover 3, June 1998. ISSN 0896-3207.

Beeton:TB19-3-234

[2186] Barbara Beeton. TUG Election Notice. TUGboat, 19(3):234, September

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-3/
tb60elec.pdf.

Beeton:TB19-3-235

[2187] Barbara Beeton. Editorial Comments — A TUG ’98 trip report. TUGboat, 19 (3):235–236, September 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-3/tb60beet.pdf.

Anonymous:TB19-3-237

[2188] Anonymous. TUG’98 Attendees. TUGboat, 19(3):237–238, September

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-3/
tub60pl1.pdf.

Nickalls:TB19-3-239

[2189] R. W. D. Nickalls. TEX in the Operating Theatre: An anaesthesia application. TUGboat, 19(3):239–241, September 1998. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb19-3/tb60nick.pdf.

Nowacki:TB19-3-242

[2190] Janusz M. Nowacki and Bogus�aw Jackowski. Antykwa Toruńska: An electronic replica of a Polish traditional type. TUGboat, 19(3):242–243, September

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-3/
tb60antyk.pdf.

Kinch:TB19-3-244

[2191] Richard J. Kinch. Belleek: A call for revival. TUGboat, 19(3): 244–249, September 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-3/tb60kinch.pdf.

Piska:TB19-3-250

[2192] Karel Píška. Georgian scripts. TUGboat, 19(3):250–255, September

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-3/
tb60pisk.pdf.

Hoekwater:TB19-3-256

[2193] Taco Hoekwater. Generating Type 1 fonts from sources. TUGboat, 19(3):256–266, September 1998. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb19-3/hoek1.pdf.

Jackowski:TB19-3-267

[2194] Bogus�aw Jackowski, Piotr Pianowski, and Piotr Strzelczyk. Threshing EPS files. TUGboat, 19(3):267–271, September 1998. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb19-3/tb60jack.pdf.

Jackowski:TB19-3-272

[2195] Bogus�aw Jackowski, Piotr Pianowski, and Piotr Strzelczyk. More TEX-PostScript links. TUGboat, 19(3):272–275, September 1998. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb19-3/tb60jps.pdf.

Bolek:TB19-3-276

[2196] Piotr Bolek. and patterns. TUGboat, 19(3):276–283, September

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-3/
tb60bolek.pdf.

Thanh:TB19-3-284

[2197] Hàn Thê´ Thành. Improving TEX’s Typeset Layout. TUGboat, 19(3):284–288, September 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-3/tb60than.pdf.

Taupin:TB19-3-289

[2198] Daniel Taupin. ltx2rtf: Exporting L^{A}TEX documents to Word addicts. TUGboat, 19(3):289–292, September 1998. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb19-3/tb60taup.pdf.

Bzyl:TB19-3-293

[2199] W�odek Bzyl. Adding native language support to the CWEB package and the TEX program. TUGboat, 19(3):293–297, September 1998. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb19-3/tb60bzyl.pdf.

Wolinski:TB19-3-298

[2200] Marcin Woliński. Pretprin — a L^{A}TEX 2ε package for pretty-printing texts in formal languages. TUGboat, 19(3):298–303, September 1998. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb19-3/tb60wolin.pdf.

Hagen:TB19-3-304

[2201] Hans Hagen. The Calculator Demo, Integrating TEX, , JavaScript and PDF. TUGboat, 19(3):304–310, September 1998. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb19-3/hagen1.pdf.

Hagen:TB19-3-311

[2202] Hans Hagen. Visual Debugging in TEX, Part 1: The Story. TUGboat, 19(3): 311–317, September 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-3/hagen2.pdf.

Hagen:TB19-3-317

[2203] Hans Hagen. Visual Debugging in TEX, Part 2: The Macros. TUGboat, 19 (3):317, September 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-3/hagen3.pdf.

Skoupy:TB19-3-318

[2204] Karel Skoupý. N T S: a New Typesetting System. TUGboat, 19(3):318–322, September 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-3/tb60nts.pdf.

NTGTFWG:TB19-3-323

[2205] NTG TEX future working group. TEX in 2003, Part I: Introduction and views on current work. TUGboat, 19(3):323–329, September 1998. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb19-3/tb60hoekI.pdf.

NTGTFWG:TB19-3-330

[2206] NTG TEX future working group. TEX in 2003, Part II: Proposal for a \special standard. TUGboat, 19(3):330–337, September 1998. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb19-3/tb60hoekII.
pdf.

Anonymous:TB19-3-338

[2207] Anonymous. Calendar. TUGboat, 19(3):338, September 1998. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb19-3/tb60prod.pdf.

Anonymous:TB19-3-339

[2208] Anonymous. Hug The Lion! TUGboat, 19(3):339, September 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-3/tb60lion.
pdf.

Burbank:TB19-3-339

[2209] Mimi Burbank. Production notes. TUGboat, 19(3):339, September 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-3/tb60prod.
pdf.

Anonymous:TB19-3-340

[2210] Anonymous. TUG’99 Announcement. TUGboat, 19(3):340, September

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-3/
Tug99.pdf.

Anonymous:TB19-3-341

[2211] Anonymous. Institutional members. TUGboat, 19(3):341, September 1998. ISSN 0896-3207.

Anonymous:TB19-3-342

[2212] Anonymous. TUG membership application. TUGboat, 19(3):342, September 1998. ISSN 0896-3207.

Anonymous:TB19-3-343

[2213] Anonymous. TEX consulting and production services. TUGboat, 19(3):343, September 1998. ISSN 0896-3207.

Anonymous:TB19-3-344

[2214] Anonymous. Y&Y Inc. TUGboat, 19(3):344, September 1998. ISSN 0896-3207.

Anonymous:1998:FCc

[2215] Anonymous. Front cover. TUGboat, 19(3):c1, September 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-3/
cover1.pdf.

Anonymous:TB19-3-Cover-3

[2216] Anonymous. Blue Sky Research. TUGboat, 19(3):Cover 3, September 1998. ISSN 0896-3207.

Anonymous:1998:TPc

[2217] Anonymous. Title page. TUGboat, 19(4):345, December 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-4/
titlepage.pdf.

Anonymous:1998:EIc

[2218] Anonymous. Editorial information. TUGboat, 19(4):346, December 1998. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-4/tb61edit.pdf.

Anonymous:TB19-4-347

[2219] Anonymous. Addresses. TUGboat, 19(4):347–348, December 1998. ISSN 0896-3207.

Carlisle:TB19-4-348

[2220] David Carlisle. A seasonal puzzle: XII. TUGboat, 19(4):348, December

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-4/
tb61carl.pdf.

Jett:TB19-4-349

[2221] Mimi Jett. From the President. TUGboat, 19(4):349–350, December

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-4/
tb61jett.pdf.

Beeton:TB19-4-351

[2222] Barbara Beeton. Editorial comments. TUGboat, 19(4):351–352, December

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-4/
tb61beet.pdf.

Flynn:TB19-4-353

[2223] Peter Flynn. Typographers’ Inn. TUGboat, 19(4):353–355, December

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-4/
tb61flynn.pdf.

Misakova:TB19-4-355

[2224] Miroslava Misáková. Typesetting with varying letter widths: New hope for your narrow columns. TUGboat, 19(4):355–365, December 1998. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb19-4/tb61misa.pdf.

Beeton:TB19-4-366

[2225] Barbara Beeton. Editorial: EncTEX, by Petr Olšák. TUGboat, 19(4):366, December 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-4/tb61olsa.pdf.

Olsak:TB19-4-366

[2226] Petr Olšák. EncTEX — A little extension of TEX. TUGboat, 19(4):366–371, December 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-4/tb61olsa.pdf.

Finston:TB19-4-372

[2227] Laurence Finston. ConcTEX: Generating a concordance from TEX input files. TUGboat, 19(4):372–403, December 1998. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb19-4/tb61fins.pdf.

Berdnikov:TB19-4-403

[2228] A. Berdnikov, O. Lapko, M. Kolodin, A. Janishevsky, and A. Burykin. Cyrillic encodings for L^{A}TEX 2ε multi-language documents. TUGboat, 19(4):403–416, December 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-4/tb61berd.pdf.

Pandey:TB19-4-417

[2229] Anshuman Pandey. Romanized Indic and L^{A}TEX. TUGboat, 19(4):417–418, December 1998. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb19-4/tb61pand.pdf.

Beccari:TB19-4-419

[2230] Claudio Beccari and Apostolos Syropoulos. New Greek fonts and the greek option of the babel package. TUGboat, 19(4):419–425, December 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-4/tb61becc.
pdf.

Gibbons:TB19-4-426

[2231] Jeremy Gibbons. ‘Hey — it works!’. TUGboat, 19(4):426–427, December

1998. ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-4/
tb61gibb.pdf.

Thiele:TB19-4-428

[2232] Christina Thiele. The Treasure Chest: A package tour from CTAN — soul.sty. TUGboat, 19(4):428–430, December 1998. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb19-4/tb61soul.pdf.

Anonymous:TB19-4-431

[2233] Anonymous. Les Cahiers GUTenberg, Contents of issue 30. TUGboat, 19 (4):431, December 1998. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb19-4/tb61cahi.pdf.

Anonymous:TB19-4-432

[2234] Anonymous. Future issues. TUGboat, 19(4):432, December 1998. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb19-4/tb61prod.pdf.

Burbank:TB19-4-432

[2235] Mimi Burbank. Production notes. TUGboat, 19(4):432, December 1998. ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-4/tb61prod.
pdf.

Anonymous:TB19-4-433

[2236] Anonymous. Calendar. TUGboat, 19(4):433, December 1998. ISSN 0896-

3207. URL https://tug.org/calendar.html.

Anonymous:TB19-4-434

[2237] Anonymous. TUG’99 Announcement. TUGboat, 19(4):434–436, December

1998. ISSN 0896-3207. URL https://www.tug.org/tug99/tug99.
html.

Anonymous:TB19-4-437

[2238] Anonymous. Institutional members. TUGboat, 19(4):437, December 1998. ISSN 0896-3207.

Anonymous:TB19-4-438

[2239] Anonymous. TUG membership application. TUGboat, 19(4):438, December

1998. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB19-4-439

[2240] Anonymous. TEX consulting and production services. TUGboat, 19(4): 439, December 1998. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB19-4-440

[2241] Anonymous. Y&Y Inc. TUGboat, 19(4):440, December 1998. ISSN 0896-3207.

Anonymous:1998:C

[2242] Anonymous. Cover. TUGboat, 19(4):c1, December 1998. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb19-4/cover1.
pdf.

Anonymous:1998:IFCc

[2243] Anonymous. Inside front cover. TUGboat, 19(4):c2, December 1998. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb19-4/
cover2.pdf.

Anonymous:TB19-4-Cover-3

[2244] Anonymous. Blue Sky Research. TUGboat, 19(4):Cover 3, December 1998. ISSN 0896-3207.

Anonymous:1999:TPa

[2245] Anonymous. Title page. TUGboat, 20(1):1, March 1999. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-1/
titlepage.pdf.

Anonymous:1999:EIa

[2246] Anonymous. Editorial information. TUGboat, 20(1):2, March 1999. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-1/
tb62edit.pdf.

Anonymous:TB20-1-3

[2247] Anonymous. Addresses. TUGboat, 20(1):3, March 1999. ISSN 0896-3207.

Anonymous:TB20-1-4

[2248] Anonymous. TUG’99 Poetry Contest. TUGboat, 20(1):4, March 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-1/tb62poet.
pdf.

Rose:TB20-1-5

[2249] Kristoffer Rose. From the Vice-President. TUGboat, 20(1):5, March

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-1/
tb62rose.pdf.

Beeton:TB20-1-6

[2250] Barbara Beeton. Editorial comments. TUGboat, 20(1):6–7, March 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-1/tb62beet.
pdf.

Andre:TB20-1-8

[2251] Jacques André and Denis Girou. Father Truchet, the typographic point, the Romain du roi, and tilings. TUGboat, 20(1):8–14, March 1999. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb20-1/tb62andr.pdf.

Toledo:TB20-1-15

[2252] Sivan Toledo. A simple technique for typesetting Hebrew with vowel points. TUGboat, 20(1):15–19, March 1999. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb20-1/tb62tole.pdf.

Rahtz:TB20-1-20

[2253] Sebastian Rahtz. The TEX Live Guide, 4^{th} edition. TUGboat, 20(1):20– 44, March 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-1/tb62cddoc.pdf.

Beeton:TB20-1-45

[2254] Barbara Beeton. TEX and the Year 2000. TUGboat, 20(1):45–50, March

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-1/
tb62y2k.pdf.

Beeton:TB20-1-50

[2255] Barbara Beeton. Hyphenation Exception Log — Update. TUGboat, 20(1):50– 51, March 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-1/Hyphen/.

LPT:TB20-1-52

[2256] L^{A}TEX Project Team. L^{A}TEX News, Issue 10, December 1998. TUGboat, 20(1): 52, March 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-1/tb62news.pdf.

Thiele:TB20-1-53

[2257] Christina Thiele. The Treasure Chest: Package tours from CTAN. TUGboat, 20(1):53–58, March 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-1/tb62chest.pdf.

Anonymous:TB20-1-59

[2258] Anonymous. Die TEXnische Komödie 9 (1997, Heft 1–4). TUGboat, 20(1):59– 64, March 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-1/tb20-1komoedie.pdf.

Anonymous:TB20-1-65

[2259] Anonymous. Calendar. TUGboat, 20(1):65–66, March 1999. ISSN

0896-3207. URL https://tug.org/calendar.html;   https:/
/tug.org/TUGboat/tb20-1/tb20-1calendar.pdf.

Anonymous:TB20-1-67

[2260] Anonymous. TUG’99 Announcement. TUGboat, 20(1):67–69, March

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-1/
tb62sched.pdf.

Anonymous:TB20-1-70

[2261] Anonymous. TUGboat web pages. TUGboat, 20(1):70, March 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-1/tb62prod.
pdf.

Anonymous:TB20-1-70-2

[2262] Anonymous. Future issues. TUGboat, 20(1):70, March 1999. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb20-1/tb62prod.pdf.

Burbank:TB20-1-70

[2263] Mimi Burbank. Production notes. TUGboat, 20(1):70, March 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-1/tb62prod.
pdf.

Thiele:TB20-1-71

[2264] Christina Thiele and Arthur Ogawa. Report: TUG 1999 Election. TUGboat, 20(1):71, March 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-1/tb62boar.pdf.

Anonymous:TB20-1-71

[2265] Anonymous. Members of the TUG Board with terms ending in 2003. TUGboat, 20(1):71–76, March 1999. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb20-1/tb62boar.pdf.

Anonymous:TB20-1-76

[2266] Anonymous. Position of President. TUGboat, 20(1):76, March 1999. ISSN 0896-3207.

Anonymous:TB20-1-77

[2267] Anonymous. Institutional members. TUGboat, 20(1):77, March 1999. ISSN 0896-3207.

Anonymous:TB20-1-78

[2268] Anonymous. TUG membership application. TUGboat, 20(1):78, March 1999.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB20-1-79

[2269] Anonymous. TEX consulting and production services. TUGboat, 20 (1):79, March 1999. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB20-1-80

[2270] Anonymous. Y&Y Inc. TUGboat, 20(1):80, March 1999. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb20-1/yandy.pdf.

Anonymous:1999:BSRa

[2271] Anonymous. Blue Sky Research. TUGboat, 20(1):c3, March 1999. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-1/
bluesky.pdf.

Anonymous:1999:Ca

[2272] Anonymous. Cover. TUGboat, 20(1):c1, March 1999. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-1/cover1.
pdf.

Anonymous:1999:IFCa

[2273] Anonymous. Inside front cover. TUGboat, 20(1):c2, March 1999. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-1/
cover2.pdf.

Anonymous:TB20-1-Cover-3

[2274] Anonymous. Blue Sky Research. TUGboat, 20(1):Cover 3, March 1999. ISSN 0896-3207.

Anonymous:1999:TPb

[2275] Anonymous. Title page. TUGboat, 20(2):81, June 1999. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-2/
titlepage.pdf.

Anonymous:TB20-2-83

[2276] Anonymous. Addresses. TUGboat, 20(2):83–84, June 1999. ISSN 0896-3207.

Jett:TB20-2-85

[2277] Mimi Jett. From the President. TUGboat, 20(2):85–86, June 1999. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb20-2/tb63gendel.
pdf.

Beeton:TB20-2-86

[2278] Barbara Beeton. Editorial comments. TUGboat, 20(2):86–87, June 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-2/tb63beet.
pdf.

Childs:TB20-2-87

[2279] Bart Childs. Norman W. Naugle — A Remembrance. TUGboat, 20(2):87–88,

June 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-
2/tb63naugle.pdf.

Langmyhr:TB20-2-89

[2280] Dag Langmyhr. Roswitha von den Schulenburg Graham. TUGboat, 20(2):89,

June 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-
2/tb63roswitha.pdf.

Burbank:TB20-2-89

[2281] Mimi Burbank. You meet the nicest people … Father Everett Larguier. TUGboat, 20(2):89–90, June 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-2/tb63larg.pdf.

Gaulle:TB20-2-91

[2282] Bernard Gaulle. The french package on and off CTAN. TUGboat, 20(2):91–92,

June 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-
2/tb63gaul.pdf.

Anonymous:TB20-2-92

[2283] Anonymous. Response from the CTAN team. TUGboat, 20(2):92, June

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-2/
tb63gaul.pdf.

Beeton:TB20-2-92

[2284] Barbara Beeton. Editor’s commentary. TUGboat, 20(2):92, June 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-2/tb63gaul.
pdf.

Fine:TB20-2-93

[2285] Jonathan Fine. The good name of TEX. TUGboat, 20(2):93, June 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-2/tb63lett.
pdf.

Olsak:TB20-2-93

[2286] Petr Olšak. Reply. TUGboat, 20(2):93, June 1999. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb20-2/tb63lett.pdf.

Flynn:TB20-2-94

[2287] Peter Flynn. Typographers’ Inn. TUGboat, 20(2):94–95, June 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-2/tb63inn.
pdf.

Gelderman:TB20-2-96

[2288] Maarten Gelderman. A short introduction to font characteristics. TUGboat, 20(2):96–104, June 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-2/tb63geld.pdf.

Jackowski:TB20-2-104

[2289] Bogus�aw Jackowski. : Practical and impractical applications. TUGboat, 20(2):104–119, June 1999. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb20-2/tb63jack.pdf.

Pandey:TB20-2-119

[2290] Anshuman Pandey. Typesetting Bengali in TEX. TUGboat, 20(2):119–126,

June 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-
2/tb63pand.pdf.

Hoppner:TB20-2-127

[2291] Klaus Höppner. The CTAN May 1999 CD ROM set by DANTE e.V. and Lehmanns bookstore. TUGboat, 20(2):127, June 1999. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb20-2/tb63cd.pdf.

Dobbelsteen:TB20-2-128

[2292] Gilbert van den Dobbelsteen. Interacting pdfTEX, PERL and ConTEXt. TUGboat, 20(2):128–133, June 1999. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb20-2/tb63dobb.pdf.

Tolksdorf:TB20-2-134

[2293] Robert Tolksdorf. NetBibTEXing. TUGboat, 20(2):134–140, June

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-2/
tb63tolk.pdf.

Preston:TB20-2-140

[2294] Roy Preston. Monk-ey business. TUGboat, 20(2):140, June 1999. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb20-2/versal.pdf.

Gibbons:TB20-2-141

[2295] Jeremy Gibbons. Hey — it works! TUGboat, 20(2):141–142, June 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-2/tb63gibb.
pdf.

Anonymous:TB20-2-143

[2296] Anonymous. Les Cahiers GUTenberg, Contents of Issues 31 (December 1998) and 32 (May 1999). TUGboat, 20(2):143–145, June 1999. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb20-2/tb63cahi.pdf.

Anonymous:TB20-2-146

[2297] Anonymous. Calendar. TUGboat, 20(2):146–147, June 1999. ISSN

0896-3207. URL https://tug.org/calendar.html;   https:/
/tug.org/TUGboat/tb20-2/tb63calendar.pdf.

Anonymous:TB20-2-147

[2298] Anonymous. Future issues. TUGboat, 20(2):147, June 1999. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb20-2/tb63prod.pdf.

Burbank:TB20-2-147

[2299] Mimi Burbank. Production notes. TUGboat, 20(2):147, June 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-2/tb63prod.
pdf.

Anonymous:TB20-2-148

[2300] Anonymous. Institutional members. TUGboat, 20(2):148, June 1999. ISSN

0896-3207. URL https://tug.org/instmem.html;   https://
tug.org/TUGboat/tb20-2/tb63instmem.pdf.

Anonymous:TB20-2-149

[2301] Anonymous. Cambridge University Press. TUGboat, 20(2):149, June

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-2/
tb63consult.pdf.

Anonymous:TB20-2-150

[2302] Anonymous. TUG membership application. TUGboat, 20(2):150, June 1999.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB20-2-151

[2303] Anonymous. TEX consulting and production services. TUGboat, 20(2):151, June 1999. ISSN 0896-3207.

Anonymous:TB20-2-152

[2304] Anonymous. Y&Y Inc. TUGboat, 20(2):152, June 1999. ISSN 0896-3207.

Anonymous:1999:BSRb

[2305] Anonymous. Blue Sky Research. TUGboat, 20(2):c3, June 1999. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-2/
bluesky.pdf.

Anonymous:TB20-2-Cover-3

[2306] Anonymous. Blue Sky Research. TUGboat, 20(2):Cover 3, June 1999. ISSN 0896-3207.

Anonymous:TB20-3-154

[2307] Anonymous. TUG2000 — The 21^{st} Annual Conference. TUGboat, 20(3):154, September 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-3/flier.pdf.

Thiele:TB20-3-155

[2308] Christina Thiele. Vancouver in August. TUGboat, 20(3):155–159, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
tb64edch.pdf.

Anonymous:TB20-3-160

[2309] Anonymous. TUG ’99 Program. TUGboat, 20(3):160–161, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
tug99sched.pdf.

Fulling:TB20-3-162

[2310] Stephen A. Fulling. Keynote: TEX and the Web in the higher education of the future: Dreams and difficulties. TUGboat, 20(3):162–166, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
tb64full.pdf.

Ion:TB20-3-167

[2311] Patrick D. F. Ion. MathML: A key to math on the Web. TUGboat, 20(3): 167–175, September 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-3/tb64ion.pdf.

Lovell:TB20-3-176

[2312] Douglas Lovell. TEXML: Typesetting XML with TEX. TUGboat, 20(3): 176–183, September 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-3/tb64love.pdf.

Topping:TB20-3-184

[2313] Paul Topping. Using MathType to create TEX and MathML equations. TUGboat, 20(3):184–188, September 1999. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb20-3/tb64topp.pdf.

Rowley:TB20-3-189

[2314] Chris Rowley. Models and languages for formatted documents. TUGboat, 20(3):189–195, September 1999. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb20-3/tb64rowl.pdf.

Story:TB20-3-196

[2315] D. P. Story. AcroTEX: Acrobat and TEX team up. TUGboat, 20(3):196–201, September 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-3/tb64story.pdf.

Doumont:TB20-3-202

[2316] Jean luc Doumont. Doing it my way: A lone TEXer in the real world. TUGboat, 20(3):202–207, September 1999. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb20-3/tb64douma.pdf.

Flynn:TB20-3-208

[2317] Peter Flynn. The vulcan package: a repair patch for L^{A}TEX. TUGboat, 20(3): 208–213, September 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-3/tb64flynn.pdf.

Carlisle:TB20-3-214

[2318] David Carlisle, Frank Mittelbach, and Chris Rowley. New interfaces for L^{A}TEX class design, Parts I and II. TUGboat, 20(3):214–216, September 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-3/tb64carl.
pdf.

Bazargan:TB20-3-217

[2319] Kaveh Bazargan. Multi-use documents: The role of the publisher. TUGboat, 20(3):217–220, September 1999. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb20-3/tb64kaveh.pdf.

Bartlett:TB20-3-221

[2320] Frederick H. Bartlett. Very like a nail: Typesetting SGML with TEX. TUGboat, 20(3):221, September 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-3/tb64bart.pdf.

Payne:TB20-3-222

[2321] Harry Payne. Making a book from contributed papers: Print and Web versions. TUGboat, 20(3):222–226, September 1999. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb20-3/tb64payne.pdf.

Kruse:TB20-3-227

[2322] Robert L. Kruse. Managing large projects with PreTEX: A preprocessor for TEX. TUGboat, 20(3):227–230, September 1999. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb20-3/tb64kruse.pdf.

Ogawa:TB20-3-231

[2323] Arthur Ogawa. Database publishing with JAVA and TEX. TUGboat, 20(3):231, September 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-3/tb64ogawa.pdf.

Mailhot:TB20-3-232

[2324] Paul A. Mailhot. Implementing dynamic cross-referencing and PDF with Pre-TEX. TUGboat, 20(3):232–236, September 1999. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb20-3/tb64mail.pdf.

Wang:TB20-3-237

[2325] Hu Wang. A Web-based submission system for meeting abstracts. TUGboat, 20(3):237–240, September 1999. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb20-3/tb64wang.pdf.

Sojka:TB20-3-241

[2326] Petr Sojka. Hyphenation on demand. TUGboat, 20(3):241–247, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
tb64sojka.pdf.

Fine:TB20-3-248

[2327] Jonathan Fine. Active TEX and the DOT input syntax. TUGboat, 20(3): 248–254, September 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-3/tb64fine.pdf.

Doumont:TB20-3-255

[2328] Jean luc Doumont. Drawing effective (and beautiful) graphs with TEX. TUGboat, 20(3):255–261, September 1999. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb20-3/tb64doumB.ps.

McKay:TB20-3-262

[2329] Wendy McKay and Ross Moore. Convenient labelling of graphics, the WARMreader way. TUGboat, 20(3):262–271, September 1999. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb20-3/tb64ross.pdf.

Lesenko:TB20-3-272

[2330] Sergey Lesenko and Laurent Siebenmann. Viewing DVI files with Acrobat Reader: DVIPDF gives birth to AcroDVI. TUGboat, 20(3):272–281, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
tb64lesen.pdf.

Hoenig:TB20-3-282

[2331] Alan Hoenig. MathKit: Alternatives to Computer Modern Mathematics. TUGboat, 20(3):282–289, September 1999. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb20-3/tb64hoenig.pdf.

Popineau:TB20-3-290

[2332] Fabrice Popineau. fpTEX: A teTEX-based distribution for Windows. TUGboat, 20(3):290–297, September 1999. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb20-3/tb64popin.pdf.

McArthur:TB20-3-298

[2333] Jeffrey McArthur. Managing TEX software development projects. TUGboat, 20(3):298–306, September 1999. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb20-3/tb64mcar.pdf.

Murphy:TB20-3-307

[2334] Timothy Murphy. JAVA and TEX. TUGboat, 20(3):307–312, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
tb64murph.pdf.

Thiele:TB20-3-313

[2335] Christina Thiele. Barbara N. Beeton: TUG Board member for 20 years. TUGboat, 20(3):313–315, September 1999. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb20-3/tb64bb20.pdf.

Thiele:TB20-3-316

[2336] Christina Thiele. Text of The Apocalypse as graphics, by Prof. Alban Grimm. TUGboat, 20(3):316, September 1999. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb20-3/tb64exhib.pdf.

Grimm:TB20-3-317

[2337] Alban Grimm. Text of The Apocalypse as graphics. TUGboat, 20(3):317–319, September 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-3/tb64grimm.pdf.

Gurari:TB20-3-320

[2338] Eitan Gurari and Sebastian Rahtz. L^{A}TEX to XML/MathML. TUGboat, 20 (3):320, September 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-3/tb64worka.pdf.

Story:TB20-3-321

[2339] D. P. Story. How to create quality interactive PDF documents for the WWW using L^{A}TEX. TUGboat, 20(3):321, September 1999. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb20-3/tb64workb.pdf.

Doob:TB20-3-322

[2340] Michael Doob. Writing class files: First steps. TUGboat, 20(3):322, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
tb64workc.pdf.

Hoover:TB20-3-323

[2341] Anita Hoover. Converting a L^{A}TEX 2.09 style to a L^{A}TEX 2ε class. TUGboat, 20(3):323, September 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-3/tb64workd.pdf.

Fulling:TB20-3-324

[2342] Stephen A. Fulling. TEX and math on the Web. TUGboat, 20(3):324, Septem-

ber 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-
3/tb64panla.pdf.

Bazargan:TB20-3-325

[2343] Kaveh Bazargan. TEX in publishing. TUGboat, 20(3):325, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
tb64panlb.pdf.

Ogawa:TB20-3-326

[2344] Arthur Ogawa. The Future of L^{A}TEX. TUGboat, 20(3):326, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
tb64panlc.pdf.

Anonymous:TB20-3-327

[2345] Anonymous. TUG ’99 Attendees. TUGboat, 20(3):327–328, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
participants.pdf.

Anonymous:TB20-3-329

[2346] Anonymous. Calendar. TUGboat, 20(3):329–330, September 1999. ISSN

0896-3207. URL https://tug.org/calendar.html.

Preston:TB20-3-330

[2347] Roy Preston. An Analogy with Web Sites. TUGboat, 20(3):330, September

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
web2.gif.

Anonymous:TB20-3-331

[2348] Anonymous. GUTenberg 2000 — L^{A}TEX and XML: Cooperating with the Internet. TUGboat, 20(3):331, September 1999. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb20-3/tb64gut00.pdf.

Anonymous:TB20-3-332

[2349] Anonymous. Institutional members. TUGboat, 20(3):332, September 1999. ISSN 0896-3207.

Anonymous:TB20-3-333

[2350] Anonymous. TUG membership application. TUGboat, 20(3):333, September

1999. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB20-3-334

[2351] Anonymous. TEX consulting and production services. TUGboat, 20(3): 334, September 1999. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB20-3-335

[2352] Anonymous. Cambridge University Press. TUGboat, 20(3):335, September 1999. ISSN 0896-3207.

Anonymous:TB20-3-336

[2353] Anonymous. Y&Y Inc. TUGboat, 20(3):336, September 1999. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb20-3/yandy.pdf.

Anonymous:1999:BSRc

[2354] Anonymous. Blue Sky Research. TUGboat, 20(3):c3, September 1999. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-
3/bluesky.pdf.

Anonymous:1999:FC

[2355] Anonymous. Front cover. TUGboat, 20(3):c1, September 1999. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-3/
cover1.pdf.

Anonymous:TB20-3-cover-3

[2356] Anonymous. Blue Sky Research. TUGboat, 20(3):Cover 3, September 1999. ISSN 0896-3207.

Anonymous:1999:TPc

[2357] Anonymous. Title page. TUGboat, 20(4):337, December 1999. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
titlepage.pdf.

Anonymous:1999:EIb

[2358] Anonymous. Editorial information. TUGboat, 20(4):338, December 1999. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-4/tb65edit.pdf.

Anonymous:TB20-4-339

[2359] Anonymous. Addresses. TUGboat, 20(4):339–340, December 1999. ISSN 0896-3207.

Preston:TB20-4-340

[2360] Roy Preston. Download free fonts! TUGboat, 20(4):340, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
freefonts.pdf.

Jett:TB20-4-341

[2361] Mimi Jett. From the President. TUGboat, 20(4):341, December 1999. ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-4/tb65jett.
pdf.

Beeton:TB20-4-342

[2362] Barbara Beeton. Editorial comments. TUGboat, 20(4):342–343, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
tb65beet.pdf.

Flynn:TB20-4-344

[2363] Peter Flynn. Typographers’ Inn. TUGboat, 20(4):344–347, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
tb65flyn.pdf.

Koutny:TB20-4-347

[2364] Vladimír Koutný. TrueType fonts in TEX. TUGboat, 20(4):347–348, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
tb65kout.pdf.

Zyka:TB20-4-348

[2365] Vít Zýka. The semaphore alphabet. TUGboat, 20(4):348–350, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
tb65zyka.pdf.

Travis:TB20-4-350

[2366] Brian E. Travis. The Paper Path: XML to paper using TEXML. TUGboat, 20 (4):350–356, December 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-4/tb65trav.pdf.

Strokov:TB20-4-356

[2367] Igor I. Strokov. A WYSIWYG TEX implementation. TUGboat, 20(4):356–359, December 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-4/tb65strok.pdf.

Casselman:TB20-4-359

[2368] Bill Casselman. The L^{A}TEX Graphics Companion and TEX Unbound — A review of two books. TUGboat, 20(4):359–364, December 1999. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb20-4/tb65reviews.pdf.

Flynn:TB20-4-364

[2369] Peter Flynn. Digital Typography, by Donald Knuth. TUGboat, 20(4):364–366, December 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-4/tb65reviews.pdf.

Fine:TB20-4-366

[2370] Jonathan Fine. The good name of TEX (TUGboat 2, no. 0(2), p. 93). TUGboat, 20(4):366, December 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-4/tb65errat.pdf.

Thiele:TB20-4-366

[2371] Christina Thiele. TUG 99 (TUGboat 2, no. 0(3)). TUGboat, 20(4):366, Decem-

ber 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-
4/tb65errat.pdf.

Hefferon:TB20-4-367

[2372] Jim Hefferon. A CTAN search page. TUGboat, 20(4):367, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
tb65heff.pdf.

Gibbons:TB20-4-367

[2373] Jeremy Gibbons. Hey — it works! TUGboat, 20(4):367–370, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
tb65gibb.pdf.

Thiele:TB20-4-370

[2374] Christina Thiele. The treasure chest. TUGboat, 20(4):370–374, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
tb65chest.pdf.

LPT:TB20-4-375

[2375] L^{A}TEX Project Team. L^{A}TEX News, Issue 12, December 1999. TUGboat, 20 (4):375, December 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-4/ltnews12.pdf.

Shawyer:TB20-4-376

[2376] Bruce Shawyer. Scaled pictures in L^{A}TEX. TUGboat, 20(4):376–377, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
tb65shaw.pdf.

Taylor:TB20-4-378

[2377] Philip Taylor. Book design for TEX users: Part 2: Practice. TUGboat, 20(4): 378–389, December 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-4/tb65tay2.pdf.

Moore:TB20-4-389

[2378] Ross Moore. Preparation of documents for multiple modes of delivery — Notes from TUG’99. TUGboat, 20(4):389–394, December 1999. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb20-4/tb65moore.pdf.

Anonymous:TB20-4-394

[2379] Anonymous. Les Cahiers GUTenberg, Contents of double issue 33/34 (November 1999). TUGboat, 20(4):394–395, December 1999. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb20-4/tb65summ.pdf.

Anonymous:TB20-4-395

[2380] Anonymous. EuroTEX ’99 Proceedings — Paperless TEX. TUGboat, 20(4): 395–398, December 1999. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb20-4/tb65summ.pdf.

Anonymous:TB20-4-399

[2381] Anonymous. Calendar. TUGboat, 20(4):399–400, December 1999. ISSN

0896-3207. URL https://tug.org/calendar.html.

Anonymous:TB20-4-400

[2382] Anonymous. Future issues. TUGboat, 20(4):400, December 1999. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb20-4/tb65calendar.
pdf.

Burbank:TB20-4-400

[2383] Mimi Burbank. Production notes. TUGboat, 20(4):400, December

1999. ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
tb65calendar.pdf.

Anonymous:TB20-4-401

[2384] Anonymous. TUG 2000 — The 21^{st} Annual Conference. TUGboat, 20(4):401, December 1999. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb20-4/tug2000.pdf.

Anonymous:TB20-4-402

[2385] Anonymous. Institutional members. TUGboat, 20(4):402, December 1999. ISSN 0896-3207.

Anonymous:TB20-4-403

[2386] Anonymous. Statement of ownership. TUGboat, 20(4):403, December 1999.

ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB20-4-403-2

[2387] Anonymous. TEX consulting and production services. TUGboat, 20(4): 403, December 1999. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB20-4-404

[2388] Anonymous. Y&Y Inc. TUGboat, 20(4):404, December 1999. ISSN 0896-3207.

Anonymous:1999:Cb

[2389] Anonymous. Cover. TUGboat, 20(4):c1, December 1999. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb20-4/cover1.
pdf.

Anonymous:1999:IFCb

[2390] Anonymous. Inside front cover. TUGboat, 20(4):c2, December 1999. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb20-4/
cover2.pdf.

Anonymous:TB20-4-Cover-3

[2391] Anonymous. Blue Sky Research. TUGboat, 20(4):Cover 3, December 1999. ISSN 0896-3207.

Anonymous:TB21-1-3

[2392] Anonymous. Addresses. TUGboat, 21(1):3, March 2000. ISSN 0896-3207.

Anonymous:TB21-1-4

[2393] Anonymous. TUG2000 — The 21^{st} Annual Conference. TUGboat, 21(1): 4, March 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb21-1/tug2000.pdf.

Jett:TB21-1-5

[2394] Mimi Jett. From the President. TUGboat, 21(1):5, March 2000. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb21-1/tb66jett.pdf.

Beeton:TB21-1-6

[2395] Barbara Beeton. Editorial comments. TUGboat, 21(1):6, March 2000. ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-1/tb66beet.
pdf.

Fujita:TB21-1-7

[2396] Shinsaku Fujita and Nobuya Tanaka. XΥMTEX (version 2.00) as implementation of the XΥM notation and the XΥM markup language. TUGboat, 21(1):7– 14, March 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb21-1/tb66fuji.pdf.

Farley:TB21-1-15

[2397] David Farley. Don Knuth finally sells out. TUGboat, 21(1):15, March

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-1/
knuth-cartoon.pdf.

Hefferon:TB21-1-15

[2398] Jim Hefferon. The TEX Users Group CTAN site makes a move. TUGboat, 21(1):15, March 2000. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb21-1/tb66heff.pdf.

TT:TB21-1-16

[2399] The TUGboat Team. TEXLive 5 and the TEX Catalogue. TUGboat, 21(1): 16, March 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb21-1/tb66cd.pdf.

Anonymous:TB21-1-17

[2400] Anonymous. Graham Williams’ TEX Catalogue. TUGboat, 21(1):17–90, March 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb21-1/tb66catal.pdf.

Eijkhout:TB21-1-91

[2401] Victor Eijkhout. The bag of tricks. TUGboat, 21(1):91, March 2000. ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-1/tb66eijk.
pdf.

Anonymous:TB21-1-92

[2402] Anonymous. Calendar. TUGboat, 21(1):92, March 2000. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB21-1-93

[2403] Anonymous. Future issues. TUGboat, 21(1):93, March 2000. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb21-1/tb66prod.pdf.

Burbank:TB21-1-93

[2404] Mimi Burbank. Production notes. TUGboat, 21(1):93, March 2000. ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-1/tb66prod.
pdf.

Anonymous:TB21-1-94

[2405] Anonymous. Institutional members. TUGboat, 21(1):94, March 2000. ISSN 0896-3207.

Anonymous:TB21-1-95

[2406] Anonymous. TUG membership application. TUGboat, 21(1):95, March 2000.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB21-1-96

[2407] Anonymous. TEX consulting and production services. TUGboat, 21 (1):96, March 2000. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:2000:TPa

[2408] Anonymous. Title page. TUGboat, 21(1):97, March 2000. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-1/
titlepage.pdf.

Anonymous:2000:EIa

[2409] Anonymous. Editorial information. TUGboat, 21(1):98, March 2000. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-1/
tb66edit.pdf.

Anonymous:2000:BSR

[2410] Anonymous. Blue Sky Research. TUGboat, 21(1):c3, March 2000. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-1/
bluesky.pdf.

Anonymous:2000:Ca

[2411] Anonymous. Cover. TUGboat, 21(1):c1, March 2000. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-1/cover1.
pdf.

Anonymous:2000:IFCa

[2412] Anonymous. Inside front cover. TUGboat, 21(1):c2, March 2000. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-1/
cover2.pdf.

Anonymous:TB21-1-Cover-3

[2413] Anonymous. Blue Sky Research. TUGboat, 21(1):Cover 3, March 2000. ISSN 0896-3207.

Anonymous:TB21-1-xx

[2414] Anonymous. Supplement: TEXLive 5 CD. TUGboat, 21(1):??, March 2000. ISSN 0896-3207.

Anonymous:2000:TPb

[2415] Anonymous. Title page. TUGboat, 21(2):97, June 2000. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-2/
titlepage.pdf.

Anonymous:2000:EIb

[2416] Anonymous. Editorial information. TUGboat, 21(2):98, June 2000. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-2/
tb67edit.pdf.

Anonymous:TB21-2-99

[2417] Anonymous. Addresses. TUGboat, 21(2):99, June 2000. ISSN 0896-3207.

Preston:TB21-2-100

[2418] Roy Preston. Font identification. TUGboat, 21(2):100, June 2000. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb21-2/cartoon.pdf.

Jett:TB21-2-101

[2419] Mimi Jett. From the President. TUGboat, 21(2):101, June 2000. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb21-2/tb67jett.pdf.

Beeton:TB21-2-102

[2420] Barbara Beeton. Editorial comments. TUGboat, 21(2):102–103, June

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-2/
tb67beet.pdf.

Levien:TB21-2-103

[2421] Advogato (Raph Levien). Interview: Donald E. Knuth. TUGboat, 21(2):103– 110, June 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb21-2/tb67advo.pdf.

Gratzer:TB21-2-111

[2422] G. Grätzer. Turbulent transition. TUGboat, 21(2):111–113, June 2000. ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-2/tb67grat.
pdf.

Lemberg:TB21-2-113

[2423] Werner Lemberg. Thai fonts. TUGboat, 21(2):113–120, June 2000. ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-2/tb67lemb.
pdf.

Toledo:TB21-2-121

[2424] Sivan Toledo. Exploiting rich fonts. TUGboat, 21(2):121–129, June

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-2/
tb67tole.pdf.

Berdnikov:TB21-2-129

[2425] Alexander Berdnikov, Hans Hagen, Taco Hoekwater, and Bogus�aw Jackowski. Even more MetaFun with : A request for permission. TUGboat, 21(2):129–130, June 2000. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb21-2/tb67metafun.pdf.

Hobby:TB21-2-131

[2426] John D. Hobby. Extending : Response to “Even more MetaFun”. TUGboat, 21(2):131–132, June 2000. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb21-2/tb67metafun.pdf.

Beeton:TB21-2-132

[2427] Barbara Beeton. Hyphenation exception log. TUGboat, 21(2):132–133, June

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-
2/hyphen;   https://tug.org/TUGboat/tb21-2/tb67hyph.
pdf.

Gibbons:TB21-2-133

[2428] Jeremy Gibbons. Hey — it works! TUGboat, 21(2):133–135, June 2000. ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-2/tb67gibb.
pdf.

Thiele:TB21-2-136

[2429] Christina Thiele. The treasure chest. TUGboat, 21(2):136–142, June

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-2/
tb67chest.pdf.

LPT:TB21-2-143

[2430] L^{A}TEX Project Team. L^{A}TEX News, Issue 13, June 2000. TUGboat, 21(2):143,

June 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-
2/tb67news.pdf.

Anonymous:TB21-2-144

[2431] Anonymous. Calendar. TUGboat, 21(2):144–145, June 2000. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb21-2/tb67calendar.
pdf.

DeLand:TB21-2-145

[2432] Donald DeLand. Report from the TUG Treasurer. TUGboat, 21(2):145–146,

June 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-
2/tb67treas.pdf.

Anonymous:TB21-2-146

[2433] Anonymous. Future issues. TUGboat, 21(2):146, June 2000. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb21-2/tb67prod.pdf.

Burbank:TB21-2-146

[2434] Mimi Burbank. Production notes. TUGboat, 21(2):146, June 2000. ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-2/tb67prod.
pdf.

Anonymous:TB21-2-147

[2435] Anonymous. 2001 TEX Users Group election. TUGboat, 21(2):147, June

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-2/
tb67elec.pdf.

Anonymous:TB21-2-148

[2436] Anonymous. TUG2000 — The 21^{st} Annual Conference. TUGboat, 21(2):148,

June 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-
2/tug2000.pdf.

Anonymous:TB21-2-149

[2437] Anonymous. Institutional members. TUGboat, 21(2):149, June 2000. ISSN 0896-3207.

Anonymous:TB21-2-150

[2438] Anonymous. TUG membership application. TUGboat, 21(2):150, June 2000.

ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB21-2-151

[2439] Anonymous. TEX consulting and production services. TUGboat, 21(2):151,

June 2000. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB21-2-152

[2440] Anonymous. IBM techexplorer. TUGboat, 21(2):152, June 2000. ISSN 0896-3207.

Anonymous:2000:Cb

[2441] Anonymous. Cover. TUGboat, 21(2):c1, June 2000. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb21-2/cover1.pdf.

Anonymous:2000:IFCb

[2442] Anonymous. Inside front cover. TUGboat, 21(2):c2, June 2000. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-2/
cover2.pdf.

Anonymous:TB21-2-

[2443] Anonymous. CTAN CDs: A 3-disk collection. TUGboat, 21(2):, June 2000. ISSN 0896-3207.

Anonymous:TB21-2-cover-3

[2444] Anonymous. Blue Sky Research. TUGboat, 21(2):Cover 3, June 2000. ISSN 0896-3207.

Anonymous:TB21-3-154

[2445] Anonymous. TUG’2001 Announcement. TUGboat, 21(3):154, Septem-

ber 2000. ISSN 0896-3207. URL https://www.tug.org/tug2001/
announcement.html.

Fairbairns:TB21-3-155

[2446] Robin Fairbairns. Editorial Comments — TUG ’2000. TUGboat, 21(3): 155–156, September 2000. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb21-3/tb68fair.pdf.

Anonymous:TB21-3-157

[2447] Anonymous. TUG ’2000 Program. TUGboat, 21(3):157–158, September

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-3/
tb68prog.pdf.

Bayart:TB21-3-159

[2448] Benjamin Bayart. The description language chosen for FDNTEX. TUGboat, 21(3):159–175, September 2000. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb21-3/tb68bay.pdf.

Beeton:TB21-3-176

[2449] Barbara Beeton. Unicode and math, a combination whose time has come — Finally! TUGboat, 21(3):176–185, September 2000. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb21-3/tb68beet.pdf.

Berdnikov:TB21-3-186

[2450] Alexander Berdnikov, Yury Yarmola, Olga Lapko, and Andrew Janishewsky. Some experience in converting LH Fonts from to Type1 format [Abstract]. TUGboat, 21(3):186, September 2000. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb21-3/tb68berd2.pdf.

Bzyl:TB21-3-187

[2451] W�odek Bzyl. Typesetting TEX documents containing computer code. TUGboat, 21(3):187–192, September 2000. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb21-3/tb68bzyl.pdf.

Carlisle:TB21-3-193

[2452] David Carlisle. xmltex: a non validating (and not 100% conforming) namespace aware XML parser implemented in TEX. TUGboat, 21(3):193–199, September 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb21-3/tb68carl.pdf.

DeLand:TB21-3-200

[2453] Donald DeLand. Developing interactive, Web-based courseware [Abstract]. TUGboat, 21(3):200, September 2000. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb21-3/tb68dela.pdf.

Downes:TB21-3-201

[2454] Michael Downes. The amsrefs L^{A}TEX package and the amsxport BibTEX style. TUGboat, 21(3):201–209, September 2000. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb21-3/tb68down.pdf.

Fine:TB21-3-210

[2455] Jonathan Fine. Line breaking and page breaking. TUGboat, 21(3):210–221, September 2000. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb21-3/tb68fine.pdf.

Goossens:TB21-3-222

[2456] Michel Goossens and Sebastian Rahtz. PassiveTEX: from XML to PDF. TUGboat, 21(3):222–234, September 2000. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb21-3/tb68goos.pdf.

Gostanza:TB21-3-235

[2457] Pedro Palao Gostanza. Fast scanners and self-parsing in TEX. TUGboat, 21(3): 235–242, September 2000. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb21-3/tb68gost.pdf.

Kakugawa:TB21-3-243

[2458] Hirotsugu Kakugawa. A device-independent DVI interpreter library for various output devices. TUGboat, 21(3):243–249, September 2000. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb21-3/tb68kaku.pdf.

Kolodin:TB21-3-250

[2459] M. Y. Kolodin, O. V. Eterevksy, O. G. Lapko, and I. A. Makhovaya. “Russian style” with L^{A}TEX and babel: what does it look like and how does it work [Abstract]. TUGboat, 21(3):250, September 2000. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb21-3/tb68kolo.pdf.

Kostin:TB21-3-251

[2460] Alex Kostin and Michael Vulis. Mixing TEX & PostScript: The GEX model. TUGboat, 21(3):251–264, September 2000. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb21-3/tb68kost.pdf.

Lavaud:TB21-3-265

[2461] Michel Lavaud. The AsTEX Assistant and Navigator [Abstract]. TUGboat, 21(3):265, September 2000. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb21-3/tb68lava.pdf.

Lipkin:TB21-3-266

[2462] Bernice Sacks Lipkin. L^{A}TEX and the personal database. TUGboat, 21(3): 266–277, September 2000. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb21-3/tb68lipk.pdf.

Mittelbach:TB21-3-278

[2463] Frank Mittelbach. Formatting documents with floats: A new algorithm for L^{A}TEX 2ε. TUGboat, 21(3):278–290, September 2000. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb21-3/tb68mittel.pdf.

Murphy:TB21-3-291

[2464] Timothy Murphy. The Penrose notation: a L^{A}TEX challenge [Abstract]. TUGboat, 21(3):291, September 2000. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb21-3/tb68murp.pdf.

Syropoulos:TB21-3-292

[2465] Apostolos Syropoulos and Richard W. D. Nickalls. A Perl port of the mathsPIC graphics package. TUGboat, 21(3):292–297, September

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-3/
tb68syro.pdf.

Nikulina:TB21-3-298

[2466] Marina Yu. Nikulina and Alexander S. Berdnikov. Chess macros for chess games and puzzles. TUGboat, 21(3):298–302, September 2000. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb21-3/tb68niku.pdf.

Plaice:TB21-3-303

[2467] John Plaice. Omega version 2 [Abstract]. TUGboat, 21(3):303, September

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-3/
tb68plai.pdf.

Taylor:TB21-3-304

[2468] Philip Taylor and Jiří Zlatuška. The N T S project: from conception to birth [Abstract]. TUGboat, 21(3):304, September 2000. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb21-3/tb68tayl.pdf.

Anonymous:TB21-3-305

[2469] Anonymous. Miscellaneous photos. TUGboat, 21(3):305, September

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-3/
page305.pdf.

Anonymous:TB21-3-306

[2470] Anonymous. Calendar. TUGboat, 21(3):306–307, September 2000. ISSN 0896-3207.

Anonymous:TB21-3-307

[2471] Anonymous. Institutional members. TUGboat, 21(3):307, September

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-3/
tb68part.pdf.

Anonymous:TB21-3-308

[2472] Anonymous. TUG ’2000 Attendees. TUGboat, 21(3):308–309, September

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-3/
tb68part.pdf.

Anonymous:TB21-3-310

[2473] Anonymous. TUG membership application. TUGboat, 21(3):310, September

2000. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB21-3-311

[2474] Anonymous. TEX consulting and production services. TUGboat, 21(3): 311, September 2000. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB21-3-312

[2475] Anonymous. IBM techexplorer. TUGboat, 21(3):312, September 2000. ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-3/ibm2000.
pdf.

Anonymous:2000:Cc

[2476] Anonymous. Cover. TUGboat, 21(3):c1, September 2000. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-3/
cover1.pdf.

Anonymous:2000:IFCc

[2477] Anonymous. Inside front cover. TUGboat, 21(3):c2, September 2000. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-3/
cover2.pdf.

Anonymous:TB21-3-cover-3

[2478] Anonymous. Blue Sky Research. TUGboat, 21(3):Cover 3, September 2000. ISSN 0896-3207.

Beeton:TB21-4-315

[2479] Barbara Beeton. Editorial Comments. TUGboat, 21(4):315, December

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-4/
tb69beet.pdf.

Thanh:2000:MTE

[2480] Hàn Thê´ Thành. Micro-typographic extensions to the TEX typesetting system. TUGboat, 21(4):317, December 2000. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb21-4/tb69thanh.pdf.

Anonymous:TB21-4-435

[2481] Anonymous. TUG ’2001 Announcement. TUGboat, 21(4):435, December

2000. ISSN 0896-3207. URL https://tug.org/TUGboat/tb21-4/
tugboat2001gray.pdf.

Anonymous:TB21-4-436

[2482] Anonymous. Calendar. TUGboat, 21(4):436, December 2000. ISSN 0896-

3207. URL https://tug.org/calendar.html.

Anonymous:TB21-4-437

[2483] Anonymous. Institutional members. TUGboat, 21(4):437, December 2000. ISSN 0896-3207.

Anonymous:TB21-4-438

[2484] Anonymous. TUG membership application. TUGboat, 21(4):438, December

2000. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB21-4-439

[2485] Anonymous. TEX consulting and production services. TUGboat, 21(4): 439, December 2000. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB21-4-440

[2486] Anonymous. IBM techexplorer. TUGboat, 21(4):440, December 2000. ISSN

0896-3207. URL https://tug.org/TUGboat/tb21-4/ibm2000.
pdf.

Anonymous:2000:CET

[2487] Anonymous. Cover, epigram, and table of contents. TUGboat, 21(4):c1, December 2000. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb21-4/cover96.pdf.

Anonymous:TB21-4-cover-3

[2488] Anonymous. Blue Sky Research. TUGboat, 21(4):Cover 3, December 2000. ISSN 0896-3207.

Preston:2001:TA

[2489] Roy Preston. Typohol Anon. TUGboat, 22(1–2):4, March 2001. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/typhohol.pdf.

Anonymous:2001:TPa

[2490] Anonymous. Title page. TUGboat, 22(1–2):97, March 2001. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/titlepage.pdf.

Anonymous:2001:EI

[2491] Anonymous. Editorial information. TUGboat, 22(1–2):98, March 2001. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-
1-2/tb70edit.pdf.

Anonymous:2001:BSRa

[2492] Anonymous. Blue Sky Research. TUGboat, 22(1–2):c3, March 2001. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-2/
bluesky.pdf.

Anonymous:2001:Ca

[2493] Anonymous. Cover. TUGboat, 22(1–2):c1, March 2001. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb22-1-2/cover1.
pdf.

Anonymous:2001:IFCa

[2494] Anonymous. Inside front cover. TUGboat, 22(1–2):c2, March 2001. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/cover2.pdf.

Anonymous:TB22-1-3

[2495] Anonymous. Addresses. TUGboat, 22(1/2):3, March 2001. ISSN 0896-3207.

Jett:TB22-1-5

[2496] Mimi Jett. From the President. TUGboat, 22(1/2):5, March 2001. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb22-1-2/tb70pres.
pdf.

Beeton:TB22-1-6

[2497] Barbara Beeton. Editorial comments. TUGboat, 22(1/2):6–7, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70beet.pdf.

Hefferon:TB22-1-8

[2498] Jim Hefferon. Why TEX? TUGboat, 22(1/2):8–14, March 2001. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb22-1-2/tb70heff.
pdf.

Anonymous:TB22-1-15

[2499] Anonymous. Question & answer session with Donald Knuth, U.K. TUG, Oxford, Sunday, 12 September 1999. TUGboat, 22(1/2):15–19, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70knut.pdf.

Anonymous:TB22-1-20

[2500] Anonymous. How (La)TEX changed the face of mathematics: An E-interview with Leslie Lamport, the author of L^{A}TEX. TUGboat, 22(1/2):20–22, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70lamp.pdf.

Flynn:TB22-1-23

[2501] Peter Flynn. Typographers’ Inn. TUGboat, 22(1/2):23, March 2001.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-2/
tb70flyn.pdf.

Mittelbach:TB22-1-24

[2502] Frank Mittelbach. Laudatio for Professor Hermann Zapf. TUGboat, 22 (1/2):24–26, March 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-1-2/tb70laud-revised.pdf.

Zapf:TB22-1-26

[2503] Hermann Zapf. My collaboration with Don Knuth and my font design work. TUGboat, 22(1/2):26–30, March 2001. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb22-1-2/tb70zapf.pdf.

Beeton:TB22-1-31

[2504] Barbara Beeton. Hyphenation exception log. TUGboat, 22(1/2):31–32, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70hyf.pdf.

Jackson:TB22-1-32

[2505] Laura Elizabeth Jackson and Herbert Voß. LYX — An Open Source document processor. TUGboat, 22(1/2):32–41, March 2001. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb22-1-2/tb70voss.pdf.

Lewenberg:TB22-1-42

[2506] Adam H. Lewenberg. DVII: A TEX dvi file information utility. TUGboat, 22 (1/2):42–45, March 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-1-2/tb70lewe.pdf.

Hobby:TB22-1-46

[2507] John D. Hobby. Drawing graphs with MetaPost. TUGboat, 22(1/2):46–58, March 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-1-2/tb70hobb.pdf.

Hagen:TB22-1-58

[2508] Hans Hagen. The status quo of the N T S project. TUGboat, 22(1/2):58– 66, March 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-1-2/tb70hans.pdf.

Adams:TB22-1-67

[2509] William Adams. The treasure chest. TUGboat, 22(1/2):67–74, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70chest.pdf.

Gratzer:TB22-1-74

[2510] George Grätzer. Publishing legacy documents on the Web. TUGboat, 22 (1/2):74–78, March 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-1-2/tb70grat.pdf.

Roegel:TB22-1-78

[2511] Denis Roegel. Anatomy of a macro. TUGboat, 22(1/2):78–82, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70roeg.pdf.

Eijkhout:TB22-1-83

[2512] Victor Eijkhout. The bag of tricks. TUGboat, 22(1/2):83–86, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70eijk.pdf.

Mauw:TB22-1-87

[2513] Sjouke Mauw and Victor Bos. Drawing message sequence charts with L^{A}TEX. TUGboat, 22(1/2):87–92, March 2001. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb22-1-2/tb70bos.pdf.

Mittelbach:TB22-1-93

[2514] Frank Mittelbach. The trace package. TUGboat, 22(1/2):93–99, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70mitt.pdf.

Anonymous:TB22-1-100

[2515] Anonymous. Les Cahiers GUTenberg, Contents of issues 35/36 (May 2000) and 37/38 (December 2000). TUGboat, 22(1/2):100–102, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70cahi.pdf.

Anonymous:TB22-1-103

[2516] Anonymous. Calendar. TUGboat, 22(1/2):103–104, March 2001. ISSN 0896-

3207. URL https://tug.org/calendar.html.

Anonymous:TB22-1-105

[2517] Anonymous. TUG 2001 Announcement. TUGboat, 22(1/2):105, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tugboat2001gray.pdf.

DeMeritt:TB22-1-106

[2518] Susan DeMeritt. Minutes of TEX Users Group Annual General Meeting, 15 August 2000, Oxford, England. TUGboat, 22(1/2):106, March 2001. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb22-1-2/tb70agm.pdf.

DeLand:TB22-1-107

[2519] Don DeLand. Financial statement, 2000. TUGboat, 22(1/2):107, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tb70treas.pdf.

Ogawa:TB22-1-108

[2520] Arthur Ogawa. TUG Election Notice. TUGboat, 22(1/2):108, March

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-1-
2/tug2003elec.pdf.

Anonymous:TB22-1-110

[2521] Anonymous. TUG membership application. TUGboat, 22(1/2):110, March

2001. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB22-1-111

[2522] Anonymous. TEX consulting and production services. TUGboat, 22 (1/2):111, March 2001. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB22-1-112

[2523] Anonymous. Just Published: TEX Reference Manual by David Bausum. TUGboat, 22(1/2):112, March 2001. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb22-1-2/kluwer-bausum.pdf.

Anonymous:TB22-1-cover-3

[2524] Anonymous. Blue Sky Research. TUGboat, 22(1/2):Cover 3, March 2001. ISSN 0896-3207.

Anonymous:2001:TPb

[2525] Anonymous. Title page. TUGboat, 22(3):113, September 2001. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71titlepage.pdf.

Anonymous:2001:EN

[2526] Anonymous. Editorial notes. TUGboat, 22(3):114, September 2001. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71editpage.pdf.

Anonymous:TB22-3-115

[2527] Anonymous. TUG 2001 Program. TUGboat, 22(3):115–116, September

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71program.pdf.

Anonymous:TB22-3-117

[2528] Anonymous. Participants at the 22nd Annual TUG Meeting. TUGboat, 22 (3):117, September 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-3/tb71participants.pdf.

Hagen:TB22-3-118

[2529] Hans Hagen. Where will the odyssey bring us? TUGboat, 22(3):118, Septem-

ber 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-
3/tb71hagen-odyssey.pdf.

Jett:TB22-3-119

[2530] Mimi Jett. Future of publishing, Part 2. TUGboat, 22(3):119, September

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71jett.pdf.

Richter:TB22-3-120

[2531] William Richter. Integrating TEX into a document imaging system. TUGboat, 22(3):120–130, September 2001. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb22-3/tb71richter.pdf.

Ogawa:TB22-3-131

[2532] Arthur Ogawa. REVTEX version 4.0, an authoring package by the American Physical Society. TUGboat, 22(3):131–133, September 2001. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb22-3/tb71ogawa.pdf.

Schwartz:TB22-3-134

[2533] Anita Schwartz. The TEX History Panel. TUGboat, 22(3):134–135, September

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71schwartz.pdf.

Hagen:TB22-3-136

[2534] Hans Hagen. Using TEX for high end typesetting. TUGboat, 22(3):136, September 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-3/tb71hagen-highend.pdf.

Flynn:TB22-3-137

[2535] Peter Flynn. TEX — a mass market product? Or just an image in need of a makeover? TUGboat, 22(3):137–139, September 2001. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb22-3/tb71flynn.pdf.

Tulett:TB22-3-140

[2536] David Tulett. L^{A}TEX for Windows: a user’s perspective. TUGboat, 22(3): 140–145, September 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-3/tb71tulett.pdf.

Thanh:TB22-3-146

[2537] Hàn Thê´ Thành. Margin kerning and font expansion with pdfTEX. TUGboat, 22(3):146–148, September 2001. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb22-3/tb71thanh.pdf.

Moore:TB22-3-149

[2538] Ross Moore. PDF presentations using the Marslide package. TUGboat, 22(3): 149–159, September 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-3/tb71moore-marslide.pdf.

Hagen:TB22-3-160

[2539] Hans Hagen. Using TEX to enhance your presentations. TUGboat, 22(3):160, September 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-3/tb71hagen-presentations.pdf.

Story:TB22-3-161

[2540] Donald P. Story. Techniques of introducing document-level JavaScript into a PDF file from a L^{A}TEX source. TUGboat, 22(3):161–167, September

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71story.pdf.

Moore:TB22-3-168

[2541] Ross Moore. Online self-marking quizzes, pdfTEX, exerquiz. TUGboat, 22(3): 168–179, September 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-3/tb71moore-quiz.pdf.

Schroder:TB22-3-180

[2542] Martin Schröder. Using pdfTEX in a PDF-based imposition tool. TUGboat, 22(3):180, September 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-3/tb71schroder.pdf.

Beebe:TB22-3-181

[2543] Nelson H. F. Beebe. pdfTEX Panel. TUGboat, 22(3):181–187, September

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71beebe-pdf.pdf.

Moore:TB22-3-188

[2544] Ross Moore. Adobe plugin for WARMreader. TUGboat, 22(3):188–196, September 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-3/tb71moore-warm.pdf.

Oliver:TB22-3-197

[2545] Stephen Oliver. The TEXspec tool for computer-aided software engineering. TUGboat, 22(3):197–203, September 2001. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb22-3/tb71oliver.pdf.

Hammond:TB22-3-204

[2546] William Hammond. GELLMU: A bridge for authors from L^{A}TEX to XML. TUGboat, 22(3):204–207, September 2001. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb22-3/tb71hammond.pdf.

Caviness:TB22-3-208

[2547] Bob Caviness. Creating Math Web Documents (Workshop). TUGboat, 22 (3):208, September 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-3/tb71caviness.pdf.

Hoenig:TB22-3-209

[2548] Alan Hoenig. Typesetting Hebrew with TEX. TUGboat, 22(3):209–215, September 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-3/tb71hoenig-hebrew.pdf.

Hoenig:TB22-3-216

[2549] Alan Hoenig. Modernizing Computer Modern. TUGboat, 22(3):216–219, September 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-3/tb71hoenig-modern.pdf.

Beebe:TB22-3-220

[2550] Nelson H. F. Beebe. Fonts Panel. TUGboat, 22(3):220–227, September

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71beebe-fonts.pdf.

Downes:TB22-3-228

[2551] Michael Downes. Managing multiple TDS trees. TUGboat, 22(3):228–237, September 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-3/tb71downes.pdf.

Doob:TB22-3-238

[2552] Michael Doob. Installing a CTAN mirror on your desktop. TUGboat, 22(3): 238–239, September 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-3/tb71doob.pdf.

Koch:TB22-3-240

[2553] Richard Koch. Installing TeXshop. TUGboat, 22(3):240–246, September

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71koch.pdf.

Adams:TB22-3-247

[2554] William Adams. Font installation: Agfa/Eaglefeather to Linotype Zapfino. TUGboat, 22(3):247–250, September 2001. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb22-3/tb71adams.pdf.

Anonymous:TB22-3-251

[2555] Anonymous. Calendar. TUGboat, 22(3):251–252, September 2001. ISSN 0896-3207.

Anonymous:TB22-3-253

[2556] Anonymous. TUG 2003 Announcement. TUGboat, 22(3):253, September

2001. ISSN 0896-3207. URL https://www.tug.org/tug2003/.

Anonymous:TB22-3-254

[2557] Anonymous. Institutional members. TUGboat, 22(3):254, September

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71inst.pdf.

Anonymous:TB22-3-255

[2558] Anonymous. TEX consulting and production services. TUGboat, 22(3):255, September 2001. ISSN 0896-3207.

Anonymous:TB22-3-256

[2559] Anonymous. Just Published: TEX Reference Manual by David Bausum. TUGboat, 22(3):256, September 2001. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb22-3/kluwer.pdf.

Anonymous:2001:BSRb

[2560] Anonymous. Blue Sky Research. TUGboat, 22(3):c3, September 2001. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-
3/bluesky.pdf.

Anonymous:2001:Cb

[2561] Anonymous. Cover. TUGboat, 22(3):c1, September 2001. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
cover1.pdf.

Anonymous:2001:IFCb

[2562] Anonymous. Inside front cover. TUGboat, 22(3):c2, September 2001. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-3/
tb71cover2.pdf.

Anonymous:TB22-3-cover-3

[2563] Anonymous. Blue Sky Research. TUGboat, 22(3):Cover 3, September 2001. ISSN 0896-3207.

Anonymous:TB22-4-259

[2564] Anonymous. Addresses. TUGboat, 22(4):259–260, December 2001. ISSN 0896-3207.

Preston:TB22-4-260

[2565] Roy Preston. Ya can’t touch us! TUGboat, 22(4):260, December 2001. ISSN

0896-3207. URL https://tug.org/TUGboat/tb22-4/m-stop.
pdf.

Ogawa:TB22-4-261

[2566] Arthur Ogawa. From the Board of Directors. TUGboat, 22(4):261–263, December 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-4/tb72ogawa.pdf.

Beeton:TB22-4-263

[2567] Barbara Beeton. Editorial comments. TUGboat, 22(4):263–265, December

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-4/
tb72ogawa.pdf.

Story:TB22-4-265

[2568] D. P. Story. execJS: A new technique for introducing discardable JavaScript into a PDF file from a L^{A}TEX source. TUGboat, 22(4):265–268, December

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-4/
tb72story.pdf.

Goossens:TB22-4-269

[2569] Michel Goossens. L^{A}TEX, SVG, fonts. TUGboat, 22(4):269–280, December

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-4/
tb72goos.pdf.

Forkosh:TB22-4-280

[2570] John Forkosh. mimeTex announcement. TUGboat, 22(4):280–281, December

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-4/
tb72forkosh.pdf.

Berry:TB22-4-281

[2571] Karl Berry. Making outline fonts from bitmap images. TUGboat, 22(4): 281–285, December 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-4/tb72berry.pdf.

Fujita:TB22-4-285

[2572] Shinsaku Fujita and Nobuya Tanaka. Size reduction of chemical structural formulas in XΥMTEX (version 3.00). TUGboat, 22(4):285–289, December

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-4/
tb72fuji.pdf.

Niepraschk:TB22-4-290

[2573] Rolf Niepraschk and Herbert Voß. The package ps4pdf: from PostScript to PDF. TUGboat, 22(4):290–292, December 2001. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb22-4/tb72voss4pdf.pdf.

Fine:TB22-4-292

[2574] Jonathan Fine. Instant Preview and the TEX daemon. TUGboat, 22(4):292–298, December 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-4/tb72fine.pdf.

Roegel:TB22-4-298

[2575] Denis Roegel. Space geometry with . TUGboat, 22(4):298–314, December 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-4/tb72roeg.pdf.

Voss:TB22-4-314

[2576] Jana Voß and Herbert Voß. The plot functions of pst-plot. TUGboat, 22(4): 314–319, December 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-4/tb72vossplot.pdf.

Voss:TB22-4-319

[2577] Herbert Voß. Three dimensional plots with pst-3dplot. TUGboat, 22(4): 319–329, December 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-4/tb72voss3d.pdf.

Perlis:TB22-4-330

[2578] Alexander R. Perlis. Axis alignment in Xy-pic diagrams. TUGboat, 22(4): 330–334, December 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-4/tb72perlax.pdf.

Obrecht:TB22-4-334

[2579] Christian Obrecht. Eukleides: A geometry drawing language. TUGboat, 22 (4):334–337, December 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-4/tb72obre.pdf.

Moye:TB22-4-338

[2580] Stephen Moye. TEX Reference Manual, by David Bausum. TUGboat, 22(4): 338–339, December 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-4/tb72moye.pdf.

Wilson:TB22-4-339

[2581] Peter Wilson. Glisterings. TUGboat, 22(4):339–341, December 2001. ISSN

0896-3207. URL https://tug.org/TUGboat/tb22-4/tb72wilson.
pdf.

Adams:TB22-4-341

[2582] William Adams. The treasure chest. TUGboat, 22(4):341–348, December

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-4/
tb72chest.pdf.

Tulett:TB22-4-349

[2583] David M. Tulett. Highlighting in the L^{A}TEX picture environment. TUGboat, 22(4):349, December 2001. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb22-4/tb72tulett.pdf.

Perlis:TB22-4-350

[2584] Alexander R. Perlis. A complement to \smash, \llap, and \rlap. TUGboat, 22(4):350–352, December 2001. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb22-4/tb72perlS.pdf.

Burt:TB22-4-353

[2585] John Burt. Typesetting critical editions of poetry. TUGboat, 22(4):353–360, December 2001. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb22-4/tb72burt.pdf.

Verna:TB22-4-361

[2586] Didier Verna. CV formatting with CurVe. TUGboat, 22(4):361–364, December

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-4/
tb72verna.pdf.

Anonymous:TB22-4-365

[2587] Anonymous. Les Cahiers GUTenberg, Contents of double issue 39/40 (May 2001). TUGboat, 22(4):365–367, December 2001. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb22-4/tb72cahi.pdf.

Anonymous:TB22-4-368

[2588] Anonymous. TUG ’2002 Announcement. TUGboat, 22(4):368, December

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-4/
tug2002flyer.pdf.

Anonymous:TB22-4-369

[2589] Anonymous. Calendar. TUGboat, 22(4):369–370, December 2001. ISSN 0896-3207.

Anonymous:TB22-4-371

[2590] Anonymous. TUG ’2003 Announcement. TUGboat, 22(4):371, December

2001. ISSN 0896-3207. URL https://tug.org/TUGboat/tb22-4/
TUG2003flyer_m8.pdf.

Anonymous:TB22-4-372

[2591] Anonymous. EuroTEX ’2003 — The 14^{th} European TEX Conference. TUGboat, 22(4):372, December 2001. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb22-4/eurotex2003-a4.pdf.

Anonymous:TB22-4-373

[2592] Anonymous. Institutional members. TUGboat, 22(4):373, December 2001. ISSN 0896-3207.

Anonymous:TB22-4-374

[2593] Anonymous. TUG membership application. TUGboat, 22(4):374, December

2001. ISSN 0896-3207. URL https://tug.org/join.html.

Anonymous:TB22-4-375

[2594] Anonymous. TEX consulting and production services. TUGboat, 22(4): 375, December 2001. ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:TB22-4-376

[2595] Anonymous. Just Published: TEX Reference Manual by David Bausum. TUGboat, 22(4):376, December 2001. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb22-4/Kluwerad.pdf.

Anonymous:2001:Cc

[2596] Anonymous. Cover. TUGboat, 22(4):c1, December 2001. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb22-4/tubcover.
pdf.

Anonymous:TB22-4-cover-3

[2597] Anonymous. Blue Sky Research. TUGboat, 22(4):Cover 3, December 2001. ISSN 0896-3207.

Christiansen:TB23-1-2

[2598] Kaja Christiansen. Editorial Comments. TUGboat, 23(1):2, 2002. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-1/kaja.pdf.

Anonymous:TB23-1-3

[2599] Anonymous. TUG 2002, Thiruvananthapuram — Report and travelogue. TUGboat, 23(1):3–7, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-1/keralaW.pdf.

Anonymous:TB23-1-8

[2600] Anonymous. TUG 2002 Program. TUGboat, 23(1):8–9, 2002. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-1/programme.pdf.

Anonymous:TB23-1-10

[2601] Anonymous. Participants at the 23^{rd} Annual TUG Meeting. TUGboat, 23 (1):10–12, 2002. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb23-1/participants.pdf.

Kumar:TB23-1-13

[2602] K. Anil Kumar. TEX and databases — TEXDBI. TUGboat, 23(1):13–16,

2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-1/
anilkumar.pdf.

Babu:TB23-1-17

[2603] Satish Babu. New horizons of free software: An Indian perspective. TUGboat, 23(1):17–20, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-1/babu.pdf.

Bujdoso:TB23-1-21

[2604] Gyöngyi Bujdosó and Ferenc Wettl. On the localization of TEX in Hungary. TUGboat, 23(1):21–26, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-1/bujdosowettl.pdf.

Bzyl:TB23-1-27

[2605] W�odzimierz Bzyl. The Tao of fonts. TUGboat, 23(1):27–40, 2002. ISSN

0896-3207. URL https://tug.org/TUGboat/tb23-1/bzyl.pdf.

Esfahbod:TB23-1-41

[2606] Behdad Esfahbod and Roozbeh Pournader. FarsiTEX and the Iranian TEX community. TUGboat, 23(1):41–45, 2002. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb23-1/farsitex.pdf.

Feng:TB23-1-46

[2607] Hong Feng. The marriage of TEX and Lojban. TUGboat, 23(1):46–48, 2002.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-1/feng.
pdf.

Hagen:TB23-1-49

[2608] Hans Hagen. ConTEXt, XML and TEX: State of the art? TUGboat, 23(1):49,

2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-1/
hagen.pdf.

Haralambous:TB23-1-50

[2609] Yannis Haralambous and John Plaice. Low-level Devanāgarī support for Omega — Adapting devnag. TUGboat, 23(1):50–56, 2002. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb23-1/haralambous.pdf.

Kastrup:TB23-1-57

[2610] David Kastrup. Revisiting WYSIWYG paradigms for authoring L^{A}TEX. TUGboat, 23(1):57–64, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-1/kastrup.pdf.

Moore:TB23-1-65

[2611] Ross Moore. serendiPDF with searchable math-fields in PDF documents. TUGboat, 23(1):65–69, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-1/moore.pdf.

Piska:TB23-1-70

[2612] Karel Píška. A conversion of public Indic fonts from into Type 1 format with TEXtrace. TUGboat, 23(1):70–73, 2002. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb23-1/piska.pdf.

Popineau:TB23-1-74

[2613] Fabrice Popineau. TEX Live under Windows: What’s new with the 7^{th} edition? TUGboat, 23(1):74–79, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-1/popineau.pdf.

Pournader:TB23-1-80

[2614] Roozbeh Pournader. Catching up to Unicode. TUGboat, 23(1):80–85,

2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-1/
roozbeh.pdf.

Rahtz:TB23-1-86

[2615] Sebastian Rahtz. PassiveTEX: An update. TUGboat, 23(1):86–89, 2002. ISSN

0896-3207. URL https://tug.org/TUGboat/tb23-1/rahtz.pdf.

Rajkumar:TB23-1-90

[2616] S. Rajkumar. Indic typesetting — Challenges and opportunities. TUGboat, 23 (1):90–92, 2002. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb23-1/rajkumar.pdf.

Roegel:TB23-1-93

[2617] Denis Roegel. METAOBJ: Very high-level objects in . TUGboat, 23(1):93–100, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-1/roegel.pdf.

Shukla:TB23-1-101

[2618] Wagish Shukla and Amitabh Trehan. Typesetting in Hindi, Sanskrit and Persian: A beginner’s perspective. TUGboat, 23(1):101–105, 2002. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-1/trehan.pdf.

Skoupy:TB23-1-106

[2619] Karel Skoupý. New typesetting language and system architecture. TUGboat, 23(1):106, 2002. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb23-1/skoupyI.pdf.

Skoupy:TB23-1-107

[2620] Karel Skoupý. TEX file server. TUGboat, 23(1):107, 2002. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb23-1/skoupyII.pdf.

Watt:TB23-1-108

[2621] Stephen M. Watt. Conserving implicit mathematical semantics in conversion between TEX and MathML. TUGboat, 23(1):108, 2002. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb23-1/watt.pdf.

Anonymous:TB23-1-109

[2622] Anonymous. Calendar. TUGboat, 23(1):109–110, 2002. ISSN 0896-3207.

Anonymous:TB23-1-110

[2623] Anonymous. Institutional members. TUGboat, 23(1):110, 2002. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-1/inst.pdf.

Anonymous:TB23-1-111

[2624] Anonymous. TEX consulting and production services. TUGboat, 23(1):111, 2002. ISSN 0896-3207.

Anonymous:TB23-1-112

[2625] Anonymous. Just Published: TEX Reference Manual by David Bausum. TUGboat, 23(1):112, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-1/kluwer.pdf.

Anonymous:2002:BSR

[2626] Anonymous. Blue Sky Research. TUGboat, 23(1):c3, March 2002. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-1/
cover3.pdf.

Anonymous:2002:C

[2627] Anonymous. Cover. TUGboat, 23(1):c1, March 2002. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb23-1/cover1.
pdf.

Anonymous:2002:IFCa

[2628] Anonymous. Inside front cover. TUGboat, 23(1):c2, March 2002. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-1/
cover2.pdf.

Anonymous:TB23-1-cover-3

[2629] Anonymous. Blue Sky Research. TUGboat, 23(1):Cover 3, 2002. ISSN 0896-3207.

Anonymous:2002:TPa

[2630] Anonymous. Title page. TUGboat, 23(2):113, 2002. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-2/titlepage.pdf.

Beeton:TB23-2-114

[2631] Barbara Beeton. Editorial Comments. TUGboat, 23(2):114, 2002. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-2/tb74beet.pdf.

Flynn:TB23-2-115

[2632] Peter Flynn. Formatting information: A beginner’s introduction to typesetting with L^{A}TEX. TUGboat, 23(2):115–237, 2002. ISSN 0896-3207.

URL https://ctan.org/pkg/beginlatex/;   https://www.
ctan.org/tex-archive/info/beginlatex/.

Anonymous:TB23-2-238

[2633] Anonymous. Calendar. TUGboat, 23(2):238–239, 2002. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB23-2-239

[2634] Anonymous. Institutional members. TUGboat, 23(2):239, 2002. ISSN 0896-3207.

Anonymous:TB23-2-240

[2635] Anonymous. TEX consulting and production services. TUGboat, 23(2):240,

2002. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2002:FCa

[2636] Anonymous. Front cover. TUGboat, 23(2):c1, 2002. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-2/cover1.pdf.

Anonymous:2002:IFCb

[2637] Anonymous. Inside front cover. TUGboat, 23(2):c2, 2002. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-2/
cover2.pdf.

Anonymous:TB23-2-c3

[2638] Anonymous. TUG 2003 Announcement. TUGboat, 23(2):c 3–c 3, 2002. ISSN 0896-3207.

Anonymous:2002:TA

[2639] Anonymous. TUG2004 announcement. TUGboat, 23(2):c3, 2002. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-2/
cover3.pdf.

Anonymous:2002:TPb

[2640] Anonymous. Title page. TUGboat, 23(3–4):1, 2002. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-3-4/titlepage.
pdf.

Anonymous:2002:FCb

[2641] Anonymous. Front cover. TUGboat, 23(3–4):c1, 2002. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb23-3-4/cover1.
pdf.

Anonymous:2002:IBC

[2642] Anonymous. Inside back cover: TUG 2004. TUGboat, 23(3–4):c3, 2002. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb23-3-4/cover3.pdf.

Anonymous:2002:IFCc

[2643] Anonymous. Inside front cover. TUGboat, 23(3–4):c2, 2002. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-
4/cover2.pdf.

Anonymous:TB23-3-243

[2644] Anonymous. Addresses. TUGboat, 23(3/4):243–244, 2002. ISSN 0896-3207.

Preston:TB23-3-244

[2645] Roy Preston. Type Design as Art. TUGboat, 23(3/4):244, 2002. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-3-4/grot.pdf.

Berry:TB23-3-245

[2646] Karl Berry. From the Board of Directors. TUGboat, 23(3/4):245, 2002.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-4/
tb75gendel.pdf.

Beeton:TB23-3-245

[2647] Barbara Beeton. Editorial comments. TUGboat, 23(3/4):245–247, 2002.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-4/
tb75beet.pdf.

Anonymous:TB23-3-247

[2648] Anonymous. Hyphenation exception log. TUGboat, 23(3/4):247–248,

2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-
4/tb75hyf.pdf.

Anonymous:TB23-3-249

[2649] Anonymous. Donald Knuth: All questions answered (University of Oslo, 30 August 2002). TUGboat, 23(3/4):249–261, 2002. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb23-3-4/tb75knuth.pdf.

Beccari:TB23-3-261

[2650] Claudio Beccari. Floating point numbers and  ,  ,

TEX, and PostScript Type 1 fonts. TUGboat, 23(3/4):261–269, 2002.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-4/
tb75beccreal.pdf.

Revets:TB23-3-269

[2651] Stefan A. Revets. The Octavo package. TUGboat, 23(3/4):269–275,

2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-
4/tb75revets.pdf.

Beccari:TB23-3-276

[2652] Claudio Beccari. The teubner L^{A}TEX package: Typesetting classical Greek philology. TUGboat, 23(3/4):276–282, 2002. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb23-3-4/tb75beccteub.pdf.

Pal:TB23-3-282

[2653] Palash B. Pal. Typesetting in Bengali script using TEX. TUGboat, 23(3/4): 282–288, 2002. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb23-3-4/tb75pal.pdf.

Kuhn:TB23-3-288

[2654] Jonathan Kuhn. interactiveworkbook: L^{A}TEX-based interactive PDF on the Web. TUGboat, 23(3/4):288–290, 2002. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb23-3-4/tb75kuhn.pdf.

Menshikov:TB23-3-291

[2655] D. Men’shikov, A. Kostin, and M. Vulis. Multiple Master math extension fonts. TUGboat, 23(3/4):291–294, 2002. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb23-3-4/tb75vulis.pdf.

Barton:TB23-3-294

[2656] Thomas H. Barton. TrueType fonts in PostScript. TUGboat, 23(3/4):294–

296, 2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-
3-4/tb75bart.pdf.

Tsolomitis:TB23-3-296

[2657] Antonis Tsolomitis. The Kerkis font family. TUGboat, 23(3/4):296–301,

2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-
4/tb75tsol.pdf.

Schmidt:TB23-3-301

[2658] Walter Schmidt. Euler-VM: Generic math fonts for use with L^{A}TEX. TUGboat, 23(3/4):301–303, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-3-4/tb75schmidt.pdf.

Saha:TB23-3-304

[2659] Prasenjit Saha. Rambutan: Literate programming in Java. TUGboat, 23(3/4): 304–309, 2002. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb23-3-4/tb75saha.pdf.

Obrecht:TB23-3-309

[2660] Christian Obrecht. Eukleides: A geometry drawing language. TUGboat, 23(3/4):309–313, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-3-4/tb75obre.pdf.

Casares:TB23-3-313

[2661] Ramón Casares. MTEX. TUGboat, 23(3/4):313–318, 2002. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb23-3-4/tb75casa.pdf.

LaPlante:TB23-3-319

[2662] Mark LaPlante and William F. Adams. The treasure chest. TUGboat, 23(3/4): 319–329, 2002. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb23-3-4/tb75chest.pdf.

Feuerstack:TB23-3-329

[2663] Thomas Feuerstack. Introduction to pdfTEX. TUGboat, 23(3/4):329–334,

2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-
4/tb75feu.pdf.

Voss:TB23-3-335

[2664] Herbert Voß. Constructing circuit diagrams with pst-circ. TUGboat, 23(3/4): 335–341, 2002. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb23-3-4/tb75voss.pdf.

Gray:TB23-3-341

[2665] Norman Gray. Absolute positioning with textpos. TUGboat, 23(3/4):341–344,

2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-
4/tb75gray.pdf.

Harders:TB23-3-344

[2666] Harald Harders. Multilingual bibliographies: Using and extending the babelbib package. TUGboat, 23(3/4):344–353, 2002. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb23-3-4/tb75harders.pdf.

Anonymous:TB23-3-354

[2667] Anonymous. Les Cahiers GUTenberg: contents of Thematic Issue 41 (November 2001). TUGboat, 23(3/4):354, 2002. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb23-3-4/tb75cahi.pdf.

Anonymous:TB23-3-355

[2668] Anonymous. Calendar. TUGboat, 23(3/4):355–357, 2002. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB23-3-357

[2669] Anonymous. Institutional members. TUGboat, 23(3/4):357, 2002. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-3-4/instmem-
list.pdf.

Anonymous:TB23-3-358

[2670] Anonymous. Practical TEX 2004: training and techniques. TUGboat, 23(3/4):358, 2002. ISSN 0896-3207. URL https://www.tug.org/

practicaltex2004.

Anonymous:TB23-3-359

[2671] Anonymous. TEX Live, 2003 Edition. TUGboat, 23(3/4):359, 2002.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-4/
texlive.pdf.

Anonymous:TB23-3-360

[2672] Anonymous. Future issues. TUGboat, 23(3/4):360, 2002. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb23-3-4/tb75boar.pdf.

Burbank:TB23-3-360

[2673] Mimi Burbank. Production notes. TUGboat, 23(3/4):360, 2002. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb23-3-4/tb75prod.
pdf.

Ogawa:TB23-3-360

[2674] Arthur Ogawa. Report of TUG election. TUGboat, 23(3/4):360–365,

2002. ISSN 0896-3207. URL https://tug.org/TUGboat/tb23-3-
4/tb75boar.pdf.

Laakso:TB23-3-365

[2675] Robin Laakso. Financial statements for 2001 and 2002. TUGboat, 23(3/4): 365–366, 2002. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb23-3-4/tb75treas.pdf.

Anonymous:TB23-3-367

[2676] Anonymous. Bugs in Computers & Typesetting, 6 July 2003. TUGboat, 23(3/4):367–374, 2002. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb23-3-4/tb75errata.pdf.

Anonymous:TB23-3-375

[2677] Anonymous. TUG membership application. TUGboat, 23(3/4):375,

2002. ISSN 0896-3207. URL https://www.tug.org/forms/2004/
memberapp.html.

Anonymous:TB23-3-376

[2678] Anonymous. TEX consulting and production services. TUGboat, 23(3/4):376,

2002. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB23-3-cover-3

[2679] Anonymous. TUG 2004 announcement. TUGboat, 23(3/4):Cover 3, 2002. ISSN 0896-3207.

Anonymous:2003:TPa

[2680] Anonymous. Title page. TUGboat, 24(1):1, 2003. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-1/titlepage.pdf.

Adams:TB24-1-2

[2681] William Adams. Editorial remarks. TUGboat, 24(1):2, 2003. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb24-1/adams.pdf.

Berry:TB24-1-3

[2682] Karl Berry. Welcome to TUG 2003. TUGboat, 24(1):3–4, 2003. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-1/berry.pdf.

Anonymous:TB24-1-5

[2683] Anonymous. Conference program and delegates. TUGboat, 24(1):5–9,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-1/
confinfo.pdf.

Rosell-Gonzalez:TB24-1-10

[2684] Pablo Rosell-González. Return to the classics. TUGboat, 24(1):10–17,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-1/
gonzalez.pdf.

Moon:TB24-1-18

[2685] Alun Moon. Digital illumination. TUGboat, 24(1):18–22, 2003. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-1/moon-celtic.
pdf.

Anonymous:TB24-1-23

[2686] Anonymous. Abstracts (Daly & Richter, Hagen, Janc, Moore). TUGboat, 24 (1):23–24, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb24-1/abs-publish.pdf.

Patashnik:TB24-1-25

[2687] Oren Patashnik. BibTEX yesterday, today, and tomorrow. TUGboat, 24(1):25–

30, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
1/patashnik.pdf.

Anonymous:TB24-1-31

[2688] Anonymous. Abstracts (Daly, Hufflen). TUGboat, 24(1):31, 2003. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-1/abs-biblio.
pdf.

Koch:TB24-1-32

[2689] Richard Koch. TeXShop in 2003. TUGboat, 24(1):32–38, 2003. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-1/koch.pdf.

Costanzo:TB24-1-39

[2690] Francesco Costanzo and Gary Gray. Creating labeled stand-alone figures in L^{A}TEX using WARMreader and Adobe Illustrator under Mac OS X. TUGboat, 24(1):39–49, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-1/gray-warm.pdf.

Saastamoinen:TB24-1-50

[2691] Kalle Saastamoinen, Jaakko Ketola, Tuukka Kurppa, and Liisa Torikka. Enabling Web access to a database of calculus problems using L^{A}TEX, PHP and L^{A}TEX2HTML. TUGboat, 24(1):50–52, 2003. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb24-1/saast.pdf.

Padovani:TB24-1-53

[2692] Luca Padovani. MathML formatting with TEX rules and TEX fonts. TUGboat, 24(1):53–61, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-1/padovani.pdf.

Anonymous:TB24-1-62

[2693] Anonymous. Abstracts (Gurari, Quirk, Hagen, Hefferon, Lehmke). TUGboat, 24(1):62–63, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-1/abs-inter.pdf.

Jackowski:TB24-1-64

[2694] Bogus�aw Jackowski and Janusz M. Nowacki. Latin Modern: Enhancing Computer Modern with accents, accents, accents. TUGboat, 24(1):64–74,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-1/
jackowski.pdf.

Volovich:TB24-1-75

[2695] Vladimir Volovich. CM-Super: Automatic creation of efficient Type 1 fonts from fonts. TUGboat, 24(1):75–78, 2003. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb24-1/volovich.pdf.

Thanh:TB24-1-79

[2696] Hàn Thê´ Thành. Making Type 1 fonts for Vietnamese. TUGboat, 24(1):79–84,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-1/
thanh.pdf.

Yiu:TB24-1-85

[2697] Candy L. K. Yiu and Wai Wong. Chinese character synthesis using . TUGboat, 24(1):85–93, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-1/yiu.pdf.

Anonymous:TB24-1-94

[2698] Anonymous. Abstracts (Adams). TUGboat, 24(1):94, 2003. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb24-1/abs-fonts.pdf.

Bilotta:TB24-1-95

[2699] Giuseppe Bilotta. eOmega: A step towards the future with a look at the past. TUGboat, 24(1):95–96, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-1/bilotta-eomega.pdf.

Hoenig:TB24-1-97

[2700] Alan Hoenig. Makor: Typesetting Hebrew with Omega. TUGboat, 24(1): 97–104, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb24-1/hoenig.pdf.

Plaice:TB24-1-105

[2701] John Plaice, Paul Swoboda, Yannis Haralambous, and Chris Rowley. A multidimensional approach to typesetting. TUGboat, 24(1):105–114, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-1/plaice.
pdf.

Anonymous:TB24-1-115

[2702] Anonymous. Abstracts (Cho). TUGboat, 24(1):115, 2003. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb24-1/abs-multi.pdf.

Moon:TB24-1-116

[2703] Alun Moon. Literate programming meets UML. TUGboat, 24(1):116–119,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-1/
moon-uml.pdf.

Bilotta:TB24-1-120

[2704] Giuseppe Bilotta. Math in ConTEXt: Bridging the gap with (AMS-)L^{A}TEX. TUGboat, 24(1):120–121, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-1/bilotta-context.pdf.

Anonymous:TB24-1-122

[2705] Anonymous. Abstracts (Bazargan et al., Esser, Lehmke, Popineau, Wierda). TUGboat, 24(1):122–123, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-1/abs-tools.pdf.

Gray:TB24-1-124

[2706] Gary Gray and Francesco Costanzo. Experiences and lessons learned teaching L^{A}TEX to university students. TUGboat, 24(1):124–131, 2003. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-1/gray-class.
pdf.

Gaudeul:TB24-1-132

[2707] Alexandre Gaudeul. The (La)TEX project: A case study of open source software. TUGboat, 24(1):132–145, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-1/gaudeul.pdf.

Odyniec:TB24-1-146

[2708] Andrzej Odyniec. A Polished TEX story. TUGboat, 24(1):146–150,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-1/
odyniec.pdf.

Anonymous:TB24-1-151

[2709] Anonymous. Abstracts (Bibby, Ranade). TUGboat, 24(1):151, 2003.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-1/abs-
socio.pdf.

Anonymous:TB24-1-152

[2710] Anonymous. Calendar. TUGboat, 24(1):152–153, 2003. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:2003:LCE

[2711] Anonymous. Book review: The L^{A}TEX Companion, 2nd edition, by Frank Mittelbach et al. TUGboat, 24(1):153, 2003. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/books/#companion.

Anonymous:TB24-1-154

[2712] Anonymous. TUG 2005 conference, Wuhan, China, August 23–25, 2005. TUGboat, 24(1):154, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/books/#guidelatex.

Anonymous:TB24-1-155

[2713] Anonymous. Recognition of support from Apple. TUGboat, 24(1):155,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-1/
appleinst.pdf.

Anonymous:TB24-1-155-2

[2714] Anonymous. Institutional members. TUGboat, 24(1):155, 2003. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-1/appleinst.pdf.

Anonymous:TB24-1-156

[2715] Anonymous. TEX consulting and production services. TUGboat, 24(1):156,

2003. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2003:FCa

[2716] Anonymous. Front cover. TUGboat, 24(1):c1, 2003. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-1/cover1.pdf.

Anonymous:2003:IBC

[2717] Anonymous. Inside back cover: TUG 2004. TUGboat, 24(1):c3, 2003. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
1/cover3.pdf.

Anonymous:2003:IFCa

[2718] Anonymous. Inside front cover. TUGboat, 24(1):c2, 2003. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-1/
cover2.pdf.

Anonymous:TB24-1-c3

[2719] Anonymous. TUG 2004 conference, Xanthi, Greece, August 30–September 3, 2004. TUGboat, 24(1):c 3–c 3, 2003. ISSN 0896-3207.

Anonymous:2003:TPb

[2720] Anonymous. Title page. TUGboat, 24(2):121, 2003. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-2/titlepage.pdf.

Berry:TB24-2-159

Beeton:TB24-2-160

[2722] Barbara Beeton. Editorial comments. TUGboat, 24(2):160–161, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-2/tb77beet.
pdf.

Plaice:TB24-2-162

[2723] John Plaice. Daniel Taupin, 1936–2003. TUGboat, 24(2):162, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-2/tb77plaice.
pdf.

Waud:TB24-2-163

[2724] Douglas Waud. What is TEX? TUGboat, 24(2):163–164, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-2/tb77waud-
what.pdf.

Lawrence:TB24-2-165

[2725] Richard Lawrence. Maths = Typography? TUGboat, 24(2):165–168,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-2/
tb77lawrence.pdf.

Garcia:TB24-2-169

[2726] Federico Garcia. On musical typesetting: Sonata for TEX and , Op. 2. TUGboat, 24(2):169–182, 2003. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb24-2/tb77garcia.pdf.

Adams:TB24-2-183

[2727] William Adams. There is no end: Omega and Zapfino. TUGboat, 24(2):183–

199, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
2/tb77adams.pdf.

Hall:TB24-2-200

[2728] Timothy Hall. The approach: Implicit, relative, and analytical font design. TUGboat, 24(2):200–205, 2003. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb24-2/tb77hall.pdf.

Fairbairns:TB24-2-205

[2729] Robin Fairbairns, Jim Hefferon, Rainer Schöpf, Joachim Schrod, Graham Williams, and Reinhard Zierke. CTAN plans. TUGboat, 24(2):205–208,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-2/
tb77heff.pdf.

Hellstrom:TB24-2-208

[2730] Lars Hellström. Some notes on templates. TUGboat, 24(2):208–210,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-2/
tb77hellstrom.pdf.

Walden:TB24-2-211

[2731] David Walden. Writing a big book — A first experience with L^{A}TEX. TUGboat, 24(2):211–215, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-2/tb77walden.pdf.

Syropoulos:TB24-2-216

[2732] Apostolos Syropoulos. Designing packages for Λ: An overview. TUGboat, 24(2):216–220, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-2/tb77syropoulos.pdf.

LPT:TB24-2-221

[2733] L^{A}TEX Project Team. L^{A}TEX news, issues 14–16, 2001–2003. TUGboat, 24(2): 221–223, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb24-2/tb77ltnews.pdf.

Luck:TB24-2-224

[2734] Uwe Lück. ednotes — critical edition typesetting with L^{A}TEX. TUGboat, 24(2):224–236, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-2/tb77lueck.pdf.

Scannell:TB24-2-236

[2735] Kevin P. Scannell. Hyphenation patterns for minority languages. TUGboat, 24(2):236–239, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-2/tb77scannell.pdf.

Caird:TB24-2-240

[2736] Andrew Caird. (La)TEX, genealogy, and the LifeLines software. TUGboat, 24(2):240–244, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-2/tb77caird.pdf.

Talole:TB24-2-245

[2737] S. E. Talole and S. B. Phadke. Generating L^{A}TEX documents through Matlab. TUGboat, 24(2):245–248, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-2/tb77seta.pdf.

Hufflen:TB24-2-249

[2738] Jean-Michel Hufflen. MlBibTEX’s Version 1.3. TUGboat, 24(2):249–262,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-2/
tb77hufflen.pdf.

Wilson:TB24-2-262

[2739] Peter Wilson. Glisterings. TUGboat, 24(2):262–265, 2003. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb24-2/tb77wilson.pdf.

LaPlante:TB24-2-265

[2740] Mark LaPlante. The treasure chest. TUGboat, 24(2):265–275, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-2/tb77chest.
pdf.

Beccari:TB24-2-275

[2741] Claudio Beccari. The L^{A}TEX Companion, Second Edition. TUGboat, 24(2): 275–277, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb24-2/tb77becc.pdf.

Waud:TB24-2-278

[2742] Douglas Waud and Mimi Burbank. Guide to L^{A}TEX, 4^{th} Edition. TUGboat, 24(2):278–281, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-2/tb77waud-kopka.pdf.

Anonymous:TB24-2-282

[2743] Anonymous. Les Cahiers GUTenberg: contents of Issue 42 (July 2003). TUGboat, 24(2):282, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-2/tb77cahi.pdf.

Anonymous:TB24-2-283

[2744] Anonymous. MAPS: contents of issues 27–28 (2002). TUGboat, 24(2):283–

285, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
2/tb77maps.pdf.

Anonymous:TB24-2-286

[2745] Anonymous. TEXemplares: contents of issues 4–6 (2003–04). TUGboat, 24(2): 286–287, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb24-2/tb77texemp.pdf.

Katre:TB24-2-288

[2746] S. A. Katre and Manjusha Joshi. Report on the Pune workshop on L^{A}TEX and free mathematical software. TUGboat, 24(2):288–290, 2003. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb24-2/tb77pune.pdf.

Beebe:TB24-2-291

[2747] Nelson Beebe, Wendy McKay, and Ross Moore. TUG at Bay. TUGboat, 24(2): 291–293, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb24-2/tb77tfaa.pdf.

Anonymous:TB24-2-293

Anonymous:TB24-2-294

[2749] Anonymous. Calendar. TUGboat, 24(2):294–295, 2003. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb24-2/tb77calendar.pdf.

Anonymous:TB24-2-295

[2750] Anonymous. TUG 2005 announcement. TUGboat, 24(2):295, 2003. ISSN

0896-3207. URL https://www.tug.org/tug2005/.

Anonymous:TB24-2-296

[2751] Anonymous. Future issues. TUGboat, 24(2):296, 2003. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb24-2/tb77treas.pdf.

Burbank:TB24-2-296

[2752] Mimi Burbank. Production notes. TUGboat, 24(2):296, 2003. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-2/tb77prod.pdf.

Laakso:TB24-2-296

[2753] Robin Laakso. Financial statements for 2003. TUGboat, 24(2):296–297,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-2/
tb77treas.pdf.

Anonymous:TB24-2-298

[2754] Anonymous. TUG 2005 election. TUGboat, 24(2):298, 2003. ISSN 0896-

3207. URL https://www.tug.org/election/.

Anonymous:TB24-2-299

[2755] Anonymous. TEX consulting and production services. TUGboat, 24(2):299,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-2/
tb77tlc2.pdf.

Anonymous:2003:ETK

[2756] Anonymous. Easy Table, Khanh Ha. TUGboat, 24(2):300, 2003. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-2/
tb77eztable.pdf.

Anonymous:2003:CODa

[2757] Anonymous. Contents ordered by difficulty. TUGboat, 24(2):c3, 2003. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
2/cover3.pdf.

Anonymous:2003:FCb

Anonymous:2003:IFCb

[2759] Anonymous. Inside front cover. TUGboat, 24(2):c2, 2003. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-2/
cover2.pdf.

Anonymous:TB24-2-cover-3

[2760] Anonymous. EuroTEX 2005 announcement. TUGboat, 24(2):Cover 3, 2003. ISSN 0896-3207.

Anonymous:2003:TPc

[2761] Anonymous. Title page. TUGboat, 24(3):121, 2003. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-3/titlepage.pdf.

Haralambous:TB24-3-302

[2762] Yannis Haralambous. The EuroTEX 2003 conference. TUGboat, 24(3):302–

305, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
3/edit.pdf.

Anonymous:TB24-3-306

[2763] Anonymous. Conference program and delegates. TUGboat, 24(3):306–310,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-3/
confinfo.pdf.

Paput:TB24-3-311

[2764] Christian Paput. French typographic patrimony, conservation and teaching. TUGboat, 24(3):311–313, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-3/paput.pdf.

Andre:TB24-3-314

[2765] Jacques André. The Cassetin project — towards an inventory of ancient types and the related standardised encoding. TUGboat, 24(3):314–318, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-3/andre.pdf.

Syropoulos:TB24-3-319

[2766] Apostolos Syropoulos. Replicating archaic documents: A typographic challenge. TUGboat, 24(3):319–322, 2003. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb24-3/syropoulos.pdf.

Lazrek:TB24-3-323

[2767] Azzeddine Lazrek. CurExt, typesetting variable-sized curved symbols. TUGboat, 24(3):323–327, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-3/lazrek.pdf.

Atanasiu:TB24-3-328

[2768] Vlad Atanasiu. Allographic biometrics and behavior synthesis. TUGboat, 24(3):328–333, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-3/atanasiu.pdf.

Mourad:TB24-3-334

[2769] Ghassan Mourad. La virgule viendrait-elle de l’écriture arabe ? (French) [Did the comma come from Arabic writing?]. TUGboat, 24(3):334–338,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-3/
mourad2.pdf.

Souchier:TB24-3-339

[2770] Emmanuël Souchier. Quelques remarques sur le sens et la servitude de la typographie. TUGboat, 24(3):339–350, 2003. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb24-3/PDF/souchier.pdf.

Maniette:TB24-3-351

[2771] Yves Maniette. Système automatisé de co-rédaction de livres. TUGboat, 24(3): 351–356, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb24-3/maniette.pdf.

Turcan:TB24-3-357

[2772] Isabelle Turcan and Viviane Berthelier. Éthique et édition scientifique d’ouvrages anciens sur support électronique: choix de traitements numériques des ornements et marques typographiques dans le programme de numérisation des collections d ouvrages anciens du laboratoire ATILF. TUGboat, 24(3): 357–368, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb24-3/PDF/turcan.pdf.

Haralambous:TB24-3-369

[2773] Yannis Haralambous and John Plaice. XL^{A}TEX, a DTD/schema which is very close to L^{A}TEX. TUGboat, 24(3):369–376, 2003. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb24-3/haralambous.pdf.

Grimm:TB24-3-377

[2774] José Grimm. Tralics, a L^{A}TEX to XML translator. TUGboat, 24(3):377–388,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-3/
grimm.pdf.

Pepping:TB24-3-389

[2775] Simon Pepping. Docbook In ConTEXt, a ConTEXt XML mapping for DocBook documents. TUGboat, 24(3):389–401, 2003. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb24-3/pepping.pdf.

Kanellos:TB24-3-402

[2776] Ioannis Kanellos. Intertextualité et typographie numérique — considérations sémantiques sur le livre électronique. TUGboat, 24(3):402–410, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-3/kanellos.
pdf.

Mourad:TB24-3-411

[2777] Ghassan Mourad. Nouveaux signes de lecture et d’écriture pour les documents électroniques. TUGboat, 24(3):411–414, 2003. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb24-3/mourad1.pdf.

Chaix:TB24-3-415

[2778] Marie-Louise Chaix and Fabrice Popineau. The XEMTEX project. TUGboat, 24(3):415–419, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-3/popineau.pdf.

Laurens:TB24-3-420

[2779] Jérôme Laurens. iTEXMac, an integrated TEX environment for Mac OS X. TUGboat, 24(3):420–426, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-3/laurens.pdf.

Vecsei:TB24-3-427

[2780] Balázs Vecsei. Description of knowledge of mathematical programs with TEX and XML. TUGboat, 24(3):427–429, 2003. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb24-3/balazs.pdf.

Turner:TB24-3-430

[2781] David Turner and Werner Lemberg. Real-time grid fitting of typographic outlines. TUGboat, 24(3):430–434, 2003. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb24-3/lemberg.pdf.

Sutto:TB24-3-435

[2782] Jean-Pierre Sutto and Pier Daniele Napolitani. L’utilisation du Mauro-TEX pour l’édition critique de Francesco Maurolico. TUGboat, 24(3):435–440,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-3/
napolitani.pdf.

Szabo:TB24-3-441

[2783] Péter Szabó. Inserting external figures with GraphicP. TUGboat, 24(3):441–

448, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
3/szabo.pdf.

Horak:TB24-3-449

[2784] Karel Horák. Geometric diversions with TEX,  and  .

TUGboat, 24(3):449–452, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-3/horak.pdf.

Boulanger:TB24-3-453

[2785] Frédéric Boulanger. Printing digital photographs with L^{A}TEX. TUGboat, 24(3): 453–461, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb24-3/boulanger.pdf.

Kastrup:TB24-3-462

[2786] David Kastrup. Output routine requirements for advanced typesetting tasks. TUGboat, 24(3):462–467, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-3/kastrup.pdf.

Widmann:TB24-3-468

[2787] Thomas Widmann. Bibulus — a Perl/XML replacement for BibTeX. TUGboat, 24(3):468–471, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-3/widmann.pdf.

Dagnat:TB24-3-472

[2788] Fabien Dagnat, Ronan Keryell, Laura Barrero Sastre, Emmanuel Donin de Rosière, and Nicolas Torneri. BibTEXc++: Toward higher-order BibTEXing. TUGboat, 24(3):472–488, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-3/keryell.pdf.

Hufflen:TB24-3-489

[2789] Jean-Michel Hufflen. European bibliography styles and MlBibTeX. TUGboat, 24(3):489–498, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-3/hufflen.pdf.

Olsak:TB24-3-499

[2790] Petr Olšák. Second version of encTEX: UTF-8 support. TUGboat, 24(3):499–

501, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
3/olsak.pdf.

Milo:TB24-3-502

[2791] Thomas Milo. ALI-BABA and the 4.0 Unicode characters — New input and output concepts under Unicode. TUGboat, 24(3):502–511, 2003. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-3/milo.pdf.

Plaice:TB24-3-512

[2792] John Plaice and Yannis Haralambous. Generating multiple outputs from OMEGA. TUGboat, 24(3):512–518, 2003. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb24-3/plaice.pdf.

Sastry:TB24-3-519

[2793] B. V. Venkata Krishna Sastry. Enhanced font features for future multilingual digital typography with sound-script-language attribute integration. TUGboat, 24(3):519–526, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-3/sastry.pdf.

Bujdoso:TB24-3-527

[2794] Gyöngyi Bujdosó. Contemporary Hungarian types and designers. TUGboat, 24(3):527–530, 2003. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb24-3/bujdoso.pdf.

Williams:TB24-3-531

[2795] George Williams. Font creation with FontForge. TUGboat, 24(3):531–544,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-3/
williams.pdf.

Peterlin:TB24-3-545

[2796] Primož Peterlin. The free UCS outline fonts project — An attempt to create a global font. TUGboat, 24(3):545–549, 2003. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb24-3/peterlin.pdf.

Mehta:TB24-3-550

[2797] Anish Mehta, Gábor Bella, and Yannis Haralambous. Adapting OMEGA to OpenType fonts. TUGboat, 24(3):550–556, 2003. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb24-3/bella.pdf.

Toledo:TB24-3-557

[2798] Sivan Toledo and Zvika Rosenberg. Experience with OpenType Font Production. TUGboat, 24(3):557–568, 2003. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb24-3/toledo.pdf.

Vakulenko:TB24-3-569

[2799] Serge Vakulenko. The METATYPE project: Creating TrueType fonts based on . TUGboat, 24(3):569–574, 2003. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb24-3/vakulenko.pdf.

Jackowski:TB24-3-575

[2800] Bogus�aw Jackowski, Janusz Nowacki, and Piotr Strzelczyk. Programming PostScript Type 1 fonts using MetaType1: Auditing, enhancing, creating. TUGboat, 24(3):575–581, 2003. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb24-3/jackowski.pdf.

Wong:TB24-3-582

[2801] Wai Wong, Candy L. K. Yiu, and Kelvin C. F. Ng. Typesetting rare Chinese characters in L^{A}TEX. TUGboat, 24(3):582–587, 2003. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb24-3/wong.pdf.

Devroye:TB24-3-588

[2802] Luc Devroye. Formatting font formats. TUGboat, 24(3):588–596, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-3/devroye.
pdf.

Tombeur:TB24-3-597

[2803] Jef Tombeur. Polices d’apprentissage de l’écriture. TUGboat, 24(3):597–604,

2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-3/
tombeur1.pdf.

Tombeur:TB24-3-605

[2804] Jef Tombeur. Réformes philanthropiques et réformes orthotypographiques, alphabets artificiels et synthétiques. TUGboat, 24(3):605–618, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-3/tombeur2.
pdf.

Anonymous:TB24-3-619

[2805] Anonymous. Calendar. TUGboat, 24(3):619–620, 2003. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB24-3-621

[2806] Anonymous. TUG 2005, 23–25 August 2005, Wuhan, China. TUGboat, 24(3):

621, 2003. ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
3/tb78elec.pdf.

Anonymous:TB24-3-621-2

[2807] Anonymous. TUG 2005 election report. TUGboat, 24(3):621, 2003. ISSN

0896-3207. URL https://tug.org/TUGboat/tb24-3/tb78elec.
pdf.

Anonymous:TB24-3-622

[2808] Anonymous. Practical TEX 2005, 14–17 June 2005, Chapel Hill, North Carolina. TUGboat, 24(3):622, 2003. ISSN 0896-3207. URL https://www.

tug.org/practicaltex2005/.

Anonymous:TB24-3-623

[2809] Anonymous. Institutional members. TUGboat, 24(3):623, 2003. ISSN 0896-

3207. URL https://tug.org/consultants.html.

Anonymous:TB24-3-623-2

[2810] Anonymous. TEX consulting and production services. TUGboat, 24(3):623,

2003. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB24-3-624

[2811] Anonymous. Easy Table, Khanh Ha. TUGboat, 24(3):624, 2003. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-3/ezad.pdf.

Anonymous:2003:CODb

[2812] Anonymous. Contents ordered by difficulty. TUGboat, 24(3):c3, 2003. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-
3/cover3.pdf.

Anonymous:2003:FCc

[2813] Anonymous. Front cover. TUGboat, 24(3):c1, 2003. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb24-3/cover1.pdf.

Anonymous:2003:IFCc

[2814] Anonymous. Inside front cover. TUGboat, 24(3):c2, 2003. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb24-3/
cover2.pdf.

Anonymous:TB24-3-c3

[2815] Anonymous. H. W. Caslon & Co. Ltd. TUGboat, 24(3):c 3–c 3, 2003. ISSN 0896-3207.

Berry:2004:ER

[2816] Karl Berry. Editorial remarks. TUGboat, 25(0):2, August 2004. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/Pres/
tb79pres.html.

Loverdos:2004:DTN

[2817] Christos KK Loverdos and Apostolos Syropoulos. Digital typography in the new millennium: Flexible documents by a flexible engine. TUGboat, 25(0):3, August 2004. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-0/loverdos.pdf.

Bzyl:2004:MXC

[2818] W�odzimierz Bzyl and Tomasz Przechlewski. Migrating to XML: The case of the GUST Bulletin archive. TUGboat, 25(0):13, August 2004. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-0/bzyl.
pdf.

Przechlewski:2004:MTR

[2819] Tomasz Przechlewski. Managing TEX resources with XML topic maps. TUGboat, 25(0):19, August 2004. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb25-0/przech.pdf.

Padovani:2004:IEM

[2820] Luca Padovani. Interactive editing of MathML markup using TEX syntax. TUGboat, 25(0):26, August 2004. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb25-0/padovani.pdf.

Holecek:2004:APG

[2821] Jan Holecek and Petr Sojka. Animations in pdfTEX-generated PDF. TUGboat, 25(0):35, August 2004. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb25-0/holecek.pdf.

Eddahibi:2004:AMD

[2822] Mustapha Eddahibi, Azzeddine Lazrek, and Khalid Sami. Arabic mathematical e-documents. TUGboat, 25(0):42, August 2004. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-0/lazrek-doc.
pdf.

Banouni:2004:DAM

[2823] Mostafa Banouni, Mohamed Elyaakoubi, and Azzeddine Lazrek. Dynamic Arabic mathematical fonts. TUGboat, 25(0):48, August 2004. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-0/
lazrek-font.pdf.

Piska:2004:CTF

[2824] Karel Píška. Creating Type 1 fonts from MF sources: Comparison of tools, techniques and results. TUGboat, 25(0):54, August 2004. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-0/piska.pdf.

Haralambous:2004:OOP

[2825] Yannis Haralambous and Gábor Bella. OpenType and OMEGA: Past, present and future. TUGboat, 25(0):66, August 2004. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb25-0/hara.pdf.

Plaice:2004:MOC

[2826] John Plaice and Paul Swoboda. Moving OMEGA to a C++-based platform. TUGboat, 25(0):68, August 2004. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb25-0/plaice.pdf.

Cho:2004:TCL

[2827] Jin-Hwan Cho and Haruhiko Okumura. Typesetting CJK languages with OMEGA. TUGboat, 25(0):71, August 2004. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb25-0/cho.pdf.

Hufflen:2004:MBL

[2828] Jean-Michael Hufflen. MlBibTeX: Beyond L^{A}TEX. TUGboat, 25(0):77, August 2004. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-0/hufflen.pdf.

Laurens:2004:IIT

[2829] Jérôme Laurens. iTeXMac: An integrated TEX environment for Mac OSX. TUGboat, 25(0):85, August 2004. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb25-0/laurens.pdf.

Schaefer:2004:SSC

[2830] Frank-Rene Schaefer. ŞäferTEX: Source code esthetics for automated typesetting. TUGboat, 25(0):92, August 2004. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb25-0/schaefer.pdf.

Goncalves:2004:FRM

[2831] L. N. Goncalves. FEATPOST and a review of 3D METAPOST packages. TUGboat, 25(0):99, August 2004. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb25-0/nobre.pdf.

Kastrup:2004:PBB

[2832] David Kastrup. The bigfoot bundle for critical editions. TUGboat, 25(0):105, August 2004. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-0/kastrup.pdf.

Likos:2004:PMJ

[2833] Johannis Likos. monopoly: Java-based conversion of monotonic to polytonic Greek. TUGboat, 25(0):111, August 2004. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb25-0/likos.pdf.

Filippou:2004:HPA

[2834] Dimitrios Filippou. Hyphenation patterns for ancient and modern Greek. TUGboat, 25(0):121, August 2004. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb25-0/filippou.pdf.

Rosell-Gonzalez:2004:PMP

[2835] Pablo Rosell-González. The mayan package and fonts. TUGboat, 25(0):127, August 2004. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-0/gonzalez.pdf.

Athale:2004:ULT

[2836] Manasi Athale and Rahul Athale. Using L^{A}TEX to typeset a Marathi–English dictionary. TUGboat, 25(0):131, August 2004. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-0/athale.pdf.

Perez:2004:BCS

[2837] Jagoba Arias Perez, Jesus Lazaro, and Juan M. Aguirregabiria. Basque: A case study in generalizing L^{A}TEX language support. TUGboat, 25(0):135, August 2004. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-0/arias.pdf.

Szabo:2004:ITH

[2838] Péter Szabó. Implementation tricks in the Hungarian babel module. TUGboat, 25(0):140, August 2004. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb25-0/szabo.pdf.

Beesley:2004:TDA

[2839] Kenneth R. Beesley. Typesetting the Deseret alphabet with L^{A}TEX and - . TUGboat, 25(0):162, August 2004. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb25-0/beesley.pdf.

Anonymous:2004:C

[2840] Anonymous. Calendar. TUGboat, 25(0):192, August 2004. CODEN ????

ISSN 0896-3207. URL https://tug.org/calendar.html.

Anonymous:2004:IM

[2841] Anonymous. Institutional members. TUGboat, 25(0):193, August 2004. CO-DEN ???? ISSN 0896-3207.

Anonymous:2004:TMA

[2842] Anonymous. TUG membership application. TUGboat, 25(0):194, August 2004. CODEN ???? ISSN 0896-3207. URL https://tug.org/join.

html.

Anonymous:2004:TCP

[2843] Anonymous. TEX consulting and production services. TUGboat, 25(0):195, August 2004. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:2004:BRG

[2844] Anonymous. Book review: Guide to L^{A}TEX, 4th Edition, by Helmut Kopka and Patrick W. Daly. TUGboat, 25(0):196, August 2004. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/books/.

Anonymous:2004:TCA

[2845] Anonymous. TUG 2005 conference announcement. TUGboat, 25(0):196, August 2004. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tug2005/.

Anonymous:2004:PTC

[2846] Anonymous. Practical TEX 2004 conference announcement. TUGboat, 25(0): c3, August 2004. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/practicaltex2004/.

Anonymous:2004:TPa

[2847] Anonymous. Title page. TUGboat, 25(1):1, 2004. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-1/titlepage.pdf.

Carnes:TB25-1-2

[2848] Lance Carnes. Highlights of the Practical TEX 2004 conference. TUGboat, 25(1):2–3, 2004. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb25-1/carnes.pdf.

Anonymous:TB25-1-4

[2849] Anonymous. Conference program and delegates and sponsors. TUGboat, 25 (1):4–5, 2004. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb25-1/confinfo.pdf.

Berry:TB25-1-6

[2850] Karl Berry. Welcome to Practical TEX 2004. TUGboat, 25(1):6, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-1/berry.pdf.

Beebe:TB25-1-7

[2851] Nelson H. F. Beebe. 25 Years of TEX and : Looking back and looking forward — TUG 2003 keynote address. TUGboat, 25(1):7–30,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-1/
beebe-2003keynote.pdf.

Flynn:TB25-1-31

[2852] Peter Flynn. TEX and the interfaces — Practical TEX 2004 keynote address. TUGboat, 25(1):31–34, 2004. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-1/flynn.pdf.

Thanh:TB25-1-35

[2853] Hàn Thê´ Thành. Micro-typographic extensions of pdfTEX in practice. TUGboat, 25(1):35–38, 2004. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-1/thanh.pdf.

Gurari:TB25-1-39

[2854] Eitan M. Gurari. TEX4ht: HTML production. TUGboat, 25(1):39–47,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-1/
gurari.pdf.

Hagen:TB25-1-48

[2855] Hans Hagen. The state of ConTEXt. TUGboat, 25(1):48–51, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-1/hagen-
context.pdf.

Grathwohl:TB25-1-52

[2856] Steve Grathwohl. A simple book design in ConTEXt. TUGboat, 25(1):52–57,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-1/
grathwohl.pdf.

Peter:TB25-1-58

[2857] Steve Peter. TEX and linguistics. TUGboat, 25(1):58–62, 2004. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-1/peter.pdf.

Moses:TB25-1-63

[2858] Brooks Moses. MetaPlot, MetaContour, and other collaborations with - . TUGboat, 25(1):63–70, 2004. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb25-1/moses.pdf.

Richter:TB25-1-71

[2859] William M. Richter. TEX and scripting languages. TUGboat, 25(1):71–88,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-1/
richter.pdf.

Beebe:TB25-1-89

[2860] Nelson H. F. Beebe. A bibliographer’s toolbox. TUGboat, 25(1):89–104,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-1/
beebe-bib.pdf.

Hoekwater:TB25-1-105

[2861] Taco Hoekwater. MetaPost developments. TUGboat, 25(1):105, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-1/hoekwater.
pdf.

Bilotta:TB25-1-105

[2862] Giuseppe Bilotta. The ℵ (Aleph) project. TUGboat, 25(1):105–107,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-1/
bilotta.pdf.

Hagen:TB25-1-108

[2863] Hans Hagen. The TEX Live 2004 collection. TUGboat, 25(1):108–111,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-1/
hagen-tl.pdf.

Anonymous:TB25-1-112

[2864] Anonymous. Calendar. TUGboat, 25(1):112–113, 2004. ISSN 0896-3207.

URL https://tug.org/calendar.html.

Anonymous:TB25-1-114

[2865] Anonymous. TUG 2005, 23–25 August 2005, Wuhan, China. TUGboat, 25(1):

114, 2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-
1/tlctug05.pdf.

Anonymous:TB25-1-114-2

[2866] Anonymous. The L^{A}TEX Companion, 2^{nd} edition, by Frank Mittelbach et al. TUGboat, 25(1):114, 2004. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-1/tlctug05.pdf.

Anonymous:TB25-1-115

[2867] Anonymous. TUG membership application. TUGboat, 25(1):115, 2004. ISSN

0896-3207. URL https://tug.org/join.html.

Anonymous:TB25-1-116

[2868] Anonymous. Institutional members. TUGboat, 25(1):116, 2004. ISSN 0896-

3207. URL https://tug.org/consultants.html.

Anonymous:TB25-1-116-2

[2869] Anonymous. TEX consulting and production services. TUGboat, 25(1):116,

2004. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB25-1-117

[2870] Anonymous. Easy Table, Khanh Ha. TUGboat, 25(1):117, 2004. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-1/ezad.pdf.

Anonymous:TB25-1-118

[2871] Anonymous. River Valley Technologies. TUGboat, 25(1):118, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-1/riversword.
pdf.

Anonymous:TB25-1-118-2

[2872] Anonymous. MacKichan Software, Inc. TUGboat, 25(1):118, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-1/riversword.
pdf.

Anonymous:TB25-1-119

[2873] Anonymous. Carleton Production Centre. TUGboat, 25(1):119, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-1/chch.pdf.

Anonymous:TB25-1-119-2

[2874] Anonymous. Cheryl Ponchin Training. TUGboat, 25(1):119, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-1/chch.pdf.

Anonymous:TB25-1-120

[2875] Anonymous. Recognition of support from Apple. TUGboat, 25(1):120,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-1/
pctexapple.pdf.

Anonymous:TB25-1-120-2

[2876] Anonymous. Personal TEX, Inc. TUGboat, 25(1):120, 2004. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb25-1/pctexapple.pdf.

Anonymous:2004:CODa

[2877] Anonymous. Contents ordered by difficulty. TUGboat, 25(1):c3, 2004. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-
1/cover3.pdf.

Anonymous:2004:FCa

[2878] Anonymous. Front cover. TUGboat, 25(1):c1, 2004. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-1/cover1.pdf.

Anonymous:2004:IFCa

[2879] Anonymous. Inside front cover. TUGboat, 25(1):c2, 2004. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-1/
cover2.pdf.

Anonymous:TB25-1-cover-3

[2880] Anonymous. Practical TEX 2005, 14–17 June 2005, Chapel Hill, North Carolina. TUGboat, 25(1):Cover 3, 2004. ISSN 0896-3207.

Anonymous:2004:TPb

[2881] Anonymous. Title page. TUGboat, 25(2):121, 2004. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-2/titlepage.pdf.

Anonymous:2004:EP

[2882] Anonymous. Editorial page. TUGboat, 25(2):122, 2004. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-2/tb81edit.
pdf.

Berry:TB25-2-123

[2883] Karl Berry. From the President. TUGboat, 25(2):123–124, 2004. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-2/tb81pres.pdf.

Beeton:TB25-2-124

[2884] Barbara Beeton. Editorial comments. TUGboat, 25(2):124–125, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-2/tb81beet.
pdf.

Hefferon:TB25-2-126

[2885] Jim Hefferon. CTAN for starters. TUGboat, 25(2):126–127, 2004. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-2/tb81heff.pdf.

Peter:TB25-2-128

[2886] Steve Peter. \starttext: Practical ConTEXt. TUGboat, 25(2):128–130,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-2/
tb81peter.pdf.

Schmitz:TB25-2-131

[2887] Thomas Schmitz. Virtual fonts — a tutorial. TUGboat, 25(2):131–133,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-2/
tb81schmitz.pdf.

Flynn:TB25-2-134

[2888] Peter Flynn. Typographers’ Inn. TUGboat, 25(2):134–136, 2004. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-2/tb81inn.pdf.

Beccari:TB25-2-136

[2889] Claudio Beccari and Cristiano Pulone. Philological facilities for the Coptic script. TUGboat, 25(2):136–140, 2004. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb25-2/tb81becc.pdf.

Lazrek:TB25-2-141

[2890] Azzeddine Lazrek. RyDArab — Typesetting Arabic mathematical expressions. TUGboat, 25(2):141–149, 2004. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb25-2/tb81lazrek.pdf.

Pakin:TB25-2-150

[2891] Scott Pakin. PerlTEX: Defining L^{A}TEX macros using Perl. TUGboat, 25(2): 150–159, 2004. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb25-2/tb81pakin.pdf.

Kroonenberg:TB25-2-159

[2892] Siep Kroonenberg. TEX and prepress. TUGboat, 25(2):159–165, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-2/tb81kroonenberg.
pdf.

Castier:TB25-2-166

[2893] Marcelo Castier and Vladimir F. Cabral. Automatic typesetting of formulas using computer algebra. TUGboat, 25(2):166–172, 2004. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb25-2/tb81castier.pdf.

Hwang:TB25-2-172

[2894] Andrew D. Hwang. ePiX: A utility for creating mathematically accurate figures. TUGboat, 25(2):172–176, 2004. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb25-2/tb81hwang.pdf.

Hagon:TB25-2-177

[2895] J. P. Hagon. L^{A}TEX in 3D: OpenDX annotations. TUGboat, 25(2):177–187,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-2/
tb81hagon.pdf.

Dominici:TB25-2-188

[2896] Massimiliano Dominici. dramatist: Another package for typesetting drama with L^{A}TEX. TUGboat, 25(2):188–192, 2004. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb25-2/tb81dominici.pdf.

Law:TB25-2-193

[2897] Simon Law. Variable width boxes in L^{A}TEX. TUGboat, 25(2):193, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-2/tb81law.
pdf.

Adriaens:TB25-2-194

[2898] Hendri Adriaens and Uwe Kern. xkeyval — new developments and mechanisms in key processing. TUGboat, 25(2):194–199, 2004. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb25-2/tb81adriaens.pdf.

Walden:TB25-2-199

[2899] David Walden. A non-expert looks at a small TEX macro. TUGboat, 25(2): 199–201, 2004. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb25-2/tb81walden.pdf.

Wilson:TB25-2-201

[2900] Peter Wilson. Glisterings: Package/package and class/package clashes. TUGboat, 25(2):201–202, 2004. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-2/tb81glister.pdf.

LaPlante:TB25-2-203

[2901] Mark LaPlante. The treasure chest. TUGboat, 25(2):203–208, 2004. ISSN

0896-3207. URL https://tug.org/TUGboat/tb25-2/tb81chest.
pdf.

Anonymous:TB25-2-209

[2902] Anonymous. Zpravodaj: contents of issues 13(1), 14(1), 14(2) (2003–2004). TUGboat, 25(2):209–210, 2004. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb25-2/tb81zprav.pdf.

Anonymous:TB25-2-210

[2903] Anonymous. Die TEXnische Komödie: contents of issues 1–4/2003. TUGboat, 25(2):210–212, 2004. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-2/tb81komo.pdf.

Anonymous:TB25-2-213

[2904] Anonymous. Biuletyn GUST: contents of issues 20–21 (2004). TUGboat, 25(2):213–215, 2004. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb25-2/tb81gust.pdf.

Anonymous:TB25-2-215

[2905] Anonymous. Les Cahiers GUTenberg: contents of issue 43 (2003). TUGboat, 25(2):215, 2004. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb25-2/tb81gust.pdf.

Anonymous:TB25-2-216

[2906] Anonymous. MAPS: contents of issues 29–31 (2003–2004). TUGboat, 25(2): 216–220, 2004. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb25-2/tb81maps.pdf.

Anonymous:TB25-2-221

[2907] Anonymous. Calendar. TUGboat, 25(2):221–222, 2004. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb25-2/tb81calendar.pdf.

Anonymous:TB25-2-223

[2908] Anonymous. TUG 2005 announcement. TUGboat, 25(2):223, 2004. ISSN

0896-3207. URL https://tug.org/tug2005/.

Bari:TB25-2-223

[2909] Onofrio de Bari and Maurizio Himmelmann. A brief report on the first GuIT meeting. TUGboat, 25(2):223, 2004. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb25-2/tb81guit.pdf.

Berry:TB25-2-224

[2910] Karl Berry and Kaja Christiansen. TEX Development Fund 2003–05 report. TUGboat, 25(2):224–225, 2004. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb25-2/tb81devfund.pdf.

Laakso:TB25-2-226

[2911] Robin Laakso. Financial statements for 2004. TUGboat, 25(2):226–227,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-2/
tb81treas.pdf.

Beeton:TB25-2-228

[2912] Barbara Beeton. TUG 2005 election report. TUGboat, 25(2):228–231,

2004. ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-2/
tb81elec.pdf.

Anonymous:TB25-2-232

[2913] Anonymous. Institutional members. TUGboat, 25(2):232, 2004. ISSN 0896-

3207. URL https://tug.org/consultants.html.

Anonymous:TB25-2-232-2

[2914] Anonymous. TEX consulting and production services. TUGboat, 25(2):232,

2004. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2004:CODb

[2915] Anonymous. Contents ordered by difficulty. TUGboat, 25(2):c3, 2004. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-
2/cover3.pdf.

Anonymous:2004:FCb

[2916] Anonymous. Front cover. TUGboat, 25(2):c1, 2004. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb25-2/cover1.pdf.

Anonymous:2004:IFCb

[2917] Anonymous. Inside front cover. TUGboat, 25(2):c2, 2004. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb25-2/
cover2.pdf.

Anonymous:2005:TPa

[2918] Anonymous. Title page. TUGboat, 26(1):1, 2005. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-1/titlepage.pdf.

Berry:TB26-1-3

Beeton:TB26-1-3

[2920] Barbara Beeton. Editorial comments. TUGboat, 26(1):3–5, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-1/tb82beet.pdf.

Beeton:TB26-1-5

[2921] Barbara Beeton. Hyphenation exception log. TUGboat, 26(1):5–6, 2005. ISSN

0896-3207. URL https://tug.org/TUGboat/tb26-1/tb82hyf.
pdf.

Quaresma:TB26-1-7

[2922] Pedro Quaresma. Stacks in TEX. TUGboat, 26(1):7–9, 2005. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb26-1/tb82quar.pdf.

Roegel:TB26-1-10

[2923] Denis Roegel. Kissing circles: A French romance in MetaPost. TUGboat, 26 (1):10–17, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-1/tb82roegel.pdf.

Miller:TB26-1-17

[2924] Tristan Miller. Using the RPM package manager for (La)TEX packages. TUGboat, 26(1):17–28, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-1/tb82miller-rpm.pdf.

Anonymous:TB26-1-29

[2925] Anonymous. Conference program, delegates, and sponsors. TUGboat, 26 (1):29–30, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-1/confinfo.pdf.

Flom:TB26-1-31

[2926] Peter Flom and Tristan Miller. Impressions from PracTEX’05. TUGboat, 26 (1):31–32, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-1/impressions.pdf.

Beebe:TB26-1-33

[2927] Nelson H. F. Beebe. The design of TEX and : A retrospective. TUGboat, 26(1):33–41, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-1/beebe.pdf.

Flom:TB26-1-52

[2928] Peter Flom. A L^{A}TEX fledgling struggles to take flight. TUGboat, 26(1):52–55,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-1/
flom.pdf.

Schwartz:TB26-1-56

[2929] Anita Schwartz. The art of L^{A}TEX problem solving. TUGboat, 26(1):56–58,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-1/
schwartz.pdf.

Hoppner:TB26-1-59

[2930] Klaus Höppner. Strategies for including graphics in L^{A}TEX documents. TUGboat, 26(1):59–62, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-1/hoeppner.pdf.

Hogg:TB26-1-63

[2931] Joseph Hogg. Making a booklet. TUGboat, 26(1):63–65, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-1/hogg.pdf.

Flynn:TB26-1-66

[2932] Peter Flynn. L^{A}TEX on the Web. TUGboat, 26(1):66–67, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-1/flynn.pdf.

Mertz:TB26-1-68

[2933] Andrew Mertz and William Slough. Beamer by example. TUGboat, 26(1):68–

73, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-
1/mertz.pdf.

Bazargan:TB26-1-74

[2934] Kaveh Bazargan. Batch Commander: A graphical user interface for TEX. TUGboat, 26(1):74–80, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-1/bazargan.pdf.

Ignat:TB26-1-81

[2935] David Ignat. Word to L^{A}TEX for a large, multi-author scientific paper. TUGboat, 26(1):81–84, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-1/ignat.pdf.

Miller:TB26-1-85

[2936] Tristan Miller. Biblet: A portable BibTEX bibliography style for generating highly customizable XHTML. TUGboat, 26(1):85–96, 2005. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb26-1/miller-biblet.pdf.

Anonymous:TB26-1-97

[2937] Anonymous. Abstracts (Allen, Burt, Fehd, Gurari, Janc, Kew, Peter). TUGboat, 26(1):97–98, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-1/abstracts.pdf.

Anonymous:TB26-1-99

[2938] Anonymous. Calendar. TUGboat, 26(1):99–100, 2005. ISSN 0896-3207. URL

      https://tug.org/TUGboat/tb26-1/tb82calendar.pdf.

Anonymous:TB26-1-101

[2939] Anonymous. Silmaril Consultants. TUGboat, 26(1):101, 2005. ISSN 0896-

      3207. URL https://tug.org/TUGboat/tb26-1/ad1.pdf.

Anonymous:TB26-1-101-2

[2940] Anonymous. Joe Hogg. TUGboat, 26(1):101, 2005. ISSN 0896-3207. URL

      https://tug.org/TUGboat/tb26-1/ad1.pdf.

Anonymous:TB26-1-101-3

[2941] Anonymous. Carleton Production Centre. TUGboat, 26(1):101, 2005. ISSN

      0896-3207. URL https://tug.org/TUGboat/tb26-1/ad1.pdf.

Anonymous:TB26-1-102

[2942] Anonymous. Personal TEX, Inc. TUGboat, 26(1):102, 2005. ISSN 0896-3207.

      URL https://tug.org/TUGboat/tb26-1/ad2.pdf.

Anonymous:TB26-1-102-2

[2943] Anonymous. River Valley Technologies. TUGboat, 26(1):102, 2005. ISSN

      0896-3207. URL https://tug.org/TUGboat/tb26-1/ad2.pdf.

Anonymous:TB26-1-102-3

[2944] Anonymous. Design Science, Inc. TUGboat, 26(1):102, 2005. ISSN 0896-

      3207. URL https://tug.org/TUGboat/tb26-1/ad2.pdf.

Anonymous:TB26-1-102-4

[2945] Anonymous. The L^{A}TEX Companion, 2^{nd} edition, by Frank Mittelbach et al. TUGboat, 26(1):102, 2005. ISSN 0896-3207. URL https://tug.org/

      TUGboat/tb26-1/ad2.pdf.

Anonymous:TB26-1-103

[2946] Anonymous. Steve Peter. TUGboat, 26(1):103, 2005. ISSN 0896-3207. URL

      https://tug.org/TUGboat/tb26-1/ad3.pdf.

Anonymous:TB26-1-103-2

[2947] Anonymous. Cheryl Ponchin Training. TUGboat, 26(1):103, 2005. ISSN

      0896-3207. URL https://tug.org/TUGboat/tb26-1/ad3.pdf.

Anonymous:TB26-1-103-3

[2948] Anonymous. MacKichan Software, Inc. TUGboat, 26(1):103, 2005. ISSN

      0896-3207. URL https://tug.org/TUGboat/tb26-1/ad3.pdf.

Anonymous:TB26-1-104

[2949] Anonymous. Institutional members. TUGboat, 26(1):104, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-1/consultants.
pdf.

Anonymous:TB26-1-104-2

[2950] Anonymous. TEX consulting and production services. TUGboat, 26(1):104,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-1/
consultants.pdf.

Anonymous:2005:CODa

[2951] Anonymous. Contents ordered by difficulty. TUGboat, 26(1):c3, 2005. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-
1/cover3.pdf.

Anonymous:2005:IFCa

[2952] Anonymous. Inside front cover. TUGboat, 26(1):c2, 2005. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-1/
cover2.pdf.

Anonymous:2005:TPb

[2953] Anonymous. Title page. TUGboat, 26(2):105, 2005. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-2/titlepage.pdf.

Anonymous:TB26-2-106

[2954] Anonymous. Conference program, delegates, and sponsors. TUGboat, 26(2): 106–107, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-2/confinfo.pdf.

Laakso:TB26-2-108

[2955] Robin Laakso. Highlights from TUG 2005. TUGboat, 26(2):108–109,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-2/
laakso.pdf.

Anonymous:TB26-2-110

[2956] Anonymous. Conference photos. TUGboat, 26(2):110, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-2/confphoto.pdf.

Wong:TB26-2-111

[2957] Wai Wong. Typesetting Chinese: A personal perspective. TUGboat, 26(2): 111–114, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-2/wong.pdf.

Kew:TB26-2-115

[2958] Jonathan Kew. XETEX, the Multilingual Lion: TEX meets Unicode and smart font technologies. TUGboat, 26(2):115–124, 2005. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb26-2/kew.pdf.

Laguna:TB26-2-125

[2959] Javier Rodríguez Laguna. Hóng-Zì: A Chinese . TUGboat, 26(2): 125–128, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-2/laguna.pdf.

Yiu:TB26-2-129

[2960] Candy L. K. Yiu and Jim Binkley. Qin notation generator. TUGboat, 26(2): 129–134, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-2/yiu.pdf.

Bagchee:TB26-2-135

[2961] Nandan Bagchee and Eitan M. Gurari. SwiExr: Spatial math exercises and worksheets, in Braille and print. TUGboat, 26(2):135–141, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-2/gurari.pdf.

Taylor:TB26-2-142

[2962] Philip Taylor. Typesetting the Byzantine Cappelli. TUGboat, 26(2):142–151,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-2/
taylor.pdf.

Hagen:TB26-2-152

[2963] Hans Hagen. LuaTEX: Howling to the moon. TUGboat, 26(2):152–157,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-2/
hagen.pdf.

Piska:TB26-2-158

[2964] Karel Píška. Converting  sources to outline fonts using  -

. TUGboat, 26(2):158–164, 2005. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb26-2/piska.pdf.

Venkatesan:TB26-2-165

[2965] S. K. Venkatesan. Moving from bytes to words to semantics. TUGboat, 26(2): 165–168, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-2/venkat.pdf.

Anonymous:TB26-2-169

[2966] Anonymous. Abstracts (Cho, Hong, Rowley, Taylor). TUGboat, 26(2):169,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-2/
abstracts.pdf.

Anonymous:TB26-2-170

[2967] Anonymous. Calendar. TUGboat, 26(2):170–171, 2005. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb26-2/tb83calendar.pdf.

Anonymous:TB26-2-172

[2968] Anonymous. TUG membership form. TUGboat, 26(2):172, 2005. ISSN 0896-

3207. URL https://tug.org/join.html.

Anonymous:TB26-2-173

[2969] Anonymous. Institutional members. TUGboat, 26(2):173, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-2/consultants.
pdf.

Anonymous:TB26-2-173-2

[2970] Anonymous. TEX consulting and production services. TUGboat, 26(2):173,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-2/
consultants.pdf.

Anonymous:TB26-2-174

[2971] Anonymous. TEX Collection 2005 (TEX Live, proTEXt, MacTEX, CTAN). TUGboat, 26(2):174, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-2/texcollection05.pdf.

Anonymous:TB26-2-175

[2972] Anonymous. EuroTEX 2006 announcement. TUGboat, 26(2):175, 2005. ISSN

0896-3207. URL https://tug.org/TUGboat/tb26-2/eurotex06.
pdf.

Anonymous:TB26-2-176

[2973] Anonymous. TUG 2006 announcement. TUGboat, 26(2):176, 2005. ISSN

0896-3207. URL https://tug.org/TUGboat/tb26-2/tug06.pdf.

Anonymous:2005:CODb

[2974] Anonymous. Contents ordered by difficulty. TUGboat, 26(2):c3, 2005. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-
2/cover3.pdf.

Anonymous:2005:IFCb

[2975] Anonymous. Inside front cover. TUGboat, 26(2):c2, 2005. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-2/
cover2.pdf.

Anonymous:TB26-2-c3

[2976] Anonymous. Table of contents, ordered by difficulty. TUGboat, 26(2):c3–c3, 2005. ISSN 0896-3207.

Anonymous:2005:TPc

[2977] Anonymous. Title page. TUGboat, 26(3):105, 2005. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-3/titlepage.pdf.

Berry:TB26-3-179

[2978] Karl Berry. From the President. TUGboat, 26(3):179, 2005. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb26-3/tb84pres.pdf.

Beeton:TB26-3-180

[2979] Barbara Beeton. Editorial comments. TUGboat, 26(3):180–182, 2005. ISSN

0896-3207. URL https://tug.org/TUGboat/tb26-3/tb84beet.
pdf.

Pignalberi:TB26-3-183

[2980] Gianluca Pignalberi. Interview with Donald E. Knuth. TUGboat, 26(3):183–

185, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-
3/tb84knut.pdf.

Tomaszewski:TB26-3-186

[2981] Andrzej Tomaszewski. Implementing editors’ ideas — lots of fun, sometimes even more trouble. TUGboat, 26(3):186–187, 2005. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb26-3/tb84tomaszewski.pdf.

Hefferon:TB26-3-188

[2982] Jim Hefferon. Minutes in less than hours: Using L^{A}TEX resources. TUGboat, 26(3):188–192, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-3/tb84heff.pdf.

Peter:TB26-3-193

[2983] Steve Peter. \starttext: Swelled rules and MetaPost. TUGboat, 26(3): 193–195, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-3/tb84peter.pdf.

Flynn:TB26-3-196

[2984] Peter Flynn. Typographers’ Inn. TUGboat, 26(3):196–198, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-3/tb84inn.pdf.

Wilson:TB26-3-199

[2985] Peter Wilson. The alphabet tree. TUGboat, 26(3):199–214, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-3/tb84wilson.
pdf.

Robertson:TB26-3-215

[2986] Will Robertson. Advanced font features with XETEX — the fontspec package. TUGboat, 26(3):215–223, 2005. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb26-3/tb84robertson.pdf.

Vamvakas:TB26-3-224

[2987] Ioannis Vamvakas and Panagiotis Kotopoulis. ByZL^{A}TEX: a package for typesetting “Byzantine” music. TUGboat, 26(3):224–232, 2005. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb26-3/tb84vamvakas.pdf.

Wierda:TB26-3-235

[2988] Gerben Wierda. i-Installer: The evolution of a TEX install on Mac OS X. TUGboat, 26(3):233–238, 2005. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb26-3/tb84wierda.pdf.

Schulz:TB26-3-239

[2989] Herbert Schulz. The MacTEX distribution. TUGboat, 26(3):239–240,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-3/
tb84schulz.pdf.

Preining:TB26-3-241

[2990] Norbert Preining. TEX Live for Debian. TUGboat, 26(3):241–242,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-3/
tb84preining.pdf.

Hagen:TB26-3-243

[2991] Hans Hagen. Hyphenation patterns in ConTEXt. TUGboat, 26(3):243–245,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-3/
tb84hagen.pdf.

Barr:TB26-3-246

[2992] Michael Barr. Diagxy, a Lego-like diagram package. TUGboat, 26(3):246–

249, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-
3/tb84barr.pdf.

Henderson:TB26-3-250

[2993] Troy Henderson. Embedding fonts in MetaPost output. TUGboat, 26(3):250–

252, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-
3/tb84henderson.pdf.

Wilson:TB26-3-253

[2994] Peter Wilson. Glisterings. TUGboat, 26(3):253–255, 2005. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb26-3/tb84glister.pdf.

Anonymous:TB26-3-256

[2995] Anonymous. Pearls of TEX programming. TUGboat, 26(3):256–263,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-3/
tb84pearls.pdf.

LaPlante:TB26-3-264

[2996] Mark LaPlante. The treasure chest. TUGboat, 26(3):264–267, 2005. ISSN

0896-3207. URL https://tug.org/TUGboat/tb26-3/tb84chest.
pdf.

Adriaens:TB26-3-268

[2997] Hendri Adriaens and Christopher Ellison. powerdot — making presentations with L^{A}TEX. TUGboat, 26(3):268–272, 2005. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb26-3/tb84adriaens.pdf.

Gregorio:TB26-3-273

[2998] Enrico Gregorio. Horrors in L^{A}TEX: How to misuse L^{A}TEX and make a copy editor unhappy. TUGboat, 26(3):273–279, 2005. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb26-3/tb84gregorio.pdf.

Berry:TB26-3-280

[2999] Karl Berry and Oleg Katsitadze. Eplain 3: Expanded plain TEX. TUGboat, 26(3):280–286, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-3/tb84katsi.pdf.

Anonymous:TB26-3-287

[3000] Anonymous. TEXemplares: Contents of issue 7 (2005). TUGboat, 26(3):287,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-3/
tb84zprav.pdf.

Anonymous:TB26-3-287-2

[3001] Anonymous. Zpravodaj: Contents of issues 14(3–4), 15(1), 15(2) (2004–05). TUGboat, 26(3):287–289, 2005. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb26-3/tb84zprav.pdf.

Anonymous:TB26-3-289

[3002] Anonymous. Les Cahiers GUTenberg: Contents of double issue 44–45 (2004). TUGboat, 26(3):289–290, 2005. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb26-3/tb84cahi.pdf.

Anonymous:TB26-3-290

[3003] Anonymous. MAPS: Contents of issues 32 (2005). TUGboat, 26(3):290–291,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-3/
tb84maps.pdf.

Anonymous:TB26-3-292

[3004] Anonymous. The PracTEX Journal: Contents of issues 2005-1–2005-4 (2005). TUGboat, 26(3):292–296, 2005. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb26-3/tb84pracjourn.pdf.

Anonymous:TB26-3-297

[3005] Anonymous. Calendar. TUGboat, 26(3):297–298, 2005. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb26-3/tb84calendar.pdf.

Laakso:TB26-3-298

[3006] Robin Laakso. Financial statements for 2005. TUGboat, 26(3):298–299,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-3/
tb84treas.pdf.

Anonymous:TB26-3-300

[3007] Anonymous. TEX consulting and production services. TUGboat, 26(3):300,

2005. ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-3/
consultants.pdf.

Berry:TB26-3-300

[3008] Karl Berry and Kaja Christiansen. TEX Development Fund 2005 report. TUGboat, 26(3):300, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-3/tb84devfund.pdf.

Anonymous:TB26-3-301

[3009] Anonymous. TUG membership form. TUGboat, 26(3):301, 2005. ISSN 0896-

3207. URL https://tug.org/join.html.

Anonymous:TB26-3-302

[3010] Anonymous. Practical TEX 2006, Piscataway, NJ, July 30–August 1, 2006. TUGboat, 26(3):302, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-3/instmem-list.pdf.

Anonymous:TB26-3-302-2

[3011] Anonymous. Institutional members. TUGboat, 26(3):302, 2005. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb26-3/instmem-list.
pdf.

Anonymous:TB26-3-303

[3012] Anonymous. EuroTEX 2006, Debrecen, Hungary, July 5–8, 2006. TUGboat, 26(3):303, 2005. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb26-3/eurotex06.pdf.

Anonymous:TB26-3-304

[3013] Anonymous. TUG 2006, Marrakesh, Morocco, November 9–11, 2006. TUGboat, 26(3):304, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-3/tug06.pdf.

Anonymous:2005:CODc

[3014] Anonymous. Contents ordered by difficulty. TUGboat, 26(3):c3, 2005. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-
3/cover3.pdf.

Anonymous:2005:IFCc

[3015] Anonymous. Inside front cover. TUGboat, 26(3):c2, 2005. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb26-3/
cover2.pdf.

Anonymous:TB26-3-c3

[3016] Anonymous. Table of contents, ordered by difficulty. TUGboat, 26(3):c3–c3, 2005. ISSN 0896-3207.

Hoppner:2005:LTT

[3017] Klaus Höppner and Fabrice Popineau. Let’s TEX together — EuroTEX 2005 proceedings. TUGboat, 27(0):iv, March 2005. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb27-0/hoeppner.pdf.

Schaa:2005:LPT

[3018] Volker R. W. Schaa. Let’s publish together — EuroTEX 2005 proceedings. TUGboat, 27(0):v, March 2005. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-0/schaa.pdf.

Anonymous:2005:S

[3019] Anonymous. Schedule. TUGboat, 27(0):vi, March 2005. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb27-0/schedule.
pdf.

Anonymous:2005:PIP

[3020] Anonymous. Pictures [introductory — photos from the conference]. TUGboat, 27(0):ix, March 2005. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb27-0/photos.pdf.

Schrod:2005:TAT

[3021] Joachim Schrod, Chris Rowley, and Christine Detig. A taxonomy of automated typesetting systems. TUGboat, 27(0):1, March 2005. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-0/
schrod.pdf.

Roegel:2005:MPO

[3022] Denis Roegel. MP2GL: Prototyping 3D objects with MetaPost and OpenGL. TUGboat, 27(0):5, March 2005. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-0/roegel.pdf.

Hoekwater:2005:MDS

[3023] Taco Hoekwater and MetaPost Team. MetaPost developments — Spring 2005. TUGboat, 27(0):25, March 2005. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-0/hoekwater.pdf.

Dierker:2005:RXL

[3024] Andre Dierker, Arne Jans, and Stephan Lehmke. From RTF to XML to L^{A}TEX. TUGboat, 27(0):33, March 2005. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-0/dierker.pdf.

Rahtz:2005:TTI

[3025] Sebastian Rahtz. The TEI/TEX interface. TUGboat, 27(0):38, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-0/rahtz.pdf.

Mittelbach:2005:LN

[3026] Frank Mittelbach and Chris Rowley. L^{A}TEX3 news. TUGboat, 27(0):50, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-0/mittelbach.pdf.

Twardoch:2005:TPO

[3027] Adam Twardoch. Typographic perfection with OpenType? TUGboat, 27(0): 55, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb27-0/twardoch.pdf.

Neugebauer:2005:NE

[3028] Gerd Neugebauer. Namespaces for exTEX. TUGboat, 27(0):67, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-0/neugebauer.pdf.

Gundlach:2005:CN

[3029] Patrick Gundlach. contextgarden.net. TUGboat, 27(0):71, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-0/gundlach.pdf.

Hoekwater:2005:PDH

[3030] Taco Hoekwater. Panel discussion with Hermann Zapf and Donald Knuth. TUGboat, 27(0):76, March 2005. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-0/knuthzapf.pdf.

Feuerstack:2005:PNT

[3031] Thomas Feuerstack and Klaus Höppner. ProTEXt, a new TEX-collection for beginners. TUGboat, 27(0):78, March 2005. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb27-0/feuerstack.pdf.

Lejay:2005:MFF

[3032] Antoine Lejay. “La machine à formulaires” (the forms’ machine), or TEX for a Kafkaian world. TUGboat, 27(0):81, March 2005. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb27-0/lejay.pdf.

Schaefer:2005:SSC

[3033] Frank-Rene Schaefer. ŞäferTEX: Source code esthetics for automated typesetters. TUGboat, 27(0):86, March 2005. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-0/schaefer.pdf.

Bezos:2005:MMP

[3034] Javier Bezos. Mem: A multilingual package for L^{A}TEX with Aleph. TUGboat, 27(0):92, March 2005. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb27-0/bezos.pdf.

Haralambous:2005:OBS

[3035] Yannis Haralambous and Gábor Bella. Omega becomes a sign processor. TUGboat, 27(0):99, March 2005. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb27-0/haralambous.pdf.

Kastrup:2005:DIL

[3036] David Kastrup. Designing an implementation language for a TEX successor. TUGboat, 27(0):111, March 2005. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-0/kastrup-lang.pdf.

Hefferon:2005:CPR

[3037] Jim Hefferon. CTAN progress report. TUGboat, 27(0):117, March 2005. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-
0/hefferon.pdf.

Szabo:2005:VPL

[3038] Péter Szabó. Verbatim phrases and listings in L^{A}TEX. TUGboat, 27(0):120, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-0/szabo.pdf.

Fine:2005:TF

[3039] Jonathan Fine. TEX forever! TUGboat, 27(0):140, March 2005. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-0/fine.
pdf.

Hagen:2005:FDM

[3040] Hans Hagen. The 16 faces of a Dutch math journal. TUGboat, 27(0):150, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-0/hagen.pdf.

Thanh:2005:EMT

[3041] Hàn Thê´ Thành. Experiences with micro-typographic extensions of pdfTEX in practice. TUGboat, 27(0):159, March 2005. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb27-0/thanh.pdf.

Kuster:2005:NU

[3042] Johannes Küster. Newmath and Unicode. TUGboat, 27(0):165, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-0/kuster.pdf.

Jackowski:2005:LMF

[3043] Bogus�aw Jackowski and Janusz M. Nowacki. Latin Modern fonts: How less means more. TUGboat, 27(0):172, March 2005. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb27-0/jackowski.pdf.

Hufflen:2005:BSE

[3044] Jean-Michel Hufflen. Bibliography styles easier with MlBibTeX. TUGboat, 27(0):179, March 2005. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb27-0/hufflen.pdf.

Laurens:2005:TWS

[3045] Jérôme Laurens. The TEX wrapper structure: A basic TEX document model implemented in iTeXmac. TUGboat, 27(0):193, March 2005. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb27-0/laurens.
pdf.

Kastrup:2005:BBC

[3046] David Kastrup. The bigfoot bundle for critical editions. TUGboat, 27(0):199, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-0/kastrup-bigfoot.pdf.

Anonymous:2005:N

[3047] Anonymous. Names. TUGboat, 27(0):206, March 2005. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb27-0/names.pdf.

Anonymous:2005:EI

[3048] Anonymous. EuroTEX 2006 invitation. TUGboat, 27(0):212, March 2005. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-0/eurotex06.pdf.

Anonymous:2006:TPa

[3049] Anonymous. Title page. TUGboat, 27(1):1, 2006. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb27-1/titlepage.pdf.

Anonymous:TB27-1-2

[3050] Anonymous. Conference delegates and sponsors. TUGboat, 27(1):2,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
tb86confinfo.pdf.

Bujdoso:TB27-1-3

[3051] Gyöngyi Bujdosó. The Hungarian TEX Rhapsody — EuroTEX 2006. TUGboat, 27(1):3, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-1/tb86bujdoso-intro.pdf.

Bella:TB27-1-4

[3052] Gábor Bella. Impressions from EuroTEX 2006. TUGboat, 27(1):4–7,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
tb86bella.pdf.

Viragvolgyi:TB27-1-8

[3053] András Virágvölgyi. Would Aldus Manutius have used TEX? TUGboat, 27 (1):8–12, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-1/tb86viragvolgyi.pdf.

Fried:TB27-1-13

[3054] Katalin Fried et al. The colourful side of TEX. TUGboat, 27(1):13–15,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
tb86fried.pdf.

Hoekwater:TB27-1-16

[3055] Taco Hoekwater. Opening up the type. TUGboat, 27(1):16–17, 2006. ISSN

0896-3207. URL https://tug.org/TUGboat/tb27-1/tb86hoekwater-
luatex.pdf.

Hoekwater:TB27-1-18

[3056] Taco Hoekwater. MetaPost developments — autumn 2006. TUGboat, 27(1): 18–21, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-1/tb86hoekwater-metapost.pdf.

Kroonenberg:TB27-1-22

[3057] Siep Kroonenberg. Managing a network TEX installation under Windows. TUGboat, 27(1):22–27, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-1/tb86kroonenberg-miktex.pdf.

Bujdoso:TB27-1-28

[3058] Gyöngyi Bujdosó. Typography based on-line help on TEX. TUGboat, 27 (1):28–31, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-1/tb86bujdoso-help.pdf.

Nemeth:TB27-1-32

[3059] László Németh. Automatic non-standard hyphenation in OpenOffice.org. TUGboat, 27(1):32–37, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-1/tb86nemeth.pdf.

Hagen:TB27-1-38

[3060] Hans Hagen. What tools do ConTEXt users have? TUGboat, 27(1):38–42,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
tb86hagen.pdf.

Miklos:TB27-1-43

[3061] Ildikó Miklós. KöMaL CD — The execution. TUGboat, 27(1):43–44,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
tb86miklos.pdf.

Bouche:TB27-1-45

[3062] Thierry Bouche. A pdfL^{A}TEX-based automated journal production system. TUGboat, 27(1):45–50, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-1/tb86bouche.pdf.

Bencze:TB27-1-51

[3063] István Bencze, Balázs Fark, László Hatala, and Péter Jeszenszky. Server side PDF generation based on L^{A}TEX templates. TUGboat, 27(1):51–56,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
tb86bencze.pdf.

Szabo:TB27-1-57

[3064] Péter Szabó and András Hraskó. Managing a math exercise database with L^{A}TEX. TUGboat, 27(1):57–61, 2006. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb27-1/tb86szabo.pdf.

Hoekwater:TB27-1-62

[3065] Taco Hoekwater and Hans Hagen. The making of a (TEX) font. TUGboat, 27 (1):62–65, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-1/tb86hoekwater-cows.pdf.

Kroonenberg:TB27-1-66

[3066] Siep Kroonenberg. Font installation the shallow way. TUGboat, 27(1):66–70,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
tb86kroonenberg-fonts.pdf.

Piska:TB27-1-71

[3067] Karel Píška. Font verification and comparison in examples. TUGboat, 27 (1):71–75, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-1/tb86piska.pdf.

Hufflen:TB27-1-76

[3068] Jean-Michel Hufflen. MlBibTeX meets ConTEXt. TUGboat, 27(1):76–82,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
tb86hufflen.pdf.

Jackowski:TB27-1-83

[3069] Bogus�aw Jackowski. Appendix G illuminated. TUGboat, 27(1):83–90,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
tb86jackowski.pdf.

Haralambous:TB27-1-91

[3070] Yannis Haralambous, Gábor Bella, and Atif Gulzar. Open-belly surgery in Ω_{2}. TUGboat, 27(1):91–97, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-1/tb86haralambous-o2.pdf.

Haralambous:TB27-1-98

[3071] Yannis Haralambous. New hyphenation techniques in Ω_{2}. TUGboat, 27(1): 98–103, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-1/tb86haralambous-hyph.pdf.

Anonymous:TB27-1-104

[3072] Anonymous. Abstracts (Fahmy, Hagen, Henkel, Hufflen, Radó et al., Schröder, Szabó). TUGboat, 27(1):104–105, 2006. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb27-1/tb86abstracts.pdf.

Anonymous:TB27-1-106

[3073] Anonymous. Calendar. TUGboat, 27(1):106, 2006. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-1/tb86calendar.pdf.

Anonymous:TB27-1-107

[3074] Anonymous. TUG 2007 election. TUGboat, 27(1):107, 2006. ISSN 0896-

3207. URL https://tug.org/TUGboat/election.

Anonymous:TB27-1-107-2

[3075] Anonymous. Institutional members. TUGboat, 27(1):107, 2006. ISSN 0896-

3207. URL https://tug.org/TUGboat/election.

Anonymous:TB27-1-108

[3076] Anonymous. TEX consulting and production services. TUGboat, 27(1):108,

2006. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2006:CODa

[3077] Anonymous. Contents ordered by difficulty. TUGboat, 27(1):c3, 2006. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-
1/cover3.pdf.

Anonymous:2006:IFCa

[3078] Anonymous. Inside front cover. TUGboat, 27(1):c1, 2006. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-1/
cover2.pdf.

Anonymous:2006:TPb

[3079] Anonymous. Title page. TUGboat, 27(2):109, 2006. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb27-2/titlepage.pdf.

Berry:TB27-2-110

[3080] Karl Berry. From the President. TUGboat, 27(2):110, 2006. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb27-2/tb87pres.pdf.

Beeton:TB27-2-111

[3081] Barbara Beeton. Editorial comments. TUGboat, 27(2):111, 2006. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb27-2/tb87beet.pdf.

Owens:TB27-2-112

[3082] John Owens. The installation and use of OpenType fonts in L^{A}TEX. TUGboat, 27(2):112–118, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-2/tb87owens.pdf.

Wilson:TB27-2-119

[3083] Peter Wilson. Glisterings: Address lists, animated books. TUGboat, 27(2): 119–120, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-2/tb87glister.pdf.

Hall:TB27-2-121

[3084] Timothy Hall. Brackets around anything. TUGboat, 27(2):121–124,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-2/
tb87hall.pdf.

Alex:TB27-2-125

[3085] Alex A.J. Typesetting Malayalam using OMEGA. TUGboat, 27(2):125–126,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-2/
tb87aj.pdf.

Anonymous:TB27-2-127

[3086] Anonymous. Conference program, delegates, and sponsors. TUGboat, 27(2): 127–130, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-2/tb87confinfo.pdf.

Hoekwater:TB27-2-131

[3087] Taco Hoekwater. TUG 2006 report. TUGboat, 27(2):131–136, 2006. ISSN

0896-3207. URL https://tug.org/TUGboat/tb27-2/tb87hoekwater.
pdf.

Benatia:TB27-2-137

[3088] Mohamed Jamal Eddine Benatia, Mohamed Elyaakoubi, and Azzeddine Lazrek. Arabic text justification. TUGboat, 27(2):137–146, 2006. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb27-2/tb87benatia.
pdf.

Jabri:TB27-2-147

[3089] Youssef Jabri. The Arabi system — TEX writes in Arabic and Farsi. TUGboat, 27(2):147–153, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-2/tb87jabri.pdf.

Eddahibi:TB27-2-154

[3090] Mustapha Eddahibi, Azzeddine Lazrek, and Khalid Sami. DadTEX — a full Arabic interface. TUGboat, 27(2):154–158, 2006. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-2/tb87eddahibi.pdf.

Fahmy:TB27-2-159

[3091] Hossam A. H. Fahmy. AlQalam for typesetting traditional Arabic texts. TUGboat, 27(2):159–166, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-2/tb87fahmy.pdf.

Haralambous:TB27-2-167

[3092] Yannis Haralambous. Infrastructure for high-quality Arabic typesetting. TUGboat, 27(2):167–175, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-2/tb87haralambous.pdf.

Wagner:TB27-2-176

[3093] Zdeněk Wagner. Babel speaks Hindi. TUGboat, 27(2):176–180, 2006. ISSN

0896-3207. URL https://tug.org/TUGboat/tb27-2/tb87wagner.
pdf.

Syropoulos:TB27-2-181

[3094] Apostolos Syropoulos. L^{A}TEX as a tool for the typographic reproduction of ancient texts. TUGboat, 27(2):181–186, 2006. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb27-2/tb87syropoulos.pdf.

Smirnova:TB27-2-187

[3095] Elena Smirnova and Stephen M. Watt. Generating TEX from mathematical content with respect to notational settings. TUGboat, 27(2):187–196,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-2/
tb87smirnova.pdf.

Frischauf:TB27-2-197

[3096] Adrian Frischauf and Paul Libbrecht. DVI2SVG: Using L^{A}TEX layout on the Web. TUGboat, 27(2):197–201, 2006. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb27-2/tb87frischauf.pdf.

Beccari:TB27-2-202

[3097] Claudio Beccari. L^{A}TEX 2ε, pict2e and complex numbers. TUGboat, 27(2): 202–212, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-2/tb87becc.pdf.

Hogholm:TB27-2-213

[3098] Morten Høgholm. Page design in L^{A}TEX3. TUGboat, 27(2):213–218,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-2/
tb87hoegholm.pdf.

Hagen:TB27-2-219

[3099] Hans Hagen. MKII–MKIV. TUGboat, 27(2):219–227, 2006. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb27-2/tb87hagen-
context.pdf.

Kew:TB27-2-228

[3100] Jonathan Kew. Unicode and multilingual typesetting with XETEX. TUGboat, 27(2):228–229, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-2/tb87kew.pdf.

Hagen:TB27-2-230

[3101] Hans Hagen, Jerzy B. Ludwichowski, and Volker R. W. Schaa. The New Font Project: TEX Gyre. TUGboat, 27(2):230–233, 2006. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-2/tb87hagen-gyre.pdf.

Piska:TB27-2-234

[3102] Karel Píška. Outline font extensions for Arabic typesetting. TUGboat, 27(2): 234–237, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-2/tb87piska.pdf.

Mounayerji:TB27-2-238

[3103] F. Mounayerji and M. A. Naal. Arabic font building for L^{A}TEX. TUGboat, 27(2): 238–240, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-2/tb87mounayerji.pdf.

Rowley:TB27-2-241

[3104] Chris Rowley. Everything we want to know about Font Resources. TUGboat, 27(2):241–242, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-2/tb87rowley.pdf.

Hufflen:TB27-2-243

[3105] Jean-Michel Hufflen. Names in BibTeX and mlBibTeX. TUGboat, 27(2): 243–253, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-2/tb87hufflen.pdf.

Anonymous:TB27-2-254

[3106] Anonymous. Abstracts (Beeton, Bujdosó, Feuerstack, Hagen, Hoekwater, Ludwichowski, Wierda). TUGboat, 27(2):254–255, 2006. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-2/tb87abstracts.pdf.

Anonymous:TB27-2-256

[3107] Anonymous. MAPS: contents of issues 33–34 (2005–06). TUGboat, 27(2): 256–257, 2006. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb27-2/tb87maps.pdf.

Anonymous:TB27-2-258

[3108] Anonymous. ArsTEXnica: contents of issue 1 (2006). TUGboat, 27(2):258,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-2/
tb87gust.pdf.

Anonymous:TB27-2-258-2

[3109] Anonymous. Biuletyn GUST: contents of issues 22–23 (2005–06). TUGboat, 27(2):258–262, 2006. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb27-2/tb87gust.pdf.

Anonymous:TB27-2-263

[3110] Anonymous. Calendar. TUGboat, 27(2):263, 2006. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb27-2/tb87calendar.pdf.

Anonymous:TB27-2-264

[3111] Anonymous. EuroBachoTEX 2007 announcement. TUGboat, 27(2):264,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-2/
tb87eurotex07.pdf.

Bari:TB27-2-265

[3112] Onofrio de Bari. The 3rd Annual GuIT Meeting. TUGboat, 27(2):265,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-2/
tb87guituktug.pdf.

Anonymous:TB27-2-266

[3113] Anonymous. Institutional members. TUGboat, 27(2):266, 2006. ISSN 0896-3207.

Goldie:TB27-2-266

[3114] Charles Goldie. UKTUG sponsors day of L^{A}TEX. TUGboat, 27(2):266,

2006. ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-2/
tb87guituktug.pdf.

Anonymous:TB27-2-267

[3115] Anonymous. TUG membership form. TUGboat, 27(2):267, 2006. ISSN 0896-3207.

Anonymous:TB27-2-268

[3116] Anonymous. TEX consulting and production services. TUGboat, 27(2):268,

2006. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2006:CODb

[3117] Anonymous. Contents ordered by difficulty. TUGboat, 27(2):c3, 2006. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-
2/cover3.pdf.

Anonymous:2006:IFCb

[3118] Anonymous. Inside front cover. TUGboat, 27(2):c2, 2006. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb27-2/
cover2.pdf.

Anonymous:2007:TPa

[3119] Anonymous. Title page. TUGboat, 28(1):1, 2007. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-1/titlepage.pdf.

Berry:TB28-1-3

[3120] Karl Berry. From the President. TUGboat, 28(1):3, 2007. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb28-1/tb88pres.pdf.

Beeton:TB28-1-4

[3121] Barbara Beeton. Editorial comments. TUGboat, 28(1):3–4, 2007. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-1/tb88beet.pdf.

Knuth:TB28-1-4

[3122] Donald E. Knuth. TEX’s infinite glue is projective. TUGboat, 28(1):4,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-1/
tb88knut.pdf.

Parashchenko:TB28-1-5

[3123] Oleg Parashchenko. TEXML: Resurrecting TEX in the XML world. TUGboat, 28(1):5–10, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-1/tb88parashchenko.pdf.

Beeton:TB28-1-11

[3124] Barbara Beeton and Idris Hamid. Oriental TEX: A new direction in scholarly complex-script typesetting. TUGboat, 28(1):11, 2007. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb28-1/tb88hamid.pdf.

Wilson:TB28-1-12

[3125] Peter Wilson. Glisterings: stringing along; loops. TUGboat, 28(1):12–14,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-1/
tb88glister.pdf.

LaPlante:TB28-1-15

[3126] Mark LaPlante. The treasure chest. TUGboat, 28(1):15–19, 2007. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-1/tb88chest.pdf.

Tortosa:TB28-1-20

[3127] Ignacio Llopis Tortosa and María José Castro Bleda. paperTEX: Creating newspapers using L^{A}TEX 2ε. TUGboat, 28(1):20–23, 2007. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb28-1/tb88tortosa.pdf.

Team:TB28-1-24

[3128] L^{A}TEX Project Team. L^{A}TEX news, issue 17. TUGboat, 28(1):24–25,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-1/
tb88ltnews.pdf.

Anonymous:TB28-1-26

[3129] Anonymous. Conference program, delegates, and sponsors. TUGboat, 28(1):

26, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-
1/tb88confinfo.pdf.

Beeton:TB28-1-29

[3130] Barbara Beeton. How to create a TEX journal: a personal journey. TUGboat, 28 (1):29–48, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-1/tb88beet-keynote.pdf.

Walden:TB28-1-49

[3131] David Walden. A lifetime as an amateur compositor. TUGboat, 28(1):49–60,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-1/
tb88walden.pdf.

Dearborn:TB28-1-61

[3132] Elizabeth Dearborn. TEX and medicine. TUGboat, 28(1):61–64, 2007. ISSN

0896-3207. URL https://tug.org/TUGboat/tb28-1/tb88dearborn.
pdf.

Breitenbucher:TB28-1-65

[3133] Jon Breitenbucher. L^{A}TEX at a liberal arts college. TUGboat, 28(1):65–69,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-1/
tb88breitenbucher.pdf.

Veytsman:TB28-1-70

[3134] Boris Veytsman. Design of presentations: Notes on principles and L^{A}TEX implementation. TUGboat, 28(1):70–76, 2007. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb28-1/tb88veytsman-presentations.
pdf.

Veytsman:TB28-1-77

[3135] Boris Veytsman and Maria Shmilevich. Automatic report generation with Web, TEX and SQL. TUGboat, 28(1):77–79, 2007. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb28-1/tb88veytsman-report.pdf.

Neveln:TB28-1-80

[3136] Bob Neveln and Bob Alps. Writing and checking complete proofs in TEX. TUGboat, 28(1):80–83, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-1/tb88neveln.pdf.

Henderson:TB28-1-84

[3137] Troy Henderson. A beginner’s guide to MetaPost for creating high-quality graphics. TUGboat, 28(1):84–90, 2007. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb28-1/tb88henderson.pdf.

Mertz:TB28-1-91

[3138] Andrew Mertz and William Slough. Graphics with PGF and TikZ. TUGboat, 28 (1):91–99, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-1/tb88mertz.pdf.

Veytsman:TB28-1-100

[3139] Boris Veytsman and Leila Akhmadeeva. Drawing medical pedigree trees with TEX and PSTricks. TUGboat, 28(1):100–109, 2007. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-1/tb88veytsman-
pedigree.pdf.

Flynn:TB28-1-110

[3140] Peter Flynn. Rolling your own Document Class: Using L^{A}TEX to keep away from the Dark Side. TUGboat, 28(1):110–123, 2007. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb28-1/tb88flynn.pdf.

Hefferon:TB28-1-124

[3141] Jim Hefferon. L^{A}TEX resources. TUGboat, 28(1):124–125, 2007. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-1/tb88heff.pdf.

Flom:TB28-1-126

[3142] Peter Flom. L^{A}TEX for academics and researchers who (think they) don’t need it. TUGboat, 28(1):126–128, 2007. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb28-1/tb88flom.pdf.

Garcia:TB28-1-129

[3143] Federico Garcia. Hypertext capabilities with pdfL^{A}TEX. TUGboat, 28(1):129–

132, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-
1/tb88garcia.pdf.

Bazargan:TB28-1-133

[3144] Kaveh Bazargan and C. V. Radhakrishnan. Removing vertical stretch — mimicking traditional typesetting with TEX. TUGboat, 28(1):133–136,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-1/
tb88bazargan.pdf.

Anonymous:TB28-1-137

[3145] Anonymous. Abstracts (Adams, Garcia, Höppner, Hummel, Moye, Peter, Wetmore). TUGboat, 28(1):137, 2007. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb28-1/tb88abstracts.pdf.

Anonymous:TB28-1-138

[3146] Anonymous. Calendar. TUGboat, 28(1):138, 2007. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb28-1/tb88calendar.pdf.

Anonymous:TB28-1-139

[3147] Anonymous. TUG 2007 announcement. TUGboat, 28(1):139, 2007. ISSN

0896-3207. URL https://tug.org/TUGboat/tb28-1/tb88tug07.
pdf.

Anonymous:TB28-1-140

[3148] Anonymous. EuroBachoTEX 2007 announcement. TUGboat, 28(1):140,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-1/
tb88eurotex07.pdf.

Peter:TB28-1-141

[3149] Steve Peter. TUG 2007 election report. TUGboat, 28(1):141–144, 2007. ISSN

0896-3207. URL https://tug.org/TUGboat/tb28-1/tb88elec.
pdf.

Walden:TB28-1-145

[3150] David Walden. Financial statements for 2006. TUGboat, 28(1):145–146,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-1/
tb88treas.pdf.

Anonymous:TB28-1-146

[3151] Anonymous. Institutional members. TUGboat, 28(1):146, 2007. ISSN 0896-3207.

Anonymous:TB28-1-147

[3152] Anonymous. TUG membership form. TUGboat, 28(1):147, 2007. ISSN 0896-

3207. URL https://tug.org/join.html.

Anonymous:TB28-1-148

[3153] Anonymous. TEX consulting and production services. TUGboat, 28(1):148,

2007. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2007:CODa

[3154] Anonymous. Contents ordered by difficulty. TUGboat, 28(1):c3, 2007. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-
1/cover3.pdf.

Anonymous:2007:FCa

[3155] Anonymous. Front cover. TUGboat, 28(1):c1, 2007. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-1/cover1.pdf.

Anonymous:2007:IFCa

[3156] Anonymous. Inside front cover. TUGboat, 28(1):c2, 2007. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-1/
cover2.pdf.

Anonymous:2007:TPb

[3157] Anonymous. Title page. TUGboat, 28(2):149, 2007. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-2/titlepage.pdf.

Berry:TB28-2-151

Beeton:TB28-2-152

[3159] Barbara Beeton. Editorial comments. TUGboat, 28(2):152–153, 2007. ISSN

0896-3207. URL https://tug.org/TUGboat/tb28-2/tb89beet.
pdf.

Moye:TB28-2-153

[3160] Stephen Moye. A wayward wayfarer’s way to TEX. TUGboat, 28(2):153–158,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-2/
tb89moye.pdf.

Miklavec:TB28-2-159

[3161] Mojca Miklavec. ConTEXt user meeting 2007: Epen, March 23–25. TUGboat, 28(2):159–163, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-2/tb89miklavec.pdf.

Guravage:TB28-2-164

[3162] Michael Guravage. EuroBachoTEX 2007. TUGboat, 28(2):164–171,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-2/
tb89guravage.pdf.

Lee:TB28-2-172

[3163] Kihwang Lee. New TEX activities in Korea. TUGboat, 28(2):172, 2007. ISSN

0896-3207. URL https://tug.org/TUGboat/tb28-2/tb89lee.
pdf.

Flynn:TB28-2-172

[3164] Peter Flynn. Typographers’ Inn. TUGboat, 28(2):172–173, 2007. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-2/tb89inn.pdf.

Hagen:TB28-2-174

[3165] Hans Hagen and Taco Hoekwater. Alphabetgeschichten by Hermann Zapf. TUGboat, 28(2):174–176, 2007. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb28-2/tb89hagen.pdf.

Robertson:TB28-2-177

[3166] Will Robertson. An exploration of the Latin Modern fonts. TUGboat, 28(2): 177–180, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-2/tb89robertson.pdf.

Hoppner:TB28-2-181

[3167] Klaus Höppner. Creation of a PostScript Type 1 logo font with MetaType 1. TUGboat, 28(2):181–185, 2007. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb28-2/tb89hoeppner.pdf.

Hellstrom:TB28-2-186

[3168] Lars Hellström. Writing ETX format font encoding specifications. TUGboat, 28(2):186–197, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-2/tb89hellstrom.pdf.

Mahajan:TB28-2-198

[3169] Aditya Mahajan. ConTEXt basics for users: Font styles. TUGboat, 28(2): 198–199, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-2/tb89mahajan.pdf.

Hamid:TB28-2-200

[3170] Idris Samawi Hamid. Installing ConTEXt expert fonts: Minion Pro. TUGboat, 28(2):200–209, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-2/tb89hamid.pdf.

Cho:TB28-2-210

[3171] Jin-Hwan Cho. Hacking DVI files: Birth of DVIasm. TUGboat, 28(2):210–217,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-2/
tb89cho.pdf.

Roegel:TB28-2-218

[3172] Denis Roegel. A complex drawing in descriptive geometry. TUGboat, 28(2): 218–228, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-2/tb89roegel.pdf.

Wilson:TB28-2-229

[3173] Peter Wilson. Glisterings: Paragraphs regular; paragraphs particular; paragraphs Russian. TUGboat, 28(2):229–232, 2007. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb28-2/tb89glister.pdf.

Berry:TB28-2-233

[3174] Karl Berry. The treasure chest. TUGboat, 28(2):233–234, 2007. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-2/tb89chest.pdf.

Garcia:TB28-2-235

[3175] Federico Garcia. L^{A}TEX and the different bibliography styles. TUGboat, 28(2): 235–240, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-2/tb89garcia.pdf.

Schmidt:TB28-2-241

[3176] Walter Schmidt. Font selection in L^{A}TEX: The most frequently asked questions. TUGboat, 28(2):241–242, 2007. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb28-2/tb89schmidt.pdf.

Wilson:2007:MC

[3177] Peter Wilson. The memoir class. TUGboat, 28(2):243, 2007. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-2/
tb89wilson.pdf.

Gregorio:TB28-2-247

[3178] Enrico Gregorio. Enjoying babel. TUGboat, 28(2):247–255, 2007. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-2/tb89gregorio.
pdf.

Reingold:TB28-2-256

[3179] Edward M. Reingold. Writing numbers in words in TEX. TUGboat, 28(2): 256–259, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-2/tb89reingold.pdf.

Anonymous:TB28-2-260

[3180] Anonymous. ArsTEXnica: contents of issues 2–3 (2006–2007). TUGboat, 28(2):260–261, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-2/tb89abs-arscahi.pdf.

Anonymous:TB28-2-261

[3181] Anonymous. Les Cahiers GUTenberg: contents of double issue 46–47 (2006). TUGboat, 28(2):261, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-2/tb89abs-arscahi.pdf.

Anonymous:TB28-2-262

[3182] Anonymous. Die TEXnische Komödie: contents of issues 2006/1–2007/1. TUGboat, 28(2):262–263, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-2/tb89komo.pdf.

Anonymous:TB28-2-263

[3183] Anonymous. The PracTEX Journal: contents of issues 2006-1–2007-2. TUGboat, 28(2):263–268, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-2/tb89pracjourn.pdf.

Anonymous:TB28-2-268

[3184] Anonymous. Institutional members. TUGboat, 28(2):268, 2007. ISSN 0896-

3207.

Anonymous:TB28-2-269

[3185] Anonymous. Calendar. TUGboat, 28(2):269, 2007. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb28-2/tb89calendar.pdf.

Anonymous:TB28-2-270

[3186] Anonymous. TUG 2007 announcement. TUGboat, 28(2):270, 2007. ISSN

0896-3207. URL https://www.tug.org/tug2007/.

Anonymous:TB28-2-271

[3187] Anonymous. TUG 2008 announcement. TUGboat, 28(2):271, 2007. ISSN

0896-3207. URL https://www.tug.org/tug2008.

Anonymous:TB28-2-272

[3188] Anonymous. TEX consulting and production services. TUGboat, 28(2):272–

282, 2007. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:2007:CODb

[3189] Anonymous. Contents ordered by difficulty. TUGboat, 28(2):c3, 2007. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-
2/cover3.pdf.

Anonymous:2007:FCb

[3190] Anonymous. Front cover. TUGboat, 28(2):c1, 2007. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-2/cover1.pdf.

Anonymous:2007:IFCb

[3191] Anonymous. Inside front cover. TUGboat, 28(2):c2, 2007. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-2/
cover2.pdf.

Anonymous:2007:TPc

[3192] Anonymous. Title page. TUGboat, 28(3):273, 2007. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-3/titlepage.pdf.

Anonymous:2007:TCI

[3193] Anonymous. TUG 2007 conference information. TUGboat, 28(3):274, 2007. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-3/tb90confinfo.pdf.

Arnold:TB28-3-276

[3194] Tim Arnold. TUG 2007: a few words. TUGboat, 28(3):276–279,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-3/
tb90arnold.pdf.

Wilson:TB28-3-280

[3195] Peter Wilson. Between then and now — a meandering memoir. TUGboat, 28(3):280–298, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-3/tb90wilson.pdf.

Beeton:TB28-3-299

[3196] Barbara Beeton. The STIX Project — From Unicode to fonts. TUGboat, 28(3): 299–304, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-3/tb90beet.pdf.

Kew:TB28-3-305

[3197] Jonathan Kew and Victor Gaultney. Fonts for every language: SIL’s font projects and the Open Font License. TUGboat, 28(3):305, 2007. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-3/tb90kew.pdf.

Hammond:TB28-3-306

[3198] William Hammond. Dual presentation with math from one source using GELLMU. TUGboat, 28(3):306–311, 2007. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb28-3/tb90hammond.pdf.

Hoekwater:TB28-3-312

[3199] Taco Hoekwater. LuaTEX. TUGboat, 28(3):312–313, 2007. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb28-3/tb90hoekwater-luatex.
pdf.

Hagen:TB28-3-314

[3200] Hans Hagen. ConTEXt MkIV: Going UTF. TUGboat, 28(3):314–316,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-3/
tb90hagen.pdf.

Hoekwater:TB28-3-317

[3201] Taco Hoekwater and Hans Hagen. MPlib: MetaPost as a reusable component. TUGboat, 28(3):317–318, 2007. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb28-3/tb90hoekwater-mplib.pdf.

Beebe:TB28-3-319

[3202] Nelson H. F. Beebe. Extending TEX and MF with floating-point arithmetic. TUGboat, 28(3):319–328, 2007. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb28-3/tb90beebe.pdf.

Koch:TB28-3-329

[3203] Richard Koch. Support for multiple TEX distributions in i-Installer and MacTEX. TUGboat, 28(3):329–334, 2007. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb28-3/tb90koch.pdf.

MacKichan:TB28-3-335

[3204] Barry MacKichan. Design decisions for a structured front end to L^{A}TEX. TUGboat, 28(3):335–339, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-3/tb90mackichan.pdf.

Topping:TB28-3-340

[3205] Paul Topping. MathType 6.0’s TEX input for MS Word and Wikipedia. TUGboat, 28(3):340–341, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-3/tb90topping.pdf.

Burgess:TB28-3-342

[3206] Robert Burgess and Emin Gün Sirer. CrossTEX: a modern bibliography management tool. TUGboat, 28(3):342–349, 2007. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb28-3/tb90burgess.pdf.

Hoppner:TB28-3-350

[3207] Klaus Höppner. Typesetting tables with L^{A}TEX. TUGboat, 28(3):350–353,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-3/
tb90hoeppner.pdf.

Mertz:TB28-3-354

[3208] Andrew Mertz and William Slough. Programming with PerlTEX. TUGboat, 28(3):354–362, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-3/tb90mertz.pdf.

Gurari:TB28-3-363

[3209] Eitan Gurari. L^{A}TEX conversion into normalized forms and speech. TUGboat, 28(3):363–368, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-3/tb90gurari.pdf.

Anonymous:TB28-3-369

[3210] Anonymous. Abstracts (Allen, DeLand, Hagen, Hamid, Hefferon, Høgholm, Ierusalimschy, Kew, Rosenthol, Rowley, Stern). TUGboat, 28(3):369–371,

2007. ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-3/
tb90abstracts.pdf.

Mahajan:TB28-3-372

[3211] Aditya Mahajan. ConTEXt basics for users: Table macros. TUGboat, 28(3): 372–374, 2007. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb28-3/tb90mahajan.pdf.

TDFC:TB28-3-375

[3212] TEX Development Fund committee. A roadmap for TEX development. TUGboat, 28(3):375–376, 2007. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb28-3/tb90devfund.pdf.

Editors:TB28-3-377

[3213] TUGboat Editors. Random comments. TUGboat, 28(3):377–378, 2007. ISSN

0896-3207. URL https://tug.org/TUGboat/tb28-3/tb90random.
pdf.

Anonymous:TB28-3-378

Anonymous:TB28-3-379

[3215] Anonymous. Calendar. TUGboat, 28(3):379, 2007. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb28-3/tb90calendar.pdf.

Anonymous:TB28-3-380

[3216] Anonymous. TUG 2008 announcement. TUGboat, 28(3):380, 2007. ISSN

0896-3207. URL https://tug.org/TUGboat/tb28-3/tb90tug08.
pdf.

Anonymous:TB28-3-381

[3217] Anonymous. Springer: More Math Into L^{A}TEX; von Hoerner & Sulger; O’Reilly. TUGboat, 28(3):381, 2007. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb28-3/tb90ads.pdf.

Anonymous:TB28-3-382

[3218] Anonymous. MacKichan: Scientific WorkPlace, Scientific Word; Addison-Wesley: The L^{A}TEX Graphics Companion. TUGboat, 28(3):382, 2007. ISSN

0896-3207. URL https://tug.org/TUGboat/tb28-3/tb90ads.
pdf.

Anonymous:TB28-3-383

[3219] Anonymous. Cheryl Ponchin training; Carleton Production Centre; design science: MathType^{6}. TUGboat, 28(3):383, 2007. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb28-3/tb90ads.pdf.

Anonymous:TB28-3-384

[3220] Anonymous. TEX consulting and production services. TUGboat, 28(3):384,

2007. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2007:CODc

[3221] Anonymous. Contents ordered by difficulty. TUGboat, 28(3):c3, 2007. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-
3/cover3.pdf.

Anonymous:2007:FCc

[3222] Anonymous. Front cover. TUGboat, 28(3):c1, 2007. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb28-3/cover1.pdf.

Anonymous:2007:IFCc

[3223] Anonymous. Inside front cover. TUGboat, 28(3):c2, 2007. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb28-3/
cover2.pdf.

Anonymous:2008:TPa

[3224] Anonymous. Title page. TUGboat, 29(1):1, 2008. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-1/titlepage.pdf.

Anonymous:TB29-1-2

[3225] Anonymous. Conference program, delegates, and sponsors. TUGboat, 29(1):

2–5, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-
1/tb91confinfo.pdf.

Anonymous:2008:EPa

[3226] Anonymous. EuroTeX 2007 program. TUGboat, 29(1):4, 2008. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91schedule.pdf.

Ludwichowski:TB29-1-6

[3227] Jerzy Ludwichowski and Petr Sojka. EuroBachoTEX 2007: Paths to the Future. TUGboat, 29(1):6–7, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91ludwichowski-intro.pdf.

Anonymous:2008:EPb

[3228] Anonymous. EuroTeX 2007 photos. TUGboat, 29(1):8, 2008. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91photos.pdf.

Guravage:TB29-1-13

[3229] Sam Guravage. Confessions of a teenage TEX user. TUGboat, 29(1):13,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91guravage.pdf.

Jackowska:TB29-1-14

[3230] Grażyna Jackowska. Handmade paper: a mixture of handcraft, art and fun. TUGboat, 29(1):14–15, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91jackowska.pdf.

Tomaszewski:TB29-1-16

[3231] Andrzej Tomaszewski. Designing a special book: With both pleasure and … fear. TUGboat, 29(1):16–19, 2008. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb29-1/tb91tomaszewski.pdf.

Cendrowska:TB29-1-20

[3232] Dorota Cendrowska. Enumerations as an interesting form of text appearance. TUGboat, 29(1):20–24, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91cendrowska.pdf.

Ludwichowski:TB29-1-25

[3233] Jerzy Ludwichowski, Bogus�aw Jackowski, and Janusz Nowacki. Five years after: Report on international TEX font projects. TUGboat, 29(1):25–26,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91ludwichowski-fonts.pdf.

Nowacki:TB29-1-27

[3234] Janusz Nowacki. Cyklop: a new font family. TUGboat, 29(1):27,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91nowacki-cyklop.pdf.

Hagen:TB29-1-28

[3235] Hans Hagen. Do we need a font system in TEX? TUGboat, 29(1):28–33,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91hagen.pdf.

Hoekwater:TB29-1-34

[3236] Taco Hoekwater. OpenType fonts in LuaTEX. TUGboat, 29(1):34–35,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91hoekwater.pdf.

Thanh:TB29-1-36

[3237] Hàn Thê´ Thành. Font-specific issues in pdfTEX. TUGboat, 29(1):36–41,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91thanh-fonts.pdf.

Horak:TB29-1-42

[3238] Karel Horák. Those obscure accents …. TUGboat, 29(1):42–44,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91horak.pdf.

Hoppner:TB29-1-45

[3239] Klaus Höppner. Creation of a PostScript Type 1 logo font with MetaType 1. TUGboat, 29(1):45–49, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91hoeppner.pdf.

Piska:TB29-1-50

[3240] Karel Píška. Procedures for font comparison. TUGboat, 29(1):50–56,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91piska-compare.pdf.

Piska:TB29-1-57

[3241] Karel Píška. Comments and suggestions about the Latin Modern fonts. TUGboat, 29(1):57–65, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91piska-lm.pdf.

Ludwichowski:TB29-1-66

[3242] Jerzy Ludwichowski and Karl Berry. The GUST Font License: an application of the L^{A}TEX Project Public License. TUGboat, 29(1):66–67,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91ludwichowski-gfl.pdf.

Reutenauer:TB29-1-68

[3243] Arthur Reutenauer. A brief history of TEX, volume II. TUGboat, 29(1):68–72,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91reutenauer.pdf.

Vieth:TB29-1-73

[3244] Ulrik Vieth. Overview of the TEX historic archive. TUGboat, 29(1):73–76,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91vieth.pdf.

Rycko:TB29-1-77

[3245] Joanna Ludmi�a Ryćko. TEX Clinic. TUGboat, 29(1):77–78, 2008. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-1/tb91rycko.pdf.

Sherif:TB29-1-79

[3246] Ameer Sherif and Hossam Fahmy. Parameterized Arabic font development for AlQalam. TUGboat, 29(1):79–88, 2008. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb29-1/tb91sherif.pdf.

Gulzar:TB29-1-89

[3247] Atif Gulzar and Shafiq ur Rahman. Nastaleeq: a challenge accepted by Omega. TUGboat, 29(1):89–94, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91gulzar.pdf.

Thanh:TB29-1-95

[3248] Hàn Thê´ Thành. Typesetting Vietnamese with VnTEX (and with the TEX Gyre fonts too). TUGboat, 29(1):95–100, 2008. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb29-1/tb91thanh-vntex.pdf.

Hufflen:TB29-1-101

[3249] Jean-Michel Hufflen. Managing order relations in MlBibTEX. TUGboat, 29(1): 101–108, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-1/tb91hufflen-order.pdf.

Hufflen:TB29-1-109

[3250] Jean-Michel Hufflen. Introducing L^{A}TEX users to XSL-FO. TUGboat, 29(1): 109–116, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-1/tb91hufflen-fo.pdf.

Luczak:TB29-1-117

[3251] Tomasz �uczak. Using TEX in a wiki. TUGboat, 29(1):117, 2008. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-1/tb91luczak-
wiki.pdf.

Sojka:TB29-1-118

[3252] Petr Sojka and Michal Růžička. Single-source publishing in multiple formats for different output devices. TUGboat, 29(1):118–124, 2008. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb29-1/tb91sojka.pdf.

Szabo:TB29-1-125

[3253] Péter Szabó. Practical journal and proceedings publication on paper and on the web. TUGboat, 29(1):125–132, 2008. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb29-1/tb91szabo-pub.pdf.

Hefferon:TB29-1-133

[3254] Jim Hefferon. An experimental CTAN upload process. TUGboat, 29(1):133–

135, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-
1/tb91heff.pdf.

Preining:TB29-1-136

[3255] Norbert Preining. TEX (Live) on Debian. TUGboat, 29(1):136–139,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91preining.pdf.

Kroonenberg:TB29-1-140

[3256] Siep Kroonenberg. Epspdf: Easy conversion between PostScript and PDF. TUGboat, 29(1):140–142, 2008. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb29-1/tb91kroonenberg.pdf.

Schroder:TB29-1-143

[3257] Martin Schröder. pdfTEX 1.40: What’s new. TUGboat, 29(1):143–145,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91schroeder.pdf.

Kew:TB29-1-146

[3258] Jonathan Kew. XETEX Live. TUGboat, 29(1):146–150, 2008. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-1/tb91kew.pdf.

Neugebauer:TB29-1-151

[3259] Gerd Neugebauer. Conventional scoping of registers — an experiment in ExTEX. TUGboat, 29(1):151–156, 2008. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb29-1/tb91neugebauer.pdf.

Hufflen:TB29-1-157

[3260] Jean-Michel Hufflen. MlBibTEX: Reporting the experience. TUGboat, 29(1): 157–162, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-1/tb91hufflen-mlbibtex.pdf.

Kastrup:TB29-1-163

[3261] David Kastrup. Writing (La)TEX documents with AUCTEX in Emacs. TUGboat, 29(1):163, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-1/tb91kastrup-emacs.pdf.

Luczak:TB29-1-164

[3262] Tomasz �uczak. LYX: An editor not just for secretaries. TUGboat, 29(1): 164–165, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-1/tb91luczak-lyx.pdf.

Szabo:TB29-1-166

[3263] Péter Szabó. Automated DVD menu authoring with pdfL^{A}TEX. TUGboat, 29(1): 166–175, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-1/tb91szabo-dvd.pdf.

Walczak:TB29-1-176

[3264] Zofia Walczak. Graphics in L^{A}TEX using TikZ. TUGboat, 29(1):176–179,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91walczak.pdf.

Murzynowski:TB29-1-180

[3265] Grzegorz Murzynowski. L^{A}TEX vs. gmL^{A}TEXplain — a modification of the logo. TUGboat, 29(1):180, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91murzynowski-logo.pdf.

Kastrup:TB29-1-181

[3266] David Kastrup. Benefits, care and feeding of the bigfoot package. TUGboat, 29(1):181–183, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91kastrup-bigfoot.pdf.

Grosse:TB29-1-184

[3267] Johannes Große. MathPSfrag: L^{A}TEX labels in Mathematica plots. TUGboat, 29(1):184–189, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91grosse.pdf.

Kastrup:TB29-1-190

[3268] David Kastrup. makematch, a L^{A}TEX package for pattern matching with wildcards. TUGboat, 29(1):190–192, 2008. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb29-1/tb91kastrup-match.pdf.

Kastrup:TB29-1-193

[3269] David Kastrup. qstest, a L^{A}TEX package for unit tests. TUGboat, 29(1):193–

198, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-
1/tb91kastrup-qstest.pdf.

Murzynowski:TB29-1-199

[3270] Grzegorz Murzynowski. gmverse and gmcontinuo — some nontrivial placement of text on a page. TUGboat, 29(1):199–200, 2008. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb29-1/tb91murzynowski-
text.pdf.

Murzynowski:TB29-1-201

[3271] Grzegorz Murzynowski. The gmdoc bundle — a new tool for documenting (La)TEX sources. TUGboat, 29(1):201–206, 2008. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb29-1/tb91murzynowski-
gmdoc.pdf.

Jackowski:TB29-1-207

[3272] Pawe� Jackowski. TEX beauties and oddities: a permanent call for TEX pearls. TUGboat, 29(1):207–215, 2008. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb29-1/tb91pearls.pdf.

Nowacki:TB29-1-216

[3273] Janusz Nowacki. Crossword. TUGboat, 29(1):216, 2008. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-1/tb91nowacki-
crossword.pdf.

Anonymous:TB29-1-217

[3274] Anonymous. Abstracts (Hagen, Hoekwater, Kastrup, Lotz, Moore, Ryćko, Tomaszewski). TUGboat, 29(1):217–218, 2008. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb29-1/tb91abstracts.pdf.

Mahajan:TB29-1-219

[3275] Aditya Mahajan. ConTEXt basics for users: Table macros II. TUGboat, 29(1): 219–222, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-1/tb91mahajan.pdf.

Anonymous:TB29-1-223

[3276] Anonymous. TUG institutional members. TUGboat, 29(1):223, 2008. ISSN 0896-3207.

Anonymous:TB29-1-223-2

[3277] Anonymous. TEX consulting and production services. TUGboat, 29(1):223, 2008. ISSN 0896-3207.

Laugier:TB29-1-224

[3278] Maurice Laugier. In memoriam Bernard Gaulle. TUGboat, 29(1):224,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91laugier.pdf.

Anonymous:2008:FCa

[3279] Anonymous. Front cover. TUGboat, 29(1):c1, 2008. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-1/cover1.pdf.

Anonymous:2008:IFCa

[3280] Anonymous. Inside front cover. TUGboat, 29(1):c2, 2008. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
cover2.pdf.

Anonymous:2008:TPb

[3281] Anonymous. Title page. TUGboat, 29(2):229, 2008. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-2/titlepage.pdf.

Berry:TB29-2-231

[3282] Karl Berry. From the President. TUGboat, 29(2):231, 2008. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb29-2/tb92pres.pdf.

Beeton:TB29-2-232

[3283] Barbara Beeton. Editorial comments. TUGboat, 29(2):232–233, 2008. ISSN

0896-3207. URL https://tug.org/TUGboat/tb29-2/tb92beet.
pdf.

Knuth:TB29-2-233

[3284] Donald Knuth. The TEX tuneup of 2008. TUGboat, 29(2):233–238,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
tb92knut.pdf.

Beeton:TB29-2-239

[3285] Barbara Beeton. Hyphenation exception log. TUGboat, 29(2):239, 2008. ISSN

0896-3207. URL https://tug.org/TUGboat/tb29-2/tb92hyf.
pdf.

Flynn:TB29-2-240

[3286] Peter Flynn. Typographers’ Inn. TUGboat, 29(2):240–241, 2008. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-2/tb92inn.pdf.

Tsagkalos:TB29-2-242

[3287] Vassilios Tsagkalos. The Greek Font Society. TUGboat, 29(2):242–245,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
tb92tsagkalos.pdf.

Beccari:TB29-2-246

[3288] Claudio Beccari and Andrea Guadagni. Designing and producing a reference book with L^{A}TEX: The Engineer’s Quick Reference Handbook. TUGboat, 29(2):246–254, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-2/tb92becc-refdesign.pdf.

Guiggiani:TB29-2-255

[3289] Massimo Guiggiani and Lapo Mori. Suggestions on how not to mishandle mathematical formulæ. TUGboat, 29(2):255–263, 2008. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb29-2/tb92guiggiani.pdf.

Rankin:TB29-2-264

[3290] John Rankin. Wikipublisher: A Web-based system to make online and print versions of the same content. TUGboat, 29(2):264–269, 2008. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-2/tb92rankin.
pdf.

Eijkhout:TB29-2-270

[3291] Victor Eijkhout. Character encoding. TUGboat, 29(2):270–277, 2008. ISSN

0896-3207. URL https://tug.org/TUGboat/tb29-2/tb92eijkhout.
pdf.

Beccari:TB29-2-278

[3292] Claudio Beccari. lxfonts: L^{A}TEX slide fonts revived. TUGboat, 29(2):278–282,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
tb92becc-slidefonts.pdf.

Hagen:TB29-2-283

[3293] Hans Hagen, Taco Hoekwater, and Volker R. W. Schaa. Reshaping Euler: a collaboration with Hermann Zapf. TUGboat, 29(2):283–287,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
tb92hagen-euler.pdf.

Bowman:TB29-2-288

[3294] John Bowman and Andy Hammerlindl. Asymptote: a vector graphics language. TUGboat, 29(2):288–294, 2008. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb29-2/tb92bowman.pdf.

Hagen:TB29-2-295

[3295] Hans Hagen. The Luafication of TEX and ConTEXt. TUGboat, 29(2):295–302,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
tb92hagen-luatex.pdf.

Barrett:TB29-2-303

[3296] Edward Barrett. Porting TEX Live to OpenBSD. TUGboat, 29(2):303–304,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
tb92barrett.pdf.

Pakin:TB29-2-305

[3297] Scott Pakin. Good things come in little packages: an introduction to writing .ins and .dtx files. TUGboat, 29(2):305–314, 2008. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb29-2/tb92pakin.pdf.

Mahajan:TB29-2-315

[3298] Aditya Mahajan. ConTEXt basics for users: Indentation. TUGboat, 29(2): 315–316, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-2/tb92mahajan.pdf.

Roegel:TB29-2-317

[3299] Denis Roegel. Kanji–Sudokus: Integrating Chinese and graphics. TUGboat, 29(2):317–319, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-2/tb92roegel.pdf.

Jackowski:TB29-2-320

[3300] Pawe� Jackowski. Interesting loops and iterations — second helping. TUGboat, 29(2):320–323, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-2/tb92jackowski.pdf.

Wilson:TB29-2-324

[3301] Peter Wilson. Glisterings: More on paragraphs regular; L^{A}TEX’s defining triumvirate; TEX’s dictator. TUGboat, 29(2):324–327, 2008. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb29-2/tb92glister.pdf.

Berry:TB29-2-328

[3302] Karl Berry. The treasure chest. TUGboat, 29(2):328–330, 2008. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-2/tb92chest.pdf.

Vieth:TB29-2-331

[3303] Ulrik Vieth. Book review: Fonts & Encodings by Yannis Haralambous. TUGboat, 29(2):331–332, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-2/tb92vieth.pdf.

Schroeder:TB29-2-333

[3304] Bernd Schroeder. Software review: TEXCAD for Windows. TUGboat, 29(2): 333–334, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-2/tb92schroeder.pdf.

Mittelbach:TB29-2-334

[3305] Frank Mittelbach. \looseness on the loose. TUGboat, 29(2):334,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
tb92mitt.pdf.

Anonymous:TB29-2-335

[3306] Anonymous. MAPS: contents of issue 35 (2007). TUGboat, 29(2):335,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
tb92maps.pdf.

Anonymous:TB29-2-336

[3307] Anonymous. Die TEXnische Komödie: contents of issues 2007/2–2008/1. TUGboat, 29(2):336, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-2/tb92komo.pdf.

Anonymous:TB29-2-337

[3308] Anonymous. Eutypon: contents of issues 16–20 (2007–2008). TUGboat, 29(2):337–338, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-2/tb92eutypon.pdf.

Anonymous:TB29-2-338

[3309] Anonymous. ArsTEXnica: contents of issue 4 (2007). TUGboat, 29(2):338–

339, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-
2/tb92arstexnica.pdf.

Anonymous:TB29-2-339

[3310] Anonymous. The PracTEX Journal: contents of issues 2007-3–2008-1. TUGboat, 29(2):339–342, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-2/tb92pracjourn.pdf.

Walden:TB29-2-343

[3311] David Walden. TUG financial statements for 2007. TUGboat, 29(2):343–344,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
tb92treas.pdf.

Anonymous:TB29-2-344

[3312] Anonymous. TUG institutional members. TUGboat, 29(2):344, 2008. ISSN 0896-3207.

Anonymous:TB29-2-345

[3313] Anonymous. TUG membership form. TUGboat, 29(2):345, 2008. ISSN 0896-

3207. URL https://tug.org/join.html.

Anonymous:TB29-2-346

[3314] Anonymous. Calendar. TUGboat, 29(2):346, 2008. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb29-2/tb92calendar.pdf.

Anonymous:TB29-2-347

[3315] Anonymous. TUG 2008 announcement. TUGboat, 29(2):347, 2008. ISSN

0896-3207. URL https://tug.org/TUGboat/tb29-2/tb92tug08.
pdf.

Anonymous:TB29-2-348

[3316] Anonymous. TEX consulting and production services. TUGboat, 29(2):348,

2008. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2008:CODa

[3317] Anonymous. Contents ordered by difficulty. TUGboat, 29(2):c3, 2008. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-
2/cover3.pdf.

Anonymous:2008:FCb

[3318] Anonymous. Front cover. TUGboat, 29(2):c1, 2008. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-2/cover1.pdf.

Anonymous:2008:IFCb

[3319] Anonymous. Inside front cover. TUGboat, 29(2):c2, 2008. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-2/
cover2.pdf.

Anonymous:2008:TPc

[3320] Anonymous. Title page. TUGboat, 29(3):349, 2008. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-3/titlepage.pdf.

Anonymous:TB29-3-350

[3321] Anonymous. Conference program, delegates, and sponsors. TUGboat, 29(3): 350–351, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-3/tb93confinfo.pdf.

Anonymous:2008:CP

[3322] Anonymous. Conference program. TUGboat, 29(3):351, 2008. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-3/
tb93program.pdf.

Flynn:TB29-3-352

[3323] Peter Flynn. TUG 2008: TEX’s 30^{th} birthday. TUGboat, 29(3):352–355,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-3/
tb93flynn.pdf.

Mansfield:TB29-3-356

[3324] Niall Mansfield. How to develop your own document class — our experience. TUGboat, 29(3):356–361, 2008. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb29-3/tb93mansfield.pdf.

Kew:TB29-3-362

[3325] Jonathan Kew. TEXworks: Lowering the barrier to entry. TUGboat, 29(3): 362–364, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-3/tb93kew.pdf.

Laurens:TB29-3-365

[3326] Jérôme Laurens. Direct and reverse synchronization with SyncTEX. TUGboat, 29(3):365–371, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-3/tb93laurens.pdf.

Schrod:TB29-3-372

[3327] Joachim Schrod. Xindy revisited: Multi-lingual index creation for the UTF-8 age. TUGboat, 29(3):372–375, 2008. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb29-3/tb93schrod.pdf.

McCool:TB29-3-376

[3328] Joe McCool. A newbie’s experiences with Lilypond, Lilypond-book, L^{A}TEX and Perl. TUGboat, 29(3):376–379, 2008. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb29-3/tb93mccool.pdf.

Hoekwater:TB29-3-380

[3329] Taco Hoekwater. MetaPost developments: MPlib project report. TUGboat, 29(3):380–382, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-3/tb93hoekwater.pdf.

Hagen:TB29-3-383

[3330] Hans Hagen. The TEX–Lua mix. TUGboat, 29(3):383–391, 2008. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-3/tb93hagen-
mix.pdf.

Pocza:TB29-3-392

[3331] Krisztián Pócza, Mihály Biczó, and Zoltán Porkoláb. docx2tex: Word 2007 to TEX. TUGboat, 29(3):392–400, 2008. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb29-3/tb93pocza.pdf.

Hufflen:TB29-3-401

[3332] Jean-Michel Hufflen. Languages for bibliography styles. TUGboat, 29(3): 401–412, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-3/tb93hufflen.pdf.

Rowley:TB29-3-413

[3333] Chris Rowley. Vistas for TEX: liberate the typography! (Part I). TUGboat, 29(3):413–417, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-3/tb93rowley.pdf.

Crossland:TB29-3-418

[3334] Dave Crossland. Why didn’t catch on? TUGboat, 29(3):418–420,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-3/
tb93crossland.pdf.

Piska:TB29-3-421

[3335] Karel Píška. Creating cuneiform fonts with MetaType1 and FontForge. TUGboat, 29(3):421–425, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-3/tb93piska.pdf.

Vieth:TB29-3-426

[3336] Ulrik Vieth. Do we need a ‘Cork’ math font encoding? TUGboat, 29(3): 426–434, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-3/tb93vieth.pdf.

Sherif:TB29-3-435

[3337] Ameer Sherif and Hossam Fahmy. Meta-designing parameterized Arabic fonts for AlQalam. TUGboat, 29(3):435–443, 2008. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb29-3/tb93sherif.pdf.

Joshi:TB29-3-444

[3338] Manjusha Joshi. Smart ways of drawing PSTricks figures. TUGboat, 29(3): 444–445, 2008. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb29-3/tb93joshi-pstricks.pdf.

Hagen:TB29-3-446

[3339] Hans Hagen. The MetaPost library and LuaTEX. TUGboat, 29(3):446–453,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-3/
tb93hagen-mplib.pdf.

Miklavec:TB29-3-454

[3340] Mojca Miklavec and Arthur Reutenauer. Putting the Cork back in the bottle — Improving Unicode support in TEX. TUGboat, 29(3):454–457,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-3/
tb93miklavec.pdf.

Sarman:TB29-3-458

[3341] Stanislav Jan Šarman. Writing Gregg Shorthand with and L^{A}TEX. TUGboat, 29(3):458–461, 2008. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb29-3/tb93sarman.pdf.

Hagen:TB29-3-462

[3342] Hans Hagen. The LuaTEX way: \framed. TUGboat, 29(3):462–463,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-3/
tb93hagen-framed.pdf.

Moore:TB29-3-464

[3343] Ross Moore. Advanced features for publishing mathematics, in PDF and on the Web. TUGboat, 29(3):464–473, 2008. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb29-3/tb93moore.pdf.

Plaice:TB29-3-474

[3344] John Plaice, Blanca Mancilla, and Chris Rowley. Multidimensional text. TUGboat, 29(3):474–479, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-3/tb93plaice.pdf.

Joshi:TB29-3-480

[3345] Manjusha Joshi. Data mining: Role of TEX files. TUGboat, 29(3):480–481,

2008. ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-3/
tb93joshi-mining.pdf.

Anonymous:TB29-3-482

[3346] Anonymous. Abstracts (Fine, Hagen, Henkel, Hoekwater, Høgholm, Küster, Mancilla et al., Mittelbach, Peter, Rahilly et al., Rhatigan, Veytsman & Akhmadeeva, Veytsman). TUGboat, 29(3):482–484, 2008. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb29-3/tb93abstracts.pdf.

Anonymous:TB29-3-484

[3347] Anonymous. TUG institutional members. TUGboat, 29(3):484, 2008. ISSN 0896-3207.

Anonymous:TB29-3-485

[3348] Anonymous. Calendar. TUGboat, 29(3):485, 2008. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb29-3/tb93calendar.pdf.

Anonymous:TB29-3-486

[3349] Anonymous. TUG 2009 announcement. TUGboat, 29(3):486, 2008. ISSN

0896-3207. URL https://tug.org/TUGboat/tb29-3/tb93tug09.
pdf.

Anonymous:TB29-3-487

[3350] Anonymous. TEX consulting and production services. TUGboat, 29(3):487,

2008. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB29-3-488

[3351] Anonymous. TUG 2009 election. TUGboat, 29(3):488, 2008. ISSN 0896-

3207. URL https://www.tug.org/election/.

Anonymous:2008:CODb

[3352] Anonymous. Contents ordered by difficulty. TUGboat, 29(3):c3, 2008. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-
3/cover3.pdf.

Anonymous:2008:FCc

[3353] Anonymous. Front cover. TUGboat, 29(3):c1, 2008. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb29-3/cover1.pdf.

Anonymous:2008:IFCc

[3354] Anonymous. Inside front cover. TUGboat, 29(3):c2, 2008. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-3/
cover2.pdf.

Anonymous:2009:TPa

[3355] Anonymous. Title page. TUGboat, 30(1):1, 2009. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb30-1/titlepage.pdf.

Berry:TB30-1-3

[3356] Karl Berry. From the President. TUGboat, 30(1):3, 2009. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb30-1/tb94pres.pdf.

Beeton:TB30-1-4

[3357] Barbara Beeton. Editorial comments. TUGboat, 30(1):4, 2009. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb30-1/tb94beet.pdf.

Daly:TB30-1-4

[3358] Patrick Daly. Helmut Kopka, 1932–2009. TUGboat, 30(1):4–5, 2009. ISSN

0896-3207. URL https://tug.org/TUGboat/tb30-1/tb94daly.
pdf.

Cho:TB30-1-6

[3359] Jin-Hwan Cho. DVI specials for PDF generation. TUGboat, 30(1):6–11,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
tb94cho.pdf.

Perry:TB30-1-12

[3360] David Perry. Ancient TEX: Using XETEX to support classical and medieval studies. TUGboat, 30(1):12, 2009. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb30-1/tb94perry.pdf.

Prichystal:TB30-1-18

[3361] Jan Přichystal. TEXonWeb. TUGboat, 30(1):18–19, 2009. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb30-1/tb94prichystal.pdf.

Flynn:TB30-1-20

[3362] Peter Flynn. Typographers’ Inn. TUGboat, 30(1):20–21, 2009. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb30-1/tb94inn.pdf.

Vieth:TB30-1-22

[3363] Ulrik Vieth. OpenType math illuminated. TUGboat, 30(1):22, 2009. ISSN

0896-3207. URL https://tug.org/TUGboat/tb30-1/tb94vieth.
pdf.

Thanh:TB30-1-32

[3364] Hàn Thê´ Thành. A closer look at TrueType fonts and pdfTEX. TUGboat, 30 (1):32–34, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-1/tb94thanh.pdf.

Crossland:TB30-1-35

[3365] Dave Crossland. The Open Font Library. TUGboat, 30(1):35, 2009. ISSN

0896-3207. URL https://tug.org/TUGboat/tb30-1/tb94crossland.
pdf.

Mori:TB30-1-36

[3366] Lapo Mori. Managing bibliographies with L^{A}TEX. TUGboat, 30(1):36–48,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
tb94mori.pdf.

Hufflen:TB30-1-49

[3367] Jean-Michel Hufflen. Managing languages within MlBibTEX. TUGboat, 30 (1):49–57, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-1/tb94hufflen.pdf.

Bowman:TB30-1-58

[3368] John Bowman and Orest Shardt. Asymptote: Lifting TEX to three dimensions. TUGboat, 30(1):58–63, 2009. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb30-1/tb94bowman.pdf.

Zheng:TB30-1-64

[3369] Wentao Zheng. Supporting layout routines in MetaPost. TUGboat, 30(1):64–

68, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-
1/tb94zheng.pdf.

Wilson:TB30-1-69

[3370] Peter Wilson. Glisterings: Reprise; MetaPost and pdfL^{A}TEX; Spidrons. TUGboat, 30(1):69–73, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-1/tb94glister.pdf.

Roegel:TB30-1-74

[3371] Denis Roegel. MetaPost macros for drawing Chinese and Japanese abaci. TUGboat, 30(1):74–79, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-1/tb94roegel-abacus.pdf.

Roegel:TB30-1-80

[3372] Denis Roegel. Spheres, great circles and parallels. TUGboat, 30(1):80–87,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
tb94roegel-spheres.pdf.

Roegel:TB30-1-88

[3373] Denis Roegel. An introduction to nomography: Garrigues’ nomogram for the computation of Easter. TUGboat, 30(1):88–104, 2009. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb30-1/tb94roegel-
garrigues.pdf.

Team:TB30-1-10

[3374] L^{A}TEX Project Team. L^{A}TEX3 news, issues 1–2. TUGboat, 30(1):105–106,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
tb94l3news.pdf.

Wright:TB30-1-107

[3375] Joseph Wright. L^{A}TEX3 programming: External perspectives. TUGboat, 30(1): 107–109, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-1/tb94wright-latex3.pdf.

Wright:TB30-1-110

[3376] Joseph Wright and Christian Feuersänger. Implementing key–value input: an introduction. TUGboat, 30(1):110–122, 2009. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb30-1/tb94wright-
keyval.pdf.

Zyka:TB30-1-123

[3377] Vít Zýka. Current typesetting position in pdfTEX. TUGboat, 30(1):123–124,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
tb94zyka.pdf.

Tai:TB30-1-125

[3378] Kaihsu Tai. In response to “mathematical formulæ”. TUGboat, 30(1):125,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
tb94tai.pdf.

Guiggiani:TB30-1-126

[3379] Massimo Guiggiani and Lapo Mori. In response to Kaihsu Tai. TUGboat, 30(1):126, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-1/tb94guiggiani.pdf.

Berry:TB30-1-127

[3380] Karl Berry. The treasure chest. TUGboat, 30(1):127, 2009. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb30-1/tb94chest.pdf.

Anonymous:TB30-1-131

[3381] Anonymous. ArsTEXnica: contents of issues 5–7 (2008–2009). TUGboat, 30(1):131–132, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-1/tb94arstexnica.pdf.

Anonymous:TB30-1-133

[3382] Anonymous. Baskerville: contents of issue 10.1 (2009). TUGboat, 30(1):133,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
tb94eutypon.pdf.

Anonymous:TB30-1-134

[3383] Anonymous. Die TEXnische Komödie: contents of issues 2008/2–2009/2. TUGboat, 30(1):134–135, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-1/tb94komo.pdf.

Anonymous:TB30-1-135

[3384] Anonymous. Eutypon: contents of issue 21 (2008). TUGboat, 30(1):135, 2009. ISSN 0896-3207.

Anonymous:TB30-1-136

[3385] Anonymous. nine MAPS: contents of issue 36–37 (2008). TUGboat, 30(1): 136–137, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-1/tb94maps.pdf.

Anonymous:TB30-1-138

[3386] Anonymous. The PracTEX Journal: contents of issues 2008-2–2008-3. TUGboat, 30(1):138, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-1/tb94pracjourn.pdf.

Anonymous:TB30-1-139

[3387] Anonymous. TEXemplares: contents of issue 8 (2006). TUGboat, 30(1):139,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
tb94texemp.pdf.

Anonymous:TB30-1-140

[3388] Anonymous. Zpravodaj: contents of issues 16(2)–18(4) (2006–2008). TUGboat, 30(1):140–143, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-1/tb94zprav.pdf.

Anonymous:TB30-1-144

[3389] Anonymous. TUG institutional members. TUGboat, 30(1):144, 2009. ISSN 0896-3207.

Beeton:TB30-1-145

[3390] Barbara Beeton. TUG 2009 election report. TUGboat, 30(1):145, 2009. ISSN

0896-3207. URL https://tug.org/TUGboat/tb30-1/tb94elec.
pdf.

TDF:TB30-1-148

[3391] TEX Development Fund committee. TEX Development Fund 2009 report. TUGboat, 30(1):148, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-1/tb94devfund.pdf.

Walden:TB30-1-149

[3392] David Walden. TUG financial statements for 2009. TUGboat, 30(1):149–150,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
tb94treas.pdf.

Anonymous:TB30-1-151

[3393] Anonymous. Calendar. TUGboat, 30(1):151, 2009. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb30-1/tb94calendar.pdf.

Anonymous:TB30-1-152

[3394] Anonymous. TEX consulting and production services. TUGboat, 30(1):152,

2009. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2009:CODa

[3395] Anonymous. Contents ordered by difficulty. TUGboat, 30(1):c3, 2009. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-
1/cover3.pdf.

Anonymous:2009:FCa

Anonymous:2009:IFCa

[3397] Anonymous. Inside front cover. TUGboat, 30(1):c2, 2009. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-1/
cover2.pdf.

Anonymous:2009:TPb

[3398] Anonymous. Title page. TUGboat, 30(2):153, 2009. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb30-2/titlepage.pdf.

Anonymous:TB30-2-154

[3399] Anonymous. Conference program, delegates, and sponsors. TUGboat, 30(2):

154, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-
2/tb95confinfo.pdf.

Walden:TB30-2-159

[3400] David Walden. Profile of Eitan Gurari (1947–2009). TUGboat, 30(2):159–162,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95gurari.pdf.

Veytsman:TB30-2-163

[3401] Boris Veytsman. L^{A}TEX class writing for wizard apprentices. TUGboat, 30(2): 163–168, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-2/tb95veytsman-classes.pdf.

Reutenauer:TB30-2-169

[3402] Arthur Reutenauer. LuaTEX for the L^{A}TEX user: an introduction. TUGboat, 30(2):169, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-2/tb95reutenauer.pdf.

Moore:TB30-2-170

[3403] Ross Moore. Ongoing efforts to generate “tagged PDF” using pdfTEX. TUGboat, 30(2):170–175, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-2/tb95moore.pdf.

Quinn:TB30-2-176

[3404] Frank Quinn. The EduTEXTUG working group. TUGboat, 30(2):176,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95quinn.pdf.

Hefferon:TB30-2-177

[3405] Jim Hefferon. Becoming a CTAN mirror. TUGboat, 30(2):177–178,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95heff-mirrors.pdf.

Berry:TB30-2-179

[3406] Karl Berry. TEX Live 2009 news. TUGboat, 30(2):179, 2009. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb30-2/tb95berry-tl.pdf.

Arnold:TB30-2-180

[3407] Tim Arnold. Getting started with plasTEX. TUGboat, 30(2):180–182,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95arnold.pdf.

Hagen:TB30-2-183

[3408] Hans Hagen. LuaTEX: Halfway to version 1. TUGboat, 30(2):183–186,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95hagen-luatex.pdf.

Hagen:TB30-2-187

[3409] Hans Hagen. LuaTEX and ConTEXt: Where we stand. TUGboat, 30(2):187–

190, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-
2/tb95hagen-stand.pdf.

Neveln:TB30-2-191

[3410] Bob Neveln and Bob Alps. ProofCheck: Writing and checking complete proofs in L^{A}TEX. TUGboat, 30(2):191–195, 2009. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb30-2/tb95neveln.pdf.

Berry:TB30-2-196

[3411] Karl Berry and David Walden. TEX People: The TUG interviews project and book. TUGboat, 30(2):196–202, 2009. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb30-2/tb95berry-interviews.pdf.

Walden:TB30-2-203

[3412] David Walden. Self-publishing: Experiences and opinions. TUGboat, 30(2): 203–208, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-2/tb95walden.pdf.

Hoppner:TB30-2-209

[3413] Klaus Höppner. Introduction to . TUGboat, 30(2):209–213,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95hoeppner.pdf.

Mertz:TB30-2-214

[3414] Andrew Mertz and William Slough. A TikZ tutorial: Generating graphics in the spirit of TEX. TUGboat, 30(2):214–226, 2009. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb30-2/tb95mertz.pdf.

Veytsman:TB30-2-227

[3415] Boris Veytsman and Leila Akhmadeeva. Medical pedigrees: Typography and interfaces. TUGboat, 30(2):227–235, 2009. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb30-2/tb95veytsman-pedigree.pdf.

Hefferon:TB30-2-236

[3416] Jim Hefferon. A first look at the TEX Gyre fonts. TUGboat, 30(2):236–240,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95heff-gyre.pdf.

Hagen:TB30-2-241

[3417] Hans Hagen. Plain TEX and OpenType. TUGboat, 30(2):241–242,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95hagen-opentype.pdf.

Mahajan:TB30-2-243

[3418] Aditya Mahajan. Integrating Unicode and OpenType math in ConTEXt. TUGboat, 30(2):243–246, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-2/tb95mahajan-cmath.pdf.

Mahajan:TB30-2-247

[3419] Aditya Mahajan. LuaTEX: a user’s perspective. TUGboat, 30(2):247–251,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95mahajan-luatex.pdf.

Beebe:TB30-2-252

[3420] Nelson H. F. Beebe. BibTEX meets relational databases. TUGboat, 30(2): 252–271, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-2/tb95beebe.pdf.

Bazargan:TB30-2-272

[3421] Kaveh Bazargan. TEX as an eBook reader. TUGboat, 30(2):272–273,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95bazargan.pdf.

Rossi:2009:DPD

[3422] Christian Rossi. From distribution to preservation of digital documents. TUGboat, 30(2):274, 2009. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb30-2/tb95rossi.pdf.

Story:TB30-2-281

[3423] D. P. Story. Rich media annotations and AcroFLeX. TUGboat, 30(2):281–284,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
tb95story.pdf.

Beccari:TB30-2-285

[3424] Claudio Beccari. Use of the \write18 feature for composing indexes. TUGboat, 30(2):285–286, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-2/tb95becc.pdf.

Wilson:TB30-2-287

[3425] Peter Wilson. Glisterings: Repetition; Rectangular text. TUGboat, 30(2):287–

289, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-
2/tb95glister.pdf.

Robertson:TB30-2-290

[3426] Will Robertson. Peter Wilson’s Herries Press packages. TUGboat, 30(2):290–

292, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-
2/tb95robertson.pdf.

Anonymous:TB30-2-293

[3427] Anonymous. Biuletyn GUST: contents of issues 25–26 (2009). TUGboat, 30(2):293–296, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-2/tb95gust.pdf.

Anonymous:TB30-2-296

[3428] Anonymous. Die TEXnische Komödie: contents of issue 2009/3. TUGboat, 30(2):296, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-2/tb95zprav.pdf.

Anonymous:TB30-2-296-2

[3429] Anonymous. Zpravodaj: contents of issue 19(1–2) (2009). TUGboat, 30(2): 296–297, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-2/tb95zprav.pdf.

Anonymous:TB30-2-298

[3430] Anonymous. TUG 2009 abstracts (Cho, de Souza, Hamid, Høgholm, Rowley). TUGboat, 30(2):298–299, 2009. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb30-2/tb95abstracts.pdf.

Anonymous:TB30-2-299

[3431] Anonymous. TUG institutional members. TUGboat, 30(2):299, 2009. ISSN 0896-3207.

Anonymous:TB30-2-300

[3432] Anonymous. Cheryl Ponchin Training; O’Reilly Media; Design Science. TUGboat, 30(2):300, 2009. ISSN 0896-3207.

Anonymous:TB30-2-301

[3433] Anonymous. MacKichan Software; River Valley Technologies; A-R Editions. TUGboat, 30(2):301, 2009. ISSN 0896-3207.

Anonymous:TB30-2-302

[3434] Anonymous. TUG 2010 announcement. TUGboat, 30(2):302, 2009. ISSN

0896-3207. URL https://www.tug.org/tug2010/.

Anonymous:TB30-2-303

[3435] Anonymous. Calendar. TUGboat, 30(2):303, 2009. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb30-2/tb95calendar.pdf.

Anonymous:TB30-2-304

[3436] Anonymous. TEX consulting and production services. TUGboat, 30(2):304,

2009. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2009:CODb

[3437] Anonymous. Contents ordered by difficulty. TUGboat, 30(2):c3, 2009. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-
2/cover3.pdf.

Anonymous:2009:FCb

[3438] Anonymous. Front cover. TUGboat, 30(2):c1, 2009. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb30-2/cover1.pdf.

Anonymous:2009:IFCb

[3439] Anonymous. Inside front cover. TUGboat, 30(2):c2, 2009. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-2/
cover2.pdf.

Hagen:TB30-3-1

[3440] Hans Hagen. EuroTEX 2009. TUGboat, 30(3):1, 2009. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb30-3/tb96hagen-e2009.pdf.

Anonymous:TB30-3-2

[3441] Anonymous. Announcement: EuroTEX conference 2010. TUGboat, 30(3):2,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-3/
tb96eurotex10.pdf.

Anonymous:TB30-3-3

[3442] Anonymous. Announcement: 4^{th} international ConTEXt meeting 2010. TUGboat, 30(3):3, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-3/tb96context10.pdf.

Anonymous:TB30-3-4

[3443] Anonymous. Announcement: TUG conference 2010. TUGboat, 30(3):4, 2009. ISSN 0896-3207.

Anonymous:TB30-3-4-2

[3444] Anonymous. TUG institutional members. TUGboat, 30(3):4, 2009. ISSN 0896-3207.

Laan:TB30-3-5

[3445] Kees van der Laan. TEX education — a neglected approach. TUGboat, 30 (3):5–33, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-3/tb96laan.pdf.

Scarso:TB30-3-34

[3446] Luigi Scarso. LuaTEX lunatic. TUGboat, 30(3):34–58, 2009. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb30-3/tb96scarso.pdf.

Egger:TB30-3-59

[3447] Willi Egger. Decorating CD-ROMs and DVDs. TUGboat, 30(3):59–71,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-3/
tb96egger.pdf.

Hagen:TB30-3-72

[3448] Hans Hagen. The language mix. TUGboat, 30(3):72–77, 2009. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb30-3/tb96hagen-
mix.pdf.

Huisman:TB30-3-78

[3449] Jelle Huisman. E16 & DEtool — typesetting language data using ConTEXt. TUGboat, 30(3):78–85, 2009. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb30-3/tb96huisman.pdf.

Kroonenberg:TB30-3-86

[3450] Siep Kroonenberg. A network TEXLive installation at the University of Groningen. TUGboat, 30(3):86–91, 2009. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb30-3/tb96kroonenberg.pdf.

Hufflen:TB30-3-92

[3451] Jean-Michel Hufflen. Using TEX’s language within a course about functional programming. TUGboat, 30(3):92–98, 2009. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb30-3/tb96hufflen-func.pdf.

Aebischer:TB30-3-99

[3452] Anne-Marie Aebischer, Bruno Aebischer, Jean-Michel Hufflen, and François Pétiard. Introducing new French-speaking users to L^{A}TEX quickly and convincingly. TUGboat, 30(3):99–104, 2009. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb30-3/tb96aebischer.pdf.

Hagen:TB30-3-105

[3453] Hans Hagen. Oriental TEX by a dummy. TUGboat, 30(3):105–106,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-3/
tb96hagen-otex.pdf.

Sarman:TB30-3-107

[3454] Stanislav Jan Šarman. Writing Pitman shorthand with and L^{A}TEX. TUGboat, 30(3):107–111, 2009. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb30-3/tb96sarman.pdf.

Szabo:TB30-3-112

[3455] Péter Szabó. Optimizing PDF output size of TEX documents. TUGboat, 30(3): 112–130, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-3/tb96szabo.pdf.

Hoekwater:TB30-3-131

[3456] Taco Hoekwater. Generating PDF for e-reader devices. TUGboat, 30(3):131–

135, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-
3/tb96hoekwater-ereader.pdf.

Hoekwater:TB30-3-136

[3457] Taco Hoekwater. LuaTEX says goodbye to Pascal. TUGboat, 30(3):136–140,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-3/
tb96hoekwater-pascal.pdf.

Hoekwater:TB30-3-141

[3458] Taco Hoekwater. The typesetting of statistics. TUGboat, 30(3):141–142,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-3/
tb96hoekwater-statistics.pdf.

Hagen:TB30-3-143

[3459] Hans Hagen and Taco Hoekwater. MetaPost 2 project goals. TUGboat, 30(3): 143–144, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-3/tb96hagen-mp2.pdf.

Trapp:TB30-3-145

[3460] John Trapp. Using L^{A}TEX as a computing language. TUGboat, 30(3):145–165,

2009. ISSN 0896-3207. URL https://tug.org/TUGboat/tb30-3/
tb96trapp.pdf.

Vieth:TB30-3-166

[3461] Ulrik Vieth. Experiences typesetting mathematical physics. TUGboat, 30(3): 166–178, 2009. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb30-3/tb96vieth.pdf.

Anonymous:TB30-3-179

[3462] Anonymous. Abstracts (Hagen, Lemmes, Walczak, Fine, Kew, Yatskovsky, Schröder, Hoekwater, Ludwichowski, Jackowski, Píška, Miklavec & Reutenauer). TUGboat, 30(3):179–181, 2009. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb30-3/tb96abstracts.pdf.

Anonymous:TB30-3-182

[3463] Anonymous. Participant list. TUGboat, 30(3):182, 2009. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb30-3/tb96names-e2009.
pdf.

Anonymous:TB30-3-183

[3464] Anonymous. TEX consulting and production services. TUGboat, 30(3):183,

2009. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2010:TPa

[3465] Anonymous. Title page. TUGboat, 31(1):1, 2010. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-1/titlepage.pdf.

Berry:TB31-1-3

[3466] Karl Berry. From the President. TUGboat, 31(1):3, 2010. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb31-1/tb97pres.pdf.

Beeton:TB31-1-4

[3467] Barbara Beeton. Editorial comments. TUGboat, 31(1):4–5, 2010. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-1/tb97beet.pdf.

Wessler:TB31-1-6

[3468] Evan Wessler. An argument for learning L^{A}TEX: The benefits of typesetting and beyond. TUGboat, 31(1):6–8, 2010. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb31-1/tb97wessler.pdf.

Mars:TB31-1-9

[3469] Nicolaas Mars. A computer scientist self-publishing in the humanities. TUGboat, 31(1):9–11, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-1/tb97mars.pdf.

Isambert:TB31-1-12

[3470] Paul Isambert. Strategies against widows. TUGboat, 31(1):12–17,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97isambert.pdf.

Miede:TB31-1-18

[3471] André Miede. Theses and other beautiful documents with classicthesis. TUGboat, 31(1):18–20, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-1/tb97miede.pdf.

Flynn:TB31-1-21

[3472] Peter Flynn. Typographers’ Inn. TUGboat, 31(1):21–22, 2010. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-1/tb97inn.pdf.

Parashchenko:TB31-1-23

[3473] Oleg Parashchenko. Minimal setup for a (Cyrillic) TrueType font. TUGboat, 31 (1):23–25, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-1/tb97parashchenko.pdf.

Hagen:TB31-1-26

[3474] Hans Hagen. LuaTEX: Microtypography for plain fonts. TUGboat, 31(1):26,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97hagen-fonts.pdf.

Dhawan:TB31-1-27

[3475] Amit Raj Dhawan. Mathematical typefaces in TEX documents. TUGboat, 31 (1):27–31, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-1/tb97dhawan.pdf.

Hagen:TB31-1-32

[3476] Hans Hagen. LuaTEX: Deeply nested notes. TUGboat, 31(1):32–35,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97hagen-nested.pdf.

Voss:TB31-1-31

[3477] Herbert Voß. The current state of the PSTricks project. TUGboat, 31(1):36–49,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97voss.pdf.

Wright:TB31-1-50

[3478] Joseph Wright. Plotting experimental data using pgfplots. TUGboat, 31(1):50–

58, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-
1/tb97wright-pgfplots.pdf.

Kmiecik:TB31-1-59

[3479] Mateusz Kmiecik. From Logo to MetaPost. TUGboat, 31(1):59–63,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97kmiecik-logo.pdf.

LPT:TB31-1-64

[3480] L^{A}TEX Project Team. L^{A}TEX news, issue 19. TUGboat, 31(1):64,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97ltnews.pdf.

Talbot:TB31-1-65

[3481] Nicola Talbot. Talbot packages: an overview. TUGboat, 31(1):65–67,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97talbot.pdf.

Kmiecik:TB31-1-68

[3482] Jacek Kmiecik. Tuning L^{A}TEX to one’s own needs. TUGboat, 31(1):68–75,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97kmiecik-latex.pdf.

Merciadri:TB31-1-76

[3483] Luca Merciadri. Some misunderstood or unknown L^{A}TEX 2ε tricks. TUGboat, 31(1):76–78, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-1/tb97merciadri.pdf.

LPT:TB31-1-79

[3484] L^{A}TEX Project Team. L^{A}TEX3 news, issue 3. TUGboat, 31(1):79,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97l3news.pdf.

Wright:TB31-1-80

[3485] Joseph Wright. Beyond \newcommand with xparse. TUGboat, 31(1):80–82,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97wright-xparse.pdf.

Wright:TB31-1-83

[3486] Joseph Wright. Programming key–value in expl3. TUGboat, 31(1):83–87,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97wright-l3keys.pdf.

Mahajan:TB31-1-88

[3487] Aditya Mahajan. ConTEXt basics for users: Conditional processing. TUGboat, 31(1):88–89, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-1/tb97mahajan.pdf.

Wilson:TB31-1-90

[3488] Peter Wilson. Glisterings: Counting; Changing the layout. TUGboat, 31 (1):90–93, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-1/tb97glister.pdf.

Hall:TB31-1-94

[3489] Timothy Hall. The exact placement of superscripts and subscripts. TUGboat, 31(1):94–95, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-1/tb97hall.pdf.

Berry:TB31-1-96

[3490] Karl Berry. The treasure chest. TUGboat, 31(1):96–98, 2010. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-1/tb97chest.pdf.

Anonymous:TB31-1-99

[3491] Anonymous. ArsTEXnica: contents of issue 8 (October 2009). TUGboat, 31(1):

99, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-
1/tb97arstexnica.pdf.

Anonymous:TB31-1-100

[3492] Anonymous. Die TEXnische Komödie: contents of issues 2009/4–2010/2. TUGboat, 31(1):100, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-1/tb97komo.pdf.

Anonymous:TB31-1-101

[3493] Anonymous. Asian Journal of TEX: contents of Volume 3 (2009). TUGboat, 31(1):101–102, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-1/tb97ajt.pdf.

Anonymous:TB31-1-102

[3494] Anonymous. Les Cahiers GUTenberg: contents of issues 48–53 (2006–2009). TUGboat, 31(1):102–103, 2010. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb31-1/tb97cahi.pdf.

Anonymous:TB31-1-104

[3495] Anonymous. Eutypon: contents of issue 22–23 (2009). TUGboat, 31(1):104,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97eutypon.pdf.

Anonymous:TB31-1-105

[3496] Anonymous. MAPS: contents of issue 38–39 (2009). TUGboat, 31(1):105,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97baskerville.pdf.

Anonymous:TB31-1-105-2

[3497] Anonymous. Baskerville: contents of issue 10.2 (2009). TUGboat, 31(1):105,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97baskerville.pdf.

Anonymous:TB31-1-106

[3498] Anonymous. Zpravodaj: contents of issue 16(1), 19(3)–19(4) (2006, 2009). TUGboat, 31(1):106–108, 2010. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb31-1/tb97zprav.pdf.

Walden:TB31-1-109

[3499] David Walden. TUG financial statements for 2009. TUGboat, 31(1):109–110,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
tb97treas.pdf.

Anonymous:TB31-1-110

[3500] Anonymous. TUG institutional members. TUGboat, 31(1):110, 2010. ISSN 0896-3207.

Anonymous:TB31-1-111

[3501] Anonymous. Calendar. TUGboat, 31(1):111, 2010. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb31-1/tb97calendar.pdf.

Anonymous:TB31-1-112

[3502] Anonymous. TEX consulting and production services. TUGboat, 31(1):112,

2010. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2010:CODa

[3503] Anonymous. Contents ordered by difficulty. TUGboat, 31(1):c3, 2010. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-
1/cover3.pdf.

Anonymous:2010:FCa

[3504] Anonymous. Front cover. TUGboat, 31(1):c1, 2010. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-1/cover1.pdf.

Anonymous:2010:IFCa

[3505] Anonymous. Inside front cover. TUGboat, 31(1):c2, 2010. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-1/
cover2.pdf.

Anonymous:2010:TPb

[3506] Anonymous. Title page. TUGboat, 31(2):113, 2010. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-2/titlepage.pdf.

Anonymous:TB31-2-114

[3507] Anonymous. Conference program, delegates, and sponsors. TUGboat, 31(2): 114–116, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-2/tb98confinfo.pdf.

Walden:TB31-2-117

[3508] David Walden. TEX Users Group 2010 conference report. TUGboat, 31(2): 117–120, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-2/tb98walden.pdf.

Knuth:TB31-2-121

[3509] Donald Knuth. An earthshaking announcement. TUGboat, 31(2):121–124,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
tb98knut.pdf.

Walden:TB31-2-125

[3510] David Walden (moderator). TUG 2010 Panel: Don Knuth & Stanford TEX Project members. TUGboat, 31(2):125–137, 2010. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb31-2/tb98panel.pdf.

Beeton:TB31-2-138

[3511] Barbara Beeton. Thoughts on TUG 2010. TUGboat, 31(2):138–139,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
tb98beet.pdf.

Hefferon:TB31-2-143

[3512] Jim Hefferon. CTAN packages get keywords. TUGboat, 31(2):143–144,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
tb98heff.pdf.

Voss:TB31-2-145

[3513] Herbert Voß. From PostScript to PDF with epstopdf, pdftricks, pst-pdf, auto-pst-pdf, pst2pdf, and more. TUGboat, 31(2):145–

147, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-
2/tb98voss.pdf.

Hicks:TB31-2-148

[3514] Stephen Hicks. Improving margin paragraphs. TUGboat, 31(2):148–150,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
tb98hicks.pdf.

Arora:TB31-2-151

[3515] Pavneet Arora. Using L^{A}TEX to generate dynamic mathematics worksheets for the web. TUGboat, 31(2):151–153, 2010. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb31-2/tb98arora.pdf.

Gander:TB31-2-154

[3516] Walter Gander. Writing the first L^{A}TEX book. TUGboat, 31(2):154–157,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
tb98gander.pdf.

Hoenig:TB31-2-158

[3517] Alan Hoenig. TEX helps you learn Chinese character meanings. TUGboat, 31(2):158–161, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-2/tb98hoenig.pdf.

Verna:TB31-2-162

[3518] Didier Verna. Classes, styles, conflicts: The biological realm of L^{A}TEX. TUGboat, 31(2):162–172, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-2/tb98verna.pdf.

Berry:TB31-2-173

[3519] Karl Berry. TEX Live 2010 news. TUGboat, 31(2):173, 2010. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb31-2/tb98berry.pdf.

Hoekwater:TB31-2-174

[3520] Taco Hoekwater and Hartmut Henkel. LuaTEX 0.60: an overview of changes. TUGboat, 31(2):174–177, 2010. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb31-2/tb98hoekwater.pdf.

Hagen:TB31-2-178

[3521] Hans Hagen. LuaTEX: PDF merging. TUGboat, 31(2):178–179, 2010. ISSN

0896-3207. URL https://tug.org/TUGboat/tb31-2/tb98hagen-
pdfmerge.pdf.

Hagen:TB31-2-180

[3522] Hans Hagen. The TEX paragraph builder in Lua. TUGboat, 31(2):180–182,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
tb98hagen-parbuilder.pdf.

Childs:TB31-2-183

[3523] Bart Childs. Thirty years of literate programming and more? TUGboat, 31(2): 183–188, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-2/tb98childs.pdf.

Ziegenhagen:TB31-2-189

[3524] Uwe Ziegenhagen. Dynamic reporting with R/Sweave and L^{A}TEX. TUGboat, 31(2):189–192, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-2/tb98ziegenhagen.pdf.

Doob:TB31-2-193

[3525] Michael Doob. A web-based TEX previewer: The ecstasy and the agony. TUGboat, 31(2):193–196, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-2/tb98doob.pdf.

Hamid:TB31-2-197

[3526] Idris Samawi Hamid. Qur’ānic typography comes of age: Æsthetics, layering, and paragraph optimization in ConTEXt. TUGboat, 31(2):197–202,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
tb98hamid.pdf.

Bowman:TB31-2-203

[3527] John Bowman. Asymptote: Interactive TEX-aware 3D vector graphics. TUGboat, 31(2):203–205, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-2/tb98bowman.pdf.

Bourgeois:TB31-2-206

[3528] Mathieu Bourgeois and Roger Villemaire. Drawing structured diagrams with SDDL. TUGboat, 31(2):206–210, 2010. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb31-2/tb98bourgeois.pdf.

Robertson:TB31-2-211

[3529] Will Robertson. Unicode mathematics in L^{A}TEX: Advantages and challenges. TUGboat, 31(2):211–220, 2010. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb31-2/tb98robertson.pdf.

Kuster:TB31-2-221

[3530] Johannes Küster. Math never seen. TUGboat, 31(2):221–229, 2010. ISSN

0896-3207. URL https://tug.org/TUGboat/tb31-2/tb98kuester.
pdf.

Veytsman:TB31-2-230

[3531] Boris Veytsman. Are virtual fonts obsolete? TUGboat, 31(2):230–235,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
tb98veytsman-vf.pdf.

Veytsman:TB31-2-236

[3532] Boris Veytsman and Leila Akhmadeeva. TEX in the GLAMP world: On-demand creation of documents online. TUGboat, 31(2):236–239, 2010. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-2/tb98veytsman-
glamp.pdf.

Hammond:TB31-2-240

[3533] William Hammond. L^{A}TEX profiles as objects in the category of markup languages. TUGboat, 31(2):240–247, 2010. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb31-2/tb98hammond.pdf.

Anonymous:TB31-2-248

[3534] Anonymous. TUG 2010 abstracts (Bazargan, Cheswick, Doumont, Grathwohl & Ruddy, Hagen & Hoekwater, Hobby, Kew, Mittelbach, Moore, Rowley, Rundell). TUGboat, 31(2):248–249, 2010. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb31-2/tb98abstracts.pdf.

Anonymous:TB31-2-250

[3535] Anonymous. Addison-Wesley; von Hoerner&Sulger. TUGboat, 31(2):250,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
tb98sponsors.pdf.

Anonymous:TB31-2-251

[3536] Anonymous. Cheryl Ponchin Training; River Valley Technologies. TUGboat, 31(2):251, 2010. ISSN 0896-3207.

Anonymous:TB31-2-252

[3537] Anonymous. Principiae; University Science Books. TUGboat, 31(2):252, 2010. ISSN 0896-3207.

Anonymous:TB31-2-253

[3538] Anonymous. Calendar. TUGboat, 31(2):253, 2010. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb31-2/tb98calendar.pdf.

Anonymous:TB31-2-254

[3539] Anonymous. TEX consulting and production services. TUGboat, 31(2):254–

255, 2010. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB31-2-255

[3540] Anonymous. TUG institutional members. TUGboat, 31(2):255, 2010. ISSN 0896-3207.

Anonymous:TB31-2-256

[3541] Anonymous. TUG 2011 election. TUGboat, 31(2):256, 2010. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb31-2/tb98elec.pdf.

Anonymous:2010:CODb

[3542] Anonymous. Contents ordered by difficulty. TUGboat, 31(2):c3, 2010. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-
2/cover3.pdf.

Anonymous:2010:FCb

[3543] Anonymous. Front cover. TUGboat, 31(2):c1, 2010. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-2/cover1.pdf.

Anonymous:2010:IFCb

[3544] Anonymous. Inside front cover. TUGboat, 31(2):c2, 2010. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-2/
cover2.pdf.

Anonymous:2010:TPc

[3545] Anonymous. Title page. TUGboat, 31(3):157, 2010. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-3/titlepage.pdf.

Berry:TB31-3-158

[3546] Karl Berry. From the President. TUGboat, 31(3):158, 2010. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb31-3/tb99pres.pdf.

Beeton:TB31-3-158

[3547] Barbara Beeton. Editorial comments. TUGboat, 31(3):158–159, 2010. ISSN

0896-3207. URL https://tug.org/TUGboat/tb31-3/tb99beet.
pdf.

Beeton:TB31-3-160

[3548] Barbara Beeton. Hyphenation exception log. TUGboat, 31(3):160, 2010. ISSN

0896-3207. URL https://tug.org/TUGboat/tb31-3/tb99hyf.
pdf.

Caignaert:TB31-3-161

[3549] Christophe Caignaert. A story of kpfonts: Reaching the limits of NFSS. TUGboat, 31(3):161–174, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-3/tb99caignaert.pdf.

Hefferon:TB31-3-175

[3550] Jim Hefferon. Giving it away. TUGboat, 31(3):175–176, 2010. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-3/tb99heff.pdf.

Wilson:TB31-3-177

[3551] Peter Wilson. Glisterings: Meandering miniature books. TUGboat, 31(3): 177–183, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-3/tb99glister.pdf.

Isambert:TB31-3-184

[3552] Paul Isambert. Three things you can do with LuaTEX that would be extremely painful otherwise. TUGboat, 31(3):184–190, 2010. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb31-3/tb99isambert.pdf.

Merciadri:TB31-3-191

[3553] Luca Merciadri. Some misunderstood or unknown L^{A}TEX 2ε tricks II. TUGboat, 31(3):191–194, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-3/tb99merciadri.pdf.

LPT:TB31-3-194

[3554] L^{A}TEX Project Team. L^{A}TEX3 news, issue 3. TUGboat, 31(3):194,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-3/
tb99l3news.pdf.

Wright:TB31-3-195

[3555] Joseph Wright. From \newcommand to \documentnewcommand with

xparse. TUGboat, 31(3):195–196, 2010. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb31-3/tb99wright.pdf.

Hagen:TB31-3-197

[3556] Hans Hagen. Tagged PDF in ConTEXt. TUGboat, 31(3):197–202,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-3/
tb99hagen.pdf.

Scarso:TB31-3-203

[3557] Luigi Scarso. Introduction to colours in ConTEXt MKiV. TUGboat, 31(3): 203–207, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-3/tb99scarso.pdf.

Parashchenko:TB31-3-208

[3558] Oleg Parashchenko. Generate TEX documents using pdfscript. TUGboat, 31(3):208–218, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-3/tb99parashchenko.pdf.

Centonza:TB31-3-213

[3559] Matteo Centonza and Vito Piserchia. illumino: an XML document production system with a TEX core. TUGboat, 31(3):213–218, 2010. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb31-3/tb99centonza.pdf.

Hufflen:TB31-3-219

[3560] Jean-Michel Hufflen. Managing printed and online versions of large educational documents. TUGboat, 31(3):219–222, 2010. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb31-3/tb99hufflen.pdf.

Barnett:TB31-3-223

[3561] Michael Barnett. Aligning text in diagrams exported by Mathematica: a question about the PostScript infrastructure. TUGboat, 31(3):223–226,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-3/
tb99barnett.pdf.

Berry:TB31-3-227

[3562] Karl Berry. The treasure chest. TUGboat, 31(3):227–228, 2010. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-3/tb99chest.pdf.

Anonymous:TB31-3-229

[3563] Anonymous. ArsTEXnica: contents of issue 9 (October 2010). TUGboat, 31(3):

229, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-
3/tb99arstexnica.pdf.

Anonymous:TB31-3-230

[3564] Anonymous. MAPS: contents of issue 40 (2010). TUGboat, 31(3):230,

2010. ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-3/
tb99maps.pdf.

Anonymous:TB31-3-231

[3565] Anonymous. The PracTEX Journal: contents of issue 2010-1. TUGboat, 31(3): 231–232, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-3/tb99pracjourn.pdf.

Anonymous:TB31-3-233

[3566] Anonymous. Zpravodaj: contents of issues 20(1–2), 20(3) (2010). TUGboat, 31(3):233–235, 2010. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb31-3/tb99zprav.pdf.

Anonymous:TB31-3-236

[3567] Anonymous. Die TEXnische Komödie: contents of issue 2010/3. TUGboat, 31(3):236, 2010. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb31-3/tb99komo.pdf.

Anonymous:TB31-3-237

[3568] Anonymous. Calendar. TUGboat, 31(3):237, 2010. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb31-3/tb99calendar.pdf.

Anonymous:TB31-3-238

[3569] Anonymous. TEX consulting and production services. TUGboat, 31(3):238–

239, 2010. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB31-3-239

[3570] Anonymous. TUG institutional members. TUGboat, 31(3):239, 2010. ISSN 0896-3207.

Anonymous:TB31-3-240

[3571] Anonymous. TUG 2011 election. TUGboat, 31(3):240, 2010. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb31-3/tb99elec.pdf.

Anonymous:2010:CODc

[3572] Anonymous. Contents ordered by difficulty. TUGboat, 31(3):c3, 2010. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-
3/cover3.pdf.

Anonymous:2010:FCc

[3573] Anonymous. Front cover. TUGboat, 31(3):c1, 2010. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb31-3/cover1.pdf.

Anonymous:2010:IFCc

[3574] Anonymous. Inside front cover. TUGboat, 31(3):c2, 2010. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb31-3/
cover2.pdf.

Anonymous:2011:TPa

[3575] Anonymous. Title page. TUGboat, 32(1):1, 2011. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-1/titlepage.pdf.

Anonymous:2011:EI

[3576] Anonymous. Editorial information. TUGboat, 32(1):2, 2011. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100edit.pdf.

Berry:TB32-1-3

[3577] Karl Berry. From the President. TUGboat, 32(1):3, 2011. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb32-1/tb100pres.pdf.

Beeton:TB32-1-4

[3578] Barbara Beeton. Editorial comments. TUGboat, 32(1):4–5, 2011. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-1/tb100beet.pdf.

Damrau:TB32-1-6

[3579] Jackie Damrau. Mimi Burbank. TUGboat, 32(1):6, 2011. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb32-1/tb100damrau.pdf.

Thiele:TB32-1-7

[3580] Christina Thiele. Missing Mimi. TUGboat, 32(1):7–8, 2011. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb32-1/tb100thiele.pdf.

Hagen:TB32-1-9

[3581] Hans Hagen. 16 years of ConTEXt. TUGboat, 32(1):9–16, 2011. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-1/tb100hagen-
16.pdf.

Walden:TB32-1-17

[3582] David Walden and Karl Berry. TUGboat’s 100 issues — basic statistics and random gleanings. TUGboat, 32(1):17–22, 2011. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb32-1/tb100walden.pdf.

Berry:TB32-1-23

[3583] Karl Berry and David Walden. TUGboat online. TUGboat, 32(1):23–26,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100berry.pdf.

Veytsman:TB32-1-27

[3584] Boris Veytsman. TEX consulting for fun and profit. TUGboat, 32(1):27–29,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100veytsman.pdf.

Hefferon:TB32-1-30

[3585] Jim Hefferon. Which way to the forum? TUGboat, 32(1):30–31,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100heff.pdf.

Hwang:TB32-1-32

[3586] Andrew Hwang. L^{A}TEX at Distributed Proofreaders and the electronic preservation of mathematical literature at Project Gutenberg. TUGboat, 32(1):32–38,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100hwang.pdf.

Farar:TB32-1-39

[3587] Pavel Farář. Introducing the PT Sans and PT Serif typefaces. TUGboat, 32 (1):39–42, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-1/tb100farar.pdf.

Hagen:TB32-1-43

[3588] Hans Hagen. Handling math: a retrospective. TUGboat, 32(1):43–46,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100hagen-math.pdf.

West:TB32-1-47

[3589] Andrew West. The rules for long s. TUGboat, 32(1):47–55, 2011. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-1/tb100west.pdf.

Gregorio:TB32-1-56

[3590] Enrico Gregorio. Installing TEX Live 2010 on Ubuntu. TUGboat, 32(1):56–61,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100gregorio.pdf.

Hoekwater:TB32-1-62

[3591] Taco Hoekwater. tlcontrib.metatex.org: a complement to TEX Live. TUGboat, 32(1):62–67, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-1/tb100hoekwater.pdf.

Isambert:TB32-1-68

[3592] Paul Isambert. LuaTEX: What it takes to make a paragraph. TUGboat, 32 (1):68–76, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-1/tb100isambert.pdf.

Jackowski:TB32-1-77

[3593] Pawe� Jackowski. Luna — my side of the moon. TUGboat, 32(1):77–82,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100jackowski.pdf.

Mittelbach:TB32-1-83

[3594] Frank Mittelbach. Reflections on the history of the L^{A}TEX Project Public License (LPPL) — A software license for L^{A}TEX and more. TUGboat, 32(1):83–

94, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-
1/tb100mitt.pdf.

Wright:TB32-1-95

[3595] Joseph Wright. siunitx: a comprehensive (SI) units package. TUGboat, 32 (1):95–98, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-1/tb100wright-siunitx.pdf.

Wilson:TB32-1-99

[3596] Peter Wilson. Glisterings: Framing, new frames. TUGboat, 32(1):99–103,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100glister.pdf.

Merciadri:TB32-1-104

[3597] Luca Merciadri. Some misunderstood or unknown L^{A}TEX 2ε tricks III. TUGboat, 32(1):104–107, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-1/tb100merciadri.pdf.

LPT:TB32-1-108

[3598] L^{A}TEX Project Team. L^{A}TEX3 news, issue 5. TUGboat, 32(1):108,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100l3news.pdf.

Veytsman:TB32-1-109

[3599] Boris Veytsman. Book review: Typesetting tables with L^{A}TEX. TUGboat, 32(1):

109, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-
1/tb100reviews.pdf.

Berry:TB32-1-110

[3600] Karl Berry. The treasure chest. TUGboat, 32(1):110–112, 2011. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-1/tb100chest.
pdf.

Wright:TB32-1-113

[3601] Joseph Wright. ‘Magic’ comments in TEXworks0.4. TUGboat, 32(1):113,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100wright-texworks.pdf.

Anonymous:TB32-1-114

[3602] Anonymous. Eutypon: contents of issue 24–25 (October 2010). TUGboat, 32(1):114, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-1/tb100eutypon.pdf.

Anonymous:TB32-1-115

[3603] Anonymous. MAPS: contents of issue 41 (2010). TUGboat, 32(1):115,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100maps.pdf.

Anonymous:TB32-1-116

[3604] Anonymous. The PracTEX Journal: contents of issue 2010-2. TUGboat, 32(1):

116, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-
1/tb100pracjourn.pdf.

Anonymous:TB32-1-117

[3605] Anonymous. Die TEXnische Komödie: contents of issues 4/2010–1/2011. TUGboat, 32(1):117, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-1/tb100komo.pdf.

Anonymous:TB32-1-118

[3606] Anonymous. ArsTEXnica: contents of issue 10 (October 2010). TUGboat, 32(1):118, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-1/tb100arstexnica.pdf.

Anonymous:TB32-1-119

[3607] Anonymous. TEX consulting and production services. TUGboat, 32(1):119–

120, 2011. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Fine:TB32-1-120

[3608] Jonathan Fine. Is TEX obsolete? TUGboat, 32(1):120, 2011. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb32-1/tb100fine.pdf.

Anonymous:TB32-1-121

[3609] Anonymous. TUG institutional members. TUGboat, 32(1):121, 2011. ISSN

0896-3207. URL https://www.tug.org/instmem.html.

Walden:TB32-1-121

[3610] David Walden. TUG financial statements for 2010. TUGboat, 32(1):121–122,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100treas.pdf.

Hefferon:2011:TE

[3611] Jim Hefferon. TUG 2011 election. TUGboat, 32(1):123, 2011. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
tb100elec.pdf.

Anonymous:TB32-1-127

[3612] Anonymous. Calendar. TUGboat, 32(1):127, 2011. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb32-1/tb100calendar.pdf.

Anonymous:TB32-1-128

[3613] Anonymous. TUG 2011 announcement. TUGboat, 32(1):128, 2011. ISSN

0896-3207. URL https://tug.org/TUGboat/tb32-1/tb100tug11.
pdf.

Anonymous:2011:CODa

[3614] Anonymous. Contents ordered by difficulty. TUGboat, 32(1):c3, 2011. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-
1/cover3.pdf.

Anonymous:2011:FCa

[3615] Anonymous. Front cover. TUGboat, 32(1):c1, 2011. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-1/cover1.pdf.

Anonymous:2011:IFCa

[3616] Anonymous. Inside front cover. TUGboat, 32(1):c2, 2011. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-1/
cover2.pdf.

Anonymous:2011:TPb

[3617] Anonymous. Title page. TUGboat, 32(2):129, 2011. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-2/titlepage.pdf.

Beeton:TB32-2-131

[3618] Barbara Beeton. Editorial comments. TUGboat, 32(2):131, 2011. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-2/tb101beet.pdf.

Berry:TB32-2-131

[3619] Karl Berry. From the President. TUGboat, 32(2):131, 2011. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb32-2/tb101pres.pdf.

TCeditors:TB32-2-132

[3620] TEX Collection editors. TEX Collection 2011 DVD. TUGboat, 32(2):132,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
tb101texcoll.pdf.

Loffler:TB32-2-133

[3621] Stefan Löffler. TEXworks — As you like it. TUGboat, 32(2):133–135,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
tb101loeffler.pdf.

Hoekwater:TB32-2-136

[3622] Taco Hoekwater. MetaPost 1.750: Numerical engines. TUGboat, 32(2):136–

138, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-
2/tb101hoekwater.pdf.

Voss:TB32-2-139

[3623] Herbert Voß. Reading and executing source code. TUGboat, 32(2):139–144,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
tb101voss.pdf.

Pszczola:TB32-2-145

[3624] Krzysztof Pszczo�a. Teaching L^{A}TEX to the students of mathematics — the experience from The Jan Kochanowski University. TUGboat, 32(2):145,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
tb101pszczola.pdf.

Isambert:TB32-2-146

[3625] Paul Isambert. Drawing tables: Graphic fun with LuaTEX. TUGboat, 32(2): 146–151, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-2/tb101isambert.pdf.

Hagen:TB32-2-152

[3626] Hans Hagen. E-books: Old wine in new bottles. TUGboat, 32(2):152–157,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
tb101hagen.pdf.

Cheswick:TB32-2-158

[3627] William Cheswick. iTEX — Document formatting in an ereader world. TUGboat, 32(2):158–163, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-2/tb101cheswick.pdf.

Sharpe:TB32-2-164

[3628] Michael Sharpe. Math alphabets and the mathalfa package. TUGboat, 32(2): 164–168, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-2/tb101sharpe.pdf.

Vieth:TB32-2-169

[3629] Ulrik Vieth and Mojca Miklavec. Another incarnation of Lucida: Towards Lucida OpenType. TUGboat, 32(2):169–176, 2011. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb32-2/tb101vieth.pdf.

Scarso:TB32-2-177

[3630] Luigi Scarso. MFLua. TUGboat, 32(2):177–184, 2011. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb32-2/tb101scarso.pdf.

Grunewald:TB32-2-185

[3631] Michael Le Barbier Grünewald. Macro interfaces and the getoptk package. TUGboat, 32(2):185–192, 2011. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb32-2/tb101grunewald.pdf.

Parashchenko:TB32-2-193

[3632] Oleg Parashchenko. The cals package: Multipage tables with decorations. TUGboat, 32(2):193–201, 2011. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb32-2/tb101parashchenko.pdf.

Wilson:TB32-2-202

[3633] Peter Wilson. Glisterings: Ornaments. TUGboat, 32(2):202–205,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
tb101glister.pdf.

Merciadri:TB32-2-206

[3634] Luca Merciadri. Merciadri packages: An overview. TUGboat, 32(2):206–210,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
tb101merciadri.pdf.

Mahajan:TB32-2-211

[3635] Aditya Mahajan. ConTEXt basics for users: Paper setup. TUGboat, 32(2): 211–212, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-2/tb101mahajan.pdf.

Walden:TB32-2-213

[3636] David Walden. Experiences with notes, references, and bibliographies. TUGboat, 32(2):213–216, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-2/tb101walden.pdf.

Shaw:TB32-2-217

[3637] Paul Shaw. Sixty years of book design at St. Gallen, Switzerland. TUGboat, 32(2):217–224, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-2/tb101shaw.pdf.

Anonymous:TB32-2-224

[3638] Anonymous. TUG institutional members. TUGboat, 32(2):224, 2011. ISSN 0896-3207.

Adams:TB32-2-225

[3639] William Adams. Book review: A Specimen Portfolio of Wood Type in the Cary Collection. TUGboat, 32(2):225–228, 2011. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb32-2/tb101reviews-woodtype.pdf.

Veytsman:TB32-2-226

[3640] Boris Veytsman. Book review: The Art of the Book in the Twentieth Century. TUGboat, 32(2):226–229, 2011. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb32-2/tb101reviews-kelly.pdf.

Veytsman:TB32-2-228

[3641] Boris Veytsman. Book review: L^{A}TEX Beginner’s Guide. TUGboat, 32(2): 228–229, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-2/tb101reviews-kottwitz.pdf.

Walden:TB32-2-230

[3642] David Walden. An appreciation: The Art of Computer Programming, Volume 4A. TUGboat, 32(2):230–232, 2011. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb32-2/tb101reviews-knuth.pdf.

Berry:TB32-2-233

[3643] Karl Berry. The treasure chest. TUGboat, 32(2):233–234, 2011. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-2/tb101chest.
pdf.

Anonymous:TB32-2-235

[3644] Anonymous. Die TEXnische Komödie: contents of issues 4/2010–1/2011. TUGboat, 32(2):235, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-2/tb101komo.pdf.

Anonymous:TB32-2-236

[3645] Anonymous. Zpravodaj: contents of issue 20(4) (2010). TUGboat, 32(2):236,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
tb101zprav.pdf.

Anonymous:TB32-2-237

[3646] Anonymous. TEX consulting and production services. TUGboat, 32(2):237–

238, 2011. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Hefferon:TB32-2-238

[3647] Jim Hefferon. Status of the American core CTAN node. TUGboat, 32(2):238,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
tb101heff.pdf.

Anonymous:TB32-2-239

[3648] Anonymous. Calendar. TUGboat, 32(2):239, 2011. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb32-2/tb101calendar.pdf.

Anonymous:TB32-2-240

[3649] Anonymous. TUG 2011 announcement. TUGboat, 32(2):240, 2011. ISSN

0896-3207. URL https://tug.org/TUGboat/tb32-2/tb101tug11.
pdf.

Anonymous:2011:CODb

[3650] Anonymous. Contents ordered by difficulty. TUGboat, 32(2):c3, 2011. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-
2/cover3.pdf.

Anonymous:2011:FCb

[3651] Anonymous. Front cover. TUGboat, 32(2):c1, 2011. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-2/cover1.pdf.

Anonymous:2011:IFCb

[3652] Anonymous. Inside front cover. TUGboat, 32(2):c2, 2011. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-2/
cover2.pdf.

Anonymous:2011:TPc

[3653] Anonymous. Title page. TUGboat, 32(3):241, 2011. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-3/titlepage.pdf.

Anonymous:TB32-3-242

[3654] Anonymous. Conference program, delegates, and sponsors. TUGboat, 32(3): 242–244, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-3/tb102confinfo.pdf.

Beeton:TB32-3-245

[3655] Barbara Beeton. TUG 2011 in India. TUGboat, 32(3):245–247, 2011. ISSN

0896-3207. URL https://tug.org/TUGboat/tb32-3/tb102beet.
pdf.

Kottwitz:TB32-3-248

[3656] Stefan Kottwitz. TEX online communities — discussion and content. TUGboat, 32(3):248–250, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-3/tb102kottwitz.pdf.

Moudgalya:TB32-3-251

[3657] Kannan Moudgalya. L^{A}TEX training through spoken tutorials. TUGboat, 32(3): 251–256, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-3/tb102moudgalya.pdf.

Wetmore:TB32-3-257

[3658] Alan Wetmore. e-Readers and L^{A}TEX. TUGboat, 32(3):257–260, 2011. ISSN

0896-3207. URL https://tug.org/TUGboat/tb32-3/tb102wetmore.
pdf.

Veytsman:TB32-3-261

[3659] Boris Veytsman and Michael Ware. Ebooks and paper sizes: Output routines made easier. TUGboat, 32(3):261–265, 2011. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb32-3/tb102veytsman-ebooks.pdf.

Rishi:TB32-3-266

[3660] Rishi T. L^{A}TEX to ePub. TUGboat, 32(3):266–268, 2011. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb32-3/tb102rishi.pdf.

Joshi:TB32-3-269

[3661] Manjusha Joshi. A dream of computing and L^{A}TEXing together: a reality with SageTEX. TUGboat, 32(3):269–271, 2011. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb32-3/tb102joshi.pdf.

Kielhorn:TB32-3-272

[3662] Axel Kielhorn. Multi-target publishing. TUGboat, 32(3):272–277,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-3/
tb102kielhorn.pdf.

Venkatesan:TB32-3-279

[3663] S. K. Venkatesan. On the use of TEX as an authoring language for HTML5. TUGboat, 32(3):278–280, 2011. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb32-3/tb102venkatesan.pdf.

Sankar:TB32-3-281

[3664] S. Sankar, S. Mahalakshmi, and L. Ganesh. An XML model of CSS3 as an XL^{A}TEX-TEXML-HTML5 stylesheet language. TUGboat, 32(3):281–284,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-3/
tb102sankar.pdf.

Veytsman:TB32-3-285

[3665] Boris Veytsman and Leyla Akhmadeeva. Towards evidence-based typography: Literature review and experiment design. TUGboat, 32(3):285–288,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-3/
tb102veytsman-typo.pdf.

Hufflen:TB32-3-289

[3666] Jean-Michel Hufflen. A comparative study of methods for bibliographies. TUGboat, 32(3):289–301, 2011. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb32-3/tb102hufflen.pdf.

Housley:TB32-3-302

[3667] Brian Housley. The hletter class and style for producing flexible letters and page headings. TUGboat, 32(3):302–308, 2011. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb32-3/tb102housley.pdf.

Verna:TB32-3-309

[3668] Didier Verna. Towards L^{A}TEX coding standards. TUGboat, 32(3):309–328,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-3/
tb102verna.pdf.

Anonymous:TB32-3-329

[3669] Anonymous. TUG 2011 abstracts (Bazargan, Crossland, Radhakrishnan, Doumont, Mittelbach, Moore, Rishi, Skoupý, Sojka, Wujastyk). TUGboat, 32(3):329–330, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-3/tb102abstracts.pdf.

LPT:TB32-3-331

[3670] L^{A}TEX Project Team. L^{A}TEX news, issue 20. TUGboat, 32(3):331–332,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-3/
tb102ltnews.pdf.

Beitzel:TB32-3-333

[3671] Brian Beitzel. The meetingmins L^{A}TEX class: Hierarchically organized meeting agendas and minutes. TUGboat, 32(3):333–334, 2011. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb32-3/tb102beitzel.pdf.

Ruiz-Agundez:TB32-3-335

[3672] Igor Ruiz-Agundez. Collaborative L^{A}TEX writing with Google Docs. TUGboat, 32(3):335–338, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-3/tb102ruiz.pdf.

Wilson:TB32-3-339

[3673] Peter Wilson. Glisterings: Verbatim arguments; Cut off in its prime. TUGboat, 32(3):339–341, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-3/tb102glister.pdf.

Merciadri:TB32-3-342

[3674] Luca Merciadri. Some L^{A}TEX 2ε tricks and tips (IV). TUGboat, 32(3):342–344,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-3/
tb102merciadri.pdf.

Isambert:TB32-3-345

[3675] Paul Isambert. TEX as you like it: The interpreter package. TUGboat, 32(3): 345–348, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-3/tb102isambert.pdf.

Dziubinski:TB32-3-349

[3676] Wiktor Dziubiński, Marcin Woliński, and Grzegorz Murzynowski. PARCAT — Applying TEX in industry. TUGboat, 32(3):349–356, 2011. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb32-3/tb102murzynowski.
pdf.

Crossland:TB32-3-357

[3677] Dave Crossland. TUG Libre Font Fund, Google Web Fonts, and Kickstarter. TUGboat, 32(3):357, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-3/tb102fontfund.pdf.

Veytsman:TB32-3-358

[3678] Boris Veytsman. Book review: Bodoni, Manual of Typography — Manuale tipografico (1818). TUGboat, 32(3):358, 2011. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb32-3/tb102reviews-bodoni.pdf.

Veytsman:TB32-3-359

[3679] Boris Veytsman. Book review: L^{A}TEX and Friends. TUGboat, 32(3):359–360,

2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-3/
tb102reviews-ltxfriends.pdf.

Berry:TB32-3-361

[3680] Karl Berry. The treasure chest. TUGboat, 32(3):361, 2011. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb32-3/tb102chest.pdf.

Anonymous:TB32-3-362

[3681] Anonymous. ArsTEXnica: contents of issues 11–12 (2011). TUGboat, 32(3): 362–363, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb32-3/tb102arstexnica.pdf.

Anonymous:TB32-3-363

[3682] Anonymous. Die TEXnische Komödie: contents of issue 3/2011. TUGboat, 32(3):363–364, 2011. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb32-3/tb102komo.pdf.

Anonymous:TB32-3-364

[3683] Anonymous. The PracTEX Journal: contents of issue 2011-1. TUGboat, 32(3):

364, 2011. ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-
3/tb102pracjourn.pdf.

Anonymous:TB32-3-365

[3684] Anonymous. TEX consulting and production services. TUGboat, 32(3):365–

366, 2011. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB32-3-366

[3685] Anonymous. TUG institutional members. TUGboat, 32(3):366, 2011. ISSN 0896-3207.

Anonymous:TB32-3-367

[3686] Anonymous. Calendar. TUGboat, 32(3):367, 2011. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb32-3/tb102calendar.pdf.

Anonymous:TB32-3-368

[3687] Anonymous. TUG 2012 announcement. TUGboat, 32(3):368, 2011. ISSN

0896-3207. URL https://tug.org/TUGboat/tb32-3/tb102tug12.
pdf.

Anonymous:2011:CODc

[3688] Anonymous. Contents ordered by difficulty. TUGboat, 32(3):c3, 2011. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-
3/cover3.pdf.

Anonymous:2011:FCc

[3689] Anonymous. Front cover. TUGboat, 32(3):c1, 2011. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb32-3/cover1.pdf.

Anonymous:2011:IFCc

[3690] Anonymous. Inside front cover. TUGboat, 32(3):c2, 2011. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb32-3/
cover2.pdf.

Anonymous:TB33-1-2

[3691] Anonymous. TUG 2012 announcement. TUGboat, 33(1):2, 2012. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-1/tb103tug12.
pdf.

Peter:TB33-1-3

[3692] Steve Peter. Ab epistulis. TUGboat, 33(1):3, 2012. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb33-1/tb103pres.pdf.

Beeton:TB33-1-3

[3693] Barbara Beeton. Editorial comments. TUGboat, 33(1):3–4, 2012. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-1/tb103beet.pdf.

Beeton:TB33-1-5

[3694] Barbara Beeton. Hyphenation exception log. TUGboat, 33(1):5–6, 2012. ISSN

0896-3207. URL https://tug.org/TUGboat/tb33-1/tb103hyf.
pdf.

Armbruster:TB33-1-7

[3695] Bruce Armbruster and Jeannie Howard Siegman. In memoriam: Tony Siegman, 1931–2011. TUGboat, 33(1):7, 2012. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb33-1/tb103siegman.pdf.

Flynn:TB33-1-8

[3696] Peter Flynn. Typographers’ Inn. TUGboat, 33(1):8–10, 2012. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-1/tb103inn.pdf.

Berry:TB33-1-11

[3697] Karl Berry. Lucida OpenType fonts available from TUG. TUGboat, 33(1):11,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103lucida.pdf.

Hosny:TB33-1-12

[3698] Khaled Hosny. The Amiri typeface. TUGboat, 33(1):12, 2012. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-1/tb103hosny.
pdf.

Kime:TB33-1-13

[3699] Philip Kime. Biber — the next generation backend processor for BibL^{A}TEX. TUGboat, 33(1):13–15, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-1/tb103kime.pdf.

Beccari:TB33-1-16

[3700] Claudio Beccari. XEL^{A}TEX and the PDF archivable format. TUGboat, 33(1):16–

20, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-
1/tb103becc-pdfa.pdf.

Madsen:TB33-1-21

[3701] Lars Madsen. Avoid eqnarray! TUGboat, 33(1):21–25, 2012. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-1/tb103madsen.
pdf.

Beccari:TB33-1-26

[3702] Claudio Beccari. The unknown picture environment. TUGboat, 33(1):26–32,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103becc-picture.pdf.

Beitzel:TB33-1-33

[3703] Brian Beitzel. The apa6 L^{A}TEX class: Challenges encountered updating to new requirements. TUGboat, 33(1):33–38, 2012. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb33-1/tb103beitzel.pdf.

Wilson:TB33-1-39

[3704] Peter Wilson. Glisterings: Timelines; Parsing a filename. TUGboat, 33(1):39–

42, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-
1/tb103glister.pdf.

Merciadri:TB33-1-43

[3705] Luca Merciadri. Some L^{A}TEX 2ε tricks and tips (V). TUGboat, 33(1):43–45,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103merciadri.pdf.

LPT:TB33-1-46

[3706] L^{A}TEX Project Team. L^{A}TEX3 news, issues 6–7. TUGboat, 33(1):46–47,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103l3news.pdf.

Henderson:TB33-1-48

[3707] Troy Henderson. User-friendly web utilities for generating L^{A}TEX output and MetaPost graphics. TUGboat, 33(1):48–50, 2012. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb33-1/tb103henderson.pdf.

Wright:TB33-1-53

[3708] Joseph Wright. TEX on Windows: MiKTEX or TEX live? TUGboat, 33(1):53,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103wright.pdf.

Gundlach:TB33-1-54

[3709] Patrick Gundlach. Generating barcodes with LuaTEX. TUGboat, 33(1):54–58,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103gundlach.pdf.

Isambert:TB33-1-59

[3710] Paul Isambert. OpenType fonts in LuaTEX. TUGboat, 33(1):59–85,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103isambert.pdf.

Hagen:TB33-1-86

[3711] Hans Hagen. ConTEXt: Updating the code base. TUGboat, 33(1):86–97,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103hagen.pdf.

Jackowski:TB33-1-98

[3712] Bogus�aw Jackowski. Computing the area and winding number for a Bézier curve. TUGboat, 33(1):98–101, 2012. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb33-1/tb103jackowski.pdf.

Wolcott:TB33-1-102

[3713] Keith Wolcott. Three-dimensional graphics with PGF/TikZ. TUGboat, 33(1): 102–113, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb33-1/tb103wolcott.pdf.

Arora:TB33-1-114

[3714] Pavneet Arora. Book review: Trees, maps, and theorems. TUGboat, 33(1): 114–115, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb33-1/tb103reviews-doumont.pdf.

Veytsman:TB33-1-116

[3715] Boris Veytsman. Book review: Design for Hackers. TUGboat, 33(1):116–117,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103reviews-kadavy.pdf.

Walden:TB33-1-118

[3716] David Walden. Book review: Companion to the Papers of Donald Knuth. TUGboat, 33(1):118, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-1/tb103reviews-knuth9-walden.pdf.

Berry:TB33-1-119

[3717] Karl Berry. The treasure chest. TUGboat, 33(1):119, 2012. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb33-1/tb103chest.pdf.

Anonymous:TB33-1-121a

[3718] Anonymous. Eutypon: contents of issue 26–27 (October 2011). TUGboat, 33 (1):121, 2012. ISSN 0896-3207.

Anonymous:TB33-1-121b

[3719] Anonymous. Die TEXnische Komödie: contents of issues 4/2011–1/2012. TUGboat, 33(1):121–122, 2012. ISSN 0896-3207.

Anonymous:TB33-1-124

[3720] Anonymous. TEX consulting and production services. TUGboat, 33(1):124–

125, 2012. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB33-1-126

[3721] Anonymous. TUG institutional members. TUGboat, 33(1):126, 2012. ISSN

0896-3207. URL https://www.tug.org/instmem.html.

Berry:TB33-1-126

[3722] Karl Berry. TUG financial statements for 2011. TUGboat, 33(1):126–127,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-1/
tb103treas.pdf.

Anonymous:TB33-1-128

[3723] Anonymous. Calendar. TUGboat, 33(1):128, 2012. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb33-1/tb103calendar.pdf.

Anonymous:2012:TPa

[3724] Anonymous. Title page. TUGboat, 33(2):129, 2012. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-2/titlepage.pdf.

Anonymous:TB33-2-130

[3725] Anonymous. Conference program, delegates, and sponsors. TUGboat, 33(2): 130–131, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb33-2/tb104confinfo.pdf.

Latchman:TB33-2-132

[3726] David Latchman. TUG 2012: a first-time attendee. TUGboat, 33(2):132–137,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
tb104latchman-tug12.pdf.

Anonymous:TB33-2-138

[3727] Anonymous. Roundtable discussion: TEX consulting. TUGboat, 33(2):138–

145, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-
2/tb104panel.pdf.

Walden:TB33-2-146

[3728] David Walden. My Boston: Some printing and publishing history. TUGboat, 33(2):146–155, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-2/tb104walden.pdf.

Veytsman:TB33-2-156

[3729] Boris Veytsman and Leyla Akhmadeeva. Towards evidence-based typography: First results. TUGboat, 33(2):156–157, 2012. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb33-2/tb104veytsman-typo.pdf.

Garcia:TB33-2-158

[3730] Federico Garcia. TEX and music: an update on TEXmuse. TUGboat, 33(2): 158–164, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb33-2/tb104garcia.pdf.

LPT:TB33-2-165

[3731] L^{A}TEX Project Team. L^{A}TEX3 news, issue 8. TUGboat, 33(2):165–166,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
tb104l3news.pdf.

Latchman:TB33-2-167

[3732] David Latchman. Preparing your thesis in L^{A}TEX. TUGboat, 33(2):167–171,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
tb104latchman-thesis.pdf.

Flynn:TB33-2-172

[3733] Peter Flynn. A university thesis class: Automation and its pitfalls. TUGboat, 33(2):172–177, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-2/tb104flynn.pdf.

Childs:TB33-2-178

[3734] Bart Childs. L^{A}TEX source from word processors. TUGboat, 33(2):178–183,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
tb104childs.pdf.

Koch:TB33-2-184

[3735] Richard Koch. The MacTEX install package for OS X. TUGboat, 33(2):184–

191, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-
2/tb104koch.pdf.

Veytsman:TB33-2-192

[3736] Boris Veytsman. TEX and friends on a Pad. TUGboat, 33(2):192–195,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
tb104veytsman-pad.pdf.

Arora:TB33-2-196

[3737] Pavneet Arora. YAWN — a TEX-enabled workflow for project estimation. TUGboat, 33(2):196–198, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-2/tb104arora.pdf.

Verna:TB33-2-199

[3738] Didier Verna. Star TEX: The Next Generation. TUGboat, 33(2):199–208,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
tb104verna.pdf.

Neveln:TB33-2-209

[3739] Bob Neveln and Bob Alps. Adapting ProofCheck to the author’s needs. TUGboat, 33(2):209–212, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-2/tb104neveln.pdf.

Doob:TB33-2-213

[3740] Michael Doob and Jim Hefferon. Approaching Asymptote. TUGboat, 33(2): 213–218, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb33-2/tb104doob.pdf.

Hendrickson:TB33-2-219

[3741] Amy Hendrickson. The joy of \csname...\endcsname. TUGboat, 33(2): 219–224, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb33-2/tb104hendrickson.pdf.

Anonymous:TB33-2-225

[3742] Anonymous. TUG 2012 abstracts (Cheswick, Garcia, Henderson, Mansour, Mittelbach, Peter, Preining, Robertson, Thiele). TUGboat, 33(2):225–226,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
tb104abstracts.pdf.

Anonymous:2012:TKC

[3743] Anonymous. Die TeXnische Komödie: Contents of issues 2–3/2012. TUGboat, 33(2):227, 2012. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb33-2/tb104komo.pdf.

Anonymous:TB33-2-228

[3744] Anonymous. ArsTEXnica: contents of issue 13 (2012). TUGboat, 33(2):228,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
tb104arstexnica.pdf.

Anonymous:TB33-2-229

[3745] Anonymous. MAPS: contents of issue 42 (2011). TUGboat, 33(2):229,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
tb104maps.pdf.

Berry:TB33-2-230

[3746] Karl Berry. The treasure chest. TUGboat, 33(2):230–231, 2012. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-2/tb104chest.
pdf.

Veytsman:TB33-2-232

[3747] Boris Veytsman. Book review: About more alphabets: The types of Hermann Zapf. TUGboat, 33(2):232, 2012. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb33-2/tb104reviews-kelly-zapf.pdf.

Anonymous:TB33-2-233

[3748] Anonymous. TEX consulting and production services. TUGboat, 33(2):233–

234, 2012. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB33-2-235

[3749] Anonymous. TUG institutional members. TUGboat, 33(2):235, 2012. ISSN 0896-3207.

Anonymous:TB33-2-235-2

[3750] Anonymous. TEX Collection 2012. TUGboat, 33(2):235, 2012. ISSN 0896-3207.

Anonymous:TB33-2-236

[3751] Anonymous. Calendar. TUGboat, 33(2):236, 2012. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb33-2/tb104calendar.pdf.

Anonymous:2012:CODa

[3752] Anonymous. Contents ordered by difficulty. TUGboat, 33(2):c3, 2012. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-
2/cover3.pdf.

Anonymous:2012:FCa

[3753] Anonymous. Front cover. TUGboat, 33(2):c1, 2012. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-2/cover1.pdf.

Anonymous:2012:IFCa

[3754] Anonymous. Inside front cover. TUGboat, 33(2):c2, 2012. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-2/
cover2.pdf.

Anonymous:2012:TPb

[3755] Anonymous. Title page. TUGboat, 33(3):237, 2012. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-3/titlepage.pdf.

Anonymous:TB33-3-238

[3756] Anonymous. TUG institutional members. TUGboat, 33(3):238, 2012. ISSN

0896-3207. URL https://tug.org/TUGboat/tb33-3/tb105edit.
pdf.

Peter:TB33-3-239

[3757] Steve Peter. Ab epistulis. TUGboat, 33(3):239, 2012. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb33-3/tb105pres.pdf.

Beeton:TB33-3-240

[3758] Barbara Beeton. Editorial comments. TUGboat, 33(3):240–241, 2012. ISSN

0896-3207. URL https://tug.org/TUGboat/tb33-3/tb105beet.
pdf.

Kohm:TB33-3-241

[3759] Markus Kohm. KOMA-Script comes of age. TUGboat, 33(3):241–244,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-3/
tb105kohm-maturity.pdf.

Thiele:TB33-3-245

[3760] Christina Thiele. Almost 30 years of using TEX. TUGboat, 33(3):245–262,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-3/
tb105thiele.pdf.

Thurnherr:TB33-3-263

[3761] Thomas Thurnherr. Changing the font size in L^{A}TEX. TUGboat, 33(3):263–264,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-3/
tb105thurnherr.pdf.

Fuster:TB33-3-265

[3762] Robert Fuster. The calculator and calculus packages: Arithmetic and functional calculations inside L^{A}TEX. TUGboat, 33(3):265–271, 2012. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb33-3/tb105fuster.pdf.

Niederberger:TB33-3-272

[3763] Clemens Niederberger. The xtemplate package: An example. TUGboat, 33(3): 272–275, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb33-3/tb105niederberger.pdf.

Pupalaikis:TB33-3-276

[3764] Peter Pupalaikis. A patent application design flow in L^{A}TEX and LYX. TUGboat, 33(3):276–281, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-3/tb105pupalaikis.pdf.

Kohm:TB33-3-282

[3765] Markus Kohm. Page style tricks with scrpage2. TUGboat, 33(3):282–284,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-3/
tb105kohm-scrpage2.pdf.

Voipio:TB33-3-285

[3766] Mari Voipio. CrafTEX: Applying TEX, MetaPost, and friends in crafts. TUGboat, 33(3):285–288, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-3/tb105voipio.pdf.

Delprat:TB33-3-289

[3767] Bruno Delprat and Stepan Orevkov. MayaPS: Maya hieroglyphics with (La)TEX. TUGboat, 33(3):289–294, 2012. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb33-3/tb105delprat.pdf.

Mansour:TB33-3-295

[3768] Sherif Mansour and Hossam Fahmy. Experiences with Arabic font development. TUGboat, 33(3):295–298, 2012. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb33-3/tb105mansour.pdf.

Veytsman:TB33-3-299

[3769] Boris Veytsman. The fonts we choose. TUGboat, 33(3):299–300,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-3/
tb105veytsman-fonts.pdf.

Voss:TB33-3-301

[3770] Herbert Voß. Using TEX GyrePagella OpenType Math. TUGboat, 33(3):301,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-3/
tb105voss.pdf.

Vieth:TB33-3-302

[3771] Ulrik Vieth. OpenType math font development: Progress and challenges. TUGboat, 33(3):302–308, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-3/tb105vieth.pdf.

Romer:TB33-3-309

[3772] Linus Romer. From drawn to filled paths. TUGboat, 33(3):309–312,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-3/
tb105romer.pdf.

Wilson:TB33-3-313

[3773] Peter Wilson. Glisterings: Cutout windows. TUGboat, 33(3):313–315,

2012. ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-3/
tb105glister.pdf.

Veytsman:TB33-3-316

[3774] Boris Veytsman. Stubborn leaders and juggling boxes: a slightly unusual table of contents. TUGboat, 33(3):316–318, 2012. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb33-3/tb105veytsman-leaders.pdf.

Berry:TB33-3-319

[3775] Karl Berry. The treasure chest. TUGboat, 33(3):319–320, 2012. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-3/tb105chest.
pdf.

Veytsman:TB33-3-320

[3776] Boris Veytsman. Book review: PSTricks: Graphics and PostScript for TEX and L^{A}TEX. TUGboat, 33(3):320–321, 2012. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb33-3/tb105reviews-voss.pdf.

Walden:TB33-3-322

[3777] David Walden. Book review: Just My Type: A book about fonts. TUGboat, 33(3):322–323, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-3/tb105reviews-garfield.pdf.

Anonymous:TB33-3-324

[3778] Anonymous. Les Cahiers GUTenberg: Contents of issue 56 (2011) (simultaneously published as MAPS 43 and Zpravodaj 2–4/2011). TUGboat, 33(3): 324–325, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb33-3/tb105cahi.pdf.

Anonymous:TB33-3-325

[3779] Anonymous. ArsTEXnica: contents of issue 14 (October 2012). TUGboat, 33(3):325–326, 2012. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb33-3/tb105arstexnica.pdf.

Anonymous:TB33-3-326

[3780] Anonymous. The PracTEX Journal: contents of issue 2012-1. TUGboat, 33(3): 326–327, 2012. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb33-3/tb105pracjourn.pdf.

Anonymous:TB33-3-328

[3781] Anonymous. TUG 2013 announcement. TUGboat, 33(3):328, 2012. ISSN

0896-3207. URL https://tug.org/TUGboat/tb33-3/tb105tug13.
pdf.

Knuth:TB33-3-328

[3782] Don Knuth. PUB and pre-TEX history. TUGboat, 33(3):328, 2012. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-3/tb105knut.pdf.

Anonymous:TB33-3-329

Anonymous:TB33-3-330

[3784] Anonymous. TEX consulting and production services. TUGboat, 33(3):330–

331, 2012. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB33-3-332

[3785] Anonymous. TUG 2013 election. TUGboat, 33(3):332, 2012. ISSN 0896-3207.

URL https://tug.org/TUGboat/election/2012/.

Anonymous:2012:CODb

[3786] Anonymous. Contents ordered by difficulty. TUGboat, 33(3):c3, 2012. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-
3/cover3.pdf.

Anonymous:2012:FCb

[3787] Anonymous. Front cover. TUGboat, 33(3):c1, 2012. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb33-3/cover1.pdf.

Anonymous:2012:IFCb

[3788] Anonymous. Inside front cover. TUGboat, 33(3):c2, 2012. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb33-3/
cover2.pdf.

Anonymous:2013:TPa

[3789] Anonymous. Title page. TUGboat, 34(1):1, 2013. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-1/titlepage.pdf.

Anonymous:TB34-1-2

[3790] Anonymous. TUGboat editorial information. TUGboat, 34(1):2, 2013. ISSN 0896-3207.

Anonymous:TB34-1-2-2

[3791] Anonymous. TUG institutional members. TUGboat, 34(1):2, 2013. ISSN 0896-3207.

Peter:TB34-1-3

[3792] Steve Peter. Ab epistulis. TUGboat, 34(1):3, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-1/tb106pres.pdf.

Beeton:TB34-1-4

[3793] Barbara Beeton. Editorial comments. TUGboat, 34(1):4–5, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-1/tb106beet.pdf.

Neugebauer:TB34-1-6

[3794] Gerd Neugebauer. CTAN: Relaunch of the Web portal. TUGboat, 34(1):6,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-1/
tb106neugebauer.pdf.

Tennent:TB34-1-10

[3795] Bob Tennent. Fonts! Fonts! Fonts! TUGboat, 34(1):10–13, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-1/tb106tennent.
pdf.

Flynn:TB34-1-14

[3796] Peter Flynn. Typographers’ Inn. TUGboat, 34(1):14–16, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-1/tb106inn.pdf.

Voipio:TB34-1-17

[3797] Mari Voipio. Entry-level MetaPost: On the grid. TUGboat, 34(1):17–20,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-1/
tb106voipio-grid.pdf.

Voipio:TB34-1-21

[3798] Mari Voipio. Recreating historical patterns with MetaPost. TUGboat, 34 (1):21–25, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb34-1/tb106voipio-patterns.pdf.

Fuster:TB34-1-26

[3799] Robert Fuster. The xpicture package. TUGboat, 34(1):26–33, 2013. ISSN

0896-3207. URL https://tug.org/TUGboat/tb34-1/tb106fuster.
pdf.

Thurnherr:TB34-1-34

[3800] Thomas Thurnherr. Side-by-side figures in L^{A}TEX. TUGboat, 34(1):34–36,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-1/
tb106thurnherr.pdf.

Wilson:TB34-1-37

[3801] Peter Wilson. Glisterings: Repetition; Verbatims; Small pages; Prefixing section heads. TUGboat, 34(1):37–39, 2013. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb34-1/tb106glister.pdf.

Messineo:TB34-1-40

[3802] Grazia Messineo and Salvatore Vassallo. The esami package for examinations. TUGboat, 34(1):40–46, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-1/tb106messineo.pdf.

Mittelbach:TB34-1-47

[3803] Frank Mittelbach. E-TEX: Guidelines for future TEX extensions — revisited. TUGboat, 34(1):47–63, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-1/tb106mitt.pdf.

Scarso:TB34-1-64

[3804] Luigi Scarso. LuaJITTEX. TUGboat, 34(1):64–71, 2013. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb34-1/tb106scarso.pdf.

Hagen:TB34-1-72

[3805] Hans Hagen. ConTEXt: Just-in-time LuaTEX. TUGboat, 34(1):72–78,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-1/
tb106hagen.pdf.

Mahajan:TB34-1-79

[3806] Aditya Mahajan. ConTEXt basics for users: Images. TUGboat, 34(1):79–82,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-1/
tb106mahajan.pdf.

Olsak:TB34-1-83

[3807] Petr Olšák. New CSplain of 2012. TUGboat, 34(1):83–87, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-1/tb106olsak-
csplain.pdf.

Olsak:TB34-1-88

[3808] Petr Olšák. OPmac: Macros for plain TEX. TUGboat, 34(1):88–95,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-1/
tb106olsak-opmac.pdf.

Berry:TB34-1-96

[3809] Karl Berry. The treasure chest. TUGboat, 34(1):96–97, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-1/tb106chest.
pdf.

Berry:TB34-1-97

[3810] Karl Berry. Production notes. TUGboat, 34(1):97, 2013. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb34-1/tb106prod.pdf.

Veytsman:TB34-1-98

[3811] Boris Veytsman. Book review: The Computer Science of TEX and L^{A}TEX. TUGboat, 34(1):98, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-1/tb106reviews-eijkhout.pdf.

Anonymous:TB34-1-99

[3812] Anonymous. Die TEXnische Komödie: contents of issues 4/2012–1/2013. TUGboat, 34(1):99, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-1/tb106komo.pdf.

Anonymous:TB34-1-100

[3813] Anonymous. Eutypon: contents of issue 28–29 (October 2012). TUGboat, 34(1):100, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb34-1/tb106eutypon.pdf.

Anonymous:TB34-1-101

[3814] Anonymous. Calendar. TUGboat, 34(1):101, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-1/tb106calendar.pdf.

Anonymous:TB34-1-102

[3815] Anonymous. TUG 2013 announcement. TUGboat, 34(1):102, 2013. ISSN

0896-3207. URL https://tug.org/TUGboat/tb34-1/tb106tug13.
pdf.

Anonymous:TB34-1-103

[3816] Anonymous. TEX consulting and production services. TUGboat, 34(1):103–

104, 2013. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB34-1-105

[3817] Anonymous. TUG membership form. TUGboat, 34(1):105, 2013. ISSN 0896-

3207. URL https://tug.org/join.html.

Berry:TB34-1-106

[3818] Karl Berry. TUG financial statements for 2012. TUGboat, 34(1):106,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-1/
tb106treas.pdf.

Anonymous:TB34-1-107

[3819] Anonymous. TUG 2013 election. TUGboat, 34(1):107, 2013. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb34-1/tb106elec.pdf.

Quinn:TB34-1-108

[3820] Daniel Quinn. Colophon. TUGboat, 34(1):108, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-1/tb106quinn.pdf.

Anonymous:2013:CODa

[3821] Anonymous. Contents ordered by difficulty. TUGboat, 34(1):c3, 2013. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-
1/cover3.pdf.

Anonymous:2013:FCa

[3822] Anonymous. Front cover. TUGboat, 34(1):c1, 2013. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-1/cover1.pdf.

Anonymous:2013:IFCa

[3823] Anonymous. Inside front cover. TUGboat, 34(1):c2, 2013. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-1/
cover2.pdf.

Anonymous:2013:TPb

[3824] Anonymous. Title page. TUGboat, 34(2):109, 2013. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-2/titlepage.pdf.

Anonymous:TB34-2-110

[3825] Anonymous. TUGboat editorial information. TUGboat, 34(2):110, 2013. ISSN 0896-3207.

Anonymous:TB34-2-110-2

[3826] Anonymous. TUG institutional members. TUGboat, 34(2):110, 2013. ISSN 0896-3207.

Peter:TB34-2-111

[3827] Steve Peter. Ab epistulis. TUGboat, 34(2):111, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-2/tb107pres.pdf.

Beeton:TB34-2-111

[3828] Barbara Beeton. Editorial comments. TUGboat, 34(2):111–112, 2013. ISSN

0896-3207. URL https://tug.org/TUGboat/tb34-2/tb107beet.
pdf.

Henderson:TB34-2-112

[3829] Doug Henderson. In memoriam: Barry Smith (1953–2012). TUGboat, 34(2):

112, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-
2/tb107smith.pdf.

Beeton:TB34-2-113

[3830] Barbara Beeton. Hyphenation exception log. TUGboat, 34(2):113–114,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107hyf.pdf.

Wilcox:TB34-2-114

[3831] Adeline Wilcox. Running TEX under Windows PowerShell. TUGboat, 34(2):

114, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-
2/tb107wilcox.pdf.

Hagen:TB34-2-115

[3832] Hans Hagen. Does TEX have a future? TUGboat, 34(2):115–119,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107hagen.pdf.

TCeditors:TB34-2-120

[3833] TEX Collection editors. TEX Collection 2013 DVD. TUGboat, 34(2):120,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107texcoll.pdf.

Hosny:TB34-2-121

[3834] Khaled Hosny. What is new in XETEX 0.9999? TUGboat, 34(2):121–122,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107hosny.pdf.

Hoekwater:TB34-2-123

[3835] Taco Hoekwater. MetaPost: PNG output. TUGboat, 34(2):123, 2013. ISSN

0896-3207. URL https://tug.org/TUGboat/tb34-2/tb107hoekwater.
pdf.

Hunniger:TB34-2-124

[3836] Dirk Hünniger. Converting Wikipedia articles to L^{A}TEX. TUGboat, 34(2):124,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107huenniger.pdf.

Sharpe:TB34-2-125

[3837] Michael Sharpe. A survey of text font families. TUGboat, 34(2):125–131,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107sharpe.pdf.

Wilson:TB34-2-132

[3838] Peter Wilson. Glisterings: a font of fleurons; Fonts, GNU/Linux, and XETEX; Mixing traditional and system fonts. TUGboat, 34(2):132–135,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107glister.pdf.

Wang:TB34-2-136

[3839] Yue Wang. Interview with Charles Bigelow. TUGboat, 34(2):136–167,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107bigelow-wang.pdf.

Bigelow:TB34-2-168

[3840] Charles Bigelow. Oh, oh, zero! TUGboat, 34(2):168–181, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-2/tb107bigelow-
zero.pdf.

Berry:TB34-2-181

[3841] Karl Berry. Production notes. TUGboat, 34(2):181, 2013. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb34-2/tb107prod.pdf.

Berry:TB34-2-182

[3842] Karl Berry. The treasure chest. TUGboat, 34(2):182–183, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-2/tb107chest.
pdf.

Lupkowski:TB34-2-184

[3843] Pawe� �upkowski and Mariusz Urbański. Preparing for scientific conferences with L^{A}TEX: a short practical how-to. TUGboat, 34(2):184–189,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107lupkowski.pdf.

Esser:TB34-2-190

[3844] Hans-Georg Eßer. LiPPGen: a presentation generator for literate-programming-based teaching. TUGboat, 34(2):190–195, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-2/tb107esser.pdf.

Voipio:TB34-2-196

[3845] Mari Voipio. Entry-level MetaPost 2: Move it! TUGboat, 34(2):196–199,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107voipio.pdf.

Dugge:TB34-2-200

[3846] Juernjakob Dugge. Creating Tufte-style bar charts and scatterplots using PGFPlots. TUGboat, 34(2):200–204, 2013. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb34-2/tb107dugge.pdf.

Jackowski:TB34-2-205

[3847] Bogus�aw Jackowski. Typographers, programmers and mathematicians, or the case of an æsthetically pleasing interpolation. TUGboat, 34(2):205–222,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107jackowski.pdf.

Donnelly:TB34-2-223

[3848] Kevin Donnelly. Representing linguistic pitch in (XE)L^{A}TEX. TUGboat, 34(2): 223–227, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb34-2/tb107donnelly.pdf.

Veytsman:TB34-2-228

[3849] Boris Veytsman. Book review: Learning L^{A}TEX. TUGboat, 34(2):228,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107reviews-learnltx.pdf.

Veytsman:TB34-2-229

[3850] Boris Veytsman. Book review: Zapf Exhibition: The Calligraphy of Hermann & Gudrun Zapf. TUGboat, 34(2):229, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-2/tb107reviews-
zapfexhibition.pdf.

Adams:TB34-2-230

[3851] Will Adams. Book review: What Our Lettering Needs. TUGboat, 34(2):230,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107reviews-zapfhallmark.pdf.

Veytsman:TB34-2-231

[3852] Boris Veytsman. Book review: L^{A}TEX Quick Reference. TUGboat, 34(2):231,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
tb107reviews-vossltxref.pdf.

Veytsman:TB34-2-232

[3853] Boris Veytsman. Book review: Presentations with L^{A}TEX. TUGboat, 34(2): 232–233, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb34-2/tb107reviews-vosspres.pdf.

Anonymous:TB34-2-233

[3854] Anonymous. TEX consulting and production services. TUGboat, 34(2):233–

234, 2013. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB34-2-235

[3855] Anonymous. Die TEXnische Komödie: contents of issue 2/2013. TUGboat, 34(2):235, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb34-2/tb107arstexnica.pdf.

Anonymous:TB34-2-235-2

[3856] Anonymous. ArsTEXnica: contents of issue 15 (April 2013). TUGboat, 34(2):

235, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-
2/tb107arstexnica.pdf.

Anonymous:TB34-2-236

[3857] Anonymous. GUST: EuroBachoTEX2013 proceedings. TUGboat, 34(2):236–

239, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-
2/tb107gust.pdf.

Anonymous:TB34-2-240

[3858] Anonymous. TUG 2013 election. TUGboat, 34(2):240–242, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-2/tb107elec.pdf.

Anonymous:TB34-2-243

[3859] Anonymous. Calendar. TUGboat, 34(2):243, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-2/tb107calendar.pdf.

Anonymous:TB34-2-244

[3860] Anonymous. TUG 2013 announcement. TUGboat, 34(2):244, 2013. ISSN

0896-3207. URL https://tug.org/TUGboat/tb34-2/tb107tug13.
pdf.

Anonymous:2013:CODb

[3861] Anonymous. Contents ordered by difficulty. TUGboat, 34(2):c3, 2013. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-
2/cover3.pdf.

Anonymous:2013:FCb

[3862] Anonymous. Front cover. TUGboat, 34(2):c1, 2013. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-2/cover1.pdf.

Anonymous:2013:IFCb

[3863] Anonymous. Inside front cover. TUGboat, 34(2):c2, 2013. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-2/
cover2.pdf.

Anonymous:2013:TPc

[3864] Anonymous. Title page. TUGboat, 34(3):245, 2013. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-3/titlepage.pdf.

Anonymous:TB34-3-246

[3865] Anonymous. Conference sponsors, participants, and program. TUGboat, 34(3):246–249, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-3/tb108confinfo.pdf.

Anonymous:2013:TCP

[3866] Anonymous. TUG 2013 conference program. TUGboat, 34(3):250–251, 2013. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb34-3/tb108program.pdf.

Preining:TB34-3-252

[3867] Norbert Preining. TUG 2013 in Tokyo. TUGboat, 34(3):252–258,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-3/
tb108preining-tug13.pdf.

Verna:TB34-3-259

[3868] Didier Verna. The incredible tale of the author who didn’t want to do the publisher’s job, . . .. TUGboat, 34(3):259–262, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-3/tb108verna.pdf.

Lewis:TB34-3-263

[3869] Jason Lewis. How to make a product catalogue that doesn’t look like a dissertation. TUGboat, 34(3):263–267, 2013. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb34-3/tb108lewis.pdf.

Ma:TB34-3-268

[3870] Clerk Ma and Jie Su. Project Fandol: GPL fonts for Chinese typesetting. TUGboat, 34(3):268, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-3/tb108ma-fandol.pdf.

Skala:TB34-3-269

[3871] Matthew Skala. Tsukurimashou: a Japanese-language font meta-family. TUGboat, 34(3):269–278, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-3/tb108skala.pdf.

Ma:TB34-3-279

[3872] Clerk Ma. Braille fonts in Project Fandol. TUGboat, 34(3):279–280,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-3/
tb108ma-braille.pdf.

Nakano:TB34-3-281

[3873] Ken Nakano and Hajime Kobayashi. Case study: Typesetting old documents of Japan. TUGboat, 34(3):281–284, 2013. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb34-3/tb108nakano.pdf.

Tanaka:TB34-3-285

[3874] Takuji Tanaka. upTEX — Unicode version of pTEX with CJK extensions. TUGboat, 34(3):285–288, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-3/tb108tanaka.pdf.

Plaice:TB34-3-289

[3875] John Plaice. Typesetting and layout in multiple directions — proposed solution. TUGboat, 34(3):289–292, 2013. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb34-3/tb108plaice.pdf.

Preining:TB34-3-293

[3876] Norbert Preining. TEX Live Manager’s rare gems: User mode and multiple repository support. TUGboat, 34(3):293–296, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-3/tb108preining-tlmgr.
pdf.

Preining:TB34-3-297

[3877] Norbert Preining. Redistributing TEX and friends. TUGboat, 34(3):297–301,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-3/
tb108preining-distro.pdf.

Mertz:TB34-3-302

[3878] Andrew Mertz and William Slough. A gentle introduction to PythonTEX. TUGboat, 34(3):302–312, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-3/tb108mertz.pdf.

Wang:TB34-3-313

[3879] Lu Wang and Wanmin Liu. Online publishing via pdf2htmlEX. TUGboat, 34(3):313–324, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-3/tb108wang.pdf.

Fujita:TB34-3-325

[3880] Shinsaku Fujita. The XΥMTEX system for publishing interdisciplinary chemistry/mathematics books. TUGboat, 34(3):325–328, 2013. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb34-3/tb108fujita.pdf.

Arora:TB34-3-329

[3881] Pavneet Arora. TANSU — a workflow for cabinet layout. TUGboat, 34(3): 329–331, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb34-3/tb108arora.pdf.

Hagen:TB34-3-332

[3882] Nathan Hagen. Bibulous — a drop-in BibTEX replacement based on style templates. TUGboat, 34(3):332–339, 2013. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb34-3/tb108hagen.pdf.

Cohen:TB34-3-340

[3883] Michael Cohen, Yannis Haralambous, and Boris Veytsman. The multibibliography package. TUGboat, 34(3):340–343, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-3/tb108cohen.pdf.

Hankus:TB34-3-344

[3884] Aleksandra Hankus and Zofia Walczak. L^{A}TEX and graphics: Basics and packages. TUGboat, 34(3):344–348, 2013. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb34-3/tb108hankus.pdf.

Veytsman:TB34-3-349

[3885] Boris Veytsman and Leyla Akhmadeeva. Plots in L^{A}TEX: Gnuplot, Octave, make. TUGboat, 34(3):349–356, 2013. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb34-3/tb108veytsman.pdf.

Voipio:TB34-3-357

[3886] Mari Voipio. Entry-level MetaPost 3: Color. TUGboat, 34(3):357–359,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-3/
tb108voipio.pdf.

Anonymous:TB34-3-360

[3887] Anonymous. TUG 2013 abstracts (Cho, Hagen, Hakuta, Maeda & Kaneko, Minoda, Mittelbach, Moore, Shikano, Takata, Terada, Verna, Wetmore, Yabe). TUGboat, 34(3):360–362, 2013. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb34-3/tb108abstracts.pdf.

Anonymous:2013:CGP

[3888] Anonymous. ConTeXt Group: Proceedings, 6th meeting (2012). TUGboat, 34 (3):363, 2013. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-3/tb108ctxg.pdf.

Anonymous:TB34-3-364

[3889] Anonymous. MAPS: contents of issue 44 (2013). TUGboat, 34(3):364–365,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-3/
tb108maps.pdf.

Anonymous:TB34-3-365

[3890] Anonymous. Die TEXnische Komödie: contents of issues 3–4/2013. TUGboat, 34(3):365, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb34-3/tb108komo.pdf.

Berry:TB34-3-366

[3891] Karl Berry. The treasure chest. TUGboat, 34(3):366–367, 2013. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-3/tb108chest.
pdf.

Bitouze:TB34-3-367

[3892] Denis Bitouzé. In memoriam: Jean-Pierre Drucbert (1947–2009). TUGboat, 34(3):367, 2013. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb34-3/tb108drucbert.pdf.

Walden:TB34-3-368

[3893] Dave Walden. Book review: Essential Knuth. TUGboat, 34(3):368–369,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-3/
tb108reviews-knuth.pdf.

Ma:TB34-3-369

[3894] Clerk Ma. Book review: Introduction to L^{A}TEX. TUGboat, 34(3):369,

2013. ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-3/
tb108reviews-liuintro.pdf.

Anonymous:TB34-3-370

[3895] Anonymous. TUG institutional members. TUGboat, 34(3):370, 2013. ISSN

0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB34-3-370-2

[3896] Anonymous. TEX consulting and production services. TUGboat, 34(3):370–

371, 2013. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB34-3-372

[3897] Anonymous. Calendar. TUGboat, 34(3):372, 2013. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb34-3/tb108calendar.pdf.

Anonymous:2013:CODc

[3898] Anonymous. Contents ordered by difficulty. TUGboat, 34(3):c3, 2013. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-
3/cover3.pdf.

Anonymous:2013:FCc

[3899] Anonymous. Front cover. TUGboat, 34(3):c1, 2013. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb34-3/cover1.pdf.

Anonymous:2013:IFCc

[3900] Anonymous. Inside front cover. TUGboat, 34(3):c2, 2013. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb34-3/
cover2.pdf.

Anonymous:2014:TPa

[3901] Anonymous. Title page. TUGboat, 35(1):1, 2014. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-1/titlepage.pdf.

Anonymous:TB35-1-2

[3902] Anonymous. TUGboat editorial information. TUGboat, 35(1):2, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-1/tb109edit.
pdf.

Peter:TB35-1-2

[3903] Steve Peter. Ab epistulis. TUGboat, 35(1):2, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-1/tb109pres.pdf.

Beeton:TB35-1-3

[3904] Barbara Beeton. Editorial comments. TUGboat, 35(1):3–4, 2014. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-1/tb109beet.pdf.

Knuth:TB35-1-5

[3905] Donald Knuth. The TEX tuneup of 2014. TUGboat, 35(1):5–8, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-1/tb109knut.
pdf.

Hunt:TB35-1-9

[3906] James Hunt. Making lists: a journey into unknown grammar. TUGboat, 35 (1):9–15, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-1/tb109hunt.pdf.

Gaspar:TB35-1-16

[3907] Jaime Gaspar. Does not suffice to run latex a finite number of times to get cross-references right. TUGboat, 35(1):16, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-1/tb109gaspar.pdf.

Hufflen:TB35-1-16

[3908] Jean-Michel Hufflen. In memoriam Jean-Pierre Drucbert. TUGboat, 35(1):16,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109drucbert.pdf.

Romer:TB35-1-17

[3909] Linus Romer. Fetamont: an extended logo typeface. TUGboat, 35(1):17–21,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109romer.pdf.

LPT:TB35-1-22

[3910] L^{A}TEX Project Team. L^{A}TEX3 news, issue 9. TUGboat, 35(1):22–26,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109l3news.pdf.

Thurnherr:TB35-1-27

[3911] Thomas Thurnherr. Introduction to presentations with beamer. TUGboat, 35 (1):27–30, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-1/tb109thurnherr.pdf.

Wright:TB35-1-31

[3912] Joseph Wright. The beamer class: Controlling overlays. TUGboat, 35(1):31–

33, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-
1/tb109wright.pdf.

Pagnossin:TB35-1-34

[3913] Ivan Pagnossin. Boxes and more boxes. TUGboat, 35(1):34–35,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109pagnossin.pdf.

Wilson:TB35-1-36

[3914] Peter Wilson. Glisterings: Glyphs, long labels. TUGboat, 35(1):36–38,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109glister.pdf.

Helvensteijn:TB35-1-39

[3915] Michiel Helvensteijn. The pkgloader and lt3graph packages: Toward simple and powerful package management for L^{A}TEX. TUGboat, 35(1):39–43,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109helvensteijn.pdf.

Dominici:TB35-1-44

[3916] Massimiliano Dominici. An overview of Pandoc. TUGboat, 35(1):44–50,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109dominici.pdf.

Montijano:TB35-1-51

[3917] Juan Montijano, Mario Pérez, Luis Rández, and Juan Luis Varona. Numerical methods with LuaL^{A}TEX. TUGboat, 35(1):51–56, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-1/tb109montijano.pdf.

Hoekwater:TB35-1-57

[3918] Taco Hoekwater. Parsing PDF content streams with LuaTEX. TUGboat, 35 (1):57–60, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-1/tb109hoekwater.pdf.

Bossard:TB35-1-61

[3919] Antoine Bossard and Takeyuki Nagao. ModernDvi: a high quality rendering and modern DVI viewer. TUGboat, 35(1):61–68, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-1/tb109bossard.pdf.

Hagen:TB35-1-69

[3920] Hans Hagen. Selection in PDF viewers and a LuaTEX bug. TUGboat, 35(1):69–

70, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-
1/tb109hagen-select.pdf.

Shibakov:TB35-1-71

[3921] Alexander Shibakov. Parsers in TEX and using CWEB for general prettyprinting. TUGboat, 35(1):71–78, 2014. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb35-1/tb109shibakov.pdf.

Voipio:TB35-1-79

[3922] Mari Voipio. Entry-level MetaPost 4: Artful lines. TUGboat, 35(1):79–81,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109voipio.pdf.

Hagen:TB35-1-82

[3923] Hans Hagen. drawdot in MetaPost: a bug, a fix. TUGboat, 35(1):82,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109hagen-drawdot.pdf.

Schlupkothen:TB35-1-83

[3924] Frederik Schlupkothen. HTML to L^{A}TEX transformation. TUGboat, 35(1):83–

90, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-
1/tb109schlupkothen.pdf.

Suarez-Burgoa:TB35-1-91

[3925] Ludger Suarez-Burgoa. Scientific documents written by novice researchers: a personal experience in Latin America. TUGboat, 35(1):91–98, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-1/tb109suarez.
pdf.

Walden:TB35-1-99

[3926] David Walden. Macro memories, 1964–2013. TUGboat, 35(1):99–110,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109walden.pdf.

Anonymous:TB35-1-111

[3927] Anonymous. Eutypon: contents of issue 30–31 (October 2013). TUGboat, 35(1):111, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-1/tb109eutypon.pdf.

Anonymous:TB35-1-111-2

[3928] Anonymous. Die TEXnische Komödie: contents of issue 1/2014. TUGboat, 35(1):111, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-1/tb109komo.pdf.

Berry:TB35-1-112

[3929] Karl Berry. The treasure chest. TUGboat, 35(1):112, 2014. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb35-1/tb109chest.pdf.

Veytsman:TB35-1-113

[3930] Boris Veytsman. Book reviews: L^{A}TEX for Complete Novices and Using L^{A}TEX to Write a PhD Thesis, by Nicola Talbot. TUGboat, 35(1):113–114,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109reviews-talbot.pdf.

Veytsman:TB35-1-115

[3931] Boris Veytsman. Book review: Dynamic Documents with R and knitr, by Yihui Xie. TUGboat, 35(1):115–119, 2014. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb35-1/tb109reviews-xie.pdf.

Berry:TB35-1-120

[3932] Karl Berry. TUG financial statements for 2013. TUGboat, 35(1):120,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
tb109treas.pdf.

Anonymous:TB35-1-121

[3933] Anonymous. TUG institutional members. TUGboat, 35(1):121, 2014. ISSN

0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB35-1-121-2

[3934] Anonymous. TEX consulting and production services. TUGboat, 35(1):121–

122, 2014. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB35-1-123

[3935] Anonymous. Calendar. TUGboat, 35(1):123, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-1/tb109calendar.pdf.

Anonymous:TB35-1-124

[3936] Anonymous. TUG 2014 announcement. TUGboat, 35(1):124, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-1/tb109tug14.
pdf.

Anonymous:2014:CODa

[3937] Anonymous. Contents ordered by difficulty. TUGboat, 35(1):c3, 2014. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-
1/cover3.pdf.

Anonymous:2014:FCa

[3938] Anonymous. Front cover. TUGboat, 35(1):c1, 2014. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-1/cover1.pdf.

Anonymous:2014:IFCa

[3939] Anonymous. Inside front cover. TUGboat, 35(1):c2, 2014. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-1/
cover2.pdf.

Anonymous:2014:TPb

[3940] Anonymous. Title page. TUGboat, 35(2):125, 2014. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-2/titlepage.pdf.

Anonymous:TB35-2-126

[3941] Anonymous. Conference sponsors, participants, program, and photos. TUGboat, 35(2):126–129, 2014. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb35-2/tb110confinfo.pdf;   https:
//tug.org/TUGboat/tb35-2/tb110photos.pdf;   https://
tug.org/TUGboat/tb35-2/tb110program.pdf.

Latchman:TB35-2-130

[3942] David Latchman. TUG 2014 in Portland. TUGboat, 35(2):130–134,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-2/
tb110latchman.pdf.

Kidder:TB35-2-134

[3943] Tracy Kidder. Visiting TUG 2014. TUGboat, 35(2):134, 2014. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-2/tb110latchman.
pdf.

Sharpe:TB35-2-135

[3944] Michael Sharpe. Recent additions to TEX’s font repertoire. TUGboat, 35(2): 135–138, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-2/tb110sharpe.pdf.

Hefferon:TB35-2-139

[3945] Jim Hefferon and Lon Mitchell. Experiences converting from PDF-only to paper. TUGboat, 35(2):139–141, 2014. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb35-2/tb110heff.pdf.

Hogg:TB35-2-142

[3946] Joseph Hogg. Texinfo visits a garden. TUGboat, 35(2):142–144,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-2/
tb110hogg.pdf.

Koch:TB35-2-145

[3947] Richard Koch. MacTEX design philosophy vs. TeXShop design philosophy. TUGboat, 35(2):145–151, 2014. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb35-2/tb110koch.pdf.

Maxwell:TB35-2-152

[3948] Adam Maxwell. TEX Live Utility: a slightly-shiny Mac interface for TEX Live Manager (tlmgr). TUGboat, 35(2):152–156, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-2/tb110maxwell.pdf.

McKenna:TB35-2-157

[3949] Doug McKenna. On tracing the trip test with JSBox. TUGboat, 35(2):157–167,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-2/
tb110mckenna.pdf;   https://www.tug.org/2014/slides/
mckenna-literac.pdf;   https://www.tug.org/tug2014/
slides/mckenna-JSBox.pdf.

Gilbey:TB35-2-168

[3950] Julian Gilbey. Creating (mathematical) jigsaw puzzles using TEX and friends. TUGboat, 35(2):168–172, 2014. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb35-2/tb110gilbey.pdf.

Arora:TB35-2-173

[3951] Pavneet Arora. SUTRA — a workflow for documenting signals. TUGboat, 35(2):173–178, 2014. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb35-2/tb110arora.pdf.

Mertz:TB35-2-179

[3952] Andrew Mertz, William Slough, and Nancy Van Cleave. Typesetting figures for computer science. TUGboat, 35(2):179–191, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-2/tb110mertz.pdf.

Allen:TB35-2-192

[3953] David Allen. Dynamic documents. TUGboat, 35(2):192–194, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-2/tb110allen.
pdf.

Akhmadeeva:TB35-2-195

[3954] Leyla Akhmadeeva and Boris Veytsman. Typography and readability: An experiment with post-stroke patients. TUGboat, 35(2):195–197,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-2/
tb110akhmadeeva.pdf.

Venkatesan:TB35-2-198

[3955] S. K. Venkatesan and C. V. Rajagopal. TEX and copyediting. TUGboat, 35(2): 198–201, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-2/tb110venkatesan.pdf.

Veytsman:TB35-2-202

[3956] Boris Veytsman. An output routine for an illustrated book: Making the FAO Statistical Yearbook. TUGboat, 35(2):202–204, 2014. ISSN 0896-3207. URL

https://github.com/filippogheri/FAOSYBLaTeXpackage;
 https://tug.org/TUGboat/tb35-2/tb110veytsman-code.
dtx;   https://tug.org/TUGboat/tb35-2/tb110veytsman-
code.pdf;   https://tug.org/TUGboat/tb35-2/tb110veytsman.
pdf.

Shikano:TB35-2-205

[3957] Keiichiro Shikano. xml2tex: an easy way to define XML-to-L^{A}TEX converters. TUGboat, 35(2):205–208, 2014. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb35-2/tb110shikano.pdf.

Beezer:TB35-2-209

[3958] Robert A. Beezer. MathBook XML. TUGboat, 35(2):209–211, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-2/tb110beezer.
pdf.

Hammond:TB35-2-212

[3959] William Hammond. Can L^{A}TEX profiles be rendered adequately with static CSS? TUGboat, 35(2):212–218, 2014. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb35-2/tb110hammond.pdf.

Anonymous:TB35-2-219

[3960] Anonymous. TUG 2014 abstracts (Bazargan, Berry, Crossland, Cunningham, de Souza, Doob, Farmer, McKenna, Mittelbach, Moore, Raies, Robertson, Tétreault, Wetmore). TUGboat, 35(2):219–221, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-2/tb110abstracts.pdf.

Parthasarathy:2014:LLL

[3961] S. Parthasarathy. Let’s Learn LaTeX: A hack-to-learn ebook. TUGboat, 35(2): 222, 2014. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb35-2/tb110partha.pdf.

Barnett:TB35-2-223

[3962] Jeffrey Barnett. Book review: Fifty Typefaces That Changed The World, by John Walters. TUGboat, 35(2):223–224, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-2/tb110reviews-walters.
pdf.

Anonymous:TB35-2-225

[3963] Anonymous. TUG 2015 election. TUGboat, 35(2):225, 2014. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb35-2/tb110elec.pdf.

Anonymous:TB35-2-226

[3964] Anonymous. TUG institutional members. TUGboat, 35(2):226, 2014. ISSN

0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB35-2-227

[3965] Anonymous. TUG 2015 announcement. TUGboat, 35(2):227, 2014. ISSN 0896-3207.

Anonymous:TB35-2-228

Anonymous:2014:CODb

[3967] Anonymous. Contents ordered by difficulty. TUGboat, 35(2):c3, 2014. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-
2/cover3.pdf.

Anonymous:2014:FCb

[3968] Anonymous. Front cover. TUGboat, 35(2):c1, 2014. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-2/cover1.pdf.

Anonymous:2014:IFCb

[3969] Anonymous. Inside front cover. TUGboat, 35(2):c2, 2014. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-2/
cover2.pdf.

Anonymous:2014:TPc

[3970] Anonymous. Title page. TUGboat, 35(3):229, 2014. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-3/titlepage.pdf.

Anonymous:TB35-3-230

[3971] Anonymous. TUGboat editorial information. TUGboat, 35(3):230,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-3/
tb111edit.pdf.

Peter:TB35-3-230

[3972] Steve Peter. Ab epistulis. TUGboat, 35(3):230, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-3/tb111pres.pdf.

Beeton:TB35-3-231

[3973] Barbara Beeton. Editorial comments. TUGboat, 35(3):231–232, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-3/tb111beet.
pdf.

Knuth:TB35-3-232

[3974] Donald Knuth. A footnote about ‘Oh, oh, zero’. TUGboat, 35(3):232–234,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-3/
tb111knut-zero.pdf.

Anonymous:TB35-3-235

[3975] Anonymous. Twenty Questions for Donald Knuth (on the occasion of the ePublication of TAOCP). TUGboat, 35(3):235–243, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-3/tb111knut-20.pdf.

Bigelow:TB35-3-244

[3976] Charles Bigelow. A letter on the persistence of (e)books. TUGboat, 35(3):244,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-3/
tb111bigelow.pdf.

Thurnherr:TB35-3-245

[3977] Thomas Thurnherr. L^{A}TEX document class options. TUGboat, 35(3):245–247,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-3/
tb111thurnherr.pdf.

Mittelbach:TB35-3-248

[3978] Frank Mittelbach. How to influence the position of float environments like figure and table in L^{A}TEX? TUGboat, 35(3):248–254, 2014. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb35-3/tb111mitt-float.
pdf.

Beeton:TB35-3-255

[3979] Barbara Beeton. Placing a full-width insert at the bottom of two columns. TUGboat, 35(3):255, 2014. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb35-3/tb111beet-banner.pdf.

Fischer:TB35-3-256

[3980] Ulrike Fischer. biblatex variations. TUGboat, 35(3):256–260, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-3/tb111fischer.
pdf.

Walden:TB35-3-261

[3981] David Walden. Every L^{A}TEX document brings new programming issues. TUGboat, 35(3):261–268, 2014. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb35-3/tb111walden.pdf.

Wilson:TB35-3-269

[3982] Peter Wilson. Glisterings: Lining up. TUGboat, 35(3):269–273, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-3/tb111glister.
pdf.

Neugebauer:TB35-3-274

[3983] Gerd Neugebauer. CTAN goes multi-lingual: Additional language support for the web portal. TUGboat, 35(3):274–275, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-3/tb111neugebauer.pdf.

Solomykov:TB35-3-276

[3984] Basil Solomykov. Obyknovennaya Novaya (Ordinary New Face) in - . TUGboat, 35(3):276, 2014. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb35-3/tb111solomykov.pdf.

Haralambous:TB35-3-277

[3985] Yannis Haralambous. A simple Arabic typesetting system for mixed Latin/ Arabic documents: d.ād. TUGboat, 35(3):277–283, 2014. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb35-3/tb111haralambous.
pdf.

Tennent:TB35-3-284

[3986] Bob Tennent. Visual editing (in a specialized case): prerex. TUGboat, 35(3): 284–286, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-3/tb111tennent.pdf.

Mittelbach:TB35-3-287

[3987] Frank Mittelbach, Will Robertson, and L^{A}TEX3 team. l3build — A modern Lua test suite for TEX programming. TUGboat, 35(3):287–293, 2014. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-3/tb111mitt-
l3build.pdf.

Hoekwater:TB35-3-294

[3988] Taco Hoekwater. MetaPost path resolution isolated. TUGboat, 35(3):294–296,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-3/
tb111hoekwater.pdf.

Wermuth:TB35-3-297

[3989] Udo Wermuth. Typeset MMIX programs with TEX. TUGboat, 35(3):297–308,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-3/
tb111wermuth.pdf.

Stender:TB35-3-309

[3990] Daniel Stender. A Citation Style Language (CSL) workshop. TUGboat, 35(3): 309–314, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-3/tb111stender.pdf.

Berry:TB35-3-315

[3991] Karl Berry. The treasure chest. TUGboat, 35(3):315–316, 2014. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-3/tb111chest.
pdf.

Adams:TB35-3-317

[3992] William Adams. Book review: Practical L^{A}TEX, by George Grätzer. TUGboat, 35(3):317, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-3/tb111reviews-gratzer.pdf.

Andre:TB35-3-318

[3993] Jacques André. Book review: Apprendre à programmer en TEX, by Christian Tellechea. TUGboat, 35(3):318, 2014. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb35-3/tb111reviews-tellechea.pdf.

Berry:TB35-3-319

[3994] Michael Berry. Book review: The Imitation Game, by Jim Ottaviani and Leland Purvis. TUGboat, 35(3):319, 2014. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb35-3/tb111reviews-ottaviani.pdf.

Talbot:TB35-3-320

[3995] Nicola Talbot. Book review: Let’s Learn L^{A}TEX, by S. Parthasarathy. TUGboat, 35(3):320–321, 2014. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb35-3/tb111reviews-partha.pdf.

Anonymous:TB35-3-322

[3996] Anonymous. Die TEXnische Komödie: contents of issues 2–3/2014. TUGboat, 35(3):322, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-3/tb111komo.pdf.

Anonymous:TB35-3-323

[3997] Anonymous. Les Cahiers GUTenberg: Contents of issue 57 (2012). TUGboat, 35(3):323, 2014. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb35-3/tb111cahi.pdf.

Anonymous:TB35-3-323-2

[3998] Anonymous. TEX Development Fund 2013 report. TUGboat, 35(3):323,

2014. ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-3/
tb111devfund.pdf.

Anonymous:TB35-3-324

[3999] Anonymous. TUG 2015 election. TUGboat, 35(3):324, 2014. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb35-3/tb111elec.pdf.

Anonymous:TB35-3-325

[4000] Anonymous. TUG membership form. TUGboat, 35(3):325, 2014. ISSN 0896-

3207. URL https://tug.org/join.html.

Anonymous:TB35-3-326

[4001] Anonymous. TUG institutional members. TUGboat, 35(3):326, 2014. ISSN

0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB35-3-326-2

[4002] Anonymous. TEX consulting and production services. TUGboat, 35(3):326–

327, 2014. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB35-3-327

[4003] Anonymous. TUG 2015 announcement. TUGboat, 35(3):327, 2014. ISSN

0896-3207. URL https://tug.org/TUGboat/tb35-3/tb111consultants.
pdf.

Anonymous:TB35-3-328

[4004] Anonymous. Calendar. TUGboat, 35(3):328, 2014. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb35-3/tb111calendar.pdf.

Anonymous:2014:CODc

[4005] Anonymous. Contents ordered by difficulty. TUGboat, 35(3):c3, 2014. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-
3/cover3.pdf.

Anonymous:2014:FCc

[4006] Anonymous. Front cover. TUGboat, 35(3):c1, 2014. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb35-3/cover1.pdf.

Anonymous:2014:IFCc

[4007] Anonymous. Inside front cover. TUGboat, 35(3):c2, 2014. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb35-3/
cover2.pdf.

Anonymous:TB36-1-1

[4008] Anonymous. Title page. TUGboat, 36(1):1, 2015. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb36-1/titlepage.pdf.

Anonymous:TB36-1-0

[4009] Anonymous. Complete issue 36:1 as one PDF (11.8MB). TUGboat, 36(1):1–

73, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-
1/tb112complete.pdf;   https://tug.org/TUGboat/tb36-
1/titlepage.pdf.

Anonymous:TB36-1-2

[4010] Anonymous. Editorial information. TUGboat, 36(1):2, 2015. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb36-1/tb112edit.pdf.

Peter:TB36-1-2

[4011] Steve Peter. Ab epistulis. TUGboat, 36(1):2, 2015. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb36-1/tb112pres.pdf.

Beeton:TB36-1-3

[4012] Barbara Beeton. Editorial comments. TUGboat, 36(1):3–6, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-1/tb112beet.pdf.

Beeton:TB36-1-7

[4013] Barbara Beeton. Hyphenation exception log. TUGboat, 36(1):7, 2015. ISSN

0896-3207. URL https://tug.org/TUGboat/tb36-1/tb112hyf.
pdf.

Phinney:TB36-1-8

[4014] Thomas Phinney. What does a typical brief for a new typeface look like? TUGboat, 36(1):8–9, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-1/tb112phinney.pdf.

Sharpe:TB36-1-10

[4015] Michael Sharpe. Inconsolata unified. TUGboat, 36(1):10, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-1/tb112sharpe.
pdf.

Wilson:TB36-1-11

[4016] Peter Wilson. A TUG postcard or, the trials of a letterpress printer. TUGboat, 36 (1):11–14, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-1/tb112wilson.pdf.

Flynn:TB36-1-15

[4017] Peter Flynn. Typographers’ Inn. TUGboat, 36(1):15–16, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-1/tb112inn.pdf.

LPT:TB36-1-17

[4018] L^{A}TEX Project Team. L^{A}TEX news, issue 21, May 2014. TUGboat, 36(1):17–18,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-1/
tb112ltnews.pdf.

Wright:TB36-1-19

[4019] Joseph Wright. Beamer overlays beyond the \visible. TUGboat, 36(1):19,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-1/
tb112wright.pdf.

Wilson:TB36-1-20

[4020] Peter Wilson. Glisterings: Here or there; parallel texts; abort the compilation. TUGboat, 36(1):20–24, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-1/tb112glister.pdf.

Lupkowski:TB36-1-25

[4021] Pawe� �upkowski. Online L^{A}TEX editors and other resources. TUGboat, 36 (1):25–27, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-1/tb112lupkowski.pdf.

Hagen:TB36-1-28

[4022] Hans Hagen. Exporting XML and ePub from ConTEXt. TUGboat, 36(1):28–31,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-1/
tb112hagen-epub.pdf.

Mittelbach:TB36-1-32

[4023] Frank Mittelbach. The box–glue-penalty algebra of TEX and its use of \prevdepth. TUGboat, 36(1):32–36, 2015. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb36-1/tb112mitt.pdf.

Cereda:TB36-1-37

[4024] Paulo Roberto Massa Cereda. The bird and the lion: arara. TUGboat, 36 (1):37–40, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-1/tb112cereda.pdf.

Scarso:TB36-1-41

[4025] Luigi Scarso. The SWIGLIB project. TUGboat, 36(1):41–47, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-1/tb112scarso.
pdf.

Hagen:TB36-1-48

[4026] Hans Hagen. Still tokens: LuaTEX scanners. TUGboat, 36(1):48–54,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-1/
tb112hagen-scan.pdf.

Berry:TB36-1-55

[4027] Karl Berry. The treasure chest. TUGboat, 36(1):55–56, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-1/tb112chest.
pdf.

Walden:TB36-1-57

[4028] David Walden. Book review: Algorithmic Barriers Falling: P = NP?, by Donald E. Knuth and Edgar Daylight. TUGboat, 36(1):57, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-1/tb112reviews-
daylight.pdf.

Veytsman:TB36-1-58

[4029] Boris Veytsman. Book review: History of the Linotype Company, by Frank Romano. TUGboat, 36(1):58–59, 2015. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb36-1/tb112reviews-romano.pdf.

Anonymous:TB36-1-60

[4030] Anonymous. BachoTEX 2014 proceedings. TUGboat, 36(1):60–62,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-1/
tb112gust.pdf.

Anonymous:TB36-1-63

[4031] Anonymous. Die TEXnische Komödie: contents of issues 4/2014–1/2015. TUGboat, 36(1):63, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-1/tb112komo.pdf.

Christiansen:TB36-1-64

[4032] Kaja Christiansen and TUG Election committee. TUG 2015 election. TUGboat, 36(1):64–67, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-1/tb112elec.pdf.

Berry:TB36-1-68

[4033] Karl Berry. TUG financial statements for 2014. TUGboat, 36(1):68,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-1/
tb112treas.pdf.

Anonymous:TB36-1-69a

[4034] Anonymous. TUG institutional members. TUGboat, 36(1):69, 2015. ISSN

0896-3207. URL https://www.tug.org/instmem.html.

Anonymous:TB36-1-69b

[4035] Anonymous. TEX consulting and production services. TUGboat, 36(1):69–70,

2015. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB36-1-71

[4036] Anonymous. TUG 2015 announcement. TUGboat, 36(1):71, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-1/tb112tug15.
pdf.

Anonymous:TB36-1-72

[4037] Anonymous. Calendar. TUGboat, 36(1):72, 2015. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb36-1/tb112calendar.pdf.

Anonymous:TB36-1-c1

[4038] Anonymous. Front cover. TUGboat, 36(1):c1, 2015. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb36-1/cover1.pdf.

Anonymous:TB36-1-c2

[4039] Anonymous. Inside front cover. TUGboat, 36(1):c2, 2015. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb36-1/cover2.pdf.

Anonymous:TB36-1-c3

[4040] Anonymous. Contents ordered by difficulty. TUGboat, 36(1):c3, 2015. ISSN

0896-3207. URL https://tug.org/TUGboat/tb36-1/cover3.
pdf.

Anonymous:2015:TPa

[4041] Anonymous. Title page. TUGboat, 36(2):73, 2015. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-2/titlepage.pdf.

Anonymous:TB36-2-74

[4042] Anonymous. Conference sponsors, participants, program, and photos. TUGboat, 36(2):74–79, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-2/tb113confinfo.pdf.

Anonymous:2015:TCP

[4043] Anonymous. TUG 2015 conference program. TUGboat, 36(2):75, 2015. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-
2/tb113program.pdf.

Anonymous:2015:TPb

[4044] Anonymous. TUG 2015 photos. TUGboat, 36(2):76–79, 2015. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113photos.pdf.

Schaa:TB36-2-80

[4045] Volker R. W. Schaa. Typographer’s banquet. TUGboat, 36(2):80–81,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113banquet.pdf.

Kottwitz:TB36-2-82

[4046] Stefan Kottwitz. TUG 2015 conference report. TUGboat, 36(2):82–88,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113kottwitz.pdf.

Beeton:TB36-2-89

[4047] Barbara Beeton. In memoriam. TUGboat, 36(2):89, 2015. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb36-2/tb113beet.pdf.

Beeton:TB36-2-90

[4048] Barbara Beeton. Pierre MacKay, 1933–2015. TUGboat, 36(2):90–91,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113mackay.pdf.

Knuth:TB36-2-92

[4049] Donald Knuth. Dedication to Hermann Zapf, 1918–2015. TUGboat, 36(2):92,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113zapf-knuth.pdf.

Thanh:TB36-2-93

[4050] Hàn Thê´ Thành. Farewell Hermann Zapf. TUGboat, 36(2):93, 2015. ISSN

0896-3207. URL https://tug.org/TUGboat/tb36-2/tb113zapf-
thanh.pdf.

Holmes:TB36-2-93

[4051] Kris Holmes. Remembering Hermann Zapf. TUGboat, 36(2):93–94,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113zapf-holmes.pdf.

Karow:TB36-2-95

[4052] Peter Karow. Digital typography with Hermann Zapf. TUGboat, 36(2):95–99,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113zapf-karow.pdf.

Andre:TB36-2-100

[4053] Jacques André and Alan Marshall. Richard Southall: 1937–2015. TUGboat, 36(2):100–102, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-2/tb113southall.pdf.

Frambach:TB36-2-103

[4054] Erik Frambach, Jerzy Ludwichowski, and Philip Taylor. Memories of Kees: C. G. van der Laan, 1943–2015. TUGboat, 36(2):103–104, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-2/tb113laan.pdf.

Wright:TB36-2-105

[4055] Joseph Wright. Development of the UK TEX FAQ. TUGboat, 36(2):105,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113wright-faq.pdf.

Robertson:TB36-2-106

[4056] Will Robertson. Single- and multi-letter identifiers in Unicode mathematics. TUGboat, 36(2):106–108, 2015. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb36-2/tb113robertson.pdf.

Veytsman:TB36-2-109

[4057] Boris Veytsman and Leyla Akhmadeeva. Trilingual templates for an educational institute in Bashkortostan, Russia. TUGboat, 36(2):109–113,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113veytsman-trilingual.pdf.

Wright:TB36-2-114

[4058] Joseph Wright. Joseph’s adventures in Unicodeland. TUGboat, 36(2):114–116,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
tb113wright-unicode.pdf.

Wright:TB36-2-117

[4059] Joseph Wright. Through the \parshape, and what Joseph found there. TUGboat, 36(2):117–118, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-2/tb113wright-parshape.pdf.

Veytsman:TB36-2-119

[4060] Boris Veytsman. TEX and controlled access to information. TUGboat, 36(2): 119–122, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-2/tb113veytsman-access.pdf.

Schrod:TB36-2-123

[4061] Joachim Schrod. DocCenter — TEXing 11 million documents a year. TUGboat, 36(2):123–127, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-2/tb113schrod.pdf.

Hejda:TB36-2-128

[4062] Tom Hejda. Preparing L^{A}TEX classes for journal articles and university theses. TUGboat, 36(2):128–129, 2015. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb36-2/tb113hejda.pdf.

Olsak:TB36-2-130

[4063] Petr Olšák. The CTUstyle template for student theses. TUGboat, 36(2):130–

132, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-
2/tb113olsak.pdf.

Veytsman:TB36-2-133

[4064] Boris Veytsman and Michael Cohen. New multibibliography package nmbib. TUGboat, 36(2):133–135, 2015. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb36-2/tb113veytsman-nmbib.pdf.

Radhakrishnan:TB36-2-136

[4065] C. V. Radhakrishnan, Hàn Thê´ Thành, Ross Moore, and Peter Selinger. Generating PDF/X- and PDF/A-compliant PDFs with pdfTEX — pdfx.sty. TUGboat, 36(2):136–142, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-2/tb113radhakrishnan.pdf.

Schulz:TB36-2-143

[4066] Herbert Schulz. TEXShop’s key bindings vs. macros vs. command completion. TUGboat, 36(2):143–144, 2015. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb36-2/tb113schulz.pdf.

Venkatesan:TB36-2-145

[4067] S. K. Venkatesan. TEX as a three-stage rocket: cookie-cutter page breaking. TUGboat, 36(2):145–148, 2015. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb36-2/tb113venkatesan.pdf.

Gregorio:TB36-2-149

[4068] Enrico Gregorio. Recollections of a spurious space catcher. TUGboat, 36(2): 149–161, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-2/tb113gregorio.pdf.

Hagen:TB36-2-162

[4069] Hans Hagen. When to stop …. TUGboat, 36(2):162–170, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-2/tb113hagen.
pdf.

Anonymous:TB36-2-171

[4070] Anonymous. TUG 2015 abstracts (Bazargan, Cretel, Drümmer, Gessler, Hagen, Jackowski). TUGboat, 36(2):171, 2015. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb36-2/tb113abstracts.pdf.

Anonymous:2015:TKC

[4071] Anonymous. Die TeXnische Komödie: Contents of issues 2–3/2015. TUGboat, 36(2):172–173, 2015. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb36-2/tb113komo.pdf.

Anonymous:TB36-2-173

[4072] Anonymous. Eutypon: contents of issue 32–33 (October 2014). TUGboat, 36(2):173, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-2/tb113eutypon.pdf.

Anonymous:TB36-2-174

[4073] Anonymous. TUG 2015 election. TUGboat, 36(2):174, 2015. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb36-2/tb113elec.pdf.

Anonymous:TB36-2-174-2

[4074] Anonymous. TUG institutional members. TUGboat, 36(2):174, 2015. ISSN 0896-3207.

Anonymous:TB36-2-175

[4075] Anonymous. TEX consulting and production services. TUGboat, 36(2):175,

2015. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB36-2-176

Anonymous:2015:CODa

[4077] Anonymous. Contents ordered by difficulty. TUGboat, 36(2):c3, 2015. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-
2/cover3.pdf.

Anonymous:2015:FCa

[4078] Anonymous. Front cover. TUGboat, 36(2):c1, 2015. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-2/cover1.pdf.

Anonymous:2015:IFCa

[4079] Anonymous. Inside front cover. TUGboat, 36(2):c2, 2015. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-2/
cover2.pdf.

Anonymous:2015:TPc

[4080] Anonymous. Title page. TUGboat, 36(3):177, 2015. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-3/titlepage.pdf.

Anonymous:2015:EI

[4081] Anonymous. Editorial information. TUGboat, 36(3):178, 2015. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
tb114edit.pdf.

Board:TB36-3-179

[4082] TEX Users Group Board. From the Board of Directors. TUGboat, 36(3):179,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
tb114board.pdf.

Beeton:TB36-3-180

[4083] Barbara Beeton. Editorial comments. TUGboat, 36(3):180–181, 2015. ISSN

0896-3207. URL https://tug.org/TUGboat/tb36-3/tb114beet.
pdf.

Preining:TB36-3-182

[4084] Norbert Preining. Adrian Frutiger, 1928–2015. TUGboat, 36(3):182–183,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
tb114frutiger.pdf.

Schrod:TB36-3-184

[4085] Joachim Schrod. Thomas Koch, 1964–2014. TUGboat, 36(3):184,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
tb114koch.pdf.

Kottwitz:TB36-3-185

[4086] Stefan Kottwitz. DANTE e.V. 2015 meeting reports. TUGboat, 36(3):185–187,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
tb114kottwitz.pdf.

Laube:TB36-3-188

[4087] Simon Laube. TEX in schools: Just Say Yes! TUGboat, 36(3):188–189,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
tb114laube.pdf.

Romer:TB36-3-190

[4088] Linus Romer. Smoky letters. TUGboat, 36(3):190, 2015. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb36-3/tb114romer.pdf.

Bigelow:TB36-3-191

[4089] Charles Bigelow. About the DK versions of Lucida. TUGboat, 36(3):191–199,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
tb114bigelow.pdf.

Hoekwater:TB36-3-200

[4090] Taco Hoekwater. History of cookbooks. TUGboat, 36(3):200–207,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
tb114hoekwater.pdf.

Flynn:TB36-3-208

[4091] Peter Flynn. Typographers’ Inn. TUGboat, 36(3):208–209, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-3/tb114inn.pdf.

Team:TB36-3-210

[4092] L^{A}TEX Project Team. L^{A}TEX news, issue 22, January 2015. TUGboat, 36(3): 210–211, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-3/tb114ltnews22.pdf.

Team:TB36-3-212

[4093] L^{A}TEX Project Team. L^{A}TEX news, issue 23, October 2015. TUGboat, 36(3): 212–213, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-3/tb114ltnews23.pdf.

Thurnherr:TB36-3-214

[4094] Thomas Thurnherr. Introduction to list structures in L^{A}TEX. TUGboat, 36(3): 214–216, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-3/tb114thurnherr.pdf.

Kumar:TB36-3-217

[4095] Anagha Kumar. gradstudentresume: A document class for graduate student CVs. TUGboat, 36(3):217–219, 2015. ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb36-3/tb114kumar.pdf.

Wilson:TB36-3-220

[4096] Peter Wilson. Glisterings: Longest string; marching along; a blank argument; a centered table of contents. TUGboat, 36(3):220–226, 2015. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb36-3/tb114glister.pdf.

Niederberger:TB36-3-227

[4097] Clemens Niederberger. Chemistry in L^{A}TEX 2ε — an overview of existing packages and possibilities. TUGboat, 36(3):227–233, 2015. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb36-3/tb114niederberger.pdf.

Wright:TB36-3-234

[4098] Joseph Wright. Automating L^{A}TEX(3) testing. TUGboat, 36(3):234–236,

2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
tb114wright.pdf.

Scarso:TB36-3-237

[4099] Luigi Scarso. Two applications of SWIGLIB: GraphicsMagick and Ghostscript. TUGboat, 36(3):237–242, 2015. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb36-3/tb114scarso.pdf.

Wermuth:TB36-3-243

[4100] Udo Wermuth. Typesetting the “Begriffsschrift” by Gottlob Frege in plain TEX. TUGboat, 36(3):243–256, 2015. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb36-3/tb114wermuth.pdf.

Murzynowski:TB36-3-257

[4101] Grzegorz Murzynowski. GMOA, the ‘general manipulation of arguments’: an extension to the l3expan package of the expl3 bundle and language. TUGboat, 36(3):257–268, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-3/tb114murzynowski.pdf.

Berry:TB36-3-268

[4102] Karl Berry. Production notes. TUGboat, 36(3):268, 2015. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb36-3/tb114prod.pdf.

Berry:TB36-3-269

[4103] Karl Berry. The treasure chest. TUGboat, 36(3):269–270, 2015. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-3/tb114chest.
pdf.

Veytsman:TB36-3-271

[4104] Boris Veytsman. An online glossary of typographic terms by Janie Kliever. TUGboat, 36(3):271, 2015. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb36-3/tb114reviews-kliever.pdf.

Anonymous:TB36-3-272

[4105] Anonymous. GUST: EuroBachoTEX 2015 proceedings. TUGboat, 36(3):272–

273, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-
3/tb114gust.pdf.

Anonymous:TB36-3-273

[4106] Anonymous. Die TEXnische Komödie: contents of issue 4/2015. TUGboat, 36(3):273, 2015. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb36-3/tb114komo.pdf.

Anonymous:TB36-3-274

[4107] Anonymous. TUG institutional members. TUGboat, 36(3):274, 2015. ISSN 0896-3207.

Benton:2015:S

[4108] Jim Benton. A summons. TUGboat, 36(3):274, 2015. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb36-3/tb114cartoon.pdf.

Anonymous:TB36-3-275

[4109] Anonymous. TEX consulting and production services. TUGboat, 36(3):275,

2015. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:2015:C

[4110] Anonymous. Calendar. TUGboat, 36(3):276, 2015. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb36-3/tb114calendar.
pdf.

Anonymous:TB36-3-276K

[4111] Anonymous. Calendar. TUGboat, 36(3):276K–276, 2015. ISSN 0896-3207.

Anonymous:2015:CODb

[4112] Anonymous. Contents ordered by difficulty. TUGboat, 36(3):c3, 2015. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-
3/cover3.pdf.

Anonymous:2015:FCb

Anonymous:2015:IFCb

[4114] Anonymous. Inside front cover. TUGboat, 36(3):c2, 2015. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb36-3/
cover2.pdf.

Anonymous:2016:TPa

[4115] Anonymous. Title page. TUGboat, 37(1):1, 2016. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-1/titlepage.pdf.

Anonymous:TB37-1-2

[4116] Anonymous. TUGboat editorial information. TUGboat, 37(1):2, 2016. ISSN

0896-3207. URL https://tug.org/TUGboat/tb37-1/tb115edit.
pdf.

Beeton:TB37-1-3

[4117] Barbara Beeton. Editorial comments: R.I.P. Sebastian Rahtz, 1955– 2016. George Greenwade, 1956–2003. Peter Breitenlohner 1940–2015. biblatex — request for feedback. L^{A}TEX courses for credit. Cooper Type. more typography videos. TUGboat, 37(1):3–5, 2016. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb37-1/tb115beet.pdf.

Lammarsch:TB37-1-5

[4118] Joachim Lammarsch and Marion Lammarsch. Peter Breitenlohner, 1940–2015. TUGboat, 37(1):5, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-1/tb115breitenlohner.pdf.

Veytsman:TB37-1-6

[4119] Boris Veytsman. The continuing TEX Users Group membership drive. TUGboat, 37(1):6–8, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-1/tb115veytsman.pdf.

Ludwichowski:TB37-1-9

[4120] Jerzy Ludwichowski. ATypI 2016 with GUST participating. TUGboat, 37(1):9,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115ludwichowski.pdf.

Fine:TB37-1-10

[4121] Jonathan Fine. The Board’s suspension of the President. TUGboat, 37(1):10–

11, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-
1/tb115fine.pdf.

Hosny:TB37-1-12

[4122] Khaled Hosny. The libertine gets mathematical. TUGboat, 37(1):12,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115hosny.pdf.

Team:TB37-1-13

[4123] L^{A}TEX Project Team. L^{A}TEX news, issue 24, January 2016. TUGboat, 37(1):13–

14, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-
1/tb115ltnews24.pdf.

Team:TB37-1-15

[4124] L^{A}TEX Project Team. L^{A}TEX news, issue 25, March 2016. TUGboat, 37(1):15,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115ltnews25.pdf.

Thurnherr:TB37-1-16

[4125] Thomas Thurnherr. On managing large documents. TUGboat, 37(1):16–17,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115thurnherr.pdf.

Kumar:TB37-1-18

[4126] Anagha Kumar. medstarbeamer: a new beamer class. TUGboat, 37(1):18–

21, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-
1/tb115kumar.pdf.

Wilson:TB37-1-22

[4127] Peter Wilson. Glisterings: Assemblies; table talk. TUGboat, 37(1):22–24,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115glister.pdf.

Al-Safi:TB37-1-25

[4128] Sabri Al-Safi. Randomising assignments with SageTEX. TUGboat, 37(1):25–

27, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-
1/tb115alsafi.pdf.

Fehd:TB37-1-28

[4129] Ron Fehd. Indexing: goals, strategies and tactics. TUGboat, 37(1):28–36,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115fehd.pdf.

Absil:TB37-1-37

[4130] Frans Absil. TEXShop review. TUGboat, 37(1):37–38, 2016. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb37-1/tb115absil.pdf.

Knypstra:TB37-1-39

[4131] Sytse Knypstra. TEXworks: a simple GUI with advanced options. TUGboat, 37 (1):39–40, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-1/tb115knypstra.pdf.

Kroonenberg:TB37-1-41

[4132] Siep Kroonenberg. TEXstudio: especially for L^{A}TEX newbies. TUGboat, 37 (1):41–44, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-1/tb115kroonenberg.pdf.

Preining:TB37-1-45

[4133] Norbert Preining. 10 years of TEX Live in Debian. TUGboat, 37(1):45–47,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115preining.pdf.

Parashchenko:TB37-1-48

[4134] Oleg Parashchenko. Paragraph designer with galley approach. TUGboat, 37 (1):48–52, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-1/tb115parashchenko.pdf.

Hagen:TB37-1-53

[4135] Hans Hagen. LuaTEX 0.90 backend changes for PDF and more. TUGboat, 37 (1):53–57, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-1/tb115hagen-pdf.pdf.

Hagen:TB37-1-58

[4136] Hans Hagen. Still expanding LuaTEX: possibly useful extensions. TUGboat, 37(1):58–65, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-1/tb115hagen-expand.pdf.

Braslau:TB37-1-66

[4137] Alan Braslau and Hans Hagen. MetaPost arrowhead variants. TUGboat, 37(1):

66, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-
1/tb115hagen-lang.pdf.

Hagen:TB37-1-66

[4138] Hans Hagen. Hyphenation languages in LuaTEX 0.90. TUGboat, 37(1):66,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115hagen-lang.pdf.

Wilson:TB37-1-67

[4139] Peter Wilson. A personal book catalogue: bookdb. TUGboat, 37(1):67–71,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115wilson.pdf.

Olsak:TB37-1-71

[4140] Petr Olšák. OPmac-bib: citations using *.bib files with no external program. TUGboat, 37(1):71–78, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-1/tb115olsak-bib.pdf.

Wright:TB37-1-79

[4141] Joseph Wright. Exploring \romannumeral and expansion. TUGboat, 37 (1):79–81, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-1/tb115wright.pdf.

Olsak:TB37-1-82

[4142] Petr Olšák. The apnum package: arbitrary precision numbers implemented in TEX macros. TUGboat, 37(1):82–88, 2016. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb37-1/tb115olsak-apnum.pdf.

Berry:TB37-1-89

[4143] Karl Berry. The treasure chest. TUGboat, 37(1):89–90, 2016. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-1/tb115chest.
pdf.

Veytsman:TB37-1-91

[4144] Boris Veytsman. L^{A}TEX for Administrative Work by Nicola Talbot. TUGboat, 37(1):91–92, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-1/tb115reviews-talbot.pdf.

Thomas:TB37-1-93

[4145] Charles Thomas. Tout ce que vous avez toujours voulu savoir sur L^{A}TEX sans jamais oser le demander by Vincent Lozano. TUGboat, 37(1):93,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115reviews-lozano.pdf.

Veytsman:TB37-1-94

[4146] Boris Veytsman. Giambattista Bodoni: His Life and His World by Valerie Lester. TUGboat, 37(1):94–96, 2016. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb37-1/tb115reviews-lester.pdf.

Walden:TB37-1-97

[4147] David Walden. Note on the publisher of the Bodoni book: David R. Godine. TUGboat, 37(1):97–98, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-1/tb115walden.pdf.

Berry:TB37-1-98

[4148] Karl Berry. Production notes. TUGboat, 37(1):98, 2016. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb37-1/tb115komo.pdf.

Anonymous:TB37-1-98

[4149] Anonymous. Die TEXnische Komödie: contents of issue 1/2016. TUGboat, 37 (1):98–99, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-1/tb115komo.pdf.

Anonymous:TB37-1-99

[4150] Anonymous. MAPS: contents of issue 46 (2015). TUGboat, 37(1):99–100,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115eutypon.pdf.

Anonymous:TB37-1-99-2

[4151] Anonymous. Eutypon: contents of issue 34–35 (October 2015). TUGboat, 37 (1):99–100, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-1/tb115eutypon.pdf.

Anonymous:TB37-1-100

[4152] Anonymous. TUG institutional members. TUGboat, 37(1):100, 2016. ISSN 0896-3207.

Hoppner:TB37-1-101

[4153] Klaus Höppner. TUG financial statements for 2014. TUGboat, 37(1):101,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
tb115treas.pdf.

Anonymous:TB37-1-102

[4154] Anonymous. TEX consulting and production services. TUGboat, 37(1):102,

2016. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB37-1-103

[4155] Anonymous. Calendar. TUGboat, 37(1):103, 2016. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb37-1/tb115calendar.pdf.

Anonymous:TB37-1-104

[4156] Anonymous. CTAN. TUGboat, 37(1):104, 2016. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb37-1/tb115ctan.pdf.

Anonymous:2016:CODa

[4157] Anonymous. Contents ordered by difficulty. TUGboat, 37(1):c3, 2016. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-
1/cover3.pdf.

Anonymous:2016:FCa

[4158] Anonymous. Front cover. TUGboat, 37(1):c1, 2016. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-1/cover1.pdf.

Anonymous:2016:IFCa

[4159] Anonymous. Inside front cover. TUGboat, 37(1):c2, 2016. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-1/
cover2.pdf.

Anonymous:2016:TPb

[4160] Anonymous. Title page. TUGboat, 37(2):105, 2016. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-2/titlepage.pdf.

Arora:TB37-2-106

[4161] Pavneet Arora. Passport to the TEX canvas. TUGboat, 37(2):106–110,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-2/
tb116arora-canvas.pdf.

Anonymous:TB37-2-111

[4162] Anonymous. Conference sponsors, participants, program, and photos. TUGboat, 37(2):111–114, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-2/tb116confinfo.pdf.

Anonymous:2016:TCP

[4163] Anonymous. TUG 2016 conference program. TUGboat, 37(2):112, 2016. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-2/tb116program.pdf.

Anonymous:2016:TPc

[4164] Anonymous. TUG 2016 photos. TUGboat, 37(2):113, 2016. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-2/
tb116photos.pdf.

Preining:TB37-2-115

[4165] Norbert Preining. TUG 2016 in Toronto. TUGboat, 37(2):115–124,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-2/
tb116preining.pdf.

Kottwitz:TB37-2-125

[4166] Stefan Kottwitz. TUG 2016 Annual General Meeting informal report. TUGboat, 37(2):125, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-2/tb116kottwitz.pdf.

Burnard:TB37-2-126

[4167] Lou Burnard. Sebastian Rahtz (1955–2016): A brief memoir. TUGboat, 37(2): 126–128, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-2/tb116rahtz-burnard.pdf.

Mittelbach:TB37-2-129

[4168] Frank Mittelbach and Joan Richmond. R.I.P. — S.P.Q.R Sebastian Patrick Quintus Rahtz (13.2.1955–15.3.2016). TUGboat, 37(2):129–130, 2016. ISSN

0896-3207. URL https://tug.org/TUGboat/tb37-2/tb116rahtz-
mitt.pdf.

Walden:TB37-2-131

[4169] David Walden. Interview with Pavneet Arora. TUGboat, 37(2):131–136,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-2/
tb116arora-interview.pdf.

Clark:TB37-2-137

[4170] Joe Clark. Type in the Toronto subway. TUGboat, 37(2):137–147,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-2/
tb116clark.pdf.

Akhmadeeva:TB37-2-148

[4171] Leila Akhmadeeva, Rinat Gizatullin, and Boris Veytsman. Are justification and hyphenation good or bad for the reader? Preliminary experimental results. TUGboat, 37(2):148–151, 2016. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb37-2/tb116akhmadeeva.pdf.

Walden:TB37-2-152

[4172] David Walden. An informal look into the history of digital typography. TUGboat, 37(2):152–153, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-2/tb116walden.pdf.

Bigelow:TB37-2-154

[4173] Charles Bigelow. A short history of the Lucida math fonts. TUGboat, 37(2): 154–160, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-2/tb116bigelow-lucidamath.pdf.

Sharpe:TB37-2-161

[4174] Michael Sharpe. New font offerings: Cochineal, Nimbus15, LibertinusT1Math. TUGboat, 37(2):161–162, 2016. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb37-2/tb116sharpe.pdf.

Choi:TB37-2-163

[4175] Jaeyoung Choi, Sungmin Kim, Hojin Lee, and Geunho Jeong. MFCONFIG: A plug-in module for the Freetype rasterizer. TUGboat, 37(2): 163–170, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-2/tb116choi.pdf.

Bayar:TB37-2-171

[4176] Abdelouahad Bayar. Towards an operational (La)TEX package supporting optical scaling of dynamic mathematical symbols. TUGboat, 37(2):171–179,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-2/
tb116bayar.pdf.

Hefferon:TB37-2-180

[4177] Jim Hefferon. A L^{A}TEX reference manual. TUGboat, 37(2):180–181,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-2/
tb116heff.pdf.

Skala:TB37-2-182

[4178] Matthew Skala. Astrological charts with horoscop and starfont. TUGboat, 37(2):182, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-2/tb116skala.pdf.

Veytsman:TB37-2-183

[4179] Boris Veytsman. Remaking the ACM L^{A}TEX styles. TUGboat, 37(2):183–186,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-2/
tb116veytsman.pdf.

Poore:TB37-2-187

[4180] Geoffrey Poore. Advances in PythonTEX with an introduction to fvextra. TUGboat, 37(2):187–192, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-2/tb116poore.pdf.

Tulett:TB37-2-193

[4181] David Tulett. Development of an e-textbook using L^{A}TEX and PSTricks. TUGboat, 37(2):193–200, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-2/tb116tulett.pdf.

Gagne:TB37-2-200

[4182] Christian Gagné. An Emacs-based writing workflow inspired by TEX and WEB, targeting the Web. TUGboat, 37(2):200–203, 2016. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb37-2/tb116gagne.pdf.

Garcia-DeCastro:TB37-2-204

[4183] Federico Garcia-De Castro. TEXcel? an unexpected use for TEX. TUGboat, 37(2):204–208, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-2/tb116garcia.pdf.

Miklavec:TB37-2-209

[4184] Mojca Miklavec and Arthur Reutenauer. Hyphenation in TEX and elsewhere, past and future. TUGboat, 37(2):209–213, 2016. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb37-2/tb116miklavec.pdf.

Cohen:TB37-2-214

[4185] Michael Cohen, Blanca Mancilla, and John Plaice. Zebrackets: A score of years and delimiters. TUGboat, 37(2):214–221, 2016. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb37-2/tb116cohen.pdf.

Banerjee:TB37-2-222

[4186] Amartyo Banerjee and S. K. Venkatesan. A Telegram bot for printing L^{A}TEX files. TUGboat, 37(2):222–228, 2016. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb37-2/tb116banerjee.pdf.

Veytsman:TB37-2-229

[4187] Boris Veytsman. Palatino: The natural history of a typeface by Robert Bringhurst. TUGboat, 37(2):229–231, 2016. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb37-2/tb116reviews-bringhurst.pdf.

Walden:TB37-2-232

[4188] David Walden. A Truck Full of Money by Tracy Kidder. TUGboat, 37(2): 232–233, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-2/tb116reviews-kidder.pdf.

Anonymous:TB37-2-234

[4189] Anonymous. TUG 2016 abstracts (Bazargan, Bigelow, Bringhurst, Claudio, Inkster, Kottwitz, Larson, Mittelbach, Preining, Reutenauer). TUGboat, 37(2): 234–237, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-2/tb116abstracts.pdf.

DeMeritt:TB37-2-238

[4190] Susan DeMeritt. TUG 2016 Annual General Meeting minutes. TUGboat, 37(2): 238–239, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-2/tb116agm.pdf.

Board:TB37-2-240

[4191] TUG Board. Report: Suspension of Kaveh Bazargan as TUG President. TUGboat, 37(2):240–248, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-2/tb116suspension.pdf.

Anonymous:TB37-2-249

[4192] Anonymous. TUG 2017 election. TUGboat, 37(2):249, 2016. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb37-2/tb116elec.pdf.

Anonymous:TB37-2-250

[4193] Anonymous. TUG institutional members. TUGboat, 37(2):250, 2016. ISSN

0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB37-2-250-2

[4194] Anonymous. TEX consulting and production services. TUGboat, 37(2):250–

251, 2016. ISSN 0896-3207. URL https://tug.org/consultants.
html.

Anonymous:TB37-2-252

[4195] Anonymous. Calendar. TUGboat, 37(2):252, 2016. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb37-2/tb116calendar.pdf.

Anonymous:2016:CODb

[4196] Anonymous. Contents ordered by difficulty. TUGboat, 37(2):c3, 2016. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-
2/cover3.pdf.

Anonymous:2016:FCb

[4197] Anonymous. Front cover. TUGboat, 37(2):c1, 2016. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-2/cover1.pdf.

Anonymous:2016:IFCb

[4198] Anonymous. Inside front cover. TUGboat, 37(2):c2, 2016. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-2/
cover2.pdf.

Anonymous:2016:TPd

[4199] Anonymous. Title page. TUGboat, 37(3):253, 2016. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-3/titlepage.pdf.

Anonymous:TB37-3-254

[4200] Anonymous. TUGboat editorial information. TUGboat, 37(3):254, 2016. ISSN 0896-3207.

Anonymous:TB37-3-254-2

[4201] Anonymous. TUG institutional members. TUGboat, 37(3):254, 2016. ISSN 0896-3207.

Hefferon:TB37-3-255

[4202] Jim Hefferon. President’s note. TUGboat, 37(3):255, 2016. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb37-3/tb117pres.pdf.

Beeton:TB37-3-256

[4203] Barbara Beeton. Editorial comments. TUGboat, 37(3):256–258, 2016. ISSN

0896-3207. URL https://tug.org/TUGboat/tb37-3/tb117beet.
pdf.

Walden:TB37-3-259

[4204] David Walden. Interview with Federico Garcia-De Castro. TUGboat, 37(3): 259–263, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-3/tb117garcia.pdf.

Flynn:TB37-3-264

[4205] Peter Flynn. Typographers’ Inn. TUGboat, 37(3):264–266, 2016. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-3/tb117inn.pdf.

Hagen:TB37-3-267

[4206] Hans Hagen. LuaTEX version 1.0.0. TUGboat, 37(3):267–268, 2016. ISSN

0896-3207. URL https://tug.org/TUGboat/tb37-3/tb117hagen-
luatex.pdf.

Hagen:TB37-3-269

[4207] Hans Hagen. LuaTEX 0.82 OpenType math enhancements. TUGboat, 37(3): 269–274, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-3/tb117hagen-otmath.pdf.

Aye:TB37-3-275

[4208] Gareth Aye. Introducing LaTeX Base. TUGboat, 37(3):275–276, 2016. ISSN

0896-3207. URL https://tug.org/TUGboat/tb37-3/tb117aye.
pdf.

Ruckert:TB37-3-277

[4209] Martin Ruckert. Computer Modern Roman fonts for ebooks. TUGboat, 37(3): 277–280, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-3/tb117ruckert.pdf.

Willadt:TB37-3-281

[4210] Peter Willadt. When (image) size matters. TUGboat, 37(3):281–283,

2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-3/
tb117willadt.pdf.

Lemberg:TB37-3-284

[4211] Werner Lemberg. A survey of the history of musical notation. TUGboat, 37(3): 284–304, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-3/tb117lemberg.pdf.

Hagen:TB37-3-305

[4212] Hans Hagen. Colorful emojis via Unicode and OpenType. TUGboat, 37(3):

305, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-
3/tb117hagen-emoji.pdf.

Hoekwater:TB37-3-306

[4213] Taco Hoekwater and Hans Hagen. Cowfont (koeieletters) update. TUGboat, 37(3):306–310, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-3/tb117hoekwater.pdf.

Romer:TB37-3-311

[4214] Linus Romer. Corrections for slanted stems in  and  .

TUGboat, 37(3):311–316, 2016. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb37-3/tb117romer.pdf.

Jackowski:TB37-3-317

[4215] Bogus�aw Jackowski, Piotr Strzelczyk, and Piotr Pianowski. GUST e-foundry font projects. TUGboat, 37(3):317–336, 2016. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb37-3/tb117jackowski.pdf.

Talbot:TB37-3-337

[4216] Nicola Talbot. Localisation of TEX documents: tracklang. TUGboat, 37(3): 337–351, 2016. ISSN 0896-3207. URL https://tug.org/TUGboat/

tb37-3/tb117talbot.pdf.

Wilson:TB37-3-352

[4217] Peter Wilson. Glisterings: Index headers; numerations; real number comparison. TUGboat, 37(3):352–356, 2016. ISSN 0896-3207. URL https://

tug.org/TUGboat/tb37-3/tb117glister.pdf.

Walden:TB37-3-357

[4218] David Walden. Messing with endnotes. TUGboat, 37(3):357, 2016. ISSN

0896-3207. URL https://tug.org/TUGboat/tb37-3/tb117walden.
pdf.

Wermuth:TB37-3-358

[4219] Udo Wermuth. Tracing paragraphs. TUGboat, 37(3):358–373, 2016. ISSN

0896-3207. URL https://tug.org/TUGboat/tb37-3/tb117wermuth.
pdf.

Berry:TB37-3-374

[4220] Karl Berry. The treasure chest. TUGboat, 37(3):374–376, 2016. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-3/tb117chest.
pdf.

Atkinson:TB37-3-376

[4221] John Atkinson. An asterisk’s lament. TUGboat, 37(3):376, 2016. ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-3/tb117chest.
pdf.

Anonymous:TB37-3-377

[4222] Anonymous. Die TEXnische Komödie: Contents of issues 2–3/2016. TUGboat, 37(3):377, 2016. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb37-3/tb117komo.pdf.

Anonymous:TB37-3-378

[4223] Anonymous. TUG 2017 election. TUGboat, 37(3):378, 2016. ISSN 0896-3207.

URL https://tug.org/TUGboat/tb37-3/tb117elec.pdf.

Anonymous:TB37-3-379

[4224] Anonymous. TEX consulting and production services. TUGboat, 37(3):379,

2016. ISSN 0896-3207. URL https://tug.org/consultants.html.

Anonymous:TB37-3-380

[4225] Anonymous. Calendar. TUGboat, 37(3):380, 2016. ISSN 0896-3207. URL

https://tug.org/TUGboat/tb37-3/tb117calendar.pdf.

Anonymous:2016:CODc

[4226] Anonymous. Contents ordered by difficulty. TUGboat, 37(3):c3, 2016. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-
3/cover3.pdf.

Anonymous:2016:FCc

[4227] Anonymous. Front cover. TUGboat, 37(3):c1, 2016. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb37-3/cover1.pdf.

Anonymous:2016:IFCc

[4228] Anonymous. Inside front cover. TUGboat, 37(3):c2, 2016. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb37-3/
cover2.pdf.

Anonymous:2017:FCa

[4229] Anonymous. Front cover. TUGboat, 38(1):c1, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-1/cover1.pdf.

Anonymous:2017:IFCa

[4230] Anonymous. Inside front cover. TUGboat, 38(1):c2, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-1/
cover2.pdf.

Anonymous:2017:CODa

[4231] Anonymous. Contents ordered by difficulty. TUGboat, 38(1):c3, 2017. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-
1/cover3.pdf.

Anonymous:2017:TPa

[4232] Anonymous. Title page. TUGboat, 38(1):1, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-1/titlepage.pdf.

Anonymous:2017:EIa

[4233] Anonymous. Editorial information. TUGboat, 38(1):2, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-1/
tb118edit.pdf.

Anonymous:2017:IMa

[4234] Anonymous. Institutional members. TUGboat, 38(1):2, 2017. CODEN ????

ISSN 0896-3207. URL https://www.tug.org/instmem.html.

Hefferon:2017:PN

[4235] Jim Hefferon. President’s note. TUGboat, 38(1):3–4, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-1/
tb118pres.pdf.

Beeton:2017:ECa

[4236] Barbara Beeton. Editorial comments. TUGboat, 38(1):5–6, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-1/
tb118beet.pdf.

Walden:2017:ISP

[4237] David Walden. Interview with Scott Pakin. TUGboat, 38(1):7–10, 2017. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-
1/tb118pakin.pdf.

Teplow:2017:WPN

[4238] David Teplow. What’s a Professor of Neurology doing using LaTeX? TUGboat, 38(1):10–15, 2017. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb38-1/tb118teplow.pdf.

Flynn:2017:TI

[4239] Peter Flynn. Typographers’ Inn. TUGboat, 38(1):16–17, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-1/
tb118inn.pdf.

Bigelow:2017:RSHa

[4240] Charles Bigelow. Review and summaries: The history of typographic writing — the 20th century. TUGboat, 38(1):18–22, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-1/
tb118bigelow.pdf.

Cozens:2017:SNT

[4241] Simon Cozens. SILE: A new typesetting system. TUGboat, 38(1):23–27, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-1/tb118cozens.pdf.

Sharpe:2017:B

[4242] Michael Sharpe. BaskervilleF. TUGboat, 38(1):28–30, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-1/
tb118sharpe.pdf.

Dunn:2017:PLS

[4243] Brian Dunn. Programming L^{A}TEX — a survey of documentation and packages. TUGboat, 38(1):31–33, 2017. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb38-1/tb118dunn-progltx.pdf.

Neugebauer:2017:CGN

[4244] Gerd Neugebauer. CTAN goes 2.0 — new community features and more. TUGboat, 38(1):34–38, 2017. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb38-1/tb118neugebauer.pdf.

Thurnherr:2017:ILC

[4245] Thomas Thurnherr. An introduction to the L^{A}TEX cross-referencing system. TUGboat, 38(1):39–40, 2017. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb38-1/tb118thurnherr.pdf.

Salimi:2017:HUB

[4246] Behzad Salimi. How to use basic color models in L^{A}TEX. TUGboat, 38(1): 41–43, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-1/tb118salimi.pdf.

Snyder:2017:SCD

[4247] Hal Snyder. SageMathCloud for collaborative document editing and scientific computing. TUGboat, 38(1):44–47, 2017. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb38-1/tb118snyder.pdf.

Dunn:2017:PHD

[4248] Brian Dunn. Producing HTML directly from L^{A}TEX — the lwarp package. TUGboat, 38(1):48–53, 2017. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb38-1/tb118dunn-lwarp.pdf.

Team:2017:LNIa

[4249] L^{A}TEX Project Team. L^{A}TEX news, issue 26, January 2017. TUGboat, 38(1): 54–55, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-1/tb118ltnews26.pdf.

Team:2017:LNIb

[4250] L^{A}TEX Project Team. L^{A}TEX3 news, issue 10, November 2016. TUGboat, 38(1): 56–57, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-1/tb118l3news.pdf.

Dunn:2017:KVI

[4251] Brian Dunn. A key/value interface for generating L^{A}TEX floats — the keyfloat package. TUGboat, 38(1):58–60, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-1/tb118dunn-
keyfloat.pdf.

Wilson:2017:GHS

[4252] Peter Wilson. Glisterings: Hanging; safety in numbers. TUGboat, 38(1):61– 64, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-1/tb118glister.pdf.

Wermuth:2017:OVC

[4253] Udo Wermuth. The optimal value for \emergencystretch. TUGboat, 38 (1):65–86, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-1/tb118wermuth.pdf.

Berry:2017:TCa

[4254] Karl Berry. The treasure chest. TUGboat, 38(1):87–88, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-1/
tb118chest.pdf.

Hefferon:2017:BRM

[4255] Jim Hefferon. Book review: More Math Into L^{A}TEX, 5th edition, by George Grätzer. TUGboat, 38(1):89, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-1/tb118reviews-
gratzer.pdf.

Veytsman:2017:BRN

[4256] Boris Veytsman. Book review: The Noblest Roman: A History of the Centaur Types … by Jerry Kelly and Misha Beletsky. TUGboat, 38(1):90–91, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-1/tb118reviews-kelly-beletsky.pdf.

Walden:2017:BRT

[4257] David Walden. Book review: Track Changes, by Matthew G. Kirschenbaum. TUGboat, 38(1):92, 2017. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb38-1/tb118reviews-kirschenbaum.
pdf.

Veytsman:2017:BRM

[4258] Boris Veytsman. Book review: Manuale Calligraphicum, David Pankow, ed. TUGboat, 38(1):93, 2017. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb38-1/tb118reviews-pankow.pdf.

Veytsman:2017:BRS

[4259] Boris Veytsman. Book review: Seminar review: Presenting data and information by Edward Tufte. TUGboat, 38(1):94–95, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-1/
tb118reviews-tufte.pdf.

Anonymous:2017:BRTa

[4260] Anonymous. Book review: Die TeXnische Komödie 4/2016–1/2017. TUGboat, 38(1):96, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-1/tb118komo.pdf.

Anonymous:2017:BRZ

[4261] Anonymous. Book review: Zpravodaj: Contents of issues 2015/3–4–2016/1–4. TUGboat, 38(1):97, 2017. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb38-1/tb118zprav.pdf.

Anonymous:2017:ECI

[4262] Anonymous. Eutypon: Contents of issue 36–37 (October 2016). TUGboat, 38 (1):98, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-1/tb118eutypon.pdf.

Hoppner:2017:TFS

[4263] Klaus Höppner. TUG financial statements for 2016. TUGboat, 38(1):99, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-1/tb118treas.pdf.

Election-committee:2017:TE

[4264] TUG Elections committee. TUG 2017 election. TUGboat, 38(1):100–106, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-1/tb118elec.pdf.

Munroe:2017:FEC

[4265] Randall Munroe. File extensions (cartoon). TUGboat, 38(1):106, 2017. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-
1/tb118elec.pdf.

Anonymous:2017:TCPa

[4266] Anonymous. TEX consulting and production services. TUGboat, 38(1): 107, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:2017:Ca

[4267] Anonymous. Calendar. TUGboat, 38(1):108, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-1/tb118calendar.
pdf.

Anonymous:2017:FCb

[4268] Anonymous. Front cover. TUGboat, 38(2):c1, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-2/cover1.pdf.

Anonymous:2017:IFCb

[4269] Anonymous. Inside front cover. TUGboat, 38(2):c2, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
cover2.pdf.

Anonymous:2017:CODb

[4270] Anonymous. Contents ordered by difficulty. TUGboat, 38(2):c3, 2017. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-
2/tb119capsule.pdf.

Anonymous:2017:TPE

[4271] Anonymous. Title page: Errata: should be 38th annual conference and not 37th. TUGboat, 38(2):109, 2017. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb38-2/titlepage.pdf.

Organizers:2017:TCI

[4272] Organizers, Sponsors, Participants, and Program. TUG 2017 conference information. TUGboat, 38(2):110–111, 2017. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb38-2/tb119confinfo.pdf.

Anonymous:2017:TCPb

[4273] Anonymous. TUG 2017 conference program. TUGboat, 38(2):112–113, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-2/tb119program.pdf.

Hufflen:2017:T

[4274] Jean-Michel Hufflen. TUG@BachoTeX 2017. TUGboat, 38(2):114–115, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-2/tb119hufflen-tb17.pdf.

Nowacki:2017:BRC

[4275] Janusz Nowacki. Book review: Calligraphy by Barbara Galińska. TUGboat, 38(2):116–117, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-2/tb119nowacki.pdf.

Rychly:2017:RS

[4276] Maciej Rych�y. Released sounds. TUGboat, 38(2):118–124, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119rychly.pdf.

Veytsman:2017:ST

[4277] Boris Veytsman. The state of TEX. TUGboat, 38(2):125, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119pres.pdf.

Hagen:2017:CT

[4278] Hans Hagen. Children of TEX. TUGboat, 38(2):126–140, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119hagen-children.pdf.

Scarso:2017:MP

[4279] Luigi Scarso. MFLua 0.8 — prologue. TUGboat, 38(2):141–144, 2017. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-
2/tb119scarso.pdf.

Gasewicz:2017:TCM

[4280] Micha� Gasewicz. Off topic (completely): Many faces (and types) of beer. TUGboat, 38(2):145–146, 2017. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb38-2/tb119gasewicz.pdf.

Hufflen:2017:HAM

[4281] Jean-Michel Hufflen. History of accidentals in music. TUGboat, 38(2):147– 156, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-2/tb119hufflen-music.pdf.

Egger:2017:BWM

[4282] Willi Egger. Bookbinding workshop: Making a portfolio. TUGboat, 38(2): 157–158, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-2/tb119egger.pdf.

Beeton:2017:DLF

[4283] Barbara Beeton. Debugging L^{A}TEX files — Illegitimi non carborundum. TUGboat, 38(2):159–164, 2017. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb38-2/tb119beet.pdf.

Sathasivam:2017:RSU

[4284] Kumaran Sathasivam, S. K. Venkatesan, and Yakov Chandy. Revealing semantics using subtle typography and punctuation. TUGboat, 38(2):165–170, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-2/tb119sathasivam.pdf.

Veytsman:2017:JJW

[4285] Boris Veytsman and Leila Akhmadeeva. To justify or not to justify? Why bad typography may be harmful for your readers. TUGboat, 38(2):171–172, 2017.

CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-2/tb119veytsman-justify.pdf.

Veytsman:2017:MLJ

[4286] Boris Veytsman. Making ltxsparklines: The journey of a CTAN contributor into the world of CRAN. TUGboat, 38(2):173–174, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119veytsman-sparklines.pdf.

Sojka:2017:TSJ

[4287] Petr Sojka and Vít Novotný. TEX in schools? Just say yes: The use of TEX at the Faculty of Informatics, Masaryk Univ. TUGboat, 38(2):175–184, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-2/tb119sojka.pdf.

Asakura:2017:IBA

[4288] Takuto Asakura. Implementing bioinformatics algorithms in TEX — the Gotoh package, a case study. TUGboat, 38(2):185–187, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119asakura.pdf.

Preining:2017:UFP

[4289] Norbert Preining. updmap and fmtutil — past and future changes (or: cleaning up the mess). TUGboat, 38(2):188–192, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119preining.pdf.

Kroonenberg:2017:TTL

[4290] Siep Kroonenberg. TLaunch, the TEX Live Launcher for Windows. TUGboat, 38(2):193–196, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-2/tb119kroonenberg.pdf.

Tolusis:2017:XDR

[4291] Sigitas Tolušis, Arūnas Povilaitis, and Valentinas Kriauačiukas. Xdvipsk: Dvips ready for OpenType fonts and more image formats. TUGboat, 38(2): 197–201, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-2/tb119tolusis.pdf.

Ludwichowski:2017:GFC

[4292] Jerzy Ludwichowski. GUST e-foundry current font projects. TUGboat, 38(2): 202, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-2/tb119ludwichowski.pdf.

Hagen:2017:VF

[4293] Hans Hagen. Variable fonts. TUGboat, 38(2):203–207, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119hagen-variable.pdf.

Jackowski:2017:PMS

[4294] Bogus�aw Jackowski, Piotr Strzelczyk, and Piotr Pianowski. Parametric math symbol fonts. TUGboat, 38(2):208–211, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119jackowski.pdf.

Team:2017:LNIc

[4295] L^{A}TEX Project Team. L^{A}TEX news, issue 27, April 2017. TUGboat, 38(2): 212, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-2/tb119ltnews27.pdf.

Mittelbach:2017:LTC

[4296] Frank Mittelbach. L^{A}TEX table columns with fixed widths. TUGboat, 38(2): 213–214, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-2/tb119mitt.pdf.

Novotny:2017:UMI

[4297] Vít Novotný. Using Markdown inside TEX documents. TUGboat, 38(2):214– 217, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-2/tb119novotny.pdf.

Murzynowski:2017:GGM

[4298] Grzegorz Murzynowski. GMS, the “General Meta-Scenarios”: A proper extension to the l3expan package of the expl3 bundle and language, two years later. TUGboat, 38(2):218–237, 2017. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb38-2/tb119murzynowski.pdf.

Luptak:2017:TBC

[4299] Dávid Lupták. Typesetting bibliographies compliant with the ISO 690 standard in L^{A}TEX. TUGboat, 38(2):238–244, 2017. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb38-2/tb119luptak.pdf.

Hufflen:2017:MNH

[4300] Jean-Michel Hufflen. MlBibTEX now handles Unicode. TUGboat, 38(2):245– 248, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-2/tb119hufflen-mlbibtex.pdf.

Tolena:2017:TUH

[4301] Lolita Tolenȧ. TEX user habits versus publisher requirements. TUGboat, 38 (2):249–254, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-2/tb119tolene.pdf.

Borkowski:2017:TYW

[4302] Marcin Borkowski. Ten years of work in Wiadomości Matematyczne — an adventure with L^{A}TEX and Emacs hacking. TUGboat, 38(2):255–263, 2017. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-
2/tb119borkowski.pdf.

Berry:2017:PN

[4303] Karl Berry. Production notes. TUGboat, 38(2):263, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119prod.pdf.

Anonymous:2017:SRN

[4304] Anonymous. Streszczenia [Reports and notices — all abstracts, in Polish]. TUGboat, 38(2):264–269, 2017. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb38-2/tb119streszczenia.pdf.

NeydeSouza:2017:TA

[4305] Paulo Ney de Souza, Willi Egger, Hans Hagen, Taco Hoekwater, Zunbeltz Izaola, Anna Beata Kwiatkowska, Jerzy Ludwichowski, Mojca Miklavec, Frank Mittelbach, Arthur Reutenauer, Przemys�aw Scherwentke, Damien T. Thiriet, Andrzej Tomaszewski, Adam Twardoch, and Ulrik Vieth. TUG@BachoTeX 2017 abstracts. TUGboat, 38(2):270–272, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119abstracts.pdf.

Anonymous:2017:BRTb

[4306] Anonymous. Book review: Die TeXnische Komödie 2/2017. TUGboat, 38(2): 273, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-2/tb119komo.pdf.

Bigelow:2017:RSHb

[4307] Charles Bigelow. Review and summaries: The history of typographic writing — the 20th century, volume 2 (ch. 1–5). TUGboat, 38(2):274–279, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-2/tb119bigelow.pdf.

Veytsman:2017:BRW

[4308] Boris Veytsman. Book review: What Is Reading For? and Some Reflections on Reading and Writing, Culture and Nature, & Sorting Things Out by Robert Bringhurst. TUGboat, 38(2):280–281, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-2/tb119reviews-
bringhurst.pdf.

Walden:2017:BRP

[4309] David Walden. Book review: Paper: Paging Through History, by Mark Kurlansky. TUGboat, 38(2):282–283, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-2/tb119reviews-
kurlansky.pdf.

Berry:2017:TCb

[4310] Karl Berry. The treasure chest. TUGboat, 38(2):284–285, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119chest.pdf.

Atkinson:2017:WS

[4311] John Atkinson. Word on the street. TUGboat, 38(2):285, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-2/
tb119chest.pdf.

Anonymous:2017:Cb

[4312] Anonymous. Calendar. TUGboat, 38(2):286, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-2/tb119calendar.
pdf.

Anonymous:2017:TCPc

[4313] Anonymous. TEX consulting and production services. TUGboat, 38(2):287– 288, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:2017:IMb

[4314] Anonymous. Institutional members. TUGboat, 38(2):288, 2017. CODEN ????

ISSN 0896-3207. URL https://www.tug.org/instmem.html.

Anonymous:2017:FCc

[4315] Anonymous. Front cover. TUGboat, 38(3):c1, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-3/cover1.pdf.

Anonymous:2017:IFCc

[4316] Anonymous. Inside front cover. TUGboat, 38(3):c2, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-3/
cover2.pdf.

Anonymous:2017:CODc

[4317] Anonymous. Contents ordered by difficulty. TUGboat, 38(3):c3, 2017. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-
3/cover3.pdf.

Anonymous:2017:TPb

[4318] Anonymous. Title page. TUGboat, 38(3):289, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-3/titlepage.pdf.

Anonymous:2017:EIb

[4319] Anonymous. Editorial information. TUGboat, 38(3):290, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-3/
tb120edit.pdf.

Anonymous:2017:IMc

[4320] Anonymous. Institutional members. TUGboat, 38(3):290, 2017. CODEN ????

ISSN 0896-3207. URL https://www.tug.org/instmem.html.

Veytsman:2017:P

[4321] Boris Veytsman. From the President. TUGboat, 38(3):291, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-3/
tb120pres.pdf.

Beeton:2017:ECb

[4322] Barbara Beeton. Editorial comments. TUGboat, 38(3):291–292, 2017. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-
3/tb120beet.pdf.

Walden:2017:CMB

[4323] David Walden. Collecting memories of the beginnings of desktop publishing. TUGboat, 38(3):293, 2017. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb38-3/tb120walden-dtp.pdf.

Walden:2017:IMS

[4324] David Walden. Interview: Michael Sharpe. TUGboat, 38(3):294–298, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-3/tb120sharpe.pdf.

Hagen:2017:AT

[4325] Hans Hagen. Advertising TEX. TUGboat, 38(3):299–300, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-3/
tb120hagen-ad.pdf.

Maggi:2017:DNT

[4326] Carla Maggi. The DuckBoat — news from TeX.SE: Asking effective questions. TUGboat, 38(3):301–305, 2017. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb38-3/tb120duck.pdf.

Bigelow:2017:RSHc

[4327] Charles Bigelow. Review and summaries: The history of typographic writing — the 20th century, volume 2 (ch. 6–8+). TUGboat, 38(3):306–311, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-3/tb120bigelow.pdf.

Tsolomitis:2017:SGT

[4328] Antonis Tsolomitis. Serifed Greek type: Is it “Greek”? TUGboat, 38(3): 312–314, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-3/tb120tsolomitis.pdf.

Egger:2017:CB

[4329] Willi Egger. ConTEXt for beginners. TUGboat, 38(3):315–323, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-3/
tb120egger.pdf.

Herbst:2017:ACB

[4330] Marcel Herbst. Art Concret, basic design and meta-design. TUGboat, 38(3): 324–328, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-3/tb120herbst.pdf.

Voss:2017:CSP

[4331] Herbert Voß. The current state of the PSTricks project, part II. TUGboat, 38 (3):329–337, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-3/tb120voss.pdf.

Wilson:2017:GRL

[4332] Peter Wilson. Glisterings: Reading lines; paragraph endings; in conclusion. TUGboat, 38(3):338–341, 2017. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb38-3/tb120glister.pdf.

Izaola:2017:DMU

[4333] Zunbeltz Izaola and Paulo Ney de Souza. DocVar: Manage and use document variables. TUGboat, 38(3):342–344, 2017. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb38-3/tb120izaola.pdf.

Walden:2017:SMP

[4334] David Walden. Set my (pdf)pages free. TUGboat, 38(3):345, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-3/
tb120walden-pdfpages.pdf.

Thackurdeen:2017:AGH

[4335] R. Sean Thackurdeen and Boris Veytsman. Automatic generation of herbarium labels from spreadsheet data using L^{A}TEX. TUGboat, 38(3):345–349, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-3/tb120thackurdeen.pdf.

Beauchemin:2017:TAS

[4336] David Beauchemin and Vincent Goulet. Typesetting actuarial symbols easily and consistently with actuarialsymbol and actuarialangle. TUGboat, 38(3): 350–353, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-3/tb120beauchemin.pdf.

Ruckert:2017:CTW

[4337] Martin Ruckert. Converting TEX from WEB to cweb. TUGboat, 38(3):353– 358, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-3/tb120ruckert.pdf.

Gieseking:2017:DGS

[4338] Martin Gieseking. dvisvgm: Generating scalable vector graphics from DVI and EPS files. TUGboat, 38(3):359–368, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-3/tb120gieseking.
pdf.

Hagen:2017:TF

[4339] Hans Hagen. Tricky fences. TUGboat, 38(3):369–372, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-3/
tb120hagen-fences.pdf.

Talbot:2017:TIT

[4340] Nicola Talbot. Testing indexes: testidx.sty. TUGboat, 38(3):373–399, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-3/tb120talbot.pdf.

Wermuth:2017:NCL

[4341] Udo Wermuth. A note on \linepenalty. TUGboat, 38(3):400–414, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-3/tb120wermuth.pdf.

Berry:2017:TCc

[4342] Karl Berry. The treasure chest. TUGboat, 38(3):415–416, 2017. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-3/
tb120chest.pdf.

Atkinson:2017:CTE

[4343] John Atkinson. Comics: Typeface; elefonts. TUGboat, 38(3):416, 2017. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb38-
3/tb120chest.pdf.

Carlisle:2017:ASP

[4344] David Carlisle. Another seasonal puzzle: XII take II. TUGboat, 38(3): 416, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb38-3/tb120xii-lat.pdf.

Wilson:2017:BRS

[4345] Peter Wilson. Book review: Shady Characters and The Book by Keith Houston. TUGboat, 38(3):417–419, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-3/tb120reviews-
houston.pdf.

Anonymous:2017:TK

[4346] Anonymous. Die TeXnische Komödie 3/2017. TUGboat, 38(3):420, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-3/tb120komo.pdf.

Anonymous:2017:MCI

[4347] Anonymous. MAPS: Contents of issue 47 (2017). TUGboat, 38(3):420, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb38-3/tb120maps.pdf.

Anonymous:2017:TCPd

[4348] Anonymous. TEX consulting and production services. TUGboat, 38(3):421– 422, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:2017:CPP

[4349] Anonymous. Conference page: Practical TEX 2018 announcement. TUGboat, 38(3):423, 2017. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb38-3/tb120practex18.pdf.

Anonymous:2017:Cc

[4350] Anonymous. Calendar. TUGboat, 38(3):424, 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb38-3/tb120calendar.
pdf.

Anonymous:2018:FCa

Anonymous:2018:IFCa

[4352] Anonymous. Inside front cover. TUGboat, 39(1):c2, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
cover2.pdf.

Anonymous:2018:CODa

[4353] Anonymous. Contents ordered by difficulty. TUGboat, 39(1):c3, 2018. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-
1/cover3.pdf.

Anonymous:2018:TPa

[4354] Anonymous. Title page. TUGboat, 39(1):1, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-1/titlepage.pdf.

Anonymous:2018:EIa

[4355] Anonymous. Editorial information. TUGboat, 39(1):2, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121edit.pdf.

Anonymous:2018:IMa

[4356] Anonymous. Institutional members. TUGboat, 39(1):2, 2018. CODEN ????

ISSN 0896-3207. URL https://www.tug.org/instmem.html.

Veytsman:2018:Pa

[4357] Boris Veytsman. From the president. TUGboat, 39(1):3, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121pres.pdf.

Beeton:2018:ECa

[4358] Barbara Beeton. Editorial comments. TUGboat, 39(1):4–5, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121beet.pdf.

Beeton:2018:HELa

[4359] Barbara Beeton. Hyphenation exception log. TUGboat, 39(1):5–6, 2018. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-
1/tb121hyf.pdf.

Preining:2018:MSW

[4360] Norbert Preining. In memoriam: Staszek Wawrykiewicz (1953–2018). TUGboat, 39(1):7, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb39-1/tb121preining-staw.pdf.

Haralambous:2018:TPT

[4361] Yannis Haralambous. TEX as a path, a talk given at Donald Knuth’s 80th birthday celebration symposium. TUGboat, 39(1):8–15, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121haralambous-knuth80.pdf.

Fine:2018:TTU

[4362] Jonathan Fine. TUG is TEX users helping each other. TUGboat, 39(1):16, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-1/tb121fine-help.pdf.

Fine:2018:LJT

[4363] Jonathan Fine. L^{A}TEX and Jupyter, TikZ and Vega. TUGboat, 39(1):16, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-1/tb121fine-jupyter.pdf.

Flynn:2018:TIa

[4364] Peter Flynn. Typographers’ Inn. TUGboat, 39(1):17–18, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121inn.pdf.

Walden:2018:TDN

[4365] David Walden. Type designer Nina Stössinger speaks at 3rd Annual Updike Prize event. TUGboat, 39(1):19, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-1/tb121walden-
updike.pdf.

Neugebauer:2018:CQ

[4366] Gerd Neugebauer. CTAN quiz. TUGboat, 39(1):20, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121neugebauer-ctan.pdf.

Maggi:2018:DNTa

[4367] Carla Maggi. The DuckBoat — news from TeX.SE: The Morse code of TikZ. TUGboat, 39(1):21–26, 2018. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb39-1/tb121duck-tikz.pdf.

Atkinson:2018:CP

[4368] John Atkinson. Comic: Prefixation. TUGboat, 39(1):27, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121hagen-lua53.pdf.

Hagen:2018:L

[4369] Hans Hagen. From Lua 5.2 to 5.3. TUGboat, 39(1):27–29, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121hagen-lua53.pdf.

Borkowski:2018:TIE

[4370] Marcin Borkowski. TEXing in Emacs. TUGboat, 39(1):30–36, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121borkowski-emacs.pdf.

Menke:2018:TUE

[4371] Henri Menke. Tutorial: Using external C libraries with the LuaTEX FFI. TUGboat, 39(1):37–40, 2018. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb39-1/tb121menke-ffi.pdf.

Hagen:2018:ETL

[4372] Hans Hagen. Executing TEX in Lua: Coroutines. TUGboat, 39(1):41–43, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-1/tb121hagen-exec.pdf.

Mittelbach:2018:NRR

[4373] Frank Mittelbach and the L^{A}TEX Project Team. New rules for reporting bugs in the L^{A}TEX core software. TUGboat, 39(1):44–47, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121mitt-latexbug.pdf.

Team:2018:LNI

[4374] L^{A}TEX Project Team. L^{A}TEX news, issue 28, April 2018. TUGboat, 39(1):48– 50, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-1/tb121ltnews28.pdf.

Gregorio:2018:TSC

[4375] Enrico Gregorio. TeX.StackExchange cherry picking: expl3. TUGboat, 39(1): 51–59, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-1/tb121gregorio-expl3.pdf.

Battaia:2018:TDG

[4376] Luciano Battaia. Three-dimensional graphics with TikZ/PSTricks and the help of Geogebra. TUGboat, 39(1):60–68, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-1/tb121battaia-
3d.pdf.

Braslau:2018:CNC

[4377] Alan Braslau, Idris Hamid, and Hans Hagen. ConTEXt nodes: commutative diagrams and related graphics. TUGboat, 39(1):69–80, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121braslau-nodes.pdf.

Wermuth:2018:TAD

[4378] Udo Wermuth. TEX’s “additional demerits” parameters. TUGboat, 39(1):81– 87, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-1/tb121wermuth-adem.pdf.

Berry:2018:TCa

[4379] Karl Berry. The treasure chest. TUGboat, 39(1):88–89, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121chest.pdf.

Anonymous:2018:TKa

[4380] Anonymous. Die TeXnische Komödie 4/2017–1/2018. TUGboat, 39(1): 90, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-1/tb121komo.pdf.

Anonymous:2018:ZCI

[4381] Anonymous. Zpravodaj: Contents of issues 2017/1–4. TUGboat, 39(1): 91, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-1/tb121zprav.pdf.

Berry:2018:TFS

[4382] Karl Berry. TUG financial statements for 2017. TUGboat, 39(1):92, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-1/tb121treas.pdf.

Anonymous:2018:TCPa

[4383] Anonymous. TEX consulting and production services. TUGboat, 39(1):93– 94, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Berry:2018:PN

[4384] Karl Berry. Production notes. TUGboat, 39(1):94, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121prod.pdf.

Anonymous:2018:TA

[4385] Anonymous. TUG 2018 announcement. TUGboat, 39(1):95, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-1/
tb121tug18.pdf.

Anonymous:2018:Ca

[4386] Anonymous. Calendar. TUGboat, 39(1):96, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-1/tb121calendar.
pdf.

Anonymous:2018:FCb

[4387] Anonymous. Front cover. TUGboat, 39(2):c1, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-2/cover1.pdf.

Anonymous:2018:IFCb

[4388] Anonymous. Inside front cover. TUGboat, 39(2):c2, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-2/
cover2.pdf.

Anonymous:2018:CODb

[4389] Anonymous. Contents ordered by difficulty. TUGboat, 39(2):c3, 2018. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-
2/cover3.pdf.

Anonymous:2018:TPE

[4390] Anonymous. Title page: Errata: should be 39th annual conference and not 38th. TUGboat, 39(2):97, 2018. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb39-2/titlepage.pdf.

Organizers:2018:TCI

[4391] Organizers, Sponsors, Participants, and Program. TUG 2018 conference information. TUGboat, 39(2):98, 2018. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb39-2/tb122confinfo.pdf.

Wright:2018:TGR

[4392] Joseph Wright. TUG goes to Rio. TUGboat, 39(2):100–103, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-2/
tb122wright-tug18.pdf.

Wright:2018:TAM

[4393] Joseph Wright. TUG 2018 annual meeting notes. TUGboat, 39(2):104, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-2/tb122tug18-agm.pdf.

Raab:2018:TP

[4394] Susanne Raab. The tikzducks package. TUGboat, 39(2):105–106, 2018. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-
2/tb122raab-ducks.pdf.

Mittelbach:2018:RCP

[4395] Frank Mittelbach. A rollback concept for packages and classes. TUGboat, 39 (2):107–112, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb39-2/tb122mitt-rollback.pdf.

Robertson:2018:FLL

[4396] Will Robertson. Font loading in L^{A}TEX using the fontspec package: Recent updates. TUGboat, 39(2):113–116, 2018. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb39-2/tb122robertson-fontspec.
pdf.

Wright:2018:SCG

[4397] Joseph Wright. Supporting color and graphics in expl3. TUGboat, 39(2):117– 118, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-2/tb122wright-l3graphics.pdf.

Wright:2018:SPP

[4398] Joseph Wright. siunitx: Past, present and future. TUGboat, 39(2):119– 121, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-2/tb122wright-siunitx.pdf.

Cereda:2018:ATA

[4399] Paulo Cereda. Arara — TEX automation made easy. TUGboat, 39(2):122– 125, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-2/tb122cereda-arara.pdf.

Robertson:2018:CLM

[4400] Will Robertson. The Canvas learning management system and LaTeXacroML. TUGboat, 39(2):126–130, 2018. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb39-2/tb122robertson-canvas.
pdf.

Moore:2018:IPS

[4401] Ross Moore. Implementing PDF standards for mathematical publishing. TUGboat, 39(2):131–135, 2018. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb39-2/tb122moore-pdf.pdf.

Choi:2018:FMU

[4402] Jaeyoung Choi, Ammar Ul Hassan, and Geunho Jeong. FreeType MF Module: A module for using Metafont directly inside the FreeType rasterizer. TUGboat, 39(2):136–142, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb39-2/tb122choi-freetype.pdf.

Venkatesan:2018:WHC

[4403] S. K. Venkatesan. WeTEX and Hegelian contradictions in classical mathematics. TUGboat, 39(2):143–146, 2018. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb39-2/tb122venkatesan-
wetex.pdf.

Behrendt:2018:TA

[4404] Doris Behrendt, S. Coriasco, D. Ahmetovic, T. Armano, C. Bernareggi, M. Berra, A. Capietto, N. Murru, A. Ruighi, E. Taranto, Joachim Heinze, Tom Hejda, Mico Loretan, Frank Mittelbach, Ross Moore, Eduardo Ochs, Boris Veytsman, and Joseph Wright. TUG 2018 abstracts. TUGboat, 39(2):147– 149, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-2/tb122abstracts.pdf.

Anonymous:2018:MCI

[4405] Anonymous. MAPS: Contents of issue 48 (2018). TUGboat, 39(2):149, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-2/tb122maps.pdf.

Anonymous:2018:TKb

[4406] Anonymous. Die TeXnische Komödie 2–3/2018. TUGboat, 39(2):150, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-2/tb122komo.pdf.

Anonymous:2018:ECI

[4407] Anonymous. Eutypon: Contents of issue 38–39 (October 2017). TUGboat, 39 (2):151, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-2/tb122eutypon.pdf.

Childs:2018:DKA

[4408] Bart Childs and Rick Furuta. Don Knuth awarded Trotter Prize. TUGboat, 39 (2):151, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-2/tb122childs-trotter.pdf.

Beeton:2018:HELb

[4409] Barbara Beeton. Hyphenation exception log. TUGboat, 39(2):152, 2018. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-
2/tb122hyf.pdf.

Veytsman:2018:BRW

[4410] Boris Veytsman. Book review: W. A. Dwiggins: A Life in Design, by Bruce Kennett. TUGboat, 39(2):153–154, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-2/tb122reviews-
kennett.pdf.

Berry:2018:TCb

[4411] Karl Berry. The treasure chest. TUGboat, 39(2):155–156, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-2/
tb122chest.pdf.

Atkinson:2018:CHN

[4412] John Atkinson. Comics: Hyphe-nation; clumsy. TUGboat, 39(2):156, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-2/tb122chest.pdf.

Anonymous:2018:TCPb

[4413] Anonymous. TEX consulting and production services. TUGboat, 39(2):157– 158, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:2018:IMb

[4414] Anonymous. Institutional members. TUGboat, 39(2):158, 2018. CODEN ????

ISSN 0896-3207. URL https://www.tug.org/instmem.html.

Election-committee:2018:TEa

[4415] TUG Elections committee. TUG 2019 election. TUGboat, 39(2):159, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-2/tb122elec.pdf.

Anonymous:2018:Cb

[4416] Anonymous. Calendar. TUGboat, 39(2):160, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-2/tb122calendar.
pdf.

Anonymous:2018:FCc

[4417] Anonymous. Front cover. TUGboat, 39(3):c1, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-3/cover1.pdf.

Anonymous:2018:IFCc

[4418] Anonymous. Inside front cover. TUGboat, 39(3):c2, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-3/
cover2.pdf.

Anonymous:2018:CODc

[4419] Anonymous. Contents ordered by difficulty. TUGboat, 39(3):c3, 2018. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-
3/cover3.pdf.

Anonymous:2018:TPb

[4420] Anonymous. Title page. TUGboat, 39(3):161, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-3/titlepage.pdf.

Anonymous:2018:EIb

[4421] Anonymous. Editorial information. TUGboat, 39(3):162, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-3/
tb123edit.pdf.

Anonymous:2018:IMc

[4422] Anonymous. Institutional members. TUGboat, 39(3):162, 2018. CODEN ????

ISSN 0896-3207. URL https://www.tug.org/instmem.html.

Veytsman:2018:Pb

[4423] Boris Veytsman. From the president. TUGboat, 39(3):163, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-3/
tb123pres.pdf.

Beeton:2018:ECb

[4424] Barbara Beeton. Editorial comments. TUGboat, 39(3):164–166, 2018. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-
3/tb123beet.pdf.

Board:2018:TOA

[4425] TUG Board. TUGboat open-access survey results. TUGboat, 39(3):167, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-3/tb123board-tubsurvey.pdf.

Preining:2018:TCJ

[4426] Norbert Preining. TEXConf 2018 in Japan. TUGboat, 39(3):168, 2018. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-
3/tb123preining-jtexconf18.pdf.

Walden:2018:CGA

[4427] David Walden. The Cary Graphic Arts Collection. TUGboat, 39(3):169– 170, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-3/tb123walden-cary.pdf.

Flynn:2018:TIb

[4428] Peter Flynn. Typographers’ Inn. TUGboat, 39(3):171–172, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-3/
tb123inn.pdf.

Schulz:2018:BGF

[4429] Herbert Schulz and Richard Koch. A beginner’s guide to file encoding and TEXShop. TUGboat, 39(3):173–176, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-3/tb123schulz-
encoding.pdf.

Maggi:2018:DNTb

[4430] Carla Maggi. The DuckBoat — News from TeX.SE: Formatting posts. TUGboat, 39(3):177–181, 2018. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb39-3/tb123duck-format.pdf.

Johansson:2018:MPT

[4431] B. Tomas Johansson. Managing the paper trail of student projects: datatool and more. TUGboat, 39(3):182–184, 2018. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb39-3/tb123johansson-datatool.
pdf.

Walden:2018:IKH

[4432] David Walden. Interview with Kris Holmes. TUGboat, 39(3):185–203, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-3/tb123holmes-walden.pdf.

Bigelow:2018:SHB

[4433] Charles Bigelow and Kris Holmes. Science and history behind the design of Lucida. TUGboat, 39(3):204–211, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-3/tb123bigelow-
lucida.pdf.

Jackowski:2018:TGT

[4434] Bogus�aw Jackowski, Piotr Pianowski, and Piotr Strzelczyk. TEX Gyre text fonts revisited. TUGboat, 39(3):212–216, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-3/tb123jackowski-
gyre.pdf.

Ruckert:2018:HRT

[4435] Martin Ruckert. HINT: Reflowing TEX output. TUGboat, 39(3):217–223, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-3/tb123ruckert-hint.pdf.

Ahmetovic:2018:ACP

[4436] D. Ahmetovic, T. Armano, C. Bernareggi, M. Berra, A. Capietto, S. Coriasco, N. Murru, and A. Ruighi. Axessibility: Creating PDF documents with accessible formulae. TUGboat, 39(3):224–227, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-3/
tb123ahmetovic-axess.pdf.

Schubotz:2018:IRC

[4437] Moritz Schubotz, André Greiner-Petter, Philipp Scharpf, Norman Meuschke, Howard S. Cohl, and Bela Gipp. Improving the representation and conversion of mathematical formulae by considering the textual context. TUGboat, 39(3):

228–240, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-3/tb123schubotz-mathtext.pdf.

Ochs:2018:DES

[4438] Eduardo Ochs. Dednat6: An extensible (semi-)preprocessor for LuaL^{A}TEX that understands diagrams in ASCII art. TUGboat, 39(3):241–245, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-3/
tb123ochs-dednat.pdf.

Mittelbach:2018:MFP

[4439] Frank Mittelbach. Managing forlorn paragraph lines (a.k.a. widows and orphans) in L^{A}TEX. TUGboat, 39(3):246–251, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-3/tb123mitt-
widows.pdf.

Mittelbach:2018:WOP

[4440] Frank Mittelbach. The widows-and-orphans package. TUGboat, 39(3):252– 262, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-3/tb123mitt-widows-code.pdf.

Mittelbach:2018:DP

[4441] Frank Mittelbach. The dashundergaps package. TUGboat, 39(3):263–274, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-3/tb123mitt-dashundergaps.pdf.

Berry:2018:SSB

[4442] Karl Berry and Oren Patashnik. State secrets in bibliography-style hacking. TUGboat, 39(3):275, 2018. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb39-3/tb123berry-bst.pdf.

Wermuth:2018:ECP

[4443] Udo Wermuth. Experiments with \parfillskip. TUGboat, 39(3):276– 303, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb39-3/tb123wermuth-parfillskip.pdf.

Berry:2018:TCc

[4444] Karl Berry. The treasure chest. TUGboat, 39(3):304, 2018. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb39-3/
tb123chest.pdf.

Anonymous:2018:TKc

[4445] Anonymous. Die TeXnische Komödie 4/2018. TUGboat, 39(3):305, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-3/tb123komo.pdf.

Anonymous:2018:TCPc

[4446] Anonymous. TEX consulting and production services. TUGboat, 39(3):305– 306, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Election-committee:2018:TEb

[4447] TUG Elections committee. TUG 2019 election. TUGboat, 39(3):307, 2018. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb39-3/tb123elec.pdf.

Anonymous:2018:Cc

[4448] Anonymous. Calendar. TUGboat, 39(3):308, 2018. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb39-3/tb123calendar.
pdf.

Anonymous:2019:FCa

[4449] Anonymous. Front cover. TUGboat, 40(1):c1, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-1/cover1.pdf.

Anonymous:2019:IFCa

[4450] Anonymous. Inside front cover. TUGboat, 40(1):c2, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-1/
cover2.pdf.

Anonymous:2019:CODa

[4451] Anonymous. Contents ordered by difficulty. TUGboat, 40(1):c3, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
1/cover3.pdf.

Anonymous:2019:TPa

[4452] Anonymous. Title page. TUGboat, 40(1):1, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-1/titlepage.pdf.

Anonymous:2019:EIa

[4453] Anonymous. Editorial information. TUGboat, 40(1):2, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-1/
tb124edit.pdf.

Anonymous:2019:IMa

[4454] Anonymous. Institutional members. TUGboat, 40(1):2, 2019. CODEN ????

ISSN 0896-3207. URL https://www.tug.org/instmem.html.

Veytsman:2019:Pa

[4455] Boris Veytsman. From the President. TUGboat, 40(1):3, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-1/
tb124pres.pdf.

Beeton:2019:ECa

[4456] Barbara Beeton. Editorial comments. TUGboat, 40(1):4–5, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-1/
tb124beet.pdf.

Lang:2019:NNC

[4457] Sarah Lang and Astrid Schmölzer. Noob to Ninja: The challenge of taking beginners’ needs into account when teaching L^{A}TEX. TUGboat, 40(1):5–9, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb40-1/tb124lang-newbie.pdf.

Maggi:2019:DNT

[4458] Carla Maggi. The DuckBoat — news from TeX.SE: Processing text files to get L^{A}TEX tables. TUGboat, 40(1):10–13, 2019. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb40-1/tb124duck-csv.pdf.

Roberts:2019:NHD

[4459] Mike Roberts. No hands — the dictation of L^{A}TEX. TUGboat, 40(1):14– 16, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-1/tb124roberts-speak.pdf.

Jolly:2019:NBM

[4460] Susan Jolly. Nemeth Braille math and L^{A}TEX source as Braille. TUGboat, 40 (1):17–21, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-1/tb124jolly-braille.pdf.

Fowler:2019:BTD

[4461] Jim Fowler. Both TEX and DVI viewers inside the web browser. TUGboat, 40(1):22–24, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-1/tb124fowler-js.pdf.

Novotn:2019:MTL

[4462] Vít Novotný. Markdown 2.7.0: Towards lightweight markup in TEX. TUGboat, 40(1):25–27, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-1/tb124novotny-markdown.pdf.

Kroonenberg:2019:NFE

[4463] Siep Kroonenberg. New front ends for TEX Live. TUGboat, 40(1):28–29, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb40-1/tb124kroonenberg-gui.pdf.

Xie:2019:TLC

[4464] Yihui Xie. TinyTeX: A lightweight, cross-platform, and easy-to-maintain L^{A}TEX distribution based on TEX Live. TUGboat, 40(1):30–32, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-1/
tb124xie-tinytex.pdf.

Wright:2019:EPC

[4465] Joseph Wright. Extending primitive coverage across engines. TUGboat, 40 (1):33, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-1/tb124wright-primitives.pdf.

Hagen:2019:CL

[4466] Hans Hagen. ConTEXt LMTX. TUGboat, 40(1):34–37, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-1/
tb124hagen-lmtx.pdf.

Hosny:2019:BWS

[4467] Khaled Hosny. Bringing world scripts to LuaTEX: The HarfBuzz experiment. TUGboat, 40(1):38–43, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-1/tb124hosny-
harfbuzz.pdf.

Team:2019:LNIa

[4468] L^{A}TEX Project Team. L^{A}TEX news, issue 29, December 2018. TUGboat, 40(1): 44–46, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-1/tb124ltnews29.pdf.

Talbot:2019:GB

[4469] Nicola Talbot. Glossaries with bib2gls. TUGboat, 40(1):47–60, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
1/tb124talbot-bib2gls.pdf.

Gregorio:2019:TSC

[4470] Enrico Gregorio. TeX.StackExchange cherry picking, part 2: Templating. TUGboat, 40(1):61–68, 2019. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb40-1/tb124gregorio-templating.pdf.

Wright:2019:RNCa

[4471] Joseph Wright. Real number calculations in L^{A}TEX: Packages. TUGboat, 40(1): 69–70, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-1/tb124wright-fpkg.pdf.

Wright:2019:RNCb

[4472] Joseph Wright. Real number calculations in TEX: Implementations and performance. TUGboat, 40(1):71–75, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-1/tb124wright-
fperf.pdf.

Hoftich:2019:TLW

[4473] Michal Hoftich. TEX4ht: L^{A}TEX to Web publishing. TUGboat, 40(1):76– 81, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-1/tb124hoftich-make4ht.pdf.

Berry:2019:TOR

[4474] Karl Berry. TUGboat online, reimplemented. TUGboat, 40(1):82–84, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb40-1/tb124berry-tubonline.pdf.

Veytsman:2019:BRN

[4475] Boris Veytsman. Book review: Never use Futura by Douglas Thomas. TUGboat, 40(1):85–87, 2019. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb40-1/tb124reviews-thomas.pdf.

Anonymous:2019:TKa

[4476] Anonymous. Die TEXnische Komödie 1/2019. TUGboat, 40(1):88, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
1/tb124komo.pdf.

Anonymous:2019:ZCI

[4477] Anonymous. Zpravodaj: Contents of issues 2018/1–4. TUGboat, 40(1): 88, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-1/tb124zprav.pdf.

Berry:2019:TCa

[4478] Karl Berry. The treasure chest. TUGboat, 40(1):89–91, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-1/
tb124chest.pdf.

Atkinson:2019:CPH

[4479] John Atkinson. Comic: Punctuation headlines. TUGboat, 40(1):91, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb40-1/tb124chest.pdf.

Anonymous:2019:TCPa

[4480] Anonymous. TEX consulting and production services. TUGboat, 40(1):91– 92, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Berry:2019:TFS

[4481] Karl Berry. TUG financial statements for 2018. TUGboat, 40(1):93, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb40-1/tb124treas.pdf.

TEC:2019:TE

[4482] TUG Elections committee. TUG 2019 election. TUGboat, 40(1):94–95, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb40-1/tb124elec.pdf.

Anonymous:2019:Ca

[4483] Anonymous. Calendar. TUGboat, 40(1):96, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-1/tb124calendar.
pdf.

Anonymous:2019:FCb

[4484] Anonymous. Front cover. TUGboat, 40(2):c1, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-2/cover1.pdf.

Anonymous:2019:IFCb

[4485] Anonymous. Inside front cover. TUGboat, 40(2):c2, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-2/
cover2.pdf.

Anonymous:2019:CODb

[4486] Anonymous. Contents ordered by difficulty. TUGboat, 40(2):c3, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
2/cover3.pdf.

Anonymous:2019:TPb

[4487] Anonymous. Title page. TUGboat, 40(2):97, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-2/titlepage.pdf.

Anonymous:2019:TCI

[4488] Anonymous. TUG 2019 conference information. TUGboat, 40(2):98, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb40-2/tb125confinfo.pdf.

Anonymous:2019:TCPb

[4489] Anonymous. TUG 2019 conference program. TUGboat, 40(2):99, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
2/tb125program.pdf.

Anonymous:2019:TPc

[4490] Anonymous. TUG 2019 photos. TUGboat, 40(2):100, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-2/
tb125photos.pdf.

Menke:2019:BRT

[4491] Henri Menke. Back to the roots: TUG2019 in Palo Alto. TUGboat, 40(2): 101–104, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-2/tb125menke-tug19.pdf.

Claudio:2019:TAG

[4492] Jennifer Claudio. TUG 2019 Annual General Meeting notes. TUGboat, 40(2): 104–105, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-2/tb125tug19-agm.pdf.

Hefferon:2019:WDT

[4493] Jim Hefferon. What do today’s newcomers want? TUGboat, 40(2):106– 108, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-2/tb125heff-newusers.pdf.

Rokicki:2019:TFP

[4494] Tomas Rokicki. Type 3 fonts and PDF search in dvips. TUGboat, 40(2): 108–111, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-2/tb125rokicki-type3search.pdf.

Reutenauer:2019:SX

[4495] Arthur Reutenauer. The state of XETEX. TUGboat, 40(2):112, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-2/
tb125reutenauer-xetex.pdf.

Reutenauer:2019:HPL

[4496] Arthur Reutenauer. Hyphenation patterns: Licensing and stability. TUGboat, 40(2):113–114, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-2/tb125reutenauer-hyph.pdf.

Koch:2019:MNH

[4497] Richard Koch. MacTeX-2019, notification, and hardened runtimes. TUGboat, 40(2):115–118, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-2/tb125koch-harden.pdf.

Verna:2019:QCL

[4498] Didier Verna. Quickref: Common Lisp reference documentation as a stress test for Texinfo. TUGboat, 40(2):119–125, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-2/tb125verna-
quickref.pdf.

Ziegenhagen:2019:CLP

[4499] Uwe Ziegenhagen. Combining L^{A}TEX with Python. TUGboat, 40(2):126– 128, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-2/tb125ziegenhagen-python.pdf.

Menke:2019:PCD

[4500] Henri Menke. Parsing complex data formats in LuaTEX with LPEG. TUGboat, 40(2):129–135, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-2/tb125menke-lpeg.pdf.

Adams:2019:DSC

[4501] William Adams. Design into 3D: A system for customizable project designs. TUGboat, 40(2):136–142, 2019. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb40-2/tb125adams-3d.pdf.

Ruckert:2019:DHF

[4502] Martin Ruckert. The design of the HINT file format. TUGboat, 40(2):143– 146, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-2/tb125ruckert-hint.pdf.

T:2019:TFT

[4503] Rishikesan Nair T., C. V. Rajagopal, and C. V. Radhakrishnan. TEXFolio — a framework to typeset XML documents using TEX. TUGboat, 40(2):147– 149, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-2/tb125rishi-texfolio.pdf.

Rajendran:2019:NPF

[4504] Aravind Rajendran, Rishikesan Nair T., and C. V. Rajagopal. Neptune — a proofing framework for L^{A}TEX authors. TUGboat, 40(2):150–152, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
2/tb125rajendran-neptune.pdf.

Mittelbach:2019:LRW

[4505] Frank Mittelbach. The L^{A}TEX release workflow and the L^{A}TEX dev formats. TUGboat, 40(2):153–156, 2019. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb40-2/tb125mitt-dev.pdf.

Rowley:2019:ALK

[4506] Chris Rowley, Ulrike Fischer, and Frank Mittelbach. Accessibility in the L^{A}TEX kernel — experiments in Tagged PDF. TUGboat, 40(2):157–158, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
2/tb125rowley-tagpdf.pdf.

Veytsman:2019:CCE

[4507] Boris Veytsman. Creating commented editions with L^{A}TEX — the commedit package. TUGboat, 40(2):159–162, 2019. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb40-2/tb125veytsman-editions.
pdf.

Ziegenhagen:2019:CAE

[4508] Uwe Ziegenhagen. Creating and automating exams with L^{A}TEX & friends. TUGboat, 40(2):163–166, 2019. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb40-2/tb125ziegenhagen-
exam.pdf.

Ramesh:2019:BBD

[4509] Sree Harsha Ramesh, Dung Thai, Boris Veytsman, and Andrew McCallum. BibTEX-based dataset generation for training citation parsers. TUGboat, 40 (2):167–169, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-2/tb125ramesh-cite.pdf.

Choi:2019:FIM

[4510] Jaeyoung Choi, Saima Majeed, Ammar Ul Hassan, and Geunho Jeong. FreeType MF Module2: Integration of , GF, and PK inside FreeType. TUGboat, 40(2):170–178, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-2/tb125choi-
freetype2.pdf.

Parhami:2019:ECP

[4511] Behrooz Parhami. Evolutionary changes in Persian and Arabic scripts to accommodate the printing press, typewriting, and computerized word processing. TUGboat, 40(2):179–186, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-2/tb125parhami-
scripts.pdf.

Sojka:2019:UEP

[4512] Petr Sojka and Ondřej Sojka. The unreasonable effectiveness of pattern generation. TUGboat, 40(2):187–193, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-2/tb125sojka-
patgen.pdf.

Park:2019:IHE

[4513] Emily Park and Jennifer Claudio. Improving Hangul to English translation for optical character recognition (OCR). TUGboat, 40(2):194–195, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-2/
tb125park-hangul.pdf.

Bossard:2019:GCS

[4514] Antoine Bossard. A glance at CJK support with XETEX and LuaTEX. TUGboat, 40(2):196–201, 2019. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb40-2/tb125bossard-cjk.pdf.

Anane:2019:TA

[4515] Anane, Asakura, Braun, Claudio, Ha, Fuchs, Garcia-De Castro, Kannan, McKenna, Mittelbach, Moore, Shreevatsa, and Terada. TUG 2019 abstracts. TUGboat, 40(2):201–203, 2019. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb40-2/tb125abstracts.pdf.

Anonymous:2019:MCI

[4516] Anonymous. MAPS: Contents of issue 49 (2019). TUGboat, 40(2):204– 205, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-2/tb125maps.pdf.

Anonymous:2019:CGJ

[4517] Anonymous. ConTEXt Group Journal 2018. TUGboat, 40(2):205, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
2/tb125ctxg.pdf.

Anonymous:2019:TCPc

[4518] Anonymous. TEX consulting and production services. TUGboat, 40(2):206– 207, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:2019:IMb

[4519] Anonymous. Institutional members. TUGboat, 40(2):207, 2019. CODEN ???? ISSN 0896-3207.

Anonymous:2019:Cb

[4520] Anonymous. Calendar. TUGboat, 40(2):208, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-2/tb125calendar.
pdf.

Anonymous:2019:FCc

[4521] Anonymous. Front cover. TUGboat, 40(3):c1, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-3/cover1.pdf.

Anonymous:2019:IFCc

[4522] Anonymous. Inside front cover. TUGboat, 40(3):c2, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
cover2.pdf.

Anonymous:2019:CODc

[4523] Anonymous. Contents ordered by difficulty. TUGboat, 40(3):c3, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
3/cover3.pdf.

Anonymous:2019:TPd

[4524] Anonymous. Title page. TUGboat, 40(3):209, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-3/titlepage.pdf.

Anonymous:2019:EIb

[4525] Anonymous. Editorial information. TUGboat, 40(3):210, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126edit.pdf.

Anonymous:2019:IMc

[4526] Anonymous. Institutional members. TUGboat, 40(3):210, 2019. CODEN ???? ISSN 0896-3207.

Veytsman:2019:Pb

[4527] Boris Veytsman. From the President. TUGboat, 40(3):211, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126pres.pdf.

Beeton:2019:ECb

[4528] Barbara Beeton. Editorial comments. TUGboat, 40(3):212, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126beet.pdf.

Strakhov:2019:UBT

[4529] Yevhen Strakhov. Ukraine at BachoTEX 2019: Thoughts and impressions. TUGboat, 40(3):213–214, 2019. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb40-3/tb126strakhov-bacho19.
pdf.

Walden:2019:ETS

[4530] David Walden. An experience of trying to submit a paper in L^{A}TEX in an XMLfirst world. TUGboat, 40(3):215–216, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-3/tb126walden-
submit.pdf.

Walden:2019:SHC

[4531] David Walden. Studying the histories of computerizing publishing and desktop publishing, 2017–19. TUGboat, 40(3):217–228, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126walden-history.pdf.

Preining:2019:TST

[4532] Norbert Preining. TEX services at texlive.info. TUGboat, 40(3):229– 230, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-3/tb126preining-services.pdf.

IT:2019:PDI

[4533] Island of TEX. Providing Docker images for TEX Live and ConTEXt. TUGboat, 40(3):231, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-3/tb126island-docker.pdf.

Hagen:2019:TRP

[4534] Hans Hagen. TEX on the Raspberry Pi. TUGboat, 40(3):232–233, 2019. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-
3/tb126hagen-rpi.pdf.

Hoekwater:2019:MT

[4535] Taco Hoekwater. MuPDF tools. TUGboat, 40(3):234–235, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126hoekwater-mupdf.pdf.

vanOostrum:2019:LR

[4536] Piet van Oostrum. L^{A}TEX on the road. TUGboat, 40(3):236–246, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126oostrum-road.pdf.

Candia:2019:BPW

[4537] Estevão Vinícius Candia. A Brazilian Portuguese work on MetaPost, and how mathematics is embedded in it. TUGboat, 40(3):247–250, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126candia-mpost.pdf.

Team:2019:LNIb

[4538] L^{A}TEX Project Team. L^{A}TEX news, issue 30, October 2019. TUGboat, 40(3): 251–254, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb40-3/tb126ltnews30.pdf.

Asakura:2019:USD

[4539] Takuto Asakura. Understanding scientific documents with synthetic analysis on mathematical expressions and natural language. TUGboat, 40(3):255– 256, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-3/tb126asakura-meaning.pdf.

Hagen:2019:MTF

[4540] Hans Hagen. Modern Type 3 fonts. TUGboat, 40(3):257–262, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126hagen-type3.pdf.

Sajib:2019:TBS

[4541] Md Qutub Uddin Sajib. Typesetting the Bangla script in Unicode TEX engines — experiences and insights. TUGboat, 40(3):263–269, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126sajib-bangla.pdf.

Flynn:2019:TI

[4542] Peter Flynn. Typographers’ Inn. TUGboat, 40(3):270–271, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126inn.pdf.

Beeton:2019:BRH

[4543] Barbara Beeton. Book review: Hermann Zapf and the World He Designed: A Biography by Jerry Kelly. TUGboat, 40(3):272–273, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126reviews-kelly-zapfbio.pdf.

Berry:2019:BRC

[4544] Karl Berry. Book review: Carol Twombly: Her brief but brilliant career in type design by Nancy Stock-Allen. TUGboat, 40(3):274, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126reviews-stock-allen.pdf.

Anonymous:2019:TKb

[4545] Anonymous. Die TEXnische Komödie 2–3/2019. TUGboat, 40(3):275, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb40-3/tb126komo.pdf.

Anonymous:2019:ECI

[4546] Anonymous. Eutypon: Contents of issue 40–41 (October 2018). TUGboat, 40 (3):276, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb40-3/tb126eutypon.pdf.

Berry:2019:TCb

[4547] Karl Berry. The treasure chest. TUGboat, 40(3):277–278, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126chest.pdf.

Munroe:2019:CHU

[4548] Randall Munroe. Comic: The history of Unicode. TUGboat, 40(3):278, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb40-3/tb126chest.pdf.

TDF:2019:TDF

[4549] TEX Development Fund committee. TEX Development Fund 2014–2019 report. TUGboat, 40(3):279, 2019. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb40-3/tb126devfund.pdf.

Anonymous:2019:TSG

[4550] Anonymous. TUG 2019 sponsors: Google; Adobe; Overleaf; Pearson; STM Document Engineering Pvt Ltd. TUGboat, 40(3):280–282, 2019. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126sponsors.pdf.

Anonymous:2019:TCPd

[4551] Anonymous. TEX consulting and production services. TUGboat, 40(3):283– 284, 2019. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Anonymous:2019:Cc

[4552] Anonymous. Calendar. TUGboat, 40(3):284, 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb40-3/tb126calendar.
pdf.

Anonymous:2020:FCa

[4553] Anonymous. Front cover. TUGboat, 41(1):c1, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-1/cover1.pdf.

Anonymous:2020:IFCa

[4554] Anonymous. Inside front cover. TUGboat, 41(1):c2, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-1/
cover2.pdf.

Anonymous:2020:CODa

[4555] Anonymous. Contents ordered by difficulty. TUGboat, 41(1):c3, 2020. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-
1/cover3.pdf.

Anonymous:2020:TPa

[4556] Anonymous. Title page. TUGboat, 41(1):1, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-1/titlepage.pdf.

Anonymous:2020:EIa

[4557] Anonymous. Editorial information. TUGboat, 41(1):2, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-1/
tb127edit.pdf.

Anonymous:2020:IMa

[4558] Anonymous. Institutional members. TUGboat, 41(1):2, 2020. CODEN ???? ISSN 0896-3207.

Veytsman:2020:Pa

[4559] Boris Veytsman. From the President. TUGboat, 41(1):3, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-1/
tb127pres.pdf.

Beeton:2020:ECa

[4560] Barbara Beeton. Editorial comments. TUGboat, 41(1):4–6, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-1/
tb127beet.pdf.

Berry:2020:RBD

[4561] Karl Berry. Reporting bugs for Don Knuth (as soon as possible). TUGboat, 41(1):7, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-1/tb127berry-knuthbug.pdf.

Fuchs:2020:BTT

[4562] David Fuchs. Beyond Trip and Trap: Testing the urtext WEB sources. TUGboat, 41(1):8–11, 2020. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb41-1/tb127fuchs-knuthbug.pdf.

Haralambous:2020:GTJ

[4563] Yannis Haralambous. Grapholinguistics, TEX, and a June 2020 conference. TUGboat, 41(1):12–19, 2020. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb41-1/tb127haralambous-
grapholinguistics.pdf.

Maggi:2020:DBP

[4564] Carla Maggi. The DuckBoat — Beginners’ Pond: You do not need to be Neo to cope with a TikZ matrix. TUGboat, 41(1):20–25, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-1/
tb127duck-matrix.pdf.

Fischer:2020:CAP

[4565] Ulrike Fischer. Creating accessible pdfs with L^{A}TEX. TUGboat, 41(1):26– 28, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-1/tb127fischer-accessible.pdf.

Team:2020:LNIa

[4566] L^{A}TEX Project Team. L^{A}TEX3 news, issue 11, February 2018. TUGboat, 41(1): 29–30, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-1/tb127l3news11.pdf.

Team:2020:LNIb

[4567] L^{A}TEX Project Team. L^{A}TEX3 news, issue 12, January 2020. TUGboat, 41(1): 31–33, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-1/tb127l3news12.pdf.

Team:2020:LNIc

[4568] L^{A}TEX Project Team. L^{A}TEX news, issue 31, February 2020. TUGboat, 41(1): 34–38, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-1/tb127ltnews31.pdf.

Wright:2020:CCT

[4569] Joseph Wright. Case changing: From TEX primitives to the Unicode algorithm. TUGboat, 41(1):39–42, 2020. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb41-1/tb127wright-case.pdf.

Gregorio:2020:TLM

[4570] Enrico Gregorio. TEX, L^{A}TEX and math. TUGboat, 41(1):43–53, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-1/
tb127gregorio-math.pdf.

Mittelbach:2020:FP

[4571] Frank Mittelbach. The fewerfloatpages package. TUGboat, 41(1):54–68, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-1/tb127mitt-fewerfloat.pdf.

Fischer:2020:TBS

[4572] Ulrike Fischer and Marcel Krüger. Typesetting Bangla script with LuaL^{A}TEX. TUGboat, 41(1):69–70, 2020. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb41-1/tb127fischer-bangla.pdf.

Flynn:2020:TI

[4573] Peter Flynn. Typographers’ Inn. TUGboat, 41(1):71–72, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-1/
tb127inn.pdf.

Wermuth:2020:ARR

[4574] Udo Wermuth. An attempt at ragged-right typesetting. TUGboat, 41(1):73– 94, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-1/tb127wermuth-ragged.pdf.

Walden:2020:WPA

[4575] David Walden. Webnotes: Practical approaches. TUGboat, 41(1):95–96, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-1/tb127walden-webnotes.pdf.

Fine:2020:RCP

[4576] Jonathan Fine. Reading 29,000 COVID-19 papers. TUGboat, 41(1):97, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-1/tb127fine-covid19.pdf.

Walden:2020:AAC

[4577] David Walden. About The Art of Computer Programming, Volume 4, Fascicle 5. TUGboat, 41(1):98–102, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-1/tb127reviews-
knuth-fascicle5.pdf.

Walden:2020:BRH

[4578] David Walden. Book review: History of Desktop Publishing, by Frank Romano. TUGboat, 41(1):102–105, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-1/tb127reviews-
romano.pdf.

Anonymous:2020:ACI

[4579] Anonymous. ArsTEXnica: Contents of issues 27–28 (April and October 2019). TUGboat, 41(1):106–107, 2020. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb41-1/tb127arstexnica.pdf.

Anonymous:2020:ZCI

[4580] Anonymous. Zpravodaj: Contents of issue 2019/1–4. TUGboat, 41(1): 108, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-1/tb127zprav.pdf.

Anonymous:2020:TKa

[4581] Anonymous. Die TEXnische Komödie 4/2019. TUGboat, 41(1):109–110, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-1/tb127komo.pdf.

Berry:2020:TCa

[4582] Karl Berry. The treasure chest. TUGboat, 41(1):110–112, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-1/
tb127chest.pdf.

Atkinson:2020:CTM

[4583] John Atkinson. Comic: Typographic map Hyphe-nation. TUGboat, 41(1): 112, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-1/tb127chest.pdf.

Berry:2020:TFS

[4584] Karl Berry. TUG financial statements for 2019. TUGboat, 41(1):113, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-1/tb127treas.pdf.

Anonymous:2020:TCPa

[4585] Anonymous. TEX consulting and production services. TUGboat, 41(1):114– 115, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

consultants.html.

Berry:2020:TLN

[4586] Karl Berry. TEX Live 2020 news. TUGboat, 41(1):115, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-1/
tb127berry-tlnews.pdf.

Anonymous:2020:Ca

[4587] Anonymous. Calendar. TUGboat, 41(1):116, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-1/tb127calendar.
pdf.

Anonymous:2020:FCb

[4588] Anonymous. Front cover. TUGboat, 41(2):c1, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-2/cover1.pdf.

Anonymous:2020:IFCb

[4589] Anonymous. Inside front cover. TUGboat, 41(2):c2, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
cover2.pdf.

Anonymous:2020:CODb

[4590] Anonymous. Contents ordered by difficulty. TUGboat, 41(2):c3, 2020. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-
2/cover3.pdf.

Anonymous:2020:TPb

[4591] Anonymous. Title page. TUGboat, 41(2):117, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-2/titlepage.pdf.

Anonymous:2020:TCI

[4592] Anonymous. TUG 2020 conference information. TUGboat, 41(2):118– 119, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-2/tb128confinfo.pdf.

Beeton:2020:RMT

[4593] Barbara Beeton. Random musings on TUG2020 online. TUGboat, 41(2): 120, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-2/tb128beet-tug20.pdf.

Walden:2020:OTU

[4594] David Walden. Observations on the TEX Users Group’s 41st Annual Conference — TUG 2020 in the COVID-19 era. TUGboat, 41(2):121–122, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-2/tb128walden-tug20.pdf.

NeydeSouza:2020:TRF

[4595] Paulo Ney de Souza. TUG 2020: A report and future recommendations. TUGboat, 41(2):123–126, 2020. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb41-2/tb128desouza-tug20.pdf.

Fine:2020:TCG

[4596] Jonathan Fine. TEX conferences and general meetings, this year and next. TUGboat, 41(2):126–127, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-2/tb128fine-
meetings.pdf.

NeydeSouza:2020:IJB

[4597] Paulo Ney de Souza. Interview with Javier Bezos. TUGboat, 41(2):127– 131, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-2/tb128bezos-desouza.pdf.

NeydeSouza:2020:IPK

[4598] Paulo Ney de Souza. Interview with Philip Kime. TUGboat, 41(2):132– 138, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-2/tb128kime-desouza.pdf.

KH:2020:BRF

[4599] Hussain KH, Rajeesh KV, and Aravind Rajendran. Beyond Roman fonts: Extra dimensions in Malayalam fonts. TUGboat, 41(2):139–144, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
tb128hussain-beyond.pdf.

Matteson:2020:RN

[4600] Steven Matteson. The road to Noto. TUGboat, 41(2):145–154, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
tb128matteson-noto.pdf.

Claudio:2020:TET

[4601] Jennifer Claudio. Typographical explorations in two unicase alphabets. TUGboat, 41(2):155–156, 2020. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb41-2/tb128claudio-unicase.pdf.

Flynn:2020:YPL

[4602] Peter Flynn. Your personal L^{A}TEX bookshelf: Improving your background in a time of lockdown. TUGboat, 41(2):157–159, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-2/tb128flynn-
bookshelf.pdf.

Walden:2020:NHP

[4603] David Walden. Noticing history — a personal view. TUGboat, 41(2):160– 167, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-2/tb128walden-noticing.pdf.

Cereda:2020:TCT

[4604] Paulo Cereda. TEX in church: A typographical adventure. TUGboat, 41(2): 168–170, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb41-2/tb128cereda-church.pdf.

Schmolzer:2020:ETL

[4605] Astrid Schmölzer and Sarah Lang. Empowerment and teaching L^{A}TEX. TUGboat, 41(2):171–172, 2020. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb41-2/tb128schmoelzer-empowerment.
pdf.

Lang:2020:DRV

[4606] Sarah Lang. Didactical reduction versus references: How to better teach L^{A}TEX. TUGboat, 41(2):173–175, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-2/tb128lang-
didactic.pdf.

Tournade:2020:LHL

[4607] Yoan Tournade. L^{A}TEX-on-HTTP: L^{A}TEX as a commodity web service for application developers. TUGboat, 41(2):175–178, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
tb128tournade-http.pdf.

Veytsman:2020:UOC

[4608] Boris Veytsman. Using Overleaf for collaborative projects: First impressions and lessons learned. TUGboat, 41(2):179–181, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
tb128veytsman-overleaf.pdf.

IoT:2020:ITD

[4609] Island of TEX. The Island of TEX: Developing abroad, your next destination. TUGboat, 41(2):182–184, 2020. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb41-2/tb128island.pdf.

Asakura:2020:DCL

[4610] Takuto Asakura. The design concept for llmk — light L^{A}TEX make. TUGboat, 41(2):185–187, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb41-2/tb128asakura-llmk.pdf.

Gundlach:2020:TPC

[4611] Patrick Gundlach. Typesetting product catalogs and other database-driven documents with the speedata Publisher. TUGboat, 41(2):188–193, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
tb128gundlach-speedata.pdf.

Hefferon:2020:FSL

[4612] Jim Hefferon. A first set of L^{A}TEX packages. TUGboat, 41(2):194–195, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-2/tb128heff-packages.pdf.

DeMeritt:2020:POL

[4613] Susan DeMeritt and Cheryl Ponchin. Presenting our L^{A}TEX workshop online. TUGboat, 41(2):196, 2020. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb41-2/tb128demeritt-workshop.pdf.

Carlisle:2020:LOT

[4614] David Carlisle, Paulo Roberto Massa Cereda, and Joseph Wright. learnlatex.org: Taking L^{A}TEX training fully interactive. TUGboat, 41(2):197–199, 2020. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-
2/tb128carlisle-learnlatex.pdf.

Claudio:2020:RLO

[4615] Jennifer Claudio. A review of learnlatex.org. TUGboat, 41(2):199– 200, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-2/tb128reviews-learnlatex.pdf.

Mittelbach:2020:QVL

[4616] Frank Mittelbach and the L^{A}TEX Project Team. Quo vadis L^{A}TEX(3) Team — a look back and at the upcoming years. TUGboat, 41(2):201–207, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
tb128mitt-quovadis.pdf.

Ruckert:2020:HPS

[4617] Martin Ruckert and Gudrun Socher. The HINT Project: Status and open questions. TUGboat, 41(2):208–211, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-2/tb128ruckert-
hint.pdf.

Carlson:2020:MSL

[4618] James Carlson. MiniLaTeX: A subset of L^{A}TEX for the Web. TUGboat, 41(2): 212–214, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-2/tb128carlson-minilatex.pdf.

Hammond:2020:WLC

[4619] William Hammond. Why the L^{A}TEX community should care about SGML. TUGboat, 41(2):215–218, 2020. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb41-2/tb128hammond-sgml.pdf.

T:2020:LTW

[4620] Rishikesan Nair T., Aravind Rajendran, Rajagopal C. V., and Radhakrishnan C. V. L^{A}TEX technologies at work — aesthetically beautiful PDFs on the fly from XML input: XML Page Composition (XPC) micro-service in the cloud. TUGboat, 41(2):219–222, 2020. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb41-2/tb128rishi-xmlfly.pdf.

Moore:2020:TLP

[4621] Ross Moore. Tagging with L^{A}TEX — Part 1: Author considerations. TUGboat, 41(2):223–242, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb41-2/tb128moore-tagging1.pdf.

Fine:2020:TA

[4622] Jonathan Fine, Amelia Hugill-Fontanel, Paul Gessler, Patrick Ion, Marcel Krüger, John MacFarlane, Ross Moore, Eric Mc Sween, Norbert Preining, Thomas Price, Brandon Rhodes, samcarter, Michael Sharpe, Paulo Ney de Souza, and Vadim Ponomarev. TUG 2020 abstracts. TUGboat, 41(2):243– 245, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-2/tb128abstracts.pdf.

Anonymous:2020:TKb

[4623] Anonymous. Die TeXnische Komödie 2–3/2020. TUGboat, 41(2):246, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-2/tb128komo.pdf.

Anonymous:2020:Z

[4624] Anonymous. Zpravodaj 2020/1–2. TUGboat, 41(2):247–248, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
tb128zprav.pdf.

Berry:2020:TCb

[4625] Karl Berry. The treasure chest. TUGboat, 41(2):249–250, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
tb128chest.pdf.

Anonymous:2020:Cb

[4626] Anonymous. Calendar. TUGboat, 41(2):251, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-2/tb128calendar.
pdf.

TEC:2020:TEa

[4627] TUG Elections committee. TUG 2021 election. TUGboat, 41(2):252, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-2/tb128elec.pdf.

Anonymous:2020:IMb

[4628] Anonymous. Institutional members. TUGboat, 41(2):253, 2020. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Anonymous:2020:TS

[4629] Anonymous. TUG 2020 sponsors. TUGboat, 41(2):253–254, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-2/
tb128sponsors.pdf.

Anonymous:2020:TCPb

[4630] Anonymous. TEX consulting and production services. TUGboat, 41(2):255– 256, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2020:FCc

[4631] Anonymous. Front cover. TUGboat, 41(3):c1, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-3/cover1.pdf.

Anonymous:2020:IFCc

[4632] Anonymous. Inside front cover. TUGboat, 41(3):c2, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-3/
cover2.pdf.

Anonymous:2020:COD

[4633] Anonymous. Contents ordered by difficulty. TUGboat, 41(3):c3, 2020. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-
3/cover3.pdf.

Anonymous:2020:TP

[4634] Anonymous. Title page. TUGboat, 41(3):257, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-3/titlepage.pdf.

Anonymous:2020:EI

[4635] Anonymous. Editorial information. TUGboat, 41(3):258, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-3/
tb129edit.pdf.

Anonymous:2020:IMc

[4636] Anonymous. Institutional members. TUGboat, 41(3):258, 2020. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Veytsman:2020:PD

[4637] Boris Veytsman. From the president. TUGboat, 41(3):259, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-3/
tb129pres.pdf.

Beeton:2020:ECD

[4638] Barbara Beeton. Editorial comments. TUGboat, 41(3):260–262, 2020. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-
3/tb129beet.pdf.

Barr:2020:HTC

[4639] Michael Barr. How TEX changed my life. TUGboat, 41(3):263–264, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-3/tb129barr-changed.pdf.

Flynn:2020:TID

[4640] Peter Flynn. Typographers’ Inn. TUGboat, 41(3):265–268, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-3/
tb129inn.pdf.

Frear:2020:ECF

[4641] Lorrie Frear. Eye charts in focus: The magic of optotypes. TUGboat, 41(3): 269–274, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-3/tb129frear-optotype.pdf.

Mansour:2020:NLS

[4642] Kamal Mansour. The non-Latin scripts & typography. TUGboat, 41(3):275– 280, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-3/tb129mansour-nonlatin.pdf.

Leingang:2020:UDM

[4643] Matthew Leingang. Using DocStrip for multiple document variants. TUGboat, 41(3):281–285, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb41-3/tb129leingang-docstrip.pdf.

Team:2020:LNI

[4644] L^{A}TEX Project Team. L^{A}TEX news, issue 32, October 2020. TUGboat, 41(3): 286–291, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb41-3/tb129ltnews32.pdf.

Mittelbach:2020:LTP

[4645] Frank Mittelbach and Chris Rowley. L^{A}TEX Tagged PDF — a blueprint for a large project. TUGboat, 41(3):292–298, 2020. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb41-3/tb129mitt-tagpdf.
pdf.

Gregorio:2020:FE

[4646] Enrico Gregorio. Functions and expl3. TUGboat, 41(3):299–307, 2020. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-
3/tb129gregorio-l3functions.pdf.

Talbot:2020:BSC

[4647] Nicola Talbot. bib2gls: selection, cross-references and locations. TUGboat, 41 (3):308–317, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb41-3/tb129talbot-bib2gls-more.pdf.

Novotny:2020:MMM

[4648] Vít Novotný. Making Markdown into a microwave meal. TUGboat, 41(3): 318–320, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb41-3/tb129novotny-frozen.pdf.

Hagen:2020:UDT

[4649] Hans Hagen. User-defined Type 3 fonts in LuaTEX. TUGboat, 41(3):320– 323, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-3/tb129hagen-type3.pdf.

Wilson:2020:DDP

[4650] Peter Wilson. Data display, plots and graphs. TUGboat, 41(3):324–326, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-3/tb129wilson-plotgraph.pdf.

Scarso:2020:SRS

[4651] Luigi Scarso. Short report on the state of LuaTEX, 2020. TUGboat, 41(3): 327–328, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb41-3/tb129scarso-luatex.pdf.

Kitagawa:2020:DBC

[4652] Hironori Kitagawa. Distinguishing 8-bit characters and Japanese characters in (u)pTEX. TUGboat, 41(3):329–334, 2020. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb41-3/tb129kitagawa-char.
pdf.

Hagen:2020:KSD

[4653] Hans Hagen. Keyword scanning. TUGboat, 41(3):335–337, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-3/
tb129hagen-keyword.pdf.

Hagen:2020:RMP

[4654] Hans Hagen. Representation of macro parameters. TUGboat, 41(3):337– 340, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-3/tb129hagen-param.pdf.

Island:2020:TOW

[4655] Island of TEX. TEXdoc online — a web interface for serving TEX documentation. TUGboat, 41(3):341–342, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-3/tb129island-
texdoc.pdf.

Vlasak:2020:MCM

[4656] Michal Vlasák. MMTEX: Creating a minimal and modern TEX distribution for GNU/Linux. TUGboat, 41(3):343–345, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-3/tb129vlasak-
mmtex.pdf.

Liferenko:2020:UIT

[4657] Igor Liferenko. UTF-8 installations of TEX. TUGboat, 41(3):346–347, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-3/tb129liferenko-utf8.pdf.

Olsak:2020:ONG

[4658] Petr Olšák. OpTEX — a new generation of Plain TEX. TUGboat, 41(3):348– 354, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-3/tb129olsak-optex.pdf.

Bigelow:2020:BRR

[4659] Charles Bigelow. Book reviews: Robert Granjon, letter-cutter, and Granjon’s Flowers, by Hendrik D.L. Vervliet. TUGboat, 41(3):355–357, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-3/
tb129reviews-vervliet.pdf.

Veytsman:2020:BRG

[4660] Boris Veytsman. Book review: Glisterings, by Peter Wilson. TUGboat, 41(3): 358–359, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb41-3/tb129reviews-wilson.pdf.

Flynn:2020:HRT

[4661] Peter Flynn. Historical review of TEX3. TUGboat, 41(3):360–367, 2020. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-
3/tb129flynn-tex3.pdf.

Berry:2020:TC

[4662] Karl Berry. The treasure chest. TUGboat, 41(3):368, 2020. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb41-3/
tb129chest.pdf.

TEC:2020:TEb

[4663] TUG Elections committee. TUG 2021 election. TUGboat, 41(3):369, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb41-3/tb129elec.pdf.

Anonymous:2020:TCP

[4664] Anonymous. TEX consulting and production services. TUGboat, 41(3):370– 371, 2020. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2020:Cc

[4665] Anonymous. Calendar. TUGboat, 41(3):372, 2020. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb41-3/tb129calendar.
pdf.

Anonymous:2021:FCa

[4666] Anonymous. Front cover. TUGboat, 42(1):c1, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-1/cover1.pdf.

Anonymous:2021:IFCa

[4667] Anonymous. Inside front cover. TUGboat, 42(1):c2, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
cover2.pdf.

Anonymous:2021:CODa

[4668] Anonymous. Contents ordered by difficulty. TUGboat, 42(1):c3, 2021. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-
1/cover3.pdf.

Anonymous:2021:TPa

[4669] Anonymous. Title page. TUGboat, 42(1):1, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-1/titlepage.pdf.

Anonymous:2021:EIa

[4670] Anonymous. Editorial information. TUGboat, 42(1):2, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130edit.pdf.

Anonymous:2021:IMa

[4671] Anonymous. Institutional members. TUGboat, 42(1):2, 2021. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Veytsman:2021:Pa

[4672] Boris Veytsman. From the president. TUGboat, 42(1):3, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130pres.pdf.

Beeton:2021:ECa

[4673] Barbara Beeton. Editorial comments. TUGboat, 42(1):4–5, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130beet.pdf.

Andre:2021:HV

[4674] Jacques André. Hendrik Vervliet: 1923–2020. TUGboat, 42(1):6, 2021. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-
1/tb130andre-vervliet.pdf.

Knuth:2021:TT

[4675] Donald Knuth. The TEX tuneup of 2021. TUGboat, 42(1):7–10, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130knuth-tuneup21.pdf.

Berry:2021:TE

[4676] Karl Berry. TEX entomology in 2021. TUGboat, 42(1):10, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130berry-filter.pdf.

Beeton:2021:HEL

[4677] Barbara Beeton. Hyphenation exception log. TUGboat, 42(1):11, 2021. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-
1/tb130hyf.pdf.

Just:2021:YG

[4678] Jérémy Just. Year 2020 at GUTenberg. TUGboat, 42(1):12–13, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130just-gut.pdf.

Beeton:2021:LTL

[4679] Barbara Beeton. Lapses in TEX — a look backward. TUGboat, 42(1):13– 17, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-1/tb130beeton-lapses.pdf.

Flynn:2021:TI

[4680] Peter Flynn. Typographers’ Inn. TUGboat, 42(1):18–19, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130inn.pdf.

Walden:2021:IAH

[4681] David Walden. Interview with Amelia Hugill-Fontanel. TUGboat, 42(1):20– 33, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-1/tb130hugill-fontanel-walden.pdf.

Maggi:2021:DBP

[4682] Carla Maggi. The DuckBoat — beginners’ pond: Crazy little thing called glue. TUGboat, 42(1):34–39, 2021. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb42-1/tb130duck-glue.pdf.

Wright:2021:CDC

[4683] Joseph Wright. Creating document commands: The good, the bad and the ugly. TUGboat, 42(1):40–41, 2021. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb42-1/tb130wright-goodbad.pdf.

Wright:2021:CNV

[4684] Joseph Wright. \NewDocumentCommand versus \newcommand ver-

sus …. TUGboat, 42(1):41–43, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-1/tb130wright-
newdoccmd.pdf.

Olsak:2021:COO

[4685] Petr Olšák. Comparison of OpTEX with other formats: L^{A}TEX and ConTEXt. TUGboat, 42(1):44–49, 2021. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb42-1/tb130olsak-fmtcmp.pdf.

Ludwichowski:2021:GFF

[4686] Jerzy Ludwichowski. GUST e-foundry font projects, closing report 2019– 2020. TUGboat, 42(1):50–51, 2021. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb42-1/tb130ludwichowski-gust.
pdf.

Tsolomitis:2021:NFF

[4687] Antonis Tsolomitis. The NewComputerModern font family. TUGboat, 42(1): 52–55, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-1/tb130tsolomitis-newcm.pdf.

Venkatesan:2021:ACF

[4688] S. K. Venkatesan. An attempt at creating font transitions. TUGboat, 42(1): 56–59, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-1/tb130venkatesan-transfont.pdf.

Hagen:2021:SFG

[4689] Hans Hagen. Scaled fonts and glyphs. TUGboat, 42(1):60–65, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130hagen-scaled.pdf.

Berry:2021:SFR

[4690] Karl Berry. Some fonts with recent TEX support. TUGboat, 42(1):65, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb42-1/tb130berry-fontlist.pdf.

Yajie:2021:IGM

[4691] Hu Yajie. Improvements to the generalized mediation macros in The Metafontbook. TUGboat, 42(1):66, 2021. CODEN ???? ISSN

0896-3207. URL https://tug.org/TUGboat/tb42-1/tb130hu-
bernstein.pdf.

Chupin:2021:AFS

[4692] Maxime Chupin. Animating Fourier series decomposition of a character with LuaTEX and MPLIB. TUGboat, 42(1):67–71, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-1/tb130chupin-
fourier.pdf.

IT:2021:WRI

[4693] Island of TEX. Working remotely from an island: arara and other tools. TUGboat, 42(1):72–73, 2021. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb42-1/tb130island-tools.pdf.

Hagen:2021:LPF

[4694] Hans Hagen. LuaMetaTEX programming features. TUGboat, 42(1):74– 80, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-1/tb130hagen-prog.pdf.

Liferenko:2021:UIC

[4695] Igor Liferenko. UTF-8 installations of CWEB. TUGboat, 42(1):81–82, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb42-1/tb130liferenko-cweb.pdf.

Veytsman:2021:BRL

[4696] Boris Veytsman. Book review: Learning L^{A}TEX, Second Edition, by David F. Griffiths and Desmond J. Higham. TUGboat, 42(1):83, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130reviews-learnltx2.pdf.

Anonymous:2021:TKa

[4697] Anonymous. Die TeXnische Komödie 4/2020–1/2021. TUGboat, 42(1):84– 85, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-1/tb130komo.pdf.

Anonymous:2021:ZCIa

[4698] Anonymous. Zpravodaj: Contents of issue 2020/3–4. TUGboat, 42(1): 85, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-1/tb130zprav.pdf.

Atkinson:2021:CFD

[4699] John Atkinson. Comic: font or dog breed? TUGboat, 42(1):85, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130zprav.pdf.

Berry:2021:TCa

[4700] Karl Berry. The treasure chest. TUGboat, 42(1):86–88, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130chest.pdf.

committee:2021:TER

[4701] TUG Elections committee. TUG 2021 election report. TUGboat, 42(1):88– 91, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-1/tb130elec.pdf.

Koch:2021:SMF

[4702] Richard Koch. Spending MacTEX funds. TUGboat, 42(1):92, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-1/
tb130koch-mactex.pdf.

Berry:2021:TFS

[4703] Karl Berry. TUG financial statements for 2020. TUGboat, 42(1):93, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb42-1/tb130treas.pdf.

Anonymous:2021:TCPa

[4704] Anonymous. TEX consulting and production services. TUGboat, 42(1):94– 95, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2021:Ca

[4705] Anonymous. Calendar. TUGboat, 42(1):96, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-1/tb130calendar.
pdf.

Anonymous:2021:FCb

[4706] Anonymous. Front cover. TUGboat, 42(2):c1, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-2/cover1.pdf.

Anonymous:2021:IFCb

[4707] Anonymous. Inside front cover. TUGboat, 42(2):c2, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
cover2.pdf.

Anonymous:2021:CODb

[4708] Anonymous. Contents ordered by difficulty. TUGboat, 42(2):c3, 2021. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-
2/cover3.pdf.

Anonymous:2021:TPb

[4709] Anonymous. Title page. TUGboat, 42(2):97, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-2/titlepage.pdf.

Anonymous:2021:TCI

[4710] Anonymous. TUG 2021 conference information and program. TUGboat, 42 (2):98–100, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb42-2/tb131confinfo.pdf.

Rosendahl:2021:TCR

[4711] Arthur Rosendahl and Jennifer Claudio. TUG 2021 conference report. TUGboat, 42(2):100, 2021. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb42-2/tb131rosendahl-tug21.pdf.

Claudio:2021:TAG

[4712] Jennifer Claudio. TUG 2021 Annual General Meeting notes. TUGboat, 42(2): 101–104, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-2/tb131tug21-agm.pdf.

NeydeSouza:2021:TCO

[4713] Paulo Ney de Souza and Jennifer Claudio. TUG 2021 conference organization reflections and recommendations. TUGboat, 42(2):105–107, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131desouza-tug21.pdf.

NeydeSouza:2021:INP

[4714] Paulo Ney de Souza. Interview with Norbert Preining, mainstay of TEX Live and much more. TUGboat, 42(2):108–113, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131preining-desouza.pdf.

NeydeSouza:2021:IFM

[4715] Paulo Ney de Souza. Interview with Frank Mittelbach, leader of the L^{A}TEX Project. TUGboat, 42(2):114–120, 2021. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb42-2/tb131mittelbach-
desouza.pdf.

NeydeSouza:2021:IJH

[4716] Paulo Ney de Souza. Interview with John Hammersley, CEO of Overleaf. TUGboat, 42(2):121–130, 2021. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb42-2/tb131hammersley-
desouza.pdf.

samcarter:2021:PTP

[4717] samcarter. The tikzlings package. TUGboat, 42(2):131–132, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131samcarter-tikzlings.pdf.

Stamerjohanns:2021:PTB

[4718] Heinrich Stamerjohanns. Texmlbus, a build system to convert documents to XML and other formats. TUGboat, 42(2):132–134, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131stamerjohanns-texmlbus.pdf.

Goswami:2021:CIT

[4719] Rohit Goswami. Continuous integration and TEX with Org-Mode. TUGboat, 42(2):135–138, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb42-2/tb131goswami-ci-org.pdf.

Ruckert:2021:WCC

[4720] Martin Ruckert. The WEB to CWEB conversion of TEX. TUGboat, 42(2): 139–142, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb42-2/tb131ruckert-cweb.pdf.

Austin:2021:PSE

[4721] Oliver Austin. Plane and simple: Exploration of machine interaction with text type for visual-based navigation systems. TUGboat, 42(2):143, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131austin-plane.pdf.

Claudio:2021:MPJ

[4722] Jennifer Claudio. My personal journey into creating word search puzzles in Cyrillic and Arabic using multilingual support in L^{A}TEX. TUGboat, 42(2):144– 146, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-2/tb131claudio-puzzles.pdf.

Bossard:2021:TEJ

[4723] Antoine Bossard. On typesetting an English–Japanese book. TUGboat, 42(2): 147–151, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-2/tb131bossard-engjpn.pdf.

Sojka:2021:NCH

[4724] Petr Sojka and Ondřej Sojka. New Czechoslovak hyphenation patterns, word lists, and workflow. TUGboat, 42(2):152–158, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131sojka-czech.pdf.

Magdowski:2021:AGP

[4725] Mathias Magdowski. Automatically generate personalized tasks and sample solutions for the fundamentals of electrical engineering with PGFPlots and CircuiTikZ. TUGboat, 42(2):159–163, 2021. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb42-2/tb131magdowski-autoee.
pdf.

Veytsman:2021:CVS

[4726] Boris Veytsman and Keiran Harcombe. Creating a VPAT statement for TEX Live. TUGboat, 42(2):164–165, 2021. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb42-2/tb131veytsman-vpat.pdf.

Papasalouros:2021:PLD

[4727] Andreas Papasalouros and Antonis Tsolomitis. latex2nemeth: A direct L^{A}TEX-to-Braille transcribing tool. TUGboat, 42(2):166–169, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131papasalouros-latex2nemeth.pdf.

Fischer:2021:RTP

[4728] Ulrike Fischer. On the road to Tagged PDF: About StructElem, marked content, PDF/A and squeezed bärs. TUGboat, 42(2):170–173, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131fischer-tagpdf.pdf.

Vaughan:2021:TXL

[4729] Nicolás Vaughan. TEI-XML to L^{A}TEX workflow: Issues and lessons. TUGboat, 42(2):174–179, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb42-2/tb131vaughan-tei.pdf.

LP:2021:LNIa

[4730] L^{A}TEX Project Team. L^{A}TEX news, issue 33, June 2021. TUGboat, 42(2):180– 185, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-2/tb131ltnews33.pdf.

Novotny:2021:PML

[4731] Vít Novotný. Markdown 2.10.0: L^{A}TEX themes & snippets, two flavors of comments, and LuaMetaTeX. TUGboat, 42(2):186–193, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131novotny-markdown.pdf.

Talbot:2021:PBS

[4732] Nicola Talbot. bib2gls: sorting. TUGboat, 42(2):194–203, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131talbot-sorting.pdf.

Hufflen:2021:PB

[4733] Jean-Michel Hufflen. Programming bibliographies. TUGboat, 42(2):204– 207, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-2/tb131hufflen-bibprog.pdf.

Alpizar:2021:TA

[4734] Alexánder Borbón Alpízar, Apu V, Robert A. Beezer, H. Andrew Black, Michael Cantino, Paulo Cereda, David Crossland, Vic van Dijk, Jonathan Fine, Michal Hoftich, Amelia Hugill-Fontanel, Rahul Jaiswal, Emílio Kavamura, Vafa Khalighi, Alexei Kolesnikov, Rahul Krishnan S, Marcel Krüger, Al Maneki, Frank Mittelbach, Ross Moore, Michael Nolan, Todd Pagano, Hugh Paterson, Simon Porter, Thomas Price, Aravind Rajendran, T. Rishi, Matheus Rocha, Volker Sorge, Paulo Ney de Souza, Suhas Chikkanaravangala Vijayakumar, and Joseph Wright. TUG 2021 abstracts. TUGboat, 42(2): 208–212, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-2/tb131abstracts.pdf.

Anonymous:2021:MI

[4735] Anonymous. MAPS: issue 51 (2021). TUGboat, 42(2):212, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131maps.pdf.

Anonymous:2021:TKI

[4736] Anonymous. Die TeXnische Komödie issues 2–3/2021. TUGboat, 42(2): 213, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-2/tb131komo.pdf.

Berry:2021:TCb

[4737] Karl Berry. The treasure chest. TUGboat, 42(2):214–215, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131chest.pdf.

Anonymous:2021:TAM

[4738] Anonymous. TUG 2021 attendee map. TUGboat, 42(2):216, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-2/
tb131sponsors.pdf.

Anonymous:2021:TSD

[4739] Anonymous. TUG 2021 sponsors: DANTE e.V.; Google; Overleaf; STM Document Engineering Pvt Ltd; U. of Adelaide. TUGboat, 42(2):216–217, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb42-2/tb131sponsors.pdf.

Anonymous:2021:TCPb

[4740] Anonymous. TeX consulting and production services. TUGboat, 42(2):218– 219, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2021:Cb

[4741] Anonymous. Calendar. TUGboat, 42(2):220, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-2/tb131calendar.
pdf.

Anonymous:2021:IMb

[4742] Anonymous. Institutional members. TUGboat, 42(2):220, 2021. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Anonymous:2021:FCc

[4743] Anonymous. Front cover. TUGboat, 42(3):c1, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-3/cover1.pdf.

Anonymous:2021:IFCc

[4744] Anonymous. Inside front cover. TUGboat, 42(3):c2, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
cover2.pdf.

Anonymous:2021:CODc

[4745] Anonymous. Contents ordered by difficulty. TUGboat, 42(3):c3, 2021. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-
3/cover3.pdf.

Anonymous:2021:TPc

[4746] Anonymous. Title page. TUGboat, 42(3):221, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-3/titlepage.pdf.

Anonymous:2021:EIb

[4747] Anonymous. Editorial information. TUGboat, 42(3):222, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
tb132edit.pdf.

Anonymous:2021:IMc

[4748] Anonymous. Institutional members. TUGboat, 42(3):222, 2021. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Veytsman:2021:Pb

[4749] Boris Veytsman. From the president. TUGboat, 42(3):223, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
tb132pres.pdf.

Beeton:2021:ECb

[4750] Barbara Beeton. Editorial comments. TUGboat, 42(3):224–225, 2021. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-
3/tb132beet.pdf.

Beeton:2021:MDS

[4751] Barbara Beeton. Michael D. Spivak, 1940–2020. TUGboat, 42(3):226– 227, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-3/tb132beeton-spivak.pdf.

Beeton:2021:HKY

[4752] Barbara Beeton. How to keep your sanity when preparing a transcript of an online interview for publication. TUGboat, 42(3):228–230, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
tb132beeton-transcribe.pdf.

Bigelow:2021:FPT

[4753] Charles Bigelow. Form, pattern & texture in the typographic image. TUGboat, 42(3):231–246, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb42-3/tb132bigelow-form.pdf.

Anane:2021:ATJ

[4754] Amine Anane. Arabic text justification using LuaL^{A}TEX and the DigitalKhatt OpenType variable font. TUGboat, 42(3):247–257, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
tb132anane-variable.pdf.

Hagen:2021:ROS

[4755] Hans Hagen. Rendering open street maps. TUGboat, 42(3):258–266, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb42-3/tb132hagen-openstreetmap.pdf.

Voss:2021:CCF

[4756] Herbert Voß. Controlling captions, fullpage and doublepage floats: hvfloat. TUGboat, 42(3):267–283, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-3/tb132voss-
hvfloat.pdf.

Mittelbach:2021:PTP

[4757] Frank Mittelbach. Preventing tofu with pdfTEX and Unicode engines. TUGboat, 42(3):284–286, 2021. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb42-3/tb132mitt-tofu.pdf.

Mittelbach:2021:PUP

[4758] Frank Mittelbach. The unicodefonttable package. TUGboat, 42(3):287– 304, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-3/tb132mitt-unicodefonttable.pdf.

LP:2021:LNIb

[4759] L^{A}TEX Project Team. L^{A}TEX news, issue 34, November 2021. TUGboat, 42(3): 305–310, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-3/tb132ltnews34.pdf.

Berry:2021:PN

[4760] Karl Berry. Production notes. TUGboat, 42(3):310, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
tb132prod.pdf.

Hagen:2021:NUL

[4761] Hans Hagen. A new unit for LMTX: The dk. TUGboat, 42(3):311, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb42-3/tb132hagen-dk.pdf.

Berry:2021:TCc

[4762] Karl Berry. The treasure chest. TUGboat, 42(3):312, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
tb132chest.pdf.

Anonymous:2021:LG

[4763] Anonymous. La Lettre GUTenberg: 41–44 (2020–2021). TUGboat, 42(3): 313–314, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb42-3/tb132lettre.pdf.

Anonymous:2021:CGC

[4764] Anonymous. Les Cahiers GUTenberg: Contents of issue 58 (2021). TUGboat, 42(3):315, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb42-3/tb132cahi.pdf.

Anonymous:2021:CGJ

[4765] Anonymous. ConTeXt Group Journal: 14th meeting (2020). TUGboat, 42(3): 315–316, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb42-3/tb132ctxg.pdf.

Atkinson:2021:CF

[4766] John Atkinson. Comic: Fontzie. TUGboat, 42(3):316, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
tb132ctxg.pdf.

Anonymous:2021:TCPc

[4767] Anonymous. TEX consulting and production services. TUGboat, 42(3):317– 318, 2021. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2021:TI

[4768] Anonymous. TeXnology Inc. TUGboat, 42(3):318, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
tb132consultants.pdf.

Anonymous:2021:PAW

[4769] Anonymous. Pearson / Addison-Wesley. TUGboat, 42(3):319, 2021. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb42-3/
tb132ad-aw.pdf.

Anonymous:2021:Cc

[4770] Anonymous. Calendar. TUGboat, 42(3):320, 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb42-3/tb132calendar.
pdf.

Anonymous:2022:FCa

[4771] Anonymous. Front cover. TUGboat, 43(1):c1, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-1/cover1.pdf.

Anonymous:2022:IFCa

[4772] Anonymous. Inside front cover. TUGboat, 43(1):c2, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-1/
cover2.pdf.

Anonymous:2022:CODa

[4773] Anonymous. Contents ordered by difficulty. TUGboat, 43(1):c3, 2022. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-
1/cover3.pdf.

Anonymous:2022:TPa

[4774] Anonymous. Title page. TUGboat, 43(1):1, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-1/titlepage.pdf.

Anonymous:2022:EIa

[4775] Anonymous. Editorial information. TUGboat, 43(1):2, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-1/
tb133edit.pdf.

Anonymous:2022:IMa

[4776] Anonymous. Institutional members. TUGboat, 43(1):2, 2022. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Veytsman:2022:Pa

[4777] Boris Veytsman. From the President. TUGboat, 43(1):3, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-1/
tb133pres.pdf.

Beeton:2022:ECD

[4778] Barbara Beeton. Editorial comments. TUGboat, 43(1):4–6, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-1/
tb133beet.pdf.

Fine:2022:RFU

[4779] Jonathan Fine. Robin Fairbairns and UK TUG. TUGboat, 43(1):6, 2022. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-
1/tb133fairbairns-fine.pdf.

Andr:2022:LDG

[4780] Jacques André, Patrick Bideault, Denis Bitouzé, Michel Bovani, Thierry Bouche, Maxime Chupin, Daniel Flipo, and Yvon Henel. The last decade at GUT. TUGboat, 43(1):7–9, 2022. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb43-1/tb133andre-gut.pdf.

Novotny:2022:MWNa

[4781] Vít Novotný. Markdown 2.15.0: What’s new? TUGboat, 43(1):10–15, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb43-1/tb133novotny-markdown.pdf.

Maggi:2022:DBP

[4782] Carla Maggi. The DuckBoat — beginners’ pond: CDs, but not compact disks. TUGboat, 43(1):16–22, 2022. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb43-1/tb133duck-tikz-cd.pdf.

Bergmann:2022:MOS

[4783] Seth Bergmann. Making open source textbooks, and diagrams with AlDraTex. TUGboat, 43(1):23–27, 2022. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb43-1/tb133bergmann-diagrams.pdf.

Chernoff:2022:ARW

[4784] Max Chernoff. Automatically removing widows and orphans with lua-widow-control. TUGboat, 43(1):28–39, 2022. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb43-1/tb133chernoff-widows.

pdf. See updates [4872].

Wright:2022:PLB

[4785] Joseph Wright. l3build: The beginner’s guide. TUGboat, 43(1):40–43, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb43-1/tb133wright-l3build.pdf.

Talbot:2022:PBS

[4786] Nicola Talbot. bib2gls: standalone entries and repeated lists (a little book of poisons). TUGboat, 43(1):44–58, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-1/tb133talbot-
bib2gls-reorder.pdf.

Wermuth:2022:TFU

[4787] Udo Wermuth. Transparent file I/O using the original TEX program and the plain TEX format. TUGboat, 43(1):59–72, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-1/tb133wermuth-
trio.pdf.

Lang:2022:BRL

[4788] Sarah Lang. Book reviews: LaTeX Beginner’s Guide, second edition, by Stefan Kottwitz. TUGboat, 43(1):73–74, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-1/tb133reviews-
kottwitz2.pdf.

Berry:2022:TCD

[4789] Karl Berry. The treasure chest. TUGboat, 43(1):75–77, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-1/
tb133chest.pdf.

Berry:2022:TLN

[4790] Karl Berry. TEX Live 2022 news. TUGboat, 43(1):77, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-1/
tb133berry-tl22news.pdf.

Anonymous:2022:TKIa

[4791] Anonymous. Die TeXnische Komödie: issue 4/2021–1/2022. TUGboat, 43(1): 78–79, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-1/tb133komo.pdf.

Anonymous:2022:ZI

[4792] Anonymous. Zpravodaj: issue 2021/1–4. TUGboat, 43(1):79, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-1/

tb133zprav.pdf. Errata: in the printed TUGboat, the ‘TEX in a nutshell ’ article is incorrectly attributed to Petr Sojka and the author is Petr Olšák.

TBC:2022:TAG

[4793] TUG Bylaws committee. TUG Annual General Meeting procedures. TUGboat, 43(1):80, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-1/tb133agmproc.pdf.

Berry:2022:TFS

[4794] Karl Berry. TUG financial statements for 2021. TUGboat, 43(1):81, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb43-1/tb133treas.pdf.

Anonymous:2022:TCPa

[4795] Anonymous. TEX consulting and production services. TUGboat, 43(1):82– 83, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2022:TNI

[4796] Anonymous. TEXnology Inc. TUGboat, 43(1):83, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-1/
tb133consultants.pdf.

Anonymous:2022:Ca

[4797] Anonymous. Calendar. TUGboat, 43(1):84, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-1/tb133calendar.
pdf.

Anonymous:2022:FCb

[4798] Anonymous. Front cover. TUGboat, 43(2):c1, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-2/cover1.pdf.

Anonymous:2022:IFCb

[4799] Anonymous. Inside front cover. TUGboat, 43(2):c2, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-2/
cover2.pdf.

Anonymous:2022:CODb

[4800] Anonymous. Contents ordered by difficulty. TUGboat, 43(2):c3, 2022. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-
2/cover3.pdf.

Anonymous:2022:TPb

[4801] Anonymous. Title page. TUGboat, 43(2):85, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-2/titlepage.pdf.

Anonymous:2022:TCI

[4802] Anonymous. TUG 2022 conference information and program. TUGboat, 43 (2):86–88, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-2/tb134confinfo.pdf.

Hefferon:2022:TCR

[4803] Jim Hefferon. TUG 2022 conference report. TUGboat, 43(2):89, 2022. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-
2/tb134hefferon-tug22.pdf.

Laakso:2022:TAG

[4804] Robin Laakso. TUG 2022 annual general meeting notes. TUGboat, 43(2): 90–92, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134tug22-agm.pdf.

Berry:2022:DCW

[4805] Karl Berry. David C. Walden, 1942–2022. TUGboat, 43(2):93–95, 2022. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-
2/tb134walden-berry.pdf.

NeydeSouza:2022:IJL

[4806] Paulo Ney de Souza. Interview with John Lees-Miller, CTO of Overleaf. TUGboat, 43(2):96–99, 2022. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb43-2/tb134lees-miller-desouza.pdf.

NeydeSouza:2022:IBV

[4807] Paulo Ney de Souza. Interview with Boris Veytsman, scientist, TEX programmer and current TUG president. TUGboat, 43(2):100–104, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-2/
tb134veytsman-desouza.pdf.

Evia:2022:FTD

[4808] Carlos Evia. The future of technical documentation starts with its recent past. TUGboat, 43(2):104–107, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-2/tb134evia-
techdoc.pdf.

Spicer:2022:STC

[4809] Dag Spicer. A stroll through computer history at the CHM. TUGboat, 43(2): 108, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134spicer-chm.pdf.

Matteson:2022:TDC

[4810] Steven Matteson. Type design: Catching up to the past. TUGboat, 43(2): 109–119, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-2/tb134matteson-adapt.pdf.

Williams:2022:TPE

[4811] Peter K. G. Williams. The Tectonic Project: Envisioning a 21stcentury TEX experience. TUGboat, 43(2):120–126, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-2/
tb134williams-tectonic.pdf.

IT:2022:ITP

[4812] Island of TEX. IoT theatre presents: The Tempest. TUGboat, 43(2):127– 129, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134island-tempest.pdf.

Veytsman:2022:UPK

[4813] Boris Veytsman. Using knitr and L^{A}TEX for literate laboratory notes. TUGboat, 43(2):130–133, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-2/tb134veytsman-labnotes.pdf.

Thurman:2022:YTA

[4814] Marnanel Thurman. yex: a TEX-alike typesetter in Python. TUGboat, 43(2): 134–135, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134thurman-yex.pdf.

Hufflen:2022:EIT

[4815] Jean-Michel Hufflen. Extracting information from (La)TEX source files. TUGboat, 43(2):136–141, 2022. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb43-2/tb134hufflen-extract.pdf.

Apu:2022:LPA

[4816] Apu V, Rishi T, and Aravind Rajendran. L^{A}TEX profiling of author submissions — completeness & usability checking. TUGboat, 43(2):142–147, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb43-2/tb134apu-submissions.pdf.

LP:2022:LNIa

[4817] L^{A}TEX Project Team. L^{A}TEX news, issue 35, June 2022. TUGboat, 43(2):148– 154, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134ltnews35.pdf.

Guichard:2022:ILW

[4818] Éric Guichard and Jean-Michel Hufflen. Introductory L^{A}TEX workshop, en français. TUGboat, 43(2):155, 2022. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb43-2/tb134guichard-ltxworkshop-
fr.pdf.

Prentice:2022:SPL

[4819] Lloyd Prentice. Self-publishing, L^{A}TEX, and Markdown. TUGboat, 43(2):156– 158, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134prentice-selfpub.pdf.

Cereda:2022:SSP

[4820] Paulo Cereda and Phelype Oleinik. The story of a silly package. TUGboat, 43 (2):159–161, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-2/tb134cereda-silly.pdf.

Wright:2022:KVS

[4821] Joseph Wright. Key–value setting handling in the L^{A}TEX kernel. TUGboat, 43 (2):162–163, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-2/tb134wright-keyval.pdf.

Wright:2022:PSL

[4822] Joseph Wright. siunitx: Launching version 3. TUGboat, 43(2):164–165, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb43-2/tb134wright-siunitx3.pdf.

Wright:2022:CCL

[4823] Joseph Wright. Case changing: L^{A}TEX reaches Unicode-land. TUGboat, 43(2): 165–166, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134wright-case.pdf.

Fischer:2022:USC

[4824] Ulrike Fischer. Using spot colors in L^{A}TEX. TUGboat, 43(2):167–172, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb43-2/tb134fischer-spotcolor.pdf.

Shirore:2022:PLL

[4825] Chetan Shirore and Ajit Kumar. The luatruthtable L^{A}TEX package. TUGboat, 43 (2):172–175, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-2/tb134shirore-luatruthtable.pdf.

Baranovskyi:2022:LCD

[4826] Oleksandr Baranovskyi. L^{A}TEX classes for doctoral theses in Ukraine: Interesting tips and painful problems. TUGboat, 43(2):176–181, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-2/
tb134baranovskyi-vakthesis.pdf.

Black:2022:XUT

[4827] H. Andrew Black and Hugh Paterson III. XLingPaper’s use of TEX technologies. TUGboat, 43(2):182–196, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-2/tb134black-
xlingpaper.pdf.

Muller:2022:LBE

[4828] Dennis Müller and Michael Kohlhase. A L^{A}TEX-based ecosystem for semantic/ active mathematical documents. TUGboat, 43(2):197–201, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-2/
tb134mueller-stex3.pdf.

Hagen:2022:PMF

[4829] Hans Hagen and Mikael Sundqvist. Pushing math forward with ConTEXt lmtx. TUGboat, 43(2):202–206, 2022. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb43-2/tb134hagen-math.pdf.

Berry:2022:TCb

[4830] Karl Berry. The treasure chest. TUGboat, 43(2):207–208, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-2/
tb134chest.pdf.

Austin:2022:TA

[4831] Oliver Austin, David Blakesley, Sarai Castañeda, Max Chernoff, Hong Dat Cheung, Jennifer Claudio, Jonathan Fine, Patrick Gundlach, Herbert Hickman, Nicolas Jimenez, Vafa Khalighi, Tia Luc, Matthew Mariano, Ross Moore, Aditya Ohri, Christopher Park, Emily Park, Norbert Preining, samcarter, Tanya Schmah, Tereza Vrabcová, and Haibin Wu. TUG 2022 abstracts. TUGboat, 43 (2):209–212, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-2/tb134abstracts.pdf.

Anonymous:2022:MI

[4832] Anonymous. MAPS: issue 52 (2022). TUGboat, 43(2):213, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-2/
tb134maps.pdf.

Anonymous:2022:LGIa

[4833] Anonymous. La Lettre GUTenberg: issue 45 (2022). TUGboat, 43(2): 214, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134lettre.pdf.

Anonymous:2022:TNK

[4834] Anonymous. Die TEXnische Komödie: issue 2/2022. TUGboat, 43(2):215, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134komo.pdf.

Anonymous:2022:IMb

[4835] Anonymous. Institutional members. TUGboat, 43(2):215, 2022. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

TEc:2022:TUGa

[4836] TUG Elections committee. 2023 TEX Users Group election. TUGboat, 43(2): 216, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-2/tb134elec.pdf.

Anonymous:2022:Cb

[4837] Anonymous. Calendar. TUGboat, 43(2):217, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-2/tb134calendar.
pdf.

Anonymous:2022:TAS

[4838] Anonymous. TUG 2022 advertisements: STM Document Engineering Pvt Ltd; Overleaf. TUGboat, 43(2):218, 2022. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb43-2/tb134tug22-ads.pdf.

Anonymous:2022:TCPb

[4839] Anonymous. TEX consulting and production services. TUGboat, 43(2):219– 220, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2022:FCc

[4840] Anonymous. Front cover. TUGboat, 43(3):c1, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-3/cover1.pdf.

Anonymous:2022:IFCc

[4841] Anonymous. Inside front cover. TUGboat, 43(3):c2, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
cover2.pdf.

Anonymous:2022:CODc

[4842] Anonymous. Contents ordered by difficulty. TUGboat, 43(3):c3, 2022. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-
3/cover3.pdf.

Anonymous:2022:TPc

[4843] Anonymous. Title page. TUGboat, 43(3):221–225, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
titlepage.pdf.

Anonymous:2022:EIb

[4844] Anonymous. Editorial information. TUGboat, 43(3):222, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135edit.pdf.

Anonymous:2022:IMc

[4845] Anonymous. Institutional members. TUGboat, 43(3):222, 2022. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Veytsman:2022:Pb

[4846] Boris Veytsman. From the President. TUGboat, 43(3):223–224, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135pres.pdf.

Beeton:2022:ECb

[4847] Barbara Beeton. Editorial comments. TUGboat, 43(3):225–226, 2022. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-
3/tb135beet.pdf.

Fine:2022:RTU

[4848] Jonathan Fine. Regarding TUG and UK-TUG. TUGboat, 43(3):226–228, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135fine-uktug.pdf.

Rosendahl:2022:R

[4849] Arthur Rosendahl. Rebuttal. TUGboat, 43(3):228, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135rosendahl-uktug.pdf.

Fine:2022:TH

[4850] Jonathan Fine. The TEX hour. TUGboat, 43(3):229, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135fine-texhour.pdf.

Kottwitz:2022:DPF

[4851] Stefan Kottwitz. DANTE project funding of TEX servers. TUGboat, 43(3): 230–231, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-3/tb135kottwitz-servers.pdf.

Blakesley:2022:RCP

[4852] David Blakesley. The residual concepts of production vs. the emergent cultures of distribution in publishing. TUGboat, 43(3):232–239, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135blakesley-production.pdf.

Flynn:2022:TI

[4853] Peter Flynn. Typographers’ Inn. TUGboat, 43(3):240–242, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135inn.pdf.

Bellows:2022:FMP

[4854] David Bellows. Formatting mesostic poems à la John Cage. TUGboat, 43(3): 243–246, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135bellows-mesostic.pdf.

Bien:2022:RPA

[4855] Janusz S. Bień. Representing Parkosz’s alphabet in the Junicode font. TUGboat, 43(3):247–251, 2022. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb43-3/tb135bien-parkosz.pdf.

Koch:2022:TSV

[4856] Richard Koch. TeXShop, version 5: HTML previews. TUGboat, 43(3):252– 253, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135koch-texshop5.pdf.

Koch:2022:ICU

[4857] Richard Koch. Interactive content using TEX4ht. TUGboat, 43(3):254– 260, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135koch-tex4ht.pdf.

Hoftich:2022:WNT

[4858] Michal Hoftich. What’s new in TEX4ht: 2022. TUGboat, 43(3):261–262, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb43-3/tb135hoftich-make4ht.pdf.

Fischer:2022:AXM

[4859] Ulrike Fischer and Frank Mittelbach. Adding XMP metadata in L^{A}TEX. TUGboat, 43(3):263–267, 2022. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb43-3/tb135fischer-xmp.pdf.

Mittelbach:2022:LTP

[4860] Frank Mittelbach and Ulrike Fischer. The L^{A}TEXTagged PDF project — a status and progress report. TUGboat, 43(3):268–272, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135mitt-tagged.pdf.

LP:2022:LNIb

[4861] L^{A}TEX Project Team. L^{A}TEX news, issue 36, November 2022. TUGboat, 43(3): 273–275, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135ltnews36.pdf.

Novotny:2022:MWNb

[4862] Vít Novotný. Markdown 2.17.1: What’s new, what’s next? TUGboat, 43(3): 276–278, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135novotny-markdown.pdf.

Wright:2022:MIC

[4863] Joseph Wright. Mapping to individual characters in expl3. TUGboat, 43(3): 279, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135wright-grapheme.pdf.

Voss:2022:TEP

[4864] Herbert Voß. Typesetting external program code and its output: hvextern. TUGboat, 43(3):280–292, 2022. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb43-3/tb135voss-extern.pdf.

Shirore:2022:PLP

[4865] Chetan Shirore and Ajit Kumar. The luamodulartables and luaset L^{A}TEX packages. TUGboat, 43(3):293–294, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-3/tb135shirore-
luamath.pdf.

Voss:2022:UOT

[4866] Herbert Voß. Using OpenType and TrueType fonts with XEL^{A}TEX and LuaL^{A}TEX. TUGboat, 43(3):295–299, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-3/tb135voss-
unifont.pdf.

Hagen:2022:NDM

[4867] Hans Hagen and Mikael Sundqvist. New directions in math fonts. TUGboat, 43(3):300–310, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-3/tb135hagen-mathchange.pdf.

Hagen:2022:PLB

[4868] Hans Hagen and Mikael Sundqvist. Patching Lucida Bright Math. TUGboat, 43(3):311–316, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb43-3/tb135hagen-lucida.pdf.

Romer:2022:VTM

[4869] Linus Romer. Ventrella’s terdragon in . TUGboat, 43(3):317– 318, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135romer-terdragon.pdf.

Finston:2022:IPG

[4870] Laurence Finston. An introduction to GNU 3DLDF. TUGboat, 43(3):319– 332, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135finston-3dldf.pdf.

Finston:2022:GEE

[4871] Laurence Finston. A graphical ellipse envelope construction with GNU 3DLDF. TUGboat, 43(3):333–339, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-3/tb135finston-
ellipse.pdf.

Chernoff:2022:UAR

[4872] Max Chernoff. Updates to “Automatically removing widows and orphans with lua-widow-control”, TUGboat 43:1. TUGboat, 43(3):340–342, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135chernoff-lwc.pdf. See [4784].

Wermuth:2022:CBP

[4873] Udo Wermuth. Can “\parfillskip = 0pt” shorten a short paragraph in plain TEX by two lines? TUGboat, 43(3):343–350, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135wermuth-shorten.pdf.

Papasalouros:2022:PLP

[4874] Andreas Papasalouros and Antonis Tsolomitis. LaTeX2Nemeth and the amsmath package. TUGboat, 43(3):351–362, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-3/tb135papasalouros-
amsmath.pdf.

Berry:2022:TCc

[4875] Karl Berry. The treasure chest. TUGboat, 43(3):363, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135chest.pdf.

Anonymous:2022:TKIb

[4876] Anonymous. Die TeXnische Komödie: issue 3/2022. TUGboat, 43(3):364, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135komo.pdf.

Anonymous:2022:LGIb

[4877] Anonymous. La Lettre GUTenberg: issue 45 (2022). TUGboat, 43(3): 364, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135lettre.pdf.

TEc:2022:TUGb

[4878] TUG Elections committee. 2023 TEX Users Group election. TUGboat, 43(3): 365, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb43-3/tb135elec.pdf.

Anonymous:2022:TCPc

[4879] Anonymous. TEX consulting and production services. TUGboat, 43(3):366– 367, 2022. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2022:TI

[4880] Anonymous. TeXnology Inc. TUGboat, 43(3):367, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb43-3/
tb135consultants.pdf.

Anonymous:2022:Cc

[4881] Anonymous. Calendar. TUGboat, 43(3):368, 2022. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb43-3/tb135calendar.
pdf.

Anonymous:2023:FCa

Anonymous:2023:IFCa

[4883] Anonymous. Inside front cover. TUGboat, 44(1):c2, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-1/
cover2.pdf.

Anonymous:2023:CODa

[4884] Anonymous. Contents ordered by difficulty. TUGboat, 44(1):c3, 2023. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-
1/cover3.pdf.

Anonymous:2023:TPa

[4885] Anonymous. Title page. TUGboat, 44(1):1, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-1/titlepage.pdf.

Anonymous:2023:EIa

[4886] Anonymous. Editorial information. TUGboat, 44(1):2, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-1/
tb136edit.pdf.

Anonymous:2023:IMa

[4887] Anonymous. Institutional members. TUGboat, 44(1):2, 2023. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Veytsman:2023:Pa

[4888] Boris Veytsman. From the President. TUGboat, 44(1):3, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-1/
tb136pres.pdf.

Beeton:2023:ECa

[4889] Barbara Beeton. Editorial comments. TUGboat, 44(1):4–8, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-1/
tb136beet.pdf.

Romano:2023:CTD

[4890] Frank Romano. A conversation with type designer Matthew Carter. TUGboat, 44(1):9–20, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-1/tb136carter-romano.pdf.

Andre:2023:PDF

[4891] Jacques André. Prehistory of digital fonts. TUGboat, 44(1):21–57, 2023. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-
1/tb136andre-prehistory.pdf.

Flynn:2023:TIa

[4892] Peter Flynn. Typographers’ Inn. TUGboat, 44(1):58–59, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-1/
tb136inn.pdf.

Wilde:2023:AJT

[4893] Tine Wilde. An artist’s journey on a TUGboat. TUGboat, 44(1):60–63, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-1/tb136wilde-art.pdf.

Maggi:2023:DBP

[4894] Carla Maggi. The DuckBoat — beginners’ pond: No more table nightmares with tabularray! TUGboat, 44(1):64–70, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-1/tb136duck-
tabularray.pdf.

Bos:2023:MJP

[4895] Joppe Bos and Kevin McCurley. Metadata in journal publishing. TUGboat, 44(1):71–76, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-1/tb136bos-metadata.pdf.

Mittelbach:2023:LAL

[4896] Frank Mittelbach. L^{A}TEX anniversaries — a look in two directions. TUGboat, 44(1):77–86, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-1/tb136mitt-history.pdf.

Peischl:2023:IE

[4897] Marei Peischl. An introduction to expl3. TUGboat, 44(1):87–93, 2023. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-
1/tb136peischl-expl3.pdf.

All:2023:LSC

[4898] Timothy All and Evan Cochrane. LuaCAS: Symbolic computation in L^{A}TEX. TUGboat, 44(1):94–98, 2023. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb44-1/tb136all-luacas.pdf.

Novotny:2023:AM

[4899] Vít Novotný. Attributes in Markdown. TUGboat, 44(1):99–101, 2023. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-
1/tb136novotny-markdown-attr.pdf.

Borja:2023:IAD

[4900] Igor Borja. An introduction to automata design with TikZ’s automata library. TUGboat, 44(1):102–107, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-1/tb136prado-
automata.pdf.

Stenborg:2023:SRP

[4901] Travis Stenborg. Styling R ggplot2 graphics with L^{A}TEX. TUGboat, 44(1): 108–109, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-1/tb136stenborg-ggplot2.pdf.

Bien:2023:CAU

[4902] Janusz S. Bień. Creating annotated Unicode-like font charts. TUGboat, 44(1): 110–115, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-1/tb136bien-unichart.pdf.

Berry:2023:PNa

[4903] Karl Berry. Production notes. TUGboat, 44(1):115, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-1/
tb136prod.pdf.

Hagen:2023:OEB

[4904] Hans Hagen and Mikael Sundqvist. OpenType extensible brace debugging. TUGboat, 44(1):116, 2023. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb44-1/tb136hagen-extensible.pdf.

Hagen:2023:CTL

[4905] Hans Hagen. ConTEXt in TeX Live 2023. TUGboat, 44(1):117–120, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-1/tb136hagen-texlive.pdf.

Olsak:2023:CMO

[4906] Petr Olšák. Creating macros in OpTeX. TUGboat, 44(1):121–126, 2023. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-
1/tb136olsak-optexmac.pdf.

Wermuth:2023:RMG

[4907] Udo Wermuth. Reflections on \globaldefs. TUGboat, 44(1):127–133, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-1/tb136wermuth-globaldefs.pdf.

Hagen:2023:PMC

[4908] Hans Hagen. Preserving the math class of variables. TUGboat, 44(1): 133, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-1/tb136hagen-classes.pdf.

Wright:2023:SUD

[4909] Joseph Wright. Storing Unicode data in TEX engines. TUGboat, 44(1):134– 137, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-1/tb136wright-unidata.pdf.

Hefferon:2023:BRD

[4910] Jim Hefferon. Book reviews: Do Not Erase, by Jessica Wynne. TUGboat, 44 (1):137–138, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-1/tb136reviews-wynne.pdf.

Lamb:2023:BRS

[4911] John Lamb. Book reviews: Stop Stealing Sheep & Find out how type works, by Erik Spiekermann. TUGboat, 44(1):139–140, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-1/tb136reviews-
spiekermann.pdf.

Berry:2023:TCa

[4912] Karl Berry. The treasure chest. TUGboat, 44(1):141–142, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-1/
tb136chest.pdf.

Anonymous:2023:TKI

[4913] Anonymous. Die TeXnische Komödie: issues 4/2022–1/2023. TUGboat, 44 (1):143, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-1/tb136komo.pdf.

Anonymous:2023:LGIa

[4914] Anonymous. La Lettre GUTenberg: issues 47–49 (2022). TUGboat, 44(1): 143–144, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-1/tb136lettre.pdf.

Anonymous:2023:ZIa

[4915] Anonymous. Zpravodaj: issue 2022/1–4. TUGboat, 44(1):145, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-1/
tb136zprav.pdf.

Anonymous:2023:TI

[4916] Anonymous. TeXnology Inc. TUGboat, 44(1):145, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-1/
tb136consultants.pdf.

Anonymous:2023:TCPa

[4917] Anonymous. TEX consulting and production services. TUGboat, 44(1): 146, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

committee:2023:TER

[4918] TUG Elections committee. TUG 2023 election report. TUGboat, 44(1):147– 150, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-1/tb136elec.pdf.

Berry:2023:TFS

[4919] Karl Berry. TUG financial statements for 2022. TUGboat, 44(1):151, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-1/tb136treas.pdf.

Anonymous:2023:Ca

[4920] Anonymous. Calendar. TUGboat, 44(1):152, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-1/tb136calendar.
pdf.

Anonymous:2023:FCb

[4921] Anonymous. Front cover. TUGboat, 44(2):c1, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/cover1.pdf.

Anonymous:2023:IFCb

[4922] Anonymous. Inside front cover. TUGboat, 44(2):c2, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
cover2.pdf.

Anonymous:2023:CODb

[4923] Anonymous. Contents ordered by difficulty. TUGboat, 44(2):c3, 2023. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-
2/cover3.pdf.

Anonymous:2023:TPb

[4924] Anonymous. Title page. TUGboat, 44(2):153, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/titlepage.pdf.

Anonymous:2023:TCI

[4925] Anonymous. TUG 2023 conference information and program. TUGboat, 44 (2):154–156, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-2/tb137confinfo.pdf.

Wagner:2023:RBT

[4926] Eileen Wagner. A rally in Bonn: TUG 2023. TUGboat, 44(2):157–162, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-2/tb137wagner-tug23.pdf.

Berry:2023:TAG

[4927] Karl Berry and Robin Laakso. TUG 2023 Annual General Meeting notes. TUGboat, 44(2):162–163, 2023. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb44-2/tb137tug23-agm.pdf.

Beeton:2023:WET

[4928] Barbara Beeton. What every (La)TeX newbie should know. TUGboat, 44(2): 164–169, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-2/tb137beeton-basic.pdf.

Ramm:2023:AGB

[4929] Henning Hraban Ramm. Architectural guides for Bonn — book production with ConTEXt. TUGboat, 44(2):170–175, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/tb137ramm-
books.pdf.

Ramm:2023:CCC

[4930] Henning Hraban Ramm. Calculating covers with ConTEXt. TUGboat, 44(2): 176–179, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-2/tb137ramm-covers.pdf.

Egger:2023:NDE

[4931] Willi Egger, Hans Hagen, Edith Sundqvist, and Mikael P. Sundqvist. New dimensions: Edith and Tove. TUGboat, 44(2):180–185, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137egger-edith.pdf.

Schmitz:2023:PDF

[4932] Thomas A. Schmitz. Producing different forms of output from XML via ConTEXt. TUGboat, 44(2):186–191, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/tb137schmitz-
xmlcontext.pdf.

LP:2023:LNI

[4933] L^{A}TEXProject Team. L^{A}TEX news, issue 37, June 2023. TUGboat, 44(2):192– 196, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-2/tb137ltnews37.pdf.

samcarter:2023:BN

[4934] samcarter and Joseph Wright. Beamer news: 2023. TUGboat, 44(2):197– 199, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-2/tb137samcarter-beamernews23.pdf.

Veytsman:2023:UPN

[4935] Boris Veytsman. Updating the nostarch class. TUGboat, 44(2):200–202, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-2/tb137veytsman-nostarch.pdf.

Kopp:2023:LTG

[4936] Oliver Kopp. The L^{A}TEX template generator: How micro-templates reduce template maintenance effort. TUGboat, 44(2):203–207, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137kopp-microtemplates.pdf.

Hagen:2023:BAO

[4937] Hans Hagen and Mikael P. Sundqvist. On bottom accents in OpenType math. TUGboat, 44(2):207–208, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/tb137hagen-
bottomaccent.pdf.

Sannier:2023:MRC

[4938] Victor Sannier. A Metafont for rustic capitals. TUGboat, 44(2):209–212, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-2/tb137sannier-rustic.pdf.

Vieth:2023:USO

[4939] Ulrik Vieth. An updated survey of OpenType math fonts. TUGboat, 44(2): 213–225, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-2/tb137vieth-otmath.pdf.

Sundqvist:2023:SOM

[4940] Mikael P. Sundqvist and Hans Hagen. Standardizing OpenType math fonts. TUGboat, 44(2):226–232, 2023. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb44-2/tb137sundqvist-otmath.
pdf.

samcarter:2023:BSG

[4941] samcarter and Gert Fischer. Behind the scenes of the Great TikZlings Christmas Extravaganza. TUGboat, 44(2):233–235, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137samcarter-tikzlings.pdf.

Romer:2023:CCH

[4942] Linus Romer. Curvature combs and harmonized paths in MetaPost. TUGboat, 44(2):236–239, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-2/tb137romer-curvetools.pdf.

Hefferon:2023:UAL

[4943] Jim Hefferon. Using Asymptote like MetaPost. TUGboat, 44(2):240–241, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-2/tb137hefferon-asymptote.pdf.

Verna:2023:IRT

[4944] Didier Verna. Interactive and real-time typesetting for demonstration and experimentation: ETAP. TUGboat, 44(2):242–248, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137verna-realtime.pdf.

IT:2023:LCT

[4945] Island of TEX. Living in containers — on TEX Live (and ConTEXt) in a Docker setting. TUGboat, 44(2):249–252, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/tb137island-
docker.pdf.

Ruckert:2023:NHP

[4946] Martin Ruckert. News from the HINT project: 2023. TUGboat, 44(2):252– 254, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-2/tb137ruckert-hint23.pdf.

Davies:2023:BRT

[4947] Ben Davies. Bumpy road towards a good L^{A}TEX visual editor at Overleaf. TUGboat, 44(2):255, 2023. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb44-2/tb137davies-visual.pdf.

Hejda:2023:OTL

[4948] Tom Hejda. Overleaf and TEX Live deployment. TUGboat, 44(2):256, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-2/tb137hejda-overleaf-tl.pdf.

T:2023:PNS

[4949] Rishi T, Apu V, Hàn Thê´ Thành, and Jan Vaněk. Primo — a new sustainable solution for publishing. TUGboat, 44(2):257–261, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137rishi-primo.pdf.

Fischer:2023:ATL

[4950] Ulrike Fischer and Frank Mittelbach. Automated tagging of L^{A}TEX documents — what is possible today, in 2023? TUGboat, 44(2):262–266, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137fischer-tagging23.pdf.

Carlisle:2023:RLT

[4951] David Carlisle, Ulrike Fischer, and Frank Mittelbach. Report on the L^{A}TEX Tagged PDF workshop, TUG 2023. TUGboat, 44(2):267–269, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137carlisle-taggedpdfworkshop23.pdf.

Moore:2023:EAS

[4952] Ross Moore. Enhancing accessibility of structured information via ‘Tagged PDF’. TUGboat, 44(2):270–274, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/tb137moore-
pdf-html.pdf.

Muller:2023:HCS

[4953] Dennis Müller. An HTML/CSS schema for TEX primitives — generating highquality responsive HTML from generic TEX. TUGboat, 44(2):275–286, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-2/tb137mueller-primitives.pdf.

Hagen:2023:CWM

[4954] Hans Hagen. Cheats (or not): When \prevdepth = -1000pt. TUGboat, 44 (2):287–288, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-2/tb137hagen-prevdepth.pdf.

Sojka:2023:RUS

[4955] Ondřej Sojka, Petr Sojka, and Jakub Máca. A roadmap for universal syllabic segmentation. TUGboat, 44(2):289–296, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/tb137sojka-
syllabic.pdf.

Radhakrishnan:2023:MMC

[4956] CV Radhakrishnan, KV Rajeesh, and KH Hussain. Metafont/MetaPost and a complex Indic script: Malayalam. TUGboat, 44(2):297–314, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137radhakrishnan-malayalam.pdf.

Gundlach:2023:TA

[4957] Patrick Gundlach, Island of TeX, Eberhard W. Lisse, Frank Mittelbach, Vít Novotný, samcarter, Jan Šustek, and Joseph Wright. TUG 2023 abstracts. TUGboat, 44(2):315–316, 2023. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb44-2/tb137abstracts.pdf.

Anonymous:2023:AIM

[4958] Anonymous. ArsTeXnica: issue 34 (May 2023). TUGboat, 44(2):317, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-2/tb137arstexnica.pdf.

Anonymous:2023:LGIb

[4959] Anonymous. La Lettre GUTenberg: issue 50 (2023). TUGboat, 44(2): 317, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-2/tb137lettre.pdf.

Anonymous:2023:CGJ

[4960] Anonymous. ConTEXt Group Journal: 15th meeting (2021). TUGboat, 44(2): 318, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-2/tb137ctxg.pdf.

Anonymous:2023:TNK

[4961] Anonymous. Die TEXnische Komödie: issue 2/2023. TUGboat, 44(2): 318, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-2/tb137komo.pdf.

Anonymous:2023:ZIb

[4962] Anonymous. Zpravodaj: issue 2023/1–2. TUGboat, 44(2):319, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137zprav.pdf.

Gratzer:2023:GSM

[4963] George Grätzer. The gods smile at me: The L^{A}TEX Companion, third edition, and ChatGPT. TUGboat, 44(2):319–321, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/tb137gratzer-
smile.pdf.

Lamb:2023:BRLa

[4964] John Lamb. Book review: The L^{A}TEX Companion, 3rd edition, by Frank Mittelbach with Ulrike Fischer. TUGboat, 44(2):322–324, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-2/
tb137reviews-mittelbach-tlc.pdf.

Anonymous:2023:TAS

[4965] Anonymous. TUG 2023 advertisements: STM Document Engineering Pvt Ltd; Overleaf. TUGboat, 44(2):325, 2023. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb44-2/tb137tug23-ads.pdf.

Anonymous:2023:TCPb

[4966] Anonymous. TEX consulting and production services. TUGboat, 44(2):326– 327, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2023:IMb

[4967] Anonymous. Institutional members. TUGboat, 44(2):327, 2023. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Anonymous:2023:Cb

[4968] Anonymous. Calendar. TUGboat, 44(2):328, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-2/tb137calendar.
pdf.

Anonymous:2023:FCc

[4969] Anonymous. Front cover. TUGboat, 44(3):c1, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-3/cover1.pdf.

Anonymous:2023:IFCc

[4970] Anonymous. Inside front cover. TUGboat, 44(3):c2, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-3/
cover2.pdf.

Anonymous:2023:CODc

[4971] Anonymous. Contents ordered by difficulty. TUGboat, 44(3):c3, 2023. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-
3/cover3.pdf.

Anonymous:2023:TPc

[4972] Anonymous. Title page. TUGboat, 44(3):329, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-3/titlepage.pdf.

Anonymous:2023:EIb

[4973] Anonymous. Editorial information. TUGboat, 44(3):330, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-3/
tb138edit.pdf.

Anonymous:2023:IMc

[4974] Anonymous. Institutional members. TUGboat, 44(3):330, 2023. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/instmem.html.

Beeton:2023:ECb

[4975] Barbara Beeton. Editorial comments. TUGboat, 44(3):331–333, 2023. CO-

DEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-
3/tb138beet.pdf.

Flynn:2023:TIb

[4976] Peter Flynn. Typographers’ Inn. TUGboat, 44(3):334–335, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-3/
tb138inn.pdf.

Matthiopoulos:2023:SHG

[4977] George D. Matthiopoulos. A short history of Greek type design before the digital era. TUGboat, 44(3):336–363, 2023. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb44-3/tb138matthiopoulos-greek.
pdf.

Bien:2023:TIO

[4978] Janusz S. Bień. Towards an inventory of old print characters: Ungler’s Rubricella, a case study. TUGboat, 44(3):364–375, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-3/
tb138bien-rubricella.pdf.

Stenborg:2023:TRA

[4979] Travis Stenborg. A TikZ rendering of the Arecibo message. TUGboat, 44(3): 375–377, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-3/tb138stenborg-arecibo.pdf.

Finston:2023:PA

[4980] Laurence Finston. A pattern from the Alhambra. TUGboat, 44(3):378– 389, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-3/tb138finston-alhambra.pdf.

Kortemeyer:2023:UAI

[4981] Gerd Kortemeyer. Using artificial-intelligence tools to make L^{A}TEX content accessible to blind readers. TUGboat, 44(3):390–399, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-3/
tb138kortemeyer-accessibility.pdf.

Team:2023:LNI

[4982] LaTeX Project Team. L^{A}TEX news, issue 38, November 2023. TUGboat, 44 (3):400–403, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.

org/TUGboat/tb44-3/tb138ltnews38.pdf.

V:2023:MTR

[4983] Vít Novotný. Markdown 3 at TUG 2023: Reflections from the Q&A session. TUGboat, 44(3):404–405, 2023. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb44-3/tb138starynovotny-markdown-
qa.pdf.

Somma:2023:SLO

[4984] Emmanuele F. Somma. Simplifying L^{A}TEX with ORG-mode in Emacs. TUGboat, 44(3):406–420, 2023. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb44-3/tb138somma-orgmode.pdf.

Wermuth:2023:PTX

[4985] Udo Wermuth. Plain TeX’s \ooalign. TUGboat, 44(3):421–426, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-3/tb138wermuth-ooalign.pdf.

Wermuth:2023:VAP

[4986] Udo Wermuth. Vertical alignments in plain TEX. TUGboat, 44(3):427– 440, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-3/tb138wermuth-valign.pdf.

Kopp:2023:JBB

[4987] Oliver Kopp, Carl Christian Snethlage, and Christoph Schwentker. JabRef: BibTEX-based literature management software. TUGboat, 44(3):441–447, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-3/tb138kopp-jabref.pdf.

Berry:2023:TCb

[4988] Karl Berry. The treasure chest. TUGboat, 44(3):447–449, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-3/
tb138chest.pdf.

Berry:2023:PNb

[4989] Karl Berry. Production notes. TUGboat, 44(3):449, 2023. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/tb44-3/
tb138prod.pdf.

Lamb:2023:BRL

[4990] John D. Lamb. Book review: The L^{A}TEX Companion, 3rd edition, by Frank Mittelbach with Ulrike Fischer (PDF edition). TUGboat, 44(3):450–451, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/TUGboat/

tb44-3/tb138reviews-mittelbach-tlc-pdf.pdf.

samcarter:2023:BRL

[4991] samcarter. Book review: L^{A}TEX Graphics with TikZ, by Stefan Kottwitz. TUGboat, 44(3):452, 2023. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb44-3/tb138reviews-kottwitz-tikz.
pdf.

Berry:2023:BRG

[4992] Karl Berry. Book review: Godine at Fifty, by David R. Godine. TUGboat, 44 (3):453, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/tb44-3/tb138reviews-godine50.pdf.

Anonymous:2023:TCP

[4993] Anonymous. TEX consulting and production services. TUGboat, 44(3):454– 455, 2023. CODEN ???? ISSN 0896-3207. URL https://tug.org/

TUGboat/consultants.html.

Anonymous:2023:TGR

[4994] Anonymous. le-tex GmbH [reports and notices — TUG 2023 sponsor]. TUGboat, 44(3):455, 2023. CODEN ???? ISSN 0896-3207. URL https://

tug.org/TUGboat/tb44-3/tb138consultants.pdf.

Anonymous:2023:Cc

[4995] Anonymous. Calendar. TUGboat, 44(3):456, 2023. CODEN ???? ISSN 0896-

3207. URL https://tug.org/TUGboat/tb44-3/tb138calendar.
pdf.